A Bibliography of Books and Articles about UNIX and UNIX Programming

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

02 May 2024 Version 4.99

Abstract

This bibliography records books and historical publications about the UNIX operating system, and UNIX programming tools. It mostly excludes networks and network programming, which are treated in a separate bibliography, internet.bib. This bibliography was started from material in postings to the sunflash list, volume 46, number 17, October 1992, and volume 57, number 29, September 1993, by Samuel Ko (e-mail: kko@fraser.sfu.ca), and then significantly extended by the present author. Entry commentaries are largely due to S. Ko.

General UNIX texts

[AL92a, AL95a, AL95b, AA86, AS93b, Ari92, Bou90a, Chr83a, Chr88, CR94, Cof90, Coh94, Dun91a, Gar91, Gt92a, Gt92b, Hah93, Hah94a, HA90, Hol92, KL94, LY93, LR89a, LL83a, MM83a, Mik89, MBL89, NS92, NH91, POLo93, PM87, RRF90, RRF93, RRH94, Rus91, Sob89, Sob91, Sob95, Sou92, TY82b, Tim93, Top90a, Top90b, Val92b, SSC93, WMP92]

Catalogs and book lists

[O’R93, Spu92, Wri93]

Communications software

[Cam87, dC87, dG93, Gia90]

Compilers

[DS88, JL78, Joh79c, JL87a, LS79a, LS79b, LMB92, MB90, SF85]

Database

[Rod90a]

Dictionaries

[MP93, Ray91, SD93, WC93]

Document preparation and typesetting

[BR87, BBM85, Geh87a, MM90a] See also the extensive bibliographies about the TEX typesetting system available via

anonymous ftp from ftp.math.utah.edu in
pub/tex/bib.

Editors

[CR91, CoL92b, Fin91, Hew90, Lam90, RB92a, SBA92, Sta93]

Electronic Mail and Netnews

[FA89, FA94, Har95, LP95]

Graphics files

[BS94, Lev94, Mv94]

History

[?, ?, Cat91, RT74, RT79a]

Linux

[BBD^{+}96, BF97, HP^{+}95, Kir95a, Kir95b, PR96b, Sob97, SU94b, SU95, SUM95, TGB95, TG96, VRJ95, VRJ96, VJR96, VRJ97, VRFJ97b, VRFJ97c, VRFJ97c, Wel94b, Wt95a, W^{+}95, Wt95b, WK95, WK96, Ygg94, Lin94]

Manuals

[?, ?, ?, ?, ?, ?, ?, ?]

Networking and communications

[ACH91, CB94, Gil93b, HH94, KW89, Kro92, NS94a, RR95b, RW92, Ste91, Ste90b, TD91, TO92] See also the extensive bibliographies in [Bee93a, Bee93b, Qua93].

Portability

[Jae89b, Lap87a, Top90a, X/O88c, X/O88a, X/O88d, X/O88b, X/O89b, X/O89c, X/O89a] See also entries in the POSIX section below.

POSIX

[BMR94, CG95a, CLJC95, DDLL94, Eck91, FGP^{+}91, GL91, Gal95, IEE86, IEE88, KMKO92, LL93a, LL93b, LL95, Lew91, Mee95, QW93, RAMP94, Rze93, Woo94, Zlo91]

Programming

[Ame89, AKW88, Ber92, Cam87, Cop92, Cur90, Dou91, ET92, ES90, HK82a, HS84, HS87a, HS91a, HS95, HS87b, Hor92, Hun85, Jae85, Jae86, Jae89a, Jae89b, Jae91, Jae93, Jae94a, Jae94b, KP84, KR88, Lap87a, Lib93, Lip91, OT91, Oua91, Paj92, Pla92, RS90, Roc85, Sch91a, Sch91b, Sch94b, SR95, Ste92b, Str86b, Str91a, TG89, Val92a, Van88, WP90, WS92, War85b, Yag91]

Programming languages: Awk

[AKW88, Dou91]

Programming languages: C

[Ame89, Cam87, Cur90, Dar88, GSN90, HK82a, HS84, HS87a, HS91a, HS95, Hor90a, Hun85, Jae85, Jae86, Jae89a, Jae89b, Jae91, Jae94a, Jae94b, KR88, Lap87a, Lev94, Lib93, Oua91, RS90, SGN88, Sch91a, TG89, Val92a, Van88, WP90, War85b]

Programming languages: C++

[Cop92, Eck93, ES90, Jae93, Lev94, Lip91, Ope94, Sch91b, SSC93, Str86b, Str91a, You92a]

Programming languages: Icon

[GG86, GG83, GG90]

Programming languages: Perl

[DP95, Ger92, Qui95, Sch93d, Til95, WS92]

Programming languages: Tcl and Tk

[Ame95, Inf95, Lib94, Ous94, SSC95, Wal95c, Wel95]

Security

[Arn93, Cur92, Far90, GS91, HM91, HR91, RG91, Sto89, WK85]

Shells

[AA86, Ari92, Art90, Art94, BK89, BK95, KW90, Olc92, Ous94, Ros91, Sou94, Val92b, Wel95]

Solaris

[BMS95, Coc94, GGM93, Gra95, HG95a, HA93, KLM92, Led94, Lei95b, OH93a, OH93b, OK94, Pew93, SSC93, Sun93b, Sun93a, Sun93c, Sun93d, Sun95, Tut93, Win93a, Win93b]

Standardization

[Ame89, IEE86, X/O88c, X/O88a, X/O88d, X/O88b, X/O89b, X/O89c, X/O89a]

System administration

[?, BS91, BMS95, BB93, Coc94, CAR93b, Cur92, FHS91, Fox85, Fri91, Gli92, HG95a, Hun92a, HH91, Kir95a, Kir95b, Lou90a, MP92, Mui95, NSS89, NSSH95, Ram94, Win93a, Win93b]

UNIX internals

[Bac86a, GC94a, Gra95, LMKQ89, MBKQ96]

X Window System

[Ado93, AS93b, AS90, Bar91a, Bea93, Ber91, Bra92, BM92, Fla94, Gra94, Gre92b, Har94a, Hel90a, Hel90b, Hel91a, JR89, JR92, JR93, Joh92, Jon89, Kel90, Kil94, Kim95, Kob91, Lee92, Man93a, Man93b, McM92, McM93, Mik90, Mik92, Mil90b, MP92, New92, Nye88a, Nye88b, NO90a, OT88, OQL88, O’R90b, Ope91a, Ope91b, Ope94, Par93, QO90, Rei93, RR92, Ros90, Ros93b, SGN88, Sch89, Seb94, She93c, Smi92, Sou94, Wom94, Yag91, You89, You90a, You90b, You92a, YP92, You95]

Title word cross-reference

+ [GS99, Ste01e]. 1 [HBH18]. 2 [HBH18]. $24.95 [Ano95b, Ano00a]. 3 [PH87]. $39.95 [Sch04, Ano96b]. $44.99 [Sta07b]. $49.95 [Jen05]. $49.99 [Hid04, Tho05]. $59.99 [Koh06]. \frac{81}{2} [Pikxx]. $89 [Gra88]. = [Ste01e]. ^{T M} [FPG89]. μ [BS90]. ϕ [Cap87]. × [Hun99]. Y_{2n} [BH1 ].

* [JN88]. *BSD [Den99].

-Compiler [PKP05, PKP05]. -dienste [WF03]. -fully [FPG89]. -mcs [CL90b]. -me [All86a, All86b, Daw86]. -mpm [KV90]. -ms [CL90b, Les76b, Les78]. -star [HBH18]. -System [BS90]. -v [Pik83b].

/70 [HR79a]. /AIXtra [AIX91]. /ARU [Hoo78]. [?]AudioHooker:1979:RTL. /bin/sh [Ree88]. /dev [Kam02]. /etc/passwd [Hoo83].

0 [Ano16b, Hid04, Koh06, Lud77b, Lyc74a, Ral78, Sch04, Sta07b, Tho05]. 0-13-101415-3 [Hid04]. 0-13-144853-6 [Tho05]. 0-13-167984-8 [Sta07b]. 0-596-00482-6 [Sch04]. 0-7821-4428-4 [Koh06]. 0.1 [Ral78, Sha79b]. 0.2 [MK84a, MK84b]. 0.3 [MK85b]. 0.6 [MK90]. ’03 [ACM03]. 046 [Sav11]. ’05 [MS05].

1 [Hub79, Jen05, LPLF91, The91, She03, Wil83a, Wil83b, ZRB^{+}93]. 1-59327-036-4 [Jen05]. 1-GHz [Ano03c]. 1.0 [Coc01b, CM78b, Hai78d, Koe78b]. 1.1 [NS94a]. 1.7 [ZB19]. 1.7-GHz [ZB19]. 1.x [KGG00]. 1/2 [Kri84a, Kri84b]. 1/2-inch [Kri84a, Kri84b]. 10 [Cal06, Cal08, Cal09, DWV06, JDB09, MM07, MMG07, SG06, San07, Vea10, VSC^{+}10]. 10.x [Pon96]. 1003.4 [GL91]. 10th [Duf90a, Kil90]. 11 [Bor14, BSW78, Bur76, Cod86, Ern11, GT78b, JF12, Jen75, Kar83, Lin79a, MH78, May78, McG79, Mue78a, Pog17, Ros78, SHLR80, Whi75b]. 11.x [Pon99b, Pon99a]. 11/44 [Tuo82]. 11/45 [LB75b, BR77, HR79a, LB75a]. 11/70 [Cla79, McG79, Sch77]. 11/780 [Tuo82, Cod86, KM79b, LR78, SHLR80]. 11D [Swa77]. 11i [Pon05a, TOML04, CM04, Pon01, Pon03, SRW04]. 11’s [GT78a]. 11th [IEE94c]. 12-step [Mil01]. 12th [MS05]. 1394 [Ale00]. 15th [NIS92, USE01a]. 16 [Arn78, Scu78]. 17 [Jos19]. 1969 [MN69]. 1973 [Fis73]. 1977 [Ric77]. 1979 [Eis79b]. 1979/80 [Lud79a]. 1982 [ACM82]. 1984 [NB84]. 1989 [Ame89, Ass89c]. 1990s [MvT^{+}90]. 1991 [Poo91]. 1992 [Won92]. 1999 [DFG99]. 19th [ACM03]. 1C300 [UNI77, Vog6 , Vog76b, Vog77c]. 1st [MM06a, Fis73]. 1Z0 [Ern11, Sav11].

1Z0-046 [Sav11]. 1Z0-821 [Ern11].

2 [Com00, Com03, Gab07, Ger92, Goe07, Har93, HK86, Jos97, LMOT02, MK04, Pon05a, Pow84a, Pow84b, SAS94b, Tho03, TNY94a, TNY94b, Wil11, Woo94]. 2-inch [Kri84a, Kri84b]. 2.* [OH93a, OH93b]. 2.0 [BF03, Böl01, Coc01c, Koc09, McN99a, PS79, Swe01a]. 2.0.1 [ISO05]. 2.1 [CV00]. 2.1.5 [Ano96c]. 2.2 [Ano00c, Bra04, VBD00, VBD01]. 2.2.6 [Ano98a]. 2.3a [Bee86a]. 2.4 [Cal00a]. 2.6 [BS05, PTS^{+}14, TCM07]. 2.x [HG95a, Gra95]. 20 [Bee87]. 200 [Cal08, Cal09, SG06]. 2000 [Bru02, FGP^{+}91, Har00, MYH00, Sha00a, USE00a, War01, Wit01]. 2003 [ACM03, Ano03e]. 2004 [ACM04]. 2006 [MM06a]. 2011 [Ano11, Gre11]. 202 [CKT80, SG06]. 21 [FR75]. 21st [IJ98]. 21st-Century [IJ98]. 22-26 [McK97b]. 22-nm [ZB19]. 23360 [ISO05]. 24th [Mil84b]. 26th [Mil84b]. 2740 [Hub79]. 2740-1 [Hub79]. 2D [GL06, Tho06b]. 2e [Str94]. 2nd [Ano86, Ano00b, USE98a, Wei93].

3 [And09, CC07, Gra05, HBG^{+}06a, HBG^{+}06b, Hid04, ML09, TV04, Wal95b, Zuc83b, Zuc83c]. 3-7908-0700-1 [Wei95]. 3-D [Wal95b]. 3.0 [ER01]. 3.1 [HHV05]. 3.5 [Lei98]. 3100 [Edd90]. 32-Bit [Ano02c, Joh79a, Ano98a, Che94, Cus92]. 32/27 [Ble83]. 3220 [RW81]. 323 [Lyo79]. 32V [Bel79c]. 354 [McI69b]. 3624 [Lud79a]. 370 [BS79a, Bur79b, Eyk86, Fel79e, FMM84, HLMR79, Pri78, Rob75]. 37th [IEE04]. 390 [Ano97b]. 3D [Gla92, Rul96, SML06]. 3rd [USE99a].

4 [AH11, Ano16a, G^{+}06a, Jen05, Koh06]. 4.0 [Ano02b, Ano03g, CBB^{+}91, Kni92, Mat07b, MF99]. 4.1BSD [CQ83a, CB83]. 4.1C [CQ83b, Lar83]. 4.2 [Cod86, Daw86, Kir84].

4.2BSD [Mor85, BH86, CKM85, JTUB85, KM83, Kri84a, Kri84b, LKM84, MK85a, McK84, O’D83a, PBL86, QSP85]. 4.3 [CTJ99, Kri91, LMKQ90, McK91, McK92a]. 4.3-BSD-Unix-Betriebssystem [LMKQ90]. 4.3bsd [Hib96, GZ89, KM86c, KTB^{+}88, LMKQ89, LM91, MKL85a, MKL85b, MK88, MKB89, MKB90b, QSP85, Tre88]. 4.3BSD-tahoe [KTB^{+}88]. 4.4 [Ano98a, Hus92, McK92b]. 4.4BSD [Com94b, Com94c, Com94d, Com94e, Com94f, Com94g, MTY01, MKS^{+}90, McK94b, MBKQ96, PW94, PM95]. 4.4BSD-Lite [PM95, Com94b]. 4.a [Gal93]. 4.x [Siv03]. 4/95 [Poz83]. 40 [Ano93-33, Dow77, Whi75b]. 40-Year-Old [Ora21]. 400 [DuC94, HR94, Tay96]. 44 [Tuo82]. 440 [R^{+}02]. 45 [LB75b, BR77, HR79a, LB75a]. 48-Bit [Rob82, Rob79a]. 486s [FJPR93]. 4BSD [JLSG84, Lan84]. 4M [CJ94]. 4th [Sch04].

5 [Ano02b, War05]. 5.2

[Hab01, JH00, Lew01, MNN04, P^{+}08]. 5.5 [Mat97]. 5.7 [dR15]. 50-and [FJPR93]. 500 [WLHC84]. 528 [FHS78a, GG99]. 5L [Mly09]. 5th [USE95a].

6 [Hon06, Lef81, Sch04, Sha81, Tho05, TU06, Web95]. 6.1 [HS99, NS94b, SPS95]. 6.2 [Mat07a]. 6000 [Cer96, CMR^{+}90, CBH^{+}92, Kel96, Lei96a, Tip92]. 622 [Lev98, RRW84]. 64 [Ano00d, Ano00f, Gig06]. 64-Bit [Cha06, ZB19, Ano03c, SGP^{+}23, Var96a]. 64-Way [Ano00d, Ano00f]. 64Bit [Ano96e]. 66 [FJPR93]. 66-MHz [FJPR93]. 68000/Unix [Gar86]. 681 [Hol75]. 6th [Ano95h, Lio96].

7 [HS00, Hon08, Lud78b, Neg01, Rob99b, Rob99c, Hag00, Win99, Ano00a]. 7.1 [Bar01c]. 7.2 [BW02]. 70 [Cla79, HR79a, McG79, Sch77]. ’73 [Fis73]. ’76 [MAT85b]. 77 [FW78, Fel79b, FW90, GG99, Mei84]. 771 [BG97]. 780 [Cod86, KM79b, LR78, SHLR80]. 7th [Ano93-31, Gue87].

8 [Ang77, Ano03a, BSW78, Bus02, Gre02, Kir77, Neg02, RS78, Rov77, Rov78, Smi99, Sob03, Sta07b, Wat02, Win01]. 8.13 [CSAJ04]. 80 [Cox83]. 8048-49 [Mul79b]. 804X [Mil79]. 8080 [Mol77]. 8080/Z80 [Vog78a]. 8086 [BM79]. 819.188 [McK91, McK92a]. 819.209 [Swi91]. 819.235 [McK92b, MK96, McK96b, McK97b, McK97a]. ’82 [ACM82]. 821 [Ern11]. 8234 [Lud77b]. 85 [Den79]. ’86 [Ano86, BBJ86, BM79, Bur76, BR77]. ’87 [BSK87]. ’89 [ACM89]. 8th [IEE83].

9 [Bai96, Cal03, Doc03, Duf90b, Dufxx, Gan08, Goo90, MGK^{+}03, PPT^{+}xx, Pikxx, Prexx, PPTTxxb, Siv03, Sko98, WN02, Win02a]. 9-Track [Rob75]. 90 [Ker93]. 9000 [Pon05a]. 90s [Hun92d]. ’91 [IEE91b]. 911 [Kir84]. ’93 [ACM93]. ’94 [Ano93-33, IEE94c]. 95 [Don95, Fos96, Fre98, LBSL97, Mer97, WS96]. 95/NT [Fre98, LBSL97]. 978 [Ano16a, Ano16b]. 978-0-12-805105-4 [Ano16a]. 978-1-59327-601-0 [Ano16b]. ’99 [DFG99, USE99b, USE99c, PKP05]. 9X [Ano95a, GMB93].

= [Ano92k, RB08].

A* [LR95]. A/UX [Mea88]. Aaron [Sch04]. Abaci [SWF^{+}22]. Abaci-finder [SWF^{+}22]. Abnormal [Ral75a]. abound [Too11]. Abraham [Ano16a]. Absolute [Luc02, Luc03a, Luc08, Luc13a, Luc18a]. Abstract [Car83a, ATS94, Hol82, LGJS91, Str80a, Str82]. Abstraction [Str84c]. Abstracts [Ede77]. Abyss [WF01]. academic [Ano95a]. Academy [ML02b].

Accelerating [Chu98]. Acceleration [IBM05]. Accelerator [OBSB23, PS22, PGW^{+}20]. Acceptance [BJ77]. Access [ALO90, AU3 , AU74, Ano92j, Bey77, FA88, Fil79, HP89a, HP89b, Kra88a, Law98, Lio79a, MC85a, MC85b, McS78, McS79, Pol93, Ric85, Roy93, RW93, Sat90b, Str90, Stu78b, Hu86, Koh99, Luc13b, MHO^{+}76, McN00, PDG^{+}87, Red97, Wei97, EM79, MC85a]. accessible [Pay84a, Pay84b]. Accessing [Far79]. Accidental [Ray99c, TD01, Ray01b]. accommodates [YHR92]. according [DD05]. Account [CKDS90, Cur91, PMD88, Moh85]. Accounting [FE77, GP88, Haz99, Rie94, Sim91, WH83]. Accounts [Arn98]. Accountworks [Arn98]. Accreditation [HG95b]. accuracy [GZ89]. Accurate [Hin88]. AccuSoft [Ano98c]. achieved [GZ89]. achieving [WTKW08]. Acid [Lew99b, Reh03]. ACLs [Str90]. ACM [ACM03, ACM04, DFG99, Fis73, IEE02, MS05, ACM75, Mei84]. ACM-SIAM [ACM97]. ACMAINT [CKDS90, Cur91]. acoustic [Kaz05]. Acquisition [BRD77, Har80, Kra78b, Stu78a]. acronyms [Lev96]. Across [MWCR90, MWCR91, MAP^{+}22, Mon00b, RT93, Xia97, IBP03, Len97, Var96b]. ACS [Ede77, Mar78g]. Action [Jan08, NR03]. activations [SS95b]. ActiveX [Kro99]. Activities [Hae90e, McC88, Sto93c, Sto93a, Sto93b, Sto93d, Sto94b, Sto94a, Sto94c, Wal92a, Wal92b, Wal92c, Wal92d, Wal93a, Wal93b, UNI93]. Activity [BG88, Eve78, Hae89, Hae90a, Hae90b, Hae90c, Hae90d, Hae91a, Hae91b, SW91a, Wal91a, Wal91b, Wal91c, Wal92e, Wal92f]. Acumen [Kro99]. Ada [ACM93, Ano95a, Ano95a, EST86, Fis86a, Fis86b, Gal91b, Gar86, Gar90, Geh84, Geh87b, GMB93, IEE92a, Lam83, Lar94, NB84, SR85a, SR85b, SHLR80, Ska94, Wan90]. ADAPT [BS79a]. adaptability [Han77]. Adaptable [AQ84, SAN08]. Adaptation [BBG85]. adapted [FRAC85, Yos85]. Adapter [ZST^{+}07]. Adapting [JAV86, Rob84b, Wat88, GG99]. Adaptive [AS93a, AS95, AS97, HK89, SJP^{+}09, XWRZ19]. ADB [MB77]. Addendum [FG80, Mat80, Sau88]. addicts [Mil01]. Adding [Ano93-32, AR07, HK79, HJAW88, Ano95d, Ano03a, DJM86]. Address [CLFL94, Che87a, Tin78c]. Addressing [FA89, FA94, Lit80]. adds [Ano03e]. Administering [Ram94]. Administration [?, Alb77b, Arc93, Bau03, BGMR91, Bor14, Cam04, CDD00, CHL16, CJ92, Cuc90, Dat88, FHS91, Fri91, Fri03, Gag01, GA04, HP98, Hes94, Hun92a, Hun00a, Hun00c, Hun00b, Hun01, Hun02b, HH91, Hun92c, JDB09, JF12, KC98, KPSS79, LH01a, McN92, NSS89, NSSH95, NSS^{+}01, NSH02, Now90, O^{+}98, Ass99a, Pon98, Pon99b, Pon99a, Pon01, Sat90a, Sat93, TSO^{+}96, USE88d, USE89d, USE91b, USE92a, USE93b, USE94d, USE96a, USE97a, USE97b, USE98c, USE98d, USE00b, USE01a, Wei04, AT07, Ano93-31, Aul01, BAC^{+}89, Bro97, Cal06, Cal08, Cal09, Col02, DeR97, FSS99, GJ08, Han01, Har00, Her93, KC00, Lev02, Max02, NSH06, NSHW10, Pon96, Pon03, Pon05a, Pon05b, Reh00, Ron05b, Sha00b, Sha01, Smi01, SS01, SS02, SS00, Tay03, TB05, WG98, Zad01, ?]. Administrative [AHPS79, Hai77, RW86a]. Administrator [Aul96, BS91, Che79b, Ern11, Hes94, Hil98, Hun92b, Kap99, Kir95a, Kir95b, Mag04, Moh99, MP92, Mui95, Sab77, SG06, San07, USE90b, Wat02, Wei98, Win01, Zie93, BDP05, BDH00, Che99, Che02, DRP01, HG95a, KD00, MM00, Par00, Pon94, Pon97c, SAS88a, SAS91b, Win93a, Win93b, Win00, Bus02, Doc03, Gre02]. Administrators [BMS95, Bia99, BDD^{+}06, Chu98, Cur92, FT09, HSM03, Lim05, Man01b, Mic08, MYH00, Sha00a, USE87c, KCTT06, Kom96, Man01a, MY03, SHN97, Bee96a, Bee96b]. administrieren [Ano03b]. Admins [Jan03]. Adobe [Mat97]. Adopt [Sta05b]. Adoption [Ksh04, CM06, Che97]. Adopts [Bar01b]. Advanced [Bor14, Cal93, CF20, Col92a, Cop92, DH05, Gar99, Gay00a, Ger02, HH91, JR90, JL16, Ker79a, MRM06, MB03, MOS01, Nor93a, OH05, Reh03, Roc85, Roc04, Ron05a, SPS93a, Smi02b, Ste92b, SR04, SR05, SR08, SR13, Ste92c, Tur87, Win01, Geh84, GS99, KPT^{+}02, LK00, Wil11, Win93a, LS87, Teo13]. Advantages [Cha06]. Adventure [Cha85, Lei96c, Lei96e, Lei96d]. Adventures [Mor88, RW92]. adversary [DLCV17, McI99]. Advice [BH86, Hai77, Mon95]. Aerospace [IEE91a]. Affecting [Rib89, SM89]. affordable [ZB84]. AFQL [CS86]. After [Ral76c, JN88, O’D83a]. Again [Gar97, PS82, Rit85]. Against [Kar92, XWRZ19]. Age [RW86b, RW86a, OS06]. AGEN [Les75b]. Agent [HP98, dSVdLT02]. Agents [Dat14, Jon02, Tho95]. Agile [PGW^{+}20]. Aglets [Jon02]. Ago [Sal97, Sal99, Sal00a, Sal00b, Sal03]. Agreement [Gar97]. Ahead [MKS^{+}90, Sim97]. AI [FES^{+}99]. Aid [AC75, Fas78, Gla79, HP98, AEG^{+}91]. Aided [Tho85a, Tho85b]. Aids [FR77, Fra78, RW86a]. Air [Rei97]. airs [Ano91b]. AIX [HS04, Kuo99, LS87, Ano96f, BGJ01, CTJ99, Cer96, CMR^{+}90, CBH^{+}92, Chu98, Coh94, DeR97, IBM88, IBM89, AIX91, Kel96, Lei96a, Mil99, Mly09, Q^{+}11, Sau17, WBSK87]. AIX/6000 [Cer96, Kel96, Lei96a]. AIXtra [AIX91]. Ajin [Ano16a]. aktuelleInformationen [Luc03b]. Alarm [Lyo76a, Vog79c]. Alaska [CLJC95]. Alberta [ACM88]. Albuquerque [ACM75]. Alcohol [LPG^{+}18]. ALEPH [CGMS89].

Alexandria [MS05]. Alfred [Wei95]. Algebra [uRH95, uRH96, Hal4 , Hal75]. algebraic [BH8 , BH79]. Algorithm [Bak76, Bak77, BK90b, BK91c, BK91b, Bro7 , Bro78, BS85, Gay87, HM75, HM76, LK73a, LK73b, Mor79a, WG00, BS84, BW94c, Che75a, HU73, MAT08, Tho68, YHR92, BG97, FHS78a, Fox79, GG99, Lev98, McI69b, RRW84]. Algorithms [ACM97, ALxx, CS90a, CS90b, KP99, Ral76b, SB97, JL96]. alike [Hol82]. all-in-one [Dul14]. Allegro [Kem91]. aller [DF00]. Alliance [Gar97, Law98]. Alloc [KSU94]. Allocation [Koc87, Mly09]. Allocator [MK88, Smi89]. Almost [Ste01c]. Alone [Eis79a, Mul79c]. Alpha [Ano00e, ST03]. Alpha 1 [Tho85a, Tho85b]. Altera [Ano03c]. Alternate [Cho98, HLMR79]. Alternative [Cod86, Wal97b, Wal98, ALP^{+}21, GLG93, Var96b]. Alternatives [Pug88c]. ALTRAN [Bro71, BH1 , BH72, Bro73, Bro77a]. Am29000 [Man92]. Amber [Car86a]. Ambiguous [AJU75]. AMD [Ano03c, PW94]. Amdahl [Wal82a, Wal82b, Wal82c]. Amendment [Mei84, Ame94]. American [You92b, Ame89]. AML [Env97a]. Amoeba [MvT^{+}90]. among [Anoxxb]. AMPL [FGK89]. Amsterdam [Gra87]. Anaheim [USE90c, USE97c]. Analysis [Aki92, Ano02c, Ano16a, ASS79, ACS79, BBH18, BP91, Ben98, Bon79a, CF20, DRN^{+}23, DB95, Eva97, GD18, HKF84, HU73, LWL05, MK93, MZA96, Per82a, Per84, PH86, SR77, She03, SHB23, Smi02a, SD95, SD87, SEKV12, Wai86, Weh77b, WB92, Yoo96, ZB19, Ano03d, BJM^{+}22, CH96, FAB^{+}07, FFPC22, HRS^{+}10, IBP03, Kaz05, KMP^{+}20, KMV^{+}09, LWL06, Mel87, MK92a, MK92b, ODH^{+}85, Pay84a, Pay84b, RKSV20, RS14, Wün13]. Analyst [Nor99a, L^{+}11]. analytic [DD05]. Analytical [Ram88]. Analyzer [Ano03f, Hol87e, Hol90b, LS79a, LS79b, LS90a, Ano03d, Les75a, Pax88]. analyzes [Ano03c]. Analyzing [CK78b, Fac02, Feu78, Kor89, Lio79d, PR96a, Ral76a]. Anarchy [TS87]. ando [DF91, MO01]. Android [Ano16a, AHA^{+}14, BBD23, HBH18, HWBH19, Jen19, KMP^{+}20, LPG^{+}18, LRB^{+}19, PRC^{+}23, SFE10, XWRZ19]. Android-Powered [SFE10]. Android-related [LRB^{+}19]. Anecdotes [TLR89, TBGE89]. Animation [BK90b, BK91c, BK91b, BS85, BS84, PH87, ZL10]. Ann [USE92c]. Anniversary [Bro19]. Annotated [ES90, McI87]. Announces [Ano03f, Bar00b, Ano03d, Ano03e]. annoyances [Jan06a]. Annual [ACM75, ACM97, Den99, USE00a, USE01b, Ano93x, Fis73, USE96b, USE97c, USE98b, USE98e]. anomaly [DFC^{+}23]. ANS [McI71]. Ansatz [Dav97]. ANSI [Ame89, PKP05, Ban88, BBD91, Kut02, PKP05, Pug88a]. Answer [BW94a, BW94b, NN92, TG89, LM91]. Answers [Ano92c, Ano92d, Ano92e, Ano92f, Ano93a, Ano93b, Ano93c]. Ant [ZK05]. anthology [HA95a, HA95b, HA96]. Antonio [USE92b]. Anwender [TBR93, Wei95]. Anwenderhandbuch [Ron05a]. Anwendung [Bor01, PKP05, WF03, GA04]. Anwendungen [PKP05]. Any [Hil93]. Anyone [Ano94b]. Anything [Hae92, SAN08]. Aonix [Kro99]. Apache [Aul01, Gab07, G^{+}06b, Hun01, LW03, Reh03, TSC02]. API [Ame94, Ano00e, IEE92a, IEE92b, Mer01, PBS98, SFR04]. APIs [Ste98]. APL [Yao83]. Apollo [SSWW83]. APP [Mil93]. Apparatus [WSS77, WSS78]. appealing [Gra84]. Appendix [Ano96a, MC75a]. AppForce [Ano03d]. Apple [Ano01, Ano02b, Ano03c, Gem05, Mea88]. AppleScript [Ros06]. AppleTalk [Lun96]. appliance [SHPP07]. appliances [SHPP07]. Application [AJ05, Ano01, Ano02b, Ano16b, Arn78, Bea93, GDFR90, Gre19, Hor82, HD98, Ima79, IK83, JT98, Kau7 , Kau79a, KSU94, KI82, Li98, McG97, Roe02, SM89, Ste90a, UTC84, Ver89, Wal97b, Wal98, ZB19, Ame94, Ame91, Ano03d, Fac02, FAB^{+}07, Fel4 , Fel75, Gab07, Gov08, IEE92a, IEE92b, JPS00, JT05, Jon05, PRC^{+}23, SS95b, Sun93a, Thi99, DD05]. Application-Class [ZB19]. application-layer [Ano03d]. Application-Level [KSU94]. Applications [?, ?, ?, ?, Abb03, Ano02a, Ano02b, Ano16a, Asb00, Bol99, Böl01, BH8 , BH79, Dun91a, GM02, GS01, HS82, Har94a, IEE91b, JR89, JR90, JR92, Joh96a, Kel90, Kep85, Les79a, Loz05b, LMT78, Lyc78a, Mar04, McS78, MCM06, Nee90, NHNW94, Noe98, Pad93, PR96a, PK93a, PK93b, Rez95, RW92, RT93, SW82, She01, Shi92, SS93, SEKV12, TBS87, USE94c, WLL96, Wal94b, Wat88, You89, You90b, YP92, Ano93r, Ano94i, Ano95h, Bro98, BGW04, Cha11, CS93, CS96b, Dal99, Dal02, DH96, FJPR93, FGP^{+}91, GS02, Gil96a, Gla91, GB01, Gue19, GMTW08, Hen92, Ivi95, Cor05a, Lar94, LZHS11, Mad94, MHS02, MCM07, PBM^{+}22, PRC^{+}23, RS14, SML06, Seb94, ST03, SZ95, Spe00, Swi91, Yeo05, ZGdDL^{+}96, Ano93h]. Applied [Kur94, Kur97, War85b]. Approach [Cox09, FKM95, GJLB77, JJ91d, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ92e, JJ92f, JJ92g, JJ92h, JJ92i, JJ92j, KmWH10, KAH83, Kur90, Mud97, New92, Pat96, PC90, Ric94b, Sin97, Sta07b, SCK86, ALP^{+}21, Afz03, AGP^{+}92, Ang97, Ang99, BBD23, Com12, DD05, Hal06, Hal07, Hon08, Joh05, JJ91o, RFS05, RFS06, SML06, Sin07, van83]. approaches [And17]. Appropriate [Jos95]. Approximate [Ben98, McI83]. Approximating [HKT05, Pol93]. approximation [Hal4 , Hal75]. Apps [Wal97a, AHA^{+}14, DLCV17, DC18, HBH18, HWBH19, KMP^{+}20, Red97, Sim96b, Tay96, Var96b, WWCW19]. April [Ano93t, IEE94b, MN69, Uni98a, USE87c, USE93d, USE94c, McK92a, NB84, Swi91]. APS [Ede77]. APT [DC18]. Arbitrary [CM78a]. Arbor [USE92c]. ARC [Env97a, Env97b]. ARC/INFO [Env97a, Env97b]. Architect [CR02]. Architecting [Blo99]. Architectural [DNQ^{+}83, Mei92]. Architecture [BBP82, Buy99, CS91a, Dun91a, Fel84, HQZ^{+}87, IB93, JZTB98, JF81a, Kle86, LBN14, LA90, MM01a, PS22, Req85, SC78, SA21, Tra85, Wat87, AGP^{+}92, CS96a, Gra95, Kel96, KS08, Mau08, MM07, Vre03, Wil98, Eyk86]. Architectures [HD02, LC90, PN04, Sch94b, Sch95b, IEE91b, RFS05, RFS06, Sun01, TF04, Luu90]. Archiving [Vog78b]. Ardent [Cod89]. area [SZ95]. arena [Ale96b]. argue [Ano95d]. Argument [Koe78c, Loi79]. Arguments [Lio79d]. ARIEL [HK83]. Arithmetic [Mor88, Sch81, SU70, TOML04, Dit93, LMOT01, LMOT02]. Arizona [USE87d]. ARM [Jor04, GNDB16, Jor04, TTB09]. ARM-based [Jor04]. ARM-basierender [Jor04]. Armin [Wei95]. Arnold [Ano00a]. Arpanet [PB21]. Array [All78, Ber85a, Col83, HCN85, JTUB85, DJM86]. array-processor [DJM86]. arrays [dVd93]. arrive [Lei90]. arsenal [Blu09]. Art [Ano93-29, Dec04, Lim05, Ray04, HK05, Hin88, WB06, Zdz05]. Articles [NCNS82, Hub96]. Artifacts [Too09]. artist [Ham07]. artists [Ham07]. Artwork [Fas78]. ARU [Hoo78, Hoo79, Hoo78]. AS/400 [HR94, Tay96]. AS/MERT [H^{+}79]. Asbestos [Lew99a]. ASCII [Mud97]. Asia [Uni98a]. Asian [CM06, PB21]. ask [Ano93r]. Asked [Tim93, Pub11, Vib11]. askew [Mon95]. asking [Tay96]. Aspect [Sat93]. Aspects [FKN85, GB92, Mei92, Set81, Set82, Set83, NHR84]. Assembler [Kir77, Mil79, Mol77, Mul79b, Rit2 , Tho79, Vog78a, Weh77a]. Assembly [Ang77, Dun00, Kap92]. Assessment [Ano01, AHPS79]. Assist [BK85]. Assistance [ASS79, GJLB77]. assistant [WT91]. Assisted [DHS78, Sha77]. Associated [Les75b]. Associates [Ano95b, Ano96b]. Association [Sof83, Sof84, SD95, Usr82, USE85c, USE86e, USE87e, KL92]. Assorted [Don96a]. Assumption [KKF^{+}84]. Assuring [AH11]. Asynchronous [Lyc74a, Mue79, SGM^{+}00, Tin78a, Yoo96]. AT&T [Bas87, AT7, Bai96, Bas87, EST86, Isl83, Jun85a, Jun85b, ?, ?]. Athena [Tre88]. Atlanta [USE86c, USE86a, USE91e]. ATLAS [Hor82]. ATM [ATS94]. Atomic [Ano03f, BLH20]. ATS [HO 78, Smi78]. Attaching [Col83]. Attack [KI99, Loz05a, Ale96a, Ras07]. attacking [PM10]. Attacks [Duf89c, Chi02, HMK^{+}20]. attribute [Wei97]. attributed [van83]. Audacity [JP09]. Audio [SW91b, TP91, Ran06, Hoo78, Hoo79]. audit [Ano93g, Gro93]. Auditing [Bis88, May97, Far92c]. Aufgaben [Die11]. Augmented [Bra79, FDMS79]. August [USE88e, USE88a, USE88b, USE90e, USE93e, USE94b, USE97a, USE97d, USE98c, USE98a, USE99b]. ausgereifte [Bra04]. Austin [USE89d]. Australia [Tob80]. Authentication [Bel18, Coh90, ZG96, Gan08]. Authmgr [Coh90]. Author [Jae86, Mor79c]. authoring [Web95]. authorization [GJJ05]. Authorized [The02, Jos97]. Autoconf [VETT00, Cal10]. autoconfiguration [JJ91a]. Automake [VETT00, Cal10]. Automated [BY96, CDD^{+}07, Des96, LL93a, Lut99, Mue79, CJBT92, Ano16a]. Automatic [Blu75, Blu77, Fol77, GJJ05, Koe84, Mar01b, NB84, FM02, JL96]. Automatically [Hob91, Hob93, Ale96a]. Automatically-Tuned [Hob91].

Automating

[Bau03, GS00a, KM95, Lib94, PLHM08]. Automation [Wat88, Lum13, Ros06, Poz83]. automatisieren [RB06a]. Automounter [PW94, Zad01]. Autotools [Cal10]. Autumn [Ano89, USE87a]. aux [RB05b]. Availability [Ado89, BS92, IBM09a, SHB23, Wey01, Ivi95, Kob99]. Available [BW94a, BW94b, McK98, Sat90b, Kop05]. Aventail [Ano98c]. Avilon [Kro99]. AVP [Ano00e]. Awards [Pin99]. Aware [XZY^{+}10, HKVR21, MHS02]. Away [Den99, Sta07a, Gil96a]. awesome [Gri14]. Awk [AKW85, Bai85, McN99b, Ora21, Sta89, Dou91, PWC00, Rob02b, Rob02a, AKW88, AKWA89, Ano85, Kaz05, Lev83, Mor89, Rob01, AKW78, AKW8 , AKW79, Art99, Bla02, Bla04, Kruxxa, Sta89, Ano00g, DF91, DR97, Pat99]. Axis [Dou02]. AXP [SAS94b].

B [JK73, Joh7x, Ker7x, MP85, MC75a, Sch76b, Tho72, Wei81]. B-splines [Sch76b]. B-trees [Wei81]. B.1 [Mas77a]. B.10 [Rit77b]. B.11 [Gla77a]. B.12 [Sab78]. B.2 [PS77]. B.5 [Mas77b]. B.9 [Ker74]. B1 [ST89a, ST89b, ST89c, SM89]. Baan [Kro99]. Back [Ano02b, AGHR89a, AGHR89b, Ell75, Wei90]. Back-Office [Ano02b]. Back-Up [Ell75]. Backbone [HIT^{+}02]. Backend [Duf82]. Background [McK02]. Backing [Jon98]. Backout [Bur79a]. backplane [IEE91b]. BackTrack [AH11]. Backup [Ano00b, Ano00e, Eng91, Hai78c, Hum88a, Man01b, McG79, Poe87, Pre99, Pre07, SKB00]. Backups [Abl93, HD97]. Bad [BK13, Cha85]. baked [Par90]. Balance [TGF88]. Balancing [HBD97, Höf01, HCG^{+}82, Kro99, Mon00a, BGW93, Jal92]. Baltimore [ITC93, IEE02, NIS92, USE89e, USE89a, USE92e]. BANCS [Lic79a]. Banff [ACM88]. banking [Ano97b]. BarbequeRTRM [BMF15]. Bargain [Rol03]. Bargain-Basement [Rol03]. Barrier [Che97]. BASAR [Tho95]. Base [Coc01b, Gel78, GR76, Hoo83, ISO05, LMT78, Nor93b, NS94b, Par78, SPS93b, SPS95, TT79, Dup05, RS14, Tin79a, Cor05a, Yeo05]. Based [AISS98, Alf94, Ano01, Ano02c, Bar86, BAKW92, BNB87, Cha79, Cho98, Chr90, Chu03, CPJ^{+}98, CBL96, Den79, FE77, GVK^{+}03, GLOR91, Gwo78, HQZ^{+}87, HP89a, HP89b, HLT^{+}21, HCG^{+}82, JH86, Koc87, KKF^{+}84, Lo79, Mac97a, MTY01, MKB90a, Mic79b, Mor96, Mue79, OT07, PNK02, RB82, Skl91, SKP94, Spi00, STT86a, STT86b, TOML04, Tra85, Ves88, Wam83a, Wam83b, Ano97b, Ano98a, Ano03d, Ano03e, Ano06a, BBD23, CM06, Cha96, Che90a, DI89, Gil96a, Gra04, Gra06b, Gra87, HMK^{+}20, Hes94, HK86, Jor04, KLP88, Lac06, LM88a, Lib94, LCH^{+}81, MHO^{+}76, MKSB22, NLR84a, NLR84b, PBM^{+}22, Pon05b, PKP05, Sun01, Sak97, Sam87, Sam89, SB08, Thi92, Van05, WN15, WT91, Wil98, YL09, LHM^{+}83, McD84a, McD84b, Per87, PY84a, PY84b, WH83, GK97, HK83, KE85, KK89b]. based [LM88b, MBC^{+}86, Mel87, SR85a, SR85b, Sti83]. Basement [Rol03]. Bases [LT80, PW79, LT80]. Bash [AVN07, May01, Die11, NR95, Smi06a, Bla02, Bla04, Bur04, Gün07, Mic08, NR98, NR05]. Basic [BF9 , BF80, Car84, CS90a, CS90b, FHS7x, Law98, Pol93, Ric85, JJ91s, JJ91b, JJ91c, JJ91a, Arn78, Scu78]. Basic-16 [Arn78, Scu78]. Basics [JR90, Lav07, RG91, Hsi01, Mus02, Rei98]. basierenden [Lac06]. basierender [Jor04]. basierte [Ric14]. Basis [Fis86c, FH88, GA04, PKP05, Tev87]. BASIX [LHM^{+}83]. BASS [Kau93a]. basshu [MO01]. Bastille [Bea04]. Batch [Bra95, Bri01, Ger92, Alb76]. Batteries [Coc01b]. Battery [McG79, Sta04]. Battery-powered [Sta04]. battle [TBG^{+}87]. battlefield [WCHS93]. Bay [DD05]. Bayesian [Zdz05]. Bazaar [Ray99c, Ray01b]. BBS [Wei88, Wei94]. BC [CM78a, Dit93]. BDP [AOAH14]. Be [Cox07, FES^{+}99, Pik88, Ale96b, Ano93v, Gil96a, Gra79b, Per93, Pub11, Vib11, Rit85, Sim96c, Tay96, Var96b, Gia99, Coc01c]. Beach [USE92a]. BeagleBone [BK13, Dat14, Gri13, Gri14, Lum13, Mol14, Ric13, Ric14, SP15, YK15]. Bean [Bea93]. bearbeiten [GGZ02]. Become [AAG^{+}16]. Becoming [Hil98, Bal96]. bedfellows [Tuc98]. Bedienung [KGG00]. been [Sch95c]. Beginner [PS77, RGRG05, Max02, Pet01a, Sha00b, Sha01]. Beginners [Ker79d, Rov77]. Beginning [Ben09, Buc06, FJWA05, Joy97, ML09, Pec08, Ron05a, See08, Tho06a, Afz03, Sau17, Ste99a]. begins [Jen19]. behandelt [Die11]. Behave [Swa00a]. Behavior [WCC^{+}12, BP91, KIT93, PB91]. Behavioral [Fra83, Gin83, Mac83]. Behaviour [LA90, TTB09]. Behind [Fro90, RP84]. Being [ASW^{+}01, Gil96a]. Beispiele [PKP05]. Believe [Wal94a]. Bell [PPTTxxa, Car88a, Car88b, Cla79, Geh03, HB82, Mil84c]. Ben [Ano00b, Gra89b, RB08]. Benchmark [MGK^{+}03, Mil86]. Benchmarking [Tuo82, Ano94a, Hin88]. Benchmarks [Sax85a]. benefit [Var96a]. Benefits [CPJ^{+}98, You93]. Bensberg [Wei95]. BeOS [Bro98]. Beowulf [Hul05, GLS03, Ste01a, Ste02b]. ber [DF00]. Berkeley [Ano24, Bra98, Gt92b, BJP04, Bas91, Cab83, CKM85, Cha85, Gil88, GZ84a, GZ84b, GZ89, Hor92, KM86a, Kir84, McK85a, MK87, MKS^{+}89, MMS86, O’D83a, Sha93, Skl91, Tre88, ZDS85]. Berkeley-Software-Distributionen [Bra98]. Berkeley-to-Boston [Cha85]. Beschreibung [Bas87]. Best [CHL16, KL94, Lav07, McI83, RRH94, FJPR93, HA95a, HA95b, HA96, Gla92]. bestselling [Ano98e, Ano00g, RRD^{+}03]. Bet [Bar00c]. Beta [Lon78].

Betriebssystem [CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, LMKQ90]. Betriebssystemen [Bas87]. Better [Coc01b, Ded09, Law98, Pla89, Ano96e, Com00, HM98, gL02]. BetterBASIC [ASW^{+}01]. Between [Joh96a, LTW77, Liu96, Mul79c, Rib89, Woe03, Dul96, Hor90b, Joh95, Pan94, SL95, SS95b, Wei92]. Beyond [BW88a, Cam91, CUH99, JR90, Kar06, Mil93, SH85, Ste88, Mus02, Sun01]. Bible [Ano00b, Blu11, Goe00, Gri00, Neg01, PM87, PR96b, Blu08, DWV06, Gri01a, HHV05, Neg02, Neg06a, Neg07a, PWN01, SJM09, W^{+}95, vGS10, vH06, vH07]. Bibliographic [AC75]. Bibliography [Bee93a, Bee93b, Bee94, Bee96a, Bee96b, Lee92, Sch80]. bid [Ano93m]. Bidirectional [BBG85]. Big [Ano84, Cha85, Mac97a, Sta04, Gar99, Miy86]. bigger [Ano96e]. biggest [Ree88, Sim96c]. Bildbearbeitung [DF00]. Bilder [DF00]. bilingual [Wal99]. bin [Ree88]. Binaries [And91, MM06b]. Binary [CWS90, Gar90, Hub02, Lew99a, Lew99b, RP84, Ame91, Gil91]. BIND [AL92b]. BIND8 [Siv01]. Binding [IEE92a, IEE92b]. Bins [Weix ]. biodata [Wün13]. biological [CH96]. biology [Wün13]. Biometric [Coc03b]. Biopolis [Ano06a]. Birth [Too11]. BISON [DS88]. Bit [Ano02c, Cha06, Joh79a, Rob82, ZB19, Ano98a, Ano03c, Ban82, Che94, Cus92, Rob79a, SGP^{+}23, Var96a]. Bitmap [BEHW86a, BEHW86b, Pik82, PLR85, PLR00, Pik83a]. Bitmaps [Hob91, Hob93, MD86]. Bits [Var96a]. Black [Bai00, BK13, Sta04, BS02, Hin88, PRC^{+}23, Gri13, Gri14]. black-box [PRC^{+}23]. BlackParrot [PGW^{+}20]. Blair [Sau88]. Blender [JP09]. Blind [But02b]. Blit [Car83b, Car83c, Car85, Pik84a, PLR85, PLR00]. Blits [Pik84b]. Block [AS93a, AS95, AS97, Rob84a, RDK75, Tin78a, BW94c]. block-sorting [BW94c]. blocks [NIS92]. Bloom [Cas94]. BLOSIM [RDK75]. Blue [Gar99]. Blueprint [Cov06]. BNR [BCP93]. Bob [Nel90, Kar84]. Bolton [ACM03]. Bolts [Ano99c]. Bon [Tho69]. Bone [BK13]. Bonkers [Ric94a]. Book [Ano95b, Ano96b, Ano00b, Ano00a, Ano16a, Ano16b, Bai00, BR82, Bea98, Chr98, Col92a, Cuc90, Elb01, Far92a, GRA89a, Gwy90, Hid04, Kad95b, Kad95a, Lea90, Lea92, Mad94, Nag05, NN92, O’R90a, Pon99b, PKP05, Ree06a, Sal03, Sch04, Sim93, Teo13, TG89, War90, Wei95, Wri93, Zwi92, Ban88, BBD91, BS02, CDM98, CC07, Coo03, Han08, Han11, Hea98, HK05, HBB^{+}06, H^{+}07a, H^{+}07b, Kir09, LD13, LM91, Pon94, Pon96, Pon97c, SAN08, War02, NN92]. Books [Bee93a, Bee93b, Bee94, Lut99, O’R93, Qua93, Sch04, Ano98e, Ano00g, RRD^{+}03]. Bookshelf [GS00b, Gri01b, JL98, Ass99b, Ros98, SWL^{+}01, Woe02, Woe03, Ano98e, Ano00g, FWS^{+}03, RRD^{+}03, WLS^{+}01]. Bookworm [Sal03]. Boolean [Bon79a]. Boon [Ano94d]. Boost [Ano92i, Kar06]. Boot [How01, Neg06a, Neg07a, Kob99]. boot/root/swap [Kob99]. Bootable [SGP^{+}23]. bootables [Neg06c, Neg07b]. Bootstrap [Reg11]. Borders [Coc03b]. Borland [Lis02]. BORPH [SB08]. BOS [Ima79]. BOS/LMOS [Ima79]. Boston [ACM82, Ano86, Ano94m, Ano95h, Fis73, Usr82, USE82, USE94e, USE01b, Cha85]. Both [Whi75b, Ale96b]. Bottlenecks [Bac92]. Boundaries [MWCR90, MWCR91]. Boundary [Blu7 a, Blu78a, CW89, CW91, Cou84, Cou85a]. Bourne [Die11, May01, Bli96, May01, Mic08]. Bourne-Shell [Die11, May01]. Boutell [Koh06]. Box [Ano00e, BS92, PRC^{+}23]. Boyer [NT05]. BPF [CF20, Gre19]. BPSS [Kau79b]. Brain [Bar01b]. Branding [Lin95]. Braun [TBGE89]. Brave [Ano97a]. breakdown [Wal99]. Breaking [Che97]. breakins [Rei86]. breakthrough [Sun01]. bred [Whe97]. Bridge [Ano00e, Ano02c, TLK99]. Bridges [UTC84]. Bridging [ACM04]. Brief [McK07, McK16]. Briefs [Ano93s, Gar97, Gar99, Gla99, Law98]. Brighton [Ano95a]. Brilliant [Ric94a]. bring [VMKB05]. Bringing [Ano93r, RW86b, RW86a, TS87]. Brings [AGHR89a, AGHR89b, Wol93]. Brother [Mac97a]. Browser [Law98, Put86, Ham90]. Browsing [Law98, FFPC22]. Brute [KI99]. BSD [Ano98a, BBJ86, Bas87, Cod86, DF00, GGZ02, Hus92, Kuo99, LMKQ90, Lin02a, Mad94, May01, McK91, McK92a, McK92b, NN92, WF03, Ame90a, ASW^{+}01, All97, AW89, Apo87, Bas87, BE88, BE89, Bra98, Cab83, CQ83b, Cha85, CL06, CS93, CS96b, Cor05b, Daw86, Fer85, Gla02, GGZ02, Guy00, Hub96, Kon07, Kri91, LKY93, Lar83, Lav04, gL02, Lin02a, Luc02, May01, McK85a, McK92c, McK98, McK07, McK16, McN99a, NC08, O’R90a, ODH^{+}85, Pla89, SSB92, Skr89, SS95b, WLC95, YLK93]. BSD-UNIX [Hub96]. BSD-Unix-Nutshell-Buch [O’R90a]. BSD/XENIX [Ame90a]. BSDCan [ZM08]. BSDCon [USE02]. Buch [DLvG07, KGG00, O’R90a]. Buddy [Koc87, Gra88]. Buffer [CKM85, LBN14, Pea80, AKB85, BRW89]. Buffered [Str79]. Buffers [McC90, Tin78c]. Bug [CPJ^{+}98, Kie99]. Bug-Free [CPJ^{+}98]. Buggy [CPJ^{+}98]. Bugs [BLH20, CCVLW23, Fil85, All97]. Bugzilla [ZK05]. Build [Asb00, HH97b, Gue19, Kop05, MD86, SML06]. Builder [Coc03b]. Building [ALO90, ACK89, Ano00b, Art03, Bau02, BK85, Don96b, Dun91c, Eck95, Fel96a, Fre91, GM02, Har94a, HL95, Hon06, Hon08, ILG19, KJG^{+}98, Cor05a, Luc05, LSC^{+}88, MLRC88, Seb94, SY00, Spe00, Tho05, TNY92a, TNY94a, TNY92b, TNY94b, Too17, UTC84, Yag03, Yeo05, ZCC00, Boo03, Gri14, Jan03, Mol14, NIS92, Neg06c, Neg07b, Sch90, SHPP07]. Bulk [Cla00]. Bulletproof [Got02]. Bunk [Ano94d]. Bus [Cer83, FJPR93, Lyc72, Paj88, Ano03c, Cer82, IEE91b]. Business [Ano93-27, Ano02b, HS82, Lat02, Lei99, SH99a, Wil82a, Wil82b, YE84, Var96b, YEB84]. Business-Oriented [Wil82a, Wil82b]. busy [SZ95]. Buttons [Hes91]. buy [Gra88]. Bye [Coc01c]. BYTE [FJPR93]. Bytes [Sem96].

’C [And80, Ame89, Ano02a, BBD91, DF00, Joh83, Kin89, LK00, Lio79e, Mar78d, Nel90, PKP05, Row79a, Swa00b, TG89, Ame94, ABS95, AA86, Ano95f, Ano01, Ano03e, Ano19b, Ari92, BBH18, Ban88, BBD91, BW96, Bec06, Bec07, Ber92, Bha12a, Bin85, Boy84, BSW78, But02a, Cam87, C^{+}78a, Car77, Cha97, Cme86, CB89, CK78b, Cop92, Cox83, Cro78b, Cur89, Cur90, Dar88, DYZC20, Dun91c, Eck93, ES90, Far79, Fed79, FGMS90, FGMS95, FG79, Feu85, Fis86a, Fis86b, FJ81, FF88, FR75, Fra79, FH96, Fra99, Fre78, GJ75, GSN90, Gil88, Gla77b, Gla79, Gra79a, Gra79c, Gru86, Hae89, Hae90a, Hae90b, Hae90c, Hae90d, Hal00, HK82a, HS84, HS87a, HS91a, HS95, HS02, Hel03, Hol12, Hor90a, Hun85, HC89, Jae85, Jae86, Jae89a, Jae89b, Jae91, Jae93, Jae94a]. C [Jae94b, JRL77, JR78, Joh79b, JR81, Jos99, Jos12, Jos19, Kar06, KLP88, KK89a, Kaz05, Ker74, KR78, KR88, Koc09, Koe78b, Koe78a, Koe78c, Koe79a, Koe79b, Koe89, KGP90, Kul06, Kut02, Lap87a, Lap87b, Lef81, Lev94, LH00, Lib93, Lio79d, Lip91, MLRC88, NH00, Ols73, Ope94, Osb92, Osm96, Oua91, Oua92, Pac00, Per79, Pla92, PKP05, Put86, RS90, Rad79, Ree06b, Ree82a, Ree82b, Rem73, Rit74, Rit77a, RJLK78, Rit79a, Rit8x, Rit93, Rov77, Row79a, Row79b, Sab76b, Sak97, Sch93a, SGN88, Sch91a, Sch91b, Sch00, Sch93b, Sch90, SML06, Sea05, Sea09, Sea14, Sea20, Set79a, Set79b, Set79c, Skj83a, Skj83b, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84f, Sny75, Ste85, Ste00, Ste01c, Ste01d, Ste01b, Ste01e, Ste01f, Ste01g]. C [Str80a, Str80b, Str82, Str84a, Str84b, Str84c, Str85, Str86b, Str91a, Str13, Swa74, TH77, Tar89, Thi92, Thoxx, TG89, USE88h, Val92a, Van88, Vol97, WP90, War85b, You92a, dG93, dG96, vdL94]. C-ISAM [Tar89]. C-Kermit [dG93, dG96]. C-Language [FR75]. C/C [PKP05, Ano03e, Vol97]. C/POSIX [Ano19b]. C/UNIX [Skj83a, Skj83b, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84f]. C11 [JP17]. C2 [Eps92, HJAW88]. CA [Ano95b, Ano96b, Ano00a, IEE03, USE93b, USE93g, USE94d, USE97b, USE00a, Ano93-31, Uni96a, USE83, Ass83, USE86d, USE88d, USE89b, USE89h, USE01a, USE02]. Cache [AKCB86, Che87a, LBN14]. caches [BRW89]. Caching [Hau94, Roy93, Sch94b, Sch95b]. CACS [Sha77]. Cactis [HK89]. CAD [HQZ^{+}87, SAL93, TLK99]. CAIS [CSA^{+}87, Fis86c, GBH86]. Calc [Hab01]. Calculated [Coc01b]. Calculates [Coc01c]. Calculator [CM78a, MC78]. Caldera [SRBB99]. California [ACM92, Ano93x, Uni96b, Uni98b, USE89f, USE90c, USE91b, USE92a, USE93f, USE94b, USE94f, USE96b, USE97c, USE92d]. Call [BS88, HG93, Kar83, Loi79, MS90, PR76, Rod86, Ryd8 , Ryd79, Whi75a, BBD23]. call-based [BBD23]. Calling [JRL77, JR81]. Calls [Can80, MMTW88, Vog79b, BHR95, Lar83, Lov07, ?, ?]. Calypso [DKM96]. CAMAC [BRD77]. Cambridge [Ste90a, USE87b, USE93e]. camel [Wal99]. Camera [Dou02]. Campaign [Coc01a]. Campbell [BH1 ]. Campus [MTT83]. Can [Ano93l, Bar00c, Cox07, Hae92, Jen00, Les88a, Lin93, O’D83c, Rit85, Ros95, Ste01d, Biv87, Rze93, SZ95, Var96a, Var96b, Ano02c, Ano93d]. Canada [ACM88, Ano00e, USE94c, Sof83]. Can’t [Mui95]. Capabilities [May79, Sma88, WALK10, WALK12]. Capability [Erl88, JKS^{+}15, Tra85]. Capable [OBSB23, VTS^{+}23]. Capacity [Cha79, Chr98, Coc01b, CW00, RWNA87, Stu78a, FSW02, Won98]. Capsicum [And17, WALK10, WALK12]. Capturing [Pal87, PDG^{+}88]. Card [Che79d, SSC95]. Cards [SSC93]. Care [Ham87, HBB13]. CAS [Fun87, GVK^{+}03]. CAS-System [GVK^{+}03]. CASE [Ano92b, CKM85, Des96, Ham87, Hei95, Kul06, Lei99, MAP^{+}22, MQ09, MT78, MT79, SG09, SA21, TTB09, Wei92, BTS09, FAB^{+}07, FFPC22, Ger03, IF10, KWDB06, MW08, War85a, Keu92]. Casper [BS88]. Cat [Strxx, Pik83b]. Catalog [Kro92, Les88b, O’R93]. Cataloging [Cal93]. Catching [SJP^{+}09]. catering [Tuc90]. Cathedral [Ray99c, USE02, Ray01b]. Cautions [Lyo76a]. CAVES [Fil79]. CB [De 78, Kau78, Mic79a]. CB-UNIX [De 78, Kau78, Mic79a]. CC [CHRG95]. CC-NUMA [CHRG95]. CCA [Com84, CCA84]. CCS [MS05]. CD [Ano00e, DF00, PKP05, AL97c, Ame95, Ano98e, Ano00g, Com94b, FWS^{+}03, GGZ02, HS99, Kom02, Kuo98a, Mei92, Ass99b, PKP05, RRD^{+}03, Ser00, SWL^{+}01, WLS^{+}01]. CD-I [Mei92]. CD-ROM [Ame95, Ano98e, Ano00g, Com94b, Kom02, RRD^{+}03, AL97c]. CD-ROMs [HS99, Kuo98a]. CDE [Sch95c, Mio97]. CDROM [HS99, HS00]. CDs [GGZ02, Neg06c, Neg07b]. CE [Ano03e]. CE-based [Ano03e]. Cedar [ADH^{+}89, Don85]. Celebrating [Ano19a]. Center [Ano98c, AHPS79, CK78a, Hoa77, IEE91a, ITC93, NIS92, Ric77, Sne01, MRM06, Cha79, MGK^{+}03, Shi77]. centralizada [VD01]. centralized [VD01]. Century [IJ98, Kol94, Sal94, Til87a, Til87b]. cert [Tom11, Sea09, Sea14]. Certifiable [ST89a, ST89b, ST89c]. Certification [Mag04, Ano95g, Dea01, Eck11, Mag00, P^{+}06, Sul02]. Certified [Bus02, CR02, Doc03, Gre02, Mag04, Pri99, SG06, San07, Sav11, DRP01, Jan04, Reh00]. Certifying [SK04]. CEST [FF88]. CFD [GVK^{+}03]. Cg [Ano03f]. CGI [BA96, DH96]. CGI/Perl [BA96]. CGO [IEE03]. Challenge [Hos83, Sta06, Wil94]. Challenges [Cha06, SEKV12]. Champion [Gre92a]. Change [BED^{+}85, Dun91b, KPSS79, Siv01]. Change-Rooted [Siv01]. Changes [Koe78b, Row79a, Row79b, Tre88]. Changing [BED^{+}85, Edw98, Mas88, Sal08]. Channel [DD05, GK97, WB92]. channels [CMR^{+}21]. Chapter [Web95]. Character [Car77, Mic79a, Sch97, Sun96, Che97, SAS93a]. Characterization [HLLP79, AJ05]. Charles [Ano00b]. Charts [McI71]. Chase [Ano93e]. Chassis [Ano00e]. Cheap [Bin85, Gra98c]. Check [Ano96d, Kow79, Sax85b, Sax85c, Bin88, Kow90]. Checked [DYZC20]. Checker [Joh78b, Kro99, MDAD^{+}14]. Checking [Dar88, HT92, HT94a, HT94b, Lio79a, Ube76]. checkpoint [BBHL08]. checkpoint/restart [BBHL08]. Checkpointing [LS92, PBKL95]. checks [Fra90b]. CHEM [BJK92, BJK87]. Chemical [BJK92, BJK87]. ch‘eng [PWC00]. Cheshire [OBSB23]. Chicago [USE96a, USE99c]. Chinese [IBC87, Lun99, PWC00]. Chinese/Japanese [IBC87]. Chip [Bar01b, Gar97, Law98, Don04]. Chips [Coc01b]. choice [Bal96, Lar94]. Choices [FJPR93, KM83]. Chon [PB21]. choose [Ale96b]. Choosing [Far91a]. CHORUS [AGP^{+}92, BCC^{+}91]. CHORUS/MiX [AGP^{+}92]. Christopher [Sta07b]. Chrome [FFPC22]. Churchill [Ste90a]. CICS [Ano94b]. Cider [AHA^{+}14]. Cin [KGP90]. Cincinnati [USE93c]. Cip [Bro90]. Circles [McI83]. Circuit [FR77, Fra78, LCH^{+}81]. Circuits [SK72]. Circular [Mur83]. circumstances [Rit85]. circumvent [DC17]. Circumventing [DC18]. Cisco [ML02b]. Citius [Ano00c]. City [Sof84, USE84a, USE95c, USE95a]. CJK [Sun96]. CJKV [Lun99]. CL [Kem91]. CL/TB [Kem91]. Clara [Ano93x, USE93g]. Clark [Pol93]. Clark-Wilson [Pol93]. Class [Bai99, Cro98b, MLRC88, WL94, ZB19, Ano95f, Ano19c, Bra95]. Classes [Str80a, Str80b, Str82]. Classic [Bri01, Lew21, MZH22, RB05a, Too11]. Classification [Vog76b, DD05, SWF^{+}22, Zdz05]. cleaner [Gat01]. Cleaning [Gat01, TS87]. Cleanup [Dwy79]. Clearly [Pet98c, Pet99, Pfa01]. ClearSight [Ano03d]. CLEO [Ano93f]. Client [Ano93-34, Ano98c, BW94a, BW94b, CS93, CS96b, Des96, Lac06, Mad94, Wid90, Bal96, EBS08, Ros94b, Wei97]. Client-Server [BW94a, BW94b, Des96, Mad94, CS93, CS96b]. Client/Server [Ano93-34, Bal96, Ros94b]. Clients [Jon98, Kos96, Ano95e, Len97, Smi92]. Clinic [San89]. CLIO [Par78]. Clist [Ger92]. Clock [Bab78, Stu78b, Vol75, GZ89]. CLOG [NB84]. Clone [Kni92, Tan87a, Tan87b]. Cloned [Yos85]. Cloning [Hun88a, Hun88b]. closer [Ano91c]. Cloud [MKSB22, BBB^{+}20, DC18]. clouded [Bal96]. CLtL2 [NSI^{+}92]. Cluster [Buy99, ER01, GVK^{+}03, GSHL03, Höf01, Hul05, Jen05, Kop00, LWL05, Rea11, RšT06, Ste01a, VBD00, VBD01, Wol04b, GK97, GLS03, Kop05, Ste02b, Vre03, Don02, EEK^{+}04, Kop05]. Clustered [Tho05, Luc05, MAT08]. Clustering [RMZ01, Boo03]. Clusters [SG09, WB96, Wey01, Boo03, LWL06, PS07, Sim97, Slo05, Spe00, YL09, Ano93-39, Nag05]. CLX [Lis02]. CM [Swi77]. CMS [Nor79]. CMU [Mac91, Mac92]. CNCC [ADJS78]. Co [Bar88a, Bar88b, Str80b, USE86b, USE88h, USE91a]. Co-Resident [Bar88a, Bar88b]. Co-Routine [Str80b]. COBOL [Gar97, Pay84a, Pay84b]. Cocoa [BY10, Chi10, GM02, Hil02, Hil08]. CoCreate [Ano03e]. Codagen [Ano03f]. Code [AQ84, BS79b, Bro19, Cou17, Cox12, Fra99, Gla78, Gra98a, Gra98b, Gra98f, Gra98d, Gra98c, Gra99b, Gra99c, Gra99d, Gra99a, Gra00a, Gra00b, Gra00c, Gra01a, Hub02, IEE03, Kul06, Lio96, LWL05, Ree82a, Ree82b, SU70, Tan87a, WCC^{+}12, WW92, AJ75, BY96, Che97, CGMS89, CP01, CP02, FRRJ07, FKL^{+}08, GHS82, Hoo05, Les88a, L^{+}11, Moo01b, Moo01a, PC01, SML06, SHLR80, Tan87b, YL09]. CodeGuide [Ano02c]. Codemesh [Ano01]. Coder [Ben88b]. Codes [Smi76, HL90, Rob8 , Rob79b]. CodeWarrior [Ano02c]. Coding [BBH18, C^{+}78a, Ver89, McS05, Sea05, Sea09, Sea14]. Coexistence [JLSG84, Raj98, WG98]. COFF [Gir88]. Cogito [BH86]. cognitive [van86]. Coherent [Kni92]. CoLiS [BJM^{+}22]. collaboration [Ano03e]. Collaborative [PK93a, PK93b]. collateral [PLM06, PLHM08]. Collect [Alb77a]. Collection [NCNS82, Ame09, Bli96, Hub96, JL96, Ngu08b]. collector [JHT^{+}07]. College [Ste90a, Osb92]. Color [Hof97, McC90, McD84a, McD84b, PH98, PH99, Wes78]. Colorado [IEE94b, USE86e, USE90b, USE91a]. Columbia [USE87e]. com [Bar00c, Sun01]. Combating [Cla00]. Combo [Var96b]. Comer [Mad94]. Comes [Ano93j, Ano96d, Cus92, Kin93, McK94a]. Coming [Hos83, McK82, Wol04b, Ano93v]. COMM [Lon78]. Command [Bai99, Blu11, BK89, BK95, BBM84, Che79b, COV79, Csa77, DM79, Gai83, Ger92, HKF84, Hol87c, Lyo79, McD87, Nel79, Per78, Pru02, RW84, Rog79, Tho76, Uhl78, Blu08, McE05, Pfa00, Sho12, SVV92, Ste99a]. Command-Line [Pru02]. Command-Software [Rog79]. Commands [Abl93, HA84, Lud77a, NC08, Ols78, Pol93, Ric98, Wee99, X/O89b, Dit93, Gol95, Hor98, Lev96, MNC09, PM87, VR00, dVd93, ?, ?]. Comment [GG99]. Commentary [Lio96, Max01]. Comments [Luu90, Rit02]. Commercial [Kin83, SW82, UTC84, Gil96b, Ivi95, JJ91r, Osm96]. Commercialization [TY82a]. CommNet [Fie88]. commodity [Kop05]. Common [CBL96, GL75, GL76, Kem91, Mac91, Mac92, NSI^{+}92, Nel79, SZ95, BKM86, ER86, FAB^{+}07, LRB^{+}19, Sch95c, BGPS91, Hen92, MC75a, Ric90, Swe01b]. Communicating [Pik93, dVd93]. Communication [Ano77, Bas87, BMS83, Bra76b, CG95b, CK06b, De 78, DHS78, DD90, Had04, Joh96a, JF81a, JLSS77, Kau78, LC78a, Lin79c, NH99, PR85b, PR85c, PR90, RR96, RR03b, RšT06, Sat93, Sha77, Sun77, TG78, Wat87, Wid90, Wit79, ZST^{+}07, Zuc77b, Bas91, CG95a, Fra5 , Fra76, GK97, IBP03, Lin02a, MS95, Sau88]. Communications [ACH91, Ano93f, Ano96d, Cam87, Fra74a, Fra74b, Fre90, Gol98, Gra97, Gra98e, HR78a, Hof79, LC77, Lon78, MS05, PR85a, Req85, Rob84b, Ste99b, Tin78b, Ung78, Whi89, dC93, And95, ZG96, Mil84c]. communities [ACM04]. Community [Cle83a, Cle83b, Var91, DD05]. Comnd [Uhl78]. como [RÓ01]. comp.lang.awk [com97]. comp.sources.unix [Wei90]. Compact [Per84]. compaction [CDD^{+}05]. Companies [Bar00c, OT07, Poo91, Won92]. Companion [Com94b, Coh94, JH00, Aul96, CSAJ04, Kom96, ML02b, Rob00, SAS90, SAS93b, SAS93c]. Company [ASW^{+}01, C^{+}78b]. Company-Wide [C^{+}78b]. Compaq [Ano99d]. Comparative [AOAH14, Hei95, PM18, Raf79, SHB23, Smi02a, Yoo96, dVd93]. Compare [Mil86, Sta04]. compares [And17]. Comparing [Ral76c]. Comparison [And17, Ano19b, CQ83a, CS90a, CS90b, FG79, GBH86, GES98, HM75, HM76, Jal92, LB96a, Luu90, PS07, SK10a, TCM07, TF04, YSC^{+}06]. Compatibility [GLOR91, Ame90a, Ano91b, Gue87, Inc83]. Compatible [Ano02b, Erm79, FH88, HMP83a, HMP83b, Hol87a, Hol87b, KM86b, Mul79b, SGP^{+}23, Swa77, Zuc83b, Zuc83c, AAB^{+}05, DS88, Guy00, MM06b]. Competition [Kat83]. Compilation [EST86, FST91]. Compile [FH96]. compiled [Inc83]. Compiler [ACM82, AU77, FW78, Fel79b, FW90, Gil88, HKS^{+}04, Hol12, Hua06, Joh78a, Joh78c, Joh79b, Joh79c, McI60, McI72, MK84a, Ols73, PKP05, Rem73, Rit8x, SF85, Sny75, Ste01c, Thoxx, WJW^{+}75, Cro82, Gra87, Joh75, Pow84a, Joh83, Lio79e, Mar78d, Row79a]. Compiler-Writing [McI72]. Compilers [ASU86, ALSU07, Ano02a, Cro78b, FST91, Gra79a, TvK83, Gou04, Pay84a, Pay84b]. Compiling [FST91, Hua06, Set81, Set82, Set83, HC89]. complaints [HBH18]. Complete [AS90, Bee96a, Bee96b, Cof90, HS99, JF12, Jos19, Kom00, Leh99, Leh03, McK98, Nor99b, O^{+}98, Ass99a, Pet98b, Pet00b, QH01, SGFR90, SGFR92, Sch91a, Sch91b, Sha00a, She93b, SU95, Wel98, ACS98, Bru02, Car89a, Chi02, Fei01, Gay99, Hub96, LD13, Luc08, McS05, Pet01b, Pet02a, Pet02b, RHK^{+}07, Sch00, Sho12, SZ95, Smi03a, War02, WS96]. complex [Gri13, Kau93b]. Complexity [Smi02a, MR67]. Component [HHLS90, Kro99, Lis02]. components [Ano03d]. Composing [Hol90a]. comprehensive [Kut02, Lev02, MSW06, Ros06]. compressed [NT05]. Compressing [Tho92]. compression [BW94c, CKV93]. Compromise [BBP99, Gar97]. COMPSEC [Ano93g]. Computation [BF9 , BF80, Bro0 , Bro81a, Bro81b, The92c, The92a, The92b, The93, MN69, Bro6 , Bro77b, BH8 , BH79, Fel4 , Fel75, PF05, Tun05]. Computational [Gar99, Wün13].

Computations [Fra79, WB96]. Compute [Ano19e, CDV^{+}94]. Computer [Alb77a, Ano93g, BED^{+}85, BRD77, BHM84, CEC^{+}96, Cla79, Coc01c, CH83a, CH83b, CBV93, DHS78, Fel78, Gor81, HM91, HB04, uRH95, uRH96, HCG^{+}82, IEE83, IEE86, IEE88, IEE91b, Isa86, KI99, KL79, Kil90, LTW77, LK76a, LK76b, LK77, Lew21, Lin65, LR78, Lyc71, Lyc73, Lyc74b, SL76, MS05, Min82a, Min82b, MC74, MC75b, Mul79c, NIS92, O’B82, Rem73, RG91, Sch81, Sha77, Sto89, Sto90, Str79, Tan88, Tho85a, Tho85b, USE86d, USE87b, USE89h, Usa78, WC98, Wei78b, WC93, AC11, And01, Ang97, Ang99, Ano00g, BSK87, FRAC85, Fel8 b, Fel79c, FL80, Fra5 , Fra76, FGP^{+}91, Gre96, KD94, May84, Nat88, Shu06, SHVP05, WLHC84, Ano01, CPJ^{+}98, MKS^{+}89, Poz83, Shi77, TBGE89]. Computer-Aided [Tho85a, Tho85b]. computer/communication [Fra5 , Fra76]. Computeralgebra [uRH95, uRH96]. Computerized [GJLB77, Mil77]. Computers [Ano82, BS90, Che76b, FT83, Gra79c, Kop00, Lin84, O’D90, SSWW83, Vol75, FHR84, Mor78a, RW87, SB08]. Computing [ACM75, ACM04, ASW^{+}01, Ano93-34, Bar00b, BDD^{+}93, BH1 , Buy99, Cer98, Cer03, DJ93, Dow93, GS01, Hah94a, HB82, Hul05, IEE94a, ILG19, KL94, Lin93, Lun99, May00, Mil16, RRN95, RR95b, RRH94, Ste01a, TSO^{+}96, Tom02, TWM86a, TWM86b, Ass88a, Ass88b, Ass89a, Ass89b, USE90a, USE93a, USE93e, USE94a, Var96b, WCM93, Bro3 , Bro74, Che94, Che90a, GS02, GLS03, HTU96, Ste02b, Wal01b, Mel95]. Compuware [Ano03f]. concentrate [SZ95]. Concentrix [Tes86]. Concept [FT83]. Conception [Dup05]. Concepts [Ber91, Dun91a, FG79, LS90b, SW99, SGG05, SGG09, Kra05, SCJ96, Sul02]. concise [FJ81]. Concludes [Coc01a, Coc03c]. Concurrency [BS90, RR96, RR03b, Bin85, Had04].

Concurrent [Cme86, HMP83b, HK89, Pik89, Ger85, Gue19, Hol82, Rob06b, Ano01, Cme86, HMP83a]. Concurrently [Har80]. Conditions [Cou84, Cou85a, Ral75a]. conducted [Ric77]. Conference [ACM75, ACM79, ACM92, ACM93, ACM04, Ano89, Ano93x, Ano93t, Ano93-31, Ano06a, BSK87, DFG99, D^{+}76, Eur91, IEE83, IEE91a, IEE91b, ITC93, IEE94b, IEE94a, IEE02, LT80, MS05, Mil84b, NIS92, Sof83, Sof84, Ste90a, Uni96b, Uni97, Uni98a, Uni98b, Usr82, USE82, USE83, Ass83, USE84a, USE84b, USE85c, USE85a, USE85b, USE86e, USE86a, USE86b, USE87a, USE87d, USE87e, USE88c, Ass88d, USE88g, USE88h, USE89e, USE89f, USE89b, USE89a, USE90b, USE90c, USE90d, USE91a, USE91b, USE91c, USE91d, USE91f, USE92a, USE92b, USE92d, USE93b, USE93c, USE93f, USE94d, USE94e, USE94f, USE95b, USE96a, USE97b, USE97c, USE98c, USE98d, USE00b, USE00a, USE01b, USE01a, Ano86, Ano94m, Ano95h, KL92, MN69, Uni96a, USE85d, USE86c, USE96b, USE98b, USE98e, Den99]. Conferencing [STT86a, STT86b, TP91]. Confessions [Ben88b]. Confidentiality [ZG96]. configurable [XWRZ19]. configuraciones [VD01]. Configuration [Ale00, Bur76, Chr98, EDK00, Eva97, Hes94, Led94, Loi77, RWNA87, Siv03, SUM95, VRJ97, Won98, Kof00, MBC^{+}86, Mag00, VRJ95, VRJ96]. Configurations [Vog79a, VD01]. Configure [BM92, Wil11]. Configuring [Pac02, Kob99]. conformance [Ame90b, IEE91c]. Confusing [Tip92]. confusion [Che94]. Connect [Ano00e, BGJ01]. Connected [Höf01, Ano97b]. Connecting [Cas93, Coc01a, Mil84a, Sch88, Gia90]. Connection [Cha85, Fel94, Liu96, Ano96f, BMR82, BMR01, KNI88]. Connectivity [Ano93-34, Hun92d, Ano93f]. connector [BCM88]. CONS [Hus92]. consider [Sim96b]. Considerations [Chr90, EGL86, Hel92, Lud79b, Pat83, Sun77, SW91b, Ver89]. Considered [LT91, Pik83b]. Considering [Tay96]. Consistency [RKSV20, HBH18, HWBH19]. Console [Cla79, LTW77, Win02b]. consolidation [R^{+}02]. Constant [Hob89]. Constants [GG99]. Constitutional [Mei84]. Constrained [Per87]. Constraints [GSHL03, HT92, HT94a, HT94b]. construct [Mar82]. Constructed [BEHW86a, BEHW86b]. Constructing [ALxx, Ryd8 , Ryd79, UTC84, SVV92]. Construction [ACM82, Hol12, SF85]. Consulting [CAG89a, CAG89b]. Consumption [LPG^{+}18]. Container [SHB23]. Containers [Gal05]. contenders [Ale96b]. Content [Rit79c, Zdz05]. Contention [Wil82a, Wil82b]. Contents [Lud78a]. Contest [Coc01a, Coc03c, Dav96]. Context [Ral80, Str90, BLH20]. Contiguous [Lyc74a, Zuc83a]. Continuous [Vol75]. Contour [Pav0 , Pav81]. Contract [KM79b]. Contractors [MH78]. contracts [Tuc90]. contribution [Hus92]. Control [ALO90, AHPS79, Bai99, BMF15, BGMR91, BS79b, CE97, Che76a, Dil02, Dou02, FA88, Gla78, Gop79, GGZ02, Hel92, Hoa77, HP89c, Ima76, Kat99, Kra88a, Kra78b, Lon77, Rei97, Rog79, Set81, Set82, Set83, Vog76a, Wei78a, Wei88, Wil82a, Wil82b, WSS77, WSS78, Yos85, Ano93g, LHM^{+}83, Luc13b, Mas05, McN00, PCSF08, TBG^{+}87, TH04, Wei97]. Controlled [BRD77, CW15a, CW15b, LTW77, Sla94]. Controller [GZ84a, GZ84b, Lon78, RS78, ZST^{+}07]. Controls [Har92a, Har01a, Har01b, Str90]. Controversy [Coc01a]. conundrum [Sim96c]. Convention [IEE91a, ITC93, NIS92]. Conventions [Wid90]. conversation [PB21]. Conversion [Alb76, Ell79a, Hai78b, Hoa77, Sny78]. conversions [Hor98]. Converter [FGMS90, FGMS95, Smi78]. Converting [Csa77, Kap79, NP78, SL95, SB96, MM06b]. Cookbook [Bor14, BM92, Hag83, Kal16, Kim95, Loz05b, Mar10, Stu01, Wil11, YK15, AVN07, BSB03, Bou97, Cev03, G^{+}06a, Ger92, HZ01, HZ02, Hun03, L^{+}11, P^{+}08, Sch05, Sch07, Sch08, Stu04]. cool [NW03, Neg06b]. Cooperating [BEHW86a, BEHW86b]. Coordination [Hol80]. Copley [USE01b]. Copy [KLK^{+}02, Lyo79, GMTW08, SM88a, SM88b, SM88c]. copy-on-write [SM88a, SM88b, SM88c]. copyright [ZB01]. Copyrights [Gar97]. Core [GT78a, ISO05, MM01a, Max01, Roo79a, SGP^{+}23, XZY^{+}10, ZL10, ZB19, Kas14, PM10, SG97a, SG97b, Sul02, Sob06]. Core-Resident [GT78a]. Corel [Dan01, Smi99]. corners [Blu09]. corporate [Mit01]. Corporation [Ano98c, Ano00e, Ano98c, Ano00e]. Correct [JP17]. Correction [CW89, CW91]. correntropy [HMK^{+}20]. Cosmology [Coc01c]. COSNIX [Kap92, Rog79]. Cost [Liu96, Rob95, Sta05a, Wai86, War84, ZB19, Ano03e, GK97, Tay96, WT91]. Could [Gar97, Wel99b, Sem96]. Count [Nis01b]. Counters [DRN^{+}23]. Counting [Mor78b]. Countries [Ksh04]. Country [TBGE89, USE01a]. Coupled [BFS89, AGP^{+}92]. coupling [FAB^{+}07]. Courier [Lar83]. Course [Kom00, Kom02, PS77, QH01, Ros77a, Wel98, Wen79a, BW02, Har93, IEE91b, Mat07a, Mat07b, Orw96, SAS94b]. Courses [BMV78, AEG^{+}91, Bai87, Mat93, Tan87b]. Courte [Sch94a]. COVER [Sim96a]. Covering [ALxx, HHH11, HHHT10]. Covert [WB92, CMR^{+}21]. CPU [Bac92, Ano96e, Ano00d, Ano00f, CS90a, CS90b, Fed79, FKN85, Hel92, Mar78a]. cq [Gra79a]. Crackberries [Sta05a]. Crackdown [Ste92a]. Cracks [Ano93u]. Craft [Fin91, EM79, Fil79]. Crafting [JP09, BK13]. Cram [Pri99, Sav11]. Crannies [Gra97, Gra98e]. Crash [AM86, Bur79a, DB95, Dun92, McK82, Ral76c, Wei78a, BW02, Bin88, Mat07a, Mat07b, SWF^{+}22]. Crashes [Ral76a]. CRAY [WW92, AO86, HK86, LR90, Mil93, Rit88b, Rit88a]. Crayons [CCVLW23]. Create [Arn98, Kuo98a, Gri13, Gri14]. Created [Lew21]. Creating [Böl01, End00, ADH^{+}89]. Creation [Coc01c, CKDS90, Cur91, Tin78d, Too10, CK06b, GGZ02]. Creative [Jam09, Ham07]. creators [Ano19c]. credit [Osb92]. Critic [Lew99a, Lew99b]. Critical [BBH18, TLR89]. Critique [BMM85]. Cross [AS88, AS91, Ale96b, Mul79b, Sim96b, HBH18, JJ91q]. cross-platform [HBH18]. Crowder [Ano00b, Ano00b]. crown [Geh03]. Crucial [Gar97]. Crude [Ker81a]. crunchers [DJM86]. Crushing [VB92, VB94]. crypt [Wei78b, RW84]. Cryptographic [CCDP01, Out82]. Cryptography [KWDB06]. crystal [Fel4 , Fel75]. CS [Bai87]. CSA [Reh05]. CSACCS [AHPS79]. Cscope [Ste85]. CSE [Kee05]. csh [DuB95]. CSL [Per82a]. CSRG [McK98]. CSWAP [Kau93b]. CTRAN [Che90a]. CTSS [AO86]. Cuckoo [Sto89, Sto90]. CUDA [SK10b]. CUG329 [dFAAW90]. CUG417 [Vol94]. CUG418 [Vol94]. Cult [Col87]. Cumulative [Rei78]. CUPS [GGZ02, GGZ02, Swe01b]. Curley [Ano00b]. Current [Gor81, MKS^{+}89]. Curses [Hol87b, Sch90, Hol87a, Str86a]. Curves [Hob89]. Custom [KM95, Sou94]. customer [Ano97b, SAS91a, AHPS79]. Customization [WCM93]. customizations [Kra07]. customizing [Neg06c, Neg07b]. Cut [Lud77a]. CVS [Fog99, TH04, Ves06, ZK05]. Cwpr [IBC87]. CX [Cal08, Cal09, SG06]. CX-310-200 [Cal08, Cal09, SG06]. CX-310-202 [SG06]. Cyberpunk [HM91]. Cyberspace [Gar97]. Cygwin32 [Noe98]. Cypress [CN88].

D [Mah89, Nag05, Sau88, Ale10, PH87, Wal95b]. DA11B [Lau77]. Daemon [DMT02, Fed88a, Fed88b, Kat99, Lau79a, Len87, Sal08]. Daemons [DiM00]. DAG [GNV88, GNV90]. Dallas [USE85c, USE85b, USE88g, USE91d, USE91f]. Dalziel [Ano16a]. Damage [MTY01]. Damaged [Wan77]. Damn [SAN08]. Danber [HNR84]. Dance [Nas97, Wol04a]. Daniel [Ano00b]. Danish [Kri91]. DAPRE [SP90]. DAPRE/UNIX [SP90]. dark [Blu09]. Darth [Ree82b, Ree82a]. Data [Alb77a, All87, ASS79, BS79a, Bai85, BI95, Ben88a, Bon79a, Bon79b, BW94a, BW94b, Cha69, DFG99, FY18, Far90, Feu78, Fra74a, Fra74b, Gel78, GVK^{+}03, Gra99d, GR76, Har80, HIT^{+}02, Hoo83, Jan08, Jon84, JN88, Kay79, Kra78b, Lev79, Lin79b, LT80, SL76, McS78, Mor78c, PR85a, Par78, PNK02, Per82a, Per84, PH86, PW79, Rit79c, Rob84b, Roc77, Roc78, SL95, SR77, Sne01, SW12, Str84c, Stu78a, Tar89, Tho92, Tin78c, Tin79a, TT79, Van88, War78, Wol93, X/O88a, BW94c, CKV93, DD05, HU73, Kaz05, Koh99, MRM06, SML06, SGM^{+}00, SAL93, Str80a, Str82, Coo85, Lin02a, McL83]. data-analytic [DD05]. Data-base [Tin79a]. Data-Compressing [Tho92]. Data-Intensive [HIT^{+}02]. Database [Ano92g, Ano02b, Bar01b, Bar86, Bur79a, BF97, CS86, Duf82, Eng91, Haw85, HL85c, HK89, JPS00, KKF^{+}84, MK84b, Pru02, Roc79, Roc82, Rod90a, Rod90b, Sti83, Sto80, Sto88, Vol97, Wei82, Bal96, Bro97, BGW04, HLHR90, PWN01, Ros94b, WCHS93]. Databases [Pac02, Wei82]. dataflow [HKVR21]. DATAKIT [Ell79b, Coh90]. dataset [DFC^{+}23]. DATASPEED [Dow77, Whi75b]. date [Jan06b]. Dateien [DF00]. Datenverschlüsselung [Lin02a]. dating [KMP^{+}20]. d’autore [ZB01]. David [Ano95j, Ano16b, Mad94]. Day [Kor97a, HMK^{+}20]. Days [MB00, CCB00, LH00, Til95]. Dayton [IEE91a]. DB2 [IBM09a]. DBA [Emm11]. DBAs [Bur01]. DBMS [Wil94, Kin83, TT79, Tin79b, TBS87]. DC [Ano85, MC78, USE84b, USE87e, Ass88c, USE90d, USE99b, Dit93]. DCE [Lei00, Ros92, RT93, Shi92]. DCE/RPC [Lei00]. DCL [Ger92]. DDA [Bai85]. DDB [SAL93]. DDD [ZL96]. Dead [DeV93, Ral76a]. Deadlock [Pea80]. deadly [Tox07]. Dealing [GSHL03, Loz05a]. deb [VD01]. Debian [Ano00c, CK06a, CK06b, DF00, Her04, Ano05, BJM^{+}22, Bra04, CV00, CK06a, CK06b, Den99, Dum05, Epp00a, GA04, Gan04, HHV05, Kle00, Kra05, McC99a, Neg06a, Neg07a, Ron05b, Ron05a, SS05, VD01, WN15]. Debian-GNU-Linux-Powerpack [Gan04]. Debug [Leh95, Reg11]. Debugger [Car83c, Car85, Har92b, Kat79, MZG14, SPS^{+}02, Zim85, Car83b]. Debuggers [ZL96]. Debugging [Ano98d, CP86, Gla79, Koe79a, PR96a, SPS^{+}02, You95, Ano03e, Bes05]. Debunker [KB01]. DEC [Ano90a, BJ77, HL95, HS83, KM79b, Kri84a, Kri84b, LR78, Sch77, Weh77a, Weh77b]. December [IEE04, USE00b, USE01a]. Decentralised [LHM^{+}83]. Decentralized [Wei78a]. Decide [Sim96c]. Decision [Bal94, GW98]. Decisions [Bel18, HK82b, Sim96c]. Declarations [Set79c]. Declarative [HLHR90]. declared [Kau93b]. Decline [You92b]. DECNET [JLSG84]. Decoder [Ale00, MYH00]. decoding [HL90]. Decomposing [JKS^{+}15]. Decrypting [Wei78b]. DeCSS [Coc03c]. DECstation [Edd90]. DECwindows [She93c]. Dedicated [GSHL03]. Deep [HLT^{+}21, vdL94]. defect [Bla86]. Deferred [CW89, CW91, Swa74]. Definition [Ano96a, Bai85, IBM88]. Definitions [Cra83, X/O89a]. Definitive [BS01, NZW^{+}02, vHW03, vHW06, BP09, Den04, Emm11, Kir09]. Defunct [DS00]. Degunking [Bal05, SD05a, SD05b]. Delegating [DM88]. Delivering [Den83, Hen92]. delivers [Sch95c]. Delphi [Jac99]. Demand [Gar97, Jun85a, Jun85b, Sha81]. denied [Chi02]. Dennis [Aho12, Ano11, Bro20, Ker12, McI12, Pat12, Sch12]. Dense [Kau7 , Kau79a]. Denver [USE86e, USE86b, USE88h]. Department [Maj77, Hal79, Lud77b, Lud79a]. Departmental [Smi76, Ano95e]. Dependencies [Flo90]. Dependency [LA90]. Deploying [BH00, BH01, End00, Hal05, MKSB22, Wid00, HSG03]. Deployment [BBD^{+}00, DeV00, MHS02]. Depot [MWCR90, MWCR91]. Depth [Doe11]. derivatives [SAS88b, SAS90, SAS91a, SAS91b]. Descendants [Lud78b]. Description [Bas87, Fol77, HR77, Gur81, Mar75, MP85, Now78, Weh77b, Win75, FJ81]. Design [AU77, Ano02c, BJP04, Bac86a, Bac86b, Bee90, Bel18, CGS96, CCDP01, DCN02, Des96, EEK^{+}04, EM00, Fas78, Fla94, FR77, Fra78, GD18, GMB93, GC94a, GC94b, HKF84, HR79a, H^{+}79, Hol80, Hol91, Hoo78, HL85c, Joh79a, KL73, KM82, Kob91, LKY93, LMKQ89, LMKQ90, Lut99, Mac83, MTY01, MRxxb, MK88, MBKQ96, MNN05, MNNW15, Mer82, Mik92, MH96, MBB^{+}91, Nee90, NN92, OH05, Pal96, PV00, PK84, RO92, SGP^{+}23, Sne01, SWKH76, Sun96, Sun77, SW91b, TW97, TLK99, Tho85a, Tho85b, Ver89, Wan02, WJW^{+}75, YLK93, Ano03c, Ano03d, BW96, Bha12b, Bor88, BY10, Col84b, Com12, Gab07, Get91, Gia99, LM91, Lov10, PDG^{+}87, Pal87, PDG^{+}88, Reg11, Sch93b, SS95a, SAL93, SHPP07, TW06, Tob80, WXX08, CS91b]. designed [Len97, Mud97]. Designers [Hae92, Ham07]. Designing [ER01, HD02, ITC93, Jou86, Kon07, MS95, Smi92, JJ91o]. Desk [Ari92, CM78a, MC78, Dys96, Haw00, Haw02, Kee05, Reh05]. DeskSet [Sun93c]. Desktop [AH04, Bar91a, Böl01, Bra04, BM92, Cos97, CBL96, Dec04, FJPR93, Gt92b, Hol92, KGG00, Kni92, Lin93, NS92, OD^{+}86, Rob99b, Rob99c, RB92a, RB92b, SZ95, Val92b, Var96b, Ano00a, Bri05a, Dea01, Hal05, Hek97, PB05, PS^{+}09, Sch95c, SH99b, S^{+}00, S^{+}05, Sun93b, SZ95]. Desktop- [Bra04]. Desqview [Rei93]. Desqview/X [Rei93]. detect [DLCV17]. detecting [CMR^{+}21, HMK^{+}20]. Detection [BLH20, BF03, MC74, MC75b, Nor99a, NZW^{+}02, Reh03, BBD23, DFC^{+}23, Ras07]. detector [CGMS89]. Detects [SEZ^{+}01]. determination [WB06]. Deterministic [AJU75]. dev [Kam02]. develop [Reg11]. developed [Moh85, Var96b]. Developer [TS97, ZM08, Chi10, Fei02, HZ01, HZ02, Jac99, Lei95b, Lei96a, Lei96b, Moh98, Sun93a, Tut93, Ano02b, LZHS11]. Developers [Ano97a, Ano06b, FT09, Sha00a, Ano93r, Bam98, Osm96, Tho06b]. Developing [BGR89, BR99, Coc03b, DH96, FN83, Kel90, Ksh04, Li98, Env97a, Hoo05, PBM^{+}22, Sea14, Ano93h]. Development [ACM79, ABB^{+}86a, ABB^{+}86b, Ano98c, Ano98d, Ano01, Ano03f, BBH18, BCGS91, Cam04, CB83, CPJ^{+}98, Coo09a, Coo09b, DH02, Dav97, Den79, Gop79, Gwo78, Hai79, HL95, Ima79, Jen75, JL78, JL87a, JT98, JL87b, Kro99, ML09, McI82, Mic79a, Mik92, NM83a, NM83b, NR03, Per87, RC83, RP02, Rog79, SNM22, SW82, SC78, Ste01f, Swe01a, SS93, Too17, Uni96b, USE94c, Wal94b, WT01, WL97, ABG^{+}86, And06, Ano95j, BV06, CM06, Fog99, Ger03, G^{+}06b, Gra87, HGS99, Joh80, JT05, LW03, Liu11, Lov05, Lov10, Pay84a, Pay84b, PB09, Raf79, Rit93, Roo79b, Swi91, Tev87, Thi99, Uni96a, VMKB05, WWCW19, War85a, Yag91, ZL10, Kro99, Lac06]. Device [Alb84, Ano02c, Coo09c, Coo09d, Don97c, ET92, IBC87, Mic79a, Paj88, Paj92, Rub98, RC01, Ung78, Wat83, Wei93, CRKH05, ET88, JJ91a, JJ92d, JJ92a, JJ92b, KPG93, Kha02, Kon12, MM06b, PLM06, PLHM08, TF04, Ven08, Fur85]. Device-Independent [IBC87]. Devices [Kam02, SFE10, SW12, Tin78a, Ano03e, DLCV17, DC17, LZHS11]. DevOps [CHL16]. DFORMAT [Ben88a]. DH [Whi75b]. DH-11 [Whi75b]. Diagnostic [Cla79, Dul94, Pek78]. diagnostics [ITC93]. Diagrams [BJK92, BJK87]. Dial [Gil93b, LP89, NL79]. Dial-Up [NL79, Gil93b, LP89]. Dialing [Fol77]. Dialogue [Lew99a, Hel03]. DIBOL [ASS85]. DIBOLIX [ASS85]. Dickens [FES^{+}99]. DICTION [CV80]. Dictionary [Jae91, Jae94a, Jae94b, Ray91, RS93, WC93, Lev96]. didactical [AEG^{+}91]. Didn’t [Wel99b]. Diego [USE83, Ass83, USE89f, USE89b, USE91b, USE93f, USE94d, USE96b, USE97b, USE00a, USE01a]. dienste [WF03]. differ [Siv90]. Difference [Ano93-42, HKT05, Mar78e, TBGE89]. Differences [Rib89, Hor90b]. Different [FES^{+}99, HR79a, MH96, And17, Mud97]. Differential [BG97, HM75, HM76, KS85, Kau90, Sch75, Sch76a, Sch76b]. differentiation [CL06]. Diffie [ASW^{+}01]. Digest [USE97a]. Digital [AC11, BBK13, FFPC22, Gar97, Hol90a, Kom02, Pal96, RC83, Vol75, JP09, Ano99d, Ano91a, Bun91, CGS96, Che96, Che99, HBMT96, PR96a, Var96a]. digitale [ZB01]. DigiTerra [Ano02b]. dilemma [Ano93r]. dimensional [MMTK94]. Direct [LWL05, Ric85, Koh99, LWL06, Cat79, CL79, HR79b, Lin79b, Lin79c, Lin79a, Roo79a]. Directed [GNV90, Set81, Set82, Set83, GNV88, IEE03]. Directions [MK87, Tag83]. directive [YL09]. directive-based [YL09]. Directories [Sin97, Smi76, Pon97a]. Directory [ASS79, FA89, FA94, GJLB77, Gol95, HSG03, Ano93-38]. diritto [ZB01]. DIS [ISO05]. Disappear [Pug90b]. Disaster [Bar00c, IBM09a]. Discipline [Hub79, PDL04, TS87]. Discover [Oua97]. Discovering [Bro20, AK^{+}07]. discovery [Lyo08]. Discrete [ACM97, Cou84, Cou85a, RDK75, GL75, GL76]. Discusses [Gar97]. Disk [How01, Koc87, KM83, LEG88, Mad97, RW93, Stu78b, TBGE89]. Diskless [ACF^{+}86, Lob85]. Disks [McK06, McK12a, McK12b]. Dispatcher [CL79]. dispersion [DD05]. Display [Ozi78, Rei77, Wes78, CGMS89, Ado93]. Display-Oriented [Rei77]. Displayed [SL76]. disruptive [ALP^{+}21]. Dissertation [Bro20]. distance [WN15]. Distributed [Ama91, And88a, Asb00, Bar86, Bar88a, Bar88b, BCC^{+}91, Bri01, BP84, CW92, Che87b, CG95b, Cme86, CJ92, CP86, CBV93, CKDS90, Cur91, Eur91, Gal00, GZ84a, GZ84b, HLW84, HHLS90, JH86, KDP83, Lee89, LS90b, LC90, Lin93, LSC^{+}88, MRK90, MMS86, MF99, MvT^{+}90, PK93a, PK93b, PPTTxxb, Sat89, Sat90b, Smi89, STT86a, STT86b, Tal89a, Tal89b, TSO^{+}96, TT79, Tom02, TWM86a, TWM86b, USE89c, USE91e, VB92, Wam83a, Wam83b, Wei98, Ano95i, BGW93, Bro94, CG95a, Lit88, LCH^{+}81, PT93]. Distributing [BFS89, Gra01b, RT93, AGHR89a, AGHR89b]. Distribution [Ano24, Koe84, LC77, LC78a, MSSvK08, Ols73, BFS07, BP09, DL08, DLP10, Fur85, PF96, THB06]. Distributionen [Bra98]. Distributions [HBD97, LZAH23, Neg06a, Neg07a]. District [USE87e]. DITROFF [BBG85]. DITROFF/FFORTID [BBG85]. Diverse [Bra93, JH86]. Divide [Ale96b, Sim96b]. Division [GS12, Wen79a]. DivX [ZB01]. DM40. [Wei95]. DM40.- [Wei95]. DMCA [Coc03c]. DMERT [Sch79a]. DNS [Luc03b, AL92b, Hun00a]. Do [Jen02, Pog17, Sim97, Wel99b, Dwy79, Hoo83, JN88, O’D83a, SF95]. Docs [Hub96]. Doctor [Jae86]. Document [BBM85, Geh86, Geh87a, GL88, KLO78, Ker84b, Ker85, Ker89, NCNS82, Ope93, Ope94, SD95, Sri93, Kan93, KFO97, KL82, YLM^{+}05, ?, ?]. Documentation [AT7, Bea93, CPJ^{+}98, Cum77, Leh03, Mar01b, Mas77a, SM79, Swi79, Vog74, Vog76a, YT83]. Documenting [PLHM08, SH85, Sin97]. Documents [Com94d, Com94g, CL90b, DHP77, Ker79g, Les77a, Les78, MK85b, MK90, SP77, Vog77b, Hub96, Les76b]. DODES [Sch75]. Does [Ano93-43, OS06, Ree82a, Ree82b, SF95, Wol04a]. doesn’t [Ano95d, Cra93]. Dog [Coc01b, WBS01]. dogs [Wis03]. doing [Ros94b]. Doings [Rit02]. Domain [BSS^{+}95, BSSW96, OBSB23, Ste96, Lei00]. Domain-Specific [OBSB23]. Domains [JKS^{+}15, Lei98]. Domino [RMZ01]. Done [Gra79b, Gra06b]. DongTing [DFC^{+}23]. données [Dup05]. Doomed [Gre82a, Gre82b]. Doors [Ano91a, Bun91]. Dorothea [Hid04]. DOS-Anwender [TBR93, Wei95]. DOS-BATCH [Alb76]. DOS-Programme [Str94]. DOS/OS/2 [Ger92]. DOS/UNIX [Hun92d]. Dot [Bar00c, Sun01]. Dot-com [Bar00c, Sun01]. Dots [Coc01a]. Double [Sch75, RR95a]. double-edged [RR95a]. DoubleVision [Ano00e]. Douglas [Mad94]. Down [JJ91p, Ang97, Ang99, RFS05, RFS06]. Download [Leh95]. Downs [Chu03]. Downsize [Ano92g]. Downsizing [Gli92]. DPD [Wag79]. DQUAD [Blu75]. Dr. [Ano94a]. Draft [Car83b, ISO05, Rit71, Ban88]. DragonFly [Ree06a]. drawbacks [Mud97]. Drawing [GNV90, Hobxxa, Bah09]. draws [GNV88]. dream [Lum13]. Driftshåndbok [Skr89]. drills [Ano19c]. Drive [Law98]. Driven [BFKS93, Bla83a, Bla83b, FRRJ07, GHS82, Ney83a, Ney83b, ODH^{+}85, RS14, Sch93c, SVV92, Wil11]. Driver [Bab78, ET92, GT79, Lau77, Mic79a, Mue78a, Paj88, Swi77, Ung78, Whi75b, ET88, TF04, Wei93]. Drivers [Alb84, Coo09c, Coo09d, Don97c, Paj92, Rub98, RC01, Ste01e, Wat83, Wel94a, CRKH05, Fur85, JJ92d, JJ92a, JJ92b, Kha02, Kon12, MM06b, PLM06, PLHM08, Ven08]. driving [Bri05b]. Dropping [VB92, VB94]. Drucksysteme [GGZ02]. Drupal [JP09]. DRX. [Wt95a]. DSM [PBS98]. DSP [Row95]. DST [Sta07a]. DTM [PS07]. DTrace [MMG07, MGMMI10]. Dublin [Mil84b, USE87a]. Dumb [McC90]. Dummies [HS99, HS00, LY93, LY98, LY04, Mea99, Mil16, Dul14, Hug98, LY97, MWM00, SP15]. Dump [DB95, Pla89, PK88]. d’une [Dup05]. DUNIX [Lit88]. DUP [Lon78]. Duplex [Lon77]. duplicating [Hun88a, Hun88b]. DV [Mue78a]. DV-11 [Mue78a]. DVD [KMV^{+}09, L^{+}11]. Dynamic [AU3 , AU74, GM11, HBD97, IBM09b, Oom13, RKPP88, Sha93, SEKV12, Jal92, JL96, MGMMI10, Mly09, PBM^{+}22]. Dynamically [MZG14, MM96, Don04]. Dynamics [PF84a, PF84b]. Dynamite [FJPR93]. DYNIX [Bad98]. DZ11 [Sha79b].

E-mail [Bur93, Cam91, Hil93]. EADAS [Stu78a]. EADAS/NM [Stu78a]. Earliest [Bro19]. Early [ASW^{+}01, Bel18, FT83, GL91, HH97b, PPT^{+}xx, Too09, Ano93m]. Easier [Ano82, Mor78a, Far91b]. Easy [BM92, Vog77a, Den04, Gra05, Mus96, Mus02]. easy-to-use [Den04]. Easysoft [Ano00e]. eclectic [HA95a, HA95b, HA96]. Eclipse [Coo85, ZK05]. Economics [Ksh04]. ed [Sch04, Ano85, Hor79]. eDesktop [Cal00a]. Edge [Sta05a, Tor99]. edged [RR95a]. Edgewatch [San98]. Edison [Thi92]. Editieren [Str94]. Editing [BBM85, Fil79, Fin91, Ker79a, eL85, SBA92, Tar80, Goe07, Son87, Com79]. Edition [Ano00b, Dep79a, DHP77, GT92c, HH86, Hul05, Kil90, KGG00, Koe78b, NO90b, NO92, OD^{+}86, PS79, PR85c, PR90, QO93b, Rag90, SRBB99, TGB95, TRS18, Too10, WIP^{+}97, You90b, YP92, Ano96b, Ano00a, BN01, Gt92b, Gue87, Kuo98b, Pog05, Pog07, PKP05, QO91, TG96, Teo13, Wei93, Bea98, Duf90a, Lio96, PR85b]. Editor [DM97, Far79, Feu79b, Fra80, Han86, Hor79, Ker79c, Lam90, McM78a, McM90, Pik87, Pik90, Rei77, RT70, Rob99a, Smi06b, Ste01c, War78, Ano00g, Hol87d, Kir09, eL85, McM79, Ube76, Uni85, McI69a, Mei84, Rob99d]. Editors [Far92b, Hew90, Ozi78, Ste01e, BB99]. EDTV [Rei77]. Education [C^{+}78b, Dia94]. EEG [CJBT92]. Effective [BLH20, BMF15, HM98, HN01, LK73a, LK73b, Rob01, Sea20, Lev02]. Effectively [LP95, May79]. Effects [KM83, Mar78a, SM88a, SM88b, SM88c, Ham07]. effektiven [Kuo99]. efficiency [BRW89]. Efficient [AC75, ALxx, Alf94, Ber98, HL90, KL70, Hen92, NSI^{+}92]. Effort [GS12]. Effortlessly [Far92c]. EFL [Fel8 a, Fel79d, Fel79a]. Egan [Wei93]. Egg [Sto89, Sto90]. EGP [Kir84]. eidaburyukei [Sas91c]. eigenvalue [Hal4 , Hal75]. Eighteen [MFBWW20]. Eighth [ACM97, USE99b, HH86, PR85b, USE94d]. Eilert [Hid04]. Einführung [CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, PKP05, dG96]. Einsatz [DLvG07, Wei04, SU94a, SE96]. Einsteiger [Ric14, Ron05a]. Einsteigerseminar [RGRG05]. Einstieg [TBR93, Wei95]. Eisenhaendler [Hid04]. Electrical [SK72]. Electronic [Ano00e, FA89, FA94, Gia90, Höf01, O’D83b, STT86a, STT86b, BK13, HH04, van86]. Electronics [IEE91a]. ELEFUNT [Cod89]. Elegant [Oua92]. Elektronik [Ric14]. Elektronik-Projekte [Ric14]. element [LWL06]. elements [Oua92]. Eleventh [USE97b]. Eliminating [MG99]. Ellen [Sch04]. Ellipses [McI92, McIxx]. ELM [Wei92]. Elmer [RW81, Tuo82]. Elmer’s [Car82]. Eloquent [Hav11]. ELXSI [Car86b]. Emacs [Bor88, CR91, Lew88, Mit84, Str94, Com84, CCA84, Gos81, Gra87, PDG^{+}87, PDG^{+}88, Sta93, Uni85, Cam98, Fin91, Pal87, RB92a, RB92b, SBA92]. Email [Cla00, Dil02, LP95, Liu96, SEZ^{+}01]. Embedded [Abb03, Ano93x, Ano01, BBH18, Cev03, DH02, Gra01a, Hal07, Hol02, Jon84, Jor04, Jos95, Kro99, Lom01, Nis01a, Nis01b, Nis01c, Nis03, Row95, TTB09, Weh01, Woe02, Yag03, Ano95i, Ano95j, Ano03d, FRRJ07, Kha02, Mol14, Reg11, SJP^{+}09, Ano02c, Coc03c, Coc03b, DH02, Fis02, Hal06, Jor04, Nis01c, Tar89, Sta07b]. Embedded-SQL [Tar89]. Embedding [ATS94, Gra99b, NMP82]. EMC [Koh99]. Emerging [BJ93, Gla92]. emission [Kaz05]. emphasis [IEE03]. Empirical [MSSvK08, MCM06, MFS90, MCM07]. Empirix [Ano03f]. Employs [Ano00d, Ano00f]. Emts [Ral80]. Emulation [CWS90, LC73, LC75, PBS98, PS82, SSWW83]. Emulator [Cap82, Cap87, MZG14]. emWare [Ano02c]. Encapsulating [Sch93c]. Encoding [CPJ^{+}98, Mud97, Lin02a]. Encryption [Gar97, LT91, Mor79a, PS22]. Encyclopedia [Mv94]. End [Bur79b, Den83, Gay93, Law98, ZL96, Gil96a]. End-User [Den83]. Ending [Zdz05]. Energy [FRRJ07, RF93, ZB19]. Energy-optimizing [FRRJ07]. Enforcement [BSS^{+}95, BSSW96, RKSV20]. Enforcing [Oom13, Pov99]. Engine [Yue01, TBGE89]. Engineer [Pri99, Jan04]. Engineering [Bel18, BM93, DP95, D^{+}76, Ela83, HKM^{+}85, IEE94b, MKB89, McK91, McK92a, McK92b, MK96, McK96b, McK97b, McK97a, Mil84c, Stu78a, Swi91, FKL^{+}08, SML06, SHVP05]. engineers [ABS95, Jam97]. England [USE88c]. English [DD05, RFJ97, RFJ99, McI73b, McI74, MC75a, VRFJ97a, WAC84]. Enhanced [Ano00e, BW88b, BW88c, Che96, MMC06, McC05, Tru83]. Enhancement [Sha79a, JF81b]. Enhancements [ADJS78, Can79, MKL85b, McK03, RW86a, Don22, MKL85a, WBSK87]. Enhances [Ano03f, Ano03d]. Enhancing [AW89, TR94]. Enlarging [GT78b]. enough [And02, Hoo83]. ensemble [BBD23]. Ensuring [RS14, ZG96]. Entering [SL76]. Enterprise [BH00, BH01, CR02, CHL16, ER01, G^{+}06a, Jen05, KJHL96, KJG^{+}98, Kop05, LBSL97, McC04, Mel95, Mil16, NBB^{+}02, Rei03, Sne01, Sob06, Ale96b, Blo99, EBS08, Gil96a, KJH^{+}97, Sim96c, Spe00, Var96b, WWCW19, Ano96e, Asb00, SKB00]. enterprise-level [Sim96c]. entertainment [NW03, Neg06b]. entropy [SNM22]. Entry [Cle83a, Cle83b, Roc77, Sab78, Lyo76b, Sab76a]. Entwicklung [Lac06]. Entwurf [Dav97]. Enumeration [Ni97]. Environment [ACM79, Ado89, Ano94f, Ano01, Ano02c, Ano03f, BS92, BL75, Bou86, Boy84, BF9 , BF80, BCGS91, Bus02, CGS96, CSA^{+}87, CMM88, Col83, Col84a, CD85, CJ92, CBL96, Don98b, Gau98, HL95, Hau94, Hel97, HCN85, Hof97, Hor92, HN01, Hub02, Jac77b, JTUB85, JH86, Kar92, KDP83, KMKO92, KLP88, KM78, KM79a, KM81, KP84, Koe78b, Koh06, Kos87, Kro99, KI82, Lam83, Len94, LC90, LSC^{+}88, Lyc72, Mer82, MK93, MZA96, NR94, NBB^{+}02, Pat83, Pek78, PK84, RC83, Ric90, Rov78, Roy93, Sch79b, Sef93, Siv01, Ste92b, SR04, SR05, Ste92c, SS93, TLK99, Tho85a, Tho85b, TSO^{+}96, Tin79b, Too17, VBD00, VBD01, Wal97b, Wal98, WL97, WCM93, AKB85, AKCB86, Bal96, Bam98, Bis89, Bla98, Bla86, GGJ89, Gra87, Gre81, JAV86, KIT93, KK89a, eL85]. environment [Lub90, MK92a, MK92b, MB05, Par93, Raf79, SAS90, SAS91a, SAS91b, San91, San94, Sch79c, Sch95c, SB08, SR08, SR13, SW91b, Tar89, War85a, Win02a, ZGdDL^{+}96, SZ95, Col92a, Nel90, Pug88b, Teo13]. EnvironmentalStats [Mil98, Mil00]. Environments [Bre86, CW15a, CW15b, Har93, IEE86, IEE88, Kie02, Rib89, SAS94b, Ste02a, dSVdLT02, CMR^{+}21, Che90a, DJM86, Nat88, Ni97, Ros94b, SAS93a, SAS93b, SAS93c, SAS94a, SAS95, WCHS93]. epilepsy [CJBT92]. Epilog [CJBT92]. Epoch [Coc01a]. EQN [Scr76, Har11]. Equation [Blu7 a, Blu78a, BG97, Sch75]. Equations [KS85, Kau90, Kau93d, Sch76a, Sch76b]. Equities [LSC^{+}88]. Equivalent [Coc01c]. ergonomic [van86]. Eric [Jon88]. ERP [Bar00c]. Error [Bra77, Bra78, Dit93, HR79a, Nad98, Per77, RW86a]. Errors [MC74, MC75b, MC75a]. ERRPT [Per78]. erstellen [GGZ02]. escalation [XWRZ19]. escape [Blu09]. eServer [R^{+}02]. Espionage [Sto89, Sto90]. ESS [Pek78]. Essay [Ano99a, Wei78b]. essays [War08]. Essential [ABS95, Fri91, Fri03, Gil93b, Smi06a, Smi06c, Ven08, Ves06, ZK05, Gre96, Hor98, Kom96, Pet00c, Wal01b]. Essentials [Bra92, JDB09, LCJH05, Rea11, Smi12, Vea10, VSC^{+}10, Wat10]. Establishing [Fer85]. Esterel [MS90]. Estimate [Stu78b]. Estimates [HR79b]. Estimation [Stu78a]. Estudio [RÓ01]. ESX [D^{+}04, OH05, R^{+}02]. ETA10 [Car89a]. etc [AC11, Hoo83, Kod82a, Kod82b]. Ethernet [Spu92, Cab83, Gor81, Sku88]. Euclid [Ger85, Hol82, HMP83a, HMP83b]. Euclidean [Blu7 b, Blu78b]. Europal [Ste90a]. Europe [Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93-35, Ano93-36, Ano93-37, Jud90, Uni97]. European [Ste90a, Gar97, HLHR90, Hus92, Ass89c, WB85a, WB85b]. EUUG [Ano89, USE87a, USE88c]. EUVL [Bar01b]. evaluating [EP05]. Evaluation [BSW78, CW89, CW91, Cha79, Cla79, DF79, HCF^{+}94, PM18, RLB08, Cha96, EEK^{+}04, GLG93, PDG^{+}87, Pal87, PDG^{+}88, Ros78, SK10a]. evasion [Blu09]. even [Len97, Wil11]. Event [RF93, Sch93c]. event-driven [Sch93c]. Events [Mor78b].

Ever [RRH94, Wal95a]. Every [RP84, Lov07, War04, War15]. Everyday [MM01b, Stu01, Gra05, Stu04]. Everyone [Coc01b, Gwy90]. Everything [Tan87c, Wal95a, Woe03, LD13]. Evidence [GS12]. Evolution [CMR^{+}90, Dun91a, Fed83, Fed84, Gla78, LADADL14, MQ09, MR17, PV00, Pir88, Rit9 , Rit80, Rit84a, Rit84b, SA21, Weh83, IF10, Kor06, MQ10, PLM06, Spi17, War08]. evolutions [PLHM08]. evoluzione [ZB01]. evolve [SS95a]. Evolving [Bal94, LL93b, MS87, MS89]. eVote [Ano98c]. ex [Hew90, eL85]. Exam [Cal09, Ern11, Pri99, San07, DRP01, Jan04, Sav11, Ern11]. Examination [Mag04, Pir88, Ste85, MKL^{+}95]. Examining [NL00]. Example [GG91, Gay00b, Qui95, Qui99, Qui02, Rob04, SK10b, Gre96, JSG00, MMC06, OP01, Qui97, Qui00, Qui05, Wal00, Wei97, Wil11]. example-driven [Wil11]. Examples [Csa77, LS90b, PKP05, QSP85, Bli96]. Exams [SG06, San07, Tom11]. Excellence [Ano93e]. Exception [BJ01, Eyk88]. Excerpts [McI87]. Exchange [Sch85a, Sch85b, Str79]. exciting [Gri13]. Exec [Loi79]. Executable [Coc03c]. Executables [SEZ^{+}01]. Execution [BBB^{+}20, JZTB98, WB96, AHA^{+}14, FH89]. Executive [LS87]. EXecutor [KE85, Sak97]. Exercises [TG89, Wan78, Zim98, Coo09a, Coo09b, Coo09c, Coo09d]. existing [LBSL97]. Expand [Skj82]. Expandable [NLR84a, NLR84b]. expanded [Ano85]. Expansion [Weix , HC89]. Expect [Lib94, Rze93]. Experience [BMR94, BFS07, BCC^{+}91, CR92, CKT80, Duf89a, Duf89b, FSS99, Fre91, HJAW88, JLSG84, MH78, Pin92, Rit88a, SK04, Tuo83, dSVdLT02, GL91, HBMT96, WWCW19]. Experienced [Jae93, RRF90]. Experiences [AAB^{+}05, ADH^{+}89, Bec84, CB83, FT83, Lio79b, Lio79c, May84, NV06, Rit88b, TP91, TBS87, USE91e, BHK^{+}84]. Experiencing [AEG^{+}91]. experiment [GHS82, Han77]. Experimental [Fra74a, Fra74b, HK83, Hof79, NHR84, SD87, PBWB84]. experimentation [GGJ89]. Experiments [IvW87, Les88b, WSS77, WSS78, WN15]. Expert [BH86, Kur90, Pet09, vdL94, FF88, Sav11]. Explained [GC94a, GC94b, Kruxxa, Kruxxb, Pet98c, Pet99, Pfa01]. Explanation [Ral75a]. Explicit [HKT05]. exploitation [PM10, Ron05b]. Exploiting [HBD97, AK^{+}07]. Exploration [Gil93b]. Exploratory [ASS79, Lud79b, SA21]. Explorer [Dum05]. Exploring [AH04, Ano93-33, Lib94, ML09, Mol14, Tho19b]. Explosion [FH92]. Expo [Ano06a]. exponential [Wal96]. exposed [HLK01]. exposition [She02]. Expression [Cox07, Cox09, Cox10, Cox12, Sab76b, AJ75, Tho68]. Expressions [Fri02, KP99, Pik06, SU70, Bro3 , Bro74, Fri97, Fri06]. ext3 [AR07]. Extend [Ano03a, Mud97]. Extended [Bee90, CFA85, Eyk86, Kol77, Str90, LGJS91, Wei97]. Extending [BP88b, BP88c, BP88a, BE00, CS96a, DJ93, FA88, Ger02, KPG93]. Extends [Ano03f]. Extensible [McN00, PRC^{+}23, GRA89a]. Extension [BMS83, Kar92, Sau88, Zuc83b, Zuc83c, Ame94, SG97a]. Extensions [AISS98, BJP04, Bec06, Cla88, CCF89, LH84a, LH84b, McI60, Sun77, Zuc77b, Bec07]. Extensive [Fie88, Bli96]. Extent [MK91]. Extent-like [MK91]. External [The92a, The92b, WLC95]. extra [Ano03a]. Extraction [Bon79b]. eyeing [Tuc90]. Eyewitness [TBGE89].

F [Fla90, DF00]. F4P [HY79]. F4P-UNIX [HY79]. Face [Edw98, PP85]. Facilities [DH77, GBH86, Jen75, Lyc74b, RW86b, Rie93, Rie94, Tin78a, CMR^{+}90, Jal92, ?, ?]. Facility [Bai85, BM87, BK84, Che79a, Flo90, Gra01a, Lic79a, Lic79b, Lio78, MF83, Mar78f, Nor79, Rit88c, Sha79a, Tin78b, TG78, TG79, Vog78b, Vog79c, CLJC95, SAS91a, Str80a, Str82, KSU94]. fact [Lon90]. factored [Bro3 , Bro74]. Factors [Fra83, Gin83, Mac83, SM89, WBSK87]. Factory [BJ77, Kro99]. Fading [Cas94]. Fail [Coc03c]. Failover [Roe02]. failure [Jen19]. fairness [WTKW08]. Faithfully [Sma88]. Fall [USE90a, You92b]. Fame [Ano19c]. Family [IBM88]. far [Var96a]. Fargo [Ano97b]. Farm [Bar00c, For00]. Fast [BS92, Cox07, FJPR93, FST91, HS90, HS91b, MDAD^{+}14, McC90, MJLF84, MG99, McK03, McK07, MQ09, McK16, McS79, Mor79a, New92, O’D90, SRTH15, SB97, Tun05, Wal96, Wes78, Ano96c, MQ10, Pax88, dBB08, BGJ01]. Fast-Access [McS79]. Fast-forward [MQ09, MQ10]. Faster [BE89, Fed78a, FH96, PK88, Ano96e, BE88]. FAStT [D^{+}04]. Fat [For00, Mar10]. Fat-free [Mar10]. Fate [JC92, Bal96]. Fault [Ano93i, BBG^{+}89, BW94a, BW94b, FES^{+}99, MRK90, KIT93, SS00]. Faults [PTS^{+}11, PTS^{+}14, KIT93]. Faust [Ham90]. Favorite [Ker81c]. FAXes [Pin92]. FDSOI [ZB19]. Fe [USE93d]. Fear [MAT85b]. Fears [Gar97]. Feasibility [Fel79e]. Feature [Loi79, PQM^{+}21, GL75, GL76]. Features [Ano00d, GMB93, HJAW88, MH88, Out82, Ano95g, Bha12b, Sim97, Wil11]. featuring [Ban88, BBD91]. February [USE89f, USE02]. Fed [IEE91b]. Federal [BSK87, Tuc90]. Fedora [HH05, McC04, Neg06a, Neg07a, Sob06, Tyl06]. feedback [IEE03]. feedback-directed [IEE03]. Feel [Car90, Var96b]. Fever [Gil96b]. few [Sim97]. FFORTID [BBG85]. Ffsck [MDAD^{+}14]. FGMC [HMK^{+}20]. FGMC-HADS [HMK^{+}20]. Fi [Goo90]. Fiberoptic [Coc01b]. fiction [Lon90]. FidoNet [Cro98a]. Field [AA86, ACS79, Gin83, Ska94, WWCW19, Zie93]. Fields [Roc78]. Fifth [Ass88c, USE89h, Ano93x, USE91b, USE95c, Far95]. Figgins [Sch04]. fighting [L^{+}11]. File [AISS98, Ama88, Ano96a, AM86, BG88, BBS95, BP88b, BP88c, BP88a, BEHW86a, BEHW86b, Bor98, BS94, Bry83a, Bry83b, CCDP01, Che79b, CE97, CFA85, dC87, DM07, Daw07, DKM96, Ell75, FG91, FE77, GGL03, GT79, Hai78b, Hai78c, Ham78, Har92b, HLW84, Hei95, HP89a, HP89b, HH86, HHLS90, Hum88a, Hum90a, HM75, HM76, IK83, Jon84, JN88, Kle86, Koc87, Kow79, Kow90, LEG88, LA90, LS90b, Lio79a, Lit80, LADADL14, Lud78b, Lyc78a, MDAD^{+}14, MLRC88, Mar75, MS92, McI63, MJLF84, McK06, McK07, McK12b, McS79, MK93, Mor79b, Mv94, Nad92, Nor79, Pat83, Pos88, Pug92, Rag90, Rea91, RW84, Rit79c, Rob75, Roc82, Rog79, RO92, Roy93, Rul96, SRTH15, Sat90b, Sax85b, Sax85c, SBMS93, SGM^{+}00, SA95, THB06, TECB03, USE92c, Vog79b, Wan77, Wan78, War78]. File [Weh83, Weix , Wei84, WF01, Wil82a, Wil82b, ZDS85, Ano00g, AR07, Bar01a, Bin88, GMSP00, Gat01, Gia99, GGZ02, Han01, Hu86, JHT^{+}07, Jan03, KAS^{+}06, Len97, LGJS91, Mad97, McK12a, MK91, Mel87, MK92a, MK92b, ODH^{+}85, dJPDD84, Wei97, Yos85, HK79]. File-System [MDAD^{+}14]. Filenames [SB96]. Files [Bis88, Bre86, Csa77, Ell79a, Fle02, HK79, Kil84, Koe79b, Lev94, Loz05a, Lyc74a, Per82b, Pug90a, Pug90c, SSB92, TTG^{+}87]. Fileserver [Kos96]. Filesystem [FM94, HAB^{+}98, IHR99, KM86b, LPLF91, MKB90a, MG99, McK03, Mel02, JJ91t, McK16]. Filesystemless [Rub79]. Filesystems [Len97, PV00, vH02, McK96a]. Filing [Mil77, MS84]. filling [Pav0 , Pav81]. Film [Gar97]. filosofia [Cor05b]. Filter [CJ94, Ree06a, SEZ^{+}01, SEZ^{+}01]. filtering [Zdz05]. Final [Sha79a, JJ92c, Ung86]. Finally [Ano93j, Gil91, Var96b]. Find [Blu7 b, Blu78b, Mui95, Ste01d]. finder [SWF^{+}22]. finding [Sim96b, Ano93i]. Fine [FAB^{+}07, LC93, KIT93]. Fine-grain [FAB^{+}07]. Finite [HKT05, LWL06]. Firefox [FFPC22]. Firewall [Don96b, Roe02, BP09, Han08, Han11]. Firewalls [Ano00b, CB94, CBR03, NZW^{+}02, SH95, WC98, Zie00, Art03, Ras07, SY00, ZC02, ZCC00]. firms [Fra90a]. firmware [Fra19, HK82b]. First [IEE94b, Mil84b, Ste90a, TLR89, USE88e, Usa78, Ano03d, KL92, SHVP05, Ano02c, Too10]. Fisherman [IEE03]. Five [OH93a, Ric93, Sal03, Skj82, IBP03]. Fix [Ava92, Wei88]. Fixing [All97, Dzo09]. Fjords [Coc01c]. FL [USE89c]. Flash [Ano02c, Ano03a]. Flash-Based [Ano02c]. Flaw [Gar97]. flex [Pax88]. Flexibility [Car82]. Flexible [CS86, Fed77, HL85c, Bor88]. Flight [PJ89a, PJ89b]. Floating [BF9 , BF80, Bro0 , Bro81a, Bro81b, Sch81, TOML04, Bro6 , Bro77b, LMOT01, LMOT02]. Floating-Point [BF9 , BF80, Bro0 , Bro81a, Bro81b, Sch81, TOML04, Bro6 , Bro77b, LMOT01, LMOT02]. FLOSS [GS12]. Flow [Set81, Set82, Set83, HU73, LHM^{+}83, RKSV20]. Flowgraphs [Bak76, Bak77]. Flows [ALxx]. Fluid [LWL05, LWL06]. FM [Rog79]. Focal [Ano96d]. Focus [Ano99a, Ano99c, Ano99b, Cas94]. Fokker [MMTK94]. fold [Ano93r]. Folding [Lut99]. folks [SF95]. Follow [MM96]. Follow-up [MM96]. fonts [Ano03a]. Footnote [NB84]. footprint [CDD^{+}07]. Force [Ano03f, FDMS79, KI99]. Forecasting [Ano98d]. Forefront [Ano00d]. forensic [DC18, FFPC22, KMP^{+}20, KMV^{+}09]. Forensics [BBK13, AC11]. Foreword [Aho84]. Fork [BS88, BAKR19, SM88a, SM88b, SM88c, dVd93]. Form [San93, Ano02c, Tun05]. Formal [Erm79, LL95, RAMP94]. Format [CL90a, Les76a, Les77c, Les79b, PT93].

Formatierung [GGZ02]. Formats

[Ben88a, BS94, Mv94, Rul96]. Formatter [KV89, Wen79b, McI73a]. Formatting [BBG85, Der90, Geh86, Geh87a, GGZ02, Smi76, Smi77, Vog77b, GL88, Com79]. Formed [Gar97, Pov99]. Former [Wei78b]. Forms [All87, Joh87, SL76]. Fort [USE89c]. ForTec [Mei84]. FORTED [Ube76]. Fortgeschritten [Ron05a]. Fortran [Mei84, Che90a, FHS7x, Hal69, HR73, IEE92b, Lou90b, McI71, Pay84a, Pay84b, Ryd72, Ryd3 , Uni82, WW92, Ano02a, Bak75, Ber85b, Blu7 b, Blu78b, BG97, Cod89, FW78, Fel79b, FGMS90, FW90, FGMS95, GG99, Ker4 , Ker75, Ker79b, Ker93, Mei84, Raf79, Ube76, Win75]. Fortran-to-C [FGMS90, FGMS95]. Forum [Lev79, Mor79c]. Forward [MPT78, MQ09, MQ10]. forwarding [SH13]. Fotos [DF00]. found [Var96b]. Foundation [ABB^{+}86a, ABB^{+}86b, IJ98, WL94, ABG^{+}86]. Foundations [EBS08, Van05]. Four [Fra90b]. Fourteenth [USE00b]. Fourth [Fow85, USE87b, USE90b]. Foxley [Jon88]. Foxtrot [Ame09]. FPGA [PBM^{+}22, SB08]. FPGA-based [SB08]. FPGAs [Ano02c, MKSB22]. FPPAC94 [MMTK94]. FR80 [Swa74]. Fragment [OH97]. Frame [Coo98, McC90]. Framemaker [Mat97]. Framemaker-5.5 [Mat97]. FRAMES [PH87]. Framework [BMF15, Cho98, FHS78a, Fox79, Lis02, MWCR90, MWCR91, RR93, GJJ05, KD94, Tho95, WXX08, XWRZ19, Sem96, Ano16a]. France [Gre92a]. Francisco [ACM92, IEE03, Ass88d, USE92d, USE94f, USE02]. Frank [Wei95]. Fraser [Ano95j]. Free [Lin02c, CPJ^{+}98, Jam09, KL94, Noe98, Pom04, Way00, ZL96, Cor05b, Cus04, Gra06b, HBH18, HWBH19, Kir09, Kop05, Lub90, Mar10, Sal08, ZB01, Lin02b, Gla97, NR03]. Free/Open [NR03]. FreeBSD [Ano00e, HS04, ZM08, And01, Ano96c, Ano98a, Cev03, CCB00, DM07, Daw07, DYZC20, GS12, GM11, Hon06, Hon08, JL16, Jud17, Kon12, KHP05, Lav07, Leh99, Leh03, Luc00, Luc02, Luc03b, Luc08, LJ15, Luc18a, Luc19, Mat07a, MGMMI10, MNN04, MNN05, McK13, MR14a, MNNW15, MR17, Mit01, NC08, PM18, Ree06a, Rob03, Sae03, Siv03, Smi03a, TU02, TU03, TU06, UT03, YSC^{+}06]. FreeBSD-4.x [Siv03]. FreeDOS [WF03]. FREENIX [USE01b, USE98b]. Freeware [Ano96c, Edw98]. Freeware-Pakete [Ano96c]. freie [Lin02b, Lin02c]. freien [Bra98]. French [Ron05b]. Frequently [Tim93]. Friend [Gra98c]. friendly [AGS90, CJ92, Thi92]. friends [DJM86]. Front [Bur79b, Gay93, ZL96]. Front-End [ZL96]. Frontier [Aud96, HM91]. FrontPage [Wid00]. FS [Rea91]. Fsck [McK02, BE88, BE89, McK13, Bin88, Kow79, Kow90]. FSDB [Har92b]. fsh [McD87]. FTP [Luc03b]. Fugue [DFV91]. full [Ano98a, Kat82a, Kat82b, Ano93j]. full-screen [Kat82a, Kat82b]. Full-Strength [Ano93j]. Fully [PS22, FPG89]. Fully-Homomorphic [PS22]. Fun [TD01]. Function [BM93, Sch91c, FF88, HC89]. Functional [ACM92, Car83a, Cou13, DFV91, KPSS79, McD87, MKL85b, Ng77, Tur87, BFS07, DL08, DLP10, DD05, MKL85a, Shu83]. Functionality [LBN14, Ano03a]. Functions [BF9 , BF80, Hol87a, Hol87b, Tho03]. Fundamentals [Rob04, Har93, ML02b, SAS94b, Swi91]. FUSE [HL95]. Future [ASW^{+}01, Ano19e, Goo90, Gor81, HNR84, Lew21, Mas87, Mas88, San98, Var96a, Var96b, Var91, WB85a, WB85b, Ano96e, Bha12a, Fra5 , Fra76, NIS92, Tev87, Tuc90]. Futurebus [PH91]. Fuzz [MKL^{+}95, MZH22]. Fuzzy [HMK^{+}20]. fwsnort [Ras07].

G [Ano02b, Gou04]. G. [Sau88]. G.1 [DH77]. G.3 [Rit7 ]. G.4 [D^{+}76]. GA [USE86a, USE91e]. Gains [Ano93n]. Galerkin [Sch76b]. Game [McS05]. Games [LH01b, Kos00]. gap [BYVF08]. Garbage [JL96, JHT^{+}07]. Garden [GC94a, GC94b]. Garfinkel [Ano96b]. Gast [WF03]. Gast-Systeme [WF03]. Gated [Fed88a, Fed88b]. Gateway [Bar86, Roe02, Kir84]. gator [Jen19]. Gaussian [HMK^{+}20]. gazillions [Var96a]. GCC [Gou04, vHW03, vHW06]. GCG [CH96]. GDB [MZG14, SPS^{+}02]. GDMO [Ano96a]. GED [Feu79b]. Geeks [Cov06, JR02, Gra04, Gra06b, Jan06a, JR04, JR05, JRR08, RJR08]. Gelato [Ano06a]. gems [Fer04, Ngu08a, PF05]. Gene [Ano96b]. General [CS86, Coo85, CH83a, CH83b, Fas78, Gla77a, H^{+}79, MH88, MK88, Ng77, SK10b, Tho79, PF05, Var96b, Wei97, McL83]. General-Purpose [Fas78, Gla77a, SK10b, PF05, Var96b]. Generalized [ALO90, Fra80, Roc77]. Generate [Bar01b, Swa77]. Generated [MM96]. Generating [AR78, Hob91, Hob93]. Generation [Bor79, Fow85, GLOR91, IEE03, MZG14, NB84, SU70, Stu78c, Thi04, AJ75, GHS82, Ros94b]. Generator [CH77, DS88, JS90, Les75b, LS79a, LS79b, LS90a, Mey18, Rob79a, Rob82, WW92, Don22, Les75a, Pax88, SHLR80, YL09, McI69b]. generators [Wal96]. Generic [Cum77, Gru86, SJ95a, SJ95b, Vog76a, Vog77b]. Genesis [Bha12b, Moh85]. gengo [AKWA89]. Genome [SMD^{+}94]. Gentoo [Thi04]. geographers [Biv87]. Geometric [Tho85a, Tho85b]. geometry [GS99]. George [ACM03]. Georgia [USE86c]. German [Lin02b, Lin02c, Bas87, Bra04, CK06a, CK06b, Dav97, GA04, GGZ02, uRH95, uRH96, Jor04, Lac06, LMKQ90, Lin02a, Mag04, May01, O’R90a, PKP05, RGRG05, Ron05a, Spr07, Sta89, TBR93]. Germany [BSK87, IEE91b, Uni97]. GESS [Kau93c]. gestalten [DF00]. gestión [VD01]. Get [Ano92m, Fro90, Gil96b, Gra98c, Hil93, Ste01c, Ano92i, Gra06b, O’D83a, Var96a, Ano94b, Nee91]. get-things-done [Gra06b]. GETLAB [Ano92h]. Gets [Far91b]. Getting [AS93b, Don97a, McI92, Ric13, Wt95b, WBS01, BA96, Wel94b]. Gex [Jac77a]. Gf [SR77]. GFS [MQ09, MQ10]. GHz [Ano03c, ZB19]. Gigabit [HIT^{+}02]. gigabytes [Var96a]. GIMP [DF00, Goe07, Ham07, JP09, LD13, Pec08, vGS10, DF00]. GIMP- [DF00]. GIS [Env97b, Li98]. Git [Cha09]. Give [BH86, Law98, Var96a]. Given [Den99, Ral77]. GLANCE [Lyc75]. Glasses [Don98a]. glimpse [GGZ02]. glimpse-Tools [GGZ02]. Global [AISS98, Uni96b, Uni97, Uni98a, Uni98b, Che97, FH89, Ger03, HU73, Uni96a, Ano98d]. Glossary [Anoxxa, MP93, Ros84]. GMAKE [Bee90]. GNOME [Ger03, Gri00, She01]. GNOME/GTK [Gri00]. GNU [Ano00c, Lin02b, Lin02c, Ano05, Bra04, CV00, DF00, GA04, Gan04, HHV05, Jor04, Kle00, Mag04, PKP05, Ron05a, Ygg94, Sal08, Ano93k, Lin02b, Lin02c, Cal10, CR91, Cam98, Cor05b, Cou17, DRP01, Don04, Dum05, FY18, GIM07, GA04, Gil88, Gou04, GS00b, Har11, Jon05, Jor04, Lew88, LO97, Mag00, Mag04, McC99a, Mec05, NR03, Lás05, PKP05, PR96b, RÓ01, Ron05b, Ron05a, SS05, SBA92, Sta93, SPS^{+}02, Sul02, Swa00b, VD01, VETT00, W^{+}95, ZB01]. GNU-Linux [Kle00]. GNU-Linux-Zertifizierung [Mag04]. GNU-Softwaretools [Jor04]. GNU/Linux [Ano00c, Lin02b, Lin02c, Ano05, CV00, GA04, HHV05, Ron05a, Lás05, RÓ01, Lin02b, Cor05b, Don04, Dum05, GA04, Jon05, Mag04, McC99a, NR03, Ron05b, Ron05a, SS05, VD01, Lin02c, Cor05b, RÓ01, ZB01]. Gnuplot [Jan08]. Go [Jos97, Red97, DK16, Gue19]. GO-Joe [Red97]. goal [SZ95]. godfather [PB21]. Goes [Bar00c, Bar01b]. Going [Bar01b, Chr90, Ste00, Uni96b, Uni97, Uni98a, Uni98b, Ale96b, O’D83a, Uni96a]. Golang [Gue19]. golden [Mee95]. Good [Cla00, Rob79a, Rob82, Ste01d, Var96b]. Google [Cox12, GGL03, Sta06]. Googlebombs [Sta07a]. Googling [Sta06]. got [And17]. GOTHIC [CMM88]. Gould [Ble83, FT83]. gourmet [HP01]. Governance [MR17]. Government [Bar01b, Coc03a, Jac91]. Governmental [CPJ^{+}98]. GPL [Guy00]. gprof [McK84]. GPU [Fer04, Ngu08a, PF05, SK10b]. GR [Bec79, Has78]. GR-Z [Bec79, Has78]. Grafik [PKP05, Str94]. Grafiken [DF00]. GRAFPAC [Swa74]. GRAFPAC/FR80 [Swa74]. grain [FAB^{+}07]. grained [LC93]. Grammars [AJU75, van83]. Grand [Cus92]. grap [GL88, BK86b, BK90a, BK91a]. Graph [DS77, DBG89, Ryd8 , Ryd79, Sto76]. Graphic [Ano94c, Kos87, Scr76]. Graphical [Bec79, Feu79b, HL95, HM90a, SR77, ZL96, AGS90]. graphically [KD94]. graphically-oriented [KD94]. Graphics [And88a, BS94, Dan83a, Dan83b, Duf90a, Has78, HB04, Jac77a, Jac77b, JR93, Ker81a, Ker82a, Ker84a, Ker90b, Ker91, Kro99, Lev94, McD84a, McD84b, MB03, Mv94, PJ89a, PJ89b, Pik82, Pik83a, PLR85, PLR00, PKP05, Rul96, Sch77, SC78, SA93a, SA93b, Sto76, TG88, USE86d, USE87b, USE89h, VK79, Van80, Vol94, Ang97, Ang99, CGMS89, Dia94, Fer04, FL80, GL06, Gla92, Ker81b, Kir09, Pav0 , Pav81, PF05, Pik84a, SML06, SVV92, Tho06b, Wal95b, Bar00b, Feu79a, SZ95]. Graphs [BK86b, BK90a, BK91a, GNV90, Hobxxa, Jan08, KL70, Stu78c, Vog77a, GNV88]. great [Gil96a]. Greed [For00]. Green [CPJ^{+}98, Ano01]. GREMLIN [Höf01]. Greps [Hum88b]. Grid [Wol04b]. Grids [McI83]. Group [ACH91, Ano93-32, Ber92, Hag83, MKS^{+}89, Sof83, Sof84, Usr82, Ass89c, WP90, WMP92, And95, Eis79b]. Groups [BMF15, Law98]. Groupware [Kin93]. Growing [Bur93, Jac91]. Grown [Gar97]. Grows [Ano93-39]. Growth [IJ98]. Grundlagen [DF00, GA04, SU94a, Str94, SE96]. gsck [Sax85b, Sax85c]. GSIP [Sto76]. GTK [DF00, Gri00]. GTK-Versionen [DF00]. Guaranteed [RT83b]. Guardians [DM88]. Guest [Rit88c, BB99]. GUI [Dal99, Dal02, Fla94, Gar99]. Guide [AT7, AA86, Ano80, Ano06b, Arn78, AS90, Bar91a, BS01, BS91, BMS95, Bia99, BBM85, BD94, BS79b, Bra93, BDD^{+}06, CR02, Cam87, C^{+}03, Cov06, Cur92, Dat88, DG00, Doc03, Ern11, Far78, Gag01, GM02, Gil93b, Gli90, Gre02, GBG97, Gwy90, HO 78, Hah93, Han86, Har11, Har95, HY79, HS83, Hew93, Hew90, Hol92, How01, Hun92b, Jos19, KM95, Kel90, KL94, KC76, KC78b, Kir95a, Kir95b, Kob99, Kro92, Lat02, Led94, Les77a, LB96b, Luc02, Lyo76b, Man93b, Mar78g, Mar10, The92c, The92a, The92b, Mat93, May97, McC02, MP92, NS92, NS94a, NDW93, NZW^{+}02, NH91, Nor93b, NS94b, The02, OQL88, The91, Ope91b, OH05, Olc91, PDG02, Pal96, Per79, Pew93, QO90, QO91, QO93a, QO93b, RRN95, RR99, RR03a, RR95b, RR96, RB08]. Guide [RB92a, RB92b, Rod87, Roo79a, Ros98, Ros90, Ros93b, Rus91, Ryd3 , SAS91a, SPS90, SPS93b, SPS95, Sab76a, Sab77, SG06, Sax85b, Sax85c, Sch75, Sha00a, Shi92, SR95, Sil84, Sob89, Sob91, Sob95, Sob97, Sob99, Sob03, Sou94, SP90, Str79, Sun95, TV04, TY82b, TNY92a, TNY94a, TNY92b, TNY94b, Top90b, Tra96, Tri90, TS97, Val92b, Wat02, Wes81, Win01, X/O88c, X/O88a, X/O88d, X/O88b, X/O89b, X/O89c, X/O89a, You90a, Zie93, Zlo91, vHW03, vHW06, Ame90b, Ame90a, APK99, Ano85, Ano96f, ACS98, Bah09, Bar04, BGJ01, BDP05, Ber85b, BM02, Bri05a, BP09, Cal10, Car89a, Cer96, Che99, Che02, Chi02, Coo09a, Coo09b, Coo09c, Coo09d, DRP01, Das01, DeR97, Den04, Eck11, Fei02, GGM93, Goo07, Gra05, Hah09, HP01, Ham07, Han08, Han11, Har93, HGS99]. guide [HG95a, Jac99, Jan04, Jos97, Kee05, KD00, Kir09, Kom96, Kon12, Kut02, LK00, LD13, Lew91, Lov10, Luc08, Lyo08, MM00, Man91, Max02, Mit01, Mon95, ML02b, Nor99b, OK94, PS94, Par00, PM10, Pet01a, QO89, Raj98, Reh05, Ric95, Ros06, SAS88a, SAS91b, SAS94b, Sha00b, Sha01, SHPP07, Sob06, SS06, Son87, SS00, Sun93a, Sun93b, Sun93d, TR99, Tom11, Top02a, Top05, Tut93, Wal01b, War02, Wil11, Win93a, Win93b, Win99, Win00, WNDO97, W^{+}99, WS96, YE84, YEB84, ?, ?, HCF^{+}94]. Guidebook [Dun91b, Sil84, Gra04, Gra06b]. Guidelines [H^{+}79, Kap79, Ols73]. Guild [Coc01a]. GUIs [Wal97a]. Guix [Cou13, CW15a, CW15b, Cou17]. gureppu [MO01]. guru [Mon95]. Gurus [Mic08]. gut [BW04, Gün07, Rob99d, Rob02b].

H [Ano95d, McI69b]. habitual [Mil01].

Hack [Chi02, Sei00]. Hacker

[CB94, CBR03, MD09, RR03a, Ray91, RS93, Sou94, Ste92a, TR99]. Hackers [HM91]. Hacking [HLK01, Kra07, Kon07]. Hacks [Lav04, ORC06, Fli03, Kra07, PB05, Ran04, Ran06, VJ06]. HADS [HMK^{+}20]. Hahn [Hah09]. Half [Lon77, Par90]. half-baked [Par90]. Half-Duplex [Lon77]. Hall [Ano19c, Hid04, Mad94, Sta07b, Tho05]. Hallinan [Sta07b]. Hanasemasu [JK86]. Handbook [Alb77b, Ano00b, Bra04, DKF94, EDK00, Hab01, Har95, HR91, HH91, KC98, Lew01, MD09, NSS89, NSSH95, NSS^{+}01, NSH02, Nor99a, Pon98, Pon99a, Pon00, Pon01, Ron05a, Smi00a, War05, AK^{+}07, Bam98, BDH00, Chi10, DuC94, Fla00, GWS94, Han01, Har01c, Har00, Jos99, Ker10, KC00, NSH06, NSHW10, Pon02a, Pon03, Cuc90, Ray08, Tho19a]. Handbuch [WF03, Bra04]. Handle [Uhl78, Len97]. Handler [Fed78a]. Handling [BJ01, Bur79b, Dul96, Eyk88, Fed78b, GVK^{+}03, HR79a, Lic79b, Rem97, RW86a, RLB08]. Hands [Ano03g, Gue19, KmWH10, Koh06, Ros99, Tom11, Gra04, MB05, Shu06, SHPP07]. Hands-On [Ros99, Gue19, KmWH10, Koh06, Tom11, Gra04, MB05, Shu06, SHPP07]. Handwriting [Kem93]. HAO [Daw86]. Happened [WCC^{+}12]. Happn [KMP^{+}20]. Happning [KMP^{+}20]. Haptic [Ano03g]. Hardening [Tur05]. Hardware [BW94a, BW94b, Bro79, Car82, Edd90, Gra98b, Hol02, Led94, Loi77, MH78, PLR85, PLR00, PKP05, Smi00a, YK15, Ano96e, Kop05, RHP^{+}07, SB08, VMKB05]. Hardware/I [Car82]. Hardware/I-O [Car82]. Hardware/Software [PLR85, PLR00, SB08]. hardwarenahe [PKP05]. Harley [Hah09]. Harmful [Pik83b]. Harris [Lef81, Sha81]. Harris/6 [Lef81, Sha81]. Hashing [Lit80, SY91, WG00, Weix ]. Hater [Ray08]. HATERS [Tho19a, GWS94]. Haviland [Gra89b]. having [Web95]. HDL [Pal96]. Head [MT74]. Headed [Ano94n]. Headers [X/O89c]. Health [Ham87, LZHS11]. Heart [Coc03a]. Heidelberg [Wei95]. held [BSK87, IEE91a, Mil84b, MN69, Tob80]. Hellman [ASW^{+}01]. Hello [PT93, Ano19c]. Help [Bro79, Dav93a, Gra98c, BK84, PDG^{+}87, Pal87, PDG^{+}88]. HEMS [Par88]. Henderson [Daw82]. Henry [Ano16a]. HEP [Wie89]. Here [Ano93-40, Ano96d, Sim96b, Spe96]. Heritage [RP84]. Heterogeneous [And91, CG95b, CBV93, Höf01, PMD88, TSO^{+}96, VTS^{+}23, BBD23, CG95a, PBM^{+}22]. Heuristic [KL70, KL73, LK73a, LK73b]. Hewlett [Cle83a, Cle83b]. Hey [Gra88]. Hi [Kay79]. Hideous [PW85]. Hiding [Flo90, Rez95]. Hierarchic [Tra85].

Hierarchical [HLT^{+}21, IK83, Kay79, SG09]. Hierarchies [Bry83a, Bry83b]. Hierarchy [HP89a, HP89b, MLRC88]. High [ACM04, Ado89, ATS94, Ano93-30, BS92, Bar01b, Buy99, CCF89, Dan83a, Dan83b, Dow93, GS01, Gig06, Gil96a, Har80, IBM09a, IEE94a, Ivi95, Kar92, The92c, The92a, The92b, The93, May00, Min82a, Min82b, MZA96, MK84a, Nag05, PBL86, RF93, SHB23, Slo05, USE94b, Wey01, Wil83a, Wil83b, AOAH14, FGP^{+}91, GS02, KPT^{+}02, Kob99, PF05, Way00, ZGdDL^{+}96, Pos88, AOAH14]. high-BDP [AOAH14]. High-Energy [RF93]. High-Level [MK84a]. High-Performance [CCF89, Gig06, IEE94a, Min82a, Min82b, The92c, The92a, The92b, The93, PBL86, KPT^{+}02, PF05, ZGdDL^{+}96]. High-Speed [MZA96, USE94b, AOAH14]. High-Tech [Bar01b, Way00]. Highly [BW94a, BW94b, ELS88, HK95, Sat90b, HBG^{+}06a, Kop05]. Highly-Available [BW94a, BW94b]. Hill [USE02, C^{+}78a]. Hills [Ano01]. Hilton [IEE83]. historic [And17]. History [Cer98, Cer03, Fel84, HH97b, HB82, Kap92, Ker19, Lév18, Mahxx, McK07, McK16, Mil84c, MT78, MT79, Pom04, Tan84, Ano90b, HH97a, McK85a, Spi17, War85a, War08]. HITAC [KAH83]. HOC6 [Pug88b]. Hodgkin [HKT05]. Holdup [Ano94k]. Hole [Sta04]. holes [AK^{+}07]. Home [FES^{+}99, EBS08, Lum13, NW03, Neg06b]. homogeneous [BBD23]. Homomorphic [PS22]. Honey [End00, HNR84]. HoneyDanBer [RW86b, RW86a]. Honeywell [Wag79]. Honor [Pin99]. Honors [Coc03b, Ano19c]. hood [McE05]. hooks [Ano93m, GJJ05]. Hope [Far92d]. Hopes [Law98]. Horizon [Ano94e]. Host [BBP99, Chu03, Mar78a, OBSB23, TSO^{+}96, Shu06]. Host-Based [Chu03]. Hosted [Coo98]. Hosts [Höf01, Mue79, SH13]. Hot [Roe02, Wis03]. Hot-Failover [Roe02]. Hotel [Mic79b, USE01b, USE01a, USE02].

Hotswapping [LC06]. Hours [CS99, Art99, Bal99, Gay99, Hsi01, SBR01, TA98, Tay03, Tay06, UT03, WT99, Wel99a]. House [Eps92, Ste01f]. Householder [Kau7 , Kau79a]. how-to [Pon97c]. HOWTO [TV04]. HP [Kuo99, Ano91b, Ano93m, HP-93, Ano93n, Ano96f, CM04, FR75, Kee05, Ket99, Kul06, Lei96b, LMOT01, LMOT02, Mad97, Olk03, Pon94, Pon96, Pon97b, Pon97c, Pon98, Pon99b, Pon99a, Pon01, Pon02b, Pon03, Pon05a, Pon05b, Reh00, Reh05, Ric93, Ric94c, Ric95, Rob00, SW99, SRW04, She06, Tho03, TOML04, WLHC84, Wey01]. HP-21 [FR75]. HP-UX [Kuo99, HP-93, Ano96f, CM04, Kee05, Ket99, Kul06, Lei96b, LMOT01, LMOT02, Mad97, Olk03, Pon94, Pon96, Pon97b, Pon97c, Pon98, Pon99b, Pon99a, Pon01, Pon02b, Pon03, Pon05a, Reh00, Reh05, Ric93, Ric94c, Ric95, Rob00, SW99, SRW04, Tho03, TOML04, WLHC84]. HP/OSLO [She06]. HP/UX [Ano93n]. HP9000 [Hib96, Lin84]. HPBSD [Hib96]. HPC [CW15a, CW15b]. HTML [Kro99, Law98]. HTTP [Ste96]. hu.man [Dys96]. HULK [VTS^{+}23]. HULK-V [VTS^{+}23]. Human [BSK87, Fra83, Gin83, Mac83, KD94, WBSK87]. Human-computer [BSK87]. Hundreds [Jae86]. Hungarian [Lás05]. Hurd [Epp00a]. Huxley [HKT05]. hybrid [HWBH19]. Hype [Gla99, KJHL96, KJH^{+}97]. Hyper [AL97a]. Hyper-Impatient [AL97a]. Hyperbolic [Cou84, Cou85a]. Hypercubes [HO93]. Hyperform [WL97]. Hypermedia [WL97]. Hypertext [Bro89a, Bro89b, FES^{+}99].

I-nodes [Len97]. I-O [Car82]. I.1 [HRW77]. I.4 [Wan77]. I.5 [Sab77]. I.9 [Pis77]. I/0 [Lyc74a]. I/O [Ano93e, Che87b, Fed77, KSU94, May00, McS79, Ric96, Rit5 , Rob84a, Sch93b, Sch93c, Tin78a, Tin78c, WMK^{+}07, Yoo96, dBB08, vMM88, vM88, Rob06b]. i80sim [Hof77]. IA [Ano00d, Ano00f, Gig06]. IA-64 [Ano00d, Ano00f, Gig06]. IA32 [ST03]. ia64 [EM00]. iAPX286 [Bar83]. IBM [Ano93f, AJ05, Ano96f, BS79a, BBB^{+}20, Coc03a, D^{+}04, EGL86, G^{+}06a, Gra79c, HR94, Hub79, AIX91, JH83, Lev88, Lyo18a, Lyo18b, Lyo18c, Mly09, P^{+}08, Q^{+}11, RH95, Ric94a, Rob75, R^{+}02, Sab76a, Sta05b, Tan87a, WMK^{+}07, Wil83a, Wil83b, Yag91, ZST^{+}07]. IBM/370 [Rob75]. ICCCCM [SGFR90, SGFR92]. ICE [Ano06a, Den79]. ICE-85 [Den79]. ICFP [Coc01a]. Icon [GG83, GG86, GG90, Mit84]. IDA [FKL^{+}08]. IDE [Ano02c, Ano02b, Ano03e]. idea [Tay96]. Ideal [GMM90, Van81, Van90]. Ideas [Lew21]. Identification [HLT^{+}21]. Identifying [Hor90b]. IDEs [Sor01]. Idioms [Cop92]. IDS [Reh03]. IEC [ISO05]. IEEE [ACM04, Ale00, Ame93, Ame94, Kuh91, LL95]. IEEE/ACM [ACM04]. If [Arn91]. IFIP [BSK87]. Ifmail [Cro98a]. Ihre [RB06a]. II [HR79b, USE91e, BGMR91, Cat79, CL79, Ell79a, HO93, HP89a, HP89b, uRH96, HA95b, JJ91c, Lin79b, Lin79c, Neg06b, Ng77, Roo79a, Sul02, USE90e, Wal95e, Zuc77b]. III [BBJ86, Mad94, Ric77, BBB^{+}20, CQ83a, Fla90, HA96, USE89d, USE92e, USE93d, Zuc83b, Zuc83c]. Illinois [USE99c]. Illustrated [Cal00b, Ste94, Ste96, WS95]. illustrating [MD86]. ILMON [BM87]. ILSG [CH77]. im [KGG00]. Image [Bee86b, Gom85, GIM07, Goe07]. Images [Hol90a, Ran06]. Imaging [Kro99]. imake [DuB93]. Impact [CKM85, HH97b, Isa86, HH97a, LRB^{+}19]. Impatient [AL92a, AL95a, AL95b, AL97a, AL97c, AL91, AL97b]. Implantación [VD01]. Implantation [VD01]. implementación [RÓ01]. Implementation [ASS85, Alf94, BJP04, Bar83, BMS95, Bee90, Bon79a, Bra76a, Car85, CSA^{+}87, CW89, CW91, CCDP01, CFA85, CS91b, CGMS89, Cro98a, Dav97, Don95, Fel79b, Fis86c, GB92, GD18, GG86, GLG93, HLMR79, Has78, HR79a, HL85c, HK89, Gur81, IB93, JHT^{+}07, Jun85a, Jun85b, Kar83, KM86a, Kep85, KN93, KM87, LKY93, LP89, Lef81, LMKQ89, LMKQ90, Lin84, Loi79, Lyc74a, MTY01, MBKQ96, MNN05, MNNW15, Mue93, NN92, Now78, NL82, NHR84, Par88, PV00, PBL86, Ral76b, RO92, Ryd72, SW88, SBMS93, SSB92, SWKH76, Sun96, TW97, Tho78, WLC95, Wil83a, Wil83b, WS95, YLK93, YSF89, Zuc77b, CLJC95, CP01, CP02, Daw82, FMM84, Gab07, HBMT96, Jal92, KAS^{+}06, LM91, Lov10, McK91, McK92a, McK92b, MK96, NSI^{+}92, PC01, RÓ01, Sch93b, TW06]. implementation [WLHC84, Wan02]. implementations [Ano19b]. Implemented [Sha79b]. Implementierung [LMKQ90]. Implementing [Car04, Dav00, D^{+}04, Dwi04, Get86, GMB93, HRO82, Lei96c, Lei96e, Lei96d, Rob79a, Rob82, Wai95, XZY^{+}10, LR95]. implements [Wei97]. Implets [CN88]. Implications [DNQ^{+}83]. Important [Sat93, CK06b, Mud97]. improve [GK97, OS06]. Improved [Jon84, Mar78f, Pea83, Ano95g]. Improvement [HR79b]. Improvements [Kra78a, MKL85b, MKL85a, Web95]. Improves [Flo90]. Improving [BRW89, LKM84, McK13, SKP94]. IMS [Ano86]. In-Core [Roo79a]. inch [Kri84a, Kri84b]. inches [Ano91c]. Incident [TLR89]. included [Fur85]. includes [SML06]. Including [Yag91, Geh84]. incomplete [Moh85]. Inconsistencies [Ols78]. Incorporating [Kra88a, Kut02]. Increasing [Sat90a]. increasingly [Bal96]. Incremental [FST91, Hum88a, Cro82]. Incrementally [QL91]. independence [Sch95c]. Independent [IBC87, KAH83, USE93e, Lin02b, Lin02c, Ker82b, Mee95]. Inder [Ano94a]. Index [Bea93, Che90b, Che92, Cox12, Mar78b, LPG^{+}18]. Indexes [BK86a, Les79a]. Indian [C^{+}78a]. Indicators [BBP99]. Indigo [SZ95]. Individual [Gil93b]. Industrial [Lav04, Fli03, Ran04]. industrial-strength [Fli03, Ran04]. Industry [Ano95c, Bar00b, Che94, Dun91a, Dun91b, Edw98]. influence [Mly09]. INFO [Env97a, Env97b]. Informal [Fie88, WC93]. Information [CBV93, HK83, Hoo83, Int90, Int93, Lun93, Lun99, RW86b, RW86a, Str79, Ame93, Ame94, IEE91c, IEE92a, IEE92b, Koh99, NIS92, Sch94a, Wal01b, War08]. Informix [IBM09b, Tar89, Fla00, Kin83]. INFORMIX-SQL [Tar89]. Infrastructure [Coc01c, Gar99, MKSB22]. Infrastructures [SHB23]. Ingol [Hid04]. INGRES [SWKH76]. INGRID [DBG89]. INIT.C [Pug88b]. Initial [AD89, BG97, Cou84, Cou85a, Hau94, Ano03c, JJ91t, JJ91u]. Initial-Boundary-Value [Cou84, Cou85a]. Initialization [Ima76, DIN05]. Initiative [Law98]. injection [KIT93]. Inkscape [Bah09, Ben09, Kir09, Tom09]. Inline [HC89]. Inlining [Tho03]. Inn [IEE83]. innovation [And80]. Innovators [Pin99]. Inodes [Lud78b]. Input [Eis79a, KMKO92, Rit77b, Rit90, Tuo83, Whi75a, Rit84d]. Input-Output [Rit90, Rit84d]. Input/Output [Eis79a]. Inquiry [Che79c]. ins [DF00, RP84]. Insight [IEE02]. Inspection [Kur90]. Installable [CE97]. InstallAnywhere [Ano02b]. Installation [BH86, BD94, Bor01, EDK00, GA04, Gay93, GES98, HR73, HS83, KGG00, Luc03b, Maj77, MWCR90, MWCR91, SU94a, Str94, SUM95, SE96, USE87c, USE88d, USE89d, USE90b, USE91b, USE98c, VRJ97, WF03, Kof00, Mag00, Ron05b, VRJ95, VRJ96, Wel94b]. Installations [HD97, KM95]. installieren [Ano03b]. Installing [AD89, KTB^{+}88, Hag00]. InstallShield [Ano02c]. Instruction [KL79, McI60, Ros78]. Instructions [Sto76].

Instrumenting [MZG14]. Instruments [Kod82a, Kod82b]. Integer [McI83]. Integral [Blu7 a, Blu78a, HCN85]. Integrate [Hur94]. Integrated [Ano01, Ano02c, BW89, Boy84, CCH^{+}77, SZ95, Ano94d]. Integrating [Ano98c, BT98, BCGS91, Hel97, Kar92, Kem93, Lee89, LBSL97, Lun96, McC01, Sat89, Wit01, TSC02, Tho95]. Integration [Ano01, Ano02b, Aud96, CPJ^{+}98, Don85, GBG97, Koh06, Lob85, Mei92, Sha00a, URK85, BHR95, Bru02, Har00, KS08, MB05, Sim96b, Sun93b, Tuc98, WG98, YHR92]. Integrity [FE77, Pon05b, McK85b, Pon05a]. INTEL [Mul79b, Mul79a, Ano02a, Ano03e, Bar00b, Bar83, BM79, Hua06, Law98, Mil79, Mol77]. INTELLEC [LT76, LTW77]. Intelligent [Ano02c, BFGK89, BNB87, HP98, Mar78a, Wes78, KD94]. Intensive [HIT^{+}02]. Inter [Lyo77, LR77, NH99, Tin78b, TG78, Wid90, Wit79]. Inter-Client [Wid90]. Inter-Machine [Wit79]. Inter-Platform [NH99]. Inter-Process [Tin78b, TG78]. Inter-UNIX [Lyo77, LR77]. INTERACT [BSK87]. Interaction [BSK87, ESSB93]. interactions [KD94]. Interactive [Ang97, Ang99, BRD77, Che76a, CP86, Fas78, FL80, HK83, Jac77a, Jac77b, Kap99, Kra78b, Lib94, LS87, McM99, Moh99, MC78, Pat99, Rad79, Ste85, Vic98, McM78a, McM90, Orw96, PWC00, CH77, Kom02]. Interactivity [Hos84, TCM07]. InterBase [BCE^{+}93, BCP93]. Interconnecting [Bra93]. interdependencies [dJPDD84]. Interface [ACM88, Ame93, Ame94, AW89, Bad98, BEHW86a, BEHW86b, BNB87, Com94a, Che87b, CS86, GSN90, HKF84, HM90a, Hug88, Int90, Int93, IEE88, IvW87, KM86b, Lal93, LT76, LTW77, The92a, The92b, MS89, Mul79c, PCD92, Per82b, Rad79, Rug83, SA93a, SS93, Wag79, WLC95, X/O89c, Ame91, AGS90, Biv87, CH96, Ewi86, IEE92a, IEE92b, Kat82a, Kat82b, Kem91, Ker10, MS87, Nat88, PH91, SNM22, Sch93c, TTG^{+}87, Ano92i, HG95b]. Interfaces [Ano94g, BFKS93, IEE92a, IEE92b, MS95, SAS93c, SH13]. Interfacing [Duf82, Hol02, PR85a]. Interference [Lyc72]. Interlinking [Joh96b]. Intermediate [Ros98]. Internals [BBD^{+}96, CS91b, GC94a, GC94b, MM01a, Pat96, Vah96, Bar00a, B^{+}98, Ben06, CM04, Gra95, Kel96, Lei00, MM07, McK91, McK92a, MK96, McK96b, McK97b, McK97a, Sin07]. International [Ano86, Ano93g, Ano95h, DFG99, D^{+}76, ITC93, IEE94b, IEE03, IEE04, ISO05, LT80, MM06a, Mil84b, SAS93a, SD95, TS97, Uni96a, Uni96b, Uni97, Uni98a, Uni98b, Tut93, YLM^{+}05]. Internationalization [Ano98d, Hiu96, Hiu97, Ric98]. Internationalizing [Sch95a]. Internet [Ano96b, Bee93a, CB94, CBR03, CK06b, DF00, GSS03, Gil93b, Gla97, KGG00, Kro92, MGK^{+}03, Qua93, Uni96a, Uni98b, And95, Ano96d, Ano96f, Bou97, CK06b, CW00, CP01, CP02, Don96c, Eck95, Gar97, GS96, HBMT96, HH97a, Hon06, Hub96, Kha02, MP93, Mat93, PB21, PC01, RR95a, Sun01, SCJ96, SG95, SH95, Sku88, Uni96b, Uni97, Uni98a, VRFJ97b, Vol97, Wan96, ZB01, ZCC00]. Internet-based [Sun01]. Internet/VB [Ano96d]. Internetdiensten [Luc03b]. Internetkommunikation [CK06b, CK06d, CK06c]. Internetprogramme [CK06b]. Internetprogrammen [CK06d, CK06c]. Internetworking [Aud96, CS91a, CS91b, Mad94, DeR99, Pab96, San91, San94, Ano93-32, Bra93, CMQ93]. Interoperability [Anoxxb, Bar00b, BM02, Sch95c, Wis03]. Interpret [MC85a, MC85b]. Interpretation [FY18, Jon84]. Interpreter [BBM84, Feu85, KLP88, Kol77, KGP90, McD87, QBdR87]. Interpreter-based [KLP88]. Interpreting [Sax85a].

Interpretor [BQd86]. Interprocess [Ano77, Bas87, BMS83, Bra76b, De 78, Fre90, Gra97, Gra98e, HR78a, Joh96a, JF81a, Kau78, Lin79c, PR85b, PR85c, PR90, Sau88, Ste99b, Sun77, Wat87, Zuc77b, Bas91, IBP03]. Interprocessor [Ung78]. Interprozess [Bas87, Bas87]. Interprozess-Kommunikations-Mechanismen [Bas87, Bas87]. interrupt [RLB08]. intervals [McK92b]. Interview [CUH99, JMR19, Mah89, Ray99b, Kar84, Pub11, Vib11]. Intervoice [Ano03d]. Intranet [McC98]. intrepid [Kon12]. Intrinsics [Kel90, NO90a, NO90b, NO92, NO93, O’R90b]. Introduces [Ano01, Ano03c, Ano03f]. Introducing [MS92, MM83a, MM83b, MW08, McL84, WALK10]. Introduction [BB99, Bou7 , Bou78b, BA96, CK06a, CK06b, Dav97, Feu79b, Ger82, HHHW91, Hum90b, Isa91, Jae93, Jon89, Ker79c, Ker7x, Kil93, MB77, Ral75b, RB05b, RRF90, Ros88, SK10b, SF85, Sch99, Sea20, Seb94, TBR93, Wat87, Wen79b, ABS95, Ber86, Geh84, GRA89a, Gou04, Hav11, Hoo05, Kar06, Kon07, GF80, Ros96, Sho12, Sym85, Wan96, Wri90, Wün13, Ros77a]. Introductory [PKP05, Sil84, Bai87, SCJ96]. Intruders [Far90]. Intrusion [BF03, MTY01, Nor99a, NZW^{+}02, Reh03]. Invariant [DBG89]. Invasion [ASW^{+}01]. Inventors [Ano19c]. Inventory [Gop79]. Inverted [Les79a]. investigation [PBWB84]. Investigations [She06, DC18, LRB^{+}19]. Invited [FK90]. Invoking [MMTW88]. IOP [Lon78]. IOP-DUP [Lon78]. iOS [Ano16a, AHA^{+}14, DLCV17, DC17, DC18, Hal11, HBH18, HWBH19, Jen19, KMP^{+}20, LPG^{+}18, LZHS11, PRC^{+}23, Pog17, SW12, SEKV12, WWCW19, Ano16b]. IOStreams [LK00]. IP [Mad94, Str94, ATS94, Ale96a, CMQ93, CS91a, CS91b, CP01, CP02, FFH^{+}86, Fer85, FES^{+}99, Hor98, Hun92a, Hun02b, JLSG84, KM86c, LP89, Mor85, PC01, SH13, San91, San94, Spu92, Ste94, Ste96, Sto95, TP91, WS95]. IP-forwarding [SH13]. IP/TCP [KM86c]. IPC [Rai89]. IPCs [Sch93c]. iPhone [KMV^{+}09, ML09, PB09, ZL10]. iPod [KMV^{+}09]. IPSIS [But02b]. iptables [Pur04, Ras07]. IPv6 [Siv03, Sta05b]. Ireland [USE87a]. Iron [Ano84, Miy86]. irreverent [Moh85]. IS/3 [Zuc83b, Zuc83c]. ISAM [Tar89, Wil82a, Wil82b]. ISBN [Ano16a, Ano16b, Hid04, Jen05, Koh06, Sch04, Sta07b, Tho05, Wei95]. ISCC [Nel79]. ISDN [KGG00, Don97b]. ISIS [Mul79c]. Isn’t [BW94a, BW94b, Hoo83]. ISO [Hus92]. ISO/IEC [ISO05]. ISO10646 [Uni98a]. isolating [ALL13]. isolation [GNDB16, WRA05]. ISP [HD02]. Issue [Ede77, Gla77b, Ng77, Row79a, UNI77, Vog76b, Vog79a, Ken90]. Issued [Gar97]. Issues [Fer82, Gar97, GZXZ23, GMB93, LR90, TB05]. Italian [Cor05b]. Itanium [Ano00d, Ano00f, Ano06a, Cha06, HKS^{+}04, Kul06, LMOT01, LMOT02, Pon05b, Tho03, TOML04]. Itanium-2 [LMOT02]. Itanium-based [Ano06a, TOML04, Pon05b]. Iterative [Mer82, WB96]. ITIL [Van05]. IUC12 [Uni98a]. IV [USE93g, Hal69, USE91a]. IX [FPG89, MRxxb, MRxxa, RM92, RMxx, Reexx, Lal93, SW88]. IZIC [FKM95].

J [Sau88, Wei93]. J2EE [Ano02b, CR02]. J2EE-Compatible [Ano02b]. JADE [Ung86]. Jails [Luc19]. James [Fla90]. Jan [Wei95]. Jan-Armin [Wei95]. Janet [Wei93]. January [ACM97, Ano95a, USE84b, USE85c, USE86e, USE87e, USE89f, USE90d, USE91d, USE92d, USE93f, USE94f, USE95b, USE96b, USE97c]. Japan [Ano11, Uni98a]. Japanese [AKWA89, DF91, IBC87, JK86, Lun93, Lun99, Sas91c]. Jargon [WC93]. Jaross [Ano95d]. JAVA [Ano98c, Gla97, SML06, Ano98d, Ano01, Ano02c, Ano03a, Ano03d, Ano03f, Che97, Cox07, Jen00, Jen02, Jon02, Kro99, Lut99, McG97, OH97, Red97]. Java/C [Ano01]. JavaScript [Ric14, Gab07, Hav11]. JDKs [Gla97]. JDS [AH04]. jedermann [Str94]. Jeopardy [Gar97]. Jermy [Koh06]. Jersey [MT93]. jetzt [KGG00]. jewel [Geh03]. jian [PWC00]. Jiao [RB08]. Job [FES^{+}99, KK89b, Lyo76b, Sab76a, Sab78]. Jobs [Bar01b]. Joe [Red97]. John [Hid04]. join [ITC93]. Joint [Usr82]. Jose [Uni96a, Uni96b, Uni98b]. Joseph [Nag05, Ano94d]. Journal [Bee96a, Bee96b, NBB^{+}02, Lin94]. Journaled [MR10, MR14b]. Journaling [LBN14, SGM^{+}00]. Journals [LK76a, LK76b, LK77]. Journey [Kor97a]. Joy [Man93a, HP01]. JPEG [SW91b]. Judy [Nel90]. Julich [IEE91b]. July [MM06a, Sof83, Usr82, USE99a, Kap92]. Jump [Lin93]. JumpStart [HN01, KM95]. June [ACM82, ACM92, Ano94m, DFG99, Eis79b, IEE91b, Sof84, USE85d, USE86c, USE87d, USE89e, USE90c, USE91c, USE92b, USE93c, USE94e, USE95c, USE95a, USE98b, USE98e, USE00a, USE01b]. Jungle [Hun92b]. JUnit [ZK05]. Just [And02, Swa03, TD01, Hoo83].

K42 [AAB^{+}05]. KAI [Ano98d]. Kanji [JK86]. Kanotix [CK06e, CK06e]. Karl [Jen05]. Karmarkar [Gay87]. Kaspersky [Ano00e]. Kaufman [Ano95b]. Kay [Nel90]. Kaydara [Ano02b]. KDE [KGG00, APK99, Böl01, Gri01a, JSG00, MWM00, Pow00, Swe01a, Thi99, Wel99a]. KDE/QT [Gri01a]. Keep [Bar01b, Spe96]. Keeping [Kar84, Sta06, Jan06b]. keeps [Ano93z]. KEIO [MTT83]. Keith [Gra89b]. Ken [Ano11, CUH99]. Kerberos [Ano93d, De 93, Wit01]. KERMIT [Gia90, dC87, dG93, dG96]. Kernel [ABB^{+}86a, ABG^{+}86, ABB^{+}86b, Alf94, Ano19e, BLH20, BBD^{+}96, BC02, BC03, CDM98, CMR^{+}21, CMM88, CWS90, DYZC20, EM00, Eyk88, FKN85, FH88, GZXZ23, GD18, GLOR91, Hal11, HHLS90, HLT^{+}21, HMP83a, HMP83b, Hua06, JKS^{+}15, JJ91p, Kam02, KLK^{+}02, Kor89, LJ84, LKY93, Lin84, LS92, Liu11, LS83, MAP^{+}22, MM01a, Max01, MFBWW20, McK84, MK88, MMTW88, NLR84a, NLR84b, Nut01, PQM^{+}21, RMF19, Req85, Ric96, RFS05, RFS06, Sch94b, Sch95b, SGWP02, SKP94, TBG^{+}87, WLL96, WB92, XZY^{+}10, YLK93, YR93, Zim85, dOCC23, ALL13, Ano94h, Ano03e, AAB^{+}05, B^{+}98, BYVF08, BC00, BC06, BV06, CDD^{+}05, CDD^{+}07, DFC^{+}23, Gai84, IF10, JGS^{+}08, JJ91s, JJ91b, JJ91c, JJ91a, Kon07, Kor06, KH07, LC06, Len97, Lov05, Lov10, Mau08, MM07, MW08, McK91, McK92a, McK92b]. kernel [MK96, McK96b, McK97b, McK97a, PM10, SWF^{+}22, ST03, SS95b, BC01, HBB13]. Kernel-Based [Alf94]. Kernel-level [CMR^{+}21, Ano03e]. kernels [YSC^{+}06]. key [Tuc90, Kri91]. Keys [Luc18b, Luc12]. Keyword [Law98]. killer [McI99]. Killing [Ano96e]. Kilnam [PB21]. Kiosk [RR95a]. Kirk [JMR19]. Kit [Ano02c, Dul06, SU95, TvK83, WIP^{+}97, Gay99, Gra87, Lei95b, Lei96a, Lei96b]. Klassische [RB06a]. KMC11 [Weh77a, Weh77b]. Knoppix [CK06f, Neg06a, Neg07a, CK06f, Ran04, Ran05]. Know [Gra03, Wal95a, Wel99b, Zie93, Lov07, SF95, Sem96, War04, War15]. Knowledge [HQZ^{+}87, Sam87, RS14, Sam89, WT91]. Knowledge-based [Sam87, Sam89]. Knows [Coc01b]. Knoxville [IEE94a]. Kommandozeilenvirtuosen [Die11]. Kommunikation [Lin02a]. Kommunikations [Bas87]. kompromisslos [Bra04]. Konfiguration [Bor01, Luc03b, WF03]. konfigurieren [Ano03b]. Kopper [Jen05]. Korean [Lun99]. Korn [Die11, May01, Ano00g, BK89, Col94, May01, Mic08, OP01, Olc91, Olc92, Olc01, Ros91, Ros93a, Ros94a, RR02b, Smi06c, TR94, Val92b]. Korn-Shell [Die11, May01]. KornShell [BK95]. KornShell93 [Ros99]. Krause [Ano00b]. ksh [Smi06c]. Kubuntu [CK06g, CK06g]. Kummerfeld [Nel90]. Kurz [Rob02b, BW04, Gün07, Rob99d]. KW11 [Bab78]. KW11-P [Bab78]. Kylix [BJ01].

L [Ano11, Mad94]. Lab [Ano96d, Ano00e, Coo09b, Coo09d, JLSS77, Tom11, FJPR93, Lyo79]. Laboratories [Ham87, HB82]. Laboratory [Gwo78, HG95b, SMD^{+}94, FES^{+}99]. Labs [Geh03, PPTTxxa]. Lake [ACM03, Sof84, USE84a, USE95c, USE95a]. LALR [Mey18]. LAM [RšT06]. LAM/MPI [RšT06]. lambda [CB83]. LAMP [G^{+}06b, LW03]. LAN [Len97, Lun96, RÓ01, dC93]. Land [Ano97a]. Landing [ACM03, Coc01a]. Langages [Bla02]. Language [Ame89, AKW78, AKW85, AKW88, Ang77, Bec84, BK86b, BK90a, BK91a, Bey87, BK89, BK95, BQd86, BSW78, Car77, Cha69, Che97, Che76a, Che76b, CM78a, CH83a, CH83b, DFV91, DM79, DK16, Dzo09, FGP64, Far78, Fed79, Fel79d, FG79, Feu85, FGK89, FR75, Fra79, Fre78, GJ75, GSN90, GG83, GG86, GG90, Hol90a, Hor82, HL85c, Jae89a, Jae89b, JK73, JRL77, JL78, Joh80, JR81, JL87a, Joh87, JL87b, JJ91q, Kap92, KLP88, KR78, Ker7x, Ker81a, Ker81c, Ker82a, Ker84a, KR88, Ker90b, Ker91, Koe90, KM87, Lef81, McI72, Ora21, Pik93, Rad79, RJLK78, Rit79a, Sab76b, Sco09, Set79a, Set79b, Set79c, Sny75, Str84a, Str86b, Str91a, Str13, TH77, Tho76, Tob80, VK79, Van80, Van90, Ame94, AKW8 , AKW79, AKWA89, Ale10, And80]. language [Cox83, Dun00, Env97a, Fel8 a, Fel79a, FF88, Geh84, Ger92, GPP68, GPP71, IEE92a, IEE92b, Jou86, Ker81b, LR95, Lar94, MM86, Mee95, QBdR87, Rit93, SAS93b, Str80a, Str82, Woo94, Zdz05, TG89]. language-independent [Mee95]. Languages [CPJ^{+}98, FG79, KP99, McI60, X/O88c, BHR95, Fis73, HLHR90, Inc83, Jou86, Mud97, Ngu08b, ?, ?]. LANs [Hur94, Hus92]. LAPACK [AD89, BD94]. Laplace [Blu7 a, Blu78a]. Lar [Ano95b]. Large [Bec84, GVK^{+}03, GES98, GP88, HD97, Hun92c, LT80, Loi79, Lyc74a, Mor78b, PMD88, PF84a, PF84b, Sat89, SSB92, SW91a, STA86, USE87c, USE88d, USE89d, USE90b, USE91b, USE97a, USE98c, CR92, DFC^{+}23, WWCW19]. Large-Scale [GVK^{+}03, GES98, GP88, Hun92c, CR92, DFC^{+}23]. LaserJet [CJ94]. Latent [Fil85]. Later [FK90, Lyc85, MFBWW20, PTS^{+}11]. LaTeX [Str94]. lattices [Pon97a]. Lauderdale [USE89c]. Launches [Ano01, Ano03g, Coc01a, Gar97, Ano03e]. Law [CPJ^{+}98, Bar01b, MSSvK08]. Layer [IB93, Noe98, Ano03d, Wei97]. Layered [Lin84]. layers [Pik83a]. Lazenby [Ano00b]. Lazy [Dat88]. LCA [Mag04]. LCP [Mag04]. LCP/LCA [Mag04]. LDAP [BH00, BH01, Cov06, HSG03]. LDP [Ray98]. Leading [Kra83]. LEAP [CH77, Mar78g, Mue78a, PW77]. Learn [Hae92, LH01b, O’D83c, PB09, Zin92, KL79]. Learned [PCD92, Coo03]. Learning [Ano00g, CR91, Gar91, HLT^{+}21, Lam90, MAP^{+}22, McC99a, McC99b, McC02, McC03, McC04, NDW93, NR95, NR98, NR05, PTS98, PTS01, Pon97b, Ros93a, Ros94a, RR02b, Sch93d, SHVP05, TPTS02, TJ03, TJ04, Tay05, TSP93a, TSP94, TPS98, TSP93b, BBD23, Cha11, Gre96, SWF^{+}22, WNDO97, W^{+}99, Gil93a]. Learning-Based [HLT^{+}21]. least [Ale96b]. Lecture [Ral77, McK91, McK92a, McK92b, Swi91].

Legacy [McG97, Che97, Gil96a]. Legal [Pom04]. legend [Jen19]. Legit [Sta06]. Lehrbuch [May01]. Leitfaden [DLvG07]. Lempel [NT05]. Length [Bey77, Koe78c]. Leopard [EBS08, Pog07]. Lern [Kuo99]. Lern- [Kuo99]. Lernprogramme [CK06h]. Lernprogrammen [CK06h]. Less [Ano93i, For00]. lesson [And80]. Lessons [Bor88, GLOR91, PCD92, Rei86]. Let [Jae85, Sau17]. Lets [Wil94]. Letter [Dav96, McI69a, San93]. Letters [ASW^{+}01, CPJ^{+}98, DM97, FES^{+}99, PLBL92]. Level [AISS98, Jac84a, Jac84b, KSU94, MK84a, Ros98, ST89a, ST89b, ST89c, SM89, SPS^{+}02, Ano03e, CMR^{+}21, Gai83, Mag00, MRxxd, RW81, Sim96c, Sul02]. Leverage [Mas87]. Lex [LS79b, Sta89, Sch90, Jac87, Les75a, LS79a, LS90a, Sta89, LMB92, MB90]. Lexical [LS79a, LS79b, LS90a, Wai86, Les75a, Pax88]. Lexicographic [Zuc77a]. Libckpt [PBKL95]. libero [Cor05b]. LIBFTP [Vol94]. libm [LMOT01, LMOT02, TOML04]. Libraries [Arn86, FHS7x, Ano95f, CB89, LRB^{+}19, PS07]. Library [Ano98d, Anoxxb, But02a, Cod86, Eis79a, Fla94, FHS77, FHS78a, FHS78b, Fox79, Gla77a, Kau93d, Koe78a, Les88b, Mor75, Mue93, Per79, Pla92, SAS91a, SGN88, SST90, TOML04, TG79, WL94, Ano03e, Ano19b, FF88, Ham90, LMOT01, LMOT02, Lov07, Pay84a, Pay84b, SG97b, ?, ?, Bec06, Bec07, Hal00, Hel03, Jos99, Jos12, Kar06, Koe79b, Ree06b]. Libtool [VETT00, Cal10]. LIBTP [SO92]. License [RP84, Guy00]. Licensing [Isl83]. Life [Gwy90, LR89a, LR89b, Lut99, PB84, Tan87c, Too11, Geh03]. Lifecycle [Kro99]. Lifetime [HBD97]. Light [BS90]. Light-weight [BS90]. Lightweight [HHLS90, JKS^{+}15, JZTB98, Kep85, OBSB23, SMK^{+}94]. Like [Hol87c, Jon84, Sta04, CKM06, CKM14, Len97, MK91, Neu00, OH97, MLRC88, HK95, Kor89, LM88a, LM88b]. likely [Pub11, Vib11]. limitation [Len97]. Limits [BAC^{+}89]. LINCS [Req85]. Linda [CG93a, Lel90]. Lindows.com [Ano03e]. Line [AW89, Blu11, CAG89a, CAG89b, Dow77, Hub79, Lon77, Pru02, Shi77, Stu78b, Uhl78, Ano96e, Ano97b, Blu08, CGMS89, McE05, NM83a, NM83b, Sho12, Ste99a, Tal89a, Tal89b]. Linear [Gay87, Kau93d, Lit80, dVd93]. Liners [Kruxxa, Kruxxb]. Lines [Kul06]. Link [Ano00e, Erm79, LC77, LC78a, Shi77, Len97]. Linkage [Mor78c, Lev79]. Linking [QL91, Sha93]. Links [Fer85]. LinkScan [Ano00e]. Linksys [Com12]. Linotron [CKT80]. Lint [Joh78b, Kor89, Dar88]. Lint-like [Kor89]. LINUS [Kra83, CPJ^{+}98]. LINUX [Bra04, KGG00, Str94, Ano00b, Ano00c, Ano00e, Lin02b, Lin02c, Ano05, Ano06a, Bar01c, BW02, CV00, CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, Cor05b, DW06, DWV06, DF00, EEK^{+}04, GA04, Gan04, G^{+}06a, Gri01b, GGZ02, HS99, HS00, HS04, Hid04, HHV05, uRH95, uRH96, Jen05, Jor04, Kad95b, Kad95a, Kle00, Koh06, Kuo99, Lin02a, Mag04, May01, Nag05, Neg01, Neg02, Ric14, RÓ01, Ron05a, Sav11, Sch04, Sob03, Sta07b, Tho05, Hag00, WF03, Wei04, Kof00, McN99a, Mea99, Lás05, RÓ01, Smi06b, SU94a, Str94, Wel94b, Abb03, ALP^{+}21, AH04, AT07, ALL13, Ale00, ASW^{+}01, AOAH14, AC11, AJ05, And17, Ano98b, Ano99a, Ano99c, Ano99b, Ano01, Lin02b, Ano02a, Ano02b, Ano02c, Ano03a]. Linux [Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano19b, AAB^{+}05, AR07, Asb00, Aul01, BLH20, Bal99, BP00, Bal00, BBD^{+}00, BN01, Bar00a, Bar01a, Bar00c, Bar01b, BSB03, BW04, Bar04, Bau02, Bau03, BsswLB05, BDP05, Bea04, BBD^{+}96, B^{+}98, Bee00, Bee16, BMF15, Bel00, BYVF08, Ben06, BFS07, Bes05, BBHL08, Bla02, Bla04, BTS09, Blu00b, Blu08, Blu11, Bok95, BR99, BJ01, BB99, Bol99, Böl01, Boo03, Bor01, BS05, BBB^{+}20, BC00, BC01, BC02, BC03, BC06, BM02, BV06, Bri05a, Bri05b, Bro06, BDD^{+}06, Bru02, Bur04, But02a, But02b, BF97, CF20, Cam04, CDM98, CDD00, CMR^{+}21, CM06, CDD^{+}05, CDD^{+}07, Cha11, Cha06, Chi02, CK06a, CK06b, C^{+}03, Chu03, Coc01b, Coc03c, Coc03b, Col02, Com00, Com03, Coo98, Coo09a, Coo09b, Coo09c, Coo09d, CRKH05]. Linux [Cor05b, CP01, CP02, DH02, Dan01, DJ01, DRP01, Dea01, Dec04, Ded09, Den99, DG00, Dik06, DIN05, DL08, DLP10, DRN^{+}23, Don04, Don22, Don96b, Don96c, Don97b, Don97c, Don98b, DFC^{+}23, Dul06, Dul14, Dum05, Dun00, Dup05, Eck11, EEMS03, EDK00, Elb01, Emm11, Epp00b, EM00, FAB^{+}07, Fel96a, Fel96b, FFPC22, FWS^{+}03, FSW02, Fis02, FM02, Fli03, FT09, FES^{+}99, FH96, Gab07, Gag01, Gag06, Gag07, GJJ05, Gan03, GA04, Gan08, Gar99, GZXZ23, Gat01, Gay99, Gay00b, GD18, G^{+}06a, GVK^{+}03, G^{+}06b, GJ08, Gig06, GSHL03, Gla99, Goe00, Gol95, Gra06a, Gra04, Gra05, Gra06b, Gre19, GLS03, GNDB16, Gue19, GMTW08, GGZ02, Hah09, HMK^{+}20, Hal05, HS99, HP01, HHJC99, Hal06, Hal07, HBB13, HST08, HLK01, Haw00, Haw02, Hek97, HH04, Her99]. Linux [uRH95, uRH96, HLT^{+}21, HKS^{+}04, Hol02, Hsi01, Hua06, Hub02, HH05, HH08, Hug98, Hun99, Hun00a, Hun00c, Hun00b, Hun01, Hun02a, HP^{+}95, IBM05, IBM09a, IBM09b, ISO05, ILG19, IF10, JKS^{+}15, JHT^{+}07, JGS^{+}08, Jan03, Jan04, Jan06a, Jan06b, JW99, JPS00, JT98, JT05, JB99, Jon05, Jud17, Kal16, Kas14, Ker10, Kha02, Kie99, KLK^{+}02, Kir95a, Kir95b, KD00, KCTT06, Kom96, KC98, Kom00, KC00, Kom02, KAS^{+}06, Kop05, Kor06, Kos00, KH07, Kro99, Ksh04, Cor05a, LH01b, Lac06, Lat02, LW03, LC06, gL02, Lei99, Len97, LH00, LZAH23, Lin02a, Lip99, LWL05, LWL06, Lom01, Lov05, Lov07, Lov10, Loz05b, LADADL14, Luc00, Luc05, LCJH05, Mag00, Mag04, Mah07, MAT08, MSSvK08, Man01a, MM00, Mar04, Mar01a, Mar06, MAP^{+}22, MS^{+}00, Mau08, Max00, Max01, May01, MMC06]. Linux [MKSB22, McC99a, McC99b, McC02, McC03, McC04, McC05, McC01, MM06b, MW08, MFBWW20, MWM00, MSW06, MZG14, MM01b, Mil01, MYH00, MY03, Min01, MOS01, Moh98, MK04, Mol14, Moo01b, Moo01a, Mor97, MB05, Mus02, NR03, NW03, Neg06a, Neg06c, Neg06b, Neg07a, Neg07b, NSH02, NSH06, NSHW10, Neu06, NV06, Nis01c, Nis03, Nor99b, Nor02, Nut01, O’G01, Olc01, Oom13, OT07, OBSB23, Oua97, PLM06, PLHM08, PBM^{+}22, PTS^{+}11, PTS^{+}14, PDG02, PS22, PS07, PNK02, Par00, P^{+}08, PQM^{+}21, Pet98b, Pet98a, Pet00b, Pet00a, Pet01b, Pet01a, Pet02a, Pet02b, PB05, Pet00c, PWN01, Pfa00, Pfa01, Phi00, PC01, Pon05b, Pri99, P^{+}06, PF96, Pub11, PR96b, Pur04, Qui00, QH01, Ran06, Ras07, Rau01, Ray01a, RRR09, RR15, Ray98, Ray99c, Ray99b, Ray01b]. Linux [RMF19, RLB08, Reg11, RP02, RGRG05, RBHB99, Rob04, Rob95, RFS05, RFS06, Roe02, Ron05b, Ron05a, Ros98, RšT06, RS14, Rub98, RC01, SGP^{+}23, SK10a, SH13, SBR01, San98, SS05, Sch05, Sch07, Sch08, SG09, SJP^{+}09, SG04, SH99a, Ser00, SK01, Sha00b, Sha01, SGWP02, SK04, She02, She03, She06, SWF^{+}22, ST03, SAN08, Sho12, Shu06, SH99b, S^{+}00, SWL^{+}01, SFW03, Sie05, S^{+}05, SHB23, Slo05, Smi99, Smi00a, Smi00b, Smi01, Smi02b, Smi02a, Smi03b, SD05a, SD05b, Smi05a, Smi05b, SHPP07, Smi12, Sob97, Sob06, SY00, Sor01, Spe00, Spr06, Spr07, SS01, SS02, Sta04, Sta05a, Sta05b, Sta06, Sta07a, Ste01a, Ste02b, Ste01c, Ste01f, Sto00, SU94b, SU95, SUM95, SU96, SE96, SE97, Stu01, Stu04, Sul02, Swa00b, TR99, TGB95, TG96, Tan00]. Linux [TSC02, Thi04, Tho06a, TTB09, TCM07, Tor99, Tot99, Tox07, Tra96, TECB03, TF04, TB05, Tur05, Tyl06, VD01, Vad02, VTS^{+}23, Ven08, VMKB05, VRJ95, VRJ96, VJR96, VRJ97, VRFJ97b, VRFJ97c, VRFJ97a, VR00, VJ06, Vre03, WWW99, Wal00, Wal01a, Wal95d, WKDD03, Wal01c, Wan02, Wan11, War01, War04, War15, Way00, WDJ04, Weh01, Wel99b, WT99, Wt95a, W^{+}95, Wt95b, WK95, WK96, WDK99, WLS^{+}01, WKDD02, WB06, Wil98, Wis03, WRA05, Woe03, Wol04a, WTKW08, WBS01, WXX08, Wün13, WM00, XZY^{+}10, Yag03, YL09, Yeo05, Ygg94, Lin94, YSC^{+}06, Zad01, ZB19, ZST^{+}07, ZB01, Zie00, ZC02, dBB08, dOCC23, vH02, vH06, vH07, Hul05, Ano95b]. Linux- [Her99]. Linux-Based [OT07, CM06, MKSB22, PBM^{+}22]. Linux-basierte [Ric14]. Linux-Capable [OBSB23]. Linux-Cluster [GVK^{+}03]. Linux-compatible [AAB^{+}05]. Linux-Hosted [Coo98]. Linux-Ready [PS22, ZB19]. Linux-Tasks [RB06a]. Linux-USB [Reg11]. Linux/FreeBSD [Ano00e]. Linux/GNU/X [Ygg94]. Linux/ia64 [EM00]. Linux/Unix [Gab07, Sor01, Ano03a]. LinuxONE [BBB^{+}20]. LinuxWorld [Bar00b]. Linuxzeitung [Lin02b, Lin02c]. Lions’ [Lio96]. LISA [USE93b, USE94d, USE96a, USE97b, USE98d, USE00b, USE01a, USE91a, USE92a]. Lisp [Lew88, Ste90a, ACM92, BGPS91, Ell80, Ama91, Hen92, Kem91, Mac91, Mac92, NSI^{+}92, Ric90]. LispWorks [Ric90]. List [Loi79, McI82, Pog17, Pug88a, Rob75, Spu92, Tim93, TTB09, Wat83, Wri93]. Listing [ACS79, Hoo79]. Listings [Cro78c]. Listproc [Sch98]. Lists [FA88, Koe78c, Kra88a, Sch98, McN00, Wei97]. LISTSERV [Sch98]. Lite [Com94b, PM95, Ano98a]. Literate [BKM86]. Live [DeV93, Neg06c, Neg07b, Lum13]. Living [Raj00]. LKOUT [Whi75a]. LMOS [Ima79]. Load [HBD97, Höf01, HLLP79, HCG^{+}82, Kro99, Mon00a, Zuc83a, BGW93, Jal92]. Loading [Gra98a]. Loathing [MAT85b]. Local [Coh79, HCG^{+}82, IEE83, RB82, Wam83a, Wam83b, MHO^{+}76]. locales [LK00]. Location [USE93e]. Location-Independent [USE93e]. Lock [Ano93u, SW88]. LOCK/ix [SW88]. lockdown [Chi02]. Locking [Bor98, CCVLW23, Pug90a]. locks [MT93]. Locus [BP84]. Log [Fle02, Gla02, RO92, SBMS93, Gat01, JHT^{+}07, KAS^{+}06]. Log-Structured [RO92, SBMS93, Gat01, JHT^{+}07, KAS^{+}06]. Logger [Per77]. Logging [BE00, Bra77, Bra78, Gra01a]. Logical [Kat82a, Kat82b, HRS^{+}10, McS79]. Login [Hau94, Hil93]. Logo [Har83]. Logon [Wat88]. London [USE88c]. Long [DeV93, Kor97a, Too11, USE92a, Ano93v, CJBT92]. long-term [CJBT92]. Look [Ano96d, Neu06, Rag90, Usa78, Hol82, Kas14, Var96b, Mil90b, Mil90c, Mil90d, YP92, Joh92]. look-alike [Hol82]. Lookalike [All83a, All83b]. Looking [ASW^{+}01, Don98a, MKS^{+}90, Var96b]. looks [Gil96a]. looming [Ano95e]. loop [MR67]. loophole [Ree88]. Loopholes [Ano92k]. loose [Sta05b]. Loosely [BFS89, AGP^{+}92]. Loosely-Coupled [BFS89]. Loosing [Ban82]. Losing [Sta05a]. lossless [BW94c]. Lost [Bro20, Coc01a]. Lösungen [Bra04, Die11]. lot [Tay96, Wal99]. Lotus [RMZ01]. Louisiana [ACM97, USE95b, USE00b]. love [Coo03]. Low [Fed77, Liu96, Rob95, Sel94, War84, Ano03e, Gil96a, GK97, VTS^{+}23, WT91]. Low-Cost [Liu96, Ano03e, GK97]. low-end [Gil96a]. lower [Tay96]. Loyal [Ded09, Gla99]. Lp [Gli90]. LPI [Dea01, P^{+}06]. LR [AJ74, JP17]. LSI [BSW78, Lyc76a, Lyc77a, Lyc77f]. LSI-11 [BSW78]. LSI-UNIX [Lyc76a, Lyc77a, Lyc77f]. LST [PF96]. LST-distribution [PF96]. LSX [Lyc77b, Lyc77d, Stu79]. Ltd [Ano00e]. Ltd. [Ree82a, Ree82b]. Lucasfilm [HL85a, HL85b, Ree82a, Ree82b]. Lucasfilms [Law83a, Law83b]. Lucke [Tho05]. Luxe [Luc03b]. LVDS [Ano02c]. LZgrep [NT05].

M [Ano11, Ral77, KAH83]. M-series [KAH83]. M. [Mah89]. M4 [KR77]. M6 [Hal72]. MA [Fis73, Ano86, Ano94m, Ano95h, USE82, USE87b]. Mac [Gla97, SML06, Ang77, BSW78, Kir77, RS78, Rov77, Rov78, AC11, And06, Bal05, BS02, Buc01, CC07, DH05, Dav06, EBS08, Fei01, Fei02, GL06, Gem05, Ger02, GM11, Har01c, Hil02, Hil08, JR02, JR04, JR05, JRR08, KHG10, KS08, Lan01, MGMMI10, McE05, MTS03, MD09, MNC09, PM18, PB09, Pog02a, Pog02b, Pog05, Pog07, RR02a, RR03a, Ros06, RJR08, Sin07, SS06, TPTS02, TJ03, TJ04, Tay05, Tho06b, Top02a, Top02b, Top05, TSM^{+}06, TECB03, ZL10]. MAC-8 [Ang77, BSW78, Kir77, Rov77, Rov78]. Mach [USE93d, ABG^{+}86, Tev87, TBG^{+}87, TTG^{+}87, ABB^{+}86a, ABB^{+}86b, AKST93, Bas91, BCC^{+}91, BKLL93, CR92, Jal92, Mor96, SJ95a, SJ95b]. Mach-Based [Mor96]. Mach-US [SJ95a, SJ95b]. Machine [Car83a, Car86a, Cox09, GG99, KAH83, LA90, Roe02, Wit79, d’A92, BBD23, CK06a, Daw82, KNI88]. Machinery [WSS77, WSS78]. Machines [Bis88, Duf82, KAH83, Hib96, Len97, MHO^{+}76, Var96a, WJ82]. Macintosh [Bra93, KMV^{+}09, Mat97, Mor97, SG95]. MacOS [May01, May01, MCM06, MCM07]. Macro [Ben98, Daw86, Hal72, KR77, KV90, McI60, Env97a, Jen75]. Macro-11 [Jen75]. macrobenthic [DD05]. Macrocommands [Stu78c]. Macromedia [Ano02b]. Macroprocessor [RRW84, Lev98]. Macros [CL90b, DS77, Les78, Lio79d, MS77, Pis77, Ros77b, SM77b, Env97a, Les76b]. Macs [Lei95a, Liu96, Sto95]. MACSYMA [FG80, Fat82, Mat80, GF80, Sym85, SM85]. made [Gra05, Mus96, Mus02]. Madison [MN69, Ric77]. Magazine [AIX91]. Magic [GC94a, GC94b, Haz99, SZ95, Hin88, SZ95]. Magnetic [Rob75, Swa77]. Mail [Blu00b, Far91b, FA89, FA94, Hil93, HH86, IvW87, Luc03b, O’D83b, PC90, SEZ^{+}01, Bur93, Cam91]. Mailing [Sch98, TTB09]. main [Var96a]. Maine [MM06a]. Mainframe [Ano96e, Bra93, HP89a, HP89b, Lyo18a, Lyo18b, Lyo18c, Ric94a, VB92, VB94, Hid04, EEMS03, Sim95]. Mainframe-Killing [Ano96e]. mainframer [Hor98]. mainframes [Ano97b, Ger92, Gil96a, Wil94]. mainstream [BBHL08]. Maintainability [YSC^{+}06]. maintainer [BJM^{+}22]. Maintaining [AS88, AS91, Fel78, Boo03, Fel8 b, Fel79c, McK85b]. Maintenance [Bry83a, Bry83b, CKDS90, Cur91, FES^{+}99]. MainWin [Fra99]. Mainz [Uni97]. Majordomo [Sch98]. Make [Ano93d, Bra79, Cme86, Fel8 b, Fel79c, FES^{+}99, Lin93, McC95, Mec05, OT91, Wal94a, Ano95d, DJM86, Len97, Bee90, TNY92a, TNY92b, Fel78, Fow85, Hum90c, TNY94a, TNY94b]. Make-Believe [Wal94a]. Makefiles [Che78]. Makes [Ano93-42, RH95]. Makeup [KV89]. Making [Ano82, BDD^{+}93, Gem05, NH99, NH00, Pac00, TvK83, Wei82, YLM^{+}05, Hus92, Kuh91, Mor78a, Ano93p]. Malicious [SEZ^{+}01, L^{+}11, SEZ^{+}01]. malloc [Pug90b]. malware [BBD23, Rau01, L^{+}11]. Man [Dat88, SH85, PB21, van83]. Manage [Ano93l, BFS07]. Management [Ano93-40, Ano94k, Ano02c, BMF15, BB93, Bre86, CKM85, Cho98, CS86, Cou13, DFG99, ELS88, Har94b, HP89a, HP89b, Hof97, How01, HK89, HH91, Jan06b, KM86c, Kay79, Kiv84, LJ84, Lim05, MF83, MCH^{+}99, McG97, Mil90a, Rod90a, Rod90b, Rog79, SM88a, SM88b, SM88c, Ste02a, Sti83, Stu78a, Tes86, Tin79a, Tra85, Win02b, X/O88a, X/O88d, YSF89, dSVdLT02, Cus04, GW98, Hor02, JL96, KPG93, LR90, Lev02, LZAH23, MBC^{+}86, Mad97, Max00, SS95b, SS00, UNI93, VD01, Van05, WCHS93]. Manager [Ano00e, Bey77, Com94e, Jac84a, Jac84b, Lun96, Sou94, Wil82a, Wil82b, Hek89, SAL93]. managers [Bam98, Gil96a, Var96b]. Managing [And91, Apo87, Cam04, CB89, Dun91b, EST86, KJHL96, Kop00, Mec05, OT88, OT91, Per87, Sch98, Ste91, TO92, Vad02, Woz82, RHP^{+}07, Sav11]. manifesto [Gra84]. manipulate [Yos85]. Manipulation [Car77, FGP64, Hoa77, Les88b, Lin79b, GIM07, Jou86]. Manipulator [DBG89]. Manual [?, ?, ?, ?, AKW85, Ang77, Bel79a, Bel79b, Bel79d, Bel79c, BK91a, BK91c, BJK92, BG77, Bro71, Bro73, Bro77a, Bro90, BR77, Com94c, Com94e, Com94f, Dep79b, Dep79a, DF00, ES90, FB93, Gas92b, Gas92a, Gra94, HS84, HS87a, HS91a, HS95, HS02, Hel90a, Hel90b, Hel91a, Hel91b, Hobxxb, Hof77, Hol87e, Hum90a, Ker79e, Ker79f, Ker79g, Ker84a, Ker91, Kir77, Kos92, McI72, McI87, Mey18, Mil00, Mul79a, MD86, NM83a, NM83b, Ng77, Nye88a, Nye88b, Nye90b, Nye90a, Nye90c, NO90a, NO90b, Nye92b, Nye92a, NO92, NO93, O’R90b, Olc92, Ope93, Ope94, Oss74, Oss76, Oss79, OK90, OK92, Pea77, PW94, Pis77, Reg78, Rit2 , Rit74, Rit77a, Rit79a, Sch89, Sta93, Str84a, Tal92, Tho69, TR71, TR72, TR73a, TR73b, TR74, TR75a, TR75b]. Manual [TRS18, Tji85, UNI77, Van91, Van81, Vog78a, Com84, CCA84, Cra90, FG80, GIM07, Geh84, Ham90, KVCS98, KFO97, Lew88, Mac91, Mac92, Mat80, Mat85a, MS87, Mil98, Olc01, Oua92, PS^{+}09, Pog02b, Pog05, Pog07, SM85, Tom11, Uni85, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, All86a, Wid90]. Manuals [Dzo09]. Manufacturing [Dix82, VMKB05]. Manuscripts [AS88, AS91]. many [TH83a, TH83b]. Map [SM79, Sta04, Isa86]. mapped [TTG^{+}87]. Mapping [SMD^{+}94]. March [IEE03, Ric77, USE91e, Uni97]. Maria [Hid04]. Mariani [Sau88]. marine [DD05]. Market [Den83, FKN85, Jac91, Gil96a, Gil96b, Tuc90, Gre93]. marketing [Ano93r]. Marketplace [Tan87c]. Markets [Ela83, Kat83]. Markov [DLCV17]. markup [YLM^{+}05]. Marriott [USE01b]. Marsh [Kar84]. Maryland [ITC93, USE89e]. Mass [Coc01b, HP89a, HP89b]. Massachusetts [ACM82, USE93e, USE94e, USE01b]. masses [Gra84, Ano92s]. Massively [Bal94, EGL86, KmWH10, ZRB^{+}93, CS96a]. Master [Bel00, Gri14]. Mastering [Dan01, DJ01, Fio04, Fri97, Fri02, Fri06, Gri14, HA90, HA93, Jae89a, KHP05, McM92, McM93, Mic03, Mic08, TNY92a, TNY94a, TNY92b, TNY94b, Wan11, WM01, Zie16, WB06]. Mastery [JL16, Luc12, LJ15, Luc18b, Luc19, Luc13b, Luc16]. match [Rob8 , Rob79b]. Matcher [Sab76b]. Matches [Car82]. Matching [AC75, BJ01, Cox07, Cox09, Cox10, Cox12, IK83, NT05]. Math [Ame09, SST90]. Mathematical [FGK89, FKM95, FHS77, FHS78b, Mor75, Tho03, IBM05, Ric77]. Mathematics [Har11, KC75, KC76, KC78b, KC79, KC90, Ric77, KC78a, uRH95, uRH96]. Mathematik [uRH95, uRH96]. Mathlab [FG80]. Mathlab/MIT [FG80, Mat80]. MATLAB [The92c, The92a, The92b, The93]. Matrix [Kau7 , Kau79a, Lon77]. Matt [Ano95b]. Matter [Ano93-43]. matters [Tun05]. Matthaeu [Hid04]. Maturation [Fun87]. Mature [Bra04]. maximize [Gri14]. Maximum [Coc01b, HKT05, Ray01a, RR03a, TR99]. MAXIUNIX [Hal79]. May [ACM75, Bar01b, Eur91, IEE91a, IEE94a, IEE94c, Ass88c, USE89g, USE92c, USE99c, Ano93v, Sim96c, Var96b, MK96]. maybe [Tay96]. Maze [Sto89, Sto90]. MBS [Shi77]. McIlroy [Mah89]. McKusick [JMR19]. MCM [KPT^{+}02]. mcs [CL90b]. MD [IEE02, NIS92, USE89a, USE92e]. MDB [MMG07]. MDS [LT76, LTW77]. Me [Bey87, Fro90, All86a, All86b, Daw86]. Mean [Stu78b]. Means [Lin95]. Measured [CEC^{+}96]. Measurement [Lyc71, DI89]. measurement-based [DI89]. Measurements [ADJS78, DD90, Hay77, Hoo79, PB91, PR76, Ral80, STA87]. Measuring [BG88, Hu86, LKM84, IEE91c]. Mechanism [Eyk88, Lo79, MSSvK08, Mor78c, GMTW08, Mly09]. Mechanismen [Bas87, Bas87]. Mechanisms [Ano94g, Bas87, Kau78, Tin78d, Wat88, DC18, Don85, IBP03, PDG^{+}87]. Mecum [Lip99]. Media [Ano93q, JP09]. Medical [Ham87, SML06]. Meeting [Ede77, Eis79b, Hos83]. Meets [Lel90, Nis01c, Spi00, CG93a, DO88]. MEF [SEZ^{+}01]. Members [Pin99]. Membership [Coc01a]. Memoir [Ker19]. Memoranda [Cro78c]. Memorandum [MS77]. Memoriam [Gre11]. Memories [AU3 , AU74]. Memory [BJP04, BS90, CHRG95, Cla88, De 76, Fra92, GSHL03, HP89c, KM86a, KLK^{+}02, KPSS79, LJ84, LBN14, Les75b, Lyc71, Lyc73, MK85a, McG79, MK88, MKB90a, Mil90a, MF99, MZA96, RRP^{+}08, SMK^{+}94, SM88a, SM88b, SM88c, AKCB86, Ber98, CDD^{+}07, CR92, GZXZ23, GNDB16, GMTW08, JL96, PPTHxx, RRP^{+}07, RHP^{+}07, RRH^{+}08, Sem96, TTG^{+}87, Var96a, WRA05, GK97]. Memory-Based [MKB90a]. Menu [Bla83a, Bla83b, Ney83a, Ney83b, War88, Pal87, PDG^{+}88]. Menu-Driven [Bla83a, Bla83b, Ney83a, Ney83b]. MENUNIX [Per82b]. menus [Env97a]. Merged [Coc03a]. Mergenthaler [CKT80]. Merging [HAB^{+}98, Pik82]. MERT [BL75, Bur78, Fed78b, Ham78, Hay77, HR78b, H^{+}79, Kap79, Lau77, Lau79b, Lic79a, Lic79b, Loi77, Lud77b, Lyc77e, Lyc77g, LB78a, LB78b, Lyc78b, LB78c, NP78, Ral78, Sha79b, SR77, Tin79a]. MERT-UNIX [SR77]. MERT/UNIX [Fed78b, Lic79a, Lic79b, Lyc78b]. Message [Hof79, PY84a, PY84b, Ral80, Ber98, HK05, PS07]. Messages [Bra76a]. Messaging [DN87]. Meta [SGM^{+}00]. Meta-data [SGM^{+}00]. MetaCreation [Kro99]. metadata [GMSP00]. MetaPost [Hobxxa, Hobxxb]. MetaTM [RRP^{+}07, RRP^{+}08, RRH^{+}08]. MetaTM/TxLinux [RRP^{+}07, RRP^{+}08]. Mether [MF99]. Method [CW89, CW91, McI63, Ric85, Sch76b, Yos85, Env97b, MHS02, Rob8 , Rob79b]. Methode [Bra98]. methodologies [DDLL94, Sun01]. Methodology [McIxx, Sne01, STA87, Stu78a, FFPC22, Tob80, Wil98]. Methods [Fla90, GES98, RAMP94, SW99, Cox83, IEE91c, Pet88]. Metro [Ano00e]. Metro-X [Ano00e]. Metrowerks [Ano02c, Ano03c, Ano03d, Ano03e, Kro99]. Mexican [Bar01b]. Mexico [ACM75, KDP83, USE93d]. mezzi [ZB01]. MGR [Uhl87]. MH [Joh7x]. MH-TSS [Joh7x]. MHz [FJPR93]. MI [USE92c]. MICA [Ng77]. Mice [Paj88, Pik93]. Michael [Ano95d]. MICRO [IEE04, GLOR91, IJ98, Lyc77h, Mat97, RS78, TT79, May84]. MICRO-37 [IEE04]. Micro-DBMS [TT79]. Micro-Kernel [GLOR91]. Micro-Processor [Lyc77h]. Microarchitecture [Bar00b, IEE04]. Microassembler [Ng77]. Microbytes [Ano93r]. Microcomputer [BCGS91, PY84a, PY84b, WSS77, WSS78]. Microcomputer-Oriented [BCGS91]. Microcomputers [Jun85a, Jun85b, Per87, Rob84a, Sti83]. Microkernel [Mor97]. Microprocessor [CCH^{+}77, Far78, Gwo78, HHLS90, Lyc78c, Lyc78d, Lyc78e, Lyc85, Man92, Mil90a, Mol77, Rub79, Weh77a, Weh77b, BY96, Gra87]. Microprocessors [BSW78, Mil79]. Microprogram [FR75]. microprogramming [RW81]. Micros [ACF^{+}86, Mei84]. Microscope [Ano03f]. Microsoft [PKP05, Ano95d, Ano96d, Gra03, Len97, Li98, Lin93, MRM06, OH97, Paj88, PH98, PH99, PKP05, SAS94b, SR95, Wid00, WL94]. Microsupercomputer [MBB^{+}91]. middleware [AJ05]. MIDI [Haw86, Haw87]. Migrating [IBM09b, Ker93, PDL04, TLK99]. Migration [BDD^{+}06, Cha06, Fre91, LS92, MS88, MK93, CDV^{+}94, HK82b, MK92a, MK92b, Raj98, Sim96b, WG98]. Migrations [PDL04]. Milk [OS06]. Millennium [HH97b]. Million [Kul06, WWCW19]. million-user [WWCW19]. Minding [BYVF08]. Mini [Lyc73, Lyc74b, Lyc76b, Lyc76c, Lyc76d, Lyc77c]. Mini-Computer [Lyc73, Lyc74b]. MINI-UNIX [Lyc76b, Lyc76c, Lyc76d, Lyc77c]. Minicomputer [JH83, Lef81, LC78b, LC78c, Sha81, Swi79]. minicomputers [Ger92]. Minimal [Mah07, WM00, BMR94]. Minimalist [HP89c]. Minimum [ALxx]. MiniRTL [WM00]. MINITAPE [Swi77]. MINIX [GLG93, HBG^{+}06a, HBG^{+}06b, Tan87a, Wai95, AEG^{+}91, CG93a, KPG93, Wil98]. Minneapolis [IEE83]. Minnesota [IEE83]. Minutes [Nel79, Ray99a, RBHB99]. MIPS [Ame91, PY84a, PY84b, Swe07]. Miranda [Tur86, Tur87]. MISRA [BBH18]. misses [ST03]. Missing [Loz05a, Pog02b, Pog05, Pog07, Sim97]. Misunderstood [DM97]. mix [Sim96b, AGP^{+}92]. Mixed [Bec84, Koh06, Bla98, MB05]. Mixed-Language [Bec84]. mixture [HMK^{+}20]. mixture-based [HMK^{+}20]. Mk [Hum90c]. MKCONF [HR77]. MkLinux [Mor97]. MKS [Mor89]. ML [Jou86, Mat85a, MT93]. MLFQ [TCM07]. MLS [HRS^{+}10]. MM [MS77, Ros77b, SM77a, SM77b, SP77, Smi78, Wen79b]. Mo [Coc01b]. Mobile [Ano03e, Ano16a, BDD^{+}93, Jon02, SFE10, Sta06, USE93e, Ano03d, Cha11, LZHS11, WWCW19]. mobile-health [LZHS11]. mobile-platform [Ano03d]. mobility [KHG10]. Mode [Dik06, GT78b, LS83, Pas88, Gue87]. Model [Bro0 , Bro81a, Bro81b, CPJ^{+}98, Coc01b, FG91, FA88, Gra01b, HKT05, Hay06, KMKO92, KM82, NR03, Ram88, SK72, SG09, Bro6 , Bro77b, DLCV17, HKVR21, Sae03]. Model-Based [CPJ^{+}98]. Modeling [AKST93, Lut99, RWNA87, WCHS93, PH87]. Modelling [TTB09]. Models [MZA96, HMK^{+}20, dJPDD84]. Modem [Far91a]. Modems [Tho92]. Modern [Cer98, Cer03, Fel79b, Fin91, Jae89a, Rob03, Sch94b, Sch95b, Sou92, Tan01, Tan08, Tan10, Gue19, Hav11, Far92b]. Modest [Hos84]. Modification [Ral78]. Modifications [Bin88, Coh79, Dea78]. modified [GO86]. mods [Kra07]. Modula [Pow84a, Pow84b, Woo94]. Modula-2 [Pow84a, Pow84b, Woo94]. Modular [BBM85, Boy84, HBG^{+}06b, SO92, Hon08]. Modularity [NR03, BTS09]. Module [Cam04, PNK02, Ano95d, KPT^{+}02]. Modules [Zuc83a, BTS09, GJJ05, LC06]. Molecular [Vol94]. Monatsblatt [Lin02b, Lin02c]. mondo [Cor05b]. Mondrian [WRA05]. Mondrix [WRA05]. Monetary [Gar97]. money [Ale96b]. Monitor [BFGK89, Bon79b, Mac97a, MMS86, Mue78b, HK82b]. Monitoring [BM87, Dil02, Ell79b, GS00a, JZTB98, LPG^{+}18, SW91a, CJBT92, KIT93]. Monitors [Gla02, Dun91c]. Monk [MK84a, MK84b, MK85b, MK90]. Monotonic [HG93]. Monterey [Ano93-31, USE86d, USE88d, USE89h, USE93b]. Monthly [Cha85, Lin02b, Lin02c]. Months [VB92]. Moon [Coc01a]. Moore [Bar01b, NT05]. Morlaix [DD05]. Morris [Gre11]. Mortem [KB01]. MOS [BP86, McG79]. Mosaic [RR95b, DKF94, VSW94]. Moses [Ano94d]. MOSIX [BW89, BGW93]. Moskowitz [Koh06]. Most [MG99, CK06b, Hub96, Pub11, Vib11, Spe96, Wil11]. Motel [Hum88a, Hum90a, Mic79b]. Motif [Bar91a, Bea93, Ber91, Bra92, FB93, Fla94, Gre92b, Hel91a, Joh92, Kel90, Kob91, McM92, McM93, Mio97, New92, NO90b, NO92, Ope91a, Ope91b, Par93, QO91, QO93b, Ros90, Ros93b, Seb94, She93c, Smi92, Sou94, Yag91, You90a, You90b, You92a, You95]. MotionBuilder [Ano02b]. Motivation [Ryd72]. Motorola [Ano02c]. Mount [Ano02b]. Mounts [PM95]. Mouse [Paj88]. Move [Bar00b, BB99, Gil96b]. movement [Way00]. Moves [Ano90a, RH95, Gil96b]. Moving [Anoxxb, Dav93b, Gag06, Gag07, Ste01b, Tin78c]. MP [Rit88b, Rit88a]. mp3 [ZB01]. MPE [Ano91b, Ano93m, Lal93]. MPE/iX [Lal93]. MPI [MK04, RšT06, Slo05, YL09]. MPI-2 [MK04]. MPICH [RšT06]. mpm [KV90]. MRCS [Gel78]. ms [GL88, FM94, Gia90, Hol87c, Kin89, Mik89, PH86, PLBL92, Pug90c, SB96, URK85, War88, CL90b, Les76b, Les78]. MS-DOS [FM94, Gia90, Hol87c, Kin89, Mik89, PLBL92, Pug90c, SB96, URK85, Ver89, War88]. MSG [Mer82]. MSS [HP89a, HP89b]. MSS-II [HP89a, HP89b]. MTA [Den04]. MTEX [TLK99]. Multi [BL75, BEHW86a, BEHW86b, Fed88a, Fed88b, Fra79, HIT^{+}02, Lyc71, Lyc72, MKSB22, MS92, McK03, Raj00, SGP^{+}23, Sch79a, ST89a, ST89b, ST89c, Tes86, XZY^{+}10, dSVdLT02, FKN85, Ano96e]. Multi-agent [dSVdLT02]. Multi-Core [SGP^{+}23, XZY^{+}10]. Multi-CPU [FKN85, Ano96e]. Multi-Environment [BL75]. Multi-Gigabit [HIT^{+}02]. Multi-Level [ST89a, ST89b, ST89c].

Multi-Operating [Raj00].

Multi-Processor [Sch79a, Tes86, Lyc72]. Multi-programming [Lyc71]. Multi-Representation [BEHW86a, BEHW86b]. Multi-Routing [Fed88a, Fed88b]. Multi-structured [MS92]. Multi-tenant [MKSB22]. Multi-Terabyte [McK03]. Multi-Thread [Fra79]. Multibug [CP86]. Multicast [Hug88]. multichip [KPT^{+}02]. Multicode [Mud97, Mud97]. Multicomputer [CGS96, Roy93]. Multicomputers [ZRB^{+}93]. Multicore [PGW^{+}20]. multidatabase [BCE^{+}93]. Multikey [Zuc77a]. Multilevel [MR92, MRxxc, MRxxd, RM92, RMxx, d’A92, dC93]. Multilevel-Secure [RM92, RMxx]. Multilingual [KMKO92, Mud97]. Multimedia [Bv92, KLK^{+}02, LKY93, Law98, Mei92, MH96, NHNW94, PNK02, Tra96, WLL96, YLK93, Ano03e, Ran06]. multiplatform [Bam98]. Multiple [HRO82, Kat99, Kle86, KM87, Lib87, Pat83, Com00, Dit93, Hoo05, SH13, Whe97]. multiple-precision-arithmetic [Dit93]. Multiple-Process [HRO82]. multiplexed [Pik84a]. multiplexing [Sch93b, Sch93c]. Multiplexor [Pik84b]. Multiprocessing [Cus92, Sch94b, Sch95b, Wal95e, HdBM75, MT93]. Multiprocessor [Bac84, BB84, BW89, BK85, BFS89, BS90, CS90a, CS90b, HCN85, LLM87, Mil86, PPTHxx, Prexx, PBL86, USE89c, USE91e, vMM88, vM88, CDV^{+}94, CR92, FGP^{+}91, GMTW08, JAV86, TGF88]. multiprocessors [Ber98, CHRG95, GGJ89]. Multiprogramming [JJ91b, JJ91c]. Multiprotocol [Bor98]. multispecies [MMTK94]. multitasking [JJ91b, JJ91c]. Multithreaded [Ano98d, Kor89, LB96b, LB98, MS89, Yoo96, MS87]. Multithreading [RR96]. Multiuser [Mil86]. Multivariate [HKF84]. MUNIX [HdBM75]. MuPAD [uRH95, uRH96, uRH95, uRH96]. Music [Haw86, Kil90, Phi00, Gem05]. Musings [Ray99c, Ray01b]. muster [Sim97]. Musterlöungen [PKP05]. mutual [Ger85]. mv [GL88]. mview [Stu78c]. MVS [Ano93u, Bam98, DuC94, Joh95, Joh96a, Joh96b, Koh99]. MX [FR75]. My [Bey87, Ker81c, WCC^{+}12, Pin92]. MySQL [BGW04, Gab07, G^{+}06b, GB01, LW03, MB00, Reh03, TSC02, WT01, Wün13, DuB01]. MySQL/PHP [BGW04, GB01]. Mysteries [Lib93]. Myths [Wal94a].

Nachschlagewerk [Kuo99]. NAECON [IEE91a]. Name [NR94, PW85, FH89]. Named [Roc78]. Names [Jon84]. namespace [WDG^{+}06]. Namespaces [Min01]. Naming [BH00, BH01, MS92]. Nan [RB08]. Nanotechnogy [Ano03f]. NAS [Har90]. Nashville [USE91c]. Nation [PP85]. National [Ame89, Hag83, IEE91a, NIS92, Ano02c, Ano19c, HG95b, Kod82a, Kod82b, NMS83, SJ83]. Native [Cov06, AHA^{+}14]. Natural [Wei88]. Navigator [Gil93b]. NC [Ano96d]. ncpfs [Len97]. NCurses [Goo07]. Near [Wol04b]. nearly [LD13]. Need [FKN85, Sem96, Var96a, Var96b]. needs [Lov07]. Negotiating [JLSG84]. nell’era [ZB01]. NEMO [GO86]. NEQN [Scr76]. Net [Kro99, Law98]. Net.BSD [Ano03b]. NetBSD [HK95, NC08, Ree06a, YSC^{+}06]. NETIX [Wam83a, Wam83b]. Netizens [HH97a]. Netlink [DeV00]. Netnews [Har95]. Netscape [Ket99]. NetView [CBH^{+}92]. NetView/6000 [CBH^{+}92]. NetVue [Ano98c]. NetVue/JAVA [Ano98c]. NetWare [Ano95e, Hur94, WF03, Ano93r, Dul96, Lun96, Raj98, Raj00, She93b]. Network [And91, ACF^{+}86, BM87, Bis88, Bla98, Bus02, BP84, CHL16, CK78a, CBV93, Dul94, Ell75, Ell79b, Fer85, FES^{+}99, FHR84, Gil93b, GZ84a, GZ84b, Har90, HIT^{+}02, Hol75, Hun92a, Hun02b, HCG^{+}82, IB93, Jan03, KM86c, Kir95a, Kir95b, Lau79a, LH01a, Lut99, Mac97a, Mar10, Mil84a, MTT83, Nad98, Nor99a, NZW^{+}02, NL79, PMD88, PC90, Reexx, RW92, RB82, Ser00, She93a, SW91a, SH95, Spu92, Ste90b, Ste98, Ste99b, SFR04, Stu78a, Wei84, YSF89, dSVdLT02, Bas91, BDP05, Ben06, Che75b, Chi02, GK97, Hun99, Hun02a, Kaz05, KD00, Lev02, Lun96, Lyo08, Man01a, MHO^{+}76, Max00, McN97, Mel87, NL82, Rag93, SH13, Smi05a, TR99, Tot99, ZG96, BBNS80, GK97, GRA89a, LBSL97, Rag90]. Network/extensible [GRA89a]. Networked [CCM^{+}90, KLK^{+}02, Pic83, Sef93, BV06]. networker [Mit01]. Networking [ACM04, Bee93a, CKM85, CF20, Cas93, Gor81, Har90, Hun92d, Gur81, KJH^{+}97, KW89, Lud79a, McK94a, Mul87, MZA96, ML02b, Ass99b, Pad93, Raj00, Req85, RW86b, SKP94, Ste98, SFR04, USE94b, Wam83a, Wam83b, X/O88b, Blo99, Cab83, Col02, DeR00, McK91, McK92a, McK92b, MK96, Pfa01, Sch07, Sch08, SG04, Smi00b, Smi02b]. Networks [Ale96a, BW94a, BW94b, Bra93, FA89, FA94, HH94, IEE83, MS88, PR85a, Tan88, AOAH14, PB21, Ano93l, Ano03d, NZW^{+}02]. Netzwerk [KGG00]. Netzwerkadministration [Kuo99]. Netzwerke [WF03]. Netzwerkkonfiguration [WF03]. Netzwerkspiele [Str94]. neural [Kaz05]. Nevada [ACM89]. Never [Ros95]. Newcastle [BMR82, BMR01]. Newnes [Hea98]. News [Ano93s, Ano94c, Bar00b, Bar00c, Bar01b, Coc01a, Coc01b, Coc01c, Coc03a, Coc03c, Coc03b, Gar97, Gar99, Gla99, Law98, Sta04, Sta05a, Sta05b, Sta06, Sta07a, GRA89a, Sal97]. Newsletter [Ass89c, Mat93]. Newsqueak [Pik93]. NEWSTRENDS [Ano96d]. Next [Per93, Thi04, Ros94b, Sim96c]. nfs [Len97, Cal00b, Fel94, Her93, Olk03, San91, San94, Ste91, Zad01]. NFTS [Roc78]. nicely [GO86]. NIDS [SK10a]. Night [Kor97a]. Nihongo [JK86, Sas91c]. Nine [VB92]. Ninth [Uni96a, Uni96b, PR85c, PR90, Rag90]. NIS [Ram94, Ste91]. NIST [Bar00b, Coc03b]. nix [Cap87, Cap82]. NixOS [DL08, DLP10]. nm [ZB19, Stu78a]. Nmap [Lyo08, Mar10]. NNTP [Ste96]. No [Ano96d, Ano16b, CPJ^{+}98, Hol12, Jen05, Kir77, McIxx, Pek78, Sim95, Sta05a, Che97, Han08, Han11, Par90, CPJ^{+}98, Dav96]. No-nonsense [Hol12, Han08, Han11]. NodeOS [SGWP02]. nodes [Len97]. Nøglen [Kri91]. Noise [dOCC23, ALL13]. Nomads [BDD^{+}93]. Non [AD89, GSHL03, McM78a, McM90, Pea83, Gra04, Gra06b, War85a]. Non-Dedicated [GSHL03]. non-geeks [Gra04, Gra06b]. Non-interactive [McM78a, McM90]. Non-Paged [Pea83]. Non-Unix [AD89, War85a]. Noncharacteristic [Cou84, Cou85a]. noninteractive [McM79]. Nonlinear [CW89, CW91, MMTK94]. nonsense [Han08, Han11, Hol12]. NonStop [Nor91]. Nonvolatile [LBN14]. Nooks [Gra97, Gra98e]. Noosphere [BV06]. Norm [Blu7 b, Blu78b]. normal [Wal96]. Norton [NH91, Nor99b]. Norway [Eur91]. Notebook [Dum05, Skj83a, Skj83b, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84f]. Notes [Ano96d, Maj77, The93, SAS91a, McK91, McK92a, McK92b, Swi91, Arn98]. Nothing [Sta04]. Notice [Erl88]. Noticeable [Kra83]. notion [Ale96a]. Novel [GJLB77]. Novell [WF03, Ano93r, NS94a, She93b]. novelty [Dia94]. November [ACM89, ACM04, Ano93-31, IEE02, McK91, MS05, USE86d, USE88d, USE89h, USE93b]. novice [Pec08, See08, Tho06a]. Nroff [Les78, All86b, Cum77, Oss76, Oss79, Rod87, Smi76, Les76b, Sun86]. NROFF/TROFF [Oss76, Oss79, Smi76, Rod87]. NS16032 [SJ83]. NSA [McC05]. Nspeak [Har79]. NT [Ano95e, HD98, Ale96b, Ano93-43, Ano95d, Ano96d, Ano97a, Ano98d, Arn98, Aud96, Bal96, Bia99, Cus92, Cus93, Dav96, DeR99, DeR00, Don97a, Don98a, Env97a, Env97b, Fos96, Fra99, Fre98, Gau98, Gil96a, Gil96b, GES98, GBG97, Har93, Hel97, HAB^{+}98, Jen00, Koh99, Kor97b, Kor97a, Kos96, Lei98, Lei00, Lin93, LBSL97, Mer97, MYH00, NH99, Nas97, O^{+}98, Ass99a, Pac00, Pon97c, Raj98, Raj00, RT93, SAS94b, SS95a, Sim96b, Sim96c, Sim97, TLK99, Tuc98, USE97a, USE97d, USE98c, USE98a, USE99a, Var96b, Wal97a, Wal98, WG98, WS96]. NT/2000 [MYH00]. NT/95 [Mer97]. NT/RISC [Var96b]. NT/UNIX [Nas97]. NTT [PS22]. NT^{T M} [Wal97b]. Nuclear [IEE91b]. Nudges [Ano91a, Bun91]. NUMA [CHRG95]. Number [Rob79a, Rob82, Don22, DJM86]. number-crunchers [DJM86]. Numbers [Mor78b]. Numeric [GR76, The92c, The92a, The92b, The93]. Numerical [ACM79, BFKS93, Blu75, Blu77, CW89, CW91, LWL05, Sch76a, Sch79b, EP05, LWL06, Sch79c]. nurtured [Rit85]. Nuts [Ano99c]. Nutshell [AL92b, CGO92, CT94, Gt92a, Gt92b, GT92c, G^{+}98a, G^{+}98b, Hek97, OD^{+}86, O’R90a, Rob99b, Rob99c, Rob05, Rob06a, Ano00a, Ano00g, Dea01, Hud06, KH07, MTS03, P^{+}06, SH99b, S^{+}00, SFW03, Sie05, S^{+}05, TSM^{+}06, Rob00, Sch04]. Nvidia [Ano03f]. NYU [EGL86]. nyumon [MO01].

o [JK86, Ano93e, Car82, Che87b, Fed77, KSU94, May00, McS79, Ric96, Rit5 , Rob06b, Rob84a, Sch93b, Sch93c, Tin78a, Tin78c, WMK^{+}07, Yoo96, dBB08, vMM88, vM88]. Oakland [USE94b]. Oberon [WF03]. Obfuscated [Lib93]. obfuscating [WWCW19]. Object [AQ84, BCC^{+}91, Dav97, Des96, Gra01b, HCF^{+}94, HK89, JGS^{+}08, KN93, NH99, NLR84a, NLR84b, Smi91, SJ95a, SJ95b, Str6 , Str88, You92a, And80, BY96, Cox83, Ewi86, LC93, Sch93b, Sch93c, SML06, SAL93, WCHS93, BJ01, Kro99, Str91b]. Object-based [NLR84a, NLR84b]. Object-Oriented [BCC^{+}91, Dav97, HK89, NH99, Str6 , Str88, JGS^{+}08, KN93, Smi91, You92a, Ewi86, Sch93b, Sch93c, SML06, Str91b]. Objective [Koc09]. Objective-C [Koc09]. Objects [NH00, Pac00]. objektorientierten [Dav97]. Observability [CF20, Gre19]. Observations [TBS87]. OCA [Ern11]. OCE [Sav11]. Octave [FY18]. October [ACM88, ACM03, Ano86, Ano93g, Ano93x, Ano95h, Ano06a, Fis73, IEE83, ITC93, Mil84b, NIS92, USE87b, USE88h, USE89c, USE90b, USE91a, USE91b, USE92a, USE93g, USE96a, USE97b, McK91, McK92b, McK96b, McK97a]. ODBC [Ano00e, Kie02]. ODBC-ODBC [Ano00e]. Odd [Rit02]. Ode [AGS90]. OdeView [AGS90]. Odyssey [McK85a]. OE [NBB^{+}02]. Off [Bal96, Ano91c, Ano93m, PH91]. off-the-shelf [Ano91c, PH91]. offer [Biv87]. Offerings [Isl83]. Office [Ano02b, Don98b, Hof79, JLSS77, Mil77, Ney83a, Ney83b, Ano03d, NW03, Neg06b, Smi00b, Tot99, van86, Gra03]. Official [NDW93, Ope93, Ope94, Sha94, TV04, Gra05, HBB^{+}06, H^{+}07a, H^{+}07b, Kee05, KFO97, Lyo08, PS^{+}09, Reh05, SAN08, WNDO97, W^{+}99]. Offs [Kri84a, Kri84b, De 78, PLR85]. OFS [Ama88]. Oh [Pin92, Swa00a]. Ohio [USE93c]. Oil [Fra90a]. oku [DF91, MO01]. OLC [CAG89a, CAG89b]. Old [Hos05, Ora21, Tho19b, Wis03]. OLTP [YR93]. Omnicore [Ano02c]. On-Line [CAG89a, CAG89b, NM83a, NM83b, Tal89a, Tal89b, Ano97b, CGMS89]. ONC [San94]. ONC/NFS [San94]. Once [LEG88]. One [Arn91, Bro90, FG79, HHLS90, Kruxxa, Kruxxb, Lib87, MZH22, Sch76a, Dul14, Sim96c, Var96b]. One-Liners [Kruxxa, Kruxxb]. Online [GW98, Gra98c, Les88b, Sta06, PDG^{+}87, Pal87, PDG^{+}88, SAS91a]. Only [Bar00b, ATS94, Che79a, Che97, Var96a]. Ontario [Sof83, USE94c]. OOoSwitch [Gra03]. Open [Adh95, Ano91a, Ano93-33, Ano94f, Ano95d, Lin02b, Lin02c, Bun91, Cas93, Cas94, Cha11, CPJ^{+}98, Coc01c, Coc03a, Dun91b, Eur91, Fog99, FT09, Fra19, GS99, Gla99, GC94a, GC94b, Hah94a, ILG19, JB99, KL94, LW03, LPG^{+}18, Lew99a, Lew99b, Lin02a, MSSvK08, Ope94, PGW^{+}20, QW93, RRN95, Ray99b, RR95b, RRH94, San98, SK04, Sor01, TTB09, X/O88c, X/O88a, X/O88d, X/O88b, X/O89b, X/O89c, X/O89a, ZST^{+}07, AC11, And01, Ano93m, Ano94i, AAB^{+}05, BV06, BP09, CM06, Cus04, DDLL94, Ger03, Goe07, HP01, JP09, LL93a, LL93b, MM00, McC05, Moo01b, PS^{+}09, Ray01b, Sal08, Siv90, Wis03, YSC^{+}06, ZK05, ZB01, Mil90b, Mil90c, Mil90d, YP92, Ano92i, Anoxxb, Lin02b, Joh92, Lin02a, Moo01a, NR03, Ray99c, Wal95a]. Open-Source [Coc03a, PGW^{+}20, Ray99b, And01, AAB^{+}05, YSC^{+}06]. OpenBSD [Ano00b, And17, Art03, Car04, DIN05, Han08, Han11, KHP05, Luc03a, Luc13a, Mat07b, NC08, PN04, Ree06a, Smi02a, SY00, YSC^{+}06, dR15]. OpenGL [Aki92, Ang97, Ang99, Ano94e, Dav93b, Fos96, GS99, Gla92, HB04, Jac99, Kil93, Kil94, Kil96, KS08, MB03, NDW93, SA93a, SA93b, Wal95b, W^{+}99, WS96]. Opening [Ano93-34, Anoxxb]. OpenLinux [Cal00a, SRBB99]. openMosix [Slo05]. OpenMP [CDK^{+}01]. Openness [HR94, RH95]. OPENNT [Wal98]. OPENNT^{T M} [Wal97b]. OpenOffice.org [Ano03d, Gra03]. OpenSolaris [FT09, Hay06, MM07, MMG07, SJM09]. OpenSSH [DLvG07, Luc12, Luc18b].

OpenSSL [WN15]. OpenSUSE [CK06d, CK06c, CK06h, CK06i, CK06i]. OpenVMS [BHM98, DuC94, Har93, SAS94b, She93c, Wis03]. OpenWindows [KLM92, Sun93c, Sun93d]. Operating [ACM03, Ame93, Ame94, AISS98, Ano88a, Ano98b, Ano01, BJP04, Bac86a, Bac86b, Bad98, Bar88a, Bar88b, Bas87, BL75, LB75b, Bec84, BBS95, BBM85, Bok95, Bri01, Bur76, BR77, Bus02, Com94a, CLFL94, CEC^{+}96, Chr83a, Chr88, CR94, Chr83b, Chr90, CN97, Com12, CBL96, Dan83a, Dan83b, DM07, Daw07, Doe11, Eve78, Fol77, Gal00, Gil93a, Gra00c, HK86, HG95b, How01, HN01, IBM88, Int90, Int93, IEE86, IEE88, JH83, KNI88, Kap92, KTB^{+}88, KM82, KN93, KK89b, Kiv84, Kra88a, LB96a, LM88a, LM88b, Lee89, LMKQ89, LMKQ90, Les77b, Lo79, LR78, LH84a, LH84b, Loz05b, LMT78, Lud79b, LB75a, LB78a, LB78b, LB78c, MC85a, MC85b, MH88, May78, MS87, MBKQ96, MNN05, MNNW15, Mic79a, MvT^{+}90, Mul79c, NP78, NBB^{+}02, Nor91, O’G01, Out82, PDL04, Raj00]. Operating [Ral75a, RRP^{+}08, Roc82, RKPP88, RB82, Sam87, Sch77, SR77, SGG05, SGG09, ST89a, ST89b, ST89c, SCK86, Sun77, TW97, TW06, Tan08, Tes86, TSP93a, TSP94, TSP93b, Ass88c, Vog6 , Vog77c, Wam83a, Wam83b, Wat87, Wei82, Won90, Zuc77b, d’A92, dOCC23, AEG^{+}91, Ale96b, And01, Ano93m, Ano98a, Ano00g, AAB^{+}05, Bac84, Bal96, BGW93, BHK^{+}84, Bun76, CE88, Cha96, Che81, CK06a, CK06b, Col84b, Com00, CKM06, CKM14, Don04, DuC94, Ewi86, FRAC85, FRRJ07, GM84, Gre81, HBG^{+}06a, HK82b, IEE91b, IEE94c, IBP03, Jal92, KWDB06, LM91, Lit88, Liu11, LS87, Lun96, MM88, Mil78, Moh85, Mue78b, Nat88, NV06, Par90, Pay84a, Pay84b, PTS98, PTS01, Pon97b, Raf79, RRP^{+}07, RLB08, RÓ01, RHP^{+}07, RRH^{+}08, SAS88b, Sam89, Sch93c, Tan87b, Tan01, Tan10]. operating [TPS98, Too11, WJ82, Win02a, Ygg94, YSC^{+}06]. operating-system [Ale96b, AAB^{+}05]. Operation [KPSS79]. Operations [CK78a, Pea77, SPS90, SM88a, SM88b, SM88c, Vog97]. operativo [RÓ01]. Operator [ASS79, Wit79, KD94]. Opportunities [Ano93-34]. Opposition [Ded09, Gla99]. Optical [Ama88, Kiv84, LEG88]. Optimal [AJ75, SU70]. Optimization [Gay90, IEE03, KL73, Kro99, Rob84a, SKP94, GL75, GL76, McC92, IEE03]. Optimizations [HKS^{+}04]. Optimizer [AQ84]. Optimizing [Ano02a, ESSB93, Fis02, GS01, Olk03, Tin78d, WJW^{+}75, Dwi04, FRRJ07, GS02, Pow84a]. Option [Sta04, Tuc98]. Options [IBM09a, Kri84a, Kri84b, Okr78, Ros95, Sim96b]. Oracle [Ern11, CG93b, Alo97, Ano02b, Bor14, Bro97, Bur01, Dav00, Emm11, GM11, IBM09b, JF12, Man01b, Mil16, Rea11, Sav11, VSC^{+}10]. Oracle8 [Alo98]. Oracle8i [Alo01]. Oracle9i [Don02]. Oral [Mahxx]. Order [SGP^{+}23, HKVR21]. order-aware [HKVR21]. Ordinary [BG97, Sch75]. Oregon [USE85d, USE90e, IEE04]. O’Reilly [Ano95b, Ano96b, Ano00a, Sch04, JW99, O’R93]. Organization [Asb00, IK83, Smi76]. Organizational [MWCR90, MWCR91]. Organizing [Pon97a]. Oriented [BCC^{+}91, BCGS91, Dav97, FR75, Hay77, HK89, Kro99, NH99, Pik84b, Rei77, Str6 , Str88, Tin78b, Wil82a, Wil82b, Cab83, Cox83, Ewi86, JGS^{+}08, KD94, KN93, MS95, Sch93b, Sch93c, SML06, SAL93, Smi91, You92a, Str91b]. orientierter [Dav97]. Origin [Ano19d, Wal99]. Original [AGHR89a, AGHR89b, Sma88]. Origins [TBGE89]. Orleans [ACM97, USE95b, USE98b, USE98e, USE00b]. OS/2 [DF00, Har93, SAS94b, TNY94a, TNY94b]. OS/400 [DuC94, Tay96]. OSCAR [Slo05]. OSDL [Ano02b]. OSes [Den99]. OSF [Bea93, Ber91, Kel90, Kob91, LPLF91, McM92, McM93, NO90b, The91, Ope91a, Ope91b, QO91, QO93b, Seb94, Yag91, You90a, You90b, You92a, ZRB^{+}93]. OSF/1 [LPLF91, The91, ZRB^{+}93]. OSF/Motif [Bea93, Ber91, Kel90, Kob91, McM92, McM93, Ope91a, Ope91b, QO93b, Seb94, Yag91, You90a, You90b, You92a, NO90b, QO91]. OSI [FFH^{+}86]. OSLO [She06]. OSx [Bot84]. Other [BBM85, CHL16, Lib93, Fri97, JP09, Neg06a, Neg07a]. Otorhinolaryngological [GVK^{+}03]. Our [CKT80, GMM90, HBB13, Sta05a]. Out-of-Order [SGP^{+}23]. Outdated [AAG^{+}16]. Outlaws [HM91]. Outlines [Hob91, Hob93]. Output [Eis79a, Hol87a, Hol87b, KMKO92, KV89, Rit77b, Rit90, Rit84d]. Outwit [Spi00]. overcomes [Mud97]. Overdog [CPJ^{+}98]. Overhead [GD18]. overlapping [Pik83a]. Overlays [Pug88c]. overriding [SZ95]. Overview [Bol99, Dav89, Feu79a, IBM88, Man93a, MK84a, MK84b, Ros88, WAC84, LS87, Tur86]. own [SML06]. Owner [Hum90a]. OWS [Dav00]. Oxford [Ano93g].

P [Ano00e, Bab78, SAS88a]. P-177 [SAS88a]. P-STAT [Ano00e]. PA [ACM04, DFG99, Ano88b, USE87c, USE88f, USE89g]. pace [Che97]. paced [Jae89a]. Pacific [Ano03f, Car88a, Car88b]. Pack [O^{+}98, Ass99a]. Package [Cou13, Dan83a, Dan83b, Daw86, Dwy79, KS85, Kau90, KV90, NB84, Oom13, Per84, Rit77b, SY91, Tar80, WH83, ZDS85, MMTK94]. Packaged [Hos84]. Packages [Sto76, BJM^{+}22, Bra04, Gru86, Spe96, VD01]. Packaging [Kro99]. Packard’s [Cle83a, Cle83b]. Packet [Ree06a, Skl91]. Page [KV89, MH96, Pis77, BI95, CDV^{+}94, ST03]. Pageable [MKB90a]. Paged [Jun85a, Jun85b, Pea83, Sha81]. Pager [GGZ02, WLC95]. Pages [SH85, Ano95b, Ano96b, Ano00a, Dys96, Yue01]. Pagination [GGZ02]. paging [BP91, PB91]. pain [Gra06b]. pain-free [Gra06b]. Pajamas [Lew99a]. Pakete [Ano96c]. PAM [Luc16]. panel [Ano19d]. Panic [DB95]. Panther [Dav06, JR04, TJ04]. paper [Woo94]. Papers [All86b, Cat91, D^{+}76, Lew21, PPT^{+}xx, ACM75, ACM82, ACM92]. paquetes [VD01]. para [RÓ01, VD01]. Paradigm [SF95]. Paradigms [Swa00a, Swa03]. Paradoxically [Tuc90]. Parallel [Bad98, Bal94, CDK^{+}01, EGL86, ELS88, GL94, JTUB85, KmWH10, LM88a, LM88b, LC90, MAT08, May00, Ung78, WB96, ZRB^{+}93, CS96a, FH89, HKVR21, LWL06, McC92, Mil93, RW87, Wil94]. parallel-processing [RW87]. Parallelism [Rol03]. Parallelization [CBB^{+}91]. Parameter [dJPDD84]. Parameters [BF9 , BF80, Wei88]. Parametric [Ano98d]. paranoid [Luc03a, Luc13a]. Parasight [PR96a]. parity [AR07]. Parlez [WB85a, WB85b]. Parlez-Vous [WB85a, WB85b]. Parse [Fre78, Fre78]. Parser [DS88, JS90, JP17, Mey18]. Parsing [AJ74, AJU75, Uhl78, van83]. Part [Ame94, Ano95f, Com94a, Cal08, FG79, HO93, Har01a, Har01b, Int93, IEE92a, IEE92b, NS94b, PH98, PJ89a, PJ89b, RW86b, RW86a, SPS95, Sei00, Set79a, Set79b, Set79c, Ame93, Lyo18a, Lyo18b, Lyo18c, PH99, She03]. Part-Task [PJ89a, PJ89b]. Partial [Col89, Col17, KS85, Kau90, Rob8 , Rob79b, Sch76a]. Partial-match [Rob8 , Rob79b]. Partially [Pov99]. Particle [IEE91b]. Partitioning [KL70, SK72]. Partitions [Pon02b]. Partners [Ano93t]. Partnership [Gar99]. Parts [Gop79, PH91]. Party [Loz05b]. paru [MO01]. PARX [LM88a, LM88b]. PASCAL [Pay84a, Pay84b, BJ01, DMM86, FG79, KE85, Ker81c, May84]. Pass [Lio79e, Sim97]. passing [Ber98, PS07]. passwd [Hoo83]. Password [And88b, Ano96a, FK90, Har92a, Har01a, Har01b, LT91, MT78, MT79, Spe96, WG00, LR89c].

Passwords [Dav98, KI99]. Past [Bha12a, Mas87, Mas88, Var91, WB85a, WB85b]. Paste [Lud77a]. Patch [HLT^{+}21, Jan06b]. Patching [Bis87, CCVLW23]. PatchNet [HLT^{+}21]. Patent [Coc01a]. Patenting [Coc01b]. Patents [FES^{+}99]. Path [HG93]. Pathname [MC85a, MC85b]. Paths [ALxx]. Pattern [AKW78, AKW85, GR76, SD95, AKW8 , AKW79]. Patterns [GS00b, Lut99, RW93, BY10, Hu86, SS95a]. PC [DeV00, FH92, Nee91, Per93, Ric94b, SU94a, Str94, SE96, Ano96d, Ano03c, Ano03e, Aud96, Bai96, Cho98, Fel94, Gia90, Gil91, Gil96b, Gra98a, Gra98b, Gra98d, JJ91u, Kad95b, Led94, LWL05, LWL06, Lun96, McK93b, Pug90b, Sau17, SG95, Sma88, SU94b, SE97, Tan87a, YL09, dFAAW90]. PC-DOS [SG95]. PC-UNIX [Cho98]. PC/VI [Sma88]. PCI [GK97]. PCI-based [GK97]. PCL11 [Ung78]. PCL11-A [Ung78]. PCs [Ano93j, Ano95f, Gra98f, Hae92, Lei95a, Yag91, Ger92]. PDEs [MAT08]. PDF [DF00, GL06]. PDF-Dateien [DF00]. PDP [LB75b, Bur76, BR77, Cla79, GT78a, GT78b, HR79a, Kar83, LB75a, MH78, May78, McG79, Ros78, Sch77, Tuo82]. PDP-11 [LB75b, Bur76, BR77, Cla79, GT78b, HR79a, Kar83, LB75a, MH78, May78, Ros78, Sch77, Tuo82]. PDP-11/44 [Tuo82]. PDP-11/45 [LB75b, BR77, HR79a, LB75a]. PDP-11/70 [Cla79, Sch77]. PDP-11’s [GT78a]. Peace [Mul87, Sch88]. Peaceful [JLSG84]. Pearls [Ben86, BKM86, Ben88b, Ben00]. PegaSoft [Ano00e]. PEM [GB92]. Penetration [AH11]. Penguin [Nis01a, Sal08]. Pentium [LB96a]. People [ASW^{+}01, BBM85, Gra84, Jam09, KJG^{+}98, LH01b, McL84, Luc13b, Mah07]. Perfect [FES^{+}99]. Perforce [Ano03f]. Perform [Dil02, Yos85]. Performance [ACM04, Alo01, Ano93-30, ADJS78, AHPS79, ACS79, BFGK89, BS88, BB93, Buy99, CKM85, CF20, CEC^{+}96, Chu98, CCF89, Coc94, Coc98, CG93b, Dan83a, Dan83b, De 78, DeV00, DRN^{+}23, Dow93, DD90, Fed79, Fed83, Fer82, FPG89, GS01, Gig06, DF79, Gra00a, Gre19, GLOR91, Hay77, HR78b, H^{+}79, Hoo79, HR79b, IEE94a, IBP03, KM86c, Kar92, Koe78a, Kra78a, KM83, LB96a, Lan84, LLM87, LKM84, LWL05, LWL06, Lou90a, Lut99, Lyc71, May00, MKL85a, MKL85b, McK13, MK91, Mel87, Min82a, Min82b, MK04, Mue94, ML02a, Nag05, Per87, Pos88, RW86b, Ros78, RšT06, SK10a, SH13, Sam87, SW99, Sim93, SKP94, SD87, STA87, Stu79, Sys97a, UNI93, WLL96, Wil83a, Wil83b, YR93, You95, ZB19, AKB85, AKCB86, Alo97, Alo98, AJ05, Ano03e, Bes05, Cab83, CHRG95, Fed84]. performance [FSW02, FM02, FGP^{+}91, GS02, GGJ89, GK97, KPT^{+}02, KS08, Lar83, The92c, The92a, The92b, The93, MMG07, Olk03, PF05, PBL86, Rze93, Sam89, SRW04, Slo05, TB05, ZGdDL^{+}96]. Performance-Constrained [Per87]. performance-setting [FM02]. Performing [Ano03f, AC11]. Perimeter [NZW^{+}02]. Peripheral [Dow77, Kra78b, LC73, LC75]. Perkin [Car82, RW81, Tuo82]. Perkin-Elmer [RW81, Tuo82]. Perkin-Elmer’s [Car82]. Perl [DF00, Sch93d, WS92, DuB01, Bla02, Bla04, BA96, Cox07, DH96, DP95, Fri97, Ger92, Inf95, JW99, LW03, Mah07, Ngu08b, Orw96, Qui95, SW91a, Til95, WIP^{+}97, Yue01]. Perly [Hes91]. Permissions [Dul96, HAB^{+}98, Lio79a, MC85a, MC85b]. Permuted [Che92, Mar78b]. Persistent [GZXZ23, HCF^{+}94]. Persistent-memory-related [GZXZ23]. Person [Yos85]. Personal [AISS98, CEC^{+}96, Haw85, O’B82, War84, And01, FRAC85]. personality [Neu00]. Perspective [Eur91, Fun87, McK06, McK12b, Nor02, WB85a, WB85b, Bha12a, LZHS11, McK12a]. Perspectives [Sat90a]. Perturbing [EP05]. Pesky [Ros95]. Peter [NH91, Nor99b]. Peters [Fla90]. PEX [Aki92, Ros88]. PEXlib [Gas92a, Gra94, Har94a, Hew93, Tal92, Wom94]. PF [Art03, Han08, Han11, Ree06a]. PFORT [Ryd74]. PfSense [Wil11, BP09, Zie16]. PG [UNI77, Vog6 , Vog76b, Vog77c]. PG-1C300 [UNI77, Vog6 , Vog76b, Vog77c]. PGI [Ano00e]. Phi [Cap82]. PHIGS [GG91, Gas92b, HHHW91, Kos92]. Philadelphia [DFG99, USE87c]. philosophies [Cor05b]. Philosophy [Fra83, Gan03, Gan95]. Phoenix [Ano93t, USE87d]. photographers [Goe07, Ham07]. Phototypeset [Vog77a]. Phototypesetter [CKT80]. Phototypesetting [Tar80, BJK87]. PHP [BGW04, Cox07, G^{+}06b, GB01, Hud06, HZ01, HZ02, LW03, Reh03, WN15, WT01]. PHP5 [Gab07]. Phrasebook [Gra06a]. PHYSCAL [DMM86]. Physica [Wei95]. Physica-Verlag [Wei95]. Physical [HP89c, DMM86]. Physics [IEE91b, RF93, Fel4 , Fel75]. Phyton [DF00]. Pi [Car90]. PIC [Ker81a, Ker81b, Ker82a, Ker84a, Ker91, Ker90b]. Picking [CPJ^{+}98]. Pico [Hol87d, Hol90a]. Picture [Bro90, Van90, Hol87d]. Picture-specification [Van90]. Pictures [Gay87]. Pieces [ZB84]. p‘ien [PWC00, PWC00]. Piepho [Luu90]. Pin [MT74]. Pining [Coc01c, FES^{+}99]. pint [Coo85]. Pipe [Sch93a, dVd93]. Pipeline [Gom85]. pipelines [HKVR21]. Pipes [Kra78a]. PipesFS [dBB08]. Pitfalls [DNQ^{+}83, Koe89]. Pithy [Ann00]. Pittsburgh [ACM04, Ano88b, USE88f, USE89g]. Place [Dia94, USE01b, Kar84]. placement [GJJ05, dJPDD84]. Plain [VRFJ97a, RFJ97, RFJ99]. Plan [Hai79, JC92, Sko98, Bai96, Duf90b, Dufxx, Gan08, Goo90, PPTTxxa, PPT^{+}xx, Pikxx, Prexx, PPTTxxb]. Planck [MMTK94]. Planet [Ano01]. Planning [Arn78, Chr98, CPJ^{+}98, CW00, RWNA87, FSW02, Har00, Won98]. Plans [Coc03a, CM78b, Gor81]. Plasma [IEE91b]. Platform [Ano01, Car04, Coc03b, Cro98a, Lac06, MWCR90, MWCR91, NH99, OBSB23, Pac02, RP02, AC11, Ano93r, Ano03c, Ano03d, BJM^{+}22, BS05, Don04, HBH18, MT93, Sch95c, Web95, Lac06, Ste01g]. Platforms [Ano92j, Ano06a, Ano19e, CBL96, Lei96c, Lei96e, Lei96d, Xia97, Hoo05, MRM06, NV06, PBM^{+}22]. Play [LH01b]. Players [Ano93-27]. PLD [Ano03c]. Please [Ano96d]. pledge [And17]. Plexus [Pic83]. Plotting [Lyc74b]. Plug [DF00, Jen02, OBSB23]. Plug-In [Jen02, OBSB23]. Plug-ins [DF00]. Plus [HHHW91, SS93, WP90, Che81, Hor02, PMW^{+}99, PMW^{+}00, WMPC83, Mil98, Mil00]. Pocket [Cam98, Fri03, Rob99a, Rob02a, Bar04, Bri05a, Bur01, Hea98, Pur04, Ran05, Top02a, Top02b, Top05]. Point [Ano96d, BF9 , BF80, Bro0 , Bro81a, Bro81b, CW89, CW91, Fer85, Sch81, TOML04, Bro6 , Bro77b, LMOT01, LMOT02]. Point-to-Point [Fer85]. Pointers [Jae86]. Policies [Oom13, WC98, RKSV20]. Policy [Mar01a, Mar06, HRS^{+}10]. Political [Tan84]. Poly [Mat85a]. PolyAWK [Pol88]. polymorphic [Car84]. Polynomials [BH1 ]. Ponders [Coc03a]. Pooh [Web95]. Poor [FJPR93, PB21, van83]. Pop [Col94]. POPL [Fis73]. Popular [Wol93, HWBH19]. popularized [Ale96a]. Port [Ble83, PF84a, PF84b, Tay96, TLR89, Hib96, Kin89, FHS77, FHS78b, Kau93d, TG78]. Portability [CE88, DuB93, Fil85, Flo90, Jae89b, JR78, Lyo77, LR77, Nae90, O’D83c, Sel94, SM89, TH77, Ver89, Wal94a, Wal97b, Wal98, Wei92, X/O88c, X/O88a, X/O88d, X/O88b, X/O89b, X/O89c, X/O89a, Che81, RS14]. Portable [Bli96, Blu7 b, Blu78b, Cus92, Dal99, Dal02, FW78, Fel79b, FW90, FHS78a, Fox79, FHS7x, Gal91a, Hal69, HK95, HMP83a, HMP83b, HG95b, Hor79, Hor90a, IEE86, IEE88, JH83, Joh78a, Kat99, Lap87a, Lap87b, RS90, Rai89, Rem97, SO92, Sny75, TvK83, Top90a, WW92, ADH^{+}89, Hoo05, Lew91, Mil78, MT93, Nat88, NSI^{+}92, Pay84a, Pay84b, Pow84a, See08, Ame93, Ame94, Com94a, Int90, Int93, Joh83, Lio79e, Mar78d, Row79a]. Portal [But02b]. Portals [McN99a]. Porters [DNQ^{+}83]. Porting [ACF^{+}86, BS79a, CB83, DM07, Daw07, Eyk86, Fra99, Gil88, Gla91, Hew93, JJ91s, JJ91b, JJ91c, JJ91a, JJ91t, JJ91q, JJ91d, JJ91o, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ91r, JJ91p, JJ91u, JJ92d, JJ92a, JJ92b, JJ92c, JJ92e, JJ92f, JJ92g, JJ92h, JJ92i, JJ92j, Jun85a, Jun85b, Kor97b, Kul06, Leh95, Lub90, MK85a, Mon00b, Noe98, O’B82, Ska94, Sun95, VB83, WJ82, WL94, BHK^{+}84, Bro98, CR92, GGM93, May84, MSW06, CT83]. Portion [Ano96a]. Portland [IEE04, MM06a, USE85d, USE85a, USE88e, USE88a, USE88b, USE90e]. Ports [Kra78a, Rib89, Jud90]. Positive [dSVdLT02]. POSIX [Ame93, Ame94, Com94a, Int90, Int93, Sko98, Ame90b, Alf94, Ano90a, Ano91b, Ano92j, Ano92k, Ano93u, Ano93v, Ano93z, Ano94d, Ano94h, Ano94f, Ano94g, Ano95g, Ano95i, Ano95j, Ano19b, Arc93, AAG^{+}16, BMR94, But97, CG95a, Coc03a, Col94, Eck91, GL91, Gal91a, Gal92, Gal93, GMB93, Har90, HG95b, IEE91c, IEE92a, IEE92b, Isa86, Isa91, IJ98, Jos95, Kut02, Lal93, LL95, Lew91, McN92, Mue93, Nat88, Nae90, QW93, RAMP94, RR93, Row95, Tip92, Wal92g, You93, Ano90b, Ano92l, Ano93m, Ano93-43, Ano94i, Ano95d, Cas94, CG95b, JC92, Kuh91, Sov92, Tuc90, Wal94a, Adh95]. POSIX-tively [Tip92]. POSIX.1 [Lew91, Zlo91]. POSIX.4 [Gal95]. POSIX.6 [Wei97]. POSIX/UNIX [IJ98]. possible [Len97, MC75a]. Possibly [JP17, Sem96]. Post [KB01]. Post-Mortem [KB01]. Postfix [Blu01, Den04, HK05]. PostgreSQL [Dup05, Wei04]. Postprocessing [KV89]. PostScript [Ado93, BQd86]. Pot [End00, Coo85]. Potato [Kle00]. Potential [Kol77]. Potpourri [Hun92c]. Power [Ano93-39, Bea98, Blo92, Edd90, GH75, GH4 , Kol93, McI90, MBL89, Ora21, POLo93, POL^{+}97, PPO^{+}02, PPO^{+}03, Rei93, RS78, Sou94, Sov92, Ano00g, Ano19c, CJ94, DF91, Gri14, HH04, KCTT06, MNC09, POL94, PF96, RRD^{+}03, Smi03b, VTS^{+}23, Mor97]. Powered [SFE10, Jen19, Sta04]. Powerful [WH83, Fri97]. PowerHA [Q^{+}11]. Powerpack [Gan04]. PowerPC [RFS05, RFS06]. PowerRAC [Ano00e]. Powers [Ano93n, Jen19]. powłoki [RB06b]. pp [Ano93-33, Wei95]. PPC [ASW^{+}01]. Practical [Arn93, BMS83, CMQ93, CQ83a, CKV93, Far92a, GS91, GS96, GSS03, Gia99, Hol12, HHHW91, JJ91d, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ92e, JJ92f, JJ92g, JJ92h, JJ92i, JJ92j, Kal16, Kel90, Kha02, Kri95, Lat02, Lea92, Loz05b, Lut99, May97, Oua91, Pat96, Pow00, RR96, Sau88, Seb94, Sel94, Sob89, Sob91, Sob95, Sob97, Sob99, Sob03, Sta07b, Ste90a, Wal94a, WALK10, WMK^{+}07, Wel95, Wel97, Wel00, WJH03, Hal06, Hal07, JJ91o, Luc03a, Luc13a, McC92, Mio97, Ree06b, Sob06, SS06, WALK12, Wil11, Wün13, Ano96b]. Practice [Bol99, CHL16, Cla00, Joh78a, LH01a, Pin92, San07, AV95]. Practices [CHL16, SKB00, Sul02]. practitioner [Cal10]. Pragmatic [Mas05, TH04]. Pragmatics [Sco09]. praktischer [SU94a, SE96]. Praxisführer [Bor01]. Praxiskurs [Die11]. Pre [Hai78a, Uni96a]. Pre-conference [Uni96a]. Pre-Release [Hai78a]. Precision [CM78a, Sch75, CLJC95, Dit93]. precompiler [Cox83]. Predictability [DI89]. Predictor [TR94]. PREEMPT RT [RMF19]. preemptive [FPG89]. Preface [Cro78a, Mar84]. Prefer [Der90]. prefetching [CKV93]. prefix [HL90]. Preliminary [ACS79, Car83b, DF79, HR78b, Lud79b]. Prelude [Set79a]. Prentice [Hid04, Mad94, Sta07b, Tho05]. Prentice-Hall [Mad94]. Preparation [BBM85, Kaz05, KLO78, Ker84b, Ker85, Ker89, Mag04, NCNS82, PS79, KL82, ?, ?]. Prepared [ASW^{+}01]. Prepares [Ano00d, Ano00f]. Preparing [Les77a]. Preprocessor [Ker4 , Ker75, Ker79b, SP90, van83]. presence [Ano95e]. Present [Mas87, Mas88, Var91, Bha12a, Fra5 , Fra76]. presented [ACM75, ACM82, ACM92]. presents [Ano93r, Len97]. preserving [YHR92]. Press [Ano16b, Jen05, Ano95e]. Presto [FES^{+}99]. Prevent [Far90]. Prevention [MTY01]. Preview [McK92b]. Price [Ano98d]. Prices [Ano93w, Ric94b]. Primary [DM79]. Prime [WJ82]. Primer [HK82a, LB96b, Lip91, LL83a, LL83b, Sta07b, WMP84, WP90, WMP92, Wey01, Fat82, Hal06, Hal07, Hor02, Man01a, Mio97, PMW^{+}99, PMW^{+}00, RFS05, RFS06, WMPC83]. primitives [Ano03e]. Princeton [Ube76]. Principle [Gla92, HKT05]. Principles [ACM03, AU77, ASU86, ALSU07, CS91a, Fis73, Get91]. Print [CJ94, Lei95a, SA95, TECB03, Jan03]. Printer [Dow77, Gli90, Hun92c, Lau79a, Pan94, Shi77, CJ94]. Printing [BK86a, GGZ02, Hof97, JH86, Swe01b]. Privacy [Coc01c, SW12]. Private [Min01, NZW^{+}02, FFPC22, ZG96]. privilege [XWRZ19]. Privileges [Gra01b]. Prize [Ano11]. Pro [Cha09, FKL^{+}08, FT09]. probe [Ale96a]. Problem [Bar00b, GH75, GH4 , Koe78a, Lin65, LK73a, LK73b, Wei88, GMSP00, Hal4 , Hal75, Joh05, WB06]. problem-solution [Joh05]. Problems [BG97, CW89, CW91, Cou84, Cou85a, Get86, HR78b, LLM87, Pug90b, YK15, Dun89, HU73]. Procedure [KL70, Lon77, BHR95, Lar83]. Procedures [Mar78c]. Proceedings [ACM79, ACM88, ACM89, ACM92, ACM97, ACM04, Ano88b, Ano89, Ano93g, Ano93x, Eur91, Fis73, IEE91a, IEE94b, IEE94a, IEE02, Mil84b, Ric77, Ste90a, Usr82, USE82, USE83, Ass83, USE84a, USE84b, USE85c, USE85a, USE85b, USE86d, USE86a, USE86b, USE87a, USE87b, USE87c, USE87d, USE88c, Ass88c, USE88d, USE88e, USE88a, Ass88d, USE88g, USE88b, USE88f, USE89c, USE89d, USE89e, USE89f, USE89b, USE89a, USE89g, USE90b, USE90c, USE90d, USE91a, USE91b, USE91c, USE91f, USE92a, USE92b, USE92c, USE93b, USE93c, USE93d, USE93e, USE93f, USE94d, USE94e, USE94b, USE94c, USE94f, USE95b, USE96a, USE96b, USE97b, USE97c, USE97d, USE98b, USE98c, USE98d, USE98e, USE98a, USE99b, USE99c, USE99a, USE02, ACM03, BSK87, ITC93, MS05, MN69, Sof84, Tob80, Uni96a, USE85d, USE86c, USE86e, USE87e, USE88h, USE90e, USE93g]. Proceedings [ACM82, Ano06a, DFG99, IEE03, IEE04, MM06a, Sof83, USE91d, USE92d, USE95c, USE00b, USE01b, USE01a]. Process [Ben98, BBS95, Ber76, Can80, DF95, Dwy79, ELS88, FG91, Fre91, GS00b, GGZ02, HBD97, Har94b, Har80, HRO82, Hun88a, Hun88b, Ima76, KK89b, Kro99, Lib87, LS92, MS88, MZG14, MZA96, PPTHxx, Tin78b, Tin78d, Ano90b, BCM88, Bra95, Cab83, DI89, GLG93, TG78, Wil98]. Process-Oriented [Kro99]. Process-Specific [BBS95]. Processes [BEHW86a, BEHW86b, BFS89, Kat99, Kep85, KJG^{+}98, Kil84, Hun88a, Hun88b, dVd93]. Processing [Ado89, AKW78, AKW85, ACK89, Bal94, Bar88a, Bar88b, BR87, Ber85a, Bri01, CCF89, Daw86, DO87, FY18, Hay77, HCN85, HKS^{+}04, IBC87, JTUB85, KM87, Lud77a, Lun93, Lun99, MCM78b, Mil84b, MS90, Nad98, Roc78, RT83b, Sha79a, Sri93, Sta89, Tal89a, Tal89b, USE89g, ZB19, vMM88, AKW8 , AKW79, Ano95h, Chr87, Gal91b, GGZ02, Kan93, Mil93, Ros84, RW87, Tar89, Wün13]. Processor [All78, Ano00d, Ano00e, Ano00f, Bur79b, Col83, Ell77, Hal72, Joh79a, KR77, Kra78b, Lyc77h, LC78b, LC78c, Pat83, RS78, SGP^{+}23, Sch79a, Tes86, Wes81, XZY^{+}10, Ame91, CLJC95, DJM86, Lyc72, MHO^{+}76, Nad92]. Processors [BS80, GL94, KmWH10, Kri84a, Kri84b, LC73, LC75, Neu06, CMR^{+}90, LR95]. Procs [MT93]. Procurement [Maj77]. produced [Rit85]. Producing [Coc01b, Cum77, Sch79b, Mee95, Sch79c]. Product [Ano93-38, Ano95f, Ano00d, BGPS91, Gra99b, Gra01a, Isl83, NMP82, Swi79, van97, Ano03c, Mat93]. Production [Fas78, Gom85, Thi92]. Productivity [GS12, Pan88]. Products [Ano93e, Ano98c, Ano00e, Ano01, Ano02b, Ano02c, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g]. Professional [Ano00e, BBD^{+}00, CK85, G^{+}06b, Hid04, JR93, Mag04, MS^{+}00, Mau08, Nor93c, PKP05, SPS93c, Sea20, Sto00, Osb92, Pec08, Raj98, See08, Tho06a, Ano02c, Gri01b]. Professional/Administrator [Mag04]. Professionals [AIX91, Mil99, Mer97, SZ95]. professionelle [PKP05]. professionellen [DF00]. Profile [BMR94, Eve78, SKP94]. Profiling [Bis87, Car86b, HG93, Rei78]. progeny [Too11]. Program [And91, Bak75, BKM86, Ben88a, BJK92, BHR95, Blu7 b, Blu78b, Bry83a, Bry83b, Che79c, CL90a, Cme86, C^{+}78b, Coo09a, Coo09b, Cum77, Far79, Fel78, GNV90, GDFR90, Gop79, Gra79a, Gra79b, HG93, Har79, Höf01, Jac77a, JZTB98, Joh78b, KL79, Kow79, Kow90, Les76a, Les77c, Les79b, Lio79d, Lo79, MC85a, MC85b, Mar78b, Mar78e, Mik92, PK84, PK88, Ral76a, Ral76c, Rei78, Rob75, Ryd8 , Ryd79, Sax85b, Sax85c, SR77, Ste85, Vog76a, Vog77b, Zuc77a, Ame94, Bah09, BJK87, Bin88, Bra04, Fel8 b, Fel79c, GIM07, GNV88, Hor90b, IEE92a, IEE92b, Mac97b, Mil01, Yag91, YHR92, HG95b, ML02b]. Programme [Str94]. Programmer [?, ?, ?, ?, AKW85, AS90, Bel79a, Bel79b, Bel79c, BR77, Com94c, Com94d, Cam87, DH77, Fla91, Goo07, GS00b, Gri01b, JL98, Ker79e, Ker79f, Ker79g, McI87, Nye95, Ope91a, Pat99, Rul96, Skj83a, Skj83b, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84f, TR71, TR72, TR73a, TR73b, TR74, TR75a, TR75b, TRS18, UNI77, Vic98, Woe02, Woe03, You92b, Zlo91, Ame90b, ACS98, Fre98, HGS99, LK00, Lew91, Lov07, MS87, PWC00, Pet98a, Pet00a, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, Coc01a, D^{+}76, DHM78, Lyo76b, MS77, Mas77a, Roo79b]. Programmers [Bea93, Chu98, Gar97, Gla79, Jae93, Lou90b, Mic08, Mik89, Sch94b, Sch95b, Zim98, GA99, Osb92, Oua92]. programmieren [Jor04]. Programmierung [Jor04, PKP05, RB06a, May01]. Programming [ACM79, ACM92, Ame89, Ado93, AKW88, ASW^{+}01, Ang77, Ano00b, Ano03f, Art90, Art94, Bad98, Bar91b, Bee93b, Bee94, Ben86, BKM86, Ben88b, Ben00, Ber91, Ber92, Böl01, BK89, BK95, Boy84, BKLL93, But97, CF20, CDK^{+}01, Coc01a, Col92a, Cop92, Cox83, Cur96, Cur97, Dal99, Dal02, DH05, DM79, DK16, DMM86, Fel79d, FG79, Fla94, Fla90, FGK89, Gal95, Gas92b, Gas92a, Gay87, Gay00b, GL06, Gla77b, Goe00, Gre92b, Gri00, GG83, GG86, GG90, Hal11, HS87b, Hel90a, Hel90b, Hel91a, Hel91b, HBG^{+}06b, Hil08, JL98, Jae86, JK73, JR89, JR90, JR92, JR93, Jor04, KLP88, Kel90, Ker74, KR78, KM78, KR79, KM79a, KM81, Ker81c, KP84, KR88, Kil96, KmWH10, KW90, KW03, Koc09, Koe90, KI82, LH01b, Lam83, Lap87a, Lap87b]. Programming [Lef81, Lev94, LB96b, LO97, McIxx, MB03, Mer01, NDW93, Nel90, Nor96, Nye88a, Nye90b, NO90a, Nye92b, NO92, NO93, Olc92, Oua91, Pow84b, PKP05, Pug88b, QH01, Ray04, Ric90, RW92, RJLK78, Rit79a, RR96, Rob01, RR03b, Rob04, Roc85, Roc04, Ros91, Ros99, SK10b, SBA92, Sco09, Sea20, Set79a, Set79b, Set79c, She93a, SR95, Spi00, Ste90b, Ste92b, Ste98, Ste99b, Ste00, Ste01c, Ste01d, Ste01b, Ste01e, Ste01f, Ste01g, SR04, SFR04, SR05, Ste92c, Str86a, Str80b, Str84a, Str86b, Str6 , Str88, Str91a, Str13, Swa00a, Swa03, Tan00, Tho68, TG89, Tur87, Val92a, Ves88, VRFJ97c, WS92, War85b, Weh01, Wel95, Wel00, WJH03, Woe02, Woo94, X/O88c, You89, You90b, YP92, vdL94, Ale10, Ama91, Art86, AB97, Bar94, BFS07, Bli96]. programming [Blo92, Bra92, Bro94, Cha97, Chi10, CS93, CS96b, DF91, Dun00, EC89, Fel8 a, Fel79a, Fer04, Fis73, FG03, Fos96, FJ97, Gay99, Gay00a, Geh84, Geh87b, Gov08, GKV21, Gri01a, GPP68, GPP71, Gue19, Had04, HM98, Hav11, HGS99, Hel03, Hil02, JPS00, JB99, Jon05, JSG00, KK89a, Kem91, Ker10, Kha02, KW85, KS08, Kur94, Kur97, Kut02, LB98, Lov07, Lyc71, Mad94, MM86, MS^{+}00, May01, McN99b, Mil93, MOS01, OP01, Olc01, Par93, Pet88, Pet01a, PF05, Rag93, Smi91, SR08, SR13, Sto00, Tee05, Teo13, TSC02, Tob80, VJR96, WWW99, Wal00, Wal01a, Wal95b, Wal01c, War01, Wel97, WNDO97, W^{+}99, WS96, You92a, ZK05, ZGdDL^{+}96, AKWA89, Gri01b, Str91b, Gra89b]. Programmpakete [Bra04]. Programowanie [RB06b]. Programs [Ano94e, Ano96d, CV80, CK78b, Cra88, Fed79, Fel78, Jen75, JR78, Kau93d, Koe79a, Kra88b, Lib87, Lib94, McIxx, MS89, MMS86, Per82a, Per82b, PH86, QL91, Rad79, Smi06a, Smi06c, Swa77, TNY92a, TNY94a, TNY92b, TNY94b, Wag79, Che90a, CK06b, Dar88, EP05, Fel8 b, Fel79c, Gue87, HM98, HC89, Kaz05, Lew91, Lub90, McC92, MS87, MR67, Oua92, SVV92, TH83a, TH83b]. Progress [HK79, Ano93t, Kuh91]. Project [ADJS78, CK85, EM00, GS12, GP88, HK95, Gur81, KM79b, PF84a, PF84b, Ger03, Gra04, Gra06b, Lyo08, Rit85, Sae03, Ube76, Yos85, But02b, HK79, Ung86]. project-based [Gra04, Gra06b]. Projections [ADJS78]. Projects [JL98, Mec05, OT91, TTB09, Gri13, Gri14, NW03, Neg06b, Nut01, Sch95a]. Projekte [Ric14]. Prolog [Ano98c, HL85c]. PROM [Mar78f]. Promise [Ora21, Sem96]. Promises [Law98]. Proof [KJHL96, KJH^{+}97]. Proofs [WG00, BY96]. Propagation [MTY01]. Proper [SK72]. Proposal [Bra76b, Gwo78, HLMR79, Hos84, Les77b, Pri78]. Proposals [Coc01c, JF81b, O’D83b]. Proposed [Che78, HA84, RAMP94, Vog78b]. proposes [Web95]. Prosecutions [Coc03c]. Prospects [War84]. Prospectus [Scu78, Wes83]. prospettive [McK93b]. Protect [Far90]. Protecting [RR03a, TR99]. Protection [BBS95, CLFL94, FA88, Rit79c, SGM^{+}00, LC93, WRA05, XWRZ19]. PROTEXT [Mil84b]. protezione [ZB01]. Protocol [dC87, DD90, Erm79, Fed88a, Fed88b, Gol79, Hol87e, Hol90b, Hub79, LKY93, Nye90a, Nye92a, SGN88, Sch89, SGFR90, SGFR92, Sha79b, Sku88, YLK93, YSF89, CP01, CP02, Len97, PC01, HBMT96]. Protocols [CS91a, FFH^{+}86, Hol91, Ste94, SG97a, Ste96]. Prototype [BSS^{+}95, BSSW96, BMR94, CSA^{+}87, Hoo78, Hoo79, RWNA87, BCE^{+}93]. Prototyping [Dav97]. Prototyping-Oriented [Dav97]. Prototyping-orientierter [Dav97].

ProtoVIew [Kro99]. Provably [GNDB16]. Provide [BS92, SZ95]. provided [Raf79]. providers [Fac02]. Provides [Sta04, Ano94h, Red97]. Providing [BS05, BS90, SC78]. Prozess [GGZ02]. Prozess-Systems [GGZ02]. PRS [Bro7 , Bro78]. Prüfungsvorbereitung [Mag04]. psad [Ras07]. pSeries [Mly09]. Pseudo [Cro98b, HO 78, Rob79a, Rob82, WN15]. Pseudo-ATS [HO 78]. Pseudo-Random [Rob82, Rob79a]. Pseudo-Terminal [Cro98b]. Pseudonymizing [Fle02]. pseudorandom [Wal96]. Pssssttt [Gra88]. PSX [Ano94h]. PTG [Kom02]. pthreads [LB98]. PTOPP [McC92]. PTR [Sta07b, Tho05]. Public [Bar01b, TG79]. Publication [Ear87]. Publications [Bee96a, Bee96b]. Publishing [Ano00e, PS^{+}09]. purely [DL08, DLP10]. Puroguramingu [AKWA89, DF91, Sas91c]. Purpose [CS86, CH83a, CH83b, Fas78, Gla77a, MH88, MK88, Ng77, Tho79, PF05, SK10b, Var96b, Wei97]. Pushes [Law98]. pushing [Che97]. Put [Wal95e]. Putting [O’D90, ZB84]. PuTTY [Luc12, Luc18b]. Puzzle [QW93]. PVM [CS96a]. PWB [BG77, Bon79a, Bon79b, Bro79, Dep79b, DHP77, DS77, Eve78, Feu79a, Gla77a, Hai77, HRW77, MS77, Mas77b, PS77, Pea77, PS79, Pis77, Ros77b, Sab76a, Sab77, Sab78, Smi76, Smi77, SM77a, SM77b, SP77, Smi78, Wan77, Wan78, Wen79b]. PWB/Graphics [Feu79a]. PWB/MM [MS77, Ros77b, SM77a, SM77b, SP77, Smi78, Wen79b]. PWB/UNIX [Bon79a, Bon79b, Dep79b, DHP77, DS77, Eve78, Gla77a, Hai77, HRW77, Mas77b, PS77, Pea77, PS79, Pis77, Sab76a, Sab77, Sab78, Smi76, Wan77, Wan78]. PYMES [RÓ01]. Python [Lac06, Ric14, SML06, Bla02, Bla04, Cox07, GJ08, Lac06, Ngu08b, TSC02]. Python-based [Lac06].

Q&A [Jen00, Jen02]. Qbytes

[Sem96, Sem96]. QED [RT70]. Qemu [CK06a, CK06b, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i, MZG14, CK06a, CK06b, CK06d, CK06g]. Qmail [Blu00a, Lev04]. QoSbox [CL06]. QT [Gri01a, Dal99, DH02, Dal02, War01]. Qt/Embedded [DH02]. Quadrature [Blu75, Blu77]. Quality [Ano01, FJPR93, GS12, Neu00]. quality-of-service [Neu00]. Quantitative [CL06]. Quantum [Coc01a, Coc01b]. quart [Coo85]. Quarter [Sal94, Kol94]. Quartz [GL06, Tho06b]. Quelltext [DF00]. Quemu [CK06h]. QUEST [Alb77b, Alb77a, Ano03d]. Questionnaire [Alb77a]. Questions [CPJ^{+}98, MM96, Tim93, Pub11, Vib11]. Queue [Lic79b]. Queueing [Cho98]. Quick [Ano00e, BD94, Gt92b, OD^{+}86, Olc91, Rob99b, Rob99c, Ros90, Ros93b, Ano00a, Dea01, Don02, Hek97, Hug98, OH93b, Ric94c, SH99b, S^{+}00, S^{+}05]. quicker [Bin88]. Quicklet [San07]. quicksort [McI99]. QuickStart [RR99]. QUIK [GJLB77]. Quintillions [Sem96].

r [DF00, Oom13, Uni98b, Wün13]. R&D [BYVF08]. R. [Luu90]. r0a [Ano05]. R5 [RMZ01]. rabbits [SJP^{+}09]. Rack [Ano02b]. Rack-Mount [Ano02b]. RAD [Ano03g]. Rage [Ste01e]. Rago [Teo13]. RAID [Ano92a, Ano93y, BV06, Vad02]. Raiding [BV06]. Raises [Ano93y]. Raleigh [Ral77]. Ralph [Ano00b]. RAMIS [Ell79a]. RAND [War08]. Random [AU3 , AU74, Bey77, Don22, MM06a, MCM06, Rob82, MCM07, Rob79a, WN15]. Ranges [Gar99]. Ransom [Sta05b]. Ransom-ware [Sta05b]. Rapid [AU3 , AU74, HP89a, HP89b]. RAppArmor [Oom13]. RASH [HP89a, HP89b]. Rasmol [Vol94]. Raster [Dan83a, Dan83b, Duf90a, McI92, McIxx, Pav0 , Pav81]. Rasterizing [Hob89]. RATFOR [Ker4 , Ker75, Ker79b, May78]. rather [SZ95]. Rating [Nis01b]. ratings [HBH18, HWBH19]. Rational [Ker4 , Ker75, Ker79b, Bro3 , Bro74]. Rationale [Mac83]. RATSNO [Han77]. Raw [McS79, Pas88, Tin78a]. RBCS [NP78]. RBCS/RCMAS [NP78]. Rc [Duf90b, Dufxx]. RCM [H^{+}79]. RCMAS [KPSS79, HR78b, NP78]. RCS [Dav00, DS00]. RCU [MFBWW20]. RDB [Pru02]. RDBMS [Ros94b]. Re [MKL^{+}95]. Re-examination [MKL^{+}95]. Reactor [Sch93b]. read [GMTW08, CK78b]. read-copy-update [GMTW08]. Readability [CK78b]. readahead [WXX08]. Reader [McI87]. readers [BI95]. Reading [Pug90c, Spu92]. Readings [?, ?, ?, ?, Sto88, Lea90]. Ready [PS22, ZB19]. Real [Abb03, All83a, All83b, Ano92l, Ano92m, Ano93z, Ano94i, BFGK89, Bar88a, Bar88b, BL75, Bla83a, Bla83b, BRD77, Bur76, Cha79, CG95b, CN97, Cra88, Dou02, Gal91a, Gal92, Gal95, Ger02, HRO82, Hor82, IEE94c, Isa86, Joh95, Kol77, LH84a, LH84b, Lyc78a, Mar04, McS78, Nad98, Nis01b, NMP82, PR96a, Pan88, PNK02, PB84, RMF19, RT83b, Rze93, She03, Sta05a, Sta07b, STT86a, STT86b, TBS87, Ass88c, Vol75, WM00, Ano95j, AKST93, CG95a, Fer04, FPG89, Gal91b, GMTW08, Hal06, Hal07, HST08, IEE91b, Luc13b, PBM^{+}22, PB21, Rau01, RLB08, SJP^{+}09, Wai95, FPG89, IEE91b, ACS79, Epp00b, FGP^{+}91, HST08, Hoo78, Hoo79, Nis03, She03]. Real-Time [Abb03, All83a, All83b, Ano92l, Ano92m, BFGK89, Bar88a, Bar88b, BL75, Bla83a, Bla83b, Bur76, Cha79, CG95b, Cra88, Dou02, Gal92, Ger02, HRO82, Joh95, Lyc78a, Mar04, McS78, NMP82, PNK02, PB84, RMF19, RT83b, Ass88c, WM00, Ano93z, Ano94i, IEE94c, LH84a, LH84b, Nad98, PR96a, Rze93, She03, STT86a, STT86b, Ano95j, AKST93, CG95a, Fer04, Gal91b, GMTW08, PBM^{+}22, RLB08, SJP^{+}09, Wai95, Epp00b, Hoo78, Nis03, She03, HST08]. Real-world [Sta07b, Hal06, Hal07]. REAL/IX [FPG89]. Real/star [FGP^{+}91]. Realistic [Bro0 , Bro81a, Bro81b, Bro6 , Bro77b]. Reality [Ano96d]. Really [Ano93d, FES^{+}99, Guy00, KMP^{+}20]. Realtime [Ano92k, Ano94a, Ano94d, BMR94, Har80, Ame94, Ano93v, Eck91]. Rearrangement [AS93a, AS95, AS97]. Rebel [Moo01a, Moo01b]. Rebooting [WS78]. Recipe [Fer85]. recipes [Joh05]. Recognition [GR76, Kem93, SD95]. Recommendations [C^{+}78b, War90]. Recommended [OT07]. recompilation [EP05]. Reconciling [Gal91b]. reconfigurable [Don04, SB08]. reconfiguration [PBM^{+}22]. Reconsidered [Ray08]. Reconstruction [RF93]. Record [Ano03f, Bey77, Lic79b, ACM75]. Record-Performing [Ano03f]. Records [Bey77]. Recoverable [SMK^{+}94]. Recovered [Coc01a]. Recovering [DS00]. Recovery [Ano00b, BS92, Bur79a, DKM96, IBM09a, KK89b, Pre99, Wei78a, WF01, Bin88, Pre07]. Rectangles [Kau90]. recursion [Ger85]. Red [Jan04, Pri99, Wan02, Ano03a, Bar01c, BW02, Bel00, Col02, G^{+}06a, HS99, HS00, HH05, Kom00, KC00, Kom02, Kro99, Max00, McC99b, McC02, McC03, McC04, Neg01, Neg02, Pet02b, SBR01, SH99a, Ser00, Sob03, Sob06, Tom11, Hag00]. RED-Linux [Wan02]. REDAS [McS78]. Redesign [KSU94]. Redmond [Dav96]. REDUCE [HS83]. reducing [ST03]. Reduction [Bon79b, KLK^{+}02, CDD^{+}07]. redundancy [Ano95g]. Reengineering [McG97]. Reepmeyer [Wei95]. Refactoring [DYZC20]. Reference [All86a, Ano02c, Ari92, Bec06, Com94c, Com94f, Cam98, Che79d, Cof90, Cos97, ES90, FB93, Fri03, Gt92b, Gra94, Gra01b, GGZ02, HS84, HS87a, HS91a, HS95, HS02, Hid04, JF12, Joh7x, Kos92, Mik90, Mor75, Nye88b, Nye90a, Nye90c, NO90b, Nye92a, OD^{+}86, O’R90b, Ope91a, Olc91, Ope93, Ope94, PF84a, PF84b, PW94, Pet98b, Pet00b, Rit2 , Rit74, Rit77a, Rit79a, Rob99b, Rob99c, Rob99a, Rob02a, Ros77b, Ros90, Ros93b, Rul96, SSC93, SSC95, SGN88, Sch89, SGFR90, SGFR92, Sch91a, Sch91b, She93b, Str84a, Tal92, TV04, Tho72, Tji85, Van91, Vog78a, You90a, Ano85, Ano00a, Bec07, Bur01, Cra90, Dea01, Don02, Dys96, Emm11, FG80, Fei01, Fre98, Geh84, Haw00, Haw02, Hek97, Hor98, Hug98, Jos12, Kee05, KFO97, LK00, Mat80, MSW06, OH93b]. reference [Pet98a, Pet00a, Pet01b, Pet02a, Pet02b, Pet00c, Pfa00, Pur04, Ran05, Reh05, Ric94c, RHK^{+}07, Sch00, SH99b, S^{+}00, S^{+}05, Smi03a, Sun93c, SM85, Top02b, Uni85, Win99, Win02a]. References [AS88, AS91, Der90]. Referencing [Vog74]. Referenz [GGZ02, dG96]. Reflections [BTS09, Boh86, Cus04, Rit84c, Rit85, Rit12, van86]. Regenerating [Lyc76b, Lyc77d, Lyc77e]. Registers [Mor78b, SM77b]. Registration [PMD88]. Regression [Cro78b]. Regular [Cox07, Cox09, Cox10, Cox12, Fri02, KP99, Pik06, Sab76b, Fri97, Fri06, Tho68]. REGULUS [All83a, All83b]. Reissue [Kir77]. Related [Can80, Hae90e, Hae91a, Hae91b, HCF^{+}94, Sto93c, Sto93a, Sto93b, Sto93d, Sto94b, Sto94c, Vog79b, Wal91a, Wal91b, Wal91c, Wal92e, Wal92f, Wal92a, Wal92b, Wal92c, Wal92d, Wal93a, Wal93b, GZXZ23, Hub96, LRB^{+}19, Sto94a]. Relation [KKF^{+}84]. Relational [CS86, McS78, Par78, Ros94b]. relational-database [Ros94b]. Relationships [GS12]. Relative [Fed79]. Relay [Coo98]. Release [All86a, AD89, Ano02b, Arn78, Bel79d, Bro19, CBB^{+}91, Cof90, CM78b, Dep79b, FHS91, Fla91, GC94a, GC94b, Hai78a, Hai78d, Har93, IF92, Lud77b, The93, MKB89, MP92, NDW93, Nor93a, Nor93b, Nor93c, NO90b, Nye95, Ope93, Ral78, RRF90, SGFR92, Sha79b, Ame90b, Ame90a, XWS95, Ano00a, GSN90, Nat90, Nat94, Ros96, SAS88a, SAS88b, SAS91b, SAS93a, SAS94a, SAS94b, SAS95, Bee85, Bee86a, Dav89, Sob91]. Released [Bar00b, Coc01b, Coc03c, Coc03b, Swa03]. Releases [Ano01, Ano03f, Den99, McK98, Ano03d, ACS98, SG97a, SG97b]. Relevance [MZH22]. Reliability [Ano93-32, MKL^{+}95, MFS90, Ano93y]. Reliable [Any85, HBG^{+}06a, Sea14]. Reloadable [Alb84]. Remark [Fox79, Lev98]. Remember [Ros95]. Remote [Cla79, GT79, HK79, Lyo76b, Sab76a, Sab78, Lar83]. Rendering [Kar92, PH87]. rendszerek [Lás05]. Reno [ACM89]. Rep [IEE91b]. Repairing [Wan77, Wan78, HBG^{+}06a]. Repelling [CB94, CBR03]. repetition [Mar82]. Replicated [Bre86, LGJS91]. Replicates [Sma88]. Reply [KB01]. Report [Bor79, Cha85, Cou85b, DF93, EM00, Far95, FDMS79, FJPR93, HK79, HR78b, Hol97, McK93a, Pin99, SAS88a, Sha79a, SST90, Ska94, TBGE89, Wal94b, BFS07, SAS94a, SAS95, Ung86, Mac86, Mac90]. Reporting [Can79, Dea78, LPG^{+}18, Mar78c, PW79]. repository [Spi17]. Representation [BEHW86a, BEHW86b, Lin79b]. Reprint [RT83a, TH83b]. reprinted [LT80]. Reproducible [CW15a, CW15b]. Republic [BSK87]. require [Ano94i]. Requirements [IEE94b, Miy86, RAMP94, Stu78a]. Rescue [ASW^{+}01]. Rescuing [JN88]. Research [CK85, HIT^{+}02, HB82, HM90b, JLSS77, Law98, Law83a, Law83b, McI87, MKS^{+}89, Ric77, Rit84c, Rit85, Rit12, Sti83, TRS18, Wil90, JJ91r, Ube76, Ano93e, LR90]. researchers [Jen19]. réseaux [Sch94a]. Reservoir [HIT^{+}02]. Resident [Bar88a, Bar88b, GT78a]. Resistant [AM86]. Resort [USE01a]. Resource [BW88a, BW88b, BW88c, BMF15, BGMR91, ESSB93, MCH^{+}99, OT07, Sim91, Smi89, WIP^{+}97, BAC^{+}89, Bru02, DI89, HH04, Mly09, Moh98]. Resources [BBM85, HG93, Mar78a, Vol97, Gol95]. Response [Hoo78, Hoo79, SM88a, SM88b, SM88c, Mor79c, Ras07]. Rest [Mea88]. restart [BBHL08]. Restoration [BBK13, Too09, Too10]. Restore [Ano00e, Poe87, SKB00]. Results [Cod89, Gin83, STA87, Tuo82]. Resume [Whi75a]. Retaining [Kra88b]. Rete [Thi92]. Rethinking [Kam02]. Retrieval [ACS79, Hoo78, Hoo79, HL85c, Rob8 , Rob79b]. Retrospection [Sto80]. Retrospective [Ker84b, Ker89, MKB90b, Rit7 , Rit78, Bor88]. Returning [Sch97]. retuschieren [DF00]. Reusable [Kri95]. Reuse [Hub02, WMK^{+}07]. Rev [Sun86]. Reveals [Gar97]. Reverse [FKL^{+}08]. Reversion [GH75, GH4 ]. Review [Ano95b, Ano96b, Ano00a, Ano16a, Ano16b, Bea98, Chr98, Col92a, Cuc90, Cur97, Edd90, Fio04, Gil93a, Gwy90, Had04, Hid04, Hul05, Jen05, Kad95b, Kad95a, Koh06, Kol93, Kol94, Lea90, Lea92, Mad94, Mat97, Nag05, NN92, Sch95b, Sch04, She93a, Sim93, Sta07b, Teo13, Tho05, Uma96, Wei95, van97, Tho19a]. Reviews [Ano93q, Ano95f, Ano00b, Far92a, Sal03, Zwi92, HBH18, HWBH19]. Revised [Lyo79, Per78]. Revision [Ope91a, Yos85]. Revisited [Fer82, MKL^{+}95]. Revisiting [BBK13, Har01a, Har01b, LRB^{+}19]. Reviving [Sch91c]. Revolution [AGHR89a, AGHR89b, Moo01a, Moo01b]. Revolutionary [Ray99c, TD01, Ray01b]. REXX [Ger92]. RFC [Hol75, Kir84]. RH302 [Jan04]. RHCE [Jan04, Tom11]. RHCSA [Tom11]. RHEL [P^{+}08]. Rich [Lac06, Yos85, Lac06]. Richard [Teo13, Ano95d]. richtig [DF00]. Riding [Gre93]. Right [McI92]. Ring [KDP83, MYH00]. Ring-Star [KDP83]. RISC [Luu90, Ano90a, CMR^{+}90, DRN^{+}23, Har90, ILG19, OBSB23, PS22, PGW^{+}20, Pin99, SGP^{+}23, SNM22, VTS^{+}23, Var96a, Var96b, ZB19]. RISC-V [DRN^{+}23, ILG19, OBSB23, PS22, PGW^{+}20, SGP^{+}23, SNM22, VTS^{+}23, ZB19]. RISC/VMS [Har90]. RISCy [Ano93-27]. Ritchie [Aho12, Ano11, Bro20, Ker12, McI12, Pat12, Sch12]. Rivalry [OH97]. RJE [Sab77]. rksuite 90 [BG97]. Road [Gil96a, McIxx, SM79, Ste01e, BAKR19, Che94, Ano94l]. Roadmap [Mas77a, Jan03]. Robbins [Ano00a]. Robert [Tho05, Gre11]. RoboMon [van97]. Robot [Coc03b, Gri14]. robot-building [Gri14]. Robotic [Coc03a, Gri13]. Robotics [Bar00b, Gri14]. Robustness [MCM06, Cha96, MCM07]. Rocks [Slo05]. Roff [McI73a]. Rogue [Ano00e, Ano03e]. Role [BBH18, DDLL94, Ano95e, LL93b, McK85b]. Roles [Mas88]. ROM [AL97c, Ale00, Ame95, Ano98e, Ano00g, Com94b, Kom02, RRD^{+}03]. ROMP [Sau17]. ROMs [HS99, Kuo98a]. Roomful [Woz82]. root [JJ91t, Kob99]. Rooted [Siv01]. rootkit [Blu09]. rootkits [Kon07]. Roots [Lyo18a, Lyo18b, Lyo18c]. Rose [Cas94]. rough [Ano93m]. Roundup [BGPS91]. Router [Fel96b, Hol97, BP09]. Routers [Cho98, NZW^{+}02, CL06, Man01a]. Routine [Fre78, Str80b, Uhl78]. Routines [Fra79, Gay90, ?, ?]. Routing [Fed88a, Fed88b, Mar06, Skl91, Mar01a]. Royal [McIxx]. RP3 [CR92, EGL86]. RPC [Blo92, Lei00, SR95]. RPTRTR [Hol97]. RS [Tip92]. RS/6000 [Tip92]. RSVP [MHS02]. RSVP-aware [MHS02]. RT [Bel79d, Bra78, Che78, Ima79, Kap79, KM79b, Loi79, Lud78a, May78, Ral80, Sau17]. RT-11 [May78]. RT/PC [Sau17]. RTLR [ACS79, Hoo78, Hoo79]. RTOSS [IEE94c]. Ruby [Bla02, Bla04, Cox07, Ngu08b, Tom02]. Rule [McC95, McI73b, Ves88, McI74]. Rule-Based [Ves88]. Rules [Sha94, Mee95, Sea14]. Rumor [Mei84]. Run [Swe07, Com00, Cra93, Ava92]. Running [Bec84, Blu00a, BS90, Cra93, DW06, Dav06, Gue87, Har80, HS83, KM83, LS83, McK02, Sch77, Tur87, WKDD03, WK95, WK96, WDK99, WKDD02, FHR84, McI69a, Ano95b]. runs [SAN08]. Runtime [BMF15, SR85a, SR85b, IEE03, Pay84a, Pay84b, SB08]. rute [She02]. RXD [Swa77]. RXD-11D [Swa77].

S [Luu90, Sau88, Ano93-33, Ano97b, KVCS98, Mil98, Mil00]. S&Tnet [MTT83]. S-40 [Ano93-33]. S-PLUS [Mil98, Mil00]. S.u.S.E [KGG00]. S/390 [Ano97b]. S7 [Kul06]. Saber [KLP88]. Saber-C [KLP88]. Safe [Bor98, Wei82, Sea14]. SafeCode [SW12]. Safeguarding [SW12]. Safety [BBH18]. Safety- [BBH18]. Sagamore [ACM03]. SAGE [Far92a, Sat93, Sim93, Zie93, Zwi92]. said [Web95]. Sair [Mag04, DRP01, Mag00, Mag04, Sul02]. Salama [Gra89b]. Salesman [Lin65, LK73a, LK73b]. Salm [Hid04]. Salt [Sof84, USE84a, USE95c, USE95a, Ano03d]. SALT-based [Ano03d]. Sam [Pik90, Pik87]. Samba [Ano98c, Bai00, BT98, BDH00, CS99, DeR99, DeR00, ECBK00, Hun00b, Lei00, Nas97, Smi01, TV04, TECB03, Wol04a]. Samba-3 [TV04]. Sample [AKDN90, SAS91a]. Sams [Art99, Bal99, CS99, CCB00, Gay99, Hsi01, LH00, MB00, Ray99a, RBHB99, SBR01, TA98, Tay03, Tay06, UT03, WT99, Wel99a]. San [ACM92, IEE03, Uni96a, Uni96b, Uni98b, USE83, Ass83, Ass88d, USE89f, USE89b, USE91b, USE92b, USE92d, USE93f, USE94d, USE94f, USE96b, USE97b, USE00a, USE01a, USE02]. Sandboxing [And17, Oom13]. Santa [Ano93x, USE93d, USE93g]. SAR [CLJC95]. Sarge [Ano05]. SAS [Har93].

SAS/SHARE [SAS94a, SAS95]. SAS/TUTOR [Har93, SAS94b]. SATAN [Ale96a, RR95a]. Satellite [BS80, Ell77, LC78b, LC78c]. Satellites [Bar01b]. Saved [FST91]. Saving [HSM03]. Savvy [Kro99]. Say [Adh95, Ano19c, Jac87, Sta05a]. Sayings [Ann00]. Says [Coc03b, Ale96b]. sbrk [Sch91c]. SC2002 [IEE02]. SCADA [Ano95g]. Scalability [She06]. Scalable [DJ93, GL94, Hes94, IEE94a, Sat90b, AAB^{+}05, Bar00b]. Scale [GVK^{+}03, GES98, GP88, Hun92c, USE97a, CR92, DFC^{+}23]. Scaling [BP86, Neu06, Spe00]. Scan [Bar01b]. Scanning [AKW78, AKW85, Mar10, AKW8 , AKW79, Lyo08]. ScanSAR [CLJC95]. ScanSoft [Ano03e]. Scattering [PQM^{+}21]. SCC [SSC93]. SCCS [BG77]. SCCS/PWB [BG77]. scene [Woo94]. scenes [PH87]. SCFS [IHR99]. Scheduler [Ben98, Boh86, Cat79, MR14a, NHNW94, Ral77, Rob03, SS95b, STA87, XZY^{+}10, AKST93, Bra95, TCM07, WTKW08]. Schedulers [Pea83, GLG93]. Scheduling [Bun76, CDV^{+}94, CN97, CS90a, CS90b, DF95, ESSB93, Hel92, Ral75b, Ral76b, STA86, MNN04, PBWB84]. Scheme [DF00, HKT05, KLK^{+}02, McI69a]. Schemes [Hol80]. Scheutz [TBGE89]. school [Hos05]. Schritt [Str94]. Sci [Goo90]. Sci-Fi [Goo90]. Science [Bar01b, BHM84, Fra83, Gin83, Lew21, Mac83, Mil84c, Pan88, Ame09, Gre96, Hin88]. Sciences [Mil84c]. Scientific [HS82, Che90a, SML06, Spe00]. scientists [ABS95]. SCM [Ano03f]. SCO [DF00, Kuo99, CT94, Ric96]. SCO-Unix [Kuo99]. Scratch [Too17, Bee00, Bee16, Hal00]. Screen [Han86, Kat82a, Kat82b, Uni85]. SCRI [Car89a]. Scribus [JP09, PS^{+}09]. Script [CH77, PW77, RB08, Sch97, Ano96d]. Scriptics [Ano00e]. Scripting [Blu11, Bur04, Mic08, Ngu08b, RB05a, Blu08, FJWA05, Joh05, Mic03, Pet09, Ros06, See08, Fio04]. Scripts [Bla04, BJM^{+}22, Bla02, BA96, RB05b]. SCSA [San07]. Sdb [Kat79]. SDE [Ano02c]. SDEF [EC89]. Sdiff [Mar78e]. SDK [Jon02, Kro99, ML09]. SDL [Kos87]. SDS [HLLP79]. SDS/UNIX [HLLP79]. SE [Far79]. Seal [JC92]. Search [AC75, Ale96a, Tho68, Cox12]. Searching [HS90, HS91b, Les88b, McI63, SB97]. Seasons [Fra90b]. Seattle [ACM93, IEE94c, USE97a, USE97d, USE98c, USE98a, USE99a]. SEBASTiAn [PRC^{+}23]. Sebastopol [Ano95b, Ano96b, Ano00a]. seccomp [And17]. SECD [Daw82]. Second [Bea98, BSK87, D^{+}76, GLOR91, Hul05, Lio79e, MT74, Ano96b]. Second-Generation [GLOR91]. Secondary [McK96a]. seconds [Bri05b]. Secret [Dat14, MYH00, Les88a]. Secretarial [Eng76, Sha79a]. Secrets [Arm96, Ray98, vdL94, Bar01c, HLK01]. Section [NB84]. sector [JJ91r]. Secure [Ano93d, Bau02, BBB^{+}20, Car99, CJL^{+}89, Car89b, Car04, Kra88b, Len94, Lin02a, Nee90, PN04, RM92, RMxx, Reexx, Rei03, Sat90b, Sea05, ST89a, ST89b, ST89c, Won90, d’A92, dC93, Ano96f, Den04, Dwi04, GNDB16, RKSV20, Sea09, Sea14, BS01, DCN02, Mel95]. SecureShare [Bor98]. Securing [Ano00b, Bea04, C^{+}03, Dav98, Loz05b, MGK^{+}03, SFE10]. Security [Ale96a, AH11, And88b, Ano93o, Ano94g, Ano06b, Ano16a, Arn93, BBH18, Bra04, Car88a, Car88b, Car93a, Che96, CB94, CBR03, Chr90, C^{+}03, Chu03, Cur92, DM88, De 93, DF93, Don96a, Don98a, Duf89c, Far92a, Far90, Far92c, Far92d, Far95, FK90, FH88, GB92, Gar97, GS91, GS96, GSS03, GS00a, GM90, Gra99c, HJAW88, HT92, HT94a, HT94b, HR91, Hsi01, IB93, Kal16, Kur90, Lin02a, Loz05b, May97, MR92, MRxxc, McK93a, MS05, Mel02, MT78, MT79, Mor79b, NIS92, NBB^{+}02, NZW^{+}02, NHR84, Oom13, RR03a, RW84, RW86b, Rie93, Rit79b, RR93, RG91, Sat90a, Sat89, Sef93, SH95, Sku88, SW12, Sys97b, TB05, USE88e, USE88a, USE88b, USE90e, USE92e, USE93g, USE95c, USE95a, USE99b, WG00, WC98, Wei78b, WN02, WK85, And17, AK^{+}07]. security [Ano93g, BSB03, BsswLB05, Bis89, Bla98, BTS09, DC18, EBS08, GJJ05, GM84, Gre00, HLK01, KHG10, KHP05, Lon90, LR89c, Lyo08, MM00, McC98, McN97, OS06, PRC^{+}23, Rau01, Ray01a, Ree88, Ros00, Shu06, Smi05a, Spe96, TR99, Tox07, Vea10, Kra83, Lin02a, MMC06, McC05, MGK^{+}03, Ano16b, Lea92, Ano96b]. Security-Critical [BBH18]. sed [Dou91, PWC00, Rob02a, Ano85, McM78a, Art99, Bla02, Bla04, Kruxxb, McM79, McM90, Rob02b, Ano00g, DF91, DR97, Pat99]. seddo [MO01]. SEDMS [USE91e]. Sedo [DF91]. See [Swe07]. Seek [Law98, Fra90a]. Seeks [Ano93o]. Seen [Arn91, Web95]. Segmentation [HR79a]. seinen [KGG00]. Seismic [FY18]. SEL [Bur76, BR77, FT83, Rem73]. SEL-86 [BR77]. SEL/86 [Bur76]. Selected [Gay90]. Selection [Maj77]. Self [GG99, HK89, Jae89a, LPG^{+}18, WB06, Env97a, Env97b, HBG^{+}06a]. Self-adapting [GG99]. Self-Adaptive [HK89]. Self-paced [Jae89a]. self-repairing [HBG^{+}06a]. Self-Reporting [LPG^{+}18]. Self-service [WB06]. self-study [Env97a, Env97b]. SELINUX [McC05, HRS^{+}10, MMC06, RKSV20, SFE10]. Sell [Ano96d]. Sema [Kor89]. semantic [Hor90b]. Semantics [Set79a, Set79b, Set79c, Set81, Set82, Set83, WDG^{+}06, YHR92, dVd93]. Semantics-Directed [Set81, Set82, Set83]. semantics-preserving [YHR92]. Semaphore [Kor89]. Semaphores [Bra76a, DF95, Gai84]. Semi [PF84a, PF84b, Car89a]. semi-complete [Car89a]. Semi-Large [PF84a, PF84b].

Semiconductor [Gar97, SJ83, Ano02c]. Seminar [Kom02, RGRG05, Ano95a, IEE91b]. Send [Che79a]. Send-only [Che79a]. Sending [For00]. Sendmail [AV95, CA97, Cos97, CA03, CSAJ04, Hun03, CAR93b]. Sensitivity [SD87]. sensor [WCHS93]. Separate [EST86, Mul87, vMM88]. separately [Inc83]. September [ACM93, Ano88b, BSK87, Tob80, Uni96a, Uni96b, Uni98b, USE88f, USE89d, USE91b, USE92e, USE94d, USE96a, McK97b]. Sequence [JRL77, JR81]. sequences [CH96, WN15]. Sequential [AU3 , AU74]. Serial [AW89, Cam87, Gol98, Wel94a, Whi89]. Series [GH75, GH4 , HS83, Lin84, McI90, WLHC84, Hib96, Wil83a, Wil83b, KAH83]. Series/1 [Wil83a, Wil83b]. Serious [Sil84, Kra07]. SERIX [Wil83a, Wil83b]. Server [AKDN90, Ano93-34, Ano93-39, Ano00d, Ano00e, Ano00f, Ano02b, Blu00b, BW94a, BW94b, Bra04, CN97, Des96, Don96c, Fel96a, FKM95, Gil96b, HK79, HR78a, Hon06, Hon08, HM90a, Hun00a, Hun00b, Hun01, IF92, Ket99, Kro99, Lei95a, Mad94, Moh99, RR03a, Ric94a, R^{+}02, Sha94, SWL^{+}01, Sim95, SA95, TECB03, VRFJ97b, Yue01, Ano96f, Aul01, Bal96, Bas91, BsswLB05, CS93, CS96b, Eck95, EBS08, FWS^{+}03, Fli03, Gil96a, Har01c, Jan03, MRM06, Ros94b, SK01, Smi01, TR99, VJ06, WLS^{+}01, AJ05, D^{+}04, IBM09b, OH05, R^{+}02, SK10a]. Server-Lösungen [Bra04]. Serveradministration [Luc03b]. Servers [Ano96e, Ano03f, Bau02, HHLS90, Hur94, McC90, Nor02, Ric94b, SJ95a, SJ95b, SKB00, Ano95e, Ano97b, BBHL08, CDV^{+}94, Gil96a, GK97, Hun99, Hun02a, KPT^{+}02, Len97, Mly09, Pon05a, Sim96c, TB05, Won98, Pon05b, Chr98]. Service [AHPS79, CL79, Coc01b, Coh90, Hes94, KLK^{+}02, MH78, NR94, Vol75, CL06, Fac02, KWDB06, Neu00, Van05, Wei97, WB06, Cha79]. Services [Ano02b, BH01, BW94a, BW94b, CW00, GJLB77, Hof79, Jen00, Lee89, Mer01, MKL^{+}95, Roe02, Shi77, Wid00, X/O88b, d’A92, dC93, Ano03c, Ano03e, HSG03, Mat93, Wai95, BH00, Liu11]. serving [Bou97]. session [Ano19d]. Set [Ano93k, FN83, McK98, Str80b, Stu78c, Sun96, Che97]. Sethi [Che75a]. Sets [ALxx]. Setting [HRW77, HW78, Len94, Lyc76d, Lyc77f, Lyc77g, Pas88, Pug92, Siv01, Wil90, FM02, Hag83]. Setting-Up [Lyc76d, Lyc77g]. SEU [Nis01c]. Seven [Mee95, Hun99, Tox07]. Seventh [ACM75, IEE91b, USE93b]. several [Mud97]. SGI [Ano03f, Var96a]. SGML [Kan93]. sh [Ree88]. Shōkai [RB06c]. Shall [Sta04]. SHARE [BGMR91, SAS94a, SAS95]. Shared [Arn86, BS90, Bur76, De 76, HCF^{+}94, LTW77, Mil91, MF99, Mul79c, PPTHxx, Ber98, CR92, GMTW08, Har80, TTG^{+}87]. Shared-Memory [BS90, PPTHxx, CR92, GMTW08]. Sharing [AO86, BW88a, BW88b, BW88c, Bel79c, Bor98, Bou78a, CLFL94, CK78a, Cro78a, DHP77, DHM78, Fra78, JR78, JL78, KLO78, Lio79b, LMT78, Lyc75, LB78c, Lyc78e, LC78c, MWCR90, MWCR91, MPT78, MCM78b, NP78, Pan94, Pek78, Rit71, RT74, Rit7 , RT78, Rit78, RJLK78, Rit9 , RT79a, RT83a, RT01, RT21, RT79b, Rov78, Swa74, RT5 , Tho78, TG79, WSS78, WS78, FL80, Koh99, Lio79c, Ney83a, Ney83b, RT73a, RT73b, Rit80, Rit84a, Rit84b, RW81, SS95b]. SHARP [BE00]. Shcvt [Sny78]. Sheet [Gla77b, Ros77b]. shelf [Ano91c, PH91]. Shell [AA86, Ari92, Art86, Art90, Art94, Blu11, BK89, Bou7 , Bou78b, Bou78a, Bur04, Com94a, Car99, Col94, Col89, Col17, Csa77, Dav97, Die11, Duf90b, Dufxx, Fio04, FG03, Hol87c, Int93, Joh05, KW90, KW03, Kol77, Lio78, Mas77b, May01, Mic08, NR95, NR98, Olc91, Olc92, QH01, Rei03, RB05a, RB06a, Ros91, Ros93a, Ros94a, RR02b, Sch97, Skj82, Smi06a, Smi06c, Sny78, Sou94, Tan00, TR94, Val92b, Vic98, Wit79, Ame93, Ano00g, AB97, Bli96, Blu08, Dwi04, Ell80, FJWA05, GKV21, KW85, MM86, Mic03, NR05, OP01, Olc01, Pet09, RB05b, Sch90, See08, Shu83, Ano85, BS01, Bla02, Bla04, May01, RB08]. Shell-Programmierung [RB06a, May01]. shellcoder [AK^{+}07]. Shells [Lut99, Qui99, Qui02, Qui97, Qui00, Qui05]. Shepherd [Sau88]. Sheru [MO01]. Shi [PWC00]. shieru [RB06c]. Shift [SF95]. Shifts [Cas94]. Ships [Ano01, Ano02b, Ano03e]. Shoot [Ano93-41]. Shoot-Out [Ano93-41]. shops [Sim96b]. short [IEE91b, Mat93]. Shortchanging [Bar01b]. Should [Pik88, Ale96b, HBB13, Sim96b, Tay96, War04, War15, Per93]. shows [Ano96e]. si [Feu85]. SIAM [ACM97]. sichere [Lin02a]. sicheren [DLvG07]. Sicherheit [Bra04]. SICPPR [Ell79a]. Side [Mar78e]. Side-by-side [Mar78e]. Siever [Sch04]. SIGACT [Fis73]. SIGACT-SIGPLAN [Fis73]. SIGGRAPH [ACM88]. SIGMOD [DFG99]. Signal [Ano95h, Bai99, KL73, Lyo76a, Rem97, She03]. Signals [Wei88]. SIGNUM [ACM79]. SIGPLAN [ACM82, Fis73]. Silicon [Ano02c, SZ95]. Silk [SGWP02]. SIM [Che76a, Che76b]. SIM48 [Mul79a]. SIMD [KI99]. Similar [IK83]. Similar-String [IK83]. similarity [SWF^{+}22]. Simple [Bro0 , Bro81a, Bro81b, Cox07, Hol97, JP17, Law98, MF83, RRW84, Rug83, War88, YSF89, CKM06, CKM14, HU73, Lev98, PCD92]. Simpler [PC90]. simplified [ZL10]. Simply [CTJ99, Ano95d]. Simson [Ano96b]. Simulate [Ale96a]. Simulating [Bad98, Che76b, HO93, LC90]. Simulation [LWL05, Lyc71, MRK90, PJ89a, PJ89b, RC83, CGMS89, LWL06]. Simulator [Hof77, Mic79b, Mul79a, Reg78, RDK75, Rob06b, SG95]. simulators [Ale96a].

Singapore [Ano06a]. Singh [Ano94a]. Single [Bot84, CLFL94, Lyc72, MRxxd, Sch97, Wal95a, DJM86, Var96b, Coc03a, Jos97, The02]. Single-Address-Space [CLFL94]. Single-level [MRxxd]. single-user [DJM86]. SINIX [BBJ86]. sins [Tox07]. SIRIO [CBV93]. sistema [RÓ01, VD01]. Site [HR91, Mac86, Mac90]. Sixteen [Strxx]. sixth [USE92a, LT80]. Size [MAP^{+}22, THB06]. Sizes [Pug92]. Sizing [DeV00]. skills [Gre96, Gri14, HTU96]. skryptów [RB06b]. Slackware [LCJH05, Wal95d]. Slashed [Ano93w]. slave [DJM86]. Sleep [BLH20, PPTHxx]. Sleep-in-atomic-context [BLH20]. Slide [DS77]. slightly [Mon95]. Slippery [SRTH15]. Sloan [Nag05]. Slope [SRTH15]. Slow [CPJ^{+}98, Cox07, Gra04]. SM80 [Reg78]. SMAL2 [Far78]. Small [Lat02, Lyc71, Mor78b, PW79, SH99a, SAN08, Tin79b, DH02]. Smalltalk [Cox83]. smart [Ale96b]. Smartcard [USE99c]. Smartcards [IHR99]. SmartList [Sch98]. smartphones [Jen19]. SMB [Lei00, Len97]. smbfs [Len97]. Smileys [SD93]. SMP [Ger02, McC95]. SMSCRIPT [QBdR87, BQd86]. SNOBOL [FGP64]. SNOBOL4 [GPP68, GPP71]. Snocone [Koe90]. SNORT [Reh03, BF03, SK10a]. Snow [SGWP02]. SoC [VTS^{+}23]. Society [CPJ^{+}98]. Socket [But02a, Gay00b, CS93, CS96b, Wal01c]. Sockets [Joh96a, Mad94, Ste98, SFR04, ZG96]. SoCs [PGW^{+}20]. Soft [Ano92k, CN97, GMSP00, MG99, MR10, MR14b, SGM^{+}00]. Soft-updates [MR10, MR14b]. Softshell [Kat82a, Kat82b]. Software [ACM79, ACM88, Ano93-32, Ano94j, Ano98d, Anoxxb, Ano00e, Ano01, Lin02b, Lin02c, Ano02c, Ano24, Arc93, BBH18, BAKW92, Bon79b, BG97, Bra98, CKM85, CPJ^{+}98, Coc01b, CW15a, CW15b, Ded09, Den79, DP95, D^{+}76, DuB93, Ell77, EM79, FES^{+}99, Fra83, Gin83, GES98, HL95, Haw86, Haw87, Hol02, HKM^{+}85, Hor90a, Hun92b, JL98, Jam09, Jor04, Kap79, KM82, KP99, Koe84, Kri95, Kro99, Lau79a, Lea90, Leh95, LPG^{+}18, LC77, LC78a, LO97, Lut99, Lyc76b, Lyc77b, Lyc77d, Lyc77e, Mac83, MWCR90, MWCR91, The92c, The92a, The92b, The93, Mor79a, Mor85, NM83a, NM83b, NR03, Nee91, O’D83c, PBS98, PF84a, PF84b, Per87, PLR85, PLR00, Pom04, Rit84c, Rit85, Rit12, Rog79, Ros77a, Sof83, Sof84, San98, Sch79b, SC78, Sel94, Smi77, Ste01f]. Software [Sti83, Uni96a, Uni96b, Uni97, Uni98a, Uni98b, Usr82, Ass88c, Wam83a, Wam83b, Wat87, YK15, Ano93r, Ano95g, Ano99d, Ano03d, Apo87, BW96, BJM^{+}22, CH96, Cus04, Don04, Eck91, FRRJ07, Gai83, Ger03, Goe07, Han77, HGS99, Hoo05, IEE94c, Inc83, IF10, JJ91o, Kop05, LL93a, MBC^{+}86, OS06, Ric77, Roo79b, Sal08, Sch95a, SS95a, Sch79c, SB08, Spe96, War85a, Way00, WMK^{+}07, WT91, ZB01, Lin02b, Ano02b, Ano03d, Ano03f, Kro99, PH87, Ano91c]. Software/Systems [Ano94j]. Softwaretools [Jor04]. SOHO [Blu00b]. Solaris [Ano00a, Bor14, Chr98, DF00, Ern11, HS04, JF12, OH93a, OH93b, San07, VSC^{+}10, WF03, Ano96f, Ano01, Ano06b, BMS95, Bia99, BH00, BH01, BDD^{+}06, Bus02, Cal03, Cal06, Cal08, Cal09, Coc94, Coc98, DM07, DeV00, Doc03, Gal05, GGM93, Gov08, Gra95, GM11, Gre00, Gre02, HSM03, HG95a, HA93, How01, HN01, JDB09, JW99, KM95, KLM92, Kuo99, Led94, Lei95b, Loz05b, MM01a, MGMMI10, MM07, MMG07, MM06b, MGK^{+}03, Mil16, NBB^{+}02, OK94, Pac02, PDL04, Pew93, Rea11, Rob99b, Rob99c, SG06, Siv01, Siv03, Sob99, Sun93a, Sun93b, Sun93c, Sun93d, Sun95, Tee05, Tut93, TS97, Vea10, Wat10, Wat02, WN02, Wil98, Win93a, Win93b, Win99, Win00, Win01, Win02b, Win02a, Won98]. Solo [Jos97]. Solution [Ano98d, KL73, Sch76a, GMSP00, Joh05, LWL06, Tot99]. Solutions [Ano00e, Blu7 a, Blu78a, Bra04, Chu03, Coo09b, Coo09d, ER01, Gig06, GH75, GH4 , HSM03, Jae86, Koh06, Lin65, PKP05, Sat90a, TG89, Wey01, YK15, Blo99, HLK01, MAT08, MB05, Ano02c]. Solved [Bar00b, MZH22, Zim98]. Solver [Sch75]. Solving [Hal4 , Hal75, KS85, Kau90, Kau93d, LLM87, Sch76b, Sim96c]. Some [GJ75, Hay77, Les79a, PLBL92, Ral80, Tuo83, War90, Sim96b, Spe96]. sophisticated [Lub90, Lum13]. Sort [BM93, Lin79a, Zuc77a]. Sorting [SB97, BW94c, Ano93-30]. SOSP [ACM03]. Sound [DFV91, Phi00, Str94]. Source [Lin02c, BS79b, Bro19, CPJ^{+}98, Coc03a, DS00, Fil85, FT09, Gla99, Gla78, Gra98a, Gra98b, Gra98f, Gra98d, Gra98c, Gra99b, Gra99c, Gra99d, Gra99a, Gra00a, Gra00b, Gra00c, Gra01a, ILG19, Leh03, LPG^{+}18, Lew99a, Lew99b, Lio96, MSSvK08, PGW^{+}20, Ray99b, San98, SK04, Sor01, SPS^{+}02, Tan87a, TS87, TTB09, AC11, And01, AAB^{+}05, BP09, CM06, Cha11, FRRJ07, Fog99, Fra19, Ger03, Goe07, Gra88, HP01, JP09, JB99, LW03, Les88a, MM00, May84, McC05, Moo01b, PS^{+}09, Ray01b, SNM22, SML06, Tan87b, YSC^{+}06, ZK05, ZB01, Lin02b, Lin02c, Lin02a, Moo01a, NR03, Ray99c]. Source-Level [SPS^{+}02]. sourcebook [Lev02]. Sourced [Coc01c]. Sources [Vol97, Wat83]. sous [Bla02, Bla04, Dup05]. Soviet [KL92]. Space [CLFL94, Coc03b, FE77, Gar97, GT78b, KS85, Nis01a, Nis01b, Nis01c, Nis03, Sch76a, Tin78c, FH89]. Spafford [Ano96b]. spam [Zdz05]. Spanish [RÓ01, VD01]. Spanning [McI69b]. SPARC [Cat91, Coc94, Coc98, Wil98]. Sparc64 [ST03]. Sparse [Kau7 , Kau79a]. spatial [DD05]. SpatialStats [KVCS98]. Spawn [Fro90]. Speaking [Bar00c, Don97a]. Speaks [JK86]. Spec [Bar00b]. Special [BN01, Ken90, Kuo98b, NB84, SRBB99, TGB95, TG96, IEE03]. specialization [CDD^{+}05]. species [DD05]. Specific [BBS95, OBSB23, PF84a, PF84b]. Specification [AS90, Can80, Coc03a, Coc03c, Coc03b, Erm79, Hei95, ISO05, MS84, Ng77, RAMP94, SA93b, Vog79b, Wal95a, ACS98, HRS^{+}10, JJ91o, Jos97, Van90, The02]. Specification/User [Ng77]. specifications [LL95]. Specifying [HT92, HT94a, HT94b, McI69a]. Specs [Nis01a]. Speech [McI73b, Tuo83, McI74]. Speed [ATS94, Ban82, Fed77, Har80, MZA96, USE94b, AOAH14]. Speeding [Hau94]. Spell [Bar00c]. Spelling [McI82, MC75a]. Spent [CKT80]. Spider [Fra74b, Fra74a]. Spin [Hol90b]. Splicing [SMD^{+}94]. splines [Sch76b]. Spooler [Gli90, Lau79a]. Spotlighting [Ano06a]. SPP [Sel94]. spread [Var96b]. Spring [Eur91, Ass88a, USE88c, Ass89a, Ass89c, USE93a, NR94]. Springs [IEE94b, USE90b, USE91a]. SPS [Sha79b]. SPSS [Nor93a, Nor93b, Nor93c]. Spy [Bar01b, Sto89, Sto90]. SQL [Arn98, Fre98, Tar89]. Square [Bal96]. Squeeze [Sta04]. squeezing [Coo85]. Squinting [McI90]. SSC [Ano98c]. SSH [DLvG07, BS01, Dwi04, Höf01, Luc12, Luc18b]. SSH-Buch [DLvG07]. SSH-Connected [Höf01]. SSL [DC17]. SSL/TLS [DC17]. Stabilität [Bra04]. Stability [Bra04]. Stable [Any85, HLT^{+}21]. Stack [Strxx, SWF^{+}22]. staging [Cou17]. Stand [Eis79a, Mul79c]. Stand-Alone [Eis79a]. Standard [Ame89, Abl93, Ano77, Arc93, BBH18, Bec06, Bec07, Coc01b, Dul94, Dzo09, Gre82a, Gre82b, Hal00, Hel03, HA84, IEE86, IEE88, ISO05, Jae89a, Jae91, Jae94a, Jae94b, Jos99, Jos12, Kar06, Koe79b, Len94, Mor75, Ree06b, RAMP94, Tho92, Til84, Vog74, Vog79a, Yeo05, Zlo91, Ame93, Ame94, Ano93z, Ano94i, Ano19b, Ban88, Gil91, GK97, Gla92, Gre81, IEE91c, IEE92a, IEE92b, Lew91, Mat85a, Sak97, Sea09, Sea14, Ano94k, BBD91, LK00, MT93, Pla92]. Standardization [Bar01b, Eck91, Vog78b, RS14]. Standardizing [Lud78b]. Standards [Ano92m, Ano93-28, BJ93, C^{+}78a, Cra83, Dun91a, EM79, Gar99, Gar90, Hae89, Hae90a, Hae90b, Hae90c, Hae90d, Hae90e, Hae91a, Hae91b, Isa86, IJ98, Law98, McC88, PKP05, Pug88a, QW93, Sto93c, Sto93a, Sto93b, Sto93d, Sto94b, Sto94c, Wal91a, Wal91b, Wal91c, Wal92e, Wal92f, Wal92a, Wal92b, Wal92c, Wal92d, Wal93a, Wal93b, Ano90b, DDLL94, Fra90a, Kut02, LL93a, LL93b, Mee95, SG97b, Woo94]. Stanford [BBNS80, Gor81, Rei86]. Star [KDP83, HBH18, HWBH19, FGP^{+}91]. Starch [Ano16b, Jen05]. StarOffice [Hab01, JH00, Lew01, Mea99, WT99]. start [Ano93m, Sau17]. Started [Wt95b, AS93b, BA96, Jen19, Ric13, Wel94b, WBS01]. Starter [Dul06, Tar80, Gay99]. Starting [Lo79]. Startup [PS07]. STAT [Ano00e, Feu78, Per84, PH86]. State [Ano93-29, FST91, KDP83, Lon77, Roe02, TWM86a, TWM86b, HK05]. state-of-the-art [HK05]. State-wide [TWM86a, TWM86b]. Statements [Set79b]. static [PRC^{+}23]. Statically [CCVLW23]. Statistical [Ano00e, Cha69, MCM78b, MN69, Pir88, WN15, Zdz05]. Statistics [GJ75, Nor93a, Nor93c, SPS93a, SPS93c]. Status [Che79c, EM00, Gor81, LZHS11, Rea91, SST90, Fra5 , Fra76]. STD [Cer82, Cer83]. Stefan [Kad95b, Kad95a]. stegomalware [CMR^{+}21]. Stellix [TG88]. Step [And06, GM02, Lut99, Dun00, JJ92c, Mil01]. Step-by-Step [GM02, Dun00]. Stephen [Sch04, Teo13]. Steps [Hil98, Sta05b, OH93a, Ric93, Sim97]. Steuerung [GGZ02]. Stevens [Mad94, Teo13]. Sticky [Ban82]. still [Mud97, Sim97]. Stockroom [Ros79]. Stolen [SW12]. stop [Coo03]. Storage [ACM04, Ano93-40, Any85, Gig06, HP89a, HP89b, Kiv84, LR90, McK03, WB92, dC93, CMR^{+}90, McK96a]. Store [Ell75, Ano96e]. Stored [Rob75]. storia [Cor05b]. stories [Cor05b, McC98, Ano93-42]. Story [Moh85, TD01, Moo01b, Wal01b, Sim96a]. Straightforward [PBL86]. Strange [Rit02, Too11, Tuc98, Tay96]. strap [Reg11]. Strategies [ZB84, Dwi04, PBWB84]. Strategy [Coc03a, HR94, Cus04]. Stream [KSU94, Mor78c, Rit84d, Rit90, ZG96, Lev79]. Streamlined [Fla94]. STREAMS [KM87, GS89, Prexx]. Street [SWG92]. Strength [Lav04, Fli03, Ran04, Ano93j]. stress [Ano94i]. StrikeIron [Ano03c]. String [AC75, Car77, FGP64, Fre78, Hoa77, HS90, HS91b, IK83, NT05]. Strings [SB97, SM77b]. Stripped [JJ91p]. Stripped-Down [JJ91p]. Strobel [Kad95b, Kad95a]. Stronger [Ano92b]. Struct [Bak75]. Structural [Pik06]. Structure [Che78, Coh79, Far79, Höf01, Roc82, BJK87]. Structured [LB75b, LB75a, RO92, SBMS93, Gat01, JHT^{+}07, KAS^{+}06, MS92]. Structures [Bak75, BED^{+}85, Far79, Lyc78a, Van88, BI95]. Structuring [Bak76, Bak77]. Student [Hah93, Mac97b]. Study [ACS79, BS88, CKM85, CR02, Cha79, Doc03, Fel79e, Gin83, Gre02, Ham87, H^{+}79, Hei95, Kul06, LADADL14, MQ09, MCM06, PQM^{+}21, RB08, RÓ01, SG06, SA21, WLL96, Wei92, AOAH14, Ano90b, BTS09, Cab83, DI89, DIN05, Env97a, Env97b, FAB^{+}07, FFPC22, Ger03, IF10, Jan04, Kee05, KWDB06, Kor06, LR89c, MW08, MFS90, MCM07, Raf79, Reh05, Ros78, SG04, dVd93]. Studying [Coc03c, HBH18, HWBH19]. Stuttgart [BSK87]. Style [C^{+}78a, Gla77b, Ope91b, Pik83b, Str80b, ER86, Oua92, CV80]. Styles [Cop92]. Subcommittee [Vog78b, Nel79]. submission [Mac97b]. Subprogram [Lio78]. Subresultant [Bro7 , Bro78]. Subroutine [Dwy79, FHS77, FHS78b]. Subroutines [BFKS93, Mor75, Skj82, Uni82]. Subset [Hal69, Ano94h]. Subsumes [LBN14]. Subsystem [AO86, Fre91, Hoo78, KM83, MK85a, Wal97b, Wal98, Mit84, IBM05]. subsystems [WMK^{+}07]. Subversion [PCSF08, Mas05]. Success [Ano93-42, Ray99b, Gil96a]. Successful [Osm96]. Successor [Hum90c]. Such [CPJ^{+}98]. Sudo [Luc13b]. SUID [Kra88b]. SUIT [PCD92]. Suite [Kro99, LKY93, YLK93, Ano03d, Ano02b, HD98]. Suited [Min82a, Min82b]. sukuriputo [RB06c]. Summary [Gay90, Lyc76c, Lyc77a, McI5 , Okr78, SM77b, Set82, ZM08]. Summer [Ano94m, BBK13, CKT80, Sof83, Sof84, USE82, USE84a, USE85d, USE85a, USE86c, USE86a, USE87d, Ass88b, Ass88d, USE89e, USE89a, USE90c, USE91c, USE92b, USE93c, USE94e]. Summit [ZM08]. Sun [Ano94e, VBD00, BBNS80, BBP82, Ano96e, BS91, Bus02, CR02, Coc94, Coc98, Cou85b, Cov06, Doc03, ER01, Gre02, Kle86, Kro99, OH97, OH93b, Rus91, SG06, San07, SKB00, VBD00, VBD01]. Sun/Microsoft [OH97]. SunOS [Kuo99, HA90, KLM92, MF99]. SunScreen [WC98]. SunSoft [OH93a, OK94]. Super [Ano93n, Bar00c, EEK^{+}04, Gra01b, Zwi92, Fox85, ST03, War04, Jon88]. super-user [War04]. Super-Users [Zwi92, Fox85, Jon88]. superbible [WS96]. Supercomputer [And88a, Bar01b, Chr90, TG88, Car89a]. Supercomputers [Ano88b, Coc03a, LM88a, LM88b, USE88f]. Supercomputing [ACM89, ACM04, MK93, Bis89, MK92a, MK92b]. superimposed [Rob8 , Rob79b]. Superminis [Bot84]. Superserver [Ano93-41]. Superuser [CCM^{+}90, HCC^{+}87, War15]. Supervisor [Bur78, GT78b, Lyc78b]. Supplement [Fla91, Nye95, Ame91]. Supplementary [Com94d, Com94g, Ker79g, X/O89a]. suppliers [Siv90]. Supply [RS78]. Support [Bal94, BDD^{+}93, Bra04, CCH^{+}77, Ell77, FES^{+}99, Gar99, Gay93, Gra79c, HCF^{+}94, HKM^{+}85, Gur81, Kos87, Lam83, Lei96c, Lei96e, Lei96d, Les77b, Mar78d, MH88, MS89, McK03, MK04, MZA96, Req85, RT83b, Rov78, Sta04, Vog76b, Whi75b, YR93, Ano03e, BS05, BV06, Don04, GW98, HK82b, JJ91q, KD94, Mat93, MS87, PBM^{+}22, SAS91a, SAS93a, SJP^{+}09, ?, ?]. Supported [Coh79]. Supporting [BCC^{+}91, BSW78, DRN^{+}23, LS92, Ran96, ER86, GMTW08, Neu00]. Supports [Ano93-43, SML06]. Surfaces [FKM95]. Surgeons [Coc03a]. Survey [Gar97, GD18, HS82, NCNS82, RMF19, Rib89, Won90]. Survival [C^{+}03, Smi06a, Smi06c, Bam98, Par00]. survive [Les88a, Ros94b]. SUS [Gra01b]. SUSE [DWV06, Neg06a, Neg07a, Bal00, Bro06, EDK00, SK01]. Suspend [Whi75a]. SVID [Fis86c]. SVR [Kuo99]. SVR4 [Bra95, Cur96, Cur97, NHNW94, Rob99b, Rob99c, WLL96]. SVR4.2 [Nad92]. Swan [Swa00b]. swap [Kob99]. Swarm [Coc03b]. SWIFI [SB02]. switch [LCH^{+}81]. Switcher [Ral77]. Switches [Ral80]. Switching [All78, Gra03, Hoa77, PY84a, PY84b, Ral75b, Ral76b, Ron05a]. sword [RR95a]. Sybase [Ano03e, Ano03g]. Sybex [Koh06]. Sydney [Tob80]. Symbolic [Kat79, KE85, BH8 , BH79, Fel4 , Fel75, Hal4 , Hal75, Sak97, Tun05]. Symbols [Scr76]. Symmetric [Sch94b, Sch95b]. Symposium [ACM75, ACM82, ACM88, ACM97, ACM03, DF93, Far95, IEE03, IEE04, McK93a, USE92e, USE93d, USE93e, USE93g, USE94b, USE94c, USE95c, USE95a, USE98a, USE99b, USE99a, Wal94b, Fis73, Ric77, Tob80, USE91e]. Synchronization [Bv92, Dun89, Gau98, Ger02, GZ89, Vog97].

Synchronized [SW91b]. Synchronous [Lau79b, Lon78, MG99, PK93a, PK93b]. Syncsort [Ano93-30]. Syngress [Ano16a]. Synopsis [DH77]. syntactic [And80, FJ81]. Syntax [COV79, HA84, And80, Ube76]. Synthesis [Bee86b, DFV91, Lon77, Pal96]. Synthetic [Ber76, McI73b, McI74]. SysAdmin [Bee96a, Bee96b, Ann00]. Syslogd [SW91a]. SYSS [Kau93e]. SYSTANT [Boy84]. System [Ame93, ?, ?, ?, Ame90b, Ame90a, Ame91, ?, ?, ?, Ado93, Aho84, Alb77a, AISS98, Ama88, ACK89, Ano82, Ano94d, Ano94f, Ano95j, XWS95, Ano98b, Ano00a, Ano00e, Ano01, Ano03f, Any85, AM86, AR78, Arn86, AHPS79, AS90, Aul96, BJP04, BB84, Bac86a, BBJ86, BG88, Bac86b, Bad98, BH86, BMR94, BFGK89, BS88, Bar88a, Bar88b, BBNS80, Bas87, BL75, LB75b, BK85, Bec84, BMS95, Bel79c, BK90b, BK91c, BK91b, BBS95, BP88b, BP88c, BP88a, BGMR91, BE00, BBM85, Bla83a, Bla83b, BEHW86a, BEHW86b, BRD77, Bok95, BS79b, Bot84, Bou78a, Bou82, Bou86, BCC^{+}91, BB93, Bro89a, BDD^{+}06, Bro89b, Bur76, BR77, Bur79a, BK84, BP84, Com94a, Com94e, Cal93, Cam04, CBB^{+}91, Can80, CGS96, CDD00, Car82, CCDP01, CHL16, CQ83a, CQ83b, CDD^{+}05]. System [Cha79, CLFL94, CB83, Che87b, CL90b, Chr83a, Chr88, CR94, Chr83b, Chr90, CN97, Chu98, Cla79, Cof90, CK78a, CFA85, CCH^{+}77, CAG89a, CAG89b, CBV93, Cou85b, Cra83, CBL96, Cro78a, Cuc90, Cur89, CKDS90, Cur90, Cur91, Cur92, Dat88, Dav89, DM07, Daw07, Dea78, Den79, DKM96, Dil02, Doc03, DHM78, DHS78, DB95, Duf89c, Duf90a, DS00, Dun89, Ell75, Ell79b, EST86, Eng76, Eps92, Ern11, Eva97, EM79, Eve78, Far90, FG91, Fed83, Fel84, FFH^{+}86, FHS91, Fil79, Fil85, Fol77, Fra74a, Fra74b, FE77, Fra78, Fri91, Fri03, Gag01, GMM90, Gau81, Geh86, Geh87a, G^{+}06a, GVK^{+}03, GBH86, GT92c, GT78a, Gla78, GT79, GC94a, GC94b, Gra89b, GM90, GP88, Gra99c, Gre19, Gre02, GGZ02, GZ84a, GZ84b, HCF^{+}94, Hai78b, Hai78c, HK83, HP98, HK95]. System [HQZ^{+}87, Har92b, Har94b, Har80, HS87b, HRO82, HLW84, Hei95, HA84, HP89a, HP89b, HSM03, HBG^{+}06b, Her93, Hil93, uRH95, uRH96, HH86, Hiu96, HK86, Hof79, HKM^{+}85, Hoo78, Hoo79, HG95b, HL85c, How01, HK89, Hum88a, HM90b, Hun00c, HH91, IBM88, Int90, Int93, IEE86, IEE88, Ima79, ILG19, IBC87, IF92, Jac77b, JDB09, JH83, JF12, JR78, JL78, KNI88, Kap92, Kap99, Kar83, KDP83, KC75, KLO78, KC79, KM82, Ker89, KN93, Kil90, KK89b, Kiv84, Kle86, Koc87, KM87, KC98, KKF^{+}84, Kow79, Kow90, Kra88a, Kra78b, Kro99, Kuo99, Lam83, LM88a, LM88b, Lan84, LEG88, Lea90, LLM87, Lee89, LMKQ89, LMKQ90, LA90, LTW77, Les76b, Les77b, Les78, Les79a, Les88b, LH01a, Lim05, Lio79b, Lio79c, Lo79, Loi79, LR78, LH84a, LH84b, Lou90a, Loz05b]. System [LADADL14, LMT78, Lud79b, LSC^{+}88, Lyc71, Lyc73, LB75a, Lyc76a, Lyc76b, Lyc77c, SL76, Lyc78a, LB78a, LB78b, LC78b, LB78c, Lyc78e, LC78c, MDAD^{+}14, Mac97a, MC85a, MC85b, MF83, Mar75, MH88, Mar84, MTY01, May78, MM83a, MM90a, MM83b, MM90b, MPT78, MS89, MJLF84, MBKQ96, MNN05, McK06, McK12b, MNNW15, MCM78b, McN92, McS78, McS79, Mer82, Mer01, Mic79a, Mic08, Mik90, Mil77, Mil84a, MF99, Min82a, Min82b, MMTW88, MS84, Mue79, Mui95, MvT^{+}90, Mul79c, MK85b, MK90, ML02a, Nat94, Nad92, NP78, NM83a, NM83b, Nee90, NSS89, NSSH95, NSS^{+}01, Ney83a, Ney83b, Nor93b, NS94b, Nor91, Nye95, O^{+}98, Ass99a, Out82, Pad93, Par78, P^{+}08, Pas88, PMD88, Pat83, PF84a, PF84b, Pek78, PDL04, PR76, PW77, PK84, Pik89, Pikxx, Pir88]. System [Pon98, Pon99b, Pon99a, Pon01, Pow84b, PM87, PR85b, PR85c, PR90, PPTTxxb, Pug90c, PY84a, PY84b, QSP85, QO90, Rag90, Rag93, Raj00, Ral75a, Ral76a, Ral76c, RRP^{+}08, Ran96, RC83, RW84, RM92, RMxx, Rib89, Ric96, Rit71, RT74, Rit5 , Rit7 , RT78, Rit78, RJLK78, Rit9 , RT79a, Rit80, RT83a, Rit90, RT01, RT21, RT79b, Roc77, Roc79, Roc82, Rod86, RKPP88, RRF90, Ros96, RO92, Rov78, RB82, SAS88b, SAS91b, SPS95, Sam87, SW82, SG06, San07, Sat90a, Sat93, Sat89, Sax85b, Sax85c, SR85a, SR85b, Sch90, Sch77, SC78, SG09, SA93b, SBMS93, Sha00a, Sha77, Sha81, SR77, She93a, SGG05, SGG09, SST90, Sil84, Sim93, Sim91, ST89a, ST89b, ST89c, Sob89, Sob91, Sob95, SA21, SD87, Sto80, Str90]. System [Stu78a, SCK86, Sun96, Sun77, STT86a, STT86b, Swa74, Swi77, Tag83, Tes86, TY82b, Tho85a, Tho85b, RT5 , Tho78, Tin78c, Tin79a, TG79, TSP93a, TSP94, TSP93b, TWM86a, TWM86b, Tur87, USE87c, USE90b, USE97a, USE98c, Uhl87, Usa78, Ves88, Vog6 , Vog77c, Vog79b, Wes83, WMP84, WMP92, Wam83a, Wam83b, War88, Wat87, WMK^{+}07, Weh83, Wei84, Wei98, Whi75a, WL97, Wil90, Win01, WSS78, WS78, WK85, X/O89c, Yos85, ZST^{+}07, ZDS85, Zie93, Zuc77b, Zuc83b, Zuc83c, dOCC23, vM88, Ame94, AT07, Ale96b, And01, Ano93m, Ano95d, Ano97b, Ano98a, Ano00g, Apo87, AAB^{+}05, AR07, BGW93, Ber86, BAC^{+}89, BBD23, Bin88, Bla86, Blu09, BHK^{+}84, Bor88, BS84, BCP93, Cal06, Cal08, Cal09, CE88, Cer96, Cha97, CHRG95, Che99, Che02, Che81, Che75b, Chr87, CK06b, Col02]. system [Col84b, CJBT92, Com12, CKM06, CKM14, DeR97, Don04, EEK^{+}04, EC89, Ewi86, FRAC85, Fed84, FRRJ07, FMM84, FL80, FSS99, FF88, FHR84, Gat01, GL88, Gia99, GJ08, GM84, Gre81, Gue19, Hal4 , Hal75, Han01, Har93, HGS99, HG95a, HBG^{+}06a, HK82b, Hor02, Hun88a, Hun88b, IEE92a, IEE92b, JHT^{+}07, Joh80, KIT93, KC78a, KWDB06, Ker10, KCTT06, Kom96, KAS^{+}06, Kra05, LM91, Lev79, LGJS91, Lit88, Liu11, LS87, Lov07, LCH^{+}81, MBC^{+}86, MM00, Max02, McI89, MS87, McK85b, McK12a, MK91, MM88, Mil78, Mor78a, Mue78b, Nat88, NSHW10, ODH^{+}85, Par90, Par00, Pay84a, Pay84b, PTS98, PTS01, dJPDD84, Pon96, Pon97b, Pon97c, Pon03, Pon05a, Pon05b, RRP^{+}07, Reh00, RT73a, RT73b, Rit84a, Rit84b, Rit84d, RÓ01, Ros00, RHP^{+}07, RRH^{+}08, SAS88a, SAS94b]. system [Sam89, Sch93c, SS95a, Siv90, SHN97, Son87, SS01, SS02, SS00, Tay03, TGF88, Thi92, TPS98, Too11, Tut83, Com79, Ube76, VD01, VSC^{+}10, VR00, Wei97, WJ82, Win93a, Win93b, Win00, WCHS93, YE84, YEB84, Ygg94, ?, ?, Ame91, ?, ?, Ano85, ACS98, AO86, BS79a, Bar91b, Bar94, Bec79, BS90, Can79, CMR^{+}90, CE97, Cra90, CGO92, DHP77, DKF94, DD90, Eyk86, FMM84, Fla91, Gel78, GKM90, Get91, GGL03, Gil93a, GRA89a, HLHR90, Hel90b, Hiu97, JR93, Joh92, Jon89, Kay79, Ken90, Kil96, KKF^{+}84, Kri91, Man91, Man93a, Man93b, MS87, McK07, McS79, Mik92, Mil84c, MP92, Nat90, New92, Nye90b, Nye90c, OD^{+}86, OQL88, PS94, Par93, PH98, PH99, Pos88, QO89, QO91, QO93a, QO93b, SGN88, SG90, SGFR90, SGFR92]. System [SG97a, SG97b, Smi91, Swe01b, Swi91, TP91, Top90b, Wes83, Wri90, YP92]. System- [Kuo99]. system-administration [FSS99]. system-driven [FRRJ07]. System-wide [CDD^{+}05]. System/370 [Eyk86, FMM84]. System/6000 [CMR^{+}90]. system/C [FF88]. System/Network [Stu78a]. System/U [KKF^{+}84]. Systemadministration [HS04]. Systeme [BBJ86, Jor04, WF03, Pri83, Sch94a, Jor04]. Systemen [Kuo99]. SystemMirror [Q^{+}11]. Systemoptimierung [Luc03b]. Systemprogrammierung [Her99]. Systems [ACM03, ACK89, Ano88a, Ano93l, Ano93x, Ano93-31, Ano93-33, Ano94j, Ano00e, AQ84, Bar86, Bas87, BS91, Bee96a, Bee96b, BD94, BFS89, Bri01, Buy99, Cal93, Cam04, CJL^{+}89, Car89b, Cas93, CW92, CEC^{+}96, Che87b, CG95b, Cla79, CS86, Coc03b, CN88, CJ92, CP86, CKDS90, Cur91, Cur96, Cur97, Dan83a, Dan83b, Dav93a, Doe11, Duf89a, Dufxx, Duf89b, Dun91b, Ear87, EGL86, ELS88, Eur91, Fie88, FES^{+}99, Gal00, Get86, GC94a, GC94b, Gra00c, GR76, GGZ02, Gwo78, Ham78, Hel92, HKS^{+}04, Hun92c, Isa86, Jor04, Jos95, JN88, KM83, Kur90, LB96a, LP89, LS90b, Lio79a, Lyc74b, MLRC88, Mat97, McC95, MKS^{+}89, McK03, Mil84b, Mil86, NCNS82, NZW^{+}02, Nor02, NL79, Ope94, Pea83, Pik88, Poe87, QW93, RWNA87, Ram88, RW86b, RW86a, RR03b, Roc82]. Systems [Rod90a, Rod90b, Rov78, SPS93b, SRTH15, Sch94b, Sch95b, SGM^{+}00, Smi77, Smi89, SD95, Sto88, STA86, TW97, Tan08, TOML04, Tin78b, TNY92a, TNY94a, TNY92b, TNY94b, Ass88a, Ass88b, Ass88c, USE88d, Ass89a, Ass89b, USE89c, Ass89c, USE89d, USE90a, USE91b, USE91e, USE92a, USE92c, USE93a, USE93b, USE94a, USE94d, USE96a, USE97b, USE98d, USE00b, USE01a, Wal95e, Wan77, Wan78, War84, Wei82, Wei92, Woe02, WH83, Won90, Woz82, Yag03, YT83, ZST^{+}07, dSVdLT02, Tho05, AEG^{+}91, AC11, Ano94i, Ano95d, Ano95i, Bac84, Bar01a, BK13, Bro97, BCE^{+}93, Bun76, Cha96, CG95a, CK06a, Com00, CP01, CP02, DH02, DDLL94, Don04, DuC94, EP87a, EP87b, GMSP00, Gue19, GMTW08, Had04, Hag83, HMK^{+}20, IEE91b, IEE94c, IBP03, Jal92, Jan06b, Jou86, KD94, Kha02, LL93b, Lev02, Luc05]. systems [Lun96, Mac97b, Mel87, MMTK94, Moh85, NIS92, NV06, O’G01, Olk03, PC01, Pon94, Pon05b, Raf79, RLB08, SAS88b, Sea14, Sin07, TH83a, TH83b, Tan87b, Tan01, THB06, TW06, Tan10, Tee05, Var96b, WLHC84, Wis03, YSC^{+}06, ZG96, Ano98d, Anoxxb, Kro99, You89, You90b]. systems-on-chip [Don04]. SYSTID [HY79]. SysView [CJ92]. Sytem [Ame94].

T. [Ral77]. T.2 [MS77, SM77b]. T.3 [SP77]. T.4 [SM77a]. T.8 [Scr76]. T.9 [DS77]. Tabellenbuch [BBJ86]. Table [GT78b, Lit80, Skl91, GHS82]. Tables [CL90a, Les76a, Les77c, Les79b, MM90a, MM90b, Roo79a, MD86, Whe97]. Tachyon [NSI^{+}92]. TAE [SS93]. tahoe [KTB^{+}88]. Tailoring [Bai87]. Take [Ano93k, Gra04]. take-it-slow [Gra04]. Takes [Ano94l, Gil96a]. Tale [Hum88b]. Tales [WF01]. Talisker [Smi02a]. Talk [NH99, NH00, Pac00]. Talking [Tuo83, WAC84]. Talks [Bey87]. Tamed [Uma96]. taming [Ran06]. Tandem [Usa78]. Tape [Bee85, Bee86a, Cal93, Coc01a, Ell75, Kri84a, Kri84b, Rob75, Fur85]. Tapes [Swa77]. TapeWare [Ano00e]. tar [Bee85, Bee86a]. Target [Ano96d, Bar00b, Ste01b, AC11, Ano95i]. Targeting [Gar86]. Task [FDMS79, PJ89a, PJ89b]. Tasking [BP84]. Tasks [RB06a, Mad97]. Taste [For00, WALK12]. Taudes [Wei95]. Taxing [Gar97]. taxonomy [DC18]. TB [Kem91]. TBL [Les79b, Okr78, CL90a, Les76a, Les77c]. TC [Die11]. TC-Shell [Die11]. TCB [SW88]. Tcl [SML06, DF00, Bla04, Ngu08b, Ame95, Bla02, HM98, Inf95, Lib94, Ous94, SSC95, Wal97a, Wal95c, Wel95, Wel97, Wel98, Wel00, WJH03, ZK05, Zim98]. Tcl-based [Lib94]. Tcl/Tk [Ame95, HM98, Inf95, Wal97a, Wal95c, Wel98, ZK05, Zim98]. TCMP [Ral76c]. TCP [Mad94, Str94, ALP^{+}21, Ale96a, All97, AOAH14, Ano80, CMQ93, CS91a, CS91b, CP01, CP02, DMT02, FFH^{+}86, Fer85, Hor98, Hun92a, Hun02b, JLSG84, KM86c, Mor85, PC01, San91, San94, Spu92, Ste94, Ste96, Sto95, TP91, WS95]. TCP/IP [Mad94, Str94, Ale96a, CMQ93, CS91a, CS91b, CP01, CP02, FFH^{+}86, Fer85, Hor98, Hun92a, Hun02b, JLSG84, Mor85, PC01, San91, San94, Ste94, Ste96, Sto95, TP91, WS95, Spu92]. tcsh [DuB95]. Teach [CS99, MB00, RBHB99, Til95, Art99, Bal99, CCB00, Gay99, Hsi01, LH00, Ray99a, SBR01, TA98, Tay03, Tay06, UT03, WT99, Wel99a]. Teaching [Gra00c, Gre96, Osb92, Shu06, CKM06, CKM14, NV06]. Teams [GS12, Ano03c]. Tech [Bar01b, Gay93, Sel94, Way00]. Technical [Cat91, Cou85b, FKN85, Hid04, Lee92, LK76a, LK76b, LK77, Mik90, OH05, SAS88a, USE95b, USE97c, USE00a, USE01b, Wal92g, FRAC85, Lar94, Mat93, MC75a, SAS94a, SAS95, USE96b, USE98b, USE98e]. Technique [Bon79a, MG99, PW77, PW79, DC17, Mac97b]. Techniques [ASU86, ALSU07, And17, BS85, GS01, GS02, HSM03, Lon77, Reh03, Smi06b, Stu01, TR94, Tho68, War85b, Bes05, Bun76, Fer04, Fri97, GL75, GL76, Ham07, Kra05, L^{+}11, MMG07, Mol14, MS95, PF05, SCJ96, Stu04, TSC02, ZL10]. TechnoFiles [Pog17]. Technologies [Ano00e, Mas88, Ano98d, Ano01, Ano03d, Ano03f]. Technology [Ano98c, Cro98a, Cus04, HN01, KJG^{+}98, Kro99, USE99c, ZB19, Ame93, Ame94, Ano95h, Cor05b, Fra5 , Fra76, Int90, Int93, IEE91c, IEE92a, IEE92b, MW08, Wal01b]. tecnologia [Cor05b]. tecnologica [ZB01]. Teixeira [Wei93]. TEK [Sto76]. TEKSTARE [Sto76]. tekunikku [Sas91c]. Telelogic [Ano01]. telemetry [FGP^{+}91]. Teletypewriter [Whi75b]. tell [Ano93r, FJPR93]. Template [BFKS93]. Template-Driven [BFKS93]. Templates [Xia97]. TEMPO [GZ84a, GZ84b, GZ89]. Ten [FK90, McK85a, PTS^{+}11, Woo94]. tenant [MKSB22]. Tennessee [IEE94a, USE91c]. Tenth [Uni97, USE96a]. Terabyte [McK03]. Terabytes [IEE02]. Teradata [Bal94]. term [CJBT92]. termcap [Flo90, SMO88]. Terminal [BBNS80, Bur79b, Cro98b, Fed78a, Fed78b, Flo90, Gol79, Lee89, MRxxd, MS90, Pas88, Pik84b, Wei88, Pik84a]. Terminals [Lau79b, Lyc75, Mar78a, Mue79, Ney83a, Ney83b, Whi75b, Mue78b]. Termination [Tin78d]. terminfo [SMO88]. Terms [Don97a, TBGE89, WC93, Lev96]. TERSE [PW79]. Test [Ano02b, BJ77, Bri05b, Coc03c, Cod89, FES^{+}99, Hor82, ITC93, Kro99, Lew99b, Lon78, Sch81, IEE91c, LL95]. Testament [PR96b, W^{+}95]. tested [FJPR93]. Testing [ALxx, AH11, AD89, Cro78b, DiM00, Dix82, Gra79a, Gra79b, Lut99, MM06a, MCM06, MZH22, Mue79, PW77, Rob79a, Rob82, SCK86, DDLL94, ITC93, LL93a, LL93b, MCM07, PRC^{+}23, WN15]. Tests [MSSvK08, Sch77]. Texas [USE85c, USE92b]. Text [BR87, BBG85, CPJ^{+}98, DO87, DF00, Fin91, Fra80, Hew90, Hor79, KMKO92, Ker79c, KV89, McM78a, MCM78b, McM90, Mil84b, Mud97, MK84a, Pik84b, Pik87, Pik90, Ral76c, Rei77, RT70, Ros84, SBA92, Smi77, Smi06b, Sta89, Tar80, Com79, Wen79b, Wes81, Chr87, GGZ02, McI73a, McM79, NT05, PT93]. Text-Oriented [Pik84b]. textbook [FG03, May01, SKS01]. Textdateien [GGZ02]. Textmusterverarbeitung [Sta89]. textual [Hor90b]. Textverarbeitung [Mau97, Str94]. Thee [Fro90, Sta04]. their [DD05, SZ95]. Them [Arn91, BED^{+}85, ITC93]. themed [Ame09]. themselves [Sim96b]. Theory [ACM75, Joh78a, AV95, Had04]. There [CPJ^{+}98, Eps92, McIxx, Sta04, O’D83a]. Thermal [XZY^{+}10]. Thermal-Aware [XZY^{+}10]. Thiel [Ano16b]. Thing [CPJ^{+}98]. Things [Gra03, Lut99, Gra06b]. Think [Tay96]. Thinking [Sov92]. Third [Loz05b, USE86d, Ano00a, Teo13, TRS18]. Third-Party [Loz05b]. Thirteenth [Uni98b]. Thomas [Kad95b, Kad95a, Koh06, Wei93].

Thompson [Ano11, CUH99, FES^{+}99]. thorough [Lov10]. Those [Ros95, Mud97]. Thou [Dec04]. though [Len97]. Thoughts [Koe79a]. Thousand [Bro90]. Thread [Fra79, MNN04, SS95b]. threaded [Lar94]. Threads [Alf94, BW88a, GMB93, LB96b, Mue93, Nor96, RR03b, But97, TBG^{+}87]. Threads.h [Ano00e]. Threaten [Gar97]. threats [XWRZ19]. Three [Dav93a, JJ91u, Sim96c, Sim97]. Tickerplants [BGH^{+}90]. Tiger [JR05, Pog05, Tay05, Top05, TSM^{+}06]. til [Kri91]. Tilde [CD85]. Tiled [Fra92]. Tilson [Ano95d]. Time [Abb03, All83a, All83b, Ano92l, Ano92m, ASS79, ACS79, AO86, BFGK89, Bar88a, Bar88b, BL75, Bel79c, Bla83a, Bla83b, BRD77, Bou78a, Bur76, Cha79, CG95b, CN97, CK78a, Cra88, Cro78a, DHP77, DHM78, Dou02, Fra78, Gal91a, Gal92, Gau98, Ger02, GZ84a, GZ84b, Har80, HRO82, HSM03, Hoo79, Hor82, Isa86, JR78, JL78, Joh95, Kau90, KLO78, Kol77, LTW77, Lim05, Lio79b, Lio79c, LMT78, Lyc75, Lyc78a, LB78c, Lyc78e, LC78c, Mar04, MPT78, MCM78b, McS78, Mul79c, NP78, Ney83a, Ney83b, Nis01b, NMP82, PNK02, Pek78, PB84, RMF19, Rez94, Rit71, RT74, Rit7 , RT78, Rit78, RJLK78, Rit9 , RT79a, Rit80, RT83a, RT01, RT21, RT79b, RT83b, Rov78, RDK75, Sch76a, SM88a, SM88b, SM88c, Stu78b, Swa74, RT5 , Tho78, TG79]. Time [Ass88c, Vol75, WSS78, WS78, WM00, Ano93v, Ano93z, Ano94i, Ano95j, AKST93, CG95a, Fer04, FL80, FPG89, Gal91b, GMTW08, HST08, IEE91b, IEE94c, LH84a, LH84b, Nad98, PBM^{+}22, PR96a, RLB08, RT73a, RT73b, Rit84a, Rit84b, Rze93, SJP^{+}09, She03, STT86a, STT86b, Wai95, IEE91b, Law98, Epp00b, Hoo78, Nis03, She03, Sim96c, HST08]. Time-Saving [HSM03]. Time-Shared [Bur76, LTW77, Har80]. Time-Sharing [Bel79c, Bou78a, CK78a, Cro78a, DHP77, DHM78, Fra78, JR78, JL78, KLO78, Lio79b, LMT78, Lyc75, LB78c, Lyc78e, LC78c, MPT78, MCM78b, NP78, Pek78, Rit71, RT74, Rit7 , RT78, Rit78, RJLK78, Rit9 , RT79a, RT83a, RT01, RT21, RT79b, Rov78, Swa74, RT5 , Tho78, TG79, WSS78, WS78, Lio79c, Ney83a, Ney83b, Rit80, FL80, RT73a, RT73b, Rit84a, Rit84b]. Time-Varying [Sch76a]. timeliness [RLB08]. Times [Lin02b, Lin02c]. Timesharing [MF83, MM88]. TimeSys [Ano03g]. Timing [Ral80, Sch77]. Tinfoil [ACF^{+}86]. Tiniest [Bar00c]. Tiny [Coc03c, Yue01, Bar00c, SAN08, Kas14]. Tip [Ros95]. Tips [Don96a, HSM03, Jae86, Lav04, RRF93, RRH94, Smi06b, Stu01, Bes05, Fer04, Fli03, Ran04, Ran06, Ree06b, Stu04]. tishisheru [MO01]. Titan [Cod89]. titans [Way00]. Title [Cro78c]. titles [McI69a]. tively [Tip92]. Tk [Ame95, Bla02, Bla04, HM98, Inf95, Ous94, SSC95, Wal97a, Wal95c, Wel95, Wel97, Wel98, Wel00, WJH03, ZK05, Zim98]. TLB [ST03]. TLS [DC17]. TMG [McI72]. To-Do [Pog17]. Today [Ano93-27, Ano19e, Jen19]. Together [ZB84, IEE91b]. Tokyo [Uni98a]. Tolerance [BBG^{+}89]. Tolerant [BW94a, BW94b]. Tom [Swa00b]. Tomorrow [Ano19e]. Too [Smi06b]. Tool [Ano98d, Ano02b, Ano02c, Ano03f, Ano03g, DM79, Feu78, Jen02, Kor89, Kro99, Lit80, MK04, RWNA87, SB02, Sny78, Spi00, TvK83, Tuo82, WB92, Ano03c, Ano03d, Ano03e, FSS99, Inc83, Lei95b, Lei96a, Lei96b, MM06b, NT05, PRC^{+}23]. Tool-Based [Spi00]. Toolbox [Hun85, NC08, MNC09]. Toolkit [JW99, Lib94, NO90a, NO90b, NO92, NO93, O’R90b, Ous94, PK93a, PK93b, Pon98, Pon99a, Pon01, Ser00, Sha00a, Wol04b, Wol04a, ACS98, Bru02, Fra99, KMV^{+}09, Pon03, SML06, SVV92, AS90, Kim95, PCD92]. toolkits [Ano03e, Smi91]. ToolMan [Sin97]. Tools [ASU86, ALSU07, Ale96b, And88a, Ano98d, BK86a, BAKW92, Bon79b, Cha06, CV80, DRN^{+}23, Dul94, Fel79b, Fla94, Gla97, Gre19, GS00b, GL94, GGZ02, HL95, HH97b, Haw85, Hol92, JL78, JL87a, JL87b, Jor04, Ker84b, Ker85, Ker89, Kol93, Kro99, Lav04, Len94, O’D83c, POLo93, POL^{+}97, PPO^{+}02, PPO^{+}03, Roc78, SW99, SHN97, Sou94, Spi94, Spi05, Sta89, Usr82, Win02b, Wol04b, Zwa95, dFAAW90, AC11, Ano95i, Ano00g, Ano03d, BW96, BGR89, DF91, ER86, FH89, Fli03, FJ97, Fri97, Gai83, JP09, Joh80, JJ91q, L^{+}11, MM00, Max00, MMG07, Mol14, Ni97, POL94, Pet88, PB09, PH87, Ran04, Ran06, RRD^{+}03, Roo79b, Ros00, Smi03b, Web95, Yos85, ?, ?, Bac92, Bea98, Fla90]. Tools.h [Ano00e]. Toolset [Ano02b, Ano03e, McC92, ZK05]. Top [Ano00d, Ano00f, Poo91, Won92, Ale96b, Ang97, Ang99, HBH18, RFS05, RFS06]. top-down [Ang97, Ang99, RFS05, RFS06]. TOP100 [EEK^{+}04]. Topaz [MS87]. TOPS [Bee87]. TOPS-20 [Bee87]. Toronto [Sof83, USE94c, Eis79b]. Totally [Ric94a]. TotalStorage [D^{+}04]. Tough [Ano96d]. Tour [Col89, Col17, Cus92, Hol12, Joh79b, Joh83, MRxxa, Rit8x]. Tower [Ano98c]. TowerJ [Ano98c]. Town [USE01a]. Toys [Zwa95, NW03, Neg06b]. Trace [MZG14, ODH^{+}85, SWF^{+}22, Cla88, Hol87e]. trace-driven [ODH^{+}85]. TRACE/UNIX [Cla88]. traceability [Ni97]. tracebacks [Dun92]. Tracer [Rod86]. Tracers [GD18]. Tracing [Gra01a, ZDS85, CMR^{+}21, GM11, KIT93, MGMMI10]. Track [New92, Rob75, USE98b, USE01b]. Tracking [She06, Sto89, Sto90, Bla86]. Trade [De 78, Kri84a, Kri84b, PLR85]. Trade-Offs [Kri84a, Kri84b, De 78, PLR85]. Tradeoffs [PLR00]. Trading [LSC^{+}88]. Tradition [MR92, MRxxc, dBB08]. Traditional [FST91, ALP^{+}21]. Traffic [AR78, Cho98, Mic79b, Rei97]. Trail [MAT85b]. Training [Hos83, Hos84, Jae89a, Kom00, Kom02, QH01, Wel98, Sta05b]. Transaction [Ado89, ACK89, Bur79a, CCF89, Hay77, HKS^{+}04, MH88, Tal89a, Tal89b, Tin78b, USE89g, Wei78a]. Transactional [RRP^{+}08, RRP^{+}07, RHP^{+}07, RRH^{+}08]. Transactions [Pov99, SO92, Ste96]. TransBase [Kem91]. Transfer [dC87, Jan03, MAP^{+}22, Nor79, HH04]. Transfers [Ral80]. Transform [Wol93]. Transformation [Kau7 , Kau79a]. transformations [FRRJ07, YHR92]. transforming [Che90a]. Transistor [Bar00c]. Transition [Dun91b, Lon77]. Translation [AKB85, BS79a]. Transmission [PNK02]. Transparent [CCDP01, Fel94, Pik88, PBKL95, URK85, ZG96, MHS02, ST03]. transport [HK05]. Transportable [SS93]. Transporting [JH83]. Transputer [LM88a, LM88b]. Transputer-based [LM88a, LM88b]. Traps [Koe89]. trauma [GW98]. TraumaTIQ [GW98]. Traveling [Lin65, LK73a, LK73b]. Traveling-Salesman [LK73a, LK73b]. Travelling [Rau01]. Tree [Skl91, Yos85]. Tree-Based [Skl91]. Trees [CD85, McI69b, AJ75, Wei81]. Trenches [Ano99b, McC98]. Trends [Ano95c, Che94, Edw98, Fra5 , Fra76, LZHS11]. TRI [ACM93]. TRI-Ada [ACM93]. Trial [IEE86]. Trial-Use [IEE86]. Trials [O’D83b, Swi79]. Tribulations [O’D83b]. Tribute [Aho12, Ker12, McI12, Pat12, Sch12]. Tricks [Smi06b, Sou94, Fer04, Wis03]. Tridia [Ano00e]. Trigram [Cox12]. Trilogy [McIxx]. trio [CM06, Jen19]. Triples [Pol93]. Troff [Ker90a, Les78, OK90, EP87b, GL88, Ker78, Ker82b, Oss74, Oss76, Oss79, Rod87, Smi76, CL90b, IBC87, OK92, EP87a, Les76b, Sun86]. Tromsø [Eur91]. TRON [BBS95]. Trouble [Can79, Dea78, Mar78c, Ger85]. troubles [HBB13]. Troubleshooting [Joh92, WF03, KCTT06, Olk03]. Tru64 [Ano99d, Che02, Don02, Han01, HS04]. Tru64-UNIX [HS04]. Tru64/DIGITAL [Ano99d]. TruCluster [CGS96]. True [Anoxxb, FES^{+}99, Jac87]. Truly [Mud97]. Trusted [CW92, SM89]. Truth [Nor81]. Try [Mor86]. TS [Bra78, Che78, CM78b, Dep79a, FDMS79, Hai78a, HW78, Hai78c, Hai78d, Hai79, Koe78b, Kow79, SM79, Vog79c]. TSPS [Mic79b]. TSS [Joh7x]. TTGR [KS85]. TTGU [Kau90]. TTY [KM87]. Tune [McK84]. Tuned [Hob91, Hob93]. TUNEX [Sam89]. Tuning [Alo01, Chu98, Coc94, Coc98, CG93b, DMT02, Gra00a, Jac87, Lou90a, ML02a, Nas97, Pac02, Ric96, Sam87, SW99, Sys97a, You95, Alo97, Alo98, Bes05, FSW02, McK91, McK92a, McK92b, MK96, Olk03, Sam89, SRW04, Sim93]. TUNIS [FH88, HMP83a, HMP83b, Hol82]. Tunnels [Luc18b, UTC84, Luc12]. TURBO [Tar89]. Turbo [Kin89]. Turbulence [LWL05, LWL06]. Turing [Coc03c, CH83a, CH83b]. Turn [Til87a, Til87b, Gil96a]. turns [Too11]. Tutor [WK92, Har93, SAS94b]. Tutorial [Arn93, Bec06, BK91a, BK91c, Feu79b, Hol91, Ker74, Ker78, Ker79c, Ker7x, Ker90a, KGP90, MB77, Mas77b, Ros91, Sch76b, SM77a, Str84b, Wom94, Bec07, Jos99, Jos12, Liu11, She02, Str85]. Tutorials [Smi06b, Uni96a]. TV [Gar97]. Twelfth [Uni98a, USE98d]. Twenty [Sal97, Sal03]. Twenty-Five [Sal03]. Twig [Tji85]. Two [Ano03g, CW89, CW91, Dou02, ER86, Hum88b, KS85, Lud77a, Pin99, SWG92, Swa77, TH83a, TH83b, All97, Hor90b, MMTK94, SJP^{+}09]. Two-Axis [Dou02]. two-dimensional [MMTK94]. Two-Point [CW89, CW91]. Two-Way [SWG92]. TX [USE88g, USE91d, USE85b, USE89d, USE91f]. TxLinux [RRP^{+}07, RRP^{+}08, RHP^{+}07, RRH^{+}08]. Type [And80, BSS^{+}95, BSSW96, Höf01, Str80a, Str82]. typechecking [Car84]. Types [Kle86, Lud78b]. Typesetter [BBK13, Ker82b]. Typesetter-independent [Ker82b]. Typesetting [BR87, BK86b, Ben88a, BK90a, BK91a, BJK92, Ede77, EP87a, EP87b, Geh86, Geh87a, Har11, KC75, KC76, KC78b, KC79, Ker81a, Ker82a, Ker84a, Ker90b, KC90, Ker91, LK76a, LK76b, LK77, MM90a, MM90b, Sri93, Tut83, VK79, Van80, GL88, KC78a, Ker81b, Com79]. Typing [CL90b, Eng76, Les76b, Les78, MK85b, MK90, SP77]. typo [MC75a]. Typographical [MC74, MC75b, MK84b].

U [KKF^{+}84]. U.S [Gre93]. U.S. [Coc03b]. Ubuntu [Gag06, Gag07, Gra06b, HHH11, HBB^{+}06, H^{+}07a, H^{+}07b, HH06, HH07, HH08, HH09, HHHT10, Kra07, Neg06a, Neg07a, ORC06, Tho06a, vH06, vH07]. UC [WAC84]. UCSB [Com79]. UCSD [Mer82]. Ufo [AISS98]. UFOS [BFGK89]. UFS [LPLF91]. Uhl [Kad95b, Kad95a]. UIO [Che87b]. UK [Ano95a, Ano93g]. ULE [MR14a, Rob03]. Ullman [Che75a]. Ultimate [Hew90, Luc02, Das01]. Ultra [VTS^{+}23]. Ultra-low-power [VTS^{+}23]. Ultracomputer [EGL86]. UltraSPARC [Ano96e]. umfassende [Kuo99]. Umgang [DF00, Kuo99]. Umgebng [CK06i]. Umgebung [CK06a, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h]. UML [Ano03f, Coc01c]. UMLinux [SB02]. Umsteiger [Ron05a]. unabhängige [Lin02b, Lin02c]. Unacceptable [NHNW94]. Unbound [Hah94a, Hah94b]. unbounded [Afz03]. Uncompromising [Bra04]. Uncovering [Bac92]. Undefined [WCC^{+}12]. Undercover [Ray98, Ste00]. undercut [Way00]. underground [Mon95]. underpants [Ame09]. Understanding [ALL13, Ben06, BC00, BC01, BC02, BC03, BC06, Env97b, Gal93, GZXZ23, Gir88, HSG03, Jan08, Lun93, Mar78d, Mue94, PLM06, Ros92]. UniCC [Mey18]. Unicode [Uni96a, Uni98a, Uni98b, Che97, DM97, Hiu96, Hiu97, Lei96c, Lei96e, Lei96d, Li98, Mud97, Ran96]. Unicode/ISO10646 [Uni98a]. UNICOS [Cra90, Rit88a, Rit88c]. UNIED [War78]. unification [WDG^{+}06]. Unified [LBN14, Sun96, Sch95c, SB08]. Uniform [NR94, Rug83, Che87b]. UniForum [Ano93-33, Ano93-33, USE84b]. UNIGRAPHIX [Sch77]. Union [Kau90, PM95, Gar97]. UniPress [Uni85]. Uniscape [Ano98d]. UNISEX [KE85, Sak97]. UNISEX-C [Sak97]. Unit [Hoo78, Hoo79, Sch81]. Unite [BMR01, BMR82]. United [Lob85]. Units [LJ84, Mil90a, DMM86]. UNITY [Tin79b]. UNIVAC [Lyo76b, DF79]. Universal [KKF^{+}84, PW77, Mey18]. UNIVERSAL FILE NAMES [Wan90]. Universe [SUM95, Tan87c, Kad95a]. University [BBNS80, BSK87, Ric77, McK85b, Gor81, KDP83, MN69]. UNIX [Ano92s, Ano93f, Ano93d, Ano93h, Ano93j, Ano93l, Ano93p, Ano93w, Ano93-27, Ano93y, Ano93-32, Ano93-30, Ano93-33, Ano93-42, Ano94b, Ano94k, Ano94n, Ano96b, Ano98c, Ano00e, BBJ86, Bac92, Bas87, Bea98, Bee96a, Bee96b, Ber85b, Cab83, CKM85, Col92a, Cuc90, Dav97, Don02, EST86, Far92c, Far92d, FH92, Fio04, Fla90, Fur85, Gil93a, Gra89b, Gre92a, Gre93, GZ89, Gwy90, Hur94, JLSG84, Jon88, Kir84, Kra83, Lea90, Lea92, Lin95, Luc03b, Mel95, Nel90, NN92, The02, Per93, Poo91, Pug88b, Sal97, Sau88, Sha94, Sim93, Sim95, SWG92, Swe01b, SM85, Teo13, Tho19a, Ass89c, USE95c, Ube76, Wei93, Wid00, Wil94, Won92, van97, Die11, DF00, GSS03, Her99, JR02, Kuo99, LMKQ90, May01, Mor85, O’R90a, Ray08, RB06a, Str94, LHM^{+}83, Mac86]. UNIX [?, ?, ?, ?, ?, ?, ?, ?, ?, ?, Ame90b, Ame90a, Ame91, ?, ?, ?, ?, ?, ?, ?, ?, Abl93, AL92a, AL95a, AL95b, AL97a, AL97b, AL97c, AL91, ALO90, ABB^{+}86a, ABG^{+}86, ABB^{+}86b, Ado89, Afz03, Aho84, ASS85, AS93a, AS97, AGP^{+}92, Alb84, Alb76, AKB85, ATS94, Ale96a, Ale96b, All83a, All83b, All87, AW89, Alo97, Alo98, Alo01, AC11, Ama88, And02, AA86, And88a, And88b, And91, ACH91, And95, ACK89, AS93b, Ann00, Ano77, Ano80, Ano82, Ano85, Ano88b, Ano91c, Ano92c, Ano92d, Ano92e, Ano92f, Ano92g, Ano92a, Ano92b, Ano92n, Ano92o, Ano92p, Ano92q, Ano92r, Ano93a, Ano93b, Ano93c, Ano93e, Ano93i, Ano93k, Ano93r, Ano93o, Ano93-34, Ano93-35, Ano93-36, Ano93-37, Ano93-38, Ano93-40]. UNIX [Ano93-39, Ano93-41, Ano94l, Ano95e, Ano95i, Ano96a, Ano96d, Ano96e, Ano97a, Ano97b, Ano98e, Ano99d, Ano00d, Ano00f, Ano00g, Ano01, Ano19a, Any85, AM86, AQ84, ACF^{+}86, Ari92, AGHR89a, AGHR89b, Arm96, Arn86, Arn91, Arn93, Arn98, Art86, Art90, Art94, AB97, ADJS78, Aud96, AO86, Aul96, Ava92, BJP04, Bac84, BB84, Bac86a, BBJ86, BG88, Bac86b, BSS^{+}95, BSSW96, Bad98, BH86, Bai85, Bai96, Bai87, BS92, Bal96, BFGK89, Bal94, Bam98, Ban82, BR82, BS80, BP86, BW89, BGW93, Bar91a, BM87, Bar88a, Bar88b, Bar83, BW96, BR87, BW88a, BW88b, BW88c, Bas91, Bas87, Bau03, BK85, Bec84, BP91, Bee85, Bee86a, Bee87, Bee90, Bee93b, Bee94, Bee86b, BGR89, Bel79a, Bel79b, Bel79d, Bel79c, BED^{+}85, BDD^{+}93, Ben98, Ber85a, BGH^{+}90, BBS95, Ber98]. UNIX [Ber76, BP88b, BP88c, BP88a, BAC^{+}89, BGMR91, Bey77, Bey87, Bha12b, BCM88, BE88, Bin88, BE89, BE00, BBM85, Bis87, Bis88, Bis89, Biv87, Bla83a, Bla83b, BMS83, BMM85, Bla86, BT98, Ble83, BEHW86a, BEHW86b, BAKW92, Blo99, BM79, BHK^{+}84, Boh86, Bon79a, Bon79b, BFS89, Bor88, BBG^{+}89, BQd86, Bor79, BW94a, BW94b, Bor98, BNB87, Bot84, Bou7 , Bou78b, Bou78a, Bou82, Bou86, Bou90a, Bou97, BHM98, Bou90b, Boy84, BCC^{+}91, Bra93, Bra76a, Bra76b, Bra77, Bra78, BMV78, BRW89, BB93, Bre86, BBM84, Bro89a, Bro94, Bro97, Bro98, Bro89b, BBP99, Bry83a, Bry83b, BBG85, BS90, Bv92, BCGS91, Bur93, BH97a, BH97b, Bur79b, BK84, BP84, Com84, CCA84, Cal93, Cam91, CBB^{+}91, CE88, Can80, Cap82, CGS96, CMQ93, Car88a, Car88b, Car93a, Car89a]. UNIX [Car86b, CJL^{+}89, CCM^{+}90, Car89b, Car82, Car04, Cas93, CW92, CCDP01, Cer82, Cer83, CMM88, CQ83a, CQ83b, Cha97, Cha79, Cha96, CHRG95, CE97, Che96, Che99, Che02, Che78, CB83, Che90a, Che87a, CWS90, Che79d, CL90b, Chi02, Cho98, Chr83a, Chr87, Chr88, CR94, Chr83b, Chr90, CN97, Chu03, CCD^{+}87, Cla88, CS86, CCF89, Cle83a, Cle83b, C^{+}78b, Coc01a, Coc01c, Coc03a, Cod86, Cod89, Cof90, CM78b, CK78a, Coh79, COV79, Col83, CFA85, Col84a, Col84b, Col87, Col89, Col17, CD85, CN88, Coo85, CAG89a, CAG89b, CJ92, CGMS89, CH83a, CH83b, Cor05b, Cou85b, CJ94, Cra93, Cra83, Cra88, CBL96, Cro98a, Cro78c, Cro78a, CS90a, CS90b, Cur89, CKDS90, Cur90, Cur91, Cur92, Cur96, Cur97, CT94, Dal99, Dan83a, Dan83b, Das88, Das01, Dat88, Dav97, DM88]. UNIX [Dav89, Dav96, Dav98, Dav00, Daw82, Daw86, DF93, De 76, De 78, DeV93, Den83, DG00, Dep79b, Dep79a, DI89, DiM00, DNQ^{+}83, Dil02, Dix82, DHP77, DS77, DHM78, Don95, Don96a, Don97a, Don98a, DO87, DF91, Dow77, DB95, DN87, DuC94, DJM86, Duf89a, Duf89c, Duf90b, Dufxx, Duf89b, Duf82, Dul94, Dul96, Dun92, Dun91a, Dun91b, Dun89, Dun91c, DF95, Dys96, Dzo09, Env97a, Env97b, Eck95, Edd90, EGL86, ELS88, ET88, ET92, Eis79b, Ela83, Ell79a, EP87a, EP87b, ER86, Eng76, Eng91, Eps92, Erl88, Eur91, ESSB93, Eva97, Eve78, Eyk86, Eyk88, FRAC85, FY18, Far92a, Far91a, Far91b, Far92b, Far90, Far95, FG80, Fat82, FDMS79, FG91, Fed77, Fed78a, Fed78b, Fed83, Fed84, Fed88a, Fed88b, Fel84, Fel94, FK90, Fel79e, FMM84, FFH^{+}86, FA88]. UNIX [Fer85, Fie88, FHS91, Fil85, FKN85, Fis86a, Fis86b, FH89, Flo90, Fol77, FM94, FG03, For00, FN83, FJ97, FJPR93, Fox85, FF88, Fra90a, Fra90b, FT83, Fra92, FES^{+}99, Fra83, FR77, FE77, Fra78, Fra99, Fre90, Fre91, Fre98, FHR84, Fro90, Fun87, FH88, FPG89, FST91, Gal91b, GB92, Gan95, Gar97, Gar91, GS91, GWS94, GS96, Gar90, GS89, Gau98, Gay00a, Gay93, Geh86, Geh87a, Geh87b, GL88, GS00a, Ger02, Ger82, Get86, GBH86, Gil91, Gil96a, Gil96b, Gt92a, Gt92b, GT92c, G^{+}98a, G^{+}98b, Gin83, Gla97, Gla91, GA99, Gla02, Gla77a, GT78a, GT78b, Gla79, GT79, Gli92, Gol98, Gol79, GDFR90, GES98, Gom85, GC94a, GC94b, Goo90, Gos81, Got02, DF79, Gra84, Gra88, GM84, GM90, Gra87, GP88, Gra97, Gra98e, Gra98a, Gra98b, Gra98f]. UNIX [Gra98d, Gra98c, Gra99b, Gra99c, Gra99d, Gra99a, Gra00a, Gra00b, Gra00c, Gra01a, Gra01b, GR76, Gre82a, Gre82b, Gre96, Gro93, Gue87, GLOR91, GBG97, GZ84a, GZ84b, HO 78, Had04, Hae89, Hae90a, Hae90b, Hae90c, Hae90d, Hae90e, Hae91a, Hae91b, Hae92, HTU96, Hag83, HK79, Hah93, Hah94a, Hah94b, Hai77, HRW77, Hai78a, HW78, Hai78b, Hai78c, Hai78d, Hai79, HK83, HG93, Ham87, HO93, HS82, HK95, HQZ^{+}87, Han86, HKF84, Har93, Har92a, Har92b, Har94b, Har96, Har01a, Har01b, Har80, HBMT96, HL95, Har83, Has78, HY79, HH97b, Hau94, HR78a, HS87b, HGS99, HdBM75, Haw85, HL85a, HL85b, Haw86, Haw87, HRO82, HD97, Haz99, HS83, Hea98, HLW84, HCC^{+}87, HJAW88, Hek89, Hel97, Hel92, HA84, HP89a, HP89b, Her93, Hes91, Hes94, HS04, HCN85, HT92, HT94a]. UNIX [HT94b, Hil93, Hil98, HH86, HHLS90, HAB^{+}98, Hiu96, Hiu97, HA95a, HA95b, HA96, Hof97, Höf01, HP89c, HLLP79, Hol75, Hol97, Hol92, HMP83a, HMP83b, HK82b, Hol87a, Hol87b, Hoo83, Hor82, Hor92, Hor79, Hos83, Hos84, Hu86, Hub96, Hub79, HM90a, Hug88, HM90b, Hun88a, Hun88b, HH91, Hun92b, Hun92c, Hun92d, HH94, HD98, IBM09b, Gur81, Ima76, Ima79, IBP03, IvW87, IB93, IBC87, IJ98, Isl83, IHR99, Jac77b, Jac84a, Jac84b, Jac87, JL98, JTUB85, Jal92, JH83, Jam97, JAV86, Jen75, JR78, JL78, Joh79b, Joh87, Joh95, Joh96a, Joh96b, JH86, JJ91s, JJ91b, JJ91c, JJ91a, JJ91t, JJ91q, JJ91d, JJ91o, JJ91e, JJ91f, JJ91g, JJ91h, JJ91i, JJ91j, JJ91k, JJ91l, JJ91m, JJ91n, JJ91r, JJ91p, JJ91u, JJ92d, JJ92a, JJ92b, JJ92c, JJ92e, JJ92f]. UNIX [JJ92g, JJ92h, JJ92i, JJ92j, Jon84, Jon98, Jos97, JF81a, JF81b, Joy97, JN88, Jud90, Jun85a, Jun85b, JK86, KNI88, KVCS98, Kam02, KIT93, Kap79, Kap92, Kap99, Kar83, KM86a, KTB^{+}88, Kar84, KDP83, Kat83, Kat82a, Kat82b, Kau78, KK89a, Kay79, KI99, KE85, Kem93, KL94, Kep85, KM78, KLO78, Ker79a, Ker79c, Ker79d, KR79, KM79a, KL82, KM82, KP84, Ker89, Ker19, Ker10, Keu92, KN93, Kil90, Kin83, Kin93, KK89b, Kiv84, Kle86, KM79b, KPT^{+}02, Kni92, KW85, KW89, KW90, KW03, Kod82a, Kod82b, Koe78b, KM87, Koh99, KAH83, Kol93, Kol94, Kor97b, Kor97c, Kor97d, Kor97a, Kor89, Kos96, Kow79, Kow90, Kra88a, Kra88b, Kra78a, KM83, Kri84a, Kri84b, Kri91, Kuo98a, Kur90, KI82, LJ84, LB96a, LKY93, Lam83, LM88a, LM88b, Lan84, LP89]. UNIX [Lap87a, Lap87b, Lar94, LEG88, Law83a, Law83b, LLM87, Lee89, LMKQ89, LMKQ90, LM91, LA90, Leh95, Lei95a, Lei90, Lei98, Lei96c, Lei96d, LT76, Len97, Len94, Len87, LT91, LK76a, Les76b, Les77a, LC77, Les78, LC78a, Les79a, Les88a, Les88b, Lev02, Lev79, LY93, Lev96, LY97, LY98, LY04, Lew91, Lew88, Li98, Lib87, LR89a, LR89b, Lic79a, Lic79b, Lin84, Lio79a, Lio79b, Lio79c, Lio96, LS92, Liu96, Lo79, LPLF91, Lob85, Loi79, LL83a, LL83b, LR78, Lon77, Lon90, LH84a, LH84b, Lou90b, LC93, Loz05a, Loz05b, eL85, Lub90, Luc13a, Lud77a, Lud78a, Lud78b, LMT78, Lud79b, LCH^{+}81, LBSL97, Lun96, LSC^{+}88, LS83, Lut99, LC73, Lyc74a, LC75, Lyc75, Lyc76a, Lyc76b, Lyc76c, Lyc76d, Lyc77a, Lyc77c, Lyc77f, Lyc77h, Lyc78b, Lyc78c, Lyc78d, LB78c, Lyc78e, LC78c]. UNIX [Lyc85, Lyo77, LR77, Lyo79, Lyo18a, Lyo18b, Lyo18c, Mac97a, Mac83, MBC^{+}86, Mac97b, MLRC88, Mah07, MC85a, MC85b, Maj77, MS88, MWCR90, MWCR91, MM86, MF83, MK85a, Man01b, Mar75, Mar78a, Mar78b, Mar82, MH88, Mar78f, Mar84, Mar01b, Mas77b, Mas87, Mas88, MTY01, Mat97, The92c, The92a, The92b, Mat93, The93, Mat80, GF80, Mau97, Max02, May79, May97, McC88, MS92, McC95, McD87, McD84a, McD84b, McG97, MM83a, MM90a, MM83b, MM90b, McI5 , MPT78, McI87, MR92, MRxxc, McI89, MS87, MS89, McK93a, McK82, MJLF84, McK85b, MK87, MK88, McK91, McK92a, McK92b, McK92c, McK93b, McK94a, MK96, McK96b, McK97b, McK97a, McL83, McL84, MCM78b, McM99, McN00, McS79, MK91, Mea88, Mei92, Mei84, Mer82, MAT85b]. UNIX [Mer01, MM88, MHS02, Mic79a, Mic79b, Mik89, Mil91, Mil98, Mil00, Mil78, Mil79, Mil84a, MMS86, MFS90, Mil90b, Mil90c, MK93, MKL^{+}95, Mil99, Mil90d, Mil86, Min82a, Min82b, MMTK94, Mit84, MMTW88, Miy86, Moh85, Moh99, Mol77, Mon00a, Mon00b, Mon95, MS84, Mor96, Mor79b, Mor81, Mor88, Mue79, Mue93, Mue94, Mue78b, Mui95, Mul87, Mul79c, MTT83, Mur83, MZA96, MK85b, MK90, MBL89, Mus96, Mus02, MNC09, ML02b, NH99, NH00, Nad92, NP78, NM83a, NM83b, Nas97, NB84, Nee91, NMS83, Nee90, NS92, NC08, NR94, NSS89, NSSH95, NSS^{+}01, NSHW10, Ney83a, Ney83b, NHNW94, NLR84a, NLR84b, Nor79, Nor81, NMP82, Nor96, NH91, Nor93a, Nor93b, Nor93c, NS94b, Nor91, NL79, NL82, O’B82, OH97, O’D83a, O’D83c, O’D83b, O’D87, O’D90, OD^{+}86, O’R90a, O^{+}98]. UNIX [Ass99a, Ols78, Osb92, Osm96, ODH^{+}85, Out82, PBS98, Pab96, Pac00, Pad93, Paj88, Paj92, PDG^{+}87, Pal87, PR96a, PDG02, Pan94, Pan88, PR85a, PB91, Par90, Par88, Pas88, Pat96, PV00, PS82, PK93a, PK93b, PMD88, Pat83, Pat99, PWC00, PF84a, PF84b, Pay84a, Pay84b, Pea80, Pea83, PBWB84, PS77, Pea77, POLo93, POL94, POL^{+}97, PTS98, PTS01, Pek78, PDL04, PR76, Per77, Per79, Per87, Per82a, Per82b, Per84, PH86, Pet88, Pet98c, Pet99, PS79, PJ89a, PJ89b, Pic83, Pik82, Pik83b, PK84, Pin92, Pir88, Pis77, Pla89, PBKL95, PLBL92, Poe87, PK88, Pol93, PB84, Pom04, Pon97a, Pon00, Pon02a, Pos88, Pov99, Pow84b, Poz83, PMW^{+}00, PR90, Pre99, Pri78, Pru02, Pug88a, Pug88c, Pug90a, Pug90c, Pug90b, Pug92, PY84a, PY84b, QSP85, QW93]. UNIX [QBdR87, Qui97, Qui99, Qui02, Qui05, Raf79, Rag93, Raj98, Raj00, Ral75a, Ral75b, Ral76a, Ral76b, Ral77, Ral80, RWNA87, Ran96, RC83, RR99, Ray04, Rea91, Red97, RP84, RW84, Ree88, RM92, RMxx, Ree82a, Ree82b, RFJ97, Rei98, RFJ99, Rei86, Rei97, Rei77, Rem97, Req85, Rez94, Rez95, Rib89, Ric94b, Ric96, Ric85, Ric98, RW92, RW86b, RW86a, Rie93, Rie94, Rit71, Rit2 , RT73a, RT73b, RT74, Rit5 , Rit7 , RT78, Rit78, RJLK78, Rit79b, RT79a, RT83a, Rit84a, Rit84b, Rit85, Rit8x, RT79b, RR96, Rob99b, Rob99c, RRD^{+}03, RR03b, Rob06b, Rob84a, Rob75, RB92a, Rob95, RB92b, Rob84b, Roc82, Roc85, Roc04, Rod87, Rod90a, Rod90b, Rod86, RKPP88, Ros95, RRF90, RRF93, RRH94, Ros96, RHK^{+}07, Ros79, Ros84, RW81, RT83b]. UNIX [Ros00, Rov78, RB82, Roy93, Rub79, RW93, Rug83, RW87, Rze93, SAS88a, SAS88b, SAS90, SAS91a, SAS91b, SAS93a, SAS93b, SAS93c, SAS94a, SAS94b, SAS95, SPS90, SPS93a, SPS93b, SPS93c, SPS95, Sab76a, Sab77, Sab78, Sal94, Sal99, Sal00a, Sal00b, Sal03, Sam87, Sam89, San93, SW82, San89, San91, San94, SF95, Sat90a, Sat93, Sax85a, Sax85b, Sax85c, Sch93a, Sch94a, SL95, SB96, Sch97, SW88, Sch88, SR85a, SR85b, Sch95a, Sch80, Sch94b, Sch95b, Sch93b, Sch93c, SS95a, Sch91c, SBA92, SF85, Sch85a, Sch85b, Sch90, SML06, Sch95c, SEZ^{+}01, SNG95, Sch77, Sch99, Sef93, Sei00, Sel94, SY91, SO92, SBMS93, Sem96, Sha00a, Sha81, SR77, Sha93, SH85, SSB92, She93a, SCJ96, SSWW83, SW91a, Shi77, SG95, SST90, SHVP05, Sim91, Sim96a, Sim96b]. UNIX [Sim96c, Sim97, Sin97, Siv90, SJ83, Skj82, Skj83a, Skj83b, Skj84a, Skj84b, Skj84c, Skj84d, Skj84e, Skj84f, Skl91, Sko98, Sku88, Sma88, ST89a, ST89b, ST89c, Smi76, Smi83, SM88a, SM88b, Smi89, SZ95, Smi06b, SM88c, SM79, SHN97, Sob89, Sob91, Sob95, Son87, SA95, Sou92, SKP94, SM89, Spe96, Spi94, Spi00, Spu92, Sri93, Sta89, SMD^{+}94, SP90, SD87, Ste90b, Ste92b, Ste96, Ste98, Ste99b, SR04, SFR04, SR08, SR13, Ste92c, SJ95a, SJ95b, Ste88, Sti83, Sto95, SS00, Sto93c, Sto93a, Sto93b, Sto93d, Sto94b, Sto94a, Sto94c, STA86, STA87, Str90, Sun96, Sun77, STT86a, STT86b, Swa74, Swa77, Swi77, Sym85, Sys97a, Sys97b, SW91b, Tag83, Tal89a, Tal89b, TH83a, TH83b, TvK83, Tan87a, Tan87b, THB06]. UNIX [TLK99, Tan84, Tan87c, Tan00, Tar89, Tar80, TBR93, Tay96, TA98, TJ04, Tay05, TG88, TP91, Tev87, TBG^{+}87, TTG^{+}87, TY82a, TY82b, Tho85a, Tho85b, TR94, TSO^{+}96, TR75b, RT5 , Tho76, Tho78, TS87, Til84, Til87a, Til87b, Tin78a, Tin78b, Tin78c, Tin78d, TG78, TG79, Tin79b, TSP93a, TSP94, TPS98, TSP93b, TBS87, TLR89, TBGE89, TNY92a, TNY94a, TNY92b, TNY94b, Too09, Top90a, Tra85, Tre88, TWM86a, TWM86b, Tuc98, Tun05, Tuo82, Tuo83, Tur87, Tut83, Com79, Uni82, USE88a, USE88b, USE88f, USE89g, USE90e, USE92e, USE93g, USE94c, USE95a, Uhl87, UTC84, Uma96, Ung78, Uni85, URK85, Vah96, Val92a, Val92b, VB92, Var96a, Var96b, van86, VB83, Ver89, Vic98, Vog74, Vog76a, Vog6 , Vog76b, Vog77a, Vog77c, Vog78b, Vog79a, Vog79b, Vog79c, Vog97, Vol94, Vol97, WMPC83]. UNIX [WMP84, WMP92, WIP^{+}97, Wal91a, Wal91b, Wal91c, Wal92e, Wal92f, Wal92a, Wal92b, Wal92c, Wal92d, Wal93a, Wal93b, Wal94b, Wal95a, Wal97b, Wal97a, Wal82a, Wal82b, Wal82c, Wam83a, Wam83b, Wan77, Wan78, WLHC84, WK92, WLC95, Wan96, War85a, War90, War84, Wat87, Wat88, WALK10, WALK12, Wat83, Wee99, Weh83, Wei78b, Wei82, WJ82, Wei88, Wei92, Wei94, Wei98, Wel94a, Wen79a, Wes81, Whi75a, Whi75b, Whi89, WF01, Wie89, Wil83a, Wil83b, WAC84, WB92, Wil82a, Wil82b, WG98, WCM93, Wil90, WL94, Win75, Wit01, WH83, Wol93, Won90, WSS77, WSS78, WS78, WB85a, WB85b, WK85, Woz82, WM01, Xia97, Yag91, Yao83, YT83, YE84, YLK93, YSF89, YR93, Yoo96, Yos85, ZRB^{+}93, ZL96, ZDS85, Zie93, Zim85, ZB84, Zin92, Zuc77b, Zuc83b, Zuc83a, Zuc83c, Zwi92, d’A92]. UNIX [dC93, dVd93, dFAAW90, vMM88, vM88, Sak97, AKCB86, Ame09, AD89, AJ^{+}02, And17, Ano95f, Ano98d, Ano00b, Ano03a, Ano19c, Ano19d, Ano19e, Bai99, BS88, Bar86, BD94, Bra95, Bro19, Bur01, Cap87, Car99, Che81, Che75b, CUH99, Coo03, CKM06, CKM14, Cro98b, Dal02, DCN02, DJ93, De 93, DeR99, DeR00, Den04, Dit93, DO88, FL80, FSS99, Fle02, Gab07, Gan03, Gar86, Gau81, GJ08, Gil88, Gre81, GKV21, GL94, Gue19, Hah09, HLMR79, HP98, Han01, HKVR21, Har00, Hei95, Hen92, Hin88, HK86, Hol82, Hol87c, Hor98, Hor02, Hos05, Hum88a, Hus92, HCG^{+}82, IBM09a, Jen19, JR04, JR05, JRR08, Joh80, Ker79e, Ker79f, Ker79g, KM81, Ker84b, Ker85, Kie02, Kri95, Kuo98b, Kuo99, Kur94, Kur97, Kut02, KL92, Law98, Lei96e]. Unix [Lel90, LK76b, LK77, Lév18, Lev88, LGJS91, Liu11, Luc03a, Mahxx, Man92, MHO^{+}76, May01, McE05, McN97, Mel87, Mel02, MSW06, Mer97, Mic03, Mic08, MK92a, MK92b, Mor78a, Nad98, Neu00, Ngu08b, Olc01, PDG^{+}88, PB21, dJPDD84, Pin99, PPO^{+}02, PPO^{+}03, PM87, PMW^{+}99, PR85b, PR85c, Pri83, Vib11, Rai89, Ram88, Ray99a, RRR09, RR15, Ric90, Rit9 , Rit80, RT01, Rit02, RT21, Rob00, Rob05, Rob06a, Ros78, Ros94b, RJR08, SKS01, SHLR80, Sil84, Skr89, SS06, Sor01, Spi05, Spi17, SA21, Ste02a, SR05, TPTS02, TJ03, Tay03, Tay06, Thi92, TR71, TR72, TR73a, TR73b, TR74, TR75a, TRS18, Too10, Too11, Too17, Top90b, TECB03, Ves88, Vog77b, Wag79, WG00, WLL96, Wal99, War88, Web95, Wei81, WDG^{+}06, Zim98, ZG96, dBB08]. Unix [dSVdLT02, ?, ?, CT83, Ano00a, Wei95]. Unix- [Her99]. UNIX-Based [McD84a, McD84b, Per87, PY84a, PY84b, WH83, HK83, KE85, KK89b, MBC^{+}86, SR85a, SR85b, Sti83, Ano01, Cha79, Cha96, Gil96a, Sak97, Mel87, Bar86, HCG^{+}82, HK86, MHO^{+}76]. UNIX-Betriebssystemen [Bas87]. UNIX-Compatible [Hol87a, Hol87b, Swa77]. Unix-Grundlagen [Str94]. UNIX-HATERS [Tho19a, GWS94]. UNIX-Like [MLRC88, HK95, LM88a, LM88b, Jon84, Hol87c, CKM06, CKM14, Neu00]. UNIX-LINUX [Smi06b]. UNIX-related [Sto94a, Hae90e, Hae91a, Hae91b, Sto93c, Sto93a, Sto93b, Sto93d, Sto94b, Sto94c, Wal91a, Wal91b, Wal91c, Wal92e, Wal92f, Wal92a, Wal92b, Wal92c, Wal92d, Wal93a, Wal93b]. UNIX-Serveradministration [Luc03b]. UNIX-Shell [Dav97, Die11]. UNIX-Systemadministration [HS04]. Unix-Systemen [Kuo99]. UNIX-Tabellenbuch [BBJ86]. UNIX-to-IBM [Ano93f]. Unix-to-NT [Ano98d]. UNIX-to-UNIX [Lyo79, PDL04].

UNIX-Type [Höf01]. UNIX-Werkzeuge [Sta89]. UNIX/32V [Bel79c]. UNIX/370 [Bur79b, Fel79e, Pri78, HLMR79]. UNIX/86 [BM79]. Unix/Linux [RB06a, Chu03]. Unix/Linux-Tasks [RB06a]. UNIX/MS [Ver89]. UNIX/MS-DOS [Ver89]. UNIX/NT [Ale96b, Gau98]. UNIX/PC [Fel94]. UNIX/POSIX [Ano95i]. UNIX/Prime [WJ82]. UNIX/RT [Bel79d, Bra78, Che78, Ima79, Kap79, KM79b, Loi79, Lud78a, Ral80]. UNIX/TS [Bra78, Che78, CM78b, Dep79a, FDMS79, Hai78a, HW78, Hai78c, Hai78d, Hai79, Koe78b, Kow79, SM79, Vog79c]. UNIX/Windows [Bor98, Don95]. UNIX/X [Aud96, Mei92]. UNIX/X-Windows [Mei92]. Unix50 [Ano19d, Ano19e]. unixdoc [Mar01b]. UNIXes [BMR82, BMR01, Tho19b]. UNIXhelp [Dav93a]. UNIXLIST [Rob75]. UNIX(R) [Noe98, Wal98]. UNIX^{T M} [Sch79b, Sch79c, Wes83, Ni97, Ano84]. UnixWare [Kuo99, Ano92s, NS94a, RRN95, Zwa95]. Unleashed [HH06, HP^{+}95, McK85b, NR03, TU02, TU03, AJ^{+}02, And09, BP00, Bal00, BH97a, BH97b, CV00, HHH11, HH05, HH07, HH08, HH09, HHHT10, RR02a, TU06, WWW99, Wal01a]. Unleashing [Kad95b, SU94b, SE97]. unofficial [Moh85]. Unsecure [LT91]. Unsolicited [Cla00]. unwired [WDJ04]. Upas [PC90]. Update [Car93a, Hae89, Hae90a, Hae90b, Hae90c, Hae90d, Hae90e, Hae91a, Hae91b, Koe79b, McC88, Sto93c, Sto93a, Sto93b, Sto93d, Sto94b, Sto94a, Sto94c, Wal91a, Wal91b, Wal91c, Wal92e, Wal92f, Wal92a, Wal92b, Wal92c, Wal92d, Wal93a, Wal93b, GMSP00, GMTW08]. Updated [TOML04]. Updates [MG99, SGM^{+}00, Ano03d, Ano03e, GMSP00, Mat93, MR10, MR14b]. Updating [HD98]. Upgrades [Ano96d, Ano02c, Ano03f, Ano03g, Ano03c, Ano03d, HBB13]. Upgrading [Luc03b, Ral78, Whe97]. Ups [Chu03]. upscale [Gil96b]. Urban [Lut99]. USA [ACM03, DFG99, USE88g, USE00a, Ano88b, ITC93, IEE02, MS05, Sof84, Uni96b, USE84b, USE85c, USE85d, USE86c, USE86e, USE87d, USE87e, USE88e, USE88a, USE89e, USE89f, USE90b, USE90c, USE90d, USE91b, USE91c, USE91d, USE92a, USE92b, USE92c, USE93b, USE93c, USE93d, USE93e, USE93f, USE93g, USE94d, USE94e, USE94b, USE94f, USE95c, USE95b, USE96a, USE96b, USE97a, USE97b, USE97c, USE99b, USE99c, USE99a, USE00b, USE01b]. usability [WBSK87]. Usage [GJ75, Gay90, HS82, Kor89, Maj77, MFBWW20, DI89, Ros78]. USB [Reg11]. Use [Ano82, GA04, HKF84, HN01, IEE86, IBC87, Jen02, Lyo79, Lyo76a, Mar78a, Mue78a, Ozi78, PKP05, Smi06b, Stu01, Ano94i, Cla79, Daw86, Den04, HO 78, Inc83, Kof00, Mor78a, Ron05b, Stu04, VD01, Var96b]. used [BY96, Gil96a]. useful [Spe96]. Usenet [OT88, TD91, TO92, Fie88, Har95, Wei92, Bel18, HH97a]. USENIX [Far95, Sof83, USE91d, USE92d, USE95c, USE00a, USE01b, Ano88b, Ano94m, Den99, Sof84, Usr82]. User [ACM88, AISS98, Ano80, Bel79d, BK91a, BK91c, BJK92, BGMR91, BG77, BS79b, Bro71, Bro73, Bro77a, Bro90, Com94f, Com94g, CJ92, CW15a, CW15b, DM88, Den83, Dep79b, Dep79a, Dik06, DF00, ESSB93, Far78, Gag01, Gil93b, Gra01b, GGZ02, Han86, Har95, HY79, Hobxxb, Hof77, Hoo83, HM90a, IvW87, Jac84a, Jac84b, Joh7x, KVCS98, KC76, KC78b, Ker84a, Ker91, Kir77, Kro92, Lat02, Len94, LS83, MP93, Man93b, Mar78g, The92c, McM99, Mey18, Mil00, Miy86, Mul79a, MD86, Ng77, Nor93b, NS94b, OQL88, Olc92, Oss74, Oss76, Oss79, OK90, OK92, PMD88, PCD92, Pon00, QO90, QO91, QO93a, QO93b, Reg78, RW86a, Rod87, Ron05a, Roo79a, Rug83, Rus91, Ryd3 , SPS93b, SPS95, Sch75, SW12, SP90, Str79, SS93, TY82b, Tho69, Tho72, Tri90, Ass89c]. User [Van81, YT83, AGS90, Biv87, Cab83, DJM86, GIM07, Ham90, HBH18, Luc13b, Mac91, Mac92, Man91, Mil98, Moh98, OK94, PS94, Pal87, PDG^{+}88, Pon02a, QO89, Ric95, RW81, SAS93c, Sun93d, WWCW19, War04, Wil98]. User-Controlled [CW15a, CW15b]. User-friendly [CJ92, AGS90]. User-Level [AISS98, Jac84a, Jac84b]. user-process [Cab83]. Users [Arn98, Bee87, Bou90a, Bou90b, Cur92, Edd90, Eis79b, Fie88, Hag83, JC92, Lyo76b, McC02, MBL89, Per79, Ron05a, RRF90, Sab76a, Sab78, Sil84, Usr82, VB92, VB94, Wes81, Zwi92, Bam98, Ber85b, BHM98, Com84, CCA84, Fox85, GA99, Gra05, HH04, KCTT06, Kom96, KL92, Lew88, Mat93, Mil01, MNC09, Red97, She02, SS06, TBR93, Jon88]. Userspace [GD18]. Uses [Ano03g]. USG [Can79, Csa77, Ros79]. Using [And88a, Bai85, BS92, BS88, BMF15, Bur79b, CWS90, CL90b, Cod89, Cur89, Cur90, Dil02, DM79, Don98b, DuB95, ECBK00, Eng76, Fel79b, Fel96b, FH89, Fol77, FT83, Fra99, FH88, Gau81, Gia90, GK97, Gir88, GT78b, HHLS90, Joh96a, Joh92, Kil94, Kuo98a, LP95, Les78, Les88b, LWL05, Lon77, Mar04, Mar06, May79, McK84, McS79, MGK^{+}03, MCM06, MK85b, MK90, Nor02, Oom13, PH98, PH99, Per77, PW77, Pow84b, Reh03, RAMP94, RS78, SAS88b, Sax85b, Sax85c, Sch93a, SS95a, Sch90, SNG95, SFE10, She93c, SRBB99, Smi99, SKP94, Stu78b, Sun86, TGB95, TP91, TR94, TD91, TECB03, VMKB05, Wei94, Wit01, Xia97, dG93, Ale96a, All86b, AC11, AJ05, BN01, BFS07, Cha97, CGMS89, DG00, GVK^{+}03, GMB93, Gro93, Gue19, Har11, Jou86, Kaz05, Kuo98b]. using [LW03, Les76b, Lum13, Mar01a, Mas05, MMC06, MCM07, NV06, PDG02, RÓ01, RHP^{+}07, Sch76b, SZ95, SB08, TG96, TH04, Com79, WRA05]. uso [VD01]. UT [USE84a, USE95a]. Utah [Sof84, USE95c, Hib96]. Utilities [Coc01c, FHS7x, KL94, Kuo98a, Lin79a, MKL^{+}95, MBL89, Ric98, X/O89b, Ame93, Com94a, Int93, JJ91u, MFS90, PM87, Ros84, ?, ?]. Utility [Rea91, Rez95, Rob75, San93]. utilizando [RÓ01]. Utilization [HIT^{+}02]. UTMOST [Ney83a, Ney83b]. Utopia [Ano02c]. Utopia-LVDS [Ano02c]. UTS [Wal82a, Wal82b, Wal82c]. UUCP [Che79a, Che79b, Che79c, HNR84, NHR84, OT88, RW86b, RW86a, TD91, TO92, Mor86, Now78, Now90, Tru83]. UUNAME [Che79b]. Uustat [Che79c]. UWIN [Kor97c, Kor97d]. UX [Kuo99, HP-93, Ano93n, Ano96f, CM04, Kee05, Ket99, Kul06, Lei96b, LMOT01, LMOT02, Mad97, Mea88, Olk03, Pon94, Pon96, Pon97b, Pon97c, Pon98, Pon99b, Pon99a, Pon01, Pon02b, Pon03, Pon05a, Reh00, Reh05, Ric93, Ric94c, Ric95, Rob00, SW99, SRW04, Tho03, TOML04, WLHC84]. üzemeltetése [Lás05].

V [?, ?, Ano00a, BBJ86, CBB^{+}91, Dav89, FHS91, Her93, KM87, Sob91, Ame90b, Ame90a, Ame91, Ano85, Arn86, BFGK89, BS88, Bas87, Ber86, Bou86, BB93, CQ83b, CWS90, Cof90, DRN^{+}23, Dun89, FG91, FHS91, GT92c, GC94a, GC94b, ILG19, Kri91, Lan84, OD^{+}86, OBSB23, Pad93, PS22, PGW^{+}20, PM87, Rag93, RRF90, Ros96, SGP^{+}23, SNM22, Sch90, She93a, Str90, Top90b, VTS^{+}23, Wes83, WMP84, WMP92, ZB19, Pik83b]. V-system [Ber86]. V.3 [Ste88]. V.42bis [Tho92]. v1.0 [Ano00e]. v1.2 [SP90]. v10.6 [KHG10]. v2 [TOML04]. v2.5.1 [NBB^{+}02]. v6.1 [van97]. Vacation [BBK13, CKT80]. Vade [Lip99]. Vader [Ree82a, Ree82b]. validate [Reg11]. Validation [FN83, Hol91, AKST93]. validations [DC17]. Value [BG97, CW89, CW91, Cou84, Cou85a].

Vanilla [Rai89]. Variable [Bey77, Koe78c, Sch76a]. Variables [KS85]. Variant [McI63]. Variants [MAP^{+}22, McK92c, Mor96, AOAH14]. variates [Wal96]. Variations [RW87]. Varying [Kau90, Sch76a]. VAX [Ber85b, Cap82, Cap87, Cod86, DJM86, FG80, Fat82, Gue87, HS83, HR79b, Gur81, JF81b, KTB^{+}88, KM79b, Kri84a, Kri84b, LR78, Mat80, GF80, Raf79, Rib89, SHLR80, SM85, Tuo82, KM83, May84]. VAX-11 [KM79b, LR78, HS83]. VAX-11/780 [KM79b, LR78]. VAX-UNIX [Gur81]. VAX/UNIX [FG80, Fat82, Mat80, GF80, Raf79, CT83]. VAX/VMS [Cap82, Cap87, Mat80, Raf79]. VAXIMA [FG80]. VB [Ano96d]. VCHK [Bry83a, Bry83b]. Vecotr [Bar00b]. Vector [Blu7 b, Blu78b, Bah09, Mil93]. Vendor [Ano92i]. Vendors [Gil96b, Tuc90, Ano96e, Ros94b]. venerable [Ano93m]. Verification [Nee90]. Verifier [HR73, Ryd74, Ryd72, Ryd3 ]. Verilog [Kro99, Pal96]. Verlag [Wei95]. Vers [HS83]. Versatile [SB02]. Versatility [WDG^{+}06]. Version [AL97c, Ano96c, Ano02c, BBJ86, BGJ01, Bec79, Bel79c, Bra79, CMR^{+}90, Daw86, FKN85, Fla91, GKM90, HS99, Hel90b, Hor79, KGP90, Luc03b, Lud78b, Mar82, Mar78f, The93, May78, NS94b, Nye88a, Nye88b, Nye90b, Nye90c, The02, OQL88, PCSF08, Pon05a, SPS95, SGFR92, SG95, Swa74, Wei84, XWS95, Ano96f, ACS98, BS02, Bou97, CS93, CS96b, Com12, DeR97, Env97a, Env97b, GSN90, HBMT96, HdBM75, IF92, Jos97, KFO97, Lew88, Mad94, Mas05, Mat80, Nat90, Nat94, Nye95, SAS90, SAS93b, SAS93c, SG97a, SG97b, TH04, WNDO97, W^{+}99, CE97, QO89]. Versionen [DF00]. Versions [AD89, MAP^{+}22, Mil00, Rob79a, Rob82, Hor90b, Sau17]. Versus [PLBL92, ASW^{+}01, Dav96, PM18, PH98, PH99, SGM^{+}00]. vertical [HK82b]. Vertigo [FM02]. Very [Höf01, LT80, O’D90, Sau17]. vfork [Kar83]. VI [Ano85, Sma88, USE92a, Rob99d, Ano00g, CoL92b, Han86, Hew90, Lam90, Rob99a, Son87, eL85, Smi06b]. Vi-Editor [Rob99d]. vi/ex [eL85]. via [CKV93, Fel94, LBN14, Liu11, Rob8 , Rob79b, Wal97b, Wal98]. Video [HK83, Mil91, Sta06, SW91b, Ran06]. Videotape [Ger82, Ros77a]. Vietnamese [Lun99]. View [Ama88, DS77, IJ98, O’D87, Stu78c, Vog77a, Wal92g]. Views [Bar00b, Bar00c, Bar01b, Coc01a, Coc01b, Coc01c, Coc03a, Coc03c, Coc03b, Sat93, Zie93]. vignettes [War08]. VII [USE93b]. VIII [USE94d]. Violations [HR79a]. Viral [Duf89c]. Virginia [MS05]. Virology [McI89]. Virtual [BJP04, BBHL08, Che87a, Cla88, Cox09, Fra92, GS12, Gol79, Ham78, KM86a, KAH83, Lut99, Lyc71, Lyc73, MK85a, Nor02, Pon02b, SMK^{+}94, TB05, Ano95f, CK06a, LCH^{+}81, MRM06, NV06, NZW^{+}02]. Virtualisierung [Spr06, Spr07]. Virtualization [Spr07, G^{+}06a, MRM06, Mly09, P^{+}08, VSC^{+}10]. virtuelle [WF03]. virtuellen [CK06a, CK06d, CK06c, CK06f, CK06e, CK06g, CK06h, CK06i]. Virtues [AGHR89a, AGHR89b, BTS09]. virus [McI89]. Viruses [Duf89a, Duf89b]. Visual [BCM88, BBM85, Kob91, MS95, Son87, Ano98c, RR99]. visualisation [Dia94]. Visualization [FKM95, The92c, The92a, The92b, The93, Wol93, SML06, ZGdDL^{+}96]. visualize [SML06]. Visualizing [Vog97]. Visually [Bel00]. VLIW [CCD^{+}87]. VLSI [BK85, CB83, SAL93]. VM [DuC94, G^{+}06a, Nor79, P^{+}08, TB05, Var91]. VM/CMS [Nor79]. VME [Isa86]. VMS [Ano91a, Bou90a, Bou90b, Bun91, Cap82, Cap87, Har90, Isa91, JLSG84, Mat93, Mat80, Raf79, Wal92g]. VMUNIX [Fer82]. VMware [CK06c, KGG00, War05, Bor01, CK06c, Com00, Com03, D^{+}04, MRM06, NL00, OH05, R^{+}02, War02, WF03]. Vnodes [Kle86]. Voice [Fil79, Law98]. VoiceGenie [Ano03d]. Voices [Pin92]. VoiceXML [Ano03d]. VoIP [Ano03f]. Vol [CS96b, GL88, VJ06, Mad94]. Volker [Len97]. Volume [Bel79c, Buy99, CHL16, HA95a, HA95b, HA96, Ker79g, Mil84c, Ste94, Ste96, WS95, ZGdDL^{+}96]. Voluntary [HG95b]. vom [SU94a, SE96]. vorbereiten [DF00]. Vorstellung [CK06b, CK06d, CK06c, CK06h]. Vote [Mei84]. Vous [WB85a, WB85b]. VPN [Ano98c]. VPNs [NZW^{+}02]. VR [GVK^{+}03]. VR-Based [GVK^{+}03]. VRML [Ano96d]. vs [Jud17, Sim96c]. VSP [FY18]. VTK [SML06]. vulnerabilities [DLCV17]. Vulnerability [Ano99d, Gar97, LZAH23, Ste02a]. vulnerable [DLCV17].

W [Sau88, Teo13, Tho05]. W. [Luu90].

W3C [Bar01b, Coc01a]. wa [JK86, ACM93, IEE94c]. Waite [ACH91, Ber92, And95]. waiting [Sch95c]. Wakeup [PPTHxx]. walk [Lev83]. WAN [Höf01]. WAN-Cluster [Höf01]. Wanna [Zin92]. want [O’D83a, SF95]. wanta [Gra88]. Wanted [Gra03, Wal95a]. Wants [Law98]. ware [Sta05b]. Warning [Gar97]. Wars [Law98]. was [Mud97, Ste99a]. Washington [USE84b, USE87e, Ass88c, USE90d, USE97a, USE97d, USE98c, USE98a, USE99b, USE99a]. Watch [Ano95e]. Watchdogs [BP88b, BP88c, BP88a]. Wave [Ano00e, ALP^{+}21, Ano03e]. Way [Ano00d, Ano00f, Bea04, BM92, CPJ^{+}98, McC95, SWG92, Com00, EP05]. Ways [Nis01b, Strxx]. WCL [Hen92]. Weakness [Mor85]. Web [Luc03b, Ano02b, Ano03c, Ano03e, Aul01, BA96, But02b, CH96, Dav00, DuB01, Fel96a, FWS^{+}03, Gab07, G^{+}06b, Har01c, Hun01, JB99, Kro99, Law98, LW03, Mac97a, McG97, Moh99, MH96, SWL^{+}01, Sla94, Sta04, TSC02, Tho95, VSW94, Web95, WT01, WLS^{+}01, Wid00, BKM86]. Web-Based [Mac97a]. Web-Savvy [Kro99]. Web/RCS [Dav00]. WebCompiler [Kro99]. Weber [Sch04]. Webifying [Wee99]. Webmaster [Ano96f]. Webmin [Cam04, Coo03]. WebSphere [AJ05]. Webster [MT74]. weekend [BW02]. weight [BS90]. Welcome [SG95]. Well [Pov99]. Well-Formed [Pov99]. Wells [Ano97b]. Welsh [Ano95b]. WEP [Coc01c]. werden [Die11]. Werkzeuge [Sta89]. Werkzeugen [KGG00]. Western [DD05]. Wharf [IEE03]. Where [Dec04, O’D83a, And17, Ano94n]. Which [Bak75, Ano93r, FJPR93, Web95, Ale96b]. While [Gil96a]. Whippany [Ral77]. White [Ste01f]. Whither [Lev88, Mei84]. Who [Sem96, Ano93r]. Whole [Kro92]. wichtigsten [CK06b]. Wide [Adh95, C^{+}78b, Tho95, VSW94, CDD^{+}05, TWM86a, TWM86b]. widely [BY96]. Widget [BAKW92]. Widgets [McM92, McM93]. Width [Hob89]. Wild [Cox10]. Wildcards [Skj82]. Will [Ano92m, BED^{+}85, FES^{+}99, JC92, OH97]. Wilson [Pol93]. Wily [CB94, CBR03]. Win32 [Dal99, Dal02, Mer01, Mon00a, Mon00b, Noe98, PBS98, Wal97a]. Winchester [TBGE89]. Windmill [TBGE89]. Window [XWS95, AS90, ACS98, Bar91b, Bar94, BNB87, CJL^{+}89, Car89b, Cra90, CGO92, DKF94, DD90, Fla91, GMM90, Get86, GKM90, GRA89a, Hel90b, IF92, Jac84a, Jac84b, JR93, Joh92, Jon89, Ken90, Kil96, MF83, Man91, Man93a, Man93b, Mik92, MP92, Nat90, Nat94, New92, Nye90b, Nye90c, Nye95, OQL88, PS94, Par93, PH98, PH99, Pik88, Pik89, Pikxx, QO89, QO90, QO91, QO93a, QO93b, SGN88, SG90, SGFR90, SGFR92, SG97a, SG97b, Smi91, Sou94, Swi91, TP91, Tra85, Uhl87, Wri90, X/O88d, You89, You90b, YP92, ALP^{+}21, Hek89, Sch90, Har93, Hiu96, Hiu97, JR89, JR90, JR92, Kel90, Mik90, RR92]. Windowing [Hol87a, Hol87b, KMKO92, ST89a, ST89b, ST89c]. Windows [Ano95d, Ano95e, Ano98c, Bor01, DF00, Get91, Gla97, Koh06, MRxxd, PKP05, SAS94b, SL95, SML06, USE97a, USE97d, USE98c, USE98a, USE99a, WF03, Wei04, AC11, Ano93-43, Ano02a, Ano03e, Aud96, Bal96, Bar00b, Bia99, BT98, Bor98, BM02, Bri05b, Bru02, Chi02, CK06a, CK06f, CK06e, CK06i, Col84a, Cus92, Cus93, Ded09, Don95, Env97a, Env97b, Fos96, FJPR93, Fre98, GBG97, Har93, Har00, HL85a, HL85b, HH04, HD98, IBM09a, IBM09b, Jan03, Jon98, KVCS98, Kor97b, Kor97c, Kor97d, Kor97a, Kos96, Lei00, Li98, Lin93, LBSL97, Mat97, McC01, Mei92, Mer97, Mil98, Mil00, Mil01, MYH00, MY03, MB05, NH99, O^{+}98, Ass99a, Pac00, Pan94, PH98, PH99, Pon97c, Raj98, RT93, SAS94b, SK10a, SF95, SS95a, SEZ^{+}01, Sha00a, Smi05b, Spi94]. Windows [Spi00, Sta07a, TF04, Var96b, Wal97b, Wal97a, Wal98, War01, WG98, Wit01, Wol04a, WS96, Zim98, ZG96]. Windows-Programmierung [PKP05]. Windows-to-Linux [Wol04a]. WINDOWS^{T M} [Ni97]. Windoze [CPJ^{+}98]. WindRiver [Ano95j]. WINDX [Col84a]. wine [OS06]. winning [Tuc90]. WinNT [NH00]. Winter [USE83, Ass83, USE85c, USE85b, USE86e, USE86b, USE87e, Ass89b, USE89f, USE89b, USE90d, USE91d, USE91f, USE92d, USE93f, USE94a, USE94f]. WIPO [Gar97]. Wired [Coc01c]. Wisconsin [MN69, Ric77]. Within [BCGS91, MMTW88, Ral75a, MKSB22, UNI93]. Without [CCM^{+}90, Pan88, VB92, EP05, HCC^{+}87, Len97, Loz05b, VB94]. wonglediff [EP05]. Word [Daw86, IBC87, Les88b, Lud77a, Sha79a]. Word-Processing [IBC87]. WordPerfect [Smi99]. Words [Bro90, Fre78, BKM86, MC75a]. Work [ASS79, Asb00, Gra79b, HCF^{+}94, Har92a, Ker85, Lud79b, Smi06a, Smi06c, Ste01d, Wal95e, Zim98, Hus92, SZ95, Ano93p]. workaround [Len97]. Workarounds [HSM03]. Workbench [DH77, MS77, SC78, Smi83, Ama91, FF88, D^{+}76, DHM78, Fra83, Gin83, Lyo76b, Mac83, Mas77a, Roo79b]. Workbenches [Ano93w]. Workbook [Jae89a, Kap99, McM99, Moh99, Pat99, Vic98, Env97a, Env97b, PWC00]. worked [Cox12]. Workgroups [Pan94]. Working [Kar92, Ket99, Lut99, Spi05]. Workloads [HKS^{+}04, BBB^{+}20]. Workplace [HHJC99]. Workplan [Hal79, Lud79a]. Works [Bar01b, Win75, War04, War15]. Workshop [Ano88b, IEE94c, MM06a, SD95, USE86d, USE87b, USE87c, Ass88c, USE88d, USE88e, USE88a, USE88b, USE88f, USE89c, USE89d, USE89g, USE92c, USE97a, USE97d, USE99c, WS78, HTU96, USE89h, USE90e]. Workstation [Ano00e, Ano03g, BBNS80, BBP82, Bor01, CK85, Kad95b, McD84a, McD84b, Mue94, Per93, PJ89a, PJ89b, RR03a, Rus91, SU94a, Str94, SU94b, SE96, Sun86, WB96, WF03, War05, SE97, SW91b, War02]. Workstations [BS91, Ear87, Lei95a, Lob85, The93, Mei92, Ric90, Tre88, BGR89, Lei90, MAT08, The92c, The92a, The92b]. World [Ano97a, BR99, BMR01, CPJ^{+}98, Das88, Fin91, Gal95, Gia90, Isa91, O’D83c, O’D87, Spi00, TBS87, Ano95e, BMR82, Cor05b, Hal06, Hal07, PT93, Sal08, SG95, Smi05b, Sta07b, HD98, Tho95, VSW94]. worrying [Coo03]. Worth [Bro90, Gra99d]. wrapper [Sch93b]. Wrappers [Rez95, JGS^{+}08]. wrestling [SZ95]. Write [Hoo05, Jen00, LH01b, LEG88, Len87, Ano93r, SM88a, SM88b, SM88c]. Write-Once [LEG88]. Writer [Fra83, Gin83, Lew01, Mac83, Mar78f, Smi83]. Writes [MG99]. Writing [All86b, Bur78, CV80, Coo09c, Coo09d, Cra88, Dal99, Dal02, Don97c, ET88, ET92, Hol87c, McC90, McI72, Paj88, Paj92, She01, Shi92, Wei93, Wel94a, YT83, BGR89, HM98, Lew91]. Written [HMP83a, HMP83b, Jen75, Hol82]. WSDL [Bar01b]. WSH [BBM84]. Wu [Luu90]. Wux [Spi94]. WVtbl [MD86]. www.deliberate.com [Ano98c]. WWW2GCG [CH96].

X

[Ano00e, Cra90, Dav06, EBS08, Har01c, JR04, JR05, May01, MNC09, Pog07, Tay05, Top05, TSM^{+}06, USE96a, Ado93, And06, AS93b, XWS95, AS90, ACS98, Aud96, Bar91a, Bar91b, Bar94, BS02, BAKW92, Buc01, CGO92, DH05, DKF94, DD90, EBS08, Fei01, Fei02, Fla91, GMM90, Gay93, GL06, Ger02, GSN90, GKM90, Get91, GM11, Hal11, Har93, Hel90b, Hes94, Hil02, Hil08, Hiu96, Hiu97, IF92, JR02, JRR08, JR89, JR90, JR92, JR93, Joh92, Jon89, Kar92, Kau93b, Kel90, Ken90, Kil93, Kil96, Kim95, KHG10, KMV^{+}09, KS08, Lan01, Lee92, Man91, Man93a, Man93b, MGMMI10, May01, McC90, McE05, MTS03, Mik90, Mik92, MP92, MNC09, Nat90, Nat94, New92, Nye90b, Nye90a, Nye90c, NO90a, NO90b, Nye92a, NO92, NO93, Nye95, OQL88, O’R90b, PS94]. X [Par93, PH98, PH99, PB09, Pog02a, Pog02b, Pog05, QO89, QO90, QO91, QO93a, QO93b, RR02a, RR03a, Red97, Rei93, RR92, Rit88b, Rit88a, Ros06, Ros90, Ros93b, RJR08, SGN88, Sch89, SG90, SGFR90, SGFR92, SG97a, SG97b, Sin07, Smi91, Smi92, SS06, Sou94, Swi91, TPTS02, TJ03, TJ04, TP91, Tho06b, Top02a, Top02b, TECB03, Wal95a, Wan96, Wri90, Ygg94, You89, You90b, YP92]. X-based [Hes94]. X-MP [Rit88a]. X-Toolkit [Kim95]. X-Window [Hiu96, Hiu97, IF92]. X-Windows [Mei92]. X.25 [Erm79, Mil84a]. X.400 [DN87]. X.deskhelp [Dav93a]. X.desktop [BM92]. X/Motif [Bar91a]. X/MP [Rit88b]. X/Open [Wal95a, X/O88c, X/O88a, X/O88d, X/O88b, X/O89b, X/O89c, X/O89a]. X11 [AKDN90, MP92, NO90b, Wid90]. X3.159 [Ame89]. X3.159-1989 [Ame89]. X86 [RFS06, RFS05]. Xandros [Gra05]. Xcode [CC07, And06, And09, Buc06, PB09]. XDE [Fra99]. Xen [Spr06, Spr07]. XENIX [BBJ86, Ame90a, Fie88, YEB84, Gre81, Top90b]. XFree86 [DF00]. XFree86/23.3.6 [DF00]. xhelp [Dav93a]. XI [USE97b, RB08]. XII [USE98d]. Xilinx [Ano02c]. Xinu [Com12]. XIV [USE00b]. XLFD [SGFR90, SGFR92]. Xlib [GSN90, Kil94, Nye88a, Nye88b, Nye90b, Nye90c, Nye92b, SGFR90, SGFR92]. XML [Ano00b, Ano01, Gar99, YLM^{+}05]. Xmt [Fla94]. XP [Woe03]. XScale [Bar00b]. xSeries [R^{+}02]. XSI [X/O89b, X/O89c]. Xt [Kel90, Smi92, You89, You90b, YP92]. Xt/Motif [Smi92]. XTI [Ste98]. Xue [RB08]. xv6 [CKM14, CKM06]. XView [Hel90a, Hel90b, Hel91b, Van91].

Y2K [Kie99, Sta07a]. Yacc [Joh79c, Sch90, DS88, GO86, Joh78c, van83, Joh75, JS90, Sta89, LMB92, MB90]. YACC-compatible [DS88]. YADS [Roc79]. YAPP [Ste01g]. Year [Ora21, Sta07a, SHVP05]. Years [FK90, Lyc85, MFBWW20, Sal97, Sal99, Sal00a, Sal00b, Sal03, Ano19a, McK85a, PTS^{+}11, Woo94]. Yellow [WBS01]. Yggdrasil [Ygg94]. York [ACM03]. Yosemite [Ano00e]. you’ll [Pub11, Vib11]. You’re [Coc01b, Tay96]. Yourself [CS99, MB00, RBHB99, Spe96, Art99, Bal99, CCB00, Gay99, Hsi01, LH00, Ray99a, Reg11, SBR01, Tay96, TA98, Tay03, Tay06, Til95, UT03, WT99, Wel99a]. You’ve [Arn91, Wal95a, Sch95c]. Yunikkusu [JK86].

Z [May01, Bec79, Dan83a, Dan83b, FHS78a, Fox79, Has78, May01, G^{+}06a, P^{+}08, TB05, WMK^{+}07]. Z-Shell [May01]. z/VM [G^{+}06a, P^{+}08, TB05]. z15 [BBB^{+}20]. Z80 [BSW78, Far78, Rub79, Vog78a]. z9 [ZST^{+}07]. zero [HMK^{+}20, Ano02b]. zero-day [HMK^{+}20]. Zertifizierung [Mag04]. ZFS [DM07, Daw07, JL16, Jud17, LJ15, Wat10]. Zhi [RB08]. Zilog [Far78]. Zipf [MSSvK08]. Ziv [NT05]. zSeries [BS05]. ZSORT [Zuc77a]. zum [Die11, Kuo99, Mag04]. zur [Jor04, KGG00, Kuo99, Luc03b, Sta89, SU94a, SE96].

References

Anderson:1986:UCS

[AA86] Gail Anderson and Paul Anderson. The UNIX C Shell Field Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-937468-X. xxi + 374 pp. LCCN QA76.76.O63 A6 1986. The C-Shell bible — everything you need to know to use UNIX.

Appavoo:2005:EKO

[AAB^{+}05] J. Appavoo, M. Auslander, M. Butrico, D. da Silva, O. Krieger, M. Mergen, M. Ostrowski, B. Rosenburg, R. W. Wisniewski, and J. Xenidis. Experiences with K42, an opensource, Linux-compatible, scalable operating-system kernel. IBM Systems Journal, 44(2): 427–440, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                                    8670.  URL http://www.
                                                    research.ibm.com/journal/
                                                    sj/442/appavoo.html;   http:
                                                    //www.research.ibm.com/journal/
                                                    sj/442/appavoo.pdf;   http:

//www.research.ibm.com/journal/
sj/442/appavoo.txt.

Atlidakis:2016:PBO

[AAG^{+}16] Vaggelis Atlidakis, Jeremy Andrus, Roxana Geambasu, Dimitris Mitropoulos, and Jason Nieh. POSIX has become outdated. ;login: the USENIX Association newsletter, 41(3):6–12, Fall 2016. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/system/files/
login/articles/login_fall16_
02_atlidakis.pdf.

Arthur:1997:USP

[AB97] Lowell Jay Arthur and Ted Burns. UNIX shell programming. Wiley, New York, NY, USA, fourth edition, 1997. ISBN 0-471-16894-7 (paper). x + 518 pp. LCCN QA76.76.O63 A76 1997.

Accetta:1986:MAN

[ABB^{+}86a] Mike Accetta, Robert Baron, William Bolosky, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young. Mach: a new kernel foundation for UNIX development. In USENIX Association [USE86c], pages 93–112.

Accetta:1986:MNKb

[ABB^{+}86b] Mike Accetta, Robert Baron, William Bolosky, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young.

Mach: a new kernel foundation for UNIX development. In USENIX [USE86a], pages 93– 112.

Abbott:2003:LER

[Abb03]  Doug Abbott.  Linux for

Embedded and Real-Time Applications. Embedded technology series. Newnes Press, Amsterdam, The Netherlands, 2003. ISBN 0-7506-7546-2 (paperback). xii + 255 pp. LCCN QA76.76.O63 A24 2003. US$49.95. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/els041/2003277020.
html;   http://www.loc.gov/
catdir/toc/els041/2003277020.
html.

Accetta:1986:MNKa

[ABG^{+}86] Mike Accetta, Robert Baron, David Golub, Richard Rashid, Avadis Tevanian, and Michael Young. MACH: a new kernel foundation for UNIX development. Technical report, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1986. 18 pp.

URL https://www.cs.ubc.
ca/~norm/508/2009W1/mach_
usenix86.pdf.

Ables:1993:BSU

[Abl93] R. King Ables. Backups with standard UNIX commands. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):6–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Andersen:1995:ECI

[ABS95] Paul K. Andersen, G. Bjedov, and M. G. Scarbrough. Essential C: an introduction for scientists and engineers. Saunders College Pub., Fort Worth, TX, USA, 1995. ISBN 0-03-004158-9. xiv + 370 pp. LCCN QA76.73.C15 A475

1995. URL http://www.loc.
gov/catdir/enhancements/
fy0640/94067128-d.html;
http://www.loc.gov/catdir/
enhancements/fy0640/94067128-
t.html.

Aho:1975:ESM

[AC75] Alfred V. Aho and Margaret J.

Corasick. Efficient string matching: an aid to bibliographic search. Communications of the Association for Computing Machinery, 18(6): 333–340, June 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Altheide:2011:DFO

[AC11] Cory Altheide and Harlan A.

Carvey. Digital forensics with open source tools: using open source platform tools for performing computer forensics on target systems: Windows, Mac, Linux, UNIX, etc. Syngress Publishing, Inc., Rockland, MA, USA, 2011. ISBN 1-59749-586-7 (paperback), 1-59749-587-5 (e-book). xvii + 264 pp. LCCN HV8079.C65 A474 2011. URL http:

                                        //www.sciencedirect.com/
                                        science/book/9781597495868.

Appelbe:1986:PUN

[ACF^{+}86] W. Appelbe, D. Coleman, A. Fratkin, J. Hutchison, and W. J. Savitch. Porting UNIX to a network of diskless micros or UNIX on tinfoil. In USENIX Association [USE86c], pages 486–498.

Anderson:1991:WGU

[ACH91] Bart Anderson, Barry Costales, and Harry Henderson. The Waite Group’s UNIX Communications. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1991. ISBN 0-672-22773-8. xxiv + 736 pp. LCCN QA76.76.O63 A5 1991. Highly recommended. A really excellent book on elm, rn and

                                        nn.

Andrade:1989:BTP

[ACK89] Juan M. Andrade, Mark T. Carges, and Kurt R. Kovach. Building a transaction processing system on UNIX systems. In USENIX Association [USE89g], pages 13–22. ISBN ???? LCCN ????

ACM:1975:CRS

[ACM75]  ACM, editor.  Conference

record of Seventh Annual ACM Symposium on Theory of Computing: papers presented at the Symposium, Albuquerque, New Mexico, May 5–May 7, 1975. ACM Press, New York, NY, USA, 1975. LCCN QA76.6 .A16 1975.

ACM:1979:PSC

[ACM79] ACM, editor. Proceedings of the SIGNUM Conference on the Programming Environment for Development of Numerical Software. ACM Press, New York, NY, USA, 1979.

ACM:1982:PSS

[ACM82] ACM, editor. Proceedings of the SIGPLAN ’82 Symposium on Compiler Construction: papers presented at the Symposium, Boston, Massachusetts, June 23–25, 1982, volume 17(6) of SIGPLAN notices. ACM Press, New York, NY, USA, June 1982. ISBN 0-89791-074-5. LCCN QA76.7 .A8 v.17, no.6. URL https://dl.

acm.org/doi/proceedings/
10.1145/800230.

ACM:1988:PAS

[ACM88] ACM, editor. Proceedings of the ACM SIGGRAPH Symposium on User Interface Software, Banff, Alberta, Canada, October 17–19, 1988. ACM Press, New York, NY, USA, 1988. ISBN 0-89791-283-7. LCCN QA76.9.U83A26 1988.

ACM:1989:PSN

[ACM89] ACM, editor. Proceedings, Supercomputing ’89: November 13–17, 1989, Reno, Nevada. ACM Press, New York, NY, USA, 1989. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

ACM:1992:PAC

[ACM92]  ACM, editor.  Proceedings

of the 1992 ACM Conference on LISP and Functional Programming: papers presented at the conference, San Francisco, California, June 22–24, 1992. ACM Press, New York, NY, USA, 1992. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

ACM:1993:TCS

[ACM93]  ACM, editor.  TRI-Ada

’93: Conference — September 1993, Seattle, WA, TRIADA - proceedings- 1993. ACM Press, New York, NY, USA, 1993. ISBN 0-89791-621-2. LCCN ???? ACM Order No. 825930.

ACM:1997:PEA

[ACM97]  ACM, editor.  Proceedings

of the Eighth Annual ACM-SIAM Symposium on Discrete Algorithms, New Orleans, Louisiana, January 5– 7, 1997. ACM Press, New York, NY, USA, 1997. CO-DEN PAAAF2. ISBN 0-89871-390-0. LCCN ???? URL

                                        http://www.acm.org/pubs/
                                        contents/proceedings/soda/
                                        314161/.

ACM:2003:SPA

[ACM03]  ACM, editor.  SOSP ’03:

proceedings of the 19th ACM Symposium on Operating Systems Principles: the Sagamore, Bolton Landing, Lake George, New York, USA, October 19– 22, 2003, volume 37(5) of Operating systems review. ACM Press, New York, NY, USA, December 2003. ISBN 1-58113-757-5. ISSN 0163-5980 (print), 1943-586X (electronic). URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://uclibs.org/PID/34720.

ACM order number 534030.

ACM:2004:SHP

[ACM04] ACM, editor. SC 2004: High Performance Computing, Networking and Storage: Bridging communities: Proceedings of the IEEE/ACM Supercomputing 2004 Conference, Pittsburgh, PA, November 6–12, 2004. ACM Press and IEEE Computer Society Press, New York, NY, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2153-3. LCCN ????

Arthurs:1979:PRT

[ACS79] E. Arthurs, D. W. Cragun, and B. W. Stuck. Preliminary Real Time Listing Retrieval (RTLR) field study performance analysis. Technical Memorandum 1415, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 9, 1979. ?? pp.

Asente:1998:XWS

[ACS98] Paul J. Asente, Donna Converse, and Ralph R. Swick. X Window System toolkit: the complete programmer’s guide and specification: X version 11, releases 6 and 6.1. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1998. ISBN 1-55558-178-1. xxxii + 1131 pp. LCCN QA76.76.W56A74 1998.

Anderson:1989:ITI

[AD89] E. Anderson and J. Dongarra. Installing and testing the initial release of LA-PACK — Unix and non-Unix versions. LAPACK Working Note 10, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, May 1989. URL http:

                                        //www.netlib.org/lapack/
                                        lawns/lawn10.ps;   http:
                                        //www.netlib.org/lapack/
                                        lawnspdf/lawn10.pdf. ANL,

MCS-TM-130, May 1989.

Atkinson:1989:ECP

[ADH^{+}89] Russ Atkinson, Alan Demers, Carl Hauser, Christian Jacobi, Peter Kessler, and Mark Weiser. Experiences creating a portable Cedar. ACM SIGPLAN Notices, 24(7):322–329, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        pldi/73141/p322-atkinson/.

Adhikari:1995:OWS

[Adh95] Richard Adhikari. Open wide and say ‘POSIX’. Software magazine, 15(9):122–??, September 1, 1995. CODEN SMWMEQ. ISSN 0279-9782 (or 0897-8085??).

Arthurs:1978:CUP

[ADJS78] E. Arthurs, P. Daily, R. Jakubek, and B. W. Stuck. CNCC UNIX project performance measurements, projections and enhancements. Technical Memorandum 1253 (MF 78-8234-10), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 24, 1978. ?? pp.

Adolph:1989:HAU

[Ado89] W. Stephen Adolph. High availability in a UNIX transaction processing environment. In USENIX Association [USE89g], pages 23–32. ISBN ???? LCCN ????

Adobe:1993:PDP

[Ado93] Adobe Systems Incorporated.

Programming the Display PostScript System with X. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-62203-3. LCCN QA76.73.P67 D57 1993. US$29.95.

Aguirre:1991:EMD

[AEG^{+}91] G. Aguirre, M. Errecalde, R. Guerrero, C. Kavka, G. Leguizamon, M. Printista, and R. Gallard. Experiencing Minix as a didactical aid for operating systems courses. Operating Systems Review, 25(3):32–39, July 1991. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Afzal:2003:UUB

[Afz03] Amir Afzal. UNIX unbounded: a beginning approach. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2003. ISBN 0-13-092736-8. xviii + 494 pp. LCCN QA76.76.O63 A366 2003.

Armand:1989:RDU

[AGHR89a] Francois Armand, Michel Gien, Frederic Herrmann, and Marc Rozier. Revolution 89 or “Distributing UNIX brings it back to its original virtues”. In USENIX Association [USE89c], pages 153–174. ISBN ???? LCCN ????

Armand:1989:RUB

[AGHR89b] Francois Armand, Michel Gien, Frederic Herrmann, and Marc Rozier. Revolution 89 or “Distributing UNIX brings it back to its original virtues”. In USENIX Association [USE89c], pages 153–174. ISBN ???? LCCN ????

Albinson:1992:ULC

[AGP^{+}92] Lawrence Albinson, Dominique Grabas, Pascal Piovesan, Michel Tombroff, Christian Tricot, and Hossein Yassaie. UNIX on a loosely coupled architecture: The CHORUS/MiX approach. Future Generation Computer Systems, 8(1–3):67–81, July 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Agrawal:1990:OUF

[AGS90] R. Agrawal, N. H. Gehani, and J. Srinivasan. OdeView. A userfriendly graphical interface to Ode. SIGMOD Record (ACM Special Interest Group on Management of Data), 19(2):389, June 1990. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Adelstein:2004:EJL

[AH04] Tom Adelstein and Sam Hiser.

Exploring the JDS Linux Desktop. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00752-3 (paperback). xiv + 390 pp. LCCN QA76.76.O63 A353 2004. US$34.95, CAN$50.95, UK£24.95. URL http:/

/www.oreilly.com/catalog/
9780596007522;   http://
www.oreilly.com/catalog/
jds/index.html.

Ali:2011:BAS

[AH11] Shakeel Ali and Tedi Heriyanto.

BackTrack 4: Assuring Security by Penetration Testing. Packt Publishing, Birmingham, UK, 2011. ISBN 1-84951-394-5. 392 (est.) pp. LCCN ????

Andrus:2014:CNE

[AHA^{+}14] Jeremy Andrus, Alexander Van’t Hof, Naser AlDuaij, Christoffer Dall, Nicolas Viennot, and Jason Nieh. Cider: native execution of iOS apps on Android. ACM SIGARCH Computer Architecture News, 42(1):

367–382, March 2014. CO-DEN CANED2. ISSN 0163-5964 (print), 1943-5851 (electronic).

Aho:1984:USF

[Aho84] A. V. Aho. The UNIX system: Foreword. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1573–1576, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Aho:2012:TDR

[Aho12] Al Aho. Tribute to Dennis Ritchie. Alcatel-Lucent 23-minute video., September 11, 2012. URL https:/

                                        /www.youtube.com/watch?v=

GfoSbffSIQ4. From an event entitled, “The Lasting Legacy of Dennis Ritchie: The Impact of Software on Society.”.

Arthurs:1979:PAC

[AHPS79] E. Arthurs, J. L. Haase, C. Perea, and B. W. Stuck. Performance assessment of Customer Service and Administrative Control Center System (CSACCS) …. Technical Memorandum 1450, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 25, 1979. ?? pp.

Alexandrov:1998:UPG

[AISS98] Albert D. Alexandrov, Maximilian Ibel, Klaus E. Schauser, and Chris J. Scheiman. Ufo: a personal global file system based on user-level extensions to the operating system. ACM Transactions on Computer Systems, 16(3):207– 233, August 1998. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1998-16-3/p207-alexandrov/

.

IBM:1991:AIM

[AIX91] /AIXtra: IBM’s Magazine for AIX Professionals, 1991. IBM Corporation, MS 01-04-60, 5 West Kirkwood Blvd, Roanoke, TX 76299, USA.

Aho:1974:LP

[AJ74] A. V. Aho and S. C. Johnson.

LR parsing. ACM Computing Surveys, 6(2):99–124, June 1974. CODEN CMSVAN. ISSN 0010-4892.

Aho:1975:OCG

[AJ75] A. V. Aho and S. C. Johnson. Optimal code generation for expression trees. In ACM [ACM75], pages 207–217. LCCN QA76.6 .A16 1975.

Anderson:2002:UU

[AJ^{+}02] Robin Anderson, Andy Johnston, et al. Unix unleashed. SAMS Publishing, Indianapolis, IN, USA, fourth edition, 2002. ISBN 0-672-32251-X. xxv + 1163 pp. LCCN QA76.76.O63 U5587 2002.

Anand:2005:MPC

[AJ05] V. K. Anand and W. C. Jamison. A middleware performance characterization of Linux using IBM WebSphere Application Server. IBM Systems Journal, 44(2):353–367, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/anand.html;   http:
//www.research.ibm.com/journal/
sj/442/anand.pdf;   http:
//www.research.ibm.com/journal/
sj/442/anand.txt.

Aho:1975:DPA

[AJU75] Alfred V. Aho, Stephen C. Johnson, and Jeffrey D. Ullman. Deterministic parsing of ambiguous grammars. Communications of the Association for Computing Machinery, 18 (8):441–452, August 1975. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anley:2007:SHD

[AK^{+}07] Chris Anley, Jack Koziol, et al., editors. The shellcoder’s handbook: discovering and exploiting security holes. Wiley, New York, NY, USA, second edition, 2007. ISBN 0-470-08023-X (paper/website). xxiv + 718 pp. LCCN QA76.9.A25 S464

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0739/2007021079-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0739/2007021079-
d.html;   http://www.loc.
gov/catdir/toc/ecip0718/
2007021079.html.

Alexander:1985:TBP

[AKB85] Cedell A. Alexander, William M.

Keshlear, and Faye Briggs. Translation buffer performance in a UNIX environment. ACM SIGARCH Computer Architecture News, 13(5):2–14, December 1985. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Alexander:1986:CMP

[AKCB86] Cedell Alexander, William Keshlear, Furrokh Cooper, and Faye Briggs. Cache memory performance in a Unix environment. ACM SIGARCH Computer Architecture News, 14(3): 41–61, June 1986. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Angebranndt:1990:XSS

[AKDN90] S. Angebranndt, P. Karlton, R. Drewry, and T. Newman. The X11 sample server. Software—Practice and Experience, 20(S2):69–81, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Akin:1992:APO

[Aki92] Allen Akin. Analysis of PEX 5.1 and OpenGL 1.0. Technical report, Silicon Graphics Computer Systems, Mountain View, CA, USA, August 3 1992.

URL  ftp://sgigate.sgi.
com/pub/opengl/doc/analysis.

Z. The manual pages for the OpenGL API, its Utility Library (GLU), and the X server extension API (GLX) and a PostScript version of the OpenGL specification are available via anonymous ftp to

sgigate.sgi.com from /pub/
opengl/doc.

Arakawa:1993:MVR

[AKST93] Hiroshi Arakawa, Daniel I. Katcher, Jay K. Strosnider, and Hideyuki Tokuda. Modeling and validation of the real-time Mach scheduler. ACM SIG-METRICS Performance Evaluation Review, 21(1):195–206, June 1993. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Aho:1978:APSa

[AKW78] A. V. Aho, Brian W. Kernighan, and Peter J. Weinberger. Awk — a pattern scanning and processing language. Technical Memorandum 1206 (TM 78-1271-12, TM 78-1273-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1, 1978. ?? pp.

Aho:1979:APS

[AKW79] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. Awk — a pattern scanning and processing language. Software—Practice and Experience, 9(4):267–279, April 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL https:/

/onlinelibrary.wiley.com/
doi/abs/10.1002/spe.4380090403.

Aho:1985:APS

[AKW85] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. Awk — a pattern scanning and processing language programmer’s manual. Computing Science Technical Report 118, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 5, 1985. ii ++ 38 pp.

Aho:1988:APL

[AKW88] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. The AWK Programming Language. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-07981-X. x + 210 pp. LCCN QA76.73.A95 A35 1988.

Aho:1978:APSb

[AKW8 ] Alfred V. Aho, Brian W. Kernighan, and Peter J. Weinberger. Awk — a pattern scanning and processing language. Computing Science Technical Report 68, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1978 (??). ?? pp.

Aho:1989:PGA

[AKWA89] Alfred V. Aho, Brian W.

Kernighan, Peter J. Weinberger, and Takanori Adachi. Puroguramingu gengo AWK. (Japanese) [Programming language AWK]. Toppan, Tokyo, Japan, 1989. ISBN 4-8101-8008-5. xvii + 299 pp. LCCN ????

Abrahams:UI91

[AL91] Paul W. Abrahams and Bruce R. Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-55703-7. xxvii + 559 pp. LCCN QA76.76.O63 A27 1992. Prepared with TEX.

Abrahams:1992:UI

[AL92a] Paul W. Abrahams and Bruce R. Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-55703-7. xxvii + 559 pp. LCCN QA76.76.O63 A27 1992. Excellent, and thorough, coverage of UNIX, with chapters on the file system, utilities, shells, editors, Emacs, data manipulation, mail, network communications and resources, the X Window System, and a comparison of MS-DOS and UNIX.

Albitz:1992:DBN

[AL92b] Paul Albitz and Cricket Liu. DNS and BIND in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-010-4. xviii + 381 pp. LCCN TK5105.875.I57 A43 1992. US$29.95.

Abrahams:1995:UIa

[AL95a] Paul W. Abrahams and Bruce R. Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, second edition, 1995. ISBN 0-201-60965-7. 704 pp. LCCN QA76.76.O63 A27 1995.

Abrahams:1995:UIb

[AL95b] Paul W. Abrahams and Bruce R.

Larson. UNIX for the Impatient. Addison-Wesley, Reading, MA, USA, second edition, 1995. ISBN 0-201-82376-4. xxxv + 824 pp. LCCN QA76.76.O63 A27 1996.

Abrahams:1997:UHI

[AL97a] Paul W. Abrahams and Bruce R.

Larson. UNIX for the Hyper-Impatient. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-41991-2. ???? pp. LCCN ???? US$29.00. Entire book comes on one CD-ROM.

Abrahams:1997:UI

[AL97b] Paul W. Abrahams and Bruce R.

Larson. UNIX for the impatient. Addison-Wesley, Reading, MA, USA, second, CD-ROM version edition, 1997. ISBN 0-201-41992-0 (CD-ROM), 0-201-41979-3 (Book/CD-ROM). xxxvi + 824 pp. LCCN QA76.76.O63 A27 1997.

Abrahams:1997:UIC

[AL97c] Paul W. Abrahams and Bruce R.

Larson. UNIX for the Impatient: CD-ROM Version. Addison-Wesley, Reading, MA, USA, second edition, 1997. ISBN 0-201-41979-3. xxxvi + 824 + CD-31 pp. LCCN QA76.76.O63 A27 1992. US$49.00.

Aho:19xx:EAC

[ALxx] Alfred V. Aho and David Lee. Efficient algorithms for constructing testing sets, covering paths, and minimum flows. Computing Science Technical Report 159, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 15 pp.

Albrecht:1976:BUC

[Alb76] E. H. Albrecht. DOS-BATCH to UNIX conversion. Technical Memorandum 1120, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1, 1976. ?? pp.

Albrecht:1977:QCS

[Alb77a] E. H. Albrecht. QUEST — a computer system to collect questionnaire data. Technical Memorandum 1172, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 16, 1977. ?? pp.

Albrecht:1977:QAH

[Alb77b] E. H. Albrecht. QUEST administration handbook. Technical Memorandum 1178 (5161-770824.02 MF), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 24, 1977. ?? pp.

Alborough:1984:RUD

[Alb84] Thomas Alborough. Reloadable UNIX device drivers. In USENIX Association [USE84b], pages 135–144. ISBN none. LCCN QA76.8.U65 U55 1984.

Alexander:1996:SSU

[Ale96a] Michael Alexander. Simulate security on UNIX networks — SATAN has popularized the notion of using attack simulators to automatically probe TCP/IP networks in search of. Datamation, 42(5):69–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Alexander:1996:TCU

[Ale96b] Michael Alexander. Tools to cross the UNIX/NT divide — the smart money says UNIX and NT are going to be the top contenders in the enterprise operating-system arena. Which should you choose? Why not both, at least for now? Datamation, 42(3):54–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Alexander:2000:LIC

[Ale00] William F. Alexander. A Linux IEEE 1394 configuration ROM decoder. Dr. Dobb’s Journal of Software Tools, 25(8): 80, 82, 84, August 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_08/
linx1394.txt;   http://www.
ddj.com/ftp/2000/2000_08/
linx1394.zip.

Alexandrescu:2010:DPL

[Ale10] Andrei Alexandrescu. The D programming language. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-321-65953-8 (hardcover), 0-321-63536-1 (paperback). xxvii + 463 pp. LCCN QA76.73.D138 A44 2010; QA76.73.D138.

Alfieri:1994:EKI

[Alf94] R. A. Alfieri. An efficient kernel-based implementation of POSIX threads. In Anonymous [Ano94m], pages 59–72. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

Alles:1978:ASP

[All78] H. G. Alles. An array and switching processor. Technical Memorandum 1254 (TM 78-1356-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 24, 1978. ?? pp.

Allen:1983:RRT

[All83a]  Bill Allen.  REGULUS, a

real-time UNIX lookalike. In USENIX [USE83], pages 268– ?? Abstract only.

Allen:1983:RRU

[All83b] Bill Allen. REGULUS, a realtime UNIX lookalike. In Association [Ass83], pages 268–?? Abstract only.

Allman:1986:MRM

[All86a] Eric P. Allman. -me Reference Manual: Release 2.27. Report USD:20, Project IN-GRES, Electronics Research Laboratory, University of California, Berkeley, Berkeley, California 94720, USA, 1986.

                                        15 pp. URL http://docs.
                                        freebsd.org/44doc/usd/20.
                                        meref/paper.pdf.

Allman:1986:WPN

[All86b] Eric P. Allman. Writing papers with NROFF using -me. Report USD:19, Project INGRES, Electronics Research Laboratory, University of California, Berkeley, Berkeley, California 94720, USA, 1986. URL http:

//docs.freebsd.org/44doc/
usd/19.memacros/paper.pdf.

Allman:1987:UDF

[All87]  Eric Allman.  UNIX: The

data forms. In USENIX Association [USE87e], pages 9– 15. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Allman:1997:FTB

[All97] Mark Allman. Fixing two BSD TCP bugs. Washington, DC, USA, 1997. ???? pp. Prepared under contract NAS3-27121. Shipping list no. 99-0025-M.

Akkan:2013:UIN

[ALL13] Hakan Akkan, Michael Lang, and Lorie Liebrock. Understanding and isolating the noise in the Linux kernel. The International Journal of High Performance Computing Applications, 27(2):136–146, May 2013. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

sagepub.com/content/27/2/
136.full.pdf+html.

Abrams:1990:BGA

[ALO90] Marshall D. Abrams, Leonard J. LaPadula, and Ingrid M. Olson. Building generalized access control on UNIX. In USENIX Association [USE90e], pages 65–70. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Alomari:1997:OUP

[Alo97] Ahmed Alomari. Oracle and UNIX performance tuning. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-849167-4. xxiii + 255 pp. LCCN QA76.9.D3 A52 1997.

Alomari:1998:OUP

[Alo98]  Ahmed Alomari.  Oracle8

and UNIX performance tuning. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-907676-X. xxiii + 315 pp. LCCN QA76.9.D3 A519

1998.  US$39.95.  URL

                                        http://www.prenhall.com/
                                        allbooks/ptr_013907676x.
                                        html.

Alomari:2001:OUP

[Alo01]  Ahmed Alomari.  Oracle8i

and UNIX Performance Tuning. Prentice Hall PTR Oracle series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-018706-2. xxix + 415 pp. LCCN QA76.9.D3 A519 2001. US$49.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130187062.html.

Abdelsalam:2021:TWL

[ALP^{+}21] A. Abdelsalam, M. Luglio, N. Patriciello, C. Roseti, and F. Zampognaro. TCP Wave over Linux: a disruptive alternative to the traditional TCP window approach. Computer Networks (Amsterdam, Netherlands: 1999), 184(??):Article 107633, January 15, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1389128620312585.

Aho:2007:CPT

[ALSU07] Alfred V. Aho, Monica S. Lam, Ravi Sethi, and Jeffrey D. Ullman, editors. Compilers: Principles, Techniques, and Tools. Pearson/Addison Wesley, Boston, MA, USA, second edition, 2007. ISBN 0-321-48681-1 (hardcover), 0-321-49169-6 (paperback). xxiv + 1009 pp. LCCN QA76.76.C65 A37 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0618/2006024333.html.

Anyanwu:1986:CRU

[AM86] J. A. Anyanwu and L. F. Marshall. A crash resistant UNIX file system. Software—Practice and Experience, 16(2):107– 118, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Amaral:1988:OOV

[Ama88] Paolo Amaral. OFS — an optical view of a UNIX file system. In USENIX Association [USE88c], pages 203–211. ISBN ???? LCCN ????

Amagai:1991:DLP

[Ama91] Yoshiji Amagai. Distributed Lisp programming workbench. NTT R&D, 40(5):679–686, 1991. CODEN NTTDEC. ISSN 0915-2326.

ANSI:1989:C

[Ame89] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. American National Standard Programming Language C, ANSI X3.159-1989, December 14 1989.

ATT:1990:USVb

[Ame90a] American Telephone and Telegraph Company. UNIX System V release 4. BSD/XENIX compatibility guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-933664-8. various pp. LCCN QA76.76.O63 U552184 1990.

ATT:1990:USVa

[Ame90b] American Telephone and Telegraph Company. UNIX System V release 4: programmer’s guide: POSIX conformance. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-933672-9. i + 44 + 3 pp. LCCN QA76.76.O63 U5521869 1990 Bar.

ATT:1991:SVA

[Ame91] American Telephone and Telegragh Company. System V application binary interface: MIPS processor supplement: UNIX System V. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-880170-3. various pp. LCCN QA76.76.O63 S9742 1991 Bar.

ANSI:1993:ISI

[Ame93] American National Standards Institute. IEEE standard for information technology: Portable Operating System Interface (POSIX): part 2, shell and utilities. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 17, 1993. ISBN 1-55937-255-9. xvii + 1195 pp. LCCN QA76.76.O63I58 1993. Two volumes. IEEE Std 1003.2-1992 (includes IEEE Std 1003.2a-1992). Approved September 17, 1992, IEEE Standards Board. Approved April 5, 1993, American National Standards Institute. The primary purpose of this standard is to define a standard interface and environment for application programs that require the services of a ‘shell’ command language interpreter and a set of common utility programs.

ANSI:1994:ISI

[Ame94] American National Standards Institute. IEEE standard for information technology:

Portable Operating Sytem Interface (POSIX). Part 1, system application program interface (API) — amendment 1

— realtime extension [C language]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 1-55937-375-X. xxiii + 590 pp. LCCN TK 153 I59i no.1003.1b 1994. IEEE Std 1003.1b-1993 (formerly known as IEEE P1003.4; includes IEEE Std 1003.1-1990). Approved September 15, 1993, IEEE Standards Board. Approved April 14, 1994, American National Standards Institute.

Amee:1995:TTC

[Ame95] ?. Amee. Tcl/Tk CD-ROM. American Software and Hardware, ????, 1995. ISBN ???? LCCN ????

Amend:2009:MSU

[Ame09] Bill Amend. Math, science, and Unix underpants: a themed Foxtrot collection. Andrews McMeel Pub., LLC, Kansas City, MO, USA, 2009. ISBN 0-7407-9140-0. ???? pp. LCCN

                                        ???? URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1013/2009934089-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1013/2009934089-
                                        d.html.

Anderson:1980:TSL

[And80] Bruce Anderson. Type syntax in the language ’C’, an object lesson in syntactic innovation. ACM SIGPLAN Notices, 15(3):21–27, March 1980. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1988:DSG

[And88a] H. Stephen Anderson. Distributed supercomputer graphics using UNIX tools. In USENIX Association [USE88f], pages 25–32. ISBN ???? LCCN ????

Anderson:1988:UPS

[And88b] Lon E. Anderson. UNIX password security. In USENIX Association [USE88e], pages 5– 7. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Anderson:1991:MPB

[And91]  Paul Anderson.  Managing

program binaries in a heterogeneous UNIX network. In USENIX [USE91b], pages 1–10.

Anderson:1995:WGU

[And95] Bart Anderson. The Waite Group’s UNIX communications and the Internet. SAMS Publishing, Indianapolis, Ind., third edition, 1995. ISBN 0-672-30537-2. xxv + 817 pp. LCCN QA76.76.O63 A5 1995. Partial Contents: UNIX communications fundamentals — UNIX mail — UNIX news: USENET — UNIX file transfer: UUCP — Internet communications — Appendixes.

Anderson:2001:FOS

[And01] Annelise Anderson. FreeBSD: an open-source operating system for your personal computer. Bit Tree Press, Portola Valley, CA, USA, 2001. ISBN 0-9712045-1-9. xx + 423 pp. LCCN QA76.76.O63

A49 2001. URL http://www.
bittreepress.com/FreeBSD/

introbook/. Includes CD-ROM.

Andersen:2002:JEU

[And02]  Paul K. Andersen.  Just

enough UNIX. McGraw-Hill, New York, NY, USA, fourth edition, 2002. ISBN 0-07-246377-5 (paperback), 0-07-115130-3. x + 466 pp. LCCN QA76.76.O63 A48 2003.

URL http://www.loc.gov/
catdir/description/mh024/
2002071419.html;   http:/
/www.loc.gov/catdir/toc/
mh023/2002071419.html.

Anderson:2006:SXM

[And06] Fritz Anderson. Step into Xcode: Mac OS X development. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-33422-1 (paperback). xxii + 463 pp. LCCN QA76.76.O63 A53 2006. URL http:

//www.loc.gov/catdir/toc/
ecip062/2005029875.html.

Anderson:2009:XU

[And09]  Fritz Anderson.  Xcode 3

unleashed. SAMS Publishing, Indianapolis, IN, USA, 2009. ISBN 0-321-55263-6 (paperback). xix + 534 pp. LCCN QA76.76.O63 A53155 2009.

Anderson:2017:CUS

[And17] Jonathan Anderson. A comparison of Unix sandboxing techniques: Why sandboxing is different from historic approaches to Unix security, how we got where we are, and how Capsicum compares with Linux’s seccomp(2) and OpenBSD’s pledge(2). The FreeBSD Journal, 4(5):16–25, September/October 2017. ISBN 0-615-88479-2. URL http:/

/issue.freebsdfoundation.
org/publication/?m=33057&
i=447089&p=18.

Angelo:1977:MAL

[Ang77] J. Angelo. MAC-8 assembly language programming manual. Technical Memorandum 1221, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 8, 1977. ?? pp.

Angel:1997:ICG

[Ang97] Edward Angel. Interactive computer graphics: a top-down approach with OpenGL. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-85571-2. xxxvi + 521 pp. LCCN T385 .A514 1997.

Angel:1999:ICG

[Ang99] Edward Angel. Interactive computer graphics: a top-down approach with OpenGL. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-38597-X. xlii + 612 pp. LCCN T385 .A514 1999.

Annis:2000:PSU

[Ann00]  William S. Annis.  Pithy

sayings for the UNIX sysadmin. ;login: the USENIX Association newsletter, 25(5): ??, August 2000. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

                                        tronic). URL http://www.
                                        usenix.org/publications/
                                        login/2000-8/features/pithy.
                                        html.

Anonymous:1977:SUI

[Ano77] Anonymous. A standard for UNIX interprocess communication. Report, Bolt Beranek and Newman Inc., Cambridge, MA, USA, August 11, 1977. URL

                                        https://www.tuhs.org/cgi-
                                        bin/utree.pl?file=BBN-V6/
                                        doc/ipc/ipc.

Anonymous:1980:UTU

[Ano80] Anonymous. UNIX TCP user’s guide. Report 3724, Bolt Beranek and Newman Inc., Cambridge, MA, USA, January 9, 1980. 40 pp. URL

                                        https://www.tuhs.org/cgi-
                                        bin/utree.pl?file=BBN-V6/
                                        doc/tcp.

Anonymous:1982:USM

[Ano82] Anonymous. The UNIX system: Making computers easier to use. AT&T 23-minute video., 1982. URL http://

                                        techchannel.att.com/play-

video.cfm/2014/1/27/ATT-
Archives-The-UNIX-System-
Making-Computers-Easier-

to-Use. Interviews with Ken Thompson, Dennis Ritchie, Brian Kernighan, John Mashey, Lorinda Cherry, Nina Macdonald, Al Aho, and others.

Anonymous:1984:UBI

[Ano84] Anonymous. Unix^{T M} on big iron. UNIX Review, 2(7): ??, October 1984. CODEN UNRED5. ISSN 0742-3136.

URL https://ia902801.us.
archive.org/3/items/Unix_
Review_1984_Oct.pdf/Unix_
Review_1984_Oct.pdf.

Anonymous:1985:URG

[Ano85] Anonymous. The UNIX reference guide for System V: with expanded AWK, DC, ED, SHELL, SED, VI. System Publications, Trenton, NJ, USA, 1985. ISBN 0-935739-00-9. 45 pp. LCCN QA76.76.O63.

Anonymous:1986:IIC

[Ano86] Anonymous, editor. IMS ’86:

2nd International conference — October 1986, Boston, MA, Official Proceedings of the International IMS Conference 1986; 2nd. Intertec Communications, ????, 1986.

Anonymous:1988:OS

[Ano88a] Anonymous. Operating systems. Datamation, 34(20):19– ??, October 15, 1988. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1988:UPW

[Ano88b] Anonymous, editor. USENIX Proceedings. Workshop on UNIX and Supercomputers, Pittsburgh, PA, USA, September 26–27, 1988. USENIX Association, Berkeley, CA, USA, 1988.

Anonymous:1989:PAE

[Ano89]  Anonymous, editor.  Pro-

ceedings of the Autumn 1989 EUUG Conference. Eur. UNIX Syst. User Group, Buntingford, UK, 1989. ISBN 0-9513181-3-6. LCCN ????

Anonymous:1990:DMR

[Ano90a] Anonymous. DEC moves to RISC, POSIX. Information Week, 293:14–??, October 29, 1990. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1990:HPS

[Ano90b] Anonymous. The history of Posix: a study in the standards process. Computer, 23(7):89– ??, July 1990. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1991:DNO

[Ano91a] Anonymous. Digital nudges open the doors to VMS. Datamation, 37(2):83–??, January 1991. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1991:HAM

[Ano91b] Anonymous. HP airs MPE with POSIX compatibility. Network World, 8(48):27–29, December 1991. ISSN 0887-7661 (print), 1944-7655 (electronic).

Anonymous:1991:UIC

[Ano91c]  Anonymous.  UNIX inches

closer to off-the-shelf software. UNIX/World, 8(9):33– ??, September 1991. ISSN 0739-5922.

Anonymous:1992:RU

[Ano92a]  Anonymous.  A RAID on

UNIX. Datamation, 38(19): 106–??, September 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:SCU

[Ano92b] Anonymous. A Stronger CASE For UNIX. Datamation, 38(17): 71–??, August 15, 1992. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:AUa

[Ano92c] Anonymous. Answers to UNIX.

UNIX/World, 9(10):117–??, October 1992. ISSN 0739-5922.

Anonymous:1992:AUb

[Ano92d] Anonymous. Answers to UNIX.

UNIX/World, 9(11):127–??, November 1992. ISSN 0739-5922.

Anonymous:1992:AUc

[Ano92e] Anonymous. Answers to UNIX.

UNIX/World, 9(8):109–??, August 1992. ISSN 0739-5922.

Anonymous:1992:AUd

[Ano92f]  Anonymous.  Answers to

UNIX. UNIX/World, 9(9):

121–??, September 1992. ISSN 0739-5922.

Anonymous:1992:DYDa

[Ano92g] Anonymous. Downsize Your Database With UNIX. Datamation, 38(21):65–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:G

[Ano92h] Anonymous. GETLAB. Computing Science Technical Report 163i, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 17, 1992. 50 pp.

Anonymous:1992:OIG

[Ano92i] Anonymous. Open Interface get vendor boost. Information Week, 358:40–??, February 3, 1992. CODEN INFWE4. ISSN 8750-6874.

Anonymous:1992:PAP

[Ano92j] Anonymous. POSIX access platforms. Security, 29(7):18– ??, July 1992. CODEN SE-CUEU. ISSN 0890-8826.

Anonymous:1992:PLS

[Ano92k]  Anonymous.  POSIX loop-

holes = soft realtime. Computer Technology Review, XII (8):4–??, July 1992. CODEN CTERES. ISSN 0278-9647.

Anonymous:1992:RP

[Ano92l] Anonymous. Real-time Posix. BYTE Magazine, 17(8):177– 186, August 1992. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1992:RWG

[Ano92m] Anonymous. Real-time will get real standards. Datamation, 38 (21):97–??, October 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1992:UEa

[Ano92n] Anonymous. UNIX in Europe.

UNIX/World, 9(10):31–??, October 1992. ISSN 0739-5922.

Anonymous:1992:UEb

[Ano92o] Anonymous. UNIX in Europe. UNIX/World, 9(11):35– ??, November 1992. ISSN 0739-5922.

Anonymous:1992:UEc

[Ano92p] Anonymous. UNIX in Europe.

UNIX/World, 9(8):33–??, August 1992. ISSN 0739-5922.

Anonymous:1992:UEd

[Ano92q] Anonymous. UNIX in Europe. UNIX/World, 9(9):31–??, September 1992. ISSN 0739-5922.

Anonymous:1992:UEe

[Ano92r] Anonymous. UNIX in Europe. UNIX/World, IX(12):31– ??, December 1992. ISSN 0739-5922.

Anonymous:1992:UUM

[Ano92s] Anonymous. UnixWare: UNIX For The Masses. Datamation, 38(??):10–??, November 15, 1992. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:AUa

[Ano93a]  Anonymous.  Answers to

UNIX. UNIX/World, 10(1):91– ??, January 1993. ISSN 0739-5922.

Anonymous:1993:AUb

[Ano93b]  Anonymous.  Answers to

UNIX. UNIX/World, 10(2):79– ??, February 1993. ISSN 0739-5922.

Anonymous:1993:AUc

[Ano93c]  Anonymous.  Answers to

UNIX. UNIX/World, 10(3):87– ??, March 1993. ISSN 0739-5922.

Anonymous:1993:CKR

[Ano93d] Anonymous. Can Kerberos Really Make UNIX Secure? Datamation, 39(1):59–??, January 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CRE

[Ano93e] Anonymous. Chase Research: Excellence in UNIX I/O products. Datamation, 39(??):S–6– S–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CCU

[Ano93f] Anonymous. CLEO Communications: UNIX-to-IBM Connectivity. Datamation, 39 (??):S–22–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:CSA

[Ano93g] Anonymous, editor. Computer security, audit and control:

Proceedings of COMPSEC International October 1993. Oxford, UK. Elsevier Advanced Technology, Oxford, UK, 1993. ISBN 1-85617-211-2. LCCN QA76.9.A25 W68 1993.

Anonymous:1993:DUA

[Ano93h] Anonymous. Developing UNIX Applications. Datamation, 39 (??):S–23–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FLF

[Ano93i]  Anonymous.  Finding Less

Fault With UNIX. Datamation, 39(20):77–??, October 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:FSU

[Ano93j]  Anonymous.  Full-Strength

UNIX Finally Comes To PCs. Datamation, 39(14):47–??, July 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:GST

[Ano93k] Anonymous. GNU set to take on UNIX. Software magazine, 13(6):33–??, April 1993. CO-DEN SMWMEQ. ISSN 0897-8085.

Anonymous:1993:HUC

[Ano93l] Anonymous. How UNIX Can Manage Systems and Networks. Datamation, 39(??):S–16–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:HEB

[Ano93m] Anonymous. HP’s early bid to open its venerable MPE operating system with Posix hooks is off to a rough start. ComputerWorld, 27(15):10–??, April 1993. CODEN CMP-WAB. ISSN 0010-4841.

Anonymous:1993:HUG

[Ano93n] Anonymous. HP/UX gains super powers. Information Week, ??(423):14–??, May 1993. CO-DEN INFWE4. ISSN 8750-6874.

Anonymous:1993:SSU

[Ano93o] Anonymous. IS Seeks Security In UNIX. Datamation, 39(15): 59–??, August 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MUW

[Ano93p] Anonymous. Making UNIX Work. Datamation, 39(??):S– 20–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:MR

[Ano93q] Anonymous. Media reviews. Computer, 26(6):134–??, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:MBU

[Ano93r]  Anonymous.  Microbytes:

Bringing UNIX into the Net-Ware fold presents Novell with a marketing dilemma: what to tell software developers who ask which platform to write applications for. BYTE Magazine, 18(6):24–??, May 1993. CO-DEN BYTEDJ. ISSN 0360-5280.

Anonymous:1993:NB

[Ano93s]  Anonymous.  News briefs.

Computer Design, 32(4):10–??, April 1993. CODEN CMP-DAM. ISSN 0010-4566.

Anonymous:1993:PPC

[Ano93t]  Anonymous, editor.  Part-

ners in progress: Conference — April 1993, Phoenix, AZ, Spring Conference — Use Incorporated 1993. USE Incorporated, Lamham, 1993.

Anonymous:1993:PCL

[Ano93u] Anonymous. POSIX cracks the lock on MVS. Datamation, 39 (7):47–??, April 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PRM

[Ano93v] Anonymous. POSIX realtime may be long time coming. Computer Design, 32(7):38–??, July 1993. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1993:PSU

[Ano93w] Anonymous. Prices Slashed For UNIX Workbenches. Datamation, 39(21):51–??, November 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:PFA

[Ano93x] Anonymous, editor. Proceedings of the fifth annual Embedded Systems Conference: Santa Clara, California, October 5–8, 1993, volume 1 of Proceedings of the Annual Embedded Systems Conference 1993. Millar Freeman, San Francisco, CA, USA, 1993. ISBN 0-87930-306-9. LCCN TK7895.E42 E53 1993.

Anonymous:1993:RRU

[Ano93y] Anonymous. RAID Raises UNIX Reliability. Datamation, 39(17):61–??, September 01, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:RKP

[Ano93z] Anonymous. Real-time OS keeps up with POSIX standard. EDN, 38(8):110–??, April 1993. CODEN EDNSBH. ISSN 0012-7515, 0364-6637.

Anonymous:1993:RBT

[Ano93-27] Anonymous. RISCy Business: Today’s UNIX Players. Datamation, 39(??):S–8–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:S

[Ano93-28] Anonymous. Standards. Computer, 26(11):81–??, November 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1993:SA

[Ano93-29] Anonymous. State of the art. Embedded Systems Programming, 6(3):73–??, March 1993. CODEN EYPRE4. ISSN 1040-3272.

Anonymous:1993:SHP

[Ano93-30] Anonymous. Syncsort: High Performance UNIX Sorting. Datamation, 39(??):S–36–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:SAC

[Ano93-31] Anonymous, editor. Systems administration: 7th Conference — November 1993, Monterey, CA, Proceedings of the Systems Administration Conference 1993; 7th. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

Anonymous:1993:SGA

[Ano93-32] Anonymous. The Software Group: Adding Reliability to UNIX Internetworking. Datamation, 39(??):S–14–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:UUE

[Ano93-33] Anonymous. UniForum: Uni-Forum ’94: Exploring UNIX and Open Systems. pp S-40. Datamation, 39(??):??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:UCO

[Ano93-34] Anonymous. UNIX Connectivity: Opening Client/ Server Computing Opportunities. Datamation, 39(??):S–12– S–??, September 15, 1993. CO-DEN DTMNAT. ISSN 0011-

Anonymous:1993:UEa

[Ano93-35] Anonymous. UNIX in Europe. UNIX/World, 10(1):39– ??, January 1993. ISSN 0739-5922.

Anonymous:1993:UEb

[Ano93-36] Anonymous. UNIX in Europe. UNIX/World, 10(2):37– ??, February 1993. ISSN 0739-5922.

Anonymous:1993:UEc

[Ano93-37] Anonymous. UNIX in Europe. UNIX/World, 10(3):37– ??, March 1993. ISSN 0739-5922.

Anonymous:1993:UPD

[Ano93-38] Anonymous. UNIX Product Directory. Datamation, 39 (??):S–28–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:USP

[Ano93-39] Anonymous. UNIX Server Power Grows In Clusters. Datamation, 39(16):57–??, August 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:USM

[Ano93-40] Anonymous. UNIX Storage Management: It’s Here Now! Datamation, 39(21):77– ??, November 01, 1993. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:USS

[Ano93-41] Anonymous. UNIX Superserver Shoot-Out! Datamation, 39(22):24–??, November 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WUM

[Ano93-42] Anonymous. When UNIX Makes a Difference: IS Success Stories. Datamation, 39 (??):S–3–S–??, September 15, 1993. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1993:WNS

[Ano93-43] Anonymous. Windows NT supports Posix, but does it matter? BYTE Magazine, 18(12): 142–??, November 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Anonymous:1994:DIS

[Ano94a] Anonymous. Dr. Inder Singh on realtime benchmarking. Computer Design, 33(2):125–??, February 1, 1994. CODEN CM-PDAM. ISSN 0010-4566.

Anonymous:1994:GYC

[Ano94b] Anonymous. Get Your CICS On Anyone’s UNIX? Datamation, 40(4):45–??, February 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:GN

[Ano94c] Anonymous. Graphic news.

IEEE Computer Graphics and Applications, 14(1):79–??, January 1994. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Anonymous:1994:ISM

[Ano94d] Anonymous. Integrated Systems’ Moses Joseph on: Realtime POSIX: Boon or bunk? Computer Design, 33(10):155– ??, September 1, 1994. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1994:OPN

[Ano94e] Anonymous. OpenGL programs a new horizon for sun. SunWorld, ??(??):15–17, January 1994. CODEN SUN-WDW. ISSN 0149-1938.

Anonymous:1994:POS

[Ano94f] Anonymous. The POSIX open system environment. NIST special publication, ??(800):7–??, ???? 1994. CODEN NSPUE2. ISSN 1048-776X.

Anonymous:1994:PSI

[Ano94g]  Anonymous.  POSIX secu-

rity interfaces and mechanisms. NIST special publication, ?? (800):19–??, ???? 1994. CO-DEN NSPUE2. ISSN 1048-776X.

Anonymous:1994:PKP

[Ano94h] Anonymous. PSX kernel provides POSIX subset. Computer Design, 33(12):90–??, November 1, 1994. CODEN CMP-DAM. ISSN 0010-4566.

Anonymous:1994:RAR

[Ano94i] Anonymous. Real-time applications require open systems and stress use of Posix standard. Federal computer week, 8(19): 34–??, July 1994. ISSN 0893-052X.

Anonymous:1994:SS

[Ano94j] Anonymous. Software/systems.

Government computer news, 13 (16):56–??, July 1994. ISSN 0738-4300.

Anonymous:1994:SUM

[Ano94k] Anonymous. Standard UNIX Management: What’s The Holdup? Datamation, 40(4): 67–??, February 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:UTR

[Ano94l] Anonymous. UNIX Takes To The Road. Datamation, 40(7): 59–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1994:USC

[Ano94m] Anonymous, editor. USENIX Summer conference: – June 1994, Boston, MA, USENIX Conference Proceedings 1994. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

Anonymous:1994:WUH

[Ano94n] Anonymous. Where’s UNIX Headed? Datamation, 40(7): 24–??, April 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:AUA

[Ano95a] Anonymous, editor. Ada UK Ada 9X academic seminar: – January 1994, Brighton, volume 16 (3) of Ada User Journal 1995. IOS Press, Amsterdam, The Netherlands, 1995. ISSN 0268-652x.

Anonymous:1995:BRRb

[Ano95b] Anonymous. Book review: Running linux: By Matt Welsh and Lar Kaufman. O’Reilly & Associates, Inc., Sebastopol, CA. (1995). 575 pages. $24.95. Computers and Mathematics with Applications, 29(11):110, June 1995. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0898122195902708.

Anonymous:1995:IT

[Ano95c] Anonymous. Industry trends. Computer, 28(4):8–??, April 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:OSR

[Ano95d] Anonymous. Open systems — Richard H. Jaross and Michael Tilson argue that simply adding a Posix module doesn’t make Microsoft’s Windows NT an open system. ComputerWorld, 29(38):39–??, ???? 1995. CODEN CMPWAB. ISSN 0010-4841.

Anonymous:1995:PWWd

[Ano95e]  Anonymous.  Press Watch

— In a world of Windows clients, NetWare departmental servers, and the looming presence of NT, what’s the role of UNIX. Datamation, 41(21):94– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1995:PRC

[Ano95f] Anonymous. Product reviews — C++ class libraries, part 2; virtual Unix for PCs. Computer, 28(8):91–??, August 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1995:SSF

[Ano95g] Anonymous. SCADA software features POSIX certification, improved redundancy. Instrumentation & control systems: I&CS, 68(4):120–??, April 1995. CODEN ICSYFX. ISSN 1074-2328.

Anonymous:1995:SPA

[Ano95h]  Anonymous, editor.  Sig-

nal processing applications and technology: 6th International conference — October 1995, Boston, MA, volume 2 of International Conference on Signal Processing Applications and Technology 1995. DSP Associates, ????, 1995.

Anonymous:1995:UTT

[Ano95i] Anonymous. UNIX/POSIX tools target distributed embedded systems. Computer Design, 34(9):56–??, September 1, 1995. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1995:WSD

[Ano95j] Anonymous. WindRiver Systems’ David Fraser on: POSIX for real-time embedded development. Computer Design, 34 (6):130–??, June 1995. CODEN CMPDAM. ISSN 0010-4566.

Anonymous:1996:APG

[Ano96a]  Anonymous.  Appendix A:

a portion of a GDMO definition for a UNIX password file. Hewlett–Packard Journal, 47(5):??, October 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/hpcc997.external.hp.com:
80/hpj/oct96/tc-10-96.htm.

Anonymous:1996:BRPi

[Ano96b] Anonymous. Book review: Practical UNIX and Internet security, second edition: By Simson Garfinkel and Gene Spafford. O’Reilly & Associates, Sebastopol, CA. (1996). 971 pages. $39.95. Computers and Mathematics with Applications, 32(6):135, September 1996. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122196902552.

Anonymous:1996:FFF

[Ano96c] Anonymous. FreeBSD: fast 400 Freeware-Pakete. — Version 2.1.5. Franzis, Feldkirchen, Germany, 1996. ISBN 3-7723-8273-8. 9 pp. LCCN ???? Includes CD-ROM.

Anonymous:1996:NNM

[Ano96d] Anonymous. NEWSTRENDS — no more upgrades, please. Internet/VB Script: a tough sell. look out, PC; here comes the NC. reality check for VRML programs. communications: The focal point. Microsoft’s target: The Internet. from NT to UNIX?. notes from the lab. and more. Windows Magazine, 7(5):32–??, ???? 1996. CODEN WINMEV. ISSN 1060-1066.

Anonymous:1996:SNM

[Ano96e]  Anonymous.  Sun’s new

mainframe-killing 64Bit Ultra-SPARC servers — Sun’s new multi-CPU UltraSPARC Enterprise line of servers shows what UNIX hardware vendors have in store for the future: bigger, better, faster, more. Datamation, 42(11):36–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1996:WGV

[Ano96f]  Anonymous.  Webmaster’s

guide: version 4.1 for AIX, HP-UX, and Solaris: IBM Internet connection server: IBM Internet connection secure server. IBM Corporation, Research Triangle Park, NC, 1996. x + 303 pp. LCCN TK5105.888 .W382 1996. Document number GC31-8287-00.

Anonymous:1997:BNW

[Ano97a] Anonymous. A brave new world: UNIX developers in an NT land. ;login: the USENIX Association newsletter, 22(7): ??, December 1997. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.

                                        usenix.org/publications/
                                        login/1997-12/brave.html.

Anonymous:1997:WFL

[Ano97b] Anonymous. Wells Fargo’s online customer banking system is based on S/390 mainframes connected to UNIX servers. Datamation, 43(4):48–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Anonymous:1998:FFB

[Ano98a] Anonymous. FreeBSD 2.2.6: a full 4.4 BSD lite based 32bit operating system. Walnut Creek CDROM, Walnut Creek, CA, USA, 1998. ISBN 1-57176-079-2. LCCN QA76.76.O63. Includes four CD-ROMs.

Anonymous:1998:LOS

[Ano98b] Anonymous. The Linux operating system. IEEE Software, 15(5):17, September/ October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://dlib.computer.
                                        org/so/books/so1998/pdf/
                                        s5017.pdf.

Anonymous:1998:NPV

[Ano98c] Anonymous. New products: Visual Prolog 5.0, Prolog Development Center; Samba: Integrating UNIX and Windows, SSC; VPN Client, Aventail Corporation; eVote 2.2, www.deliberate.com; TowerJ 2.0, Tower Technology Corporation; NetVue/JAVA, Accu-Soft Corporation. Linux Journal, 49:??, May 1998. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:1998:NTS

[Ano98d] Anonymous. New tools: Software development: Uniscape’s internationalization library; Global Technologies’ Unix-to-NT solution; KAI’s multithreaded Java debugging tool; Price Systems’ parametric forecasting tool. Computer, 31(6): 98, 102, June 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1998/pdf/
r6098.pdf.

Anonymous:1998:UCB

[Ano98e] Anonymous. The UNIX CD bookshelf: 6 bestselling books on CD-ROM. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-001-5. ???? pp. LCCN QA76.76.O63 U5453 1992. Includes: UNIX in a nutshell: a desktop quick reference for System V Release 4 and Solaris 2.0, UNIX power tools, Sed & awk, Learning the Korn shell, Learning Vi, and Learning the UNIX operating system.

Anonymous:1999:FLE

[Ano99a] Anonymous. Focus: Linux — essay. IEEE Software, 16(1):??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLT

[Ano99b] Anonymous. Focus: Linux — from the trenches. IEEE Software, 16(1):??, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:FLN

[Ano99c] Anonymous. Focus: Linux — nuts & bolts. IEEE Software, 16(1):??, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Anonymous:1999:VCTb

[Ano99d]  Anonymous.  Vulnerability

in Compaq’s Tru64/DIGITAL UNIX software. Network Security, 1999(7):2, July 1999. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S1353485899900004.

Anonymous:19xx:G

[Anoxxa] Anonymous. Glossary. Computing Science Technical Report 163h, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 2 pp.

Anonymous:19xx:OOS

[Anoxxb]  Anonymous.  Opening Up

“Open Systems”: Moving toward True Interoperability among Library Software, volume 1 of DataResearch Automation Guide Series. DataResearch, ????, 19xx. ???? pp.

Anonymous:2000:BRUc

[Ano00a] Anonymous. Book review:

UNIX in a nutshell: a desktop quick reference for System V release 4 and Solaris 7, third edition: By Arnold Robbins. O’Reilly, Sebastopol, CA. (1999). 598 pages. $24.95. Computers and Mathematics with Applications, 39(3–4):265, February 2000. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0898122100900750.

Anonymous:2000:BRUb

[Ano00b] Anonymous. Book reviews:

Unix Backup and Recovery by Charles Curley; L^{A}TEX for Linux by Ben Crowder; The XML Handbook 2nd Edition, by Daniel Lazenby; Securing Linux, by Charles Curley; Building Linux and OpenBSD Firewalls, by Ralph Krause; Linux Programming Bible by Ben Crowder. Linux Journal, 78:??, October 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL http:

//noframes.linuxjournal.
com/lj-issues/issue78/3839.
html;   http://noframes.
linuxjournal.com/lj-issues/
issue78/3872.html.

Anonymous:2000:CDG

[Ano00c] Anonymous. Citius Debian GNU/Linux 2.2. Investigación y Desarrolo Agora, Navarra, Spain, 2000. ISBN 84-607-1193-5. LCCN ???? One volume and ten CD-ROMs.

Anonymous:2000:FPF

[Ano00d] Anonymous. Forefront: Product features: Top UNIX server employs 64-way processor and prepares for the IA-64 Itanium CPU. Electronic Design, 48 (22):62–63, 2000. CODEN ELODAW. ISSN 0013-4872.

Anonymous:2000:NPAa

[Ano00e]  Anonymous.  New prod-

ucts: AVP for Linux/FreeBSD UNIX, Kaspersky Lab Ltd.; API PowerRAC Chassis 320, Alpha Processor Inc.; ODBC-ODBC Bridge, Easysoft Ltd.; LinkScan 6.1, Electronic Software Publishing Corporation; Metro-X Enhanced Server CD, Metro Link, Inc.; P-STAT Statistical Software, P-STAT, Inc.; System Manager in a Box v1.0, PegaSoft Canada; PGI Workstation 3.1, PGI; Quick Restore 2.6, Workstation Solutions, Inc.; Threads.h++ and Tools.h++ Professional, Rogue Wave Software; Scriptics Connect 1.0, 1.1, Scriptics Corporation; TapeWare 6.2 Backup Software, Yosemite Technologies, Inc.; DoubleVision for Linux Systems, Tridia Corporation. Linux Journal, 71:??, March 2000. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Anonymous:2000:TUS

[Ano00f] Anonymous. Top UNIX server employs 64-way processor and prepares for the IA-64 Itanium CPU. Electronic Design, 48 (22):62–63, ???? 2000. CODEN ELODAW. ISSN 0013-4872.

Anonymous:2000:UCB

[Ano00g] Anonymous. The UNIX CD bookshelf [computer file]: 6 bestselling books on CD-ROM: UNIX power tools; UNIX in a nutshell; Learning the UNIX operating system; sed and awk; Learning the vi editor; Learning the Korn shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, version 2.1. edition, February 2000. ISBN 0-596-00000-6. 624 pp. LCCN QA76.76.O63. US$69.95. URL http:/

/www.oreilly.com/catalog/
9780596000004;   http://
www.oreilly.com/catalog/

unixcdbs2. Includes CD-ROM.

Anonymous:2001:PPT

[Ano01] Anonymous. Products: Planet 7 Technologies’ new XML development software; Apple Computer launches UNIXbased operating system; Codemesh releases Java/C++ integration software for Solaris; Telelogic AB’s C++ quality assessment environment; Green Hills ships embedded integrated development platform; Concurrent Computer introduces Linux application development environment. Computer, 34(5): 85–86, May 2001. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co2001/pdf/
r5085.pdf.

Anonymous:2002:OAI

[Ano02a] Anonymous. Optimizing applications with the Intel C++ and Fortran compilers for Windows and Linux. Technical report, Hewlett–Packard Corporation, Palo Alto, CA, USA, 2002. URL http:

//www.intel.com/software/
products/compilers/c60/techtopics/
Compiler_Optimization_6.
pdf.

Anonymous:2002:PNR

[Ano02b] Anonymous. Products: New rack-mount server from Apple; OSDL’s Linux database test tool; DigiTerra ships backoffice integration toolset; Oracle’s Developer Suite for business applications and Web services; Macromedia release J2EE-compatible application server and IDE; Motion-Builder 4.0 from Kaydara; Zero G Software’s InstallAnywhere 5. Computer, 35(7):96– 97, July 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/07/r7096.
htm;   http://csdl.computer.
org/dl/mags/co/2002/07/r7096.
pdf.

Anonymous:2002:POU

[Ano02c] Anonymous. Products: Omnicore upgrades Java IDE CodeGuide emWare’s SDE for intelligent device management; Metrowerks’ CodeWarrior for Embedded Linux; integrated software environment form Xilinx; new version of InstallShield Professional; Motorola’s 32-bit CAN reference design; Utopia-LVDS bridge reference design kit from National Semiconductor; First Silicon Solutions’ analysis tool for flash-based FPGAs. Computer, 35(11):78–79, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/11/ry078.
htm;   http://csdl.computer.
org/dl/mags/co/2002/11/ry078.
pdf.

Anonymous:2003:LUE

[Ano03a] Anonymous. Linux/Unix extend Red Hat 8’s functionality by adding Flash, extra fonts and Java. Personal computer world, 26(3):202–207, 2003. CODEN PCWODU. ISSN 0142-0232.

Anonymous:2003:NBI

[Ano03b] Anonymous. Net.BSD 1.6: installieren, konfigurieren, administrieren. Computer-und-Literatur-Verlag, Böblingen, Germany, 2003. ISBN 3-936546-00-2. 848 pp. LCCN ???? EUR 49.90. Includes CD-ROM.

Anonymous:2003:PAI

[Ano03c] Anonymous. Products: Apple introduces 64-bit PC with 1-GHz bus; Altera upgrades PLD design tool; StrikeIron’s initial product analyzes Web services; Metrowerks teams up with AMD on Linux platform. Computer, 36(8):88, August 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/08/r8088.
htm;   http://csdl.computer.
org/dl/mags/co/2003/08/r8088.
pdf.

Anonymous:2003:PCN

[Ano03d] Anonymous. Products: Clear-Sight Networks releases applicationlayer analyzer; Intervoice announces first SALT-based components; VoiceGenie Technologies upgrades VoiceXML platform; AppForce enhances mobile-platform design software; Metrowerks upgrades tools for embedded Linux products; OpenOffice.org updates Linux office tool suite; Quest Software releases Java analysis tools. Computer, 36(11):86–87, November 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/11/ry086.
htm;   http://csdl.computer.
org/dl/mags/co/2003/11/ry086.
pdf.

Anonymous:2003:PIU

[Ano03e] Anonymous. Products: Intel updates multimedia performance primitives library; Lindows.com launches lowcost Linux PC; Metrowerks adds Linux kernel-level debugging; Rogue Wave announces C/C++ tool for Web services; ScanSoft announces toolkits to support Windows Mobile 2003; CoCreate launches collaboration toolset; Sybase ships IDE for Windows CE-based devices. Computer, 36(10):100–101, October 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/10/rx100.
htm;   http://csdl.computer.
org/dl/mags/co/2003/10/rx100.
pdf.

Anonymous:2003:PSA

[Ano03f] Anonymous. Products: SGI announces record-performing Linux servers; Pacific Nanotechnogy releases atomic force microscope; Nvidia extends Cg programming environment; Empirix introduces VoIP analyzer; Codagen Technologies upgrades UML tool; Perforce Software enhances SCM system; Compuware upgrades Java development tool. Computer, 36(2):80–81, February 2003. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2003/02/r2080.

                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/02/r2080.
                                        pdf.

Anonymous:2003:PSU

[Ano03g] Anonymous. Products: Sybase upgrades RAD tool; haptic workstation uses two hands; TimeSys launches Linux 4.0. Computer, 36(4):89, April 2003. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

                                        http://csdl.computer.org/
                                        dl/mags/co/2003/04/r4089.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/co/2003/04/r4089.
                                        pdf.

Anonymous:2005:DGL

[Ano05] Anonymous. Debian GNU/ Linux 3.1 r0a “Sarge”. Lehmann, Berlin, Germany, 2005. ISBN 3-86541-064-2. LCCN ???? EUR9.95. Two DVD-ROMs.

Anonymous:2006:PGI

[Ano06a] Anonymous, editor. Proceedings of Gelato ICE: Itanium Conference and Expo: Spotlighting Linux on Itanium-based Platforms, October 1–4, 2006, Biopolis, Singapore. ????, ????, 2006. ISBN ???? LCCN

                                        ????  URL http://www.
                                        ice.gelato.org/;   http:/
                                        /www.ice.gelato.org/about/
                                        oct06_presentations.php.

Anonymous:2006:SSD

[Ano06b] Anonymous. Solaris Security for Developers Guide. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 2006. ISBN 0-595-28558-9. ???? pp. LCCN

???? URL http://docs.sun.
com/app/docs/doc/816-4863.

Anonymous:2011:JPD

[Ano11] Anonymous. 2011 Japan Prize to Dennis M. Ritchie and Ken L. Thompson. Web site, 2011. URL http://www.

japanprize.jp/en/prize_past_
2011_prize01.html.

Anonymous:2016:BRA

[Ano16a] Anonymous. Book review:

Automated Security Analysis of Android and iOS Applications with Mobile Security Framework, by Henry Dalziel and Ajin Abraham. Syngress. ISBN 978-0-12-805105-4. Network Security, 2016(5):4, May 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1353485816300447.

Anonymous:2016:BRI

[Ano16b] Anonymous. Book review: iOS Application Security, by David Thiel. No Starch Press. ISBN 978-1-59327-601-0. Network Security, 2016(2):4, February 2016. CODEN NTSCF5. ISSN 1353-4858 (print), 1872-9371 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1353485816300150.

Anonymous:2019:CYU

[Ano19a] Anonymous. Celebrating 50 years of UNIX: 1969–2019. Web site, 2019. URL https:// unix50.org/. Offers shell access for historical versions of UNIX.

Anonymous:2019:CCP

[Ano19b] Anonymous. Comparison of C/POSIX standard library implementations for Linux. Web site., 2019. URL http:/

                                        /www.etalabs.net/compare_
                                        libcs.html.

Anonymous:2019:NIH

[Ano19c] Anonymous. National Inventors Hall of Fame honors creators of Unix, power drills and more: Say hello to the class of 2019. Web site., January

                                        8, 2019. URL https://www.
                                        engadget.com/2019/01/08/
                                        national-inventors-hall-
                                        of-fame-class-of-2019/.

Anonymous:2019:UOU

[Ano19d] Anonymous. Unix50 — the origin of Unix — panel session. Conference video (38m52s)., October 22, 2019. URL https:

                                        //www.youtube.com/watch?
                                        v=dUt3L3fLFt4.

Anonymous:2019:UUT

[Ano19e] Anonymous. Unix50 — Unix today and tomorrow: Future of compute & platforms: The kernel. Conference video (58m13s)., October 22, 2019.

                                        URL https://www.youtube.
                                        com/watch?v=dUt3L3fLFt4.

Anonymous:2024:BSD

[Ano24] Anonymous. The berkeley software distribution. Web site, February 4, 2024. URL https:

//www.abortretry.fail/p/
the-berkley-software-distribution.

Anyanwu:1985:RSS

[Any85] J. A. Anyanwu. A reliable stable storage system for UNIX. Software—Practice and Experience, 15(10):973–990, October 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Auerbach:1986:USC

[AO86] Karl Auerbach and Robin O’Neill. A UNIX subsystem on the Cray Time Sharing System (CTSS). In USENIX Association [USE86e], pages 211–218. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Alrshah:2014:CSH

[AOAH14] Mohamed A. Alrshah, Mohamed Othman, Borhanuddin Ali, and Zurina Mohd Hanapi. Comparative study of high-speed Linux TCP variants over high-BDP networks. Journal of Network and Computer Applications, 43(??):66– 75, August 2014. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1084804514000903.

Anderson:1999:KG

[APK99] Todd Anderson, Laurie Petrycki, and Sarah Kearns. KDE guide. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 0-7357-0015-7 (??invalid ISBN??). ???? pp. LCCN ???? Edited by Laurie Petrycki and Sarah Kearns.

Apollo:1987:MBS

[Apo87] Apollo Computer Inc., Chelmsford, MA, USA. Managing BSD system software, 1987.

Appelbe:1984:AOC

[AQ84] Bill Appelbe and Bob Querido. An adaptable object code optimizer for UNIX systems. In Software Tools Users Group [Sof84], pages 111–118. LCCN QA76.8.U65 U83 1984.

Arnold:1978:TGS

[AR78] S. L. Arnold and B. E. Redman. A traffic generating system. Technical Memorandum 1333 (TM 78-4533-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 29, 1978. ?? pp.

Apte:2007:APL

[AR07] Himani Apte and Meenali Rungta. Adding parity to the Linux ext3 file system. Operating Systems Review, 41(1):56– 65, January 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Archer:1993:TPS

[Arc93] B. Archer. Towards a POSIX standard for software administration. In Anonymous [Ano93-31], pages 67–79. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

Arick:1992:UCS

[Ari92] Martin Arick. UNIX C Shell — Desk Reference. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1992. ISBN 0-89435-328-4. xiv + 204 pp. LCCN QA76.76.O63 A75 1992. US$34.95. A more recent text on C-shell.

Armstrong:1996:US

[Arm96] James C. Armstrong, Jr. UNIX Secrets. I D G Books Worldwide, Indianapolis, IN, USA, 1996. ISBN 1-56884-499-9. xxxviii + 936 pp. LCCN QA76.76.O63 A757 1996 Sci-Eng.

Arnold:1978:APG

[Arn78] J. Arnold. Application planning guide, release 0 of Basic-16. Technical Memorandum 5 (TM 78-2522-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 15, 1978. ?? pp.

Arnold:1986:SLU

[Arn86] James Q. Arnold. Shared libraries on UNIX system V. In USENIX Association [USE86c], pages 395–404.

Arnold:1991:IYS

[Arn91] Bob Arnold. If you’ve seen one UNIX, you’ve seen them all. In USENIX [USE91b], pages 11– 20.

Arnold:1993:USP

[Arn93] Derek Arnold. UNIX Security — a Practical Tutorial. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-002560-

6. xix + 386 pp. LCCN QA76.76.O63 A758. Covering the tools and techniques for managing and controlling security.

Arnold:1998:AUC

[Arn98] Bob Arnold. Accountworks: Users create accounts on SQL, Notes, NT, and UNIX. In USENIX [USE98d], page ?? ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL

http://www.usenix.org/publications/
library/proceedings/lisa98/
arnold.html.

Arthur:1986:USP

[Art86] Lowell Jay Arthur. UNIX Shell programming. Wiley-Intersci-ence, New York, NY, USA, 1986. ISBN 0-471-84932-4, 0-471-83900-0 (paperback). xv + 261 pp. LCCN QA76.76.O63 A765 1986.

Arthur:1990:USP

[Art90] Lowell Arthur. UNIX Shell Programming. Wiley-Intersci-ence, New York, NY, USA, second edition, 1990. ISBN 0-471-51821-2. xii + 272 pp. LCCN QA76.76.O63 A765 1990. This covers not only common shells but also general software tool concepts.

Arthur:1994:USP

[Art94] Lowell Arthur. UNIX Shell Programming. Wiley-Intersci-ence, New York, NY, USA, third edition, 1994. ISBN 0-471-59941-7. xiv + 462 pp. LCCN QA76.76.O63 A765 1994. US$29.95. This covers not only common shells but also general software tool concepts.

Artymiak:1999:STY

[Art99] Jacek Artymiak. Sams teach yourself Sed and Awk in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31737-0. 400 (est.) pp. LCCN ????

Artymiak:2003:BFO

[Art03] Jacek Artymiak. Building firewalls with OpenBSD and PF. devGuide.net, Lublin, Poland, second edition, 2003. ISBN 83-916651-1-9. 321 pp. LCCN ????

Aho:1988:MCR

[AS88] Alfred V. Aho and Ravi Sethi.

Maintaining cross references in manuscripts. Software—Practice and Experience, 18(1):1– 13, January 1988. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Asente:1990:XWS

[AS90] Paul J. Asente and Ralph R.

Swick. X Window System Toolkit: The Complete Programmer’s Guide and Specification. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 0-13-972191-6 (Prentice Hall), 0-13-973173-3, 0-13-975491-6 (Prentice Hall), 1-55558-051-3 (DP paperback). xxxv + 967 pp. LCCN QA76.76.W56 A74 1990.

Aho:1991:MCR

[AS91] Alfred V. Aho and Ravi Sethi. Maintaining cross references in manuscripts. Computing Science Technical Report 127, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 20, 1991. 10 pp.

URL https://web.archive.
org/web/*/http://cm.bell-
labs.com/cm/cs/cstr/127.
ps.gz.

Akyurek:1993:ABR

[AS93a] Sedat Akyurek and Kenneth Salem. Adaptive block rearrangement under UNIX. In USENIX [USE93c], pages 307– 321. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/cinci93/akyurek.
html.

Andreasson:1993:GSU

[AS93b] Torbjörn Andréasson and Jan Skansholm. Getting started with UNIX and X. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-63170-9. xi + 271 pp. LCCN QA76.76.O63 A5347 1993.

Akyurek:1995:ABR

[AS95] Sedat Akyürek and Kenneth Salem. Adaptive block rearrangement. ACM Transactions on Computer Systems, 13(2): 89–121, May 1995. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1995-13-2/p89-akyurek/

.

Akyurek:1997:ABR

[AS97] Sedat Akyürek and Kenneth Salem. Adaptive block rearrangement under UNIX. Software—Practice and Experience, 27(1):1–23, January 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
7254;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=7254&PLACEBO=IE.pdf.

Asbury:2000:ELW

[Asb00] Steve Asbury. Enterprise Linux at Work: How to Build 10 Distributed Applications for Your Organization. Wiley, New York, NY, USA, 2000. ISBN 0-471-36349-9. xxiii + 418 pp. LCCN QA76.76.O63 A78 2000. US$49.99; CDN$77.50. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/onix06/99059493.html.

Arthurs:1979:DAO

[ASS79] E. Arthurs, N. Sharko, and B. W. Stuck. Directory assistance operator work time exploratory data analysis. Technical Memorandum 1414, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 6, 1979. ?? pp.

USENIX:1983:UCPb

[Ass83] USENIX Association, editor. USENIX Conference Proceedings, Winter, 1983. San Diego, CA. USENIX Association, Berkeley, CA, USA, Winter 1983.

Aitken:1985:DID

[ASS85] Gary Aitken, Christine Scott, and Kenneth Scott. DIBOLIX — an implementation of DI-BOL under UNIX. In USENIX Association [USE85c], pages 30–33.

USENIX:1988:CSSa

[Ass88a] USENIX Association, editor. Computing Systems, Spring, 1988. USENIX Association, Berkeley, CA, USA, Spring 1988.

USENIX:1988:CSSb

[Ass88b] USENIX Association, editor. Computing Systems, Summer, 1988. USENIX Association, Berkeley, CA, USA, Summer 1988.

USENIX:1988:FRT

[Ass88c] USENIX Association, editor. Fifth Real-Time Software and Operating Systems Workshop Proceedings, May 12–13, 1988. Washington, DC. USENIX Association, Berkeley, CA, USA, May 12–13, 1988.

USENIX:1988:UCPa

[Ass88d] USENIX Association, editor.

USENIX Conference Proceedings, Summer, 1988. San Francisco. USENIX Association, Berkeley, CA, USA, Summer 1988.

USENIX:1989:CSSa

[Ass89a] USENIX Association, editor.

Computing Systems, Spring, 1989. USENIX Association, Berkeley, CA, USA, Spring 1989.

USENIX:1989:CSW

[Ass89b] USENIX Association, editor.

Computing Systems, Winter, 1989. USENIX Association, Berkeley, CA, USA, Winter 1989.

USENIX:1989:EUS

[Ass89c] USENIX Association, editor. European UNIX Systems User Group Newsletter, Spring, 1989. EUUG, Berkeley, CA, USA, Spring 1989. CODEN EONLE8. ISSN 1011-4211.

OReilly:1999:CWN

[Ass99a] O’Reilly and Inc. Associates.

Complete Windows NT & UNIX System Administration Pack. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ???? pp. LCCN ???? US$149.00. URL http:/

/www.oreilly.com/catalog/
hp12pack.

OReilly:1999:NCB

[Ass99b] O’Reilly and Inc. Associates. The Networking CD Bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-523-8. 456 pp. LCCN ???? US$79.95. URL http:/

/www.oreilly.com/catalog/

netcd/. Six O’Reilly networking books on CD-ROM. Issued in container and accompanied by a print ed. of DNS and BIND (3rd ed.). Contents: DNS and BIND (3rd ed.) – TCP/IP network administration (2nd ed.)

– Building Internet firewalls – Practical UNIX and Internet security (2nd ed.) – Sendmail desktop reference – Sendmail (2nd ed.).

Aho:1986:CPC

[ASU86] Alfred V. Aho, Ravi Sethi, and Jeffrey D. Ullman. Compilers: Principles, Techniques, and Tools. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10088-6 (hardcover), 0-201-10194-7 (paperback). x + 796 pp. LCCN QA76.76.C65 A371 1986. See [AU77].

Allison:2001:LLE

[ASW^{+}01] Dennis Allison, Randy Schrickel, Reid Womack, Jeremy C. Reed, Ashley Tate, and Paul Munsey. Letters: Looking for early PPC [People’s Computing Company] people; being prepared for invasion; BetterBASIC; Linux versus BSD; Diffie–Hellman to the rescue; the future of programming. Dr. Dobb’s Journal of Software Tools, 26(6):10, 12, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

ATT:1987:ADG

[AT7] AT&T Customer Information Center, 2855 N Franklin Road, Indianapolis, IN 46219, USA. AT&T Documentation Guide, November 1987. 95

pp. URL https://bitsavers.
org/pdf/att/000-111_ATT_
Documentation_Guide_Nov87.
pdf.

Adelstein:2007:LSA

[AT07] Tom Adelstein and Falko Timme. Linux system administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-00952-6 (paperback). xiii + 279 pp. LCCN QA76.76.O63; QA76.76.O63 A34 2007eb. URL http:/

/www.oreilly.com/catalog/
9780596009526.

Alexander:1994:EHS

[ATS94] D. Scott Alexander, C. Brendan S. Traw, and Jonathan M. Smith. Embedding high speed ATM in UNIX IP (abstract only). In USENIX [USE94b], pages 119–121. ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994. URL http://www.

                                        usenix.org/publications/
                                        library/proceedings/hispeed/
                                        index.html.

Aho:1974:DMR

[AU74] Alfred V. Aho and Jeffrey D. Ullman. Dynamic memories with rapid random and sequential access. IEEE Transactions on Computers, C-23(3): 272–276, March 1974. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1672515;   https:
                                        //ieeexplore.ieee.org/abstract/
                                        document/1672515.

Aho:1977:PCD

[AU77] Alfred V. Aho and Jeffrey D. Ullman. Principles of Compiler Design. Addison-Wesley, Reading, MA, USA, 1977. ISBN 0-201-00022-9 (hardcover), 0-201-10073-8. x + 604 pp. LCCN QA76.6 .A285 1977. See also the much expanded subsequent book [ASU86].

Aho:1973:DMR

[AU3 ] A. V. Aho and J. D. Ullman. Dynamic memories with rapid random and sequential access. Computing Science Technical Report 7, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973 (??). ?? pp.

Auditore:1996:PIN

[Aud96] Peter Auditore. PC internetworking: a new frontier

— Windows NT and UNIX/X Windows integration. The X Journal, 5(5):70–??, May 1996. CODEN XJOUEA. ISSN 1056-7003.

Ault:1996:USA

[Aul96] Michael R. Ault. UNIX System Administrator’s companion. Wiley, New York, NY, USA, 1996. ISBN 0-471-11144-9 (paperback). xii + 356 pp. LCCN QA76.76.O63 A89 1996.

Aulds:2001:LAW

[Aul01] Charles Aulds. Linux Apache Web server administration. The Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2734-7. xxiii + 615 pp. LCCN TK5105.8885.A63 A95 2001.

Avolio:1995:STP

[AV95] Frederick M. Avolio and Paul A. Vixie. Sendmail: theory and practice. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1995. ISBN 1-55558-127-7. xv + 262 pp. LCCN HE6239.E54 A96 1995.

Avant:1992:RFU

[Ava92] Harry Avant. Run DOS to Fix UNIX. UNIX/World, 9(10):95– ??, October 1992. ISSN 0739-5922.

Albing:2007:BC

[AVN07] Carl Albing, J. P. Vossen, and Cameron Newham. Bash cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-52678-4 (paperback). xxi + 598 pp. LCCN QA76.76.O63 A39885 2007.

Almada:1989:EBU

[AW89] Alfredo Almada and David H. Williams. Enhancing the 4.3 BSD UNIX serial line interface. ;login: the USENIX Association newsletter, 14(1):6–29, January/February 1989. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Beck:1998:LKI

[B^{+}98] Michael Beck et al. Linux kernel internals. Addison-Wesley, Reading, MA, USA, second edition, 1998. ISBN 0-201-33143-8. xvi + 480 pp. LCCN QA76.76.O63L54813 1998. Includes CD-ROM. Translation of the German edition Linux-Kernel-Programmierung.

Brenner:1996:ICP

[BA96] Steven Brenner and Edwin Aoki. Introduction to CGI/ Perl: getting started with Web scripts. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1996. ISBN 1-55851-478-3. xv + 151 pp. LCCN TK5105.888 .B75 1996.

Babu:1978:KPC

[Bab78] R. R. Babu. KW11-P clock driver. Technical Memorandum 1309 (TM 78-3112-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 22, 1978. ?? pp.

Bach:1984:MUO

[Bac84] M. J. Bach. Multiprocessor UNIX operating systems. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1733–1749, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Bach:1986:DUO

[Bac86a] Maurice J. Bach. The Design of the UNIX Operating System. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-201799-7. xiv + 471 pp. LCCN QA76.76.O63 B32 1986. See also [GC94a].

Bach:DUO86

[Bac86b] Maurice J. Bach. The Design of the UNIX Operating System. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1986. ISBN 0-13-201799-7. xiv + 471 pp. LCCN QA76.76.O63 B33 1986.

Bettison:1989:LSU

[BAC^{+}89] A. Bettison, F. Adcock, P. Chubb, A. Gollan, C. Maltby, and N. Russell. Limits — a system for UNIX resource administration. In ACM [ACM89], pages 686–692. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Backman:1992:UCB

[Bac92] Rex Backman. Uncovering CPU Bottlenecks with UNIX Tools. UNIX/World, 9(10): 101–??, October 1992. ISSN 0739-5922.

Badii:1998:SDO

[Bad98] Mehdi Badii. Simulating the DYNIX operating system parallel programming interface on a UNIX system. Software— Practice and Experience, 28 (5):463–480, May 25, 1998. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
1791;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=1791&PLACEBO=IE.pdf.

Bah:2009:IHE

[Bah09] Taymjong Bah. Inkscape: guide to a vector drawing program. Shortcut. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2009. ISBN 0-13-700628-4 (paperback). xviii + 383 pp. LCCN T385 .B34 2009.

Bailes:1985:DDD

[Bai85] Paul A. Bailes. DDA — a data definition facility for UNIX using awk. Software—Practice and Experience, 15(10):1011– 1020, October 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baird:1987:TUI

[Bai87] W. Gus Baird. Tailoring UNIX for introductory CS courses.

SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19 (2):24–28, June 1987. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Bailey:1996:PUP

[Bai96] David Bailey. PC UNIX — Plan 9 from AT&T. UNIX Review, 14(1):27–??, ???? 1996. CODEN UNRED5. ISSN 0742-3136.

Bailey:1999:SCC

[Bai99] Patrick Bailey. A signal command and control class for Unix. C/C++ Users Journal, 17(3):??, March 1999. CODEN CCUJEX. ISSN 1075-2838.

Baines:2000:SBB

[Bai00] Dominic Baines. SAMBA Black Book. Coriolis Group Books, Scottsdale, AZ, USA, 2000. ISBN 1-57610-455-9. xxiv + 582 pp. LCCN QA76.76.O63 B3446 2000. US$49.99.

Baker:1975:SPW

[Bak75] Brenda S. Baker. Struct — a program which structures Fortran. Technical Memorandum 1084 (TM 75-1271-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 22, 1975. ?? pp. See also [Bak77].

Baker:1976:ASF

[Bak76] Brenda S. Baker. An algorithm for structuring flowgraphs. Technical Memorandum 1103 (TM 76-1271-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 14, 1976. ?? pp.

Baker:1977:ASF

[Bak77] Brenda S. Baker. An algorithm for structuring flowgraphs. Journal of the ACM, 24 (1):98–120, January 1977. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Baumann:2019:FR

[BAKR19] Andrew Baumann, Jonathan Appavoo, Orran Krieger, and Timothy Roscoe. A fork() in the road. In ????, editor, Proceedings of the 17th Workshop on Hot Topics in Operating Systems, May 2019, pages 1–9. ACM Press, New York, NY, USA, 2019. URL https:/

/www.microsoft.com/en-us/
research/uploads/prod/2019/
04/fork-hotos19.pdf.

Blewett:1992:XWB

[BAKW92] Doug Blewett, Scott Anderson, Meg Kilduff, and Mike Wish. X widget based software tools for UNIX. In USENIX [USE92d], pages 111–124.

Ballinger:1994:ETD

[Bal94] Carrie Ballinger. Evolving teradata decision support for massively parallel processing with UNIX. SIGMOD Record (ACM Special Interest Group on Management of Data), 23(2):490, June 1994. CODEN SRECD8.

ISSN 0163-5808 (print), 1943-5835 (electronic).

Ball:1996:UWN

[Bal96] Cheryl Ball. UNIX and Windows NT square off — the fate of UNIX as the database client/ server operating environment of choice is becoming increasingly clouded. Datamation, 42(5):75– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Ball:1999:STY

[Bal99] Bill Ball. Sams’ teach yourself Linux in 24 hours. SAMS Publishing, Indianapolis, IN, USA, second edition, 1999. ISBN 0-672-31526-2. xv + 574 pp. LCCN QA76.76.O63 B358 1999. Edited by Jeff Koch and Jane Brownlow.

Ball:2000:SLU

[Bal00] Bill Ball. SuSE Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31780-X. xxix + 1096 pp. LCCN QA76.76.O63 B355 2000.

Ballew:2005:DYM

[Bal05] Joli Ballew. Degunking your Mac. Paraglyph Press, Scottsdale, AZ, USA, tiger edition, 2005. ISBN 1-933097-05-1. xxvii + 396 pp. LCCN QA76.76.S64 B35 2005.

Bambara:1998:MUS

[Bam98] Richard Bambara. MVS and UNIX: a survival handbook for users, developers, and managers in a multiplatform environment. McGraw-Hill, New York, NY, USA, 1998. ISBN 0-07-006663-9. xxii + 521 pp. LCCN QA76.76.O63B362 1998. US$65.00.

Banahan:1982:LSB

[Ban82] Mike Banahan. The loosing of the sticky bit, or how to speed up your UNIX. EUUG Newsletter, 2(4):125–128, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Banahan:1988:CBF

[Ban88] Mike Banahan. The C book, featuring the draft ANSI C standard. The Instruction set. Addison-Wesley, Reading, MA, USA, 1988. ISBN 0-201-17370-0 (paperback). xi + 268 pp. LCCN QA76.73.C15 B36 1988.

Barrett:1983:IUI

[Bar83] P. L. Barrett. An implementation of UNIX for the Intel iAPX286. In Software Tools Users Group [Sof83], pages 79– ?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Barr:1986:UGD

[Bar86] W. J. Barr. Unix-based gateway to distributed database systems. The Journal of Systems and Software, 6(3):225– 235, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Barr:1988:COS

[Bar88a] John R. Barr. Co-resident operating system: UNIX and realtime distributed processing. In Association [Ass88c], pages 47– 53.

Barr:1988:CRO

[Bar88b] John R. Barr. Co-resident operating system: UNIX and realtime distributed processing. In Association [Ass88c], pages 47– 53.

Barkakati:1991:UDG

[Bar91a] Nabajyoti Barkakati. UNIX Desktop Guide to X/Motif. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1991. ISBN 0-672-22836-X. xviii + 406 pp. LCCN QA76.76.W56 B357 1991. US$27.95.

Barkakati:1991:XWS

[Bar91b] Nabajyoti Barkakati. X Window System Programming. SAMS Publishing, Indianapolis, IN, USA, 1991. ISBN 0-672-22750-9. xxxii + 750 pp. LCCN QA76.76.W56 B36 1991. US$29.95, CDN$38.95.

Barkakati:1994:XWS

[Bar94]  Nabajyoti Barkakati.  X

Window System programming. SAMS Publishing, Indianapolis, IN, USA, second edition, 1994. ISBN 0-672-30542-9. xxiv + 980 pp. LCCN QA76.76.W56 B36 1994. System requirements for computer disk: PC with 486 microprocessor; 8MB RAM; DOS; X Window System; SVGA graphics capabilities; hard disk with 300MB free space.

Bar:2000:LI

[Bar00a] Moshe Bar. Linux internals. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212598-5. xv + 351 pp. LCCN QA76.76.O63 B362 2000.

Baran:2000:NVC

[Bar00b] Nicholas Baran. News and views: Computing problem solved; Scalable Vecotr Graphics spec released; Intel announces XScale microarchitecture; Windows not the only target at LinuxWorld; NIST, robotics industry move towards interoperability. Dr. Dobb’s Journal of Software Tools, 25 (11):18, November 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVE

[Bar00c] Nicholas Baran. News and views: ERP can spell disaster; Linux goes super; dot-com companies bet the farm; the tiniest transistor; and speaking of tiny. … Dr. Dobb’s Journal of Software Tools, 25 (2):18, February 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ieee.
                                        org/conference/iedm.

Bar:2001:LFS

[Bar01a] Moshe Bar. Linux file systems. Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212955-7 (set), 0-07-212954-9 (book), 0-07-212953-0 (CD-ROM). xiv + 348 pp. LCCN QA76.76.O63 B3626 2001. Includes CD-ROM.

Baran:2001:NVW

[Bar01b] Nicholas Baran. News and views: WSDL goes to W3C for standardization; shortchanging science; EUVL may keep Moore’s Law going; spy satellites to generate high-tech jobs; Mexican government adopts Linux; supercomputer on a chip in the works; brain scan database goes public. Dr. Dobb’s Journal of Software Tools, 26(6):18, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Barkakati:2001:RLS

[Bar01c]  Nabajyoti Barkakati.  Red

Hat Linux 7.1 secrets. Hungry Minds, Indianapolis, IN, 2001. ISBN 0-7645-4771-2 (paperback). xxxiv + 1135 pp. LCCN QA76.76.O63 B366166 2001.

Barrett:2004:LPG

[Bar04]  Daniel J. Barrett.  Linux

pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00628-4. viii + 191 pp. LCCN QA76.76.O63 B3685 2004. URL http:/

/www.oreilly.com/catalog/
9780596006280.

Basler:1987:IKM

[Bas87] Werner Basler. Interprozess-Kommunikations-Mechanismen:

Beschreibung der Interprozess-Kommunikations-Mechanismen in den UNIX-Betriebssystemen 4.2 BSD und AT&T System V. (German) [interprocess communication mechanisms: Description of the interprocess communication mechanisms in the UNIX 4.2 BSD and AT&T System V operating systems]. Semesterarbeit, Institut für Informatik der Universität Zürich, Zürich, Switzerland, 1987. 65 pp.

Basavaiah:1991:MIC

[Bas91] Muralidhar Basavaiah. Mach interprocess communication server and network server on Berkeley UNIX. Thesis (M.S.), Arizona State University, Tempe, AZ, USA, 1991. viii + 91 pp.

Bauer:2002:BSS

[Bau02] Michael D. Bauer. Building Secure Servers with Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00217-3. xv + 430 pp. LCCN TK5105.59 .B38 2002. US$44.95. URL http:/

/www.oreilly.com/catalog/
9780596002176;   http://
www.oreilly.com/catalog/
bssrvrlnx.

Bauer:2003:AUL

[Bau03]  Kirk Bauer.  Automating

UNIX and Linux Administration. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-212-3. 600 (est.) pp. LCCN ????

Bach:1984:MUS

[BB84] Maurice J. Bach and Steven J.

Buroff. A multiprocessor UNIX system. In Software Tools Users Group [Sof84], pages 174–177. LCCN QA76.8.U65 U83 1984.

Bregman:1993:USV

[BB93] Phyllis Bregman and Sally Browning. UNIX System V Performance Management. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-016429-1. xvi + 360 pp. LCCN QA76.76.O63 B74 1993. Incorporating the material in the official SVR4 manuals, this book offers tips on tuning your system for better performance …. This is one title in the P. H. Open Systems Library, others covers subjects like NFS administration, print services ….

Bollinger:1999:GEI

[BB99] Terry Bollinger and Peter H.

Beckman. Guest Editors’ introduction: Linux on the move. IEEE Software, 16(1):30–35, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1999/pdf/
s1030.pdf.

Borntrager:2020:SYC

[BBB^{+}20] C. Bornträger, J. D. Bradbury, R. Bündgen, F. Busaba, L. C. Heller, and V. Mihajlovski. Secure your cloud workloads with IBM Secure Execution for Linux on IBM z15 and LinuxONE III. IBM Journal of Research and Development, 64(5/6):2:1–2:11, 2020. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Banahan:1991:CBF

[BBD91] Mike Banahan, Declan Brady, and Mark Doran. The C book, featuring the ANSI C Standard. The Instruction set. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-54433-4. xiii + 310 pp. LCCN QA76.73.C15 B36 1991.

Beck:1996:LKI

[BBD^{+}96] Michael Beck, Harold Bohme, Mirko Dzladzka, Ulrich Kunitz, Robert Magnus, and Dirk Verworner. Linux Kernel Internals. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-87741-4. xvii + 438 pp. LCCN QA76.76.O63L54813 1996. US$38.68. Forward by Linus Torvalds, creator of LINUX.

Banahan:2000:PLD

[BBD^{+}00] Mike Banahan, Michael Boerner, Ian Dickson, Jonathon Kelly, Nikhilesh Kumar Mandalay, Richard Ollerenshaw, Jonathan Pinnock, Ganesh Prasad, Joel Rowbottom, Geoff Sherlock, and Mark Wilcox. Professional Linux Deployment. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-287-4. xvii + 653 pp. LCCN QA76.76.O63 P76124 2000. US$49.99.

Bhat:2023:SCB

[BBD23] Parnika Bhat, Sunny Behal, and Kamlesh Dutta. A system call-based Android malware detection approach with homogeneous and heterogeneous ensemble machine learning. Computers & Security, 130(??):??, July 2023. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167404823001876.

Buchman:1985:DFA

[BBG85] C. Buchman, D. M. Berry, and J. Gonczarowski. DITROFF/ FFORTID, an adaptation of the UNIX DITROFF for formatting bidirectional text. ACM Transactions on Office Information Systems, 3(4):380– 397, October 1985. CODEN ATOSDO. ISSN 0734-2047.

URL http://www.acm.org:
80.

Borg:1989:FTU

[BBG^{+}89] Anita Borg, Wolfgang Blau, Wolfgang Graetsch, Ferdinand Herrmann, and Wolfgang Oberle. Fault tolerance under UNIX. ACM Transactions on Computer Systems, 7(1):1– 24, February 1989. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1989-7-1/p1-borg/.

Bagnara:2018:MCC

[BBH18] Roberto Bagnara, Abramo Bagnara, and Patricia M. Hill. The MISRA C coding standard and its role in the development and analysis of safetyand security-critical embedded software. In Static Analysis, pages 5–23. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-99725-4. LCCN ????

Bhattiprolu:2008:VSC

[BBHL08] Sukadev Bhattiprolu, Eric W. Biederman, Serge Hallyn, and Daniel Lezcano. Virtual servers and checkpoint/restart in mainstream Linux. Operating Systems Review, 42(5):104– 113, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bach:1986:UTS

[BBJ86] Fred Bach, Adalbert Baur, and Christian Jansen, editors. UNIX-Tabellenbuch für die Systeme UNIX Version 7, UNIX System III, UNIX System V, SINIX, XENIX 286, 4.2 BSD, XENIX 86. Hanser, Munich, Germany; Vienna, Austria, 1986. ISBN 3-446-14622-9. 287 (est.) pp. LCCN ????

Bagley:2013:RSVb

[BBK13] Steven R. Bagley, David F. Brailsford, and Brian W. Kernighan. Revisiting a summer vacation: Digital restoration and typesetter forensics.

In Proceedings of the 2013 ACM Symposium on Document Engineering, pages 3– 12. ACM Press, New York, NY, USA, 2013. URL https:

//dl.acm.org/doi/10.1145/
2494266.2494275.

Bresnahan:1984:WNC

[BBM84] J. Barrie Bresnahan, David T.

Barnard, and Ian A. Macleod. WSH: a new command interpreter for UNIX. Software—Practice and Experience, 14(12):1197–1205, December 1984. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Birns:1985:UPM

[BBM85] Peter M. Birns, Patrick B.

Brown, and John C. C. Muster. UNIX for People — a Modular Guide to the UNIX Operating System — Visual Editing, Document Preparation, & Other Resources. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-937459-0 (hardcover), 0-13-937442-6 (paperback). xiii + 528 pp. LCCN QA76.6 .B5725 1985.

Baskett:1980:SWT

[BBNS80] Forest Baskett, Andreas Bechtolsheim, Bill Nowicki, and John Seamon. The SUN workstation: a terminal system for the Stanford University Network. Draft report, Computer Science Department, Stanford University, Stanford, CA 94305-2192, USA, March 30,

                                        1980. URL https://stacks.
                                        stanford.edu/file/druid:
                                        gg867qx3134/gg867qx3134.
                                        pdf.

Bechtolsheim:1982:SWA

[BBP82] Andreas Bechtolsheim, Forest Baskett, and Vaughan Pratt. The SUN workstation architecture. Technical Report 229, Computer Systems Laboratory, Department of Electrical Engineering and Computer Science, Stanford University, Stanford, CA 94305-2192, USA, March 1982.

                                        URL  http://i.stanford.
                                        edu/pub/cstr/reports/csl/
                                        tr/82/229/CSL-TR-82-229.
                                        pdf.

Brutch:1999:IUH

[BBP99] Paul C. Brutch, Tasneem G. Brutch, and Udo Pooch. Indicators of UNIX host compromise. ;login: the USENIX Association newsletter, 24(5s): ??, September 1999. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

                                        tronic). URL http://www.
                                        usenix.org/publications/
                                        login/1999-9/features/compromise.

html. Special issue on intrusion

detection.

Berman:1995:TPF

[BBS95] Andrew Berman, Virgil Bourassa, and Erik Selberg. TRON: Process-specific file protection for the UNIX operating system. In USENIX [USE95b], pages 165–175. ISBN 1-880446-67-7.

LCCN QA 76.76 O63 U88 1995.

URL  http://www.usenix.
org/publications/library/
proceedings/neworl/bourassa.
html.

Bovet:2000:ULK

[BC00] D. (Daniele) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 0-596-00002-2. xvi + 684 pp. LCCN QA76.76.O63 B665 2001.

Bovet:2001:ULK

[BC01] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00002-2. xvi + 684 pp. LCCN QA76.76.O63 B665 2001; QA76.76.O63 B674 2001; QA76.76.O63 B683 2001.

URL http://www.oreilly.
com/catalog/9780596000028.

Bovet:2002:ULK

[BC02] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00213-0. xv + 765 pp. LCCN QA76.76.O63 B683 2003. US$49.95. URL http:/

/www.oreilly.com/catalog/
linuxkernel2.

Bovet:2003:ULK

[BC03] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux Kernel. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00213-0. xv + 765 pp. LCCN QA76.76.O63 B683 2003. US$49.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596002138;   http://
                                        www.oreilly.com/catalog/
                                        linuxkernel2.

Bovet:2006:ULK

[BC06] Daniel P. (Daniel Pierre) Bovet and Marco Cesati. Understanding the Linux kernel. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2006. ISBN 0-596-00565-2 (paperback). xvi + 923 pp. LCCN QA76.76.O63 B683 2006. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596005658.

Boyer:1991:SOO

[BCC^{+}91] F. Boyer, J. Cayuela, P. Y. Chevalier, A. Freyssinet, and Daniel Hagimont. Supporting an object-oriented distributed system: Experience with UNIX, Mach and Chorus. In USENIX [USE91e], page ?? ISBN ???? LCCN ????

Bukhres:1993:IMP

[BCE^{+}93] Omran A. Bukhres, Jiansan Chen, Ahmed K. Elmagarmid, Xiangning Liu, and James G. Mullen. InterBase: a multidatabase prototype systems. SIGMOD Record (ACM Special Interest Group on Management of Data), 22(2):534–539, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Bumbulis:1991:IUW

[BCGS91] Peter Bumbulis, Donald Cowan, Eric Giguere, and Terry Stepien. Integrating UNIX within a microcomputer-oriented development environment. In USENIX [USE91b], pages 29– 36.

Bhattacharyya:1988:VPC

[BCM88] Mitali Bhattacharyya, David Cohrs, and Barton Miller. Visual process connector for UNIX. IEEE Software, 5(4): 43–50, July 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Bukhres:1993:ISB

[BCP93] Omran Bukhres, Jiansan Chen, and Rob Pezzoli. An Inter-Base system at BNR. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):426–429, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Blackford:1994:QIG

[BD94] S. Blackford and J. Dongarra. Quick installation guide for LAPACK on Unix systems. LAPACK Working Note 81, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 1994. URL http:

//www.netlib.org/lapack/
lawns/lawn81.ps;   http:
//www.netlib.org/lapack/
lawnspdf/lawn81.pdf. UT-

CS-94-249, September, 1994.

Bender:1993:UNM

[BDD^{+}93] Michael Bender, Alexander Davidson, Clark Dong, Steven Drach, Anthony Glenning, Karl Jacob, Jack Jia, James Kempf, Nachiappan Periakaruppan, Gale Snow, and Becky Wong. UNIX for nomads: Making UNIX support mobile computing. In USENIX [USE93e], pages 53– 67. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/mobile93/bender.
html.

Brown:2006:SLM

[BDD^{+}06] Mark Brown, Chuck Davis, William Dy, Paul Ionescu, Jeff Richardson, Kurt Taylor, and Robbie Williamson. Solaris to Linux Migration: a Guide for System Administrators. IBM redbooks. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9608-1. xx + 424 pp. LCCN ???? URL

http://www.redbooks.ibm.
com/redbooks/pdfs/sg247186.

pdf. IBM order number SG24-7186-00.

Brooksbank:2000:SAH

[BDH00] Ed (Edward) Brooksbank, Lisa A. Doyle, and George Haberberger. Samba administrator’s handbook. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-4636-8. xxii + 518 pp. LCCN QA76.9.C55 B76 2000.

Bautts:2005:LNA

[BDP05] Tony Bautts, Terry Dawson, and Gregor N. Purdy. Linux network administrator’s guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00548-2. xxii + 338 pp. LCCN QA76.76.O63; QA76.76.O63 K566 2005; QA76.76.O63 K566 2005eb. URL http:/

/www.oreilly.com/catalog/
9780596005481.

Bina:1988:FFB

[BE88] Eric Jon Bina and Perry A.

Emrath. A faster fsck for BSD UNIX. Technical Report CSRD 823, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, October 1988. 12 pp.

Bina:1989:FFB

[BE89] Eric J. Bina and Perry A. Emrath. A faster fsck for BSD UNIX. In USENIX Association [USE89f], pages 173–185.

Bing:2000:EUS

[BE00] Matthew Bing and Carl Erickson. Extending UNIX system logging with SHARP. In USENIX [USE00b], page ?? ISBN 1-880446-13-8. LCCN

????  URL http://www.
usenix.org/publications/
library/proceedings/lisa2000/
bing.html.

Bean:1993:BIO

[Bea93] Gary M. C. Bean. Bean’s Index to OSF/Motif Documentation for Application Programmers. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1993. ISBN 0-89435-438-8. LCCN QA76.76.W56 B4 1993. US$29.95.

Beardsley:1998:BRU

[Bea98] Reginald Beardsley. Book review: UNIX Power Tools, Second Edition. ;login: the USENIX Association newsletter, 23(1):??, February 1998. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Beale:2004:SLB

[Bea04] Jay Beale. Securing Linux the Bastille Way. Addison-Wesley, Reading, MA, USA, 2004. ISBN ????

????  pp.  LCCN ????

URL http://www.bastille-
linux.org/;   http://www.
linuxsecurity.com/feature_
stories/feature_story-59.

html. To appear.

Becker:1979:VGZ

[Bec79] R. A. Becker. Version 3 of the GR-Z Graphical System. Technical Memorandum 1464 (TM 79-1215-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 4, 1979. ?? pp.

Becker:1984:ELM

[Bec84] Richard A. Becker. Experiences with a large mixedlanguage system running under the UNIX operating system. In Software Tools Users Group [Sof84], pages 326–331. LCCN QA76.8.U65 U83 1984.

Becker:2006:CSL

[Bec06] Pete Becker. The C++ Standard Library Extensions: a Tutorial and Reference. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-41299-0 (hardback). ???? pp. LCCN QA76.73.C153 B43 2006. URL http:

//www.awprofessional.com/
bookstore/product.asp?isbn=
0321412990;   http://www.
loc.gov/catdir/toc/ecip0613/
2006014959.html;   http:/
/www.unixreview.com/documents/
s=10094/ur0608k/ur0608k.
html.

Becker:2007:CSL

[Bec07] Pete Becker. The C++ Standard Library extensions: a tutorial and reference. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-321-41299-0 (hardback). xv + 595 pp. LCCN QA76.73.C153 B43

                                        2007. URL http://www.loc.
                                        gov/catdir/toc/ecip0613/
                                        2006014959.html.

Bell:1985:CSC

[BED^{+}85] C. Gordon Bell, Steve Emmerich, Ivor Durham, Daniel P. Siewiorek, and Andrew Wilson. Computer structures are changing: Will UNIX change with them? In USENIX Association [USE85d], pages 1–4. LCCN QA76.8.U65 U8 1985.

Beebe:1985:PUTb

[Bee85] Nelson H. F. Beebe. The <PLOT79> UNIX tar tape [Release 2.3]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, October 2, 1985. 50 pp.

Beebe:1986:PUT

[Bee86a] Nelson H. F. Beebe. The <PLOT79> UNIX tar tape [Release 2.3a]. Technical report, College of Science Computer, University of Utah, Salt Lake City, UT 84112, USA, January 11, 1986. 58 pp.

Beeker:1986:ISU

[Bee86b] Etienne Beeker. Image synthesis with UNIX. EUUG Newsletter, 6(1):4–9, Spring 1986. CO-DEN EONLE8. ISSN 1011-4211.

Beebe:1987:UTU

[Bee87] Nelson H. F. Beebe. UNIX for TOPS-20 users. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, May 28, 1987. 29 pp.

Beebe:1990:DGE

[Bee90] Nelson H. F. Beebe. The design of GMAKE—an extended implementation of UNIX MAKE. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, February 5, 1990. 32 pp.

Beebe:1993:BBAa

[Bee93a] Nelson H. F. Beebe. A Bibliography of Books about the Internet and Networking, 1993. This bibliography in BibTEX format includes the Quarterman list [Qua93], and many others. It, and many others maintained by the same author, are updated frequently, and available for anonymous ftp

on ftp.math.utah.edu in /

pub/tex/bib. Via e-mail, send

mail to tuglib@math.utah.

edu with the text help and

send index from tex/bib.

Beebe:1993:BBAb

[Bee93b] Nelson H. F. Beebe. A bibliography of books about UNIX and UNIX programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, September 30, 1993. 54 pp. URL https:

                                        //www.math.utah.edu/pub/
                                        tex/bib/index-table-u.html#

unix. This report is updated frequently.

Beebe:1994:BBAb

[Bee94] Nelson H. F. Beebe. A bibliography of books about UNIX and UNIX programming. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, 1994. URL https:

                                        //www.math.utah.edu/pub/
                                        tex/bib/unix.*. This report

is updated frequently.

Beebe:1996:CBPd

[Bee96a] Nelson H. F. Beebe. A complete bibliography of publications in SysAdmin: The Journal for UNIX Systems Administrators. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 31, 1996. 27 pp. This report is updated frequently.

Beebe:1996:CBPe

[Bee96b] Nelson H. F. Beebe. A complete bibliography of publications in SysAdmin: The Journal for UNIX Systems Administrators. Technical report, Center for Scientific Computing, Department of Mathematics, University of Utah, Salt Lake City, UT 84112, USA, August 31, 1996. 27 pp. URL https:

                                        //www.math.utah.edu/pub/

tex/bib/index-table-s.html#

sysadmin. This report is updated frequently.

Beekmans:2000:LS

[Bee00]  Gerard Beekmans.  Linux

from scratch. IUniverse.com, Inc., San Jose, CA, USA, 2000. ISBN 0-595-13765-2. xvi + 95 pp. LCCN QA76.76.O63. URL http://

www.linuxfromscratch.org/.

Beekmans:2016:LS

[Bee16]  Gerard Beekmans.  Linux

from scratch. Web site.,

2016.  URL http://www.
linuxfromscratch.org//lfs/
faq.html.

Blewett:1986:MBI

[BEHW86a] C. D. Blewett, J. T. Edmark, J. I. Helfman, and M. Wish. A multi-representation, bitmap interface to the UNIX file system constructed from cooperating processes. In USENIX Association [USE86d], pages 41– 48. ISBN ???? LCCN ????

Blewett:1986:MRB

[BEHW86b] C. D. Blewett, J. T. Edmark, J. I. Helfman, and M. Wish. A multi-representation, bitmap interface to the UNIX file system constructed from cooperating processes. In USENIX Association [USE86d], pages 41– 48. ISBN ???? LCCN ????

Bell:1979:UPMa

[Bel79a] Bell Laboratories. UNIX programmer’s manual. Technical Memorandum 1400, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Bell:1979:UPMb

[Bel79b] Bell Laboratories. UNIX programmer’s manual. Technical Memorandum 1401, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Bell:1979:UTS

[Bel79c] Bell Laboratories. UNIX/32V time-sharing system: UNIX programmer’s manual, version 1.0, volume 1. Technical Memorandum 1402, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1979. ?? pp.

Bell:1979:URU

[Bel79d] Bell Laboratories. UNIX/RT user’s manual — release 1. Technical Memorandum 1070, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1979. ?? pp.

Bellomo:2000:MRL

[Bel00]  Michael Bellomo.  Mas-

ter Red Hat Linux Visually. IDG Books, San Mateo, CA, USA, 2000. ISBN 0-7645-3436-X. 56 pp. LCCN ?QA76.76.O63 B44982 2000??? US$34.99. URL

                                        http://catalog.idgbooks.
                                        com/product.asp?isbn=076453436X.

Bellovin:2018:UAE

[Bel18] Steven M. Bellovin. Usenet, authentication, and engineering (or: Early design decisions for Usenet). Web article., February 23, 2018. URL https://

www.cs.columbia.edu/~smb/
blog/2018-02/2018-02-23.
html.

Bentley:1986:PP

[Ben86]  Jon Louis Bentley.  Pro-

gramming Pearls. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-10331-1 (paperback). viii + 195 pp. LCCN QA76.6.B453 1986. Reprinted with corrections.

Bentley:1988:DPT

[Ben88a] J. L. Bentley. DFORMAT — a program for typesetting data formats. Computing Science Technical Report 142, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1988. i + 13 pp.

URL https://github.com/
arnoldrobbins/dformat;
https://web.archive.org/
web/*/http://cm.bell-labs.
com/cm/cs/cstr/142.ps.gz.

142.

Bentley:1988:MPP

[Ben88b] Jon Louis Bentley. More Programming Pearls: Confessions of a Coder. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-11889-0. viii + 207 pp. LCCN QA76.6 .B452 1988. US$18.75.

Benhenni:1998:AAU

[Ben98] Robert A. Benhenni. An approximate analysis of a UNIX macro process scheduler. IN-FORMS Journal on Computing, 10(3):351–357, Summer 1998. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

                                        //pubsonline.informs.org/
                                        doi/abs/10.1287/ijoc.10.
                                        3.351.

Bentley:2000:PP

[Ben00] Jon Louis Bentley. Programming Pearls. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-65788-0 (paperback), 0-13-449802-X (ebook), 0-13-449805-4. xi + 239 pp. LCCN QA76.6 .B454 2000. US$24.95.

Benvenuti:2006:ULN

[Ben06] Christian Benvenuti. Understanding Linux network internals. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00255-6. xxiv + 1035 pp. LCCN QA76.76.O63; QA76.76.O63 B46 2006eb; QA76.76.O63 B46 2006. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596002558.

Benjamin:2009:BI

[Ben09] Donna Benjamin. Beginning Inkscape. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-2513-0. LCCN ????

Bernstein:1976:SPU

[Ber76] D. R. Bernstein. Synthetic process for UNIX. Technical Memorandum 1114 (TM 76-8234-17, TM 76-9156-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 24, 1976. 20 pp.

Berens:1985:APU

[Ber85a] Peter H. Berens. Array processing under UNIX. In USENIX Association [USE85d], pages 175–181. LCCN QA76.8.U65 U8 1985.

Berkman:1985:GVU

[Ber85b] Jerry Berkman. A guide for VAX UNIX Fortran users. Unx; 2.3.3 unx; 2.3.3., University of California, Berkeley, Academic Computing Services, Berkeley, CA, USA, September 19, 1985. iii + 62 pp.

Berglund:1986:IV

[Ber86] Eric J. Berglund. An introduction to the V-system. IEEE Micro, 6(4):35–52, July/August 1986. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berlage:1991:OMC

[Ber91] Thomas Berlage. OSF/Motif — Concepts and Programming. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-55792-4. xiv + 487 pp. LCCN QA76.76.W56 B4613 1991.

Berry:1992:WGC

[Ber92] John Berry. The Waite Group’s C++ Programming. Howard W. Sams, Indianapolis, IN 46268, USA, second edition, 1992. ISBN 0-672-22771-1. xvii + 408 pp. LCCN QA76.73.C153B47 1992. A fine book on OOP with C++ (for UNIX and DOS).

Bernaschi:1998:EMP

[Ber98]  Massimo Bernaschi.  Effi-

cient message passing on UNIX shared memory multiprocessors. Future Generation Computer Systems, 13(6):443–449, May 20, 1998. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.elsevier.com/
gej-ng/10/19/19/28/20/18/
abstract.html.

Best:2005:LDP

[Bes05] Steve (Steve Francis) Best. Linux debugging and performance tuning: tips and techniques. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-149247-0. ???? pp. LCCN QA76.76.O63 B4756 2005.

URL ftp://uiarchive.cso.
uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/
toc/ecip0514/2005017192.
html;   http://www.phptr.
com/title/0131492470.

Beyer:1977:RMR

[Bey77] J. D. Beyer. A record manager for random access to variable length records under UNIX. Technical Memorandum 1171 (TM 77-1353-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1, 1977. ?? pp.

Beyls:1987:NUT

[Bey87] Pascal Beyls. Now UNIX talks to me in my language. In USENIX Association [USE87d], pages 307–321.

Brown:1980:EPB

[BF80] W. S. Brown and S. I. Feldman. Environment parameters and basic functions for floatingpoint computation. ACM Transactions on Mathematical Software, 6(4):510–523, December 1980. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/doi/10.1145/
355921.355924.

Butzen:1997:LD

[BF97] Fred Butzen and Dorothy Forbes. The Linux Database. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-491-5. xiv + 561 pp. LCCN QA76.9.D3B893 1997. US$39.95, CDN$55.95. URL

http://www.mispress.com/.

Beale:2003:SID

[BF03] Jay Beale and James C. Foster.

Snort 2.0 Intrusion Detection. Syngress, Rockland, MA, USA, 2003. ISBN 1-931836-74-4, 3-8266-1304-X. xxviii + 523 pp. LCCN ???? US$49.95. Jeffrey Posluns, technical advisor and Brian Caswell, technical editor.

Brown:1979:EPB

[BF9 ] W. S. Brown and S. I. Feldman. Environment parameters and basic functions for floatingpoint computation. Computing Science Technical Report 72, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1979 (??). ?? pp.

Ballance:1989:UIR

[BFGK89] Charles Ballance, Sean Fleming, Jay Goldberg, and Nelly Karasik. UFOS: An intelligent real-time performance monitor for UNIX System V. In USENIX Association [USE89e], pages 277–286. LCCN QA 76.76 O63 U83 1989.

Bentley:1993:TDI

[BFKS93] Jon L. Bentley, Mary F. Fernandez, Brian W. Kernighan, and Norman L. Schryer. Template-driven interfaces for numerical subroutines. ACM Transactions on Mathematical Software, 19(3):265–287, September 1993. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL http://www.acm.org/
                                        pubs/citations/journals/
                                        toms/1993-19-3/p265-bentley/

.

Bonomi:1989:DPL

[BFS89] F. Bonomi, P. J. Fleming, and P. D. Steinberg. Distributing processes in looselycoupled UNIX multiprocessor systems. In USENIX Association [USE89e], pages 61–72. LCCN QA 76.76 O63 U83 1989.

Beshers:2007:ERU

[BFS07] Clifford Beshers, David Fox, and Jeremy Shaw. Experience report: using functional programming to manage a Linux distribution. ACM SIGPLAN Notices, 42(9):213– 218, September 2007. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonanni:1977:SPU

[BG77] L. E. Bonanni and A. L.

Glasser. SCCS/PWB user’s manual. Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1977. 2 + 22 pp.

Bach:1988:MFS

[BG88] Maurice J. Bach and Ron Gomes. Measuring file system activity in the UNIX system. In USENIX Association [USE88c], pages 43–52. ISBN ???? LCCN ????

Brankin:1997:ARF

[BG97] R. W. Brankin and I. Gladwell.

Algorithm 771. rksuite 90: Fortran software for ordinary differential equation initial value problems. ACM Transactions on Mathematical Software, 23(3):402–415, September 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1997-23-3/p402-brankin/.

Berkley:1990:TU

[BGH^{+}90] Robert Berkley, Skip Gilbrech, Timothy Hunt, Mark Luppi, and Richard Plevin. Tickerplants on UNIX. In USENIX [USE90d], pages 65–74. LCCN QA76.8.U65 U82 1990.

Batagini:2001:FCA

[BGJ01] Elton Costa Batagini, Mariusz Gieparda, and KyeongWon Jeong. Fast Connect for AIX Version 3.1 guide. Redbooks. IBM Corporation, Austin, TX, USA, second edition, September 2001. ISBN 0-7384-2305-X (paperback). xiv + 211 pp. LCCN QA76.76.O63 B378 2001. Publication number SG24-5527-01.

Bettison:1991:SIU

[BGMR91] Andrew Bettison, Andrew Gollan, Chris Maltby, and Neil Russell. SHARE II — a user administration and resource control system for UNIX. In USENIX [USE91b], pages 51– 60.

Brinsmead:1991:CLP

[BGPS91] Mark Brinsmead, Ken Gamble, Michael Pazzani, and Glenn Silverstein. Common LISP product roundup. AI Expert, 6(6):48–??, June 1991. ISSN 0888-3785.

Beer:1989:DWT

[BGR89] M. D. Beer, S. M. George, and R. Rada. Developing writing tools for UNIX workstations. In Anonymous [Ano89], pages 31– 36. ISBN 0-9513181-3-6. LCCN ????

Barak:1993:MDO

[BGW93] Amnon Barak, Shai Guday, and Richard G. Wheeler. The MOSIX distributed operating system: load balancing for UNIX, volume 672 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. CO-DEN LNCSD9. ISBN 3-540-56663-5 (Berlin), 0-387-56663-5 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). x + 221 pp. LCCN QA76.76.O63 B363 1993.

Bulger:2004:MPD

[BGW04] Brad Bulger, Jay Greenspan, and Dave Wall. MySQL/PHP database applications. Wiley, New York, NY, USA, second edition, 2004. ISBN 0-7645-3799-7 (electronic book). xliii + 764 pp. LCCN QA7673.S67. URL http://

www.netLibrary.com/urlapi.
asp?action=summary&v=1&bookid=
100883.

Brown:1972:A

[BH72] W. S. Brown and A. D. Hall.

ALTRAN. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), ??(24):4–7, October 1972. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Brown:1979:ASA

[BH79] W. S. Brown and A. C. Hearn.

Applications of symbolic algebraic computation. Computer Physics Communications, 17(1–2):207–215, April/ May 1979. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0010465579900821.

Bahill:1986:CES

[BH86] A. Terry Bahill and Pat Harris. Cogito, an expert system to give installation advice for UNIX 4.2BSD. ;login: the USENIX Association newsletter, 11(5):4–8, September/October 1986. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Burk:1997:UUa

[BH97a] Robin Burk and David B. Horvath. UNIX unleashed. SAMS Publishing, Indianapolis, IN, USA, second Internet edition, 1997. ISBN 0-672-31205-0. li + 1114 pp. LCCN QA76.76.O63 B867 1997.

Burk:1997:UUb

[BH97b] Robin Burk and David B. Horvath. UNIX unleashed. SAMS Publishing, Indianapolis, IN, USA, second system administrator’s edition, 1997. ISBN 0-672-30952-1. li + 1342 pp. LCCN QA76.76.O63 B867 1997b. CD-ROM includes: an assortment of third-party tools and product demos; entire text of this book in electronic format; and an electronic version of UNIX unleashed, Internet edition.

Bialaski:2000:SLN

[BH00] Tom Bialaski and Michael Haines. Solaris and LDAP Naming Services: Deploying LDAP in the Enterprise. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-030678-9. xxvii + 372 pp. LCCN QA76.76.O63B518

2001. URL books/ldap.pdf;
 http://www.sun.com/books/
catalog/haines/. Part No.

806-2893-10 October 2000.

Bialaski:2001:SLN

[BH01] Tom Bialaski and Michael Haines. Solaris and LDAP Naming Services: Deploying LDAP in the Enterprise. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-030678-9. xxvii + 372 pp. LCCN QA76.76.O63 B518 2001. US$39.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0130306789.html;   http:
//www.sun.com/books/catalog/
haines/index.html.

Brown:1971:CCP

[BH1 ] W. S. Brown and A. D. Hall.

Computing Campbell’s Y_{2n} polynomials with ALTRAN. Computing Science Technical Report 6, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1971 (??). ?? pp.

Brown:1978:ASA

[BH8 ] W. S. Brown and A. C. Hearn.

Applications of symbolic algebraic computation. Computing Science Technical Report 66, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1978 (??). ?? pp.

Bhaskar:2012:CPP

[Bha12a] K. Bhaskar. C — past, present, and future — a perspective. Resonance, 17(8):748–758, August 2012. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Bhatt:2012:UGD

[Bha12b] Pramod Chandra P. Bhatt. UNIX: Genesis and design features. Resonance, 17(8):727– 747, August 2012. CODEN RE-SOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Bodenstab:1984:UOS

[BHK^{+}84] D. E. Bodenstab, Thomas F. Houghton, Keith A. Kelleman, George Ronkin, and Edward P. Schan. UNIX operating system porting experiences. AT&T Bell Laboratories Technical Journal, 63 (8 part 2):1769–1790, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic). URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/otherports/newp.
                                        pdf.

Brown:1984:CS

[BHM84] W. Stanley Brown, Bernard D. Holbrook, and M. Douglas McIlroy. Computer science. In Millman [Mil84c], chapter 9, pages 351–398. ISBN 0-932764-06-1. LCCN ????

Bourne:1998:UOU

[BHM98] Philip E. Bourne, Richard Holstein, and Joseph Mc-Mullen. UNIX for Open-VMS users. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1998. ISBN 1-55558-155-2 (paperback). xix + 428 pp. LCCN QA76.76.O63 B669

1998. URL http://www.loc.
gov/catdir/description/els033/
98035518.html;   http://
www.loc.gov/catdir/toc/els032/
98035518.html.

Binkley:1995:PIL

[BHR95] David Binkley, Susan Horwitz, and Thomas Reps. Program integration for languages with procedure calls. ACM Transactions on Software Engineering and Methodology, 4(1):3–35, January 1995. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
1995-4-1/p3-binkley/p3-binkley.
pdf;   http://www.acm.org/
pubs/citations/journals/
tosem/1995-4-1/p3-binkley/

.

Baird:1995:DSP

[BI95] H. S. Baird and D. J. Ittner.

Data structures for page readers. In Spitz and Dengel [SD95], pages 3–15. ISBN 981-02-2122-3, 981-02-2122-3. LCCN TK7882.P3I55 1994.

Bialaski:1999:SGW

[Bia99] Tom Bialaski. Solaris Guide for Windows NT Administrators. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1999. ISBN 0-13-025854-7. xiv + 15–168 pp. LCCN QA76.76.O63B52

                                        1999.  URL books/ssg.
                                        pdf;   http://www.phptr.
                                        com/ptrbooks/ptr_0130258547.
                                        html;   http://www.sun.com/
                                        books/catalog/bialaski/. Part

No. 805-7622-10 June 1999, Revision A.

Binding:1985:CCC

[Bin85] C. Binding. Cheap concurrency in C. ACM SIGPLAN Notices, 20(9):21–26, September 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bina:1988:MUF

[Bin88] Eric Jon Bina. Modifications to the UNIX file system check program FSCK for quicker crash recovery. Thesis (M.S.), University of Illinois at Urbana-Champaign, Urbana, IL 61801, USA, August 1988. iv + 51 pp.

Bishop:1987:PUU

[Bis87] Matt Bishop. Profiling under UNIX by patching. Software—Practice and Experience, 17(10):729–739, October 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bishop:1988:AFN

[Bis88] Matthew A. Bishop. Auditing files on a network of UNIX machines. In USENIX Association [USE88e], pages 51– 52. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Bishop:1989:USS

[Bis89] M. Bishop. UNIX security in a supercomputing environment. In ACM [ACM89], pages 693–698. ISBN 0-89791-341-8. LCCN QA 76.5 S87 1989. IEEE 89CH2802-7.

Bivand:1987:UIG

[Biv87] Roger Bivand. A user interface for geographers — what can UNIX offer? In USENIX Association [USE87a], pages 183– 190. ISBN ???? LCCN ????

Brandt:1977:DFA

[BJ77] R. B. Brandt and C. D. January. DEC factory acceptance test. Technical Memorandum 1145 (MF 77-8234-001), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 19, 1977. ?? pp.

Broadbent:1993:ES

[BJ93] Craig Broadbent and Hal Jespersen. Emerging standards. UNIX Review, 11(3): 30–??, March 1993. CODEN UNRED5. ISSN 0742-3136.

Boling:2001:EHK

[BJ01] Eli Boling and Chuck Jazdzewski.

Exception handling in Kylix:

Matching up Object Pascal, Linux, and exception handling. Dr. Dobb’s Journal of Software Tools, 26(11):66, 68–71, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_11/
kylix.txt.

Bentley:1987:CPP

[BJK87] Jon L. Bentley, Lynn W. Jelinski, and Brian W. Kernighan. Chem — a program for phototypesetting chemical structure diagrams. Computers and Chemistry, 11(4):281–297, January 1987. CODEN COCHDK. ISSN 0097-8485 (print), 1879-0763 (electronic).

Bentley:1992:CPT

[BJK92] Jon L. Bentley, Lynn W. Jelinski, and Brian W. Kernighan. CHEM — a program for typesetting chemical diagrams: User manual. Computing Science Technical Report 122, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, October 6, 1992. ii + 31

pp. URL https://9p.io/cm/
cs/cstr/122.ps.gz.

Becker:2022:CPA

[BJM^{+}22] Benedikt Becker, Nicolas Jeannerod, Claude Marché, Yann Régis-Gianas, Mihaela Sighireanu, and Ralf Treinen. The CoLiS platform for the analysis of maintainer scripts in Debian software packages. International Journal on Software Tools for Technology Transfer (STTT), 24(5):717– 733, October 2022. CO-DEN ???? ISSN 1433-2779 (print), 1433-2787 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10009-022-00671-1.

Babaoglu:2004:DIB

[BJP04] Özalp Babaoğlu, William Joy, and Juan Porca. Design and implementation of the Berkeley virtual memory extensions to the UNIX operating system. Report, Computer Science Division, Department of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, California 9472, May 18, 2004. 1–12 pp.

URL http://roguelife.org/
~fujita/COOKIES/HISTORY/
3BSD/design.pdf.

Butler:1984:USH

[BK84] T. W. Butler and L. A.

Kennedy. The UNIX system help facility. In Software Tools Users Group [Sof84], pages 253–257. LCCN QA76.8.U65 U83 1984.

Beck:1985:VAB

[BK85] Bob Beck and Bob Kasten. VLSI assist in building a multiprocessor UNIX system. In USENIX Association [USE85d], pages 255–275. LCCN QA76.8.U65 U8 1985.

Bentley:1986:TPI

[BK86a] Jon L. Bentley and Brian W. Kernighan. Tools for printing indexes. Computing Science Technical Report 128, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1986. ii + 13 pp.

                                        URL https://web.archive.
                                        org/web/*/http://cm.bell-
                                        labs.com/cm/cs/cstr/128.
                                        ps.gz.

Bentley:1986:GLT

[BK86b] Jon Louis Bentley and Brian W. Kernighan. GRAP — a language for typesetting graphs. Communications of the Association for Computing Machinery, 29(8):782–792, August 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0001-0782/
                                        6429.html.

Bolsky:1989:KSC

[BK89] Morris Bolsky and David Korn. The Korn Shell Command and Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-516972-0. xvi + 356 pp. LCCN QA76.73.K67 B64 1989. The authoritative reference …. See also [Val92b].

Bentley:1990:GLT

[BK90a] J. L. Bentley and B. W. Kernighan. Grap — a language for typesetting graphs. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Bentley:1990:SAA

[BK90b] J. L. Bentley and B. W.

Kernighan. A system for algorithm animation. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Bentley:1991:GLT

[BK91a] Jon L. Bentley and Brian W.

Kernighan. Grap — a language for typesetting graphs tutorial and user manual. Computing Science Technical Report 114, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1991. ii + 37 pp.

URL https://web.archive.
org/web/*/http://cm.bell-
labs.com/cm/cs/cstr/114.

ps.gz. Revised edition of August 1984 version.

Bentley:1991:SAAb

[BK91b] Jon L. Bentley and Brian W. Kernighan. A system for algorithm animation. Computing Systems, 4(1):5–30, Winter 1991. CODEN CMSYE2. ISSN 0895-6340. URL http://www.

                                        usenix.org/publications/
                                        compsystems/1991/win_bentley.
                                        pdf;   https://www.usenix.
                                        org/publications/compsystems/
                                        computing-systems-1991.

Bentley:1991:SAAa

[BK91c] Jon L. Bentley and Brian W. Kernighan. A system for algorithm animation tutorial and user manual. Computing Science Technical Report 132, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 6, 1991. ii + 24 pp.

                                        URL https://web.archive.
                                        org/web/*/http://cm.bell-
                                        labs.com/cm/cs/cstr/132.
                                        ps.gz.

Bolsky:1995:NKC

[BK95] Morris I. Bolsky and David G. Korn. The New KornShell Command And Programming Language. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1995. ISBN 0-13-182700-6. 416 pp. LCCN QA76.73.K67 B64 1995. ????

Barrett:2013:BBC

[BK13] Steven F. Barrett and Jason Kridner. Bad to the Bone: crafting electronic systems with BeagleBone and BeagleBone Black, volume 41 of Synthesis lectures on digital circuits and systems. Morgan and Claypool, San Rafael, CA, USA, 2013. ISBN 1-62705-137-6 (paperback), 1-62705-138-4. ISSN 1932-3174. xxv + 397 pp. LCCN QA76.5 .B37 2013; TJ223.P76.

Boykin:1993:PUM

[BKLL93] Joseph Boykin, David Kirschan, Alan Langerman, and Susan LoVerso. Programming under Mach. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52739-1. xvii + 490 pp. LCCN QA76.8.N49 P76 1993.

Bentley:1986:PPLb

[BKM86] Jon Bentley, Donald E. Knuth, and M. Douglas McIlroy. Programming pearls: a literate program: a web program for common words. Communications of the Association for Computing Machinery, 29(6): 471–483, June 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=315653.

See another solution to the problem of printing the k most common words of a text file in [?]. Reprinted in [?, 151–177].

Bayer:1975:MME

[BL75] D. L. Bayer and Heinz Lycklama. MERT — a multienvironment real-time operating system. Technical Memorandum 1062 (TM-75-1352-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 18, 1975. 36 pp.

                                        URL https://www.tuhs.org/
                                        Archive/Documentation/TechReports/
                                        Heinz_Tech_Memos/TM-75-1352-
                                        7_MERT_A_Multi-Environment_
                                        Real-Time_Operating_System_
                                        19751118.pdf.

Blackett:1983:MDR

[Bla83a] Kent Blackett. A menu-driven real-time UNIX system. In USENIX [USE83], pages 279– 283.

Blackett:1983:MRU

[Bla83b] Kent Blackett. A menu-driven real-time UNIX system. In Association [Ass83], pages 279– 283.

Blair:1986:DTS

[Bla86] Steven R. Blair. A defect tracking system for the UNIX environment. Hewlett–Packard Journal, 37(3):15–18, March 1986. CODEN HPJOAX. ISSN 0018-1153.

Blacharski:1998:NSM

[Bla98] Dan Blacharski. Network security in a mixed environment. I D G Books Worldwide, Indianapolis, IN, USA, 1998. ISBN 0-7645-3152-2. xxiv + 408 pp. LCCN TK5105.59 .B55 1998. Covers Windows NT, UNIX, and NetWare Networks.

Blaess:2002:LSS

[Bla02] Christophe Blaess. Langages de scripts sous Linux: Shell Bash, Sed, Awk, Perl, Tcl, Tk, Python, Ruby. Eyrolles, Paris, France, 2002. ISBN 2-212-11028-6. xx + 733 pp. LCCN QA76.7 B4 2002.

Blaess:2004:SSL

[Bla04] Christophe Blaess. Scripts sous Linux: Shell Bash, Sed, Awk, Perl, TCL, Tk, Python, Ruby. Eyrolles, Paris, France, second edition, 2004. ISBN 2-212-11405-2. xxi + 761 pp. LCCN QA76.76O63; QA76.7.

Blevins:1983:PUG

[Ble83] Jack Blevins. The port of UNIX to the Gould 32/27. In Association [Ass83], pages 273–278.

Bai:2020:EDS

[BLH20] Jia-Ju Bai, Julia Lawall, and Shi-Min Hu. Effective detection of sleep-in-atomic-context bugs in the Linux kernel. ACM Transactions on Computer Systems, 36(4):10:1–10:30, June 2020. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
3381990.

Blinn:1996:PSP

[Bli96] Bruce Blinn. Portable shell programming: an extensive collection of Bourne shell examples. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-451494-7. xxii + 281 pp. LCCN QA76.76.O63 B593 1996.

Bloomer:1992:PPR

[Blo92] John Bloomer. Power programming with RPC. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 1992. ISBN 0-937175-77-3. xxxii + 459 pp. LCCN QA76.9.D5 B55 1991. US$29.95.

Blommers:1999:AES

[Blo99] John Blommers. Architecting enterprise solutions with UNIX networking. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-792706-1. 352 pp. LCCN QA76.76.O63B595 1998. US$44.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0137927061.html.

Blue:1975:ANQ

[Blu75]  James L. Blue.  Auto-

matic numerical quadrature — DQUAD. Computing Science Technical Report 25, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1975. ?? pp.

Blue:1977:ANQ

[Blu77] James L. Blue. Automatic numerical quadrature. The Bell System Technical Journal, 56(9):1651–1678, November 1977. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol56/bstj56-9-1651.
pdf;   http://www.alcatel-
lucent.com/bstj/vol56-1977/
articles/bstj56-9-1651.pdf;

https://archive.org/details/
bstj56-9-1651/mode/2up;
https://onlinelibrary.wiley.
com/doi/abs/10.1002/j.1538-
7305.1977.tb00586.x.

Blue:1978:BIS

[Blu78a] J. L. Blue. Boundary integral solutions of Laplace’s equation. The Bell System Technical Journal, 57(8):2797–2822, October 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-8-2797.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-8-2797.pdf;
https://archive.org/details/
bstj57-8-2797/mode/2up.

Blue:1978:PFP

[Blu78b] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. ACM Transactions on Mathematical Software, 4(1):15–23, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/doi/10.1145/
355769.355771.

Blum:2000:RQ

[Blu00a] Richard Blum. Running Qmail.

SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31945-4. xi + 540 pp. LCCN TK5105.74.Q53 B58 2000.

Blum:2000:SLM

[Blu00b] Richard Blum. SOHO Linux Mail Server. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31834-2. x + 526 pp. LCCN TK5105.73 .B58 2000. US$34.95. Edited by Angie Wethington, Neil Rowe, and Tony Amico.

Blum:2001:P

[Blu01] Richard Blum. Postfix. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32114-9. xix + 593 pp. LCCN TK5105.73 .B578 2001. EUR$56.65.

Blum:2008:LCL

[Blu08] Richard Blum. Linux command line and shell scripting bible. Wiley, New York, NY, USA, 2008. ISBN 0-470-25128-X (paperback). xxx + 809 pp. LCCN QA76.76.O63 B598

2008. URL http://www.loc.
gov/catdir/enhancements/
fy0827/2008012238-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0827/2008012238-
t.html;   http://www.loc.
gov/catdir/enhancements/
fy0828/2008012238-b.html.

Blunden:2009:RAE

[Blu09] Bill Blunden. The rootkit arsenal: escape and evasion in the dark corners of the system. Wordware Publishing, Plano, TX, USA, 2009. ISBN 1-59822-061-6 (paperback). xxvii + 908 pp. LCCN QA76.9.A25 2009.

Blum:2011:LCL

[Blu11] Richard S. Blum. Linux Command Line and Shell Scripting Bible. Wiley, New York, NY, USA, second edition, 2011. ISBN 1-118-00442-6 (paperback), 1-118-08777-1 (e-book), 1-118-08778-X (e-book), 1-118-08779-8 (e-book). xxvii + 812 pp. LCCN QA76.76.O63 B598 2011.

Blue:1977:BIS

[Blu7 a] J. L. Blue. Boundary integral solutions of Laplace’s equation. Computing Science Technical Report 60, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Blue:1977:PFP

[Blu7 b] James L. Blue. A portable Fortran program to find the Euclidean norm of a vector. Computing Science Technical Report 45, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Blumer:1979:UUI

[BM79] T. P. Blumer and R. Muha.

UNIX/86: UNIX on the Intel 8086. Technical Memorandum 1483, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1, 1979. ?? pp.

Barnett:1987:IUN

[BM87] Lewis Barnett and Michael K.

Malloy. ILMON: A UNIX network monitoring facility. In USENIX Association [USE87e], pages 133–144. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Burgard:1992:XDC

[BM92] Michael Burgard and Mike Moore. X.desktop Cookbook — An Easy Way to Configure your Desktop. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-978537-X. x + 377 pp. LCCN QA76.76.O63 B858 1992. US$46.00.

Bentley:1993:ESF

[BM93] Jon Louis Bentley and M. Douglas McIlroy. Engineering a sort function. Software—Practice and Experience, 23(11):1249– 1265, November 1993. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bradford:2002:LWI

[BM02] Ed Bradford and Lou Mauget. Linux and Windows interoperability guide. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-032477-9. xv + 600 pp. LCCN QA76.76.O63 B7168 2002 Bar. Configuring, programming, and administering mixed Linux/Windows environments.

Bellasi:2015:ERR

[BMF15] Patrick Bellasi, Giuseppe Massari, and William Fornaciari. Effective runtime resource management using Linux control groups with the BarbequeRTRM framework. ACM Transactions on Embedded Computing Systems, 14(2): 39:1–39:??, March 2015. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Blair:1985:CU

[BMM85] Gordon S. Blair, Jon R. Malone, and John A. Mariani. A critique of UNIX. Software—Practice and Experience, 15(12):1125–1139, December 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brownbridge:1982:NCU

[BMR82] David R. Brownbridge, L. F.

Marshall, and B. Randell. The Newcastle connection or UNIXes of the world unite! Software—Practice and Experience, 12(12):1147–1162, December 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Baker:1994:EPP

[BMR94] T. P. Baker, Frank Mueller, and Viresh Rustagi. Experience with a prototype of the POSIX “Minimal realtime system profile”. In IEEE [IEE94c], pages 12–17. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994.

Brownbridge:2001:NCU

[BMR01] David R. Brownbridge, Lindsay F. Marshall, and Brian Randell. The Newcastle Connection or Unixes of the world unite. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01], pages 528– 549. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425

2001. URL https://link.
springer.com/chapter/10.
1007/978-1-4757-3510-9_24.

Blair:1983:PEU

[BMS83] Gordon S. Blair, John A. Mariani, and W. D. Shepherd. A practical extension to UNIX for interprocess communication. Software—Practice and Experience, 13(1):45– 58, January 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Becker:1995:SIG

[BMS95] George Becker, Mary E. S. Morris, and Kathy Slattery. Solaris Implementation: a Guide for System Administrators. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-353350-6. xvi + 345 pp. LCCN QA76.76.O63B433 1995. US$35.00.

Brandt:1978:UC

[BMV78] R. B. Brandt, J. F. Maranzano, and G. C. Vogel. UNIX courses. Technical Memorandum 1277 (MF 78-8234-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 9, 1978. ?? pp.

Bandel:2001:SEU

[BN01] David A. (David Allan) Bandel and Robert Napier. Special edition using Linux. Que Corporation, Indianapolis, IN, USA, sixth edition, 2001. ISBN 0-7897-2543-6. xv + 817 pp. LCCN QA76.76.O63 B3625 2001.

Borthwick:1987:IWB

[BNB87] Stuart Borthwick, John R.

Nicol, and Gordon S. Blair. An intelligent, window based interface to UNIX. In USENIX Association [USE87a], pages 225– 241. ISBN ???? LCCN ????

Bohannon:1986:RUS

[Boh86] Thomas A. Bohannon. Reflections on A UNIX scheduler. ;login: the USENIX Association newsletter, 11(4):39– 45, July/August 1986. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Bokhari:1995:LOS

[Bok95] Shahid H. Bokhari. The Linux operating system. Computer, 28(8):74–79, August 1995. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bollinger:1999:LPO

[Bol99] Terry Bollinger. Linux in practice: An overview of applications. IEEE Software, 16(1):72–79, January/ February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s1072abs.
htm;   http://dlib.computer.

                                        org/so/books/so1999/pdf/
                                        s1072.pdf.

Boloni:2001:PKC

[Böl01] Lotzi Bölöni. Programming KDE 2.0: Creating Linux Desktop Applications. CMP Books, Lawrence, KS, 2001. ISBN 1-929629-13-3. ix + 265 pp. LCCN QA76.9.U83B65 2001. US$49.95. Includes CD-ROM.

Bonanni:1979:IBD

[Bon79a] L. E. Bonanni. Implementation of a Boolean data analysis technique for PWB/UNIX. Technical Memorandum 1378 (TM 79-9441-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 21, 1979. ?? pp.

Bonanni:1979:TER

[Bon79b] L. E. Bonanni. Tools for the extraction and reduction of PWB/UNIX software monitor data. Technical Memorandum 1474 (TM 79-3781-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 15, 1979. ?? pp.

Bookman:2003:LCB

[Boo03] Charles Bookman. Linux clustering: building and maintaining Linux clusters. New Riders Publishing, Carmel, IN, USA, 2003. ISBN 1-57870-274-7. xv + 265 pp. LCCN QA76.76.O63 B427 2003.

Borison:1979:RGU

[Bor79] E. Borison. Report generation on UNIX. Technical Memorandum 1337, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1, 1979. ?? pp.

Borenstein:1988:UER

[Bor88]  N. S. Borenstein.  UNIX

emacs: a retrospective. lessons for flexible system design. In ACM [ACM88], pages 95–101. ISBN 0-89791-283-7. LCCN QA76.9.U83A26 1988.

Borr:1998:SSU

[Bor98] Andrea J. Borr. SecureShare:

Safe UNIX/Windows file sharing through multiprotocol locking. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

usenix.org/publications/
library/proceedings/usenix-
nt98/borr.html.

Born:2001:VWP

[Bor01] Günter Born. VMware Workstation Praxisführer: Installation, Konfiguration, Anwendung unter Windows und Linux. SuSE-PRESS, Nürnberg, Germany, 2001. ISBN 3-934678-81-5. xxii + 372 pp. LCCN ???? DM 79.00, EUR 40.00.

Borges:2014:OSA

[Bor14] Alexandre Borges. Oracle Solaris 11 Advanced Administration Cookbook. Packt Publishing, Birmingham, UK, 2014. ISBN 1-84968-827-3, 1-84968-826-5. 478 pp. LCCN ????

Bott:1984:OTS

[Bot84] Ross Bott. OSx: Towards a single UNIX system for superminis. In USENIX Association [USE84b], pages 145–167. ISBN none. LCCN QA76.8.U65 U55 1984.

Bourne:1978:UTS

[Bou78a] S. R. Bourne. UNIX timesharing system: The UNIX shell. The Bell System Technical Journal, 57(6):1971–1990, July/August 1978. CO-DEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (elec-

tronic). URL http://bstj.
bell-labs.com/BSTJ/images/
Vol57/bstj57-6-1971.pdf;
 http://www.alcatel-lucent.
com/bstj/vol57-1978/articles/
bstj57-6-1971.pdf;   https:
//archive.org/details/bstj57-
6-1971.

Bourne:1978:IUS

[Bou78b] Stephen R. Bourne. An introduction to the UNIX shell. Technical Memorandum 1239, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 12, 1978. ?? pp.

Bourne:1982:US

[Bou82] Stephen R. Bourne. The UNIX System, volume 6 of International computer science series. Addison-Wesley, Reading, MA, USA, 1982. ISBN 0-201-13791-7 (paperback). xiii + 351 pp. LCCN QA76.8.U65 B68 1982.

URL http://books.google.
com/books?id=5bFQAAAAMAAJ.

Bourne:1986:USV

[Bou86]  Stephen R. Bourne.  The

UNIX System V Environment. Addison-Wesley, UK, Wokingham, UK, 1986. ISBN 0-201-18484-2 (paperback). xiii + 378 pp. LCCN QA76.76.O63 B68 1987.

Bourne:1990:UVU

[Bou90a] Philip E. Bourne. UNIX for VMS Users. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 1-55558-034-3. xvi + 368 pp. LCCN QA76.76.O63 B67 1990.

Bourne:UVU90

[Bou90b] Philip E. Bourne. UNIX for VMS Users. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 1-55558-034-3. xvi + 368 pp. LCCN QA76.76.O63 B67 1989.

Bourne:1997:CSI

[Bou97] Philip E. Bourne. A cookbook for serving the Internet: UNIX version. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-519992-1. xxix + 303 pp. LCCN QA76.9.C55 B68 1997.

Bourne:1977:IUS

[Bou7 ] Stephen R. Bourne. An introduction to the UNIX shell. Computing Science Technical Report 70, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Boyd:1984:SIP

[Boy84] Stowe Boyd. SYSTANT: An integrated programming environment for modular C under UNIX. In Software Tools Users Group [Sof84], pages 141–150. LCCN QA76.8.U65 U83 1984.

Butterfield:1984:NTL

[BP84] David A. Butterfield and Gerald J. Popek. Network tasking in the locus distributed UNIX system. In Software Tools Users Group [Sof84], pages 62–71. LCCN QA76.8.U65 U83 1984.

Barak:1986:MSU

[BP86] Amnon Barak and On G. Paradise. MOS — scaling up UNIX. In USENIX Association [USE86c], pages 414–418.

Bershad:COMPSYS-1-2-169

[BP88a] Brian N. Bershad and C. Brian Pinkerton. Watchdogs — extending the UNIX file system. Computing Systems, 1(2):169– 188, Spring 1988. CODEN CM-SYE2. ISSN 0895-6340.

Bershad:1988:WEUa

[BP88b] Brian N. Bershad and C. Brian Pinkerton. Watchdogs — extending the UNIX file system. In USENIX Association [USE88g], pages 267–275. ISBN ???? LCCN ????

Bershad:1988:WEUb

[BP88c] Brian N. Bershad and C. Brian Pinkerton. Watchdogs — extending the UNIX file system.

In Association [Ass88a], pages 169–188.

Becker:1991:APB

[BP91] Jeffrey C. Becker and Arvin Park. Analysis of the paging behavior of UNIX. ACM SIG-METRICS Performance Evaluation Review, 19(2):36–43, August 1991. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Ball:2000:LU

[BP00] Bill Ball and David Pitts. Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, fourth edition, 2000. ISBN 0-672-31688-9. xl + 1468 pp. LCCN QA76.76.O63 B35 2000.

Buechler:2009:PDG

[BP09] Chrisopher M. Buechler and Jim Pingle. pfSense: the definitive guide: the definitive guide to the pfSense open source firewall and router distribution. Reed Media Services, Marysville, WA, USA, 2009. ISBN 0-9790342-8-0 (paperback). xxvi + 489 pp. LCCN QA76.76.S95 B84 2009.

Borghi:1986:SIP

[BQd86] Bruno Borghi, Stephane Querel, and Daniel deRauglaudre. Sm-Script: An interpretor for the PostScript language under UNIX. In USENIX Association [USE86c], pages 284–293.

Burnette:1977:PMS

[BR77] W. A. Burnette and J. R.

Remde. Programmer’s manual for the SEL-86 PDP-11/ 45 operating system. Technical Memorandum 1155 (TM 77-8231-2, TM 77-1228-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 7, 1977. ?? pp.

Banahan:1982:UB

[BR82] Michael F. Banahan and Andrew Rutter. UNIX: the Book. Sigma Technical Press, Wilmslow, Cheshire, UK, 1982. ISBN 0-905104-21-8. 265 pp. LCCN QA76.76.O63.

Barron:1987:TPT

[BR87] David Barron and Mike Rees. Text Processing and Typesetting with UNIX. International computer science series. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-14219-8. xiv + 447 pp. LCCN Z286.D47 B37 1987. US$24.95; UK£15.95.

Bokhari:1999:LDW

[BR99] Shahid H. Bokhari and Rafeequr Rehman. Linux and the developing world. IEEE Software, 16(1):58–64, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s1058abs.
htm;   http://dlib.computer.
org/so/books/so1999/pdf/
s1058.pdf.

Brandt:1976:ISM

[Bra76a] R. B. Brandt. Implementation of semaphores and messages in UNIX. Technical Memorandum 1136 (MF 76-8234-76), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 7, 1976. ?? pp.

Brandt:1976:PUI

[Bra76b] R. B. Brandt. Proposal for UNIX interprocess communication. Technical Memorandum 1090 (TM 76-8234-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 17, 1976. ?? pp.

Brandt:1977:UEL

[Bra77] R. B. Brandt. UNIX error logging. Technical Memorandum 1163 (TM 77-8234-10), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 6, 1977. ?? pp.

Brandt:1978:ELU

[Bra78] R. B. Brandt. Error logging in UNIX/TS and UNIX/ RT. Technical Memorandum 1284 (TM 78-8234), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 26, 1978. ?? pp.

Bradford:1979:AVM

[Bra79] E. G. Bradford. An augmented version of make. Technical Memorandum 1458, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1, 1979. ?? pp.

Brain:1992:MPE

[Bra92] Marshall Brain. Motif programming — The Essentials — and More. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1992. ISBN 1-55558-089-0. xviii + 601 pp. LCCN QA76.76.W56 B73 1992.

Brambert:1993:GI

[Bra93] Dave Brambert, ed. Guide to Internetworking: Interconnecting Diverse Networks: DOS, Macintosh, UNIX, Mainframe. Miller Freeman Publications, San Francisco, CA, USA, 1993. ISBN 0-87930-262-3. 186 pp. LCCN TK5105.7 .G85 1993. US$26.95.

Braams:1995:BCP

[Bra95] Jan Braams. Batch class process scheduler for Unix SVR4. ACM SIGMETRICS Performance Evaluation Review, 23 (1):301–302, May 1995. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Braun:1998:BMF

[Bra98] Jörg Braun. BSD mit Methode: [die freien Berkeley-Software-Distributionen]. C&L, Computer- und Literaturverlag, Vaterstetten, Germany, 1998. ISBN 3-932311-31-0. 856 (est.) pp. LCCN ????

Bramer:2004:DGL

[Bra04] Michael Bramer. Debian GNU LINUX 2.2: kompromisslos in Sicherheit & Stabilität: über 4000 Programmpakete: ausgereifte Desktop- & Server-Lösungen: mit Support & Handbuch. (German) [Debian GNU LINUX 2.2: Uncompromising in Security and Stability: Over 4000 program packages: Mature Desktop and Server Solutions: With Support and Handbook]. LinuxLand Internationale, Müenchen, Germany, 2004. ISBN 3-00-005785-4. ???? pp. LCCN ????

Boie:1977:ICC

[BRD77] R. A. Boie, M. Robbins, and R. Data. An interactive computer controlled data acquisition system with a CAMAC real time …. Technical Memorandum 1169 (TM 77-1131-17), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 12, 1977. ?? pp.

Brereton:1986:MRF

[Bre86] O. P. Brereton. Management of replicated files in a UNIX environments. Software—Practice and Experience, 16(8):771– 780, August 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

BrinchHansen:2001:COS

[Bri01] Per Brinch Hansen. Classic Operating Systems: from Batch Processing to Distributed Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 0-387-95113-X. x + 597 pp. LCCN QA76.76.O63 B7425

2001. URL http://www.loc.
gov/catdir/enhancements/
fy0816/00045036-d.html;
http://www.loc.gov/catdir/
enhancements/fy0816/00045036-

                                        t.html;   https://link.springer.
                                        com/book/10.1007/978-1-4757-
                                        3510-9.

Brickner:2005:LDP

[Bri05a] David Brickner. Linux desktop pocket guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10104-X. xvi + 181 pp. LCCN QA76.76.O63 B7423 2005.

Brickner:2005:TDL

[Bri05b] David Brickner. Test driving Linux: from Windows to Linux in 60 seconds. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00754-X. xvii + 341 pp. LCCN QA76.76.O63 B744 2005. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596007546.

Brown:1971:AUM

[Bro71] W. S. Brown. ALTRAN user’s manual. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1971. ?? pp.

Brown:1973:AUM

[Bro73] W. S. Brown. ALTRAN user’s manual. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973. ?? pp.

Brown:1974:CFR

[Bro74] W. S. Brown. On computing with factored rational expressions. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 8(3):26–34, August 1974. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
1086837.1086842.

Brown:1977:AUM

[Bro77a] W. S. Brown. ALTRAN user’s manual. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977. ?? pp.

Brown:1977:RMF

[Bro77b] W. S. Brown. A realistic model of floating-point computation. In Rice [Ric77], pages 343–360. ISBN 0-12-587260-7. LCCN QA3 .U45 no. 39; QA297 .M36 1977. URL https:

//www.sciencedirect.com/
science/article/abs/pii/
B9780125872607500170.

Brown:1978:SPA

[Bro78] W. S. Brown. The subresultant PRS algorithm. ACM Transactions on Mathematical Software, 4(3):237–249, September 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

//dl.acm.org/doi/10.1145/
355791.355795.

Brooks:1979:PHH

[Bro79] C. A. Brooks. PWB hardware help. Technical Memorandum 1423, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1979. ?? pp.

Brown:1981:SRMa

[Bro81a]  W. S. Brown.  A simple

but realistic model of floatingpoint computation. Computing Science Technical Report 83, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1981.

Brown:1981:SRMb

[Bro81b] W. S. Brown. A simple but realistic model of floating-point computation. ACM Transactions on Mathematical Software, 7(4):445–480, December 1981. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        355972.355975.

Brown:1989:HSU

[Bro89a] P. J. Brown. A hypertext system for UNIX. In Association [Ass89b], pages 37–53.

Brown:COMPSYS-2-1-37

[Bro89b] P. J. Brown. A hypertext system for UNIX. Computing Systems, 2(1):37–53, Winter 1989. CODEN CMSYE2. ISSN 0895-6340.

Browning:1990:CUM

[Bro90] S. A. Browning. Cip user’s manual: One picture is worth a thousand words. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

                                        1990. URL http://doc.cat-
                                        v.org/unix/v10/10thEdMan/
                                        v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Brown:1994:UDP

[Bro94] Chris Brown. UNIX distributed programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-075896-5. xix + 385 pp. LCCN QA76.76.O63 B745 1994.

Brown:1997:ODA

[Bro97]  Lynnwood Brown.  Oracle

database administration on UNIX systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-244666-9. xxi + 189 pp. LCCN QA76.9.D3 B7826 1997.

Brown:1998:BPU

[Bro98] Martin C. Brown. BeOS: porting UNIX applications. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1998. ISBN 1-55860-532-0 (paperback). xii + 484 pp. LCCN QA76.76.O63 B756 1998.

Brown:2006:SL

[Bro06] Chris Brown. SuSE Linux.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10183-X. xiii + 430 pp. LCCN QA76.76.O63 B76 2006. US$39.99. URL http:/

/www.oreilly.com/catalog/
9780596101831.

Brock:2019:EUC

[Bro19] David C. Brock. The earliest Unix code: An anniversary source code release. Computer History Museum Web site, October 17, 2019. URL

                                        http://bit.ly/31pWcvM;
                                        https://computerhistory.
                                        org/blog/the-earliest-unix-
                                        code-an-anniversary-source-
                                        code-release/.

Brock:2020:DDR

[Bro20] David C. Brock. Discovering Dennis Ritchie’s lost dissertation. Web site, June 19, 2020.

Brown:1973:CFR

[Bro3 ] W. S. Brown. On computing with factored rational expressions. Computing Science Technical Report 22, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973 (??). ?? pp.

Brown:1976:RMF

[Bro6 ] W. S. Brown. A realistic model of floating-point computation. Computing Science Technical Report 58, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1976 (??). ?? pp.

Brown:1977:SPA

[Bro7 ] W. S. Brown. The subresultant PRS algorithm. Computing Science Technical Report 56, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Brown:1980:SRM

[Bro0 ]  W. S. Brown.  A simple

but realistic model of floatingpoint computation. Computing Science Technical Report 83, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1980 (??). ?? pp.

Brunson:2002:LWI

[Bru02] Ross Brunson. Linux and Windows 2000 integration toolkit: a complete resource. Wiley, New York, NY, USA, 2002. ISBN 0-471-41746-7 (paperback). xii + 371 pp. LCCN QA76.6 .B7778 2002.

Braunstein:1989:IEU

[BRW89] A. Braunstein, M. Riley, and John Wilkes. Improving the efficiency of UNIX buffer caches. Operating Systems Review, 23 (5):71–82, December 1989. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Bryan:1983:VAM

[Bry83a] Scott Bryan. VCHK — a maintenance program for UNIX file hierarchies. In Software Tools Users Group [Sof83], pages 41– 44. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Bryan:1983:VMP

[Bry83b] Scott Bryan. VCHK — a maintenance program for UNIX file hierarchies. In Association [Ass83], pages 41–44.

Bach:1979:PAD

[BS79a] M. J. Bach and B. L. Smith. Porting the ADAPT Data Translation System to the IBM 370. Technical Memorandum 1373 (TM 79-9157-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 8, 1979. ?? pp.

Bonanni:1979:SCC

[BS79b] L. E. Bonanni and C. A. Salemi. Source code control system user’s guide. Technical Memorandum 1367 (TM 79-9441-4, TM 79-9442-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1, 1979. ?? pp.

Barak:1980:USP

[BS80] Amnon B. Barak and Amos Shapir. UNIX with satellite processors. Software—Practice and Experience, 10(5):383–392, May 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Brown:1984:SAA

[BS84] Marc H. Brown and Robert Sedgewick. A system for algorithm animation. In ????, editor, Proceedings of the 11th Annual Conference on Computer Graphics and Interactive Techniques — SIGGRAPH ’84. ACM Press, New York, NY, USA, 1984.

Brown:1985:TAA

[BS85] Marc H. Brown and Robert Sedgewick. Techniques for algorithm animation. IEEE Software, 2(1):28–39, January 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Barkley:1988:PSU

[BS88] Ronald E. Barkley and Curt F.

Schimmel. A performance study of the Unix System V fork system call using Casper. AT&T Technical Journal, 67(5):100–109, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Buhr:1990:SPL

[BS90] Peter A. Buhr and Richard A.

Stroobosscher. The μ-System: Providing light-weight concurrency on shared-memory multiprocessor computers running UNIX. Software—Practice and Experience, 20(9):929–963, September 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Becker:1991:SAG

[BS91] George Becker and Kathy Slattery. A Systems Administrator’s Guide to Sun Workstations. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97250-1 (New York), 3-540-97250-1 (Berlin). xiii + 288 pp.

LCCN QA76.525 .B43 1991. US$44.00.

Baker:1992:RBU

[BS92] Mary Baker and Mark Sullivan. The recovery box: Using fast recovery to provide high availability in the UNIX environment. In USENIX [USE92b], pages 31–44. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Brown:1994:GFF

[BS94] C. Wayne Brown and Barry J. Shepherd. Graphics File Formats. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 1-884777-00-

7. 488 pp. LCCN T385 .B777 1994. US$35.00.

Barrett:2001:SSS

[BS01] Daniel J. Barrett and Richard E. Silverman. SSH: The Secure Shell: The Definitive Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00011-1. xv + 540 pp. LCCN QA76.76.O63 B369 2001. US$39.95. URL http:/

/www.oreilly.com/catalog/
sshtdg/;   http://www.snailbook.
com/.

Bell:2002:MXV

[BS02] Mark R. Bell and Debrah D. Suggs. Mac OS X version 10.1 black book. Coriolis Group Books, Scottsdale, AZ, USA, 2002. ISBN 1-57610-606-3. xxxii + 654 pp.

LCCN QA76.76.O63 B44973 2002 Computer. CD-ROM contains a selection of shareware, freeware and demos for use with Mac OS X and later versions.

Borntrager:2005:PLS

[BS05] C. Bornträger and M. Schwidefsky. Providing Linux 2.6 support for the zSeries platform. IBM Systems Journal, 44(2): 331–340, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/borntraeger.html;
 http://www.research.ibm.
com/journal/sj/442/borntraeger.
pdf;   http://www.research.
ibm.com/journal/sj/442/borntraeger.
txt.

Barrett:2003:LSC

[BSB03] Daniel J. Barrett, Richard E.

Silverman, and Robert G. Byrnes. Linux security cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00391-9. xvii + 311 pp. LCCN TK5105.59 .B36 2003. URL http:/

/www.oreilly.com/catalog/
9780596003913.

Bullinger:1987:HII

[BSK87] H.-J. Bullinger, B. Shackel, and K. Kornwachs, editors. Human-computer Interaction

— INTERACT ’87: proceedings of the Second IFIP Conference on Human-Computer Interaction, held at the University of Stuttgart, Federal Republic of Germany, 1– 4 September, 1987. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987. ISBN 0-444-70304-7. LCCN QA76.9.S88I325 1987.

Badger:1995:DTE

[BSS^{+}95] Lee Badger, Daniel F. Sterne, David L. Sherman, Kenneth M. Walker, and Sheila A. Haghighat. A domain and type enforcement UNIX prototype. In USENIX Association [USE95c], pages 127–140. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-

                                        1995(5). URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/security95/
                                        badger.html.

Badger:COMPSYS-9-1-47

[BSSW96] Lee Badger, Daniel F. Sterne, David L. Sherman, and Kenneth M. Walker. A domain and type enforcement UNIX prototype. Computing Systems, 9(1): 47–83, Winter 1996. CODEN CMSYE2. ISSN 0895-6340.

Bauer:2005:LSS

[BsswLB05] Michael D. Bauer and Michael D. Building secure servers with Linux Bauer. Linux server security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2005. ISBN 0-596-00670-5. xvii + 522 pp. LCCN TK5105.59 .B38 2005; TK5105.59 .B38 2005eb; TK5105.59. URL http:/

/www.oreilly.com/catalog/
9780596006709.

Bradley:1978:EMS

[BSW78] M. H. Bradley, W. E. Sherbon, and J. S. Waldhuter. Evaluation of microprocessors supporting the C language: LSI-11, MAC-8, Z80. Technical Memorandum 1347 (TM 78-3323-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 5, 1978. ?? pp.

Blair:1998:SIU

[BT98] John D. Blair and The Samba Team. Samba: Integrating UNIX and Windows. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, June 1998. ISBN 1-57831-006-7. xviii + 298 pp. LCCN ???? US$29.95. URL http:

//www.clbooks.com/sqlnut/
SP/search/gtsumt?source=&
isbn=1578310067;   http:
//www.ssc.com/ssc/samba/.

From the publisher: “Includes CD-ROM containing version 1.9.18 of the Samba server, a library of useful tools and scripts, the Samba mailing list archives, and all examples discussed in the book.”.

Blaich:2009:RVM

[BTS09] Andrew Blaich, Douglas Thain, and Aaron Striegel. Reflections on the virtues of modularity: a case study in Linux security modules. Software—Practice and Experience, 39(15):1235– 1251, October ??, 2009. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bucki:2001:MX

[Buc01]  Lisa Bucki.  Mac OS X.

Fast and easy. Prima Publishing, Roseville, CA, USA, 2001. ISBN 0-7615-1984-X (paperback). xvi + 368 pp. LCCN QA76.8.M3 B83 2001 Computer.

Bucanek:2006:BX

[Buc06] James Bucanek. Beginning Xcode. Wrox beginning guides. Wrox/Wiley Publishers, Indianapolis, IN, USA, 2006. ISBN 0-471-75479-X (paper/ website). xxviii + 590 pp. LCCN QA76.76.O63 B826

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0623/2005037953-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0740/2005037953-
                                        b.html;   http://www.loc.
                                        gov/catdir/toc/ecip066/2005037953.
                                        html.

Bunt:1976:STO

[Bun76] Rick B. Bunt. Scheduling techniques for operating systems. Computer, 9(10):10–18, October 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bunker:1991:DNO

[Bun91] Ted Bunker. Digital nudges open the doors to VMS. Datamation, 37(2):83–??, January 15, 1991. CODEN DTMNAT. ISSN 0011-6963.

Burnette:1976:RTT

[Bur76] W. A. Burnette. A real-time time-shared operating system for an SEL/86–PDP-11 configuration. Technical Memorandum 1122 (TM 76-8231-9, TM 76-1228-11), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 29, 1976. ?? pp.

Burnette:1978:WMS

[Bur78] W. A. Burnette. Writing a MERT supervisor. Technical Memorandum 1258 (TM 78-9156-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 24, 1978. ?? pp.

Burnette:1979:DTB

[Bur79a] W. A. Burnette. A database transaction backout and crash recovery system. Technical Memorandum 1432 (TM 79-9457-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 16, 1979. ?? pp.

Buroff:1979:THU

[Bur79b] S. J. Buroff. Terminal handling for UNIX/370 using a front end processor. Technical Memorandum 1466 (TM 79-3628-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 18, 1979. ?? pp.

Burgard:1993:UEI

[Bur93] Michael Burgard. UNIX E-mail is Growing Up. UNIX/World, 10(3):111–??, March 1993. ISSN 0739-5922.

Burleson:2001:UOD

[Bur01] Donald K. Burleson. Unix for Oracle DBAs: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00066-9. v + 104 pp. LCCN QA76.9.D3 B8754 2001. URL http:/

/www.oreilly.com/catalog/
9780596000660.

Burtch:2004:LSS

[Bur04] Ken O. Burtch. Linux Shell Scripting with Bash. SAMS Publishing, Indianapolis, IN, USA, 2004. ISBN 0-672-32642-6. xvi + 412 pp. LCCN QA76.76.O63 B8725 2004. US$34.99.

Bushnell:2002:SCN

[Bus02] Rick Bushnell. Sun Certified Network Administrator for Solaris 8 Operating Environment. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-064669-5. xxviii + 465 pp. LCCN QA76.3.B874 2002.

URL http://www.sun.com/
books/catalog/bushnell.

Butenhof:1997:PPT

[But97] David R. Butenhof. Programming with POSIX threads. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63392-2. xviii + 381 pp. LCCN QA76.76.T55B88 1997. US$31.95. URL http://www.

amazon.com/exec/obidos/ASIN/

0201633922/ref=sim_books/
002-4892305-5599452.

But:2002:CSL

[But02a] Jason But. A C++ socket library for Linux. Dr. Dobb’s Journal of Software Tools, 27 (6):19–22, 24, June 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_06/
socketcc.txt;   http://www.
ddj.com/ftp/2002/2002_06/
socketcc.zip.

Butorac:2002:PIW

[But02b] Danko Butorac. Project IP-SIS — Web portal and Linux for the blind. Lecture Notes in Computer Science, 2398:585– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2398/23980585.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2398/23980585.
pdf.

Buyya:1999:HPC

[Buy99] Rajkumar Buyya, editor. High Performance Cluster Computing, Volume 1: Architecture and Systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-013784-7. 881 pp. LCCN QA76.88.H489

1999. URL http://www.dgs.
monash.edu.au/~rajkumar/
cluster/index.html;   http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130137847.html.

Bulterman:1992:MSU

[Bv92] D. C. A. Bulterman and R. van Liere. Multimedia synchronization and UNIX. Lecture Notes in Computer Science, 614:108– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breuer:2006:RNO

[BV06] Peter T. Breuer and Marisol García Valls. Raiding the Noosphere: the open development of networked RAID support for the Linux kernel. Software—Practice and Experience, 36(4):365– 395, ???? 2006. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Barton:1988:BTR

[BW88a] J. M. Barton and J. C. Wagner. Beyond threads: Resource sharing in UNIX. In USENIX Association [USE88g], pages 259–266. ISBN ???? LCCN ????

Barton:1988:ERS

[BW88b] J. M. Barton and J. C. Wagner. Enhanced resource sharing in UNIX. In Association [Ass88a], pages 111–133.

Barton:COMPSYS-1-2-111

[BW88c] J. M. Barton and J. C. Wagner. Enhanced resource sharing in UNIX. Computing Systems, 1(2):111–133, Spring 1988. CO-DEN CMSYE2. ISSN 0895-6340.

Barak:1989:MIM

[BW89] Amnon Barak and Richard Wheeler. MOSIX: An integrated multiprocessor UNIX. In USENIX Association [USE89f], pages 101–112.

Borr:1994:HAD

[BW94a] Andrea Borr and Carol Wilhelmy. Highly-available data services for UNIX client-server networks: Why fault tolerant hardware isn’t the answer. Lecture Notes in Computer Science, 774:285–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Borr:1994:HDS

[BW94b] Andrea Borr and Carol Wilhelmy. Highly-available data services for UNIX client-server networks: Why fault tolerant hardware isn’t the answer. Lecture Notes in Computer Science, 774:285–??, 1994. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burrows:1994:BSL

[BW94c] Michael Burrows and D. J.

Wheeler. A block-sorting lossless data compression algorithm. Research Report 124, Digital SRC, Palo Alto, CA, USA, May 10, 1994.

URL  ftp://ftp.digital.
com/pub/DEC/SRC/research-
reports/SRC-124.ps.gz. This

is the second of four key papers behind the bzip2 compression tools. The others are [HL90, Whe97, SB97].

Barrett:1996:CUT

[BW96] Martin L. Barrett and Clifford H. Wagner. C and UNIX: tools for software design. Wiley, New York, NY, USA, 1996. ISBN 0-471-30927-3. xvii + 446 pp. LCCN QA76.73.C15 B38 1996.

Barkakati:2002:RLW

[BW02] Naba Barkakati and Kurt Wall. Red Hat Linux 7.2 weekend crash course. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, 2002. ISBN 0-7645-3642-7. xix + 347 pp. LCCN QA76.76.O63 B366167 2002. URL http:

                                        //www.loc.gov/catdir/toc/
                                        wiley021/2001092935.html.

Barrett:2004:LKG

[BW04] Daniel J. Barrett and Torsten Wilhelm. Linux — kurz und gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 3-89721-501-2 (paperback). 204 pp. LCCN ???? EUR 9.90. URL http:

                                        //www.gbv.de/dms/ilmenau/
                                        toc/388679999.PDF.

Boyer:1996:APO

[BY96] Robert S. Boyer and Yuan Yu. Automated proofs of object code for a widely used microprocessor. Journal of the ACM, 43(1):166–192, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Buck:2010:CDP

[BY10] Erik M. Buck and Donald A.

Yacktman. Cocoa design patterns. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-321-53502-2 (paperback). xxv + 427 pp. LCCN QA76.64 .B82 2010.

Ben-Yehuda:2008:MGR

[BYVF08] Muli Ben-Yehuda, Eric Van Hensbergen, and Marc Fiuczynski. Minding the gap: R&D in the Linux kernel. Operating Systems Review, 42(5): 1–3, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cannon:1978:IHC

[C^{+}78a] L. W. Cannon et al. Indian Hill C style and coding standards. Technical Memorandum 1356 (TM 78-5512-1, TM 78-5514-1, TM 78-5522-3, TM 78-5326-1, TM 78-5521-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 29, 1978. ?? pp.

Coben:1978:RCW

[C^{+}78b] R. M. Coben [sic] et al. Recommendations for a companywide UNIX education program. Technical Memorandum 1291, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 6, 1978. ?? pp.

Chuvakin:2003:SLS

[C^{+}03] Anton Chuvakin et al. Securing Linux: a Survival Guide for Linux Security. The SANS Institute, Bethesda, MD, USA, 2003. ISBN ???? 104 pp. LCCN ???? US$39.00.

Costales:1997:S

[CA97] Bryan Costales and Eric Allman. sendmail. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-222-0. xxv + 1021 pp. LCCN HE6239.E54 C67 1997. US$39.95. URL http://www.

                                        ora.com/catalog/sendmail2/
                                        ;   http://www.oreilly.com/
                                        catalog/sendmail2.

Costales:2003:S

[CA03] Bryan Costales and Eric Allman. Sendmail. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 1-56592-839-3. xxiv + 1205 pp. LCCN TK5105.74.S44+. US$59.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        sendmail3.

Cabrera:1983:UPO

[Cab83] Luis Felipe Cabrera. A userprocess oriented performance study of Ethernet networking under Berkeley UNIX 4.2 BSD. Progress report 84/19, Computer Science Division (EELS), University of California, Berkeley, Berkeley, CA, USA, 1983. 37 pp.

Coppeto:1989:OLC

[CAG89a] Thomas J. Coppeto, Beth L.

Anderson, and Daniel E. Geer, Jr. OLC: An online consulting system for UNIX. In USENIX [USE89a], pages 83–94. FTP - aeneas.mit.edu:/pub/usenix/olc.PS; local - olc.ps.

Coppeto:1989:OOC

[CAG89b] Thomas J. Coppeto, Beth L.

Anderson, and Daniel E. Geer, Jr. OLC: An online consulting system for UNIX. In USENIX Association [USE89e], pages 83–94. LCCN QA 76.76 O63 U83 1989.

URL ftp://aeneas.mit.edu/
pub/usenix/olc.PS;   ftp:
//ftp.uu.net/networking/
athena/usenix/olc.PS.Z. FTP

- aeneas.mit.edu:/pub/usenix/olc.PS; local - olc.ps.

Calabrese:1993:ATC

[Cal93] Christopher J. Calabrese. An advanced tape cataloging system for UNIX systems. In USENIX Association [USE93f], pages 289–293. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

CS:2000:LSA

[Cal00a]  Caldera Systems.  Caldera

OpenLinux eDesktop 2.4. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, fourth edition, 2000. ISBN 0-13-088247-X. 450 pp. LCCN ???? US$59.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_013088247X.html.

Callaghan:2000:NI

[Cal00b] Brent Callaghan. NFS Illustrated. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-32570-5. xix + 513 pp. LCCN TK5105.574 .C35 2000.

Calkins:2003:IS

[Cal03] Bill Calkins. Inside Solaris 9. New Riders Publishing, Carmel, IN, USA, 2003. ISBN 0-7357-1101-1. xix + 910 pp. LCCN QA76.76.O63 C3515 2003.

Calkins:2006:SSA

[Cal06] Bill Calkins. Solaris 10 system administration. Exam prep. Que Corporation, Indianapolis, IN, USA, 2006. ISBN 0-7897-3461-3. xx + 1077 pp. LCCN QA76.76.O63 C3518 2006.

Calkins:2008:SSA

[Cal08] Bill Calkins. Solaris 10 system administration: CX-310-200, Part 1. Exam prep. Que Corporation, Indianapolis, IN, USA, 2008. ISBN 0-7897-3790-6 (paperback). xix + 738 pp. LCCN X08.E03772.

Calkins:2009:SSA

[Cal09] Bill Calkins. Solaris 10 system administration: (Exam CX-310-200). Que Corporation, Indianapolis, IN, USA, 2009. ISBN 0-7897-3790-6 (part 1: paperback), 0-7897-3817-1 (part 2: paperback). xix + 738 pp. LCCN JL-21-29.

Calcote:2010:APG

[Cal10] John Calcote. Autotools: a practitioner’s guide to GNU Autoconf, Automake, and Libtool. No Starch Press, San Francisco, CA, USA, 2010. ISBN 1-59327-206-5 (paperback). xxiv + 332 pp. LCCN QA76.76.D47 C335 2010.

Campbell:1987:CPG

[Cam87] Joe Campbell. C Programmer’s Guide to Serial Communications. Howard W. Sams, Indianapolis, IN 46268, USA, 1987. ISBN 0-672-22584-0. xii + 655 pp. LCCN QA76.73.C15 C36 1987. US$22.95.

Campbell:1991:EBU

[Cam91] Art Campbell. E-Mail Beyond UNIX. UNIX/World, 8(11):77– ??, November 1991. ISSN 0739-5922.

Cameron:1998:GEP

[Cam98] Debra Cameron. GNU Emacs Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-496-7. iii + 58 pp. LCCN QA76.76.T49 C348 1998. US$6.95. URL http:/

/www.oreilly.com/catalog/
gnupr.

Cameron:2004:MLS

[Cam04] Jamie Cameron. Managing Linux Systems with Webmin: System Administration and Module Development. Bruce Perens’ Open source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-140882-8. xvi + 792 pp. LCCN QA76.76.O63 2004. US$44.99.

Cannon:1979:EUT

[Can79] L. W. Cannon. Enhancements to the USG Trouble Reporting System. Technical Memorandum 1338, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 19, 1979. ?? pp.

Cannata:1980:USP

[Can80] P. E. Cannata. UNIX specification — process related system calls. Technical Memorandum 1495, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 28, 1980. ?? pp.

Caplinger:1982:PUE

[Cap82] Michael Caplinger. [Phi]NIX: a UNIX emulator for VAX/VMS. In Usr Group [Usr82], pages 249–255.

Caplinger:1987:UEV

[Cap87] Michael Caplinger. ϕnix: a Unix emulator for VAX/VMS. Technical report TR82-8, Rice University, Houston, TX, USA, August 10, 1987. URL http:

                                        //www.caam.rice.edu/caam/
                                        trs/82/TR82-08.pdf.

Carr:1977:CSM

[Car77] R. G. Carr. Character string manipulation in the C language. Technical Memorandum 1208, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1977. ?? pp.

Carter:1982:PEH

[Car82] Joel R. Carter. Perkin-Elmer’s hardware/I-O system: Flexibility that matches UNIX. In USENIX [USE82], pages 108– ?? Abstract only.

Cardelli:1983:FAM

[Car83a] Luca Cardelli. The functional abstract machine. Technical Report TR-107, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 4, 1983. 45 pp.

URL http://lucacardelli.
name/Papers/FAM.pdf.

Cargill:1983:BDa

[Car83b]  T. A. Cargill.  The Blit

debugger (preliminary draft). ACM SIGPLAN Notices, 18(8): 190–200, August 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cargill:1983:BDb

[Car83c] Thomas A. Cargill. Blit debugger. The Journal of Systems and Software, 3(4):277– 284, December 1983. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Cardelli:1984:BPT

[Car84] Luca Cardelli. Basic polymorphic typechecking. Computing Science Technical Report 112, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? 1984. 17 pp. URL

http://lucacardelli.name/
Papers/BasicTypechecking%
20(TR%201984).pdf.

Cargill:1985:IBD

[Car85] Thomas A. Cargill. Implementation of the Blit debugger. Software—Practice and Experience, 15(2):153–168, February 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cardelli:1986:AM

[Car86a] Luca Cardelli. The Amber machine. Technical Report 119, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1986. 28 pp.

URL http://lucacardelli.
name/Papers/AmberMachine.
A4.pdf.

Carrington:1986:PUE

[Car86b] David A. Carrington. Profiling under ELXSI UNIX. Software —Practice and Experience, 16 (9):865–873, September 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Carlin:1988:USA

[Car88a] Jerry M. Carlin. UNIX security at Pacific Bell. In USENIX Association [USE88e], pages 86– 87. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Carlin:1988:USP

[Car88b] Jerry M. Carlin. UNIX security at Pacific Bell. In USENIX [USE88b], pages 86– 87. LCCN QA76.8.U65 U55 1988(1)-1990(2)//. Abstract only.

Carr:1989:SSG

[Car89a] J. A. Carr. SCRI’s semicomplete guide to UNIX on the ETA10 supercomputer. Technical Report FSU-SCRI-89-28, Florida State University, Tallahassee, FL, USA, February 13, 1989. 64 pp.

Carson:SWS89

[Car89b] Mark Carson, et.al. Secure window systems for UNIX. In Proceedings of the USENIX Winter 1989 Conference. USENIX Association, Berkeley, CA, USA, January 1989. An architecture for a CMW based on Trusted XENIX and a text-based windowing system. Also mentions some X related issues.

Cargill:1990:FP

[Car90] T. A. Cargill. The feel of pi.

In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Carlin:1993:USU

[Car93a] Jerry M. Carlin. UNIX security update. In USENIX Association [USE93g], pages 119–130. ISBN 1-880446-55-3. LCCN QA 76.9 A25

U54 1993. URL http://www.
usenix.org/publications/
library/proceedings/sec4/.

Costales:1993:TS

[CAR93b] Bryan Costales, Eric Allman, and Neil Rickert. sendmail. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-056-2. xxxvi + 792 pp. US$32.95. From Samuel Ko: Highly recommended. This is really an encyclopedia on sendmail. The authors did a great job in de-mysterifying this complex email routing and delivering program …. The book is divided into four parts. Part One is a tutorial on understanding sendmail from the ground up …. Part Two deals with practical topics in sendmail administration …. Part Three is a comprehensive reference section …. Part Four consists of appendices with details ….

Carasik:1999:USS

[Car99] Anne H. Carasik. Unix Secure Shell. McGraw-Hill tools series. McGraw-Hill, New York, NY, USA, 1999. ISBN 0-07-134933-2 (paperback). xxv + 339 pp. LCCN QA76.76.O63 C37294 1999.

Carter:2004:OIS

[Car04] Brian Carter. OpenBSD: Implementing the Secure UNIX Platform. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, 2004. ISBN 0-7645-3933-7. 360 (est.) pp. LCCN ???? US$40.00.

Cashin:1993:NUC

[Cas93] Jerry Cashin. Networking with UNIX — Connecting Open Systems. Computer Technology Research Corp., 6 N. Atlantic Wharf, Charleston, SC 29401-2150, USA, 1993. ISBN 1-56607-003-1. iv + 171 pp. LCCN TK5105.5 .C37 1992. US$220.00.

Cashin:1994:BFP

[Cas94] Jerry Cashin. Bloom fading from Posix rose as open focus shifts. Software magazine, 14 (3):87–??, March 1994. CO-DEN SMWMEQ. ISSN 0897-8085.

Cato:1979:DIS

[Cat79] H. E. Cato. The DIRECT II scheduler. Technical Memorandum 1462 (TM 79-9456-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 30, 1979. ?? pp.

Catanzaro:1991:STP

[Cat91] Ben J. Catanzaro, editor.

The SPARC Technical Papers.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97634-5 (New York), 3-540-97634-5 (Berlin). xvi + 501 pp. LCCN QA76.9.A73 S65 1991. US$43.00.

Chen:1983:EPU

[CB83] Paul Chen and Chet Britten. Experiences in porting 4.1BSD UNIX to the [lambda]750 VLSI development system. In Association [Ass83], pages 132–?? Abstract only.

Coggins:1989:MCL

[CB89] James M. Coggins and Gregory Bollella. Managing C++ libraries. ACM SIGPLAN Notices, 24(6):37–48, June 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheswick:1994:FIS

[CB94] William R. Cheswick and Steven M. Bellovin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63357-4. xiv + 306 pp. LCCN TK5105.875.I57C44 1994. US$24.95.

Campbell:1991:PUS

[CBB^{+}91] Mark Campbell, Richard Barton, Jim Browning, Dennis Cervenka, Ben Curry, Tod Davis, Tracy Edmonds, Russ Holt, John Slice, Tucker Smith, and Rich Wescott. The parallelization of UNIX System V Release 4.0. In USENIX [USE91d], pages 307– 324. LCCN QA 76.76 O63 U84 1992.

Chou:1992:AN

[CBH^{+}92] J. H. Chou, C. R. Buckman, T. Hemp, A. Himwich, and F. Niemi. AIX NetView/ 6000. IBM Systems Journal, 31(2):270–285, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5473.

Cripe:1996:CDE

[CBL96] Brian E. Cripe, Jon A. Brewster, and Dana E. Laursen. A common desktop environment for platforms based on the UNIX operating system. Hewlett–Packard Journal, 47(2):6–13, April 1996. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/hpcc997.external.hp.com:
80/hpj/apr96/apr96.htm;
http://www.hp.com/hpj/apr96/
ap96a1.pdf.

Cheswick:2003:FIS

[CBR03] William R. Cheswick, Steven M.

Bellovin, and Aviel D. Rubin. Firewalls and Internet Security: Repelling the Wily Hacker. Addison-Wes-ley, Reading, MA, USA, second edition, 2003. ISBN 0-201-63466-X. xiv + 433 pp. LCCN TK5105.875.I57C44 2003. US$49.99, CAN$77.99.

Costilla:1993:SDI

[CBV93] C. R. Costilla, M. J. Bas, and J. Villamor. SIRIO: a distributed information system over a heterogeneous computer network. SIGMOD Record (ACM Special Interest Group on Management of Data), 22 (1):28–33, March 1993. CO-DEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Cohen:2007:MXB

[CC07] Michael E. Cohen and Dennis R. Cohen. The Mac Xcode 3 book. Wiley, New York, NY, USA, 2007. ISBN 0-470-05339-9 (paperback). 352 (est.) pp. LCCN ????

                                        URL http://www.loc.gov/
                                        catdir/enhancements/fy0741/
                                        2007926015-d.html.

CCA:1984:CEMb

[CCA84] CCA Uniworks, Cambridge, MA, USA. CCA EMACS: manual for UNIX users, version 162.45z edition, October 1984. various pp.

Clayton:2000:STY

[CCB00] Nik Clayton, Chris Coleman, and Sue Blake. Sams teach yourself FreeBSD in 21 days. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31854-7. ???? pp. LCCN ???? Edited by Angie Wethington, Neil Rowe, and Tony Amico.

Clancy:1987:UV

[CCD^{+}87] Patrick Clancy, Benjamin F.

Cutler, J. Christopher Dodd, Douglas W. Gilmore, Robert P. Nix, John J. O’Donnell, and Christopher P. Ryland. UNIX on a VLIW. In USENIX Association [USE87d], pages 225– 241.

Cattaneo:2001:DIT

[CCDP01] Giuseppe Cattaneo, Luigi Catuogno, Aniello Del Sorbo, and Pino Persiano. The design and implementation of a transparent cryptographic file system for UNIX. In USENIX [USE01b], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001. URL

http://www.usenix.org/publications/
library/proceedings/usenix01/
freenix01/cattaneo.html.

Clay:1989:UEH

[CCF89] Larry Clay, George Copeland, and Mike Franklin. UNIX extensions for high-performance transaction processing. In USENIX Association [USE89g], pages 73–80. ISBN ???? LCCN ????

Copp:1977:ISM

[CCH^{+}77] D. H. Copp, R. J. Council, D. J.

Hunsberger, J. J. Molinelli, and J. H. VanOrnum. An integrated system for microprocessor support. Technical Memorandum 1153 (TM 77-8234-1, TM 77-3233-6, TM 77-3231-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 17, 1977. ?? pp.

Carson:1990:NUS

[CCM^{+}90] Mark Carson, Janet Cugini, Sohail Malik, Mythili Kannan, and Wen-Der Jiang. Networked UNIX without the superuser. In USENIX Association [USE90e], pages 137–140. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Cruz-Carlon:2023:PLB

[CCVLW23] Juan Cruz-Carlon, Mahsa Varshosaz, Claire Le Goues, and Andrzej Wasowski. Patching locking bugs statically with crayons. ACM Transactions on Software Engineering and Methodology, 32(3):56:1–56:??, July 2023. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3548684.

Comer:1985:TTU

[CD85] Douglas Comer and Ralph E. Droms. Tilde trees in the UNIX environment. In USENIX Association [USE85c], pages 23–29.

Carling:2000:LSA

[CDD00] M. Carling, Stephen Degler, and James T. Dennis. Linux System Administration. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-562-05934-3. x + 337 pp. LCCN QA76.76.O63 C3745 2000. US$29.99.

Chanet:2005:SWC

[CDD^{+}05] Dominique Chanet, Bjorn De Sutter, Bruno De Bus, Ludo Van Put, and Koen De Bosschere. System-wide compaction and specialization of the Linux kernel. ACM SIGPLAN Notices, 40(7):95–104, July 2005. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chanet:2007:ARM

[CDD^{+}07] Dominique Chanet, Bjorn De Sutter, Bruno De Bus, Ludo Van Put, and Koen De Bosschere. Automated reduction of the memory footprint of the Linux kernel. ACM Transactions on Embedded Computing Systems, 6(4):23:1–23:??, September 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Chandra:2001:PPO

[CDK^{+}01] Rohit Chandra, Leonardo Dagum, David Kohr, Dror Maydan, Jeff McDonald, and Ramesh Menon. Parallel Programming in OpenMP. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2001. ISBN 1-55860-671-8. xvi + 230 pp. LCCN QA76.642 .P38 2001. US$39.95. URL

http://www.mkp.com/books_
catalog/catalog.asp?ISBN=
1-55860-671-8.

Card:1998:LKB

[CDM98] Remy Card, Eric Dumas, and Franck Mevel. The Linux Kernel book. Wiley, New York, NY, USA, 1998. ISBN 0-471-98141-9. xxx + 518 pp. LCCN QA76.76.O63 C374 1998. US$44.99. Translated by Chris Skrimshire.

Chandra:1994:SPM

[CDV^{+}94] Rohit Chandra, Scott Devine, Ben Verghese, Anoop Gupta, and Mendel Rosenblum. Scheduling and page migration for multiprocessor compute servers. ACM SIGPLAN Notices, 29 (11):12–24, November 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
asplos/195473/p12-chandra/

.

Canas:1988:PUO

[CE88] Daniel A. Cañas and Laura M. Esquivel. Portability and the UNIX operating system. Operating Systems Review, 22(2): 6–23, April 1988. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Chee:1997:IVC

[CE97] Chye-Lin L. Chee and Sevki S. Erdogan. An installable Version Control File System for UNIX. Software—Practice and Experience, 27(6):725–746, June 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
7313;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=7313&PLACEBO=IE.pdf.

Chen:1996:MPP

[CEC^{+}96] J. Bradley Chen, Yasuhiro Endo, Kee Chan, David Mazières, Antonio Dias, Margo Seltzer, and Michael D. Smith. The measured performance of personal computer operating systems. ACM Transactions on Computer Systems, 14(1): 3–40, February 1996. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1996-14-1/p3-chen/.

Cerofolini:1982:USB

[Cer82] Luigi Cerofolini. UNIX for the STD bus. EUUG Newsletter, 2(4):49–51, Winter 1982. CO-DEN EONLE8. ISSN 1011-4211.

Cerofolini:1983:USB

[Cer83] Luigi Cerofolini. UNIX for the STD bus. In Association [Ass83], pages 185–?? Abstract only.

Cervone:1996:ASG

[Cer96] Frank Cervone. AIX/6000 system guide. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-024129-5. xv + 501 pp. LCCN QA76.76.O63 C4 1996.

Ceruzzi:1998:HMC

[Cer98] Paul E. Ceruzzi. A History of Modern Computing. History of computing. MIT Press, Cambridge, MA, USA, 1998. ISBN 0-262-03255-4. x + 398 pp. LCCN QA76.17 .C47 1998.

Ceruzzi:2003:HMC

[Cer03] Paul E. Ceruzzi. A History of Modern Computing. History of computing. MIT Press, Cambridge, MA, USA, second edition, 2003. ISBN 0-262-53203-4 (paperback). xi + 445 pp. LCCN QA76.17.C47 2003.

Cevoli:2003:EFC

[Cev03]  Paul Cevoli.  Embedded

FreeBSD cookbook. Newnes, Oxford, UK, 2003. ISBN 1-58995-004-6 (paperback). x + 233 pp. LCCN QA76.76.O63 C482 2003.

Calavera:2020:LOB

[CF20] David Calavera and Lorenzo Fontana. Linux Observability with BPF: Advanced Programming for Performance Analysis and Networking. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2020. ISBN 1-4920-5020-2. xiii + 162 pp. LCCN QA76.774.L46. URL http://

proquest.safaribooksonline.
com/?fpi=9781492050193. Fore-

word by Jessie Frazelle.

Cole:1985:IEF

[CFA85] Clement T. Cole, Perry B. Flinn, and Alan B. Atlas. An implementation of an extended file system for UNIX. In USENIX Association [USE85d], pages 131–149. LCCN QA76.8.U65 U8 1985.

Ciancarini:1993:LMM

[CG93a] P. Ciancarini and N. Guerrini.

Linda meets Minix. Operating Systems Review, 27(4):76– 92, October 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Corrigan:1993:OPT

[CG93b] Peter Corrigan and Mark Gurry. ORACLE Performance Tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 1993. ISBN 1-56592-048-1. 650 pp. LCCN QA76.9.D3 C67 1993. US$34.95.

Chastek:1995:THC

[CG95a] Gary Chastek and Anthony Gargaro. Towards heterogeneous communication for POSIX real-time distributed systems. Ada User Journal, 16 (3):163–176, September 1995. CODEN AUJOET. ISSN 0268-652X.

Chestek:1995:THC

[CG95b] Chestek and Gargaro. Towards heterogeneous communication for Posix real-time distributed systems. In Anonymous [Ano95a], pages 163–175. ISSN 0268-652x.

Corden:1989:IAD

[CGMS89] M. J. Corden, C. H. Georgiopoulos, M. E. Mermikides, and J. Streets. Implementation of the ALEPH detector simulation code using UNIX with online graphics display. Computer Physics Communications, 57 (1–3):260–262, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0010465589902245.

Cutler:1992:XWS

[CGO92] Ellie Cutler, Daniel Gilly, and Tim O’Reilly, editors. The X Window System in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1992. ISBN 1-56592-017-1. 424 pp. LCCN QA76.76.W56 X2 1992. US$9.95.

Cardoza:1996:DTM

[CGS96] Wayne M. Cardoza, Frederick S. Glover, and William E. Snaman, Jr. Design of the TruCluster multicomputer system for the Digital UNIX environment. Digital Technical Journal of Digital Equipment Corporation, 8(1):5–17, May 1996. CODEN DTJOEL. ISSN 0898-901X. URL ftp://ftp.

digital.com/pub/Digital/
info/DTJ/v8n1/Design_of_
the_TruCluster_Multi_04jun1996DTJL01P8.
ps;   http://www.digital.
com:80/info/DTJL01/DTJL01.
HTM;   http://www.digital.
com:80/info/DTJL01/DTJL01AH.
HTM;   http://www.digital.
com:80/info/DTJL01/DTJL01P8.

PS;   http://www.digital.
com:80/info/DTJL01/DTJL01PF.
PDF;   http://www.digital.
com:80/info/DTJL01/DTJL01SC.
TXT.

Cykiert:1977:ILS

[CH77] H. M. Cykiert and B. E. Hess.

An Interactive LEAP Script Generator (ILSG). Technical Memorandum 1223 (TM 77-9423-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 2, 1977. ?? pp.

Cordy:1983:TAN

[CH83a] J. R. Cordy and R. C. Holt.

Turing: a new general purpose computer language under UNIX. In Software Tools Users Group [Sof83], pages 249–254. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Cordy:1983:TNG

[CH83b] J. R. Cordy and R. C. Holt.

Turing: a new general purpose computer language under UNIX. In Association [Ass83], pages 249–254.

Colet:1996:WWI

[CH96] Marc Colet and Robert Herzog. WWW2GCG, a Web interface to the GCG biological sequences analysis software. Computers and Graphics, 20 (3):445–450, May–June 1996. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cgi-bin/cas/
                                        tree/store/cag/cas_sub/browse/
                                        browse.cgi?year=1996&volume=
                                        20&issue=3&aid=9600014.

Chambers:1969:SDL

[Cha69] J. M. Chambers. A statistical data language. In Milton and Nelder [MN69], pages 179–199. ISBN 0-12-498150-X. LCCN QA276.4 .C6 1969.

Chang:1979:CSS

[Cha79] S.-J. Chang. A capacity study for the Service Evaluation Center — a UNIX-based realtime system. Technical Memorandum 1362 (TM 79-5223-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 9, 1979. ?? pp.

Chandler:1985:MRB

[Cha85] David Chandler. The monthly report: BSD’s big adventure: the bad Berkeley-to-Boston connection. UNIX Review, 3(1):8–15, January 1985. CODEN UNRED5. ISSN 0742-3136.

Chang:1996:UOS

[Cha96] Yu-Ming Chang. UNIX-based operating systems robustness evaluation. Washington, DC, USA, 1996. ???? pp. Shipping list no. 98-0846-M.

Chan:1997:USP

[Cha97] Terrence Chan. UNIX system programming using C++. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-331562-2. xv + 598 pp. LCCN QA76.73.C153 C46 1997.

Chatterjee:2006:BML

[Cha06] Soumitra Chatterjee. 64-bit migration to Linux on Itanium: Challenges, advantages, and tools. In Anonymous [Ano06a], page ?? ISBN ???? LCCN

????  URL http://www.
ice.gelato.org/;   http:/
/www.ice.gelato.org/about/
oct06_presentations.php.

Chacon:2009:PG

[Cha09] Scott Chacon. Pro Git. The expert’s voice in software development. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1833-9 (paperback). xxi + 265 pp. LCCN QA76.9.D5 C42 2009.

Chao:2011:OSM

[Cha11] Lee Chao, editor. Open source mobile learning: mobile Linux applications. Information Science Reference, Hershey, PA, USA, 2011. ISBN 1-60960-613-2. ???? pp. LCCN LB1044.87 .O64 2011.

Chen:1975:SUA

[Che75a] Stephen Chen. On the Sethi– Ullman algorithm. International Journal of Computer Mathematics, 5(1–4):37–55, January 1975. CODEN IJC-MAT. ISSN 0020-7160.

Chesson:1975:NUS

[Che75b] Gregory L. Chesson. The network Unix system. Operating Systems Review, 9(5):60– 66, November 1975. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Cherry:1976:ICL

[Che76a] Lorinda L. Cherry. An interactive control language for SIM. Technical Memorandum 1105 (TM 76-1271-8), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 28, 1976. ?? pp.

Cherry:1976:SLS

[Che76b] Lorinda L. Cherry. SIM — a language for simulating computers. Technical Memorandum 1097 (TM 76-1271-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 5, 1976. ?? pp.

Chellis:1978:PSU

[Che78] A. L. Chellis. Proposed structure for UNIX/TS and UNIX/ RT makefiles. Technical Memorandum 1285 (MF 78-8234-73), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 26, 1978. ?? pp.

Che:1979:SOU

[Che79a] H. Che. A send-only UUCP facility. Technical Memorandum 1492, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1979. ?? pp.

Che:1979:UAF

[Che79b] H. Che. UUCP administrators’ file and the UUNAME command. Technical Memorandum 1490, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1979. ?? pp.

Che:1979:UUS

[Che79c] H. Che. Uustat — a UUCP status inquiry program. Technical Memorandum 1491, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1979. ?? pp.

Cherry:1979:URC

[Che79d] Lorinda Cherry. UNIX reference card. Technical Memorandum 1021, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1979. ?? pp.

Cherlin:1981:UOS

[Che81] Edward Cherlin. The Unix operating system: portability A plus. Mini-Micro Systems, 14(4):153–154, 156, 159, April 1981. CODEN XWJXEH. ISSN 1000-1220.

Cheng:1987:VAC

[Che87a] Ray Cheng. Virtual address cache in UNIX. In USENIX Association [USE87d], pages 217– 224.

Cheriton:1987:UUS

[Che87b]  David R. Cheriton.  UIO:

a Uniform I/O system interface for distributed systems. ACM Transactions on Computer Systems, 5(1):12– 46, February 1987. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/

tocs/1987-5-1/p12-cheriton/

.

Chen:1990:CTS

[Che90a] Qi Chen. CTRAN: transforming scientific FORTRAN programs to UNIX based computing environments. Thesis (M.S.), Western Michigan University, Kalamazoo, MI, USA, 1990. vi + 124 pp.

Cherry:1990:I

[Che90b] L. L. Cherry. Index. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Cheswick:1992:PIT

[Che92] Bill Cheswick. A permuted index for TEX and L^{A}TEX. Computing Science Technical Report 145, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 13, 1992. ii + 111 pp.

Cheek:1994:ITB

[Che94] M. Cheek. Industry trends: 32bit computing: road to confusion? Computer, 27(7):13–14, July 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheek:1996:ESD

[Che96] Matthew Cheek. Enhanced security on Digital UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 31, 32, 35, 36, 38, 40, 42, 44, 46, November 1996. CODEN SYADE7. ISSN 1061-2688.

Cherlin:1997:BLB

[Che97] Edward Cherlin. Breaking the language barrier: Java, with no legacy code, is pushing the pace of adoption of Unicode, the only global character set. UNIX Developer, 1(1):56–59, January/February 1997. ISSN 1090-2279.

Cheek:1999:DUS

[Che99]  Matthew Cheek.  Digital

UNIX system administrator’s guide. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1999. ISBN 1-55558-199-4. x + 402 pp. LCCN QA76.76.O63C4573 1999. US$39.95.

Cheek:2002:TUS

[Che02] Matthew Cheek. Tru64 UNIX system administrator’s guide. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-255-9 (paperback). xiii + 470 pp. LCCN QA76.76.O63 T77 2002.

Chirillo:2002:HAD

[Chi02] John Chirillo. Hack attacks denied: a complete guide to network lockdown for UNIX, Windows, and Linux. Wiley, New York, NY, USA, second edition, 2002. ISBN 0-471-23283-1. xiii + 689 pp. LCCN QA76.9.A25 C47 2002.

Chisnall:2010:CPD

[Chi10] David Chisnall. Cocoa programming developer’s handbook. Developer’s library. Addison-Wes-ley, Reading, MA, USA, 2010. ISBN 0-321-63963-4 (paperback). xxviii + 896 pp. LCCN QA76.64 .C485 2010.

Chalup:2016:PSN

[CHL16] Strata R. Chalup, Christina J. Hogan, and Tom Limoncelli. The Practice of System and Network Administration: Volume 1: DevOps and Other Best Practices for Enterprise IT. Addison-Wesley, Reading, MA, USA, third edition, 2016. ISBN 0-13-341507-4 (e-book), 0-13-341508-2, 0-13-341510-4 (e-book), 0-321-91916-5 (paperback). li + 1168 pp. LCCN TK5105.5 .L53 2017.

Cho:1998:FAQ

[Cho98]  Kenjiro Cho.  A frame-

work for alternate queueing: Towards traffic management by PC-UNIX based routers. In USENIX [USE98e], page ?? ISBN 1-880446-94-4. LCCN QA76.76.O63 U88 1998.

URL  http://www.usenix.
org/publications/library/
proceedings/usenix98/cho.
html.

Christian:1983:UOS

[Chr83a] Kaare Christian. The UNIX Operating System. Wiley-Inter-science, New York, NY, USA, 1983. ISBN 0-471-87542-2 (hardcover) and 0-471-89052-9 (paperback). xviii + 318 pp. LCCN QA76.8.U65 C45 1983.

Christian:UOS83

[Chr83b] Kaare Christian. The UNIX Operating System. Wiley-Inter-science, New York, NY, USA, 1983. ISBN 0-471-87542-2 (hardcover) and 0-471-89052-9 (paperback). xviii + 318 pp. LCCN QA76.8.U65 C45 1983.

Christian:1987:UTP

[Chr87] Kaare Christian. The UNIX text processing system. Wiley, New York, NY, USA, 1987. ISBN 0-471-85581-2. xii + 250 pp. LCCN QA76.76.O63 C47 1987. US$24.95.

Christian:1988:UOS

[Chr88] Kaare Christian. The UNIX Operating System. Wiley-Inter-science, New York, NY, USA, second edition, 1988. ISBN 0-471-84782-8 (hardcover), 0-471-84781-X (paperback). xxii + 455 pp. LCCN QA76.8.U65 C45 1988. See [Sob89].

Christoph:1990:SCG

[Chr90] G. Christoph. Security considerations of going to a UNIX based supercomputer operating system. In USENIX Association [USE90e], pages 129–130.

ISBN ???? LCCN QA 76.9 A25 U55 1990.

Christenson:1998:BRCb

[Chr98]  Nick Christenson.  Book

review: Configuration and Capacity Planning for Solaris Servers. ;login: the USENIX Association newsletter, 23(4):??, June 1998. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1998-6/wong.html.

Chapin:1995:MSP

[CHRG95] John Chapin, A. Herrod, Mendel Rosenblum, and Anoop Gupta. Memory system performance of UNIX on CC-NUMA multiprocessors. ACM SIG-METRICS Performance Evaluation Review, 23(1):1–13, May 1995. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Chukran:1998:AAP

[Chu98] Rudy Chukran. Accelerating AIX: Performance Tuning for Programmers and System Administrators. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-63382-5. xix + 217 pp. LCCN QA76.76.O63C495 1998. US$49.95.

Chuvakin:2003:UDU

[Chu03] Anton Chuvakin. Ups and downs of UNIX/Linux hostbased security solutions. ;login: the USENIX Association newsletter, 28(2):??, April 2003. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic). URL http:/

/www.usenix.org/publications/
login/2003-04/pdfs/chuvakin.
pdf.

Coq:1992:SUF

[CJ92] Philippe Coq and Sylvie Jean.

SysView: a user-friendly environment for administration of distributed UNIX systems. In USENIX [USE92a], pages 145–151. ISBN 1-880446-47-2. LCCN ????

Covington:1994:PFU

[CJ94] Michael A. Covington and Mark Juric. A print filter for UNIX: More power for your LaserJet 4M printer. Dr. Dobb’s Journal of Software Tools, 19(10):64–65, 67, 95– 96, September 1994. CODEN DDJOEB. ISSN 1044-789X.

Collura:1992:ESL

[CJBT92] Thomas F. Collura, Ernest C.

Jacobs, Richard C. Burgess, and John P. Turnbull. The Epilog system — automated long-term EEG monitoring for epilepsy. Computer, 25(9):5–14, September 1992. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carson:1989:SWS

[CJL^{+}89] Mark E. Carson, Wen-Der Jiang, Jeremy G. Liang, Gary L. Luckenbaugh, and Debra H. Yakov. Secure window systems for UNIX. In USENIX Association [USE89f], pages 441–455. An architecture for a CMW based on Trusted XENIX and a text-based windowing system. Also mentions some X related issues.

Cohen:1978:UTS

[CK78a] H. Cohen and J. C. Kaufeld, Jr. UNIX time-sharing system: The network operations center system. The Bell System Technical Journal, 57(6): 2289–2304, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2289.
                                        pdf.

Coke:1978:RCP

[CK78b] E. U. Coke and M. E. Koether. READ — C programs for analyzing readability. Technical Memorandum 1331 (TM 78-1222-9), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 22, 1978. ?? pp.

Campbell:1985:PWR

[CK85] Roy H. Campbell and William J. Kubitz. Professional workstation research project. IEEE Computer Graphics and Applications, 6(5):17–24, May 1985. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Chryselius:2006:DQE

[CK06a] Toralf Chryselius and Andrea Kuntz. Debian unter Qemu Einführung in das Betriebssystem Debian Linux in der virtuellen Umgebung Qemu unter Windows. (German) [Debian under Qemu: Introduction in the Debian Linux operating systems in the Qemu virtual machine under Windows], volume 17 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-116-3 (book), 3-86768-716-1 (DVD). 159 pp. LCCN ????

Chryselius:2006:IDQ

[CK06b] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in Debian unter Qemu Einführung in das Betriebssystem Debian Linux in Qemu und Vorstellung der wichtigsten Internetprogramme. (German) [Internet Communication in Debian under Qemu: Introduction in the Debian Linux operating system in Qemu and creation of the most important Internet programs], volume 18 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-117-1 (book), 3-86768-717-X (DVD). 109 pp. LCCN ????

Chryselius:2006:IOV

[CK06c] Toralf Chryselius and Andrea Kuntz. Internetkommunikation in OpenSUSE unter VMware [Qemu] Einführung in das Betriebssystem Open-SUSE Linux und Vorstellung von Internetprogrammen in der virtuellen Umgebung VMware, volume 66 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-165-1 (book). 117 pp. LCCN ????

Chryselius:2006:IOQ

[CK06d] Toralf Chryselius and Andrea Kuntz. Internetkommunikation mit OpenSUSE unter Qemu: Einführung in das Betriebssystem OpenSUSE Linux und Vorstellung von Internetprogrammen in der virtuellen Umgebung Qemu, volume 66 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-165-1. 104 pp. LCCN ????

Chryselius:2006:KLQb

[CK06e] Toralf Chryselius and Andrea Kuntz. Kanotix Linux unter Qemu Einführung in das Betriebssystem Kanotix Linux in der virtuellen Umgebung Qemu unter Windows, volume 33 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-132-5 (book), 3-86768-732-3 (DVD). 156 pp. LCCN ????

Chryselius:2006:KLQa

[CK06f] Toralf Chryselius and Andrea Kuntz. Knoppix Linux unter Qemu Einführung in das Betriebssystem Knoppix Linux in der virtuellen Umgebung Qemu unter Windows, volume 49 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-148-1 (book), 3-86768-748-X (DVD). 142 pp. LCCN ????

Chryselius:2006:KQE

[CK06g] Toralf Chryselius and Andrea Kuntz. Kubuntu unter Qemu Einführung in das Betriebssystem Kubuntu Linux in der virtuellen Umgebung Qemu, volume 5 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-104-X (book), 3-86768-704-8 (DVD). 158 pp. LCCN ????

Chryselius:2006:LOL

[CK06h] Toralf Chryselius and Andrea Kuntz. Lernprogramme mit OpenSUSE Linux unter Qemu: Einführung in das Betriebssystem, OpenSUSE Linux und Vorstellung von Lernprogrammen in der virtuellen Umgebung Quemu, volume 63 of Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-166-X, 3-86768-766-8. 147 pp. LCCN ????

Chryselius:2006:OLQ

[CK06i] Toralf Chryselius and Andrea Kuntz. OpenSuSE Linux unter Qemu Einführung in das Betriebssystem Open-SUSE Linux in der virtuellen Umgebng Qemu unter Windows, volume 65 of Schriftenreihe Grenzgänger - Linux leicht verständlich; Schriftenreihe Grenzgänger - Linux leicht verständlich. CVTD, Bergfelde bei Berlin, Germany, 2006. ISBN 3-86768-164-3 (book). 168 pp. LCCN ????

Curry:1990:AAC

[CKDS90] David A. Curry, Samuel D. Kimery, Kent C. De La Croix, and Jeffrey R. Schwab. AC-MAINT: An account creation and maintenance system for distributed UNIX systems. In USENIX [USE90b], pages 1–10. LCCN QA76.76.O63 L37 1990.

Cabrera:1985:IBM

[CKM85] Luis Felipe Cabrera, Michael J. Karels, and David Mosher. The impact of buffer management on networking software performance in Berkeley UNIX 4.2BSD: a case study. In USENIX Association [USE85d], pages 507–518. LCCN QA76.8.U65 U8 1985.

Cox:2006:XSU

[CKM06] Russ Cox, Frans Kaashoek, and Robert Morris. Xv6, a simple Unix-like teaching operating system. MIT course 6.828 Operating System Engineer-

ing, 2006. URL http://pdos.
csail.mit.edu/6.828/2014/
xv6.html.

Cox:2014:XSU

[CKM14] Russ Cox, Frans Kaashoek, and Robert Morris. xv6: a simple, Unix-like teaching operating system. Technical report, ????, ????, September 3, 2014.

101 pp. URL https://pdos.
csail.mit.edu/6.828/2014/
xv6/book-rev8.pdf.

Condon:1980:EML

[CKT80] Joe Condon, Brian Kernighan, and Ken Thompson. Experience with the Mergenthaler Linotron 202 phototypesetter, or, how we spent our summer vacation. Technical Memorandum TM-80-1270-1, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 6, 1980. 11 + 2 pp. URL http:

//www.cs.princeton.edu/~

bwk/202/. Publication of this internal memo of 1980 was suppressed for 33 years, but both the original, and the reconstructed memo are now available on the Web. See [BBK13] for the story of its reconstruction and Web publication.

Curewitz:1993:PPD

[CKV93] Kenneth M. Curewitz, P. Krishnan, and Jeffrey Scott Vitter. Practical prefetching via data compression. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):257–266, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Chang:1979:DIS

[CL79] C. W. Chang and J. P. Linderman. DIRECT II service dispatcher. Technical Memorandum 1425 (TM 79-9156-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 12, 1979. ?? pp.

Cherry:1990:TPF

[CL90a] L. L. Cherry and M. E. Lesk. Tbl — a program to format tables. In Hume and McIlroy [HM90b], pages 35– 51. ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Cherry:1990:TDU

[CL90b] L. L. Cherry and M. E. Lesk. Typing documents on the UNIX system: Using the -ms and -mcs macros with Troff. In Hume and McIlroy [HM90b], pages 7–17. ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990.

URL http://doc.cat-v.org/
unix/v10/10thEdMan/msdoc.
pdf;   http://doc.cat-v.
org/unix/v10/10thEdMan/v2contents.

html. Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Christin:2006:QQS

[CL06] Nicolas Christin and Jörg Liebeherr. The QoSbox: Quantitative service differentiation in BSD routers. Computer Networks (Amsterdam, Netherlands: 1999), 50(17):3353– 3374, December 5, 2006. CO-DEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic).

Clark:1979:ERD

[Cla79] D. L. Clark. Evaluation of the remote diagnostic console for use on Bell System PDP-11/ 70 computer systems. Technical Memorandum 1392, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 11, 1979. ?? pp.

Clancy:1988:VME

[Cla88] Patrick Clancy. Virtual memory extensions in TRACE/ UNIX. In USENIX Association [USE88f], pages 137–150. ISBN ???? LCCN ????

Clayton:2000:COE

[Cla00] Richard Clayton. Good practice for combating unsolicited bulk email. RIPE/Demon Internet document, 2000. URL

                                        http://www.ripe.net/ripe/
                                        docs/ripe-206.html.

Clegg:1983:HEU

[Cle83a] Frederick W. Clegg. Hewlett– Packard’s entry into the UNIX community. In Association [Ass83], pages 119–131.

Clegg:1983:HPE

[Cle83b] Frederick W. Clegg. Hewlett– Packard’s entry into the UNIX community. In USENIX [USE83], pages 119–131.

Chase:1994:SPS

[CLFL94] Jeffrey S. Chase, Henry M. Levy, Michael J. Feeley, and Edward D. Lazowska. Sharing and protection in a singleaddress-space operating system. ACM Transactions on Computer Systems, 12(4):271– 307, November 1994. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tocs/1994-12-4/p271-chase/

.

Cheng:1995:SPP

[CLJC95] T. Cheng, K. Leung, M. Jin, and E. Chu. ScanSAR and precision processor implementation at the Alaska SAR facility. International Geoscience and Remote Sensing Symposium (IGARSS), 3:2302–2306, 1995. CODEN IGRSE3. IEEE catalog number 95CH35770.

Cherry:1978:BAP

[CM78a] Lorinda L. Cherry and Robert Morris. BC — an arbitrary precision desk calculator language. Technical Memorandum 1053, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 12, 1978. 2 + 14 pp.

Cohen:1978:UTR

[CM78b] A. S. Cohen and J. F. Maranzano. UNIX/TS Release 1.0 plans. Technical Memorandum 1232 (MF 78-8234-11), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 30, 1978. ?? pp.

Cooper:2004:HUI

[CM04] Chris Cooper and C. G.

(Chris G.) Moore. HP-UX 11i internals. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-032861-8. xxxiii + 386 pp. LCCN QA76.76.O63 C6644 2004.

Chae:2006:ATA

[CM06] Bongsug (Kevin) Chae and Roger McHaney. Asian trio’s adoption of Linux-based open source development. Communications of the Association for Computing Machinery, 49(9): 95–99, September 2006. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cmelik:1986:CMD

[Cme86] R. F. Cmelik. Concurrent make: A distributed program in Concurrent C. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1986. ?? pp.

Certen:1988:UEG

[CMM88] Pascale Le Certen, Beatrice Michel, and Gilles Muller. A UNIX environment for the GOTHIC kernel. In USENIX Association [USE88c], pages 219–229. ISBN ???? LCCN ????

Carl-Mitchell:1993:PIT

[CMQ93] Smoot Carl-Mitchell and John S. Quarterman. Practical Internetworking with TCP/IP and UNIX. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-58629-0. xviii + 476 pp. LCCN TK5105.5 .C36 1993. US$43.25.

Chang:1990:ESF

[CMR^{+}90] A. Chang, M. F. Mergen, R. K. Rader, J. A. Roberts, and S. L. Porter. Evolution of storage facilities in AIX version 3 for RISC System/6000 processors. IBM Journal of Research and Development, 34(1): 105–110, January 1990. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Caviglione:2021:KLT

[CMR^{+}21] Luca Caviglione, Wojciech Mazurczyk, Matteo Repetto, Andreas Schaffhauser, and Marco Zuppelli. Kernel-level tracing for detecting stegomalware and covert channels in Linux environments. Computer Networks (Amsterdam, Netherlands: 1999), 191(??): ??, May 22, 2021. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1389128621001249.

Comer:1988:USC

[CN88] Douglas Comer and Thomas Narten. UNIX systems as Cypress implets. In USENIX Association [USE88g], pages 55– 62. ISBN ???? LCCN ????

Chu:1997:SRT

[CN97] H.-H. Chu and K. Nahrstedt. A soft real time scheduling server in UNIX operating system. Lecture Notes in Computer Science, 1309:153–??, 1997. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cockcroft:1994:SPT

[Coc94] Adrian Cockcroft. Sun Performance and Tuning: SPARC and Solaris. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-149642-5. 254 pp. LCCN QA76.8.S86 C63 1994. US$38.00.

Cockcroft:1998:SPT

[Coc98] Adrian Cockcroft. Sun Performance and Tuning: SPARC and Solaris. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-095249-4. xxxvi + 587 pp. LCCN QA76.8.S86C63

1998. URL http://www.fdds.
com/books/catalog/books_
comingsoon.html;   http:/
/www.sun.com/books/catalog/
Cockcroft/index.html.

Cochran:2001:NVI

[Coc01a] Shannon Cochran. News and views: ICFP programming contest concludes; lost Moon landing tape recovered; W3C patent controversy; the UNIX epoch; connecting the quantum dots; Programmer’s Guild launches membership campaign. Dr. Dobb’s Journal of Software Tools, 26(12):20, December 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Cochran:2001:NVL

[Coc01b] Shannon Cochran. News and views: Linux Standard Base 1.0 released; now everyone knows you’re a dog; maximum fiberoptic capacity calculated; mass producing quantum chips; mo’ better batteries; patenting the software service model. Dr. Dobb’s Journal of Software Tools, 26(10):18, October 2001. CODEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Cochran:2001:NVW

[Coc01c] Shannon Cochran. News and views: WEP [wired equivalent privacy]: Pining for the fjords?; bye bye Be; UNIX utilities open sourced; cosmology computer calculates creation; UML 2.0 infrastructure proposals. Dr. Dobb’s Journal of Software Tools, 26(11):18, November 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Cochran:2003:NVGb

[Coc03a]  Shannon Cochran.  News

and views: Government ponders open-source strategy; POSIX, Single UNIX specification merged; IBM plans new supercomputers; robotic surgeons have a heart. Dr. Dobb’s Journal of Software Tools, 28 (2):14, February 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7790/
ddj0302o/.

Cochran:2003:NVe

[Coc03b] Shannon Cochran. News and views: MIT honors builder of robot swarm; developing space; Embedded Linux platform specification released; biometric systems at U.S. borders, says NIST. Dr. Dobb’s Journal of Software Tools, 28(5):14, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7927/ddj0305o/.

Cochran:2003:NVT

[Coc03c] Shannon Cochran. News and views: Tiny executable contest concludes; DeCSS, DMCA prosecutions fail; studying for the Turing test; Embedded Linux specification released. Dr. Dobb’s Journal of Software Tools, 28(3):14, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/documents/s=7826/
                                        ddj0303n/.

Cody:1986:ALB

[Cod86] W. J. Cody. An alternative library under 4.2 BSD UNIX on a VAX 11/780. Technical Report ANL-86-10, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, February 1986. iii + 30 pp.

Cody:1989:ETR

[Cod89] W. J. Cody. ELEFUNT test results using Titan Fortran under Ardent UNIX 2.0 on the Titan. Technical Report MCS-TM-129, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, March 1989. iii + 14 pp.

Coffin:1990:USV

[Cof90] Stephen Coffin. UNIX System V Release 4: The Complete Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881653-X. xxxviii + 905 pp. LCCN QA76.76.O63 C6415 1990. Another good book on UNIX fundamentals and related subjects.

Cohen:1979:SSL

[Coh79] A. S. Cohen. Supported structure for local modifications to UNIX. Technical Memorandum 1420, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 26, 1979. ?? pp.

Cohrs:1990:AAS

[Coh90]  D. Cohrs.  Authmgr —

an authentication service for Datakit. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Cohn:1994:AC

[Coh94] David L. Cohn. An AIX Companion. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-291220-1. xvii + 476 pp. LCCN QA76.76.O63 C64 1994. US$38.00.

Cole:1983:AAP

[Col83] Clement T. Cole. Attaching an array processor in the UNIX environment. In Software Tools Users Group [Sof83], pages 135–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Collins:1984:WWU

[Col84a] Peter E. Collins. WINDX — Windows for the UNIX environment. In Software Tools Users Group [Sof84], pages 159–165. LCCN QA76.8.U65 U83 1984.

Collinson:1984:DUO

[Col84b] Peter Collinson. On the design of the UNIX operating system. AUUGN Australian Unix User Group Newsletter, 5(4):21–23, July 24, 1984.

                                        URL http://www.tuhs.org/
                                        Archive/Documentation/AUUGN/
                                        AUUGN-V05.4.pdf.

Collinson:1987:UC

[Col87] Peter Collinson. UNIX: The cult. In USENIX Association [USE87e], pages 22–28. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Collyer:1989:PTT

[Col89] Geoff Collyer. A partial tour through the UNIX shell. In USENIX Association [USE89f], pages 343–353.

Collinson:1992:BRA

[Col92a] Peter Collinson. Book review: Advanced Programming in the UNIX Environment. ;login: the USENIX Association newsletter, 17(5):44–45, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Colon:1992:VTV

[CoL92b] Robert Colon (Tutor), others, and Maarten Litmaati (Reference). vi Tutor and vi Reference, 2.1 (tutor), 8 (reference) edition, 1992. The latest interactive tutorial (vitutor2. 1.shar(.Z)) can be obtained by anonymous ftp from ftp.

mines.colorado.edu (in /
pub/tutorials) . . . . The ref-

erence and other vi stuff are obtainable by anonymous ftp

from ftp.uwp.edu (in /pub/
vi).

Collinson:1994:PKP

[Col94] P. Collinson. Pop Korn — the POSIX shell. .EXE: the software developers’ magazine, 8 (9):64–??, ???? 1994. CODEN EXEEE5. ISSN 0268-6872.

Collings:2002:RLN

[Col02] Terry Collings. Red Hat Linux networking and system administration. Transworld, London, UK, 2002. ISBN 0-7645-3632-X. xxix + 843 pp. LCCN QA76.76.O63 C625 2002.

Collyer:2017:PTT

[Col17] Geoff Collyer. A partial tour through the UNIX shell. Report, Department of Statistics, University of Toronto, Toronto, ON M5S 1A1, Canada, July 20, 2017. 11 pp. URL http://www.

collyer.net/who/geoff/sh.
tour.pdf.

UC:1979:TEF

[Com79] Computer Center, University of California, Santa Barbara. Text editing, formatting and typesetting using the UCSB UNIX system, 1979. 244 pp.

CCA:1984:CEMa

[Com84] Computer Corporation of America, Cambridge, MA, USA. CCA EMACS manual for UNIX users, version 162.43z edition, 1984. various pp.

CSL:1994:POS

[Com94a] Computer Systems Laboratory (U.S.). Portable Operating System Interface (POSIX). Part 2, Shell and utilities. Washington, DC, USA, October 11, 1994. 7 pp. Shipping list no.: 95-0036-P.

CSRG:1994:CRC

[Com94b] Computer Systems Research Group, UC Berkeley. 4.4BSD-Lite CD-ROM Companion. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1994. ISBN 1-56592-081-3 (domestic edition), 1-56592-092-9 (international edition). 112 pp. LCCN QA76.8.U65F67 .F684 1994. US$40.00.

CSRG:1994:PRM

[Com94c] Computer Systems Research Group, UC Berkeley. 4.4BSD Programmer’s Reference Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1994. ISBN 1-56592-078-3. 886 pp. LCCN QA76.8.U65 F682 1994. US$30.00.

CSRG:1994:PSD

[Com94d] Computer Systems Research Group, UC Berkeley. 4.4BSD Programmer’s Supplementary Documents. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1994. ISBN 1-56592-079-1. 596 pp. LCCN QA76.8.U65 F683 1994. US$30.00.

CSRG:1994:SMM

[Com94e] Computer Systems Research Group, UC Berkeley. 4.4BSD System Manager’s Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1994. ISBN 1-56592-080-5. 804 pp. LCCN QA76.8.U65 F68 1994. US$30.00.

CSRG:1994:URM

[Com94f] Computer Systems Research Group, UC Berkeley. 4.4BSD User’s Reference Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1994. ISBN 1-56592-075-9. 905 pp. LCCN QA76.8.U65 F67 1994. US$30.00.

CSRG:1994:USD

[Com94g] Computer Systems Research Group, UC Berkeley. 4.4BSD User’s Supplementary Documents. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1994. ISBN 1-56592-076-7. 712 pp. LCCN QA76.8.U65 F672 1994. US$30.00.

comp-lang-awk:1997:CLA

[com97] comp.lang.awk. Internet mailing list, 1997. URL news:comp.

                                        lang.awk.

Compton:2000:VLB

[Com00] Jason Compton. VMware 2 for Linux: [a better way to run multiple operating systems on Linux]. Prima Tech, Rocklin, CA, USA, 2000. ISBN 0-7615-2764-8. xxii + 406 pp. LCCN QA76.76.O63 C656 2000.

Compton:2003:VL

[Com03] Jason Compton. VMware 2 for Linux. Premier Press, Rocklin, CA, USA, 2003. ISBN 0-7615-2764-8. xxii + 406 pp. LCCN QA76.76.O63 C656 2000. US$39.99. Includes CD-ROM.

Comer:2012:OSD

[Com12] Douglas E. Comer. Operating system design: the Xinu approach, Linksys version. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-8109-X (hardcover). xxii + 604 pp. LCCN QA76.76.O63 C65 2012.

Cooper:1985:SUQ

[Coo85] R. E. M. Cooper. On squeezing the UNIX quart into a Data General Eclipse pint pot. Software—Practice and Experience, 15(6):571–581, June 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Cook:1998:LHF

[Coo98] Anthony L. Cook. Linuxhosted frame relay. Sys Admin: The Journal for UNIX Systems Administrators, 7(1): 8, 10, 14, 16, 18–19, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Cooper:2003:BWH

[Coo03] Joe (R. Joe) Cooper. The book of Webmin, or, How I learned to stop worrying and love Unix. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-92-1. xxviii + 281 pp. LCCN QA76.9.U83 C6597 2003eb. URL http:/

/www.oreilly.com/catalog/
9781886411920.

Cooperstein:2009:LPDa

[Coo09a] Jerry Cooperstein. Linux Program Development: a guide with exercises. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4499-0602-8. 290 (est.) pp. LCCN ????

Cooperstein:2009:LPDb

[Coo09b] Jerry Cooperstein. Linux Program Development: Lab Solutions: a guide with exercises. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4499-0604-4. 204 (est.) pp. LCCN ????

Cooperstein:2009:WLDa

[Coo09c] Jerry Cooperstein. Writing Linux Device Drivers: a guide with exercises, volume 3. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4486-7238-4. 394 (est.) pp. LCCN ????

Cooperstein:2009:WLDb

[Coo09d] Jerry Cooperstein. Writing Linux Device Drivers: Lab Solutions: a guide with exercises. CreateSpace, Scotts Valley, CA, USA, 2009. ISBN 1-4495-3124-5. 270 (est.) pp. LCCN ????

Coplien:1992:ACP

[Cop92] James Coplien. Advanced C++ Programming Styles and Idioms. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-54855-0. xxiv + 520 pp. LCCN QA76.73.C153 C67 1992. An advanced book for any C++ expert-wanna-be.

LSBT:2005:BAL

[Cor05a] Core Members of the Linux Standard Base Team. Building applications with the Linux Standard Base. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-145695-4. xxvi + 246 pp. LCCN QA76.76.O63 B8375 2004. URL http:

                                        //www.freestandards.org/;
                                         http://www.lanana.org/;
                                        http://www.linuxbase.org/

;   http://www.linuxbase.
org/test/registered.html;
 http://www.phptr.com/title/
0131456954;   https://www.
linux-foundation.org/en/

Book. Foreword by Theodore Ts’o. Includes CD-ROM.

Cornelio:2005:MLG

[Cor05b] Pietro Cornelio. Il mondo libero di GNU/Linux e UNIX BSD: storia, filosofia, tecnologia. (Italian) [The free world of GNU/Linux and BSD UNIX: stories, philosophies, technology]. Duke Italia, Milano, Italy, 2005. ISBN 88-86460-10-4. 104 pp. LCCN ???? Suppement to Linux Journal (Italian edition), number 59, October 2005.

Costales:1997:SDR

[Cos97] Bryan Costales. sendmail Desktop Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-278-6. viii + 68 pp. LCCN ???? US$6.95.

URL http://www.ora.com/
catalog/sendmailqr/;   http:
//www.oreilly.com/catalog/
sendmailqr.

Coughran:1984:NBC

[Cou84] W. M. Coughran, Jr. On noncharacteristic boundary conditions for discrete hyperbolic initial-boundary-value problems. Computing Science Technical Report 94, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1984. ?? pp.

Coughran:1985:NBC

[Cou85a] W. M. Coughran, Jr. On noncharacteristic boundary conditions for discrete hyperbolic initial-boundary-value problems. Journal of Computational Physics, 60(1):135–154, August 1985. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        002199918590021X.

Courington:1985:USS

[Cou85b] Bill Courington. The UNIX system: a Sun technical report. Report, Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1985. xi + 113 pp. URL https:

                                        //www.rcsri.org/library/
                                        80s/UNIX-A-Sun-Tech-Report.
                                        pdf.

Courtes:2013:FPM

[Cou13] Ludovic Courtès. Functional package management with Guix. arxiv.org, ??(??):1–11, May 20, 2013. URL https:/

                                        /arxiv.org/abs/1305.4584.

Courtes:2017:CSG

[Cou17] Ludovic Courtès. Code staging in GNU Guix. ACM SIGPLAN Notices, 52(12):41–48, December 2017. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1979:UCS

[COV79] A. S. Cohen, S. B. Olsson, and G. C. Vogel. UNIX command syntax. Technical Memorandum 1358, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 16, 1979. ?? pp.

Covington:2006:SGG

[Cov06] Jim Covington. The Sun Geeks Guide To Native LDAP: a Native LDAP Blueprint. Book-Surge Publishing, ????, 2006. ISBN 1-4196-3028-8. 126 (est.) pp. LCCN ???? US$34.95.

Cox:1983:OOP

[Cox83] Brad J. Cox. The object oriented precompiler: Programming Smalltalk 80 methods in C language. ACM SIGPLAN Notices, 18(1):15–22, January 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cox:2007:REM

[Cox07] Russ Cox. Regular expression matching can be simple and fast (but is slow in Java, Perl, PHP, Python, Ruby, …). Report, swtch.com, Cambridge, MA, USA, January

2007. URL http://swtch.
com/~rsc/regexp/regexp1.

html. See also [Tho68, KP99, Cox09, Cox10, Cox12].

Cox:2009:REM

[Cox09] Russ Cox. Regular expression matching: the virtual machine approach. Report, swtch.com, Cambridge, MA, USA, Decem-

ber 2009. URL http://swtch.
com/~rsc/regexp/regexp2.

html. See also [Tho68, KP99, Cox07, Cox10, Cox12].

Cox:2010:REM

[Cox10] Russ Cox. Regular expression matching in the wild. Report, swtch.com, Cambridge, MA, USA, March 2010. URL http:

//swtch.com/~rsc/regexp/
regexp3.html.  See also

[Tho68, KP99, Cox07, Cox09, Cox12].

Cox:2012:REM

[Cox12] Russ Cox. Regular expression matching with a trigram index, or how Google Code Search worked. Report, swtch.com, Cambridge, MA, USA, January

2012. URL http://swtch.
com/~rsc/regexp/regexp4.

html. See also [Tho68, KP99, Cox07, Cox09, Cox10].

Corsini:1986:MID

[CP86] Paolo Corsini and Cosimo Antonio Prete. Multibug: Interactive debugging in distributed systems. IEEE Micro, 6(3): 26–33, May/June 1986. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Crowcroft:2001:TIL

[CP01] Jon Crowcroft and Iain Phillips. TCP/IP and the Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. Wiley, New York, NY, USA, 2001.

ISBN 0-471-40882-4. 925 pp. LCCN TK5105.585 .T34 2001.

Crowcroft:2002:TIL

[CP02] Jon Crowcroft and Iain Phillips.

TCP/IP and Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. Wiley, New York, NY, USA, 2002. ISBN 0-471-40882-4 (cloth). xlix + 925 pp. LCCN TK5105.585 .T34 2002.

Clinton:1998:LBM

[CPJ^{+}98] Terry Clinton, Tom Parsons, Capers Jones, William Adams, Garth Klatt, Eric Haines, Ted Lewis, Philip Machanik, Stig Nilsson, Karl Reed, Howard R. Stearns, Neville Holmes, and John Brownie. Letters: The benefits of model-based integration; documentation is not green; picking on the overdog; buggy, slow windoze; there’s no such thing as free software; Linus’ law of open source development; bug-free development? No way; governmental IT planning and the Computer Society; text encoding questions; encoding the world’s languages. Computer, 31(11):4, 5– 7, 11, November 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1998/pdf/

ry004.pdf. Two letters discuss Unicode and Multicode [Mud97].

Chambers:1983:USI

[CQ83a] John Chambers and John Quarterman. UNIX System III and 4.1BSD; a practical comparison. In Association [Ass83], pages 25–38.

Chambers:1983:USV

[CQ83b] John Chambers and John Quarterman. UNIX system V and 4.1C BSD. In Software Tools Users Group [Sof83], pages 265–291. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Cameron:1991:LGE

[CR91] Debra Cameron and Bill Rosenblatt. Learning GNU Emacs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-84-6. xxvii + 411 pp. LCCN QA76.76.T49 C35 1991.

                                        URL http://www.oreilly.
                                        com/catalog/9780937175842.

Chang:1992:EPMa

[CR92] Henry H. Y. Chang and Bryan Rosenburg. Experience porting Mach to the RP3 large-scale shared-memory multiprocessor. Future Generation Computer Systems, 7(2–3):259–267, April 1992. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Christian:1994:UOS

[CR94] Kaare Christian and Susan Richter. The UNIX Operating System. Wiley-Interscience, New York, NY, USA, third edition, 1994. ISBN 0-471-58684-6. LCCN QA76.76.O63 C465 1994.

Cade:2002:SCE

[CR02] Mark Cade and Simon Roberts.

Sun Certified Enterprise Architect for J2EE Study Guide. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-044916-4. xviii + 199 pp. LCCN QA76.3.C23 2002.

Cragun:1983:USD

[Cra83] Don Cragun. UNIX system definitions and standards. In Association [Ass83], pages 112–?? Abstract only.

Cramer:1988:WRT

[Cra88] William D. Cramer. Writing real-time programs under UNIX. Dr. Dobb’s Journal of Software Tools, 13(6):18–??, June 1988. CODEN DDJOEB. ISSN 1044-789X. CrayResearch:1990:UXW

[Cra90] Cray Research, Inc. UNICOS X Window System reference manual. Mendota Heights, MN, version 6.0 edition, 1990. ix + 54 pp.

Crabb:1993:RUI

[Cra93] Don Crabb. Running UNIX so it doesn’t run you. Ziff-Davis Press, Emeryville, CA, USA, 1993. ISBN 1-56276-061-0. xix + 268 pp. LCCN QA76.76.O63 C724 1993. US$29.95.

Corbet:2005:LDD

[CRKH05] Jonathan Corbet, Alessandro Rubini, and Greg Kroah-Hartman. Linux device drivers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00590-3. xviii + 615 pp. LCCN QA76.76.D49 R92 2005; QA76.76.D49 R92 2005eb; QA76.76.D49. URL http:/

/www.oreilly.com/catalog/
9780596005900.

Crowley:1978:UTS

[Cro78a] T. H. Crowley. UNIX timesharing system: Preface. The Bell System Technical Journal, 57(6):1897–1898, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1897.
pdf.

Crowley:1978:RTC

[Cro78b] T. R. Crowley. Regression testing for C compilers. Technical Memorandum 1295 (MF 78-8234-78), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 10, 1978. ?? pp.

Crowley:1978:TLU

[Cro78c] T. R. Crowley. Title listings for UNIX memoranda. Technical Memorandum 1296 (MF 78-8234-79), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 7, 1978. ?? pp.

Crowe:1982:IC

[Cro82] M. K. Crowe. An incremental compiler. ACM SIGPLAN Notices, 17(10):13–22, October 1982. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crosser:1998:IFT

[Cro98a] Eugene G. Crosser. Ifmail:

FidoNet — technology implementation of UNIX platform. In USENIX [USE98b], page ?? ISBN ???? LCCN ????

URL  http://www.usenix.
org/publications/library/
proceedings/usenix98/freenix/
crosser.ps.

Crowe:1998:PTC

[Cro98b]  William L. Crowe.  A

pseudo-terminal class for Unix. C/C++ Users Journal, 16(3): ??, March 1998. CODEN CCU-JEX. ISSN 1075-2838.

Clark:1986:AAF

[CS86] Paul Clark and Andrew Simms.

AFQL — a flexible, general purpose interface to relational database management systems under UNIX. EUUG Newsletter, 6(3):7–25, Winter 1986. CODEN EONLE8. ISSN 1011-4211.

Curran:1990:CBC

[CS90a] Stephan Curran and Michael Stumm. A comparison of basic CPU scheduling algorithms for multiprocessor UNIX. In USENIX [USE90a], pages 551–

Curran:COMPSYS-3-4-551

[CS90b] Stephan Curran and Michael Stumm. A comparison of basic CPU scheduling algorithms for multiprocessor UNIX. Computing Systems, 3(4):551–579, Fall 1990. CODEN CMSYE2. ISSN 0895-6340.

Comer:1991:ITIa

[CS91a] Douglas Comer and David Stevens. Internetworking with TCP/IP: Principles, Protocols and Architecture. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-468505-9. LCCN TK5105.5 .C59 1991.

Comer:1991:ITIb

[CS91b] Douglas E. Comer and David L. Stevens. Internetworking with TCP/IP. Design, Implementation, and Internals, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-465378-5, 0-13-472242-6. LCCN TK5105.5 .C59 1991-92.

Comer:1993:CSP

[CS93] Douglas E. Comer and David L. Stevens. 3: Client-server programming and applications: BSD socket version. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-020272-X. xxiv + 498 pp. LCCN ????

Ciampolini:1996:EPM

[CS96a] A. Ciampolini and C. Stefanelli. Extending PVM to a massively parallel architecture. Future Generation Computer Systems, 12(1):13–23, May 1996. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Comer:1996:VCS

[CS96b] Douglas E. Comer and David L.

Stevens. Vol. 3: Client-server programming and applications: BSD socket version. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1996. ISBN 0-13-262148-7. xxvi + 519 pp. LCCN ????

Carter:1999:STY

[CS99] Gerald Carter and Richard Sharpe. Sams Teach Yourself Samba in 24 Hours. SAMS Publishing, Indianapolis, IN, USA, April 1999. ISBN 0-672-31609-9. xv + 490 pp. LCCN QA76.9.C55C36 1999. US$24.99.

Csaszar:1977:ECU

[Csa77] M. A. Csaszar. Examples of converting USG shell command files for the new shell. Technical Memorandum 1210 (MF 77-8234-100), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 22, 1977. ?? pp.

Carr:1987:IPC

[CSA^{+}87] P. Carr, R. Stevenson, J. Alea, J. Berthold, G. Groucher, M. Davis, G. Dobbins, D. Law, V. Szarek, and W. Webster. Implementation of a prototype CAIS environment. ACM SIGADA Ada Letters, 7(2):58– 72, March/April 1987. CO-DEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Costales:2004:SC

[CSAJ04] Bryan Costales, Gregory Neil Shapiro, Claus Assmann, and George Jansen. Sendmail 8.13 companion. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00845-7 (paperback). x + 179 pp. LCCN TK5105.74.S44 C67 2004.

Curtis:TB4-1-18

[CT83] Pavel Curtis and Howard Trickey. Porting TEX to /UNIX. TUGboat, 4(1):18– 20, April 1983. ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb04-1/tb07site.
                                        pdf.

Cutler:1994:SUN

[CT94] Ellie Cutler and The staff of O’Reilly and Associates. SCO UNIX in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 1994. ISBN 1-56592-037-6. xix + 568 pp. LCCN QA76.76.O63 C93 1994. US$19.95. URL http://www.

                                        oreilly.com/catalog/sco.

Cannon:1999:SA

[CTJ99] Casey Cannon, Scott Trent, and Carolyn Jones. Simply AIX 4.3. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-021344-6. 368 pp. LCCN QA76.76.O63C3725

1999.  URL http://www.
phptr.com/ptrbooks/ptr_0130213446.
html.

Cuccia:1990:BRU

[Cuc90] Nichlos H. Cuccia. Book review: UNIX System Administration Handbook. ;login: the USENIX Association newsletter, 15(4):13–??, July/August 1990. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Cooke:1999:UBI

[CUH99] Daniel Cooke, Joseph Urban, and Scott Hamilton. Unix and beyond: An interview with Ken Thompson. Computer, 32(5): 58–64, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r5058.pdf;   http://www.
computer.org/computer/co1999/
r5058abs.htm.

Cummingham:1977:NPG

[Cum77] N. J. Cummingham. NROFF for producing generic program documentation. Technical Memorandum 1149 (TM 77-5222-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 18, 1977. ?? pp.

Curry:1989:UCU

[Cur89] Dave Curry. Using C on the UNIX System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1989. ISBN 0-937175-23-4. 250 pp. LCCN QA76.73.C15 C87 1989. US$24.95.

Curry:1990:UCU

[Cur90] David Curry. Using C on the UNIX System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-23-4. xii + 379 pp. LCCN QA76.76.O63 H68 1992. US$24.95. This book is directed to (would-be) system programmers.

Curry:1991:AAC

[Cur91] David Curry. ACMAINT: An account creation and maintenance system for distributed UNIX systems. In USENIX [USE91a], page ?? ISBN ???? LCCN ????

Curry:1992:USS

[Cur92] David Curry. UNIX System Security — a Guide for Users and System Administrators. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56327-4. xiii + 279 pp. LCCN QA76.9.A25 C87 1992. Comprehensive coverage … with pointers to further information.

Curry:1996:USP

[Cur96] David A. (Allan) Curry. UNIX Systems Programming for SVR4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1996. ISBN 1-56592-163-1. xxi + 596 pp. LCCN QA76.76.O63C865 1996. US$34.95. URL http:/

/www.oreilly.com/catalog/
usp.

Curry:1997:RUS

[Cur97] David Curry. Review: UNIX systems programming for SVR4. ;login: the USENIX Association newsletter, 22(1):58– 59, February 1997. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Custer:1992:GTW

[Cus92] Helen Custer. A grand tour of Windows NT: Portable 32-bit multiprocessing comes to Windows. Microsoft Systems Journal, 7(4):17–31, July 1992. CO-DEN MSJOED. ISSN 0889-9932.

Custer:1993:IWN

[Cus93] Helen Custer. Inside Windows NT. Microsoft Press, Bellevue, WA, USA, 1993. ISBN 1-55615-481-X. xxiv + 385 pp. LCCN QA76.76.O63 C89 1993. The authoritative technical reference on Windows NT (the ultimate challenge to UNIX?) ….

Cusumano:2004:TSMc

[Cus04] Michael A. Cusumano. Technology strategy and management: Reflections on free and open software. Communications of the Association for Computing Machinery, 47(10):

25–27, October 2004. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cherry:1980:WTS

[CV80] Lorinda L. Cherry and W. Vesterman. Writing tools — the STYLE and DICTION programs. Technical Memorandum 1497, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 16, 1980. ?? pp.

Camou:2000:DGL

[CV00] Mario Camou and Aaron Von Cowenberghe. Debian GNU/Linux 2.1 unleashed. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31700-1. xxvii + 1119 pp. LCCN QA76.76.O63 C355 2000.

Cash:1989:DCM

[CW89] J. R. Cash and M. H. Wright. A deferred correction method for nonlinear two-point boundary value problems: Implementation and numerical evaluation. Computing Science Technical Report 146, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1989. ?? pp.

Cash:1991:DCM

[CW91] J. R. Cash and M. H. Wright. A deferred correction method for nonlinear two-point boundary value problems: Implementation and numerical evaluation. SIAM Journal on Scientific and Statistical Computing, 12(4):971–989, July 1991.

CODEN SIJCD4. ISSN 0196-

5204. URL https://epubs.
siam.org/doi/abs/10.1137/
0912052.

Castagnoli:1992:DTU

[CW92] Charisse Castagnoli and Charles Watt. Distributed trusted UNIX systems. In USENIX [USE92e], pages 347–?? ISBN 1-880446-46-4. LCCN ????

Cockcroft:2000:CPI

[CW00] Adrian Cockcroft and Bill Walker. Capacity Planning for Internet Services. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-089402-8. xx + 222 pp. LCCN TK5105.5.C557 2001.

URL  books/caphi.pdf;
http://www.sun.com/books/
catalog/capplan/. Part No.

806-3684-10 May 2000, Revision 01.

Courtes:2015:RUCa

[CW15a] Ludovic Courtès and Ricardo Wurmus. Reproducible and user-controlled software environments in HPC with Guix. arxiv.org, pages 1–13,

2015. URL https://arxiv.
org/abs/1506.02822.

Courtes:2015:RUCb

[CW15b] Ludovic Courtès and Ricardo Wurmus. Reproducible and user-controlled software environments in HPC with Guix. Lecture Notes in Computer Science, 9523:579–591, 2015. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheriton:1990:BEU

[CWS90] David R. Cheriton, Gregory R. Whitehead, and Edward W. Sznyter. Binary emulation of UNIX using the V kernel. In USENIX [USE90c], pages 73– 86. LCCN QA76.8.U65 U81 1990.

Dolotta:1976:PWP

[D^{+}76] T. A. Dolotta et al. Programmer’s Workbench papers from The Second International Conference on Software Engineering. (G.4). Technical Memorandum 1107 (TM 76-9141-6, TM 76-9144-3, TM 76-9411-1, TM 76-9421-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 30, 1976. ?? pp.

Dufrasne:2004:IVE

[D^{+}04] Bertrand Dufrasne et al., editors. Implementing VMware ESX Server with IBM Total-Storage FAStT. IBM redbooks. IBM, International Support Organization, ????, September 2004. ISBN 0-7384-9134-9. xvi + 246 pp. LCCN TK5105.86 I475 2004. Publication number SG24-6434-00.

dAusbourg:1992:UOS

[d’A92] Bruno d’Ausbourg. UNIX operating services on a multilevel secure machine. In USENIX [USE92e], pages 329–346. ISBN 1-880446-46-4. LCCN ????

Dalheimer:1999:PQW

[Dal99] Matthias Kalle Dalheimer. Programming with Qt: Writing Portable GUI applications on UNIX and Win32. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-588-2. xix + 361 pp. LCCN QA76.9.U83 D355 1999. US$32.95. URL http:/

/www.oreilly.com/catalog/
prowqt/.

Dalheimer:2002:PQW

[Dal02] Matthias Kalle Dalheimer. Programming with Qt: Writing Portable GUI applications on Unix and Win32. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00064-2. xviii + 499 pp. LCCN QA76.9.U83 D355 2002 Stacks. US$39.95.

URL http://www.oreilly.
com/catalog/prowqt2.

Daniel:1983:ZAH

[Dan83a] Steve Daniel. Z — a high performance raster graphics package for UNIX operating systems. In Software Tools Users Group [Sof83], pages 135–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Daniel:1983:ZHP

[Dan83b] Steve Daniel. Z — a high performance raster graphics package for UNIX operating systems. In Association [Ass83], pages 135–?? Abstract only.

Danesh:2001:MCL

[Dan01] Arman Danesh. Mastering Corel Linux. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2852-1. xxix + 734 pp. LCCN QA76.76.O63 D3429 2001. Includes index.

Darwin:1988:CCP

[Dar88] Ian F. Darwin. Checking C programs with lint. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1988. ISBN 0-937175-30-7. x + 72 pp. LCCN QA76.73.C15 D37 1990. US$12.95.

Das:1988:UAW

[Das88] Sunil K. Das. UNIX around the world. In USENIX Association [USE88c], pages 1–6. ISBN ???? LCCN ????

Das:2001:YUU

[Das01] Sumitabha Das. Your UNIX: the ultimate guide. McGraw-Hill, New York, NY, USA, 2001. ISBN 0-07-240500-7 (paperback). liv + 918 pp. LCCN QA76.76.O63 D3495 2001.

Datdeva:1988:LMG

[Dat88] Bjorn Datdeva. Lazy man’s guide to UNIX system administration. In USENIX Association [USE88d], pages 25–?? ISBN ???? LCCN ???? Abstract only.

Datko:2014:BSA

[Dat14] Josh Datko. BeagleBone for Secret Agents. Community Experience Distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78398-604-2, 1-78398-605-0 (e-book), 1-322-15302-7 (e-book). LCCN QA76.774.L46. URL http:

//lib.myilibrary.com?id=
646557;   http://public.
eblib.com/choice/PublicFullRecord.
aspx?p=1797268.

Davis:1989:OUS

[Dav89] Janet Davis. Overview of UNIX System V Release 4.0. EUUG Newsletter, 9(1):74–78, Spring 1989. CODEN EONLE8. ISSN 1011-4211.

Davey:XR-6-1-141

[Dav93a] Paul Davey. Three help systems: xhelp, X.deskhelp, and UNIXhelp. The X Resource, 6 (1):141–151, March 1993. CO-DEN XRESEA. ISBN 1-56592-021-X. ISSN 1058-5591.

Davis:1993:MO

[Dav93b]  Tom Davis.  Moving to

OpenGL. IRIS Universe, 25: 76, 1993.

Davis:1996:LRNa

[Dav96] Dwight B. Davis. Letter from Redmond — NT versus UNIX? No contest. Datamation, 42 (14):140–??, ???? 1996. CO-DEN DTMNAT. ISSN 0011-

David:1997:EIO

[Dav97]  Olaf David.  Entwurf und

Implementation einer objektorientierten UNIX-Shell: ein Prototyping-orientierter Ansatz. (German) [Development and Implementation of an Object-Oriented UNIX Shell: a Prototyping-Oriented Introduction], volume 6 of Jenaer geographische Schriften. Friedrich-Schiller-Universität, Institut für Geographie, Selbstverlag, Jena, Germany, 1997. ISSN 0944-9795. ix + 66 pp. LCCN QA76.64 .D38 1997.

Davis:1998:SUP

[Dav98] Eric Davis. Securing UNIX passwords. Sys Admin: The Journal for UNIX Systems Administrators, 7(8):8, 10, 12, 14– 16, August 1998. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Davis:2000:IWR

[Dav00] Alan C. Davis. Implementing Web/RCS under Oracle OWS for UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 9(2):8, 10, 12, 14, February 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Davidson:2006:RMX

[Dav06] James Duncan Davidson. Running Mac OS X Panther. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-00913-5 (paperback). 372 pp. LCCN QA76.76.O63 D3498

2006.  URL http://www.
oreillynet.com/catalog/runmacx2/

.

Dawson:1982:IHS

[Daw82] Mark Dawson. An implementation of Henderson’s SECD machine under UNIX. EUUG Newsletter, 2(4):33–39, Winter 1982. CODEN EONLE8. ISSN 1011-4211.

Dawson:1986:UWP

[Daw86] N. Dawson. UNIX word processing at HAO (for use with the 4.2 BSD version and -me macro package). Report ANL-86-10, National Science Foundation, Washington, DC, USA, 1986. Contract No. W-31-109-ENG-38.

Dawidek:2007:PZF

[Daw07] Paweł Jakub Dawidek. Porting the ZFS file system to the FreeBSD operating system. In ????, editor, Proceedings of Asia BSDCON 2007, University of Tokyo, Tokyo, Japan, 8–11 March 2007, pages 97–

103. ????, ????, 2007. URL

https://2007.asiabsdcon.
org/papers/P16-paper.pdf.

Drake:1995:PUS

[DB95] Chris Drake and Kimberley Brown. Panic! UNIX System Crash Dump Analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-149386-8. xx + 492 pp. LCCN QA76.76.O63D7556

1995¡.  US$44.00.  URL

                                        http://www.sun.com/books/
                                        catalog/drake/index.html.

Includes CD ROM.

deBruijn:2008:PFL

[dBB08] Willem de Bruijn and Herbert Bos. PipesFS: fast Linux I/O in the Unix tradition. Operating Systems Review, 42(5):55– 63, July 2008. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dutton:1989:IGI

[DBG89] R. D. Dutton, R. C. Brigham, and F. Gomez. INGRID: a graph invariant manipulator. Journal of Symbolic Computation, 7(2):163–178 (or 163– 177??), February 1989. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Cruz:1987:KFT

[dC87] Frank da Cruz. Kermit — a File Transfer Protocol. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1987. ISBN 0-932376-88-6. xvii + 379 pp. LCCN TK5105.5 .D23 1987.

dAusbourg:1993:USM

[dC93] Bruno d’Ausbourg and Christel Calas. UNIX services for multilevel storage and communications over a secure LAN. In USENIX Association [USE93f], pages 157–168. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

DOrazio:2017:TCS

[DC17] Christian J. D’Orazio and Kim-Kwang Raymond Choo. A technique to circumvent SSL/ TLS validations on iOS devices. Future Generation Computer Systems, 74(??):366–374, September 2017. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167739X16302801.

DOrazio:2018:CIS

[DC18] Christian J. D’Orazio and Kim-Kwang Raymond Choo. Circumventing iOS security mechanisms for APT forensic investigations: A security taxonomy for cloud apps. Future Generation Computer Systems, 79 (part 1)(??):247–261, 2018. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167739X16305647.

Dalton:2002:DSU

[DCN02] C. I. Dalton, T. H. Choo, and A. P. Norman. Design of Secure Unix. Information Security Technical Report, 7(1):37–56, March 31, 2002. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S136341270200105X.

Droms:1990:PMX

[DD90] R. Droms and W. R. Dyksen. Performance measurements of the X Window System communication protocol. Software—Practice and Experience, 20(S2):119–136, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Durbec:2005:FDA

[DD05] C. MantéJ. P. Durbec and J. C. Dauvin. A functional dataanalytic approach to the classification of species according to their spatial dispersion. Application to a marine macrobenthic community from the Bay of Morlaix (Western English Channel). Journal of Applied Statistics, 32(8):831–840, 2005. CODEN ???? ISSN 0266-4763 (print), 1360-0532 (electronic).

Deshmukh:1994:RTM

[DDLL94] G. Deshmukh, R. Deshpande, J. Leathrum, and K. Liburdy. Role of testing methodologies in open systems standards. Proceedings — International Conference on Software Engineering, pages 233–240, 1994. CODEN PCSEDE. ISSN 0270-5257 (print), 1558-1225 (electronic). IEEE catalog number 94CH3409-0.

DeJager:1976:SMU

[De 76] D. S. De Jager. Shared memory in UNIX. Technical Memorandum 1127 (TM 76-5223-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 19, 1976. ?? pp.

DeJager:1978:ICP

[De 78] D. S. De Jager. Interprocess communication performance trade-offs in CB-UNIX and UNIX …. Technical Memorandum 1288 (TM 78-5223-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 18, 1978. ?? pp.

DeDecker:1993:USK

[De 93] B. De Decker. Unix security and Kerberos. Lecture Notes in Computer Science, 741:257– 274, 1993. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1978:MTR

[Dea78] J. Dean. Modifications to the trouble reporting system. Technical Memorandum 4 (MF 78-8234-77), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 30, 1978. ?? pp.

Dean:2001:LLC

[Dea01] Jeffrey Dean. LPI Linux certification in a nutshell: a desktop quick reference. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 2001. ISBN 1-56592-748-6. xvi + 551 pp. LCCN QA76.76.O63 D435 2001. US$39.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        9781565927483;   http://
                                        www.oreilly.com/catalog/

lpicertnut. Covers exams 101 and 102 for LPI level 1.

Decrem:2004:DLW

[Dec04] Bart Decrem. Desktop Linux: Where art thou? ACM Queue: Tomorrow’s Computing Today, 2(3):48–56, May 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Dedeke:2009:LOL

[Ded09] Adenekan (Nick) Dedeke. Loyal opposition: Is Linux better than Windows software? IEEE Software, 26(3):104, 103, May/ June 2009. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Densmore:1979:IBS

[Den79] W. Densmore. An ICE-85 based 8085 software development system. Technical Memorandum 1417 (TM 79-3231-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 31, 1979. ?? pp.

Denny:1983:DUE

[Den83] Michael Denny. Delivering UNIX to the end-user market. In Association [Ass83], pages 311–312. Abstract and summary.

Deno:1999:NRB

[Den99] Cynthia Deno. New releases of *BSD and Debian Linux OSes given away at USENIX Annual Conference. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1999-6/newreleases.
html.

Dent:2004:PDG

[Den04] Kyle D. Dent. Postfix: the definitive guide: a secure and easy-to-use MTA for Unix. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00212-2. xv + 260 pp. LCCN TK5105.74.P66 D46 2004. US$. URL http:/

/www.oreilly.com/catalog/
postfix/.

Dept:1979:UTU

[Dep79a] Department 3624. UNIX/TS user’s manual — edition 1.1. Technical Memorandum 1440, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Dept:1979:PUU

[Dep79b] Department 3782. PWB/UNIX user’s manual — release 2.0. Technical Memorandum 1407, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1979. ?? pp.

Derr:1990:FRP

[Der90]  M. A. Derr.  Formatting

references with prefer. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

                                        1990. URL http://doc.cat-
                                        v.org/unix/v10/10thEdMan/
                                        v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

DeRoest:1997:AVS

[DeR97] James W. DeRoest. AIX version 4: system and administration guide. McGraw-Hill, New York, NY, USA, 1997. ISBN 0-07-036688-8. xxv + 591 pp. LCCN QA76.76.O63D4725 1997. US$44.95.

DeRoest:1999:SUN

[DeR99] James DeRoest. Samba: Unix and NT internetworking. Mc-Graw-Hill, New York, NY, USA, 1999. ISBN 0-07-135104-3. xv + 303 pp. LCCN QA76.76.O63 D4727 1999.

DeRoest:2000:SUN

[DeR00] James W. DeRoest. Samba: Unix and NT networking. Mc-Graw-Hill, New York, NY, USA, 2000. ISBN 0-07-135104-3. xv + 303 pp. LCCN QA76.76.O63 D4727 2000 Bar.

Desfray:1996:AOD

[Des96] Philippe Desfray. Automated object design: The client-server case. Computer, 29(2):62–66, February 1996. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeVaney:1993:UDL

[DeV93]  Mike DeVaney.  UNIX is

dead; long live UNIX. ;login: the USENIX Association newsletter, 18(1):23–24, January/February 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

DeVitt:2000:SPN

[DeV00]  Don DeVitt.  Solaris PC

Netlink: Performance, Sizing, and Deployment. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-026686-8. xxxiv + 444 pp. LCCN QA76.8.S86D48

2000. URL books/pcnlbody.
pdf;   http://www.phptr.
com/ptrbooks/ptr_0130266868.
html;   http://www.sun.com/
books/catalog/devitt/.

Graaf:1979:PPE

[DF79] D. A. De Graaf and Jerome Feder. Preliminary performance evaluation of Univac UNIX. Technical Memorandum 1487, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 4, 1979. ?? pp.

Dougherty:1991:SAO

[DF91] Dale Dougherty and Toshihiro Fukuzaki. Sedo ando oku puroguramingu: UNIX power tools (Japanese) [sed and awk programming: UNIX power tools]. Asuki, Tokyo, Japan, 1991. ISBN 4-7561-0091-0. 508 pp. LCCN ????

DeHart:1993:USS

[DF93] Edward DeHart and Barbara Fraser. 1992 UNIX security symposium report. ;login: the USENIX Association newsletter, 18(1):4–??, January/February 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Dunstan:1995:PSU

[DF95] Neil Dunstan and Ivan Fris. Process scheduling and UNIX semaphores. Software—Practice and Experience, 25(10): 1141–1153, October 1995. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Dutt:2000:GBG

[DF00] Christoph Dutt and Joachim Freiburg. GiMP: Bilder gestalten, Fotos retuschieren; [Grundlagen der professionellen Bildbearbeitung, der Umgang mit Fotos, Grafiken und Text, Bilder f?r das Internet richtig vorbereiten; auf der CD: GIMP f?r Windows, SCO Unix, Debian GNU Linux, Solaris, OS/2 und BSD, Quelltext aller GIMP- und GTK-Versionen, ?ber 300 Plug-ins in C, Perl, tcl, Phyton und Scheme, XFree86/23.3.6, GIMP User Manual als PDF-Dateien]. C und L, Böblingen, Germany, 2000. ISBN 3-932311-64-7. 522 + 98 pp. LCCN ???? Includes CD-ROM.

deFeraudy:1990:CUT

[dFAAW90] Henri de Feraudy, Robert Artigas, Jr., Arkin Asaf, and Martin D. Winnick. CUG329 UNIX tools for PC. C Users Journal, 8(11):126–??, November 1990. ISSN 0898-9788.

Duan:2023:DLS

[DFC^{+}23] Guoyun Duan, Yuanzhi Fu, Minjie Cai, Hao Chen, and Jianhua Sun. DongTing: a large-scale dataset for anomaly detection of the Linux kernel. The Journal of Systems and Software, 203(??):??, September 2023. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0164121223001401.

Delis:1999:PAS

[DFG99] Alex Delis, Christos Faloutsos, and Shahram Ghandeharizadeh, editors. Proceedings of the 1999 ACM SIG-MOD International Conference on Management of Data: SIG-MOD ’99, Philadelphia, PA, USA, June 1–3, 1999, volume 28(2) of SIGMOD Record (ACM Special Interest Group on Management of Data). ACM Press, New York, NY, USA, 1999. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87.

Dannenberg:1991:FFL

[DFV91] Roger B. Dannenberg, Christopher Lee Fraley, and Peter Velikonja. Fugue: a functional language for sound synthesis. Computer, 24(7):36–42, July 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

daCruz:1993:UCK

[dG93] Frank da Cruz and Christine M. Gianone. Using C-Kermit. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Upper Saddle River, NJ 07458, USA, 1993. ISBN 1-55558-108-0 (DP), 0-13-037490-3 (PH). xxi + 514 pp. LCCN TK5105.9.D33 1993. Available in German translation [dG96].

daCruz:1996:CKE

[dG96] Frank da Cruz and Christine M. Gianone. C-Kermit: Einführung und Referenz. Verlag Heinz Heise, Helsdorfer Straße 7, D-30625, Hannover, Germany, 1996. ISBN 3-88229-023-4. 557 pp. LCCN ???? 88,00 DM. URL

                                        http://www.emedia.de/bin/
                                        bookshop?show=3401&id=. Ger-

man translation of [dG93].

Dent:2000:GUU

[DG00] Jack Dent and Tony Gaddis. Guide to UNIX: using Linux. Course Technology, Cambridge, MA, USA, 2000. ISBN 0-7600-1096-X. xiv + 552 pp. LCCN QA76.76.O63 D4 2000.

Dolotta:1977:PWS

[DH77] T. A. Dolotta and R. C.

Haight. Programmer’s workbench — synopsis of facilities (g.1). Technical Memorandum 1109, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 15, 1977. ?? pp.

Deep:1996:DCA

[DH96] John Deep and Peter Holfelder.

Developing CGI applications with Perl. Wiley, New York, NY, USA, 1996. ISBN 0-471-14158-5. xiv + 299 pp. LCCN TK5105.888 .D45 1996.

Dalheimer:2002:EDQ

[DH02] Matthias Kalle Dalheimer and Steffen Hansen. Embedded development with Qt/ Embedded: Linux for small systems. Dr. Dobb’s Journal of Software Tools, 27(3):48, 50, 52–54, March 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_03/qtembed.
txt;   http://www.ddj.com/
ftp/2002/2002_03/qtembed.
zip.

Dalrymple:2005:AMX

[DH05] Mark Dalrymple and Aaron Hillegass. Advanced Mac OS X Programming. Big Nerd Ranch, Atlanta, GA, USA, second edition, 2005. ISBN 0-9740785-1-

4. 646 (est.) pp. LCCN ????

Dolotta:1978:UTS

[DHM78] T. A. Dolotta, R. C. Haight, and John R. Mashey. UNIX time-sharing system: The Programmer’s Workbench. The Bell System Technical Journal, 57(6):2177–2200, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2177.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2177.pdf.

Dolotta:1977:DPU

[DHP77] T. A. Dolotta, R. C. Haight, and E. M. Piskorik. Documents for the PWB/UNIX Time-Sharing System edition 1.0. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977. 484 pp. URL https:

//bitsavers.org/pdf/att/
unix/PWB_UNIX/Documents_
for_the_PWB_UNIX_Time-Sharing_
System_Edition_1.0_197710.
pdf.

Dorris:1978:CAC

[DHS78] H. N. Dorris, D. W. Hagelbarger, and M. J. Shannon. Computer assisted communication system. Technical Memorandum ?? (TM 78-1356-7, TM 78-1352-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1978. ?? pp.

Devarakonda:1989:PPR

[DI89] M. V. Devarakonda and R. K. Iyer. Predictability of process resource usage: a measurement-based study on UNIX. IEEE Transactions on Software Engineering, 15(12):1579–1586, December 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=58769.

Diaz:1994:PNG

[Dia94] Bernard M. Diaz. Place for novelty in graphics and visualisation education. Computers and Graphics, 18(3):281–286, May– June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Dietze:2011:PUS

[Die11]  Martin Dietze.  Praxiskurs

Unix-Shell: [werden Sie zum Kommandozeilenvirtuosen, behandelt Bourne-Shell, Korn-Shell, bash and TC-Shell, mit Aufgaben und Lösungen]. O’Reilly basics. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2011. ISBN 3-89721-565-9. xxiii + 282 pp. LCCN ???? EUR 19.90.

Dike:2006:UML

[Dik06] Jeff Dike. User Mode Linux. Bruce Perens Open Source series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-186505-6 (paperback). ???? pp. LCCN QA76.76.O63 D545 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip068/2006004225.html.

Dilworth:2002:UEP

[Dil02] Bob Dilworth. Using email to perform UNIX system monitoring and control. Sys Admin: The Journal for UNIX Systems Administrators, 11(5): 31–32, 34–36, May 2002. CO-DEN SYADE7. ISSN 1061-2688.

DiMaggio:2000:TUD

[DiM00] Len DiMaggio. Testing UNIX daemons. Dr. Dobb’s Journal of Software Tools, 25(3):38, 40, 42, 44, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Dodge:2005:SIL

[DIN05] Catherine Dodge, Cynthia Irvine, and Thuy Nguyen. A study of initialization in Linux and OpenBSD. Operating Systems Review, 39(2):79–93, April 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Dittmer:1993:EUC

[Dit93] Ingo Dittmer. Error in Unix commands dc and bc for multiple-precision-arithmetic.

ACM SIGNUM Newsletter, 28 (2):8–11, April 1993. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dixon:1982:UMT

[Dix82] Jack Dixon. UNIX and manufacturing testing. In Usr Group [Usr82], pages 177–?? Abstract only.

DeBenedictis:1993:EUS

[DJ93] Erik P. DeBenedictis and Stephen C. Johnson. Extending Unix for scalable computing. Computer, 26(11):43–53, November 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Danesh:2001:ML

[DJ01] Arman Danesh and Michael H.

Jang. Mastering Linux. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 2001. ISBN 0-7821-2915-3. xxxv + 955 pp. LCCN QA76.76.O63 D3428 2001.

Dudek:1986:HMF

[DJM86] Gregory Dudek, Michael Jenkin, and Howard Marcus. How to make friends with numbercrunchers: adding single-user array-processor slave environments to VAX UNIX. In USENIX Association [USE86c], pages 200–208.

Perez-Davila:1984:PIF

[dJPDD84] Alfredo de J. Perez-Davila and Lawrence W. Dowdy. Parameter interdependencies of file placement models in a Unix system. ACM SIGMET-RICS Performance Evaluation Review, 12(3):15–26, August 1984. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Donovan:2016:GPL

[DK16] Alan A. A. Donovan and Brian W. Kernighan. The Go Programming Language. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, 2016. ISBN 0-13-419044-0 (paperback). xvii + 380 pp. LCCN QA76.73.G63 D66 2016.

URL https://www.pearson.
com/us/higher-education/
program/Donovan-Go-Programming-
Language-The/PGM234922.html.

Dougherty:1994:MHX

[DKF94] Dale Dougherty, Richard Koman, and Paula Ferguson. The Mosaic Handbook for the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1994. ISBN 1-56592-095-3. xxiii + 262 pp. LCCN TK5105.875.I57 D687 1994. US$29.95.

Devarakonda:1996:RCF

[DKM96] Murthy Devarakonda, Bill Kish, and Ajay Mohindra. Recovery in the Calypso file system. ACM Transactions on Computer Systems, 14(3):287– 310, August 1996. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1996-14-3/p287-devarakonda/

.

Dolstra:2008:NPF

[DL08] Eelco Dolstra and Andres Löh.

NixOS: a purely functional Linux distribution. ACM SIG-PLAN Notices, 43(9):367–378, September 2008. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DOrazio:2017:MAM

[DLCV17] Christian J. D’Orazio, Rongxing Lu, Kim-Kwang Raymond Choo, and Athanasios V. Vasilakos. A Markov adversary model to detect vulnerable iOS devices and vulnerabilities in iOS apps. Applied Mathematics and Computation, 293(??):523–544, January 15, 2017. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0096300316305549.

Dolstra:2010:NPF

[DLP10] Eelco Dolstra, Andres Löh, and Nicolas Pierron. NixOS: A purely functional Linux distribution. Journal of Functional Programming, 20(5– 6):577–615, November 2010. CODEN JFPRES. ISSN 0956-7968 (print), 1469-7653 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        product/C1ACBA2A51D2E5466820F5B5086EA2CE.

Dotzauer:2007:SBL

[DLvG07] Timo Dotzauer, Tobias Lütticke, and Alexander von Gernler. Das SSH-Buch: Leitfaden für den sicheren Einsatz von OpenSSH. Millin, Lohmar, Germany, 2007. ISBN 3-938626-03-8. xviii + 602 pp. LCCN ???? URL http:

                                        //www.gbv.de/dms/ilmenau/
                                        toc/508392802.PDF.

Dolotta:1979:UCL

[DM79] T. A. Dolotta and John R. Mashey. Using a command language as the primary programming tool. Technical Memorandum 1452 (TM 79-3624-2, TM 79-9322-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1979. ?? pp.

Davida:1988:UGD

[DM88] George I. Davida and Brian J. Matt. UNIX guardians: Delegating security to the user. In USENIX Association [USE88e], pages 14–

23. LCCN QA76.8.U65 U55

1988(1)-1990(2)//.

Do:1997:LEU

[DM97] James Do and Muhammed Mudawwar. Letters to the editor: Unicode misunderstood. Computer, 30(6):6, 9, June 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Response and rebuttal to [Mud97].

Dawidek:2007:PSZ

[DM07] Paweł Jakub Dawidek and Marshall Kirk McKusick. Porting the Solaris ZFS file system to the FreeBSD operating system. ;login: the USENIX Association newsletter, 32(3):19–24, June 2007. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic). URL https:/

/papers.freebsd.org/2007/
dawidek-porting_the_solaris_
zfs_file_system_to_the_freebsd_
operating_system/.

Dreiheller:1986:PPP

[DMM86] A. Dreiheller, M. Moerschbacher, and B. Mohr. PHYSCAL: Programming Pascal with physical units. ACM SIGPLAN Notices, 21(12):114– 123, December 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dunigan:2002:TTD

[DMT02] Tom Dunigan, Matt Mathis, and Brian Tierney. A TCP tuning daemon. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.
org/paperpdfs/pap.pap151.
pdf.

Draskoy:1987:XMU

[DN87] Andrew Draskoy and Gerald Neufeld. X.400 messaging on UNIX. In USENIX Association [USE87d], pages 111–115.

Dickey:1983:AIU

[DNQ^{+}83] Matt Dickey, Greg Noel, Bob Querido, Bill Appelbe, and Jim McGinness. Architectural implications of UNIX (or pitfalls for UNIX porters!). In Association [Ass83], pages 307–?? Abstract only.

Dougherty:1987:UTP

[DO87] Dale Dougherty and Tim O’Reilly, editors. UNIX Text Processing. Hayden Books UNIX library system. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1987. ISBN 0-672-46291-5 (paperback). xiv + 665 pp. LCCN Z286.D47 D68 1987. US$26.95. URL ftp:

                                        //ftp.ffii.org/pub/groff/
                                        contrib/documentation/utp/
                                        ;   http://books.google.
                                        com/books?id=H148AQAAIAAJ;
                                         http://catalog.hathitrust.
                                        org/api/volumes/oclc/15722288.
                                        html;   http://proquest.
                                        safaribooksonline.com/?fpi=
                                        9780810462915;   http://
                                        www.oreilly.com/openbook/
                                        utp/UnixTextProcessing.pdf
                                         http://proquest.tech.safaribooksonline.
                                        de/9780810462915.

Dougherty:1988:MU

[DO88] D. Dougherty and T. O’Reilly. DOS meets Unix. BYTE Magazine, 13(11):117–126, November 1988. CODEN BYTEDJ. ISSN 0360-5280.

Docter:2003:SSC

[Doc03] Quentin Docter. Solaris 9:

Sun Certified System Administrator Study Guide. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4181-1. xxxix + 597 pp. LCCN QA76.3 D635 2003.

deOliveira:2023:OSN

[dOCC23] Daniel Bristot de Oliveira, Daniel Casini, and Tommaso Cucinotta. Operating system noise in the Linux kernel. IEEE Transactions on Computers, 72(1):196–207, January 2023. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

Doeppner:2011:OSD

[Doe11] Thomas W. Doeppner. Operating Systems in Depth. Wiley, New York, NY, USA, 2011. ISBN 0-471-68723-5 (hardcover). xvi + 444 pp. LCCN QA76.76.O63 D64

2011. URL http://www.loc.
gov/catdir/toc/fy11pdf04/
2010034669.html.

Donahue:1985:IMC

[Don85] James Donahue. Integration mechanisms in Cedar. ACM SIGPLAN Notices, 20(7):245– 251, July 1985. CODEN SIN-ODQ. ISBN 0-89791-165-2. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donkers:1995:UWI

[Don95] Arthur Donkers. A UNIX/ Windows 95 implementation.

Sys Admin: The Journal for UNIX Systems Administrators, 4(5):27–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:AST

[Don96a] Arthur Donkers. Assorted security tips for UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 5(11): 48, 50, 52, 54–58, November 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:BLF

[Don96b] Arthur Donkers. Building a Linux firewall. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):12–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1996:LIS

[Don96c] Arthur Donkers. Linux as an Internet server. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):52– ??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:GUN

[Don97a] Arthur Donkers. Getting UNIX and NT on speaking terms. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):8–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:IL

[Don97b] Arthur Donkers. ISDN and Linux. Sys Admin: The Journal for UNIX Systems Administrators, 6(10):51–53, 55, 57– 60, October 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1997:WLD

[Don97c] Arthur Donkers. Writing Linux device drivers. Sys Admin: The Journal for UNIX Systems Administrators, 6(1):8–??, January 1997. CODEN SYADE7. ISSN 1061-2688.

Donkers:1998:LNS

[Don98a] Arthur Donkers. Looking at NT security through UNIX glasses. Sys Admin: The Journal for UNIX Systems Administrators, 7(9):NT–17–NT– 20, September 1998. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Donkers:1998:ULO

[Don98b] Arthur Donkers. Using Linux in an office environment. Sys Admin: The Journal for UNIX Systems Administrators, 7(1): 45–46, 48, 50, 51, January 1998. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Donar:2002:TUO

[Don02] Tim Donar. Tru64 UNIX– Oracle9i Cluster quick reference. Compact solutions. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-272-9 (paperback). xiv + 298 pp. LCCN QA76.76.D63 D66 2002.

Donato:2004:SPS

[Don04] Alberto Donato. A software platform to support dynamically reconfigurable systems-on-chip under the GNU/Linux operating system. Politecnico, Milano, Italy, 2004. ISBN ???? xvi + 95 pp. LCCN ????

Donenfeld:2022:RNG

[Don22] Jason A. Donenfeld. Random number generator enhancements for Linux 5.17 and 5.18. Web document, March

                                        18, 2022. URL https://www.
                                        zx2c4.com/projects/linux-
                                        rng-5.17-5.18/.

Dougherty:1991:SA

[Dou91] Dale Dougherty. sed & awk. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-59-5. xxii + 394 pp. LCCN QA76.76.U84 D69 1991.

                                        URL http://www.oreilly.
                                        com/catalog/9780937175590.

Dougan:2002:TAR

[Dou02] Cort Dougan. Two-axis, realtime camera control. Dr. Dobb’s Journal of Software Tools, 27(10):60, 62, 64– 67, October 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        rtlcam.txt;   http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        rtlcam.zip.

Dowd:1977:DLP

[Dow77] P. G. Dowd. DATASPEED 40 line printer as a UNIX peripheral. Technical Memorandum 1255 (TM 77-8234-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 13, 1977. ?? pp.

Dowd:1993:HPC

[Dow93] Kevin Dowd. High Performance Computing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-032-5. xxv + 371 pp. LCCN QA76.88 .D6 1993. US$25.95.

Dichter:1995:SEP

[DP95] Carl Dichter and Mark Pease.

Software Engineering with Perl. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-016965-X. 282 pp. LCCN QA76.758 .D53 1995. US$30.00. Includes disk.

Dougherty:1997:SA

[DR97] Dale Dougherty and Arnold Robbins. sed and awk. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, February 1997. ISBN 1-56592-225-5. xxii + 407 pp. LCCN QA76.76.U84 D69 1997. US$29.95. URL http://www.

oreilly.com/catalog/sed2.

deRaadt:2015:O

[dR15] Theo de Raadt. OpenBSD 5.7.

????, ????, 2015. ISBN 0-9881561-5-6. ???? pp. LCCN

????  URL http://www.
openbsd.org/57.html.

Domingos:2023:SRV

[DRN^{+}23] Joao Mario Domingos, Tiago Rocha, Nuno Neves, Nuno Roma, Pedro Tomás, and Leonel Sousa. Supporting RISC-V performance counters through Linux performance analysis tools. In IEEE, editor, 2023 IEEE 34th International Conference on Applicationspecific Systems, Architectures and Processors (ASAP): ASAP 2023, 19–21 July 2023, Porto, Portugal. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, July 2023.

Danesh:2001:SLG

[DRP01] Arman Danesh, James Russell, and Richard Petersen. SAIR Linux & GNU certified administrator: exam guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213205-1. xxv + 1008 pp. LCCN QA76.76.O63 D342855 2001. URL ftp:

//uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/description/
mh022/2002277480.html;
http://www.loc.gov/catdir/
toc/mh021/2002277480.html.

Dolotta:1977:PUV

[DS77] T. A. Dolotta and D. W. Smith. PWB/UNIX view graph and slide macros (T.9). Technical Memorandum 1192, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1977. ?? pp.

Donnelly:1988:BYC

[DS88] Charles Donnelly and Richard M.

Stallman. BISON — the YACC-compatible parser generator. Technical report, Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1988. Bison was largely written by Robert Corbett, and made yacc-compatible by Richard Stallman. Electronic

mail: rms@prep.ai.mit.edu.

Software also available via ANONYMOUS FTP to prep. ai.mit.edu. See also [Pax88].

Dunham:2000:RSD

[DS00] Mike Dunham and Ed Schaefer. Recovering source from a defunct RCS system. Sys Admin: The Journal for UNIX Systems Administrators, 9(4): 55–58, April 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

dosSantos:2002:MAS

[dSVdLT02] Nelson dos Santos, Flávio Miguel Varejão, and Orivaldo de Lira Tavares. Multi-agent systems and network management — a positive experience on Unix environments. Lecture Notes in Computer Science, 2527:616– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2527/25270616.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2527/25270616.pdf.

DuBois:1993:SPT

[DuB93] Paul DuBois. Software Portability with imake. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-055-4. 390 pp. US$27.95.

DuBois:1995:UCT

[DuB95]  Paul DuBois.  Using csh

and tcsh. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1995. ISBN 1-56592-132-1. xviii + 221 pp. LCCN QA76.76.O63D82 1995. US$29.95. URL http://www.

                                        oreilly.com/catalog/tcsh.

DuBois:2001:MPW

[DuB01] Paul DuBois. Mysql and perl for the Web. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1054-6. xvi + 520 pp. LCCN QA76.73.S67 D585 2002. Edited by Stephanie Wall, Gina Brown, and Chris Zahn.

DuCharme:1994:OSH

[DuC94] Bob DuCharme. The operating systems handbook: UNIX, OpenVMS, OS/400, VM and MVS. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-017891-7. xvii + 390 pp.

LCCN QA76.76.O63 D83 1994. US$49.50.

Duffy:1982:IUB

[Duf82] Michael E. Duffy. Interfacing UNIX to backend database machines. In Usr Group [Usr82], pages 19–28.

Duff:1989:EVU

[Duf89a] Tom Duff. Experience with viruses on UNIX systems. In Association [Ass89a], pages 155–171.

Duff:COMPSYS-2-2-155

[Duf89b] Tom Duff. Experience with viruses on UNIX systems. Computing Systems, 2(2):155– 171, Spring 1989. CODEN CM-SYE2. ISSN 0895-6340.

Duff:1989:VAU

[Duf89c] Tom Duff. Viral attacks on UNIX system security. In USENIX Association [USE89f], pages 165–171.

Duff:1990:ERG

[Duf90a] T. Duff. The 10th Edition raster graphics system. In Hume and McIlroy [HM90b], pages 483–495. ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/graphics.pdf;
http://doc.cat-v.org/unix/
v10/10thEdMan/v2contents.

html. Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Duff:1990:RSP

[Duf90b]  T. Duff.  Rc — a shell

for Plan 9 and UNIX. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Duff:19xx:RSP

[Dufxx] Tom Duff. Rc — a shell for Plan 9 and UNIX systems. Computing Science Technical Report 158g, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 14 pp.

Dulaney:1994:SUN

[Dul94] Emmett Dulaney. Standard UNIX network diagnostic tools. Sys Admin: The Journal for UNIX Systems Administrators, 3(5):53–??, September/October 1994. CODEN SYADE7. ISSN 1061-2688.

Dulaney:1996:HPB

[Dul96] Emmett Dulaney. Handling permissions between UNIX and NetWare. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):69–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Dulaney:2006:LSK

[Dul06]  Emmett Dulaney.  Linux

Starter Kit. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32887-9. ???? pp. LCCN ???? US$31.99.

Dulaney:2014:LAO

[Dul14] Emmett A. Dulaney. Linux allin-one for dummies. For dummies. Wiley, New York, NY, USA, fifth edition, 2014. ISBN 1-118-84435-1. xviii + 550 pp. LCCN QA76.76.O63 D8557

2014. URL http://www.loc.
gov/catdir/enhancements/
fy1506/2013958402-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1506/2013958402-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1506/2013958402-t.html.

Dumbill:2005:DGL

[Dum05] Edd Dumbill. Debian GNU/ Linux: An Explorer’s Notebook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00883-X. 256 (est.) pp. LCCN ???? EUR 24.00. URL http:/

/www.oreilly.com/catalog/
9780596008833.

Dunstan:1989:SPU

[Dun89] N. Dunstan. Synchronization problems and UNIX System V. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 21 (4):15–19, December 1989. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dunphy:1991:UIE

[Dun91a] Ed Dunphy. The UNIX Industry — Evolution, Concepts, Architecture, Applications, and Standards. QED Information Sciences, Inc., P. O. Box 82-181, Wellesley, MA 02181, USA, 1991. ISBN 0-89435-390-X. xxii + 338 pp. LCCN QA76.76.O63 D8598 1991. Covering UNIX technology and the structure of the UNIX marketplace.

Dunphy:1991:UIO

[Dun91b] Ed Dunphy. The UNIX Industry and Open Systems in Transition: a Guidebook for Managing Change. Wiley-Intersci-ence, New York, NY, USA, second edition, 1991. ISBN 0-471-60608-1. xxx + 616 pp. LCCN QA76.76.O63 D8599 1994. Covering UNIX technology and the structure of the UNIX marketplace.

Dunstan:1991:BMU

[Dun91c] Neil Dunstan. Building monitors with UNIX and C. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 23 (3):7–9, September 1991. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Dunham:1992:CTU

[Dun92] Alan Dunham. Crash tracebacks in UNIX. Dr. Dobb’s Journal of Software Tools, 17 (9):80, 82, 84, 86, 89, 113, September 1992. CODEN DDJOEB. ISSN 1044-789X.

Duntemann:2000:ALS

[Dun00] Jeff Duntemann. Assembly language step-by-step: programming with DOS and Linux. Wiley, New York, NY, USA, second edition, 2000. ISBN 0-471-37523-3. xxv + 613 pp. LCCN QA76.73.A8 D88 2000. Includes CD-ROM.

Dupuy:2005:CBD

[Dup05] Stéphane Dupuy. Conception d’une base de données PostgreSQL sous Linux. Mémoire de diplôme d’ingénieur, Centre d’enseignement, Université de soutenance, Bordeaux, France, 2005. 99 pp.

deBakker:1993:CSL

[dVd93] J. W. de Bakker, F. Van Breugel, and A. de Bruin. Comparative semantics for linear arrays of communicating processes, a study of the UNIX fork and pipe commands. Lecture Notes in Computer Science, 711:252–??, 1993. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dalheimer:2006:RL

[DW06] Matthias Kalle Dalheimer and Matt Welsh. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2006. ISBN 0-596-00760-4 (paperback). xviii + 951 pp. LCCN QA76.76.O63 R855

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007270110-d.html;
 http://www.oreilly.com/
catalog/9780596007607.

Dwivedi:2004:ISS

[Dwi04] Himanshu Dwivedi. Implementing SSH: strategies for optimizing the secure shell. Wiley, New York, NY, USA, 2004. ISBN 0-471-45880-5. xxvi + 376 pp. LCCN QA76.76.O63 D895 2004. UK£24.50. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
bios/wiley046/2004297174.
html;   http://www.loc.gov/
catdir/description/wiley041/
2004297174.html;   http:/
/www.loc.gov/catdir/toc/
wiley041/2004297174.html.

Davies:2006:SLB

[DWV06] Justin Davies, Roger Whittaker, and William Von Hagen. SUSE Linux 10 bible. Wiley, New York, NY, USA, 2006. ISBN 0-471-75488-9 (paperback/DVD). xlvii + 852 pp. LCCN QA76.76.O63 D34992 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0519/2005027885.html.

Dwyer:1979:SPD

[Dwy79] T. J. Dwyer. A subroutine package to do process cleanup. Technical Memorandum 1494 (TM 79-9457-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 14, 1979. ?? pp.

Dyson:1996:UDR

[Dys96] Peter John Dyson. The UNIX desk reference: the hu.man pages. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1996. ISBN 0-7821-1658-2. xi + 523 pp. LCCN QA76.76.O63 D99 1996. US$29.99.

Duan:2020:RFK

[DYZC20] Junhan Duan, Yudi Yang, Jie Zhou, and John Criswell. Refactoring the FreeBSD kernel with Checked C. Report ??, Department of Computer Science, University of Rochester, Rochester, NY, USA, July 10, 2020. 8

pp. URL https://www.cs.
rochester.edu/u/jzhou41/
papers/freebsd_checkedc.
pdf.

Dzonsons:2009:FSL

[Dzo09] Kristaps Dzonsons. Fixing on a standard language for UNIX manuals. ;login: the USENIX Association newsletter, 34(5): ??, October 2009. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL https://www.
usenix.org/publications/
login/october-2009-volume-
34-number-5/fixing-standard-
language-unix-manuals.

Earnshaw:1987:WPS

[Ear87] Rae A. Earnshaw, editor.

Workstations and Publication Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. ISBN 0-387-96527-0, 3-540-96527-0. viii + 229 pp. LCCN Z286.E43 W67 1987. US$29.50.

Edge:2008:FMX

[EBS08] Charles S. Edge, William Barker, and Zack Smith. Foundations of Mac OS X Leopard security: [Mac OS X client and server security, from the home to the enterprise]. The expert’s voice in Mac OS X. Apress, Berkeley, CA, USA, 2008. ISBN 1-59059-989-6 (paperback). xxix + 455 pp. LCCN ????

Engstrom:1989:SPS

[EC89] Bradley R. Engstrom and Peter R. Cappello. The SDEF programming system. Journal of Parallel and Distributed Computing, 7(2):201–231, October 1989. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Eckstein:2000:US

[ECBK00] Robert Eckstein, David Collier-Brown, and Peter Kelly. Using Samba. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-449-5. xi + 398 pp. LCCN QA76.9.C55 E267 2000; QA76.9.C55 E267 2000; TK5105.5 .E35 2000. US$34.95.

URL http://www.oreilly.
com/catalog/9781565924499.

Includes CD-ROM with complete mirror of Samba site with Samba 2.0.5a, including sources, documentation, binaries, and supplementary tools.

Eck:1991:SRS

[Eck91]  Christoph Eck.  Standard-

ization of realtime software POSIX 1003.4. In IEEE [IEE91b], pages 149–152. ISBN 0-7803-0458-6, 0-7803-0459-4. LCCN QA76.54.C66 1991. IEEE catalog number 92TH0404-4.

Eckel:1993:CI

[Eck93] Bruce Eckel. C++ Inside & Out. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-881809-5. xxviii + 754 pp. LCCN QA76.73.C153 E24 1993.

Eckel:1995:BUI

[Eck95] George Eckel. Building a UNIX Internet server. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-494-5. xvii + 325 pp. LCCN QA76.76.O63 E235 1995.

Eckert:2011:LGL

[Eck11] Jason W. Eckert. Linux+ guide to Linux certification. Cengage Learning - Delmar, Clifton Park, NY, USA, third edition, 2011. ISBN 1-4188-3721-0, 1-111-54153-1. ???? pp. LCCN ????

Eddy:1990:HRD

[Edd90] William F. Eddy. Hardware review: The DECstation 3100 — UNIX for power users. Chance, 3(2):42–47, Spring 1990. CO-DEN CNDCE4. ISSN 0933-2480 (print), 1867-2280 (electronic).

Edelson:1977:TAA

[Ede77] D. Edelson. Typesetting ACS and APS meeting abstracts — issue 2. Technical Memorandum 1075 (TM 77-1351-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 3, 1977. ?? pp.

El-Dirghami:2000:SLI

[EDK00] Nazeeh Amin El-Dirghami and Youssef A. Abu Kwaik. SuSE Linux Installation and Configuration Handbook. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2355-7. xx + 756 pp. LCCN QA76.76.O63 E4 2000. US$39.99. Includes CD-ROM.

Edwards:1998:ITC

[Edw98]  John Edwards.  Industry

trends: The changing face of freeware. Computer, 31(10): 11–13, October 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1998/pdf/
rx011.pdf.

Edmundsson:2004:DET

[EEK^{+}04] Niklas Edmundsson, Erik Elmroth, Bo Kågström, Markus Mårtensson, Mats Nylén, Åke Sandgren, and Mattias Wadenstein. Design and evaluation of a TOP100 Linux Super Cluster system. Concurrency and Computation: Practice and Experience, 16(8):735–750, July 2004. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Eilert:2003:LM

[EEMS03] John Eilert, Maria Eisenhaendler, Dorothea Matthaeu, and Ingol Salm, editors. Linux on the mainframe. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-101415-3 (paperback). xxviii + 430 pp. LCCN QA76.76.O63 L54553 2003.

Edler:1986:CMP

[EGL86] Jan Edler, Allan Gottlieb, and Jim Lipkis. Considerations for massively parallel UNIX systems on the NYU Ultracomputer and IBM RP3. In USENIX Association [USE86e], pages 193–210. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Eisen:1979:SAI

[Eis79a] S. R. Eisen. A stand-alone input/output library. Technical Memorandum 1475 (TM 79-3624-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 4, 1979. ?? pp.

Eisen:1979:UUG

[Eis79b] S. R. Eisen. UNIX Users Group meeting — Toronto, June 9–23, 1979. Technical Memorandum 1396, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 25, 1979. ?? pp.

Lozy:1985:EUE

[eL85] Mohamed el Lozy. Editing in a UNIX environment — the vi/ ex editor. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-235599-X. xiv + 226 pp. LCCN QA76.6 .L69 1985. US$18.95.

Elahian:1983:NUM

[Ela83] Camran Elahian. New UNIX markets in engineering. In Association [Ass83], pages 313–?? Abstract only.

Elboth:2001:LB

[Elb01] David Elboth. The Linux Book. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-032765-4. xxii + 501 pp. LCCN QA76.76.O63 E42 2001. US$34.99.

Elliott:1975:TBS

[Ell75] R. J. Elliott. A tape backup system for the network file store. Technical Memorandum 1081 (TM 75-1273-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 24, 1975. ?? pp.

Elliott:1977:SSS

[Ell77] R. J. Elliott. Software support for satellite processor. Technical Memorandum 1218 (TM 77-1273-16), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 7, 1977. ?? pp.

Ellington:1979:CSF

[Ell79a] M. B. Ellington. Conversion of SICPPR files from UNIX to RAMIS II. Technical Memorandum 1399 (TM 79-2174-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 21, 1979. ?? pp.

Elliott:1979:DNM

[Ell79b] R. J. Elliott. A DATAKIT network monitoring system. Technical Memorandum 1460 (TM 79-1273-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 2, 1979. ?? pp.

Ellis:1980:LS

[Ell80] John R. Ellis. A LISP shell.

ACM SIGPLAN Notices, 15 (5):24–34, May 1980. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edler:1988:PMH

[ELS88] Jan Edler, Jim Lipkis, and Edith Schonberg. Process management for highly parallel UNIX systems. In USENIX Association [USE88f], pages 1–17. ISBN ???? LCCN ????

Evenson:1979:CAS

[EM79] E. K. Evenson and N. March.

Craft Access system software standards. Technical Memorandum 1366 (TM 79-4563-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 8, 1979. ?? pp.

Eranian:2000:LIP

[EM00] Stéphane Eranian and David Mosberger. The Linux/ ia64 project: Kernel design and status report. Technical Report HPL-2000-85, Hewlett–Packard Corporation, Palo Alto, CA, USA, 2000.

URL  http://www.hpl.hp.
com/techreports/.

Emmons:2011:LOD

[Emm11] Jon Emmons. Linux for the Oracle DBA: the definitive reference, volume 40 of Oracle infocus series. Rampant Tech-Press, Kittrell, NC, USA, 2011. ISBN 0-9823061-9-9. ???? pp. LCCN ????

Endler:2000:CDH

[End00] Dave Endler. Creating and deploying a honey pot. Sys Admin: The Journal for UNIX Systems Administrators, 9(1): 8, 10, 12, 14, 16, 18, 20–22, 24, January 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Engelbert:1976:STS

[Eng76] E. F. Engelbert. A secretarial typing system using UNIX. Technical Memorandum 1091, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 27, 1976. ?? pp.

Engquist:1991:DUB

[Eng91] Jim Engquist. A database for UNIX backup. In USENIX [USE91b], pages 89–96.

ESRI:1997:AML

[Env97a] Environmental Systems Research Institute (Redlands, CA). ARC macro language: developing ARC/INFO menus and macros with AML: selfstudy workbook: version 7.1.1 for UNIX and Windows NT. Wiley, New York, NY, USA, second edition, 1997. ISBN 1-86242-044-0. various pp. LCCN G70.212 .A73 1997. CD-ROM includes sample programs, data and exercise solutions for the workbook.

ESRI:1997:UGA

[Env97b] Environmental Systems Research Institute (Redlands, CA). Understanding GIS: the ARC/INFO method: selfstudy workbook: version 7.1 for UNIX and Windows NT. Wiley, New York, NY, USA, fourth edition, 1997. ISBN 1-86242-033-5. various pp. LCCN G70.2 .U535 1997. Includes index.

Emerson:1987:TTT

[EP87a] Sandra L. Emerson and Karen Paulsell. troff Typesetting for UNIX systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-930959-4. xx + 359 pp. LCCN Z253.4.U53 E45 1987. US$29.95.

Emerson:1987:TTU

[EP87b] Sandra L. Emerson and Karen Paulsell. troff Typesetting for UNIX systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-930959-4. xx + 359 pp. LCCN Z253.4.U53 E45 1987. US$29.95.

Eggert:2005:PEN

[EP05] P. R. Eggert and D. S. Parker. Perturbing and evaluating numerical programs without recompilation — the wonglediff way. Software—Practice and Experience, 35(4):313–322, April 10, 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Epplin:2000:IDH

[Epp00a] Jerry Epplin. Inside Debian Hurd. Dr. Dobb’s Journal of Software Tools, 25(12):21–22, 24, 26, December 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Epplin:2000:IRT

[Epp00b] Jerry Epplin. Inside Real-Time Linux. Dr. Dobb’s Journal of Software Tools, 25(3):72, 74, 76, 78, March 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_03/rtlinux.
                                        txt;   http://www.ddj.com/
                                        ftp/2000/2000_03/rtlinux.
                                        zip.

Epstein:1992:TCU

[Eps92] Jeremy Epstein. Is there a C2 UNIX system in the house? In USENIX [USE92e], pages 243–252. ISBN 1-880446-46-4. LCCN ????

Emery:1986:TUT

[ER86] Dave Emery and Steve Rosen.

Two UNIX tools supporting a common style. ACM SIGADA Ada Letters, 6(2):84, March/ April 1986. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Elling:2001:DES

[ER01] Richard Elling and Tim Read.

Designing Enterprise Solutions with Sun Cluster 3.0. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-008458-1. xxviii + 270 pp. LCCN TK5105.52.E45 2002.

URL  books/SC3_0.pdf;
http://www.sun.com/books/
catalog/elling/.

Erlinger:1988:NCU

[Erl88] Michael A. Erlinger. A notice capability for UNIX. In USENIX Association [USE88d], pages 21–22. ISBN ???? LCCN ????

Ermann:1979:FSX

[Erm79] R. Ermann. Formal specification of X.25 compatible link protocol. Technical Memorandum 1389 (TM 79-5251-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 21, 1979. ?? pp.

Ernest:2011:OOS

[Ern11] Michael Ernest. OCA Oracle Solaris 11 System Administrator Exam Guide (Exam 1Z0-821). McGraw-Hill Osborne Media, ????, 2011. ISBN 0-07-177574-9. ???? pp. LCCN ????

Ellis:1990:ACR

[ES90] Margaret A. Ellis and Bjarne Stroustrup. The Annotated C++ Reference Manual. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-51459-

1. x + 447 pp. LCCN QA76.73.C153 E35 1990.

Evans:1993:OUR

[ESSB93] Steve Evans, Bart Smaalders, Dave Singleton, and Jeff Bonwick. Optimizing UNIX resource scheduling for user interaction. In USENIX [USE93c], pages 205–218. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/cinci93/evans.
html.

Elsesser:1986:MSC

[EST86] G. W. Elsesser, M. S. Safran, and T. Tieger. Managing separate compilation in AT&T’s UNIX Ada system. In USENIX Association [USE86e], pages 252–260. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Egan:1988:WUD

[ET88] Janet I. Egan and Thomas J. Teixeira. Writing a UNIX device driver. Wiley, New York, NY, USA, 1988. ISBN 0-471-62859-X (paperback), 0-471-62811-5. viii + 357 pp. LCCN QA76.76.O63 E35 1988.

Egan:1992:WUD

[ET92] Janet I. Egan and Thomas J.

Teixeira. Writing a UNIX Device Driver. Wiley, New York, NY, USA, second edition, 1992. ISBN 0-471-53575-3 (cloth), 0-471-53574-5 (paper). ix + 358 pp. LCCN QA76.76.D49 E42 1992. A classic book on writing programs to control hardware devices. See book review [Wei93].

EurOpen:1991:EUD

[Eur91] EurOpen, editor. EurOpen.

UNIX Distributed Open Systems in Perspective. Proceedings of the Spring 1991 EurOpen Conference, Tromsø, Norway, May 20–24, 1991. EurOpen, Buntingford, Herts, UK, 1991. ISBN 1-873611-00-5. LCCN ????

Evard:1997:AUS

[Eva97] Remy Evard. An analysis of UNIX system configuration. In USENIX [USE97b], page ?? ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL  http://www.usenix.
org/publications/library/
proceedings/lisa97/20.evard.
html.

Everett:1978:APP

[Eve78] W. W. Everett. Activity profile of the PWB/UNIX operating system. Technical Memorandum 1290, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 15, 1978. ?? pp.

Ewing:1986:OOO

[Ewi86]  J. J. Ewing.  An object-

oriented operating system interface. ACM SIGPLAN Notices, 21(11):46–53, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Eykholt:1986:PUS

[Eyk86] Joseph R. Eykholt. Porting UNIX to the System/370 Extended Architecture. In USENIX Association [USE86e], pages 157–164. URL https://

                                        bitsavers.org/pdf/usenix/
                                        USENIX_1986_Winter_Technical_
                                        Conference_Proceedings_198601.
                                        pdf.

Eykholt:1988:NEH

[Eyk88] Joseph R. Eykholt. A new exception handling mechanism for the UNIX kernel. In USENIX Association [USE88g], pages 291–295. ISBN ???? LCCN ????

Fernandez:1988:EUP

[FA88] Gary Fernandez and Larry Allen. Extending the UNIX protection model with access control lists. In Association [Ass88d], pages 119–132.

Frey:1989:VDE

[FA89] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1989. ISBN 0-937175-39-0. xv + 284 pp. LCCN HE6239.E54 F73 1989.

Frey:1994:VDE

[FA94] Donnalyn Frey and Rick Adams. !%@:: a Directory of Electronic Mail Addressing and Networks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 1994. ISBN 1-56592-046-5. 662 pp. LCCN HE6239.E54 F73 1989. US$9.95.

Feitelson:2007:FGA

[FAB^{+}07] Dror G. Feitelson, Tokunbo O. S. Adeshiyan, Daniel Balasubramanian, Yoav Etsion, Gabor Madl, Esteban P. Osses, Sameer Singh, Karlkim Suwanmongkol, Minhui Xie, and Stephen R. Schach. Fine-grain analysis of common coupling and its application to a Linux case study. The Journal of Systems and Software, 80(8): 1239–1255, August 2007. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Factor:2002:AAS

[Fac02] Alexander Factor. Analyzing application service providers. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-089425-7. xxiv + 326 pp. LCCN HF5548.32 .F33

                                        2002. URL http://www.sun.
                                        com/books/catalog/factor/.

Farrell:1978:UGS

[Far78] J. W. Farrell. User’s guide to the SMAL2 language for the Zilog Z80 microprocessor. Technical Memorandum 1262 (TM 78-8234-8), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 21, 1978. ?? pp.

Fargo:1979:SSE

[Far79] P. A. Fargo. SE — a structure editor program for accessing C program structures. Technical Memorandum 1468 (TM 79-5354-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 17, 1979. ?? pp.

Farrow:1990:USS

[Far90] Rik Farrow. UNIX System Security — How to Protect your Data and Prevent Intruders. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-57030-0. vii + 278 pp. LCCN QA76.76.O63 F38 1991. US$22.95, CDN$29.95. Fine … slightly not as technical as other books.

Farris:1991:CMU

[Far91a] Rick Farris. Choosing a Modem for UNIX. UNIX/World, 8 (10):85–??, October 1991. ISSN 0739-5922.

Farris:1991:UMG

[Far91b] Rick Farris. UNIX Mail Gets Easier. UNIX/World, 8(11): 111–??, November 1991. ISSN 0739-5922.

Farmer:1992:SBR

[Far92a] Dan Farmer. SAGE book reviews: Practical UNIX security. ;login: the USENIX Association newsletter, 17(6):13– 14, November/December 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Farris:1992:MEU

[Far92b] Rick Farris. Modern Editors for UNIX. UNIX/World, 9(2):73– 80, February 1992. ISSN 0739-5922.

Farrow:1992:AUS

[Far92c] Rik Farrow. Auditing UNIX Security, Effortlessly. UNIX/World, 9(8):85–??, August 1992. ISSN 0739-5922.

Farrow:1992:NHU

[Far92d] Rik Farrow. New Hope for UNIX Security? UNIX/World, 9(10):73–??, October 1992. ISSN 0739-5922.

Farrow:1995:RFU

[Far95] Rik Farrow. Report on the Fifth USENIX UNIX security symposium. ;login: the USENIX Association newsletter, 20(4):6–10, August 1995. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Fasciano:1978:IDA

[Fas78] V. A. Fasciano. Interactive design aid for general-purpose artwork production. Technical Memorandum 1266 (TM 78-8254-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1, 1978. ?? pp.

Fateman:1982:MPV

[Fat82] Richard J. Fateman. MAC-SYMA primer for VAX/UNIX. Computing Services, University of California, Berkeley, Berkeley, CA, USA, 1982. 32 pp.

Ferguson:1993:MRM

[FB93] Paula Ferguson and David Brennan. Motif Reference Manual, volume 6B. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1993. ISBN 1-56592-038-4. 920 pp. LCCN QA76.76.W56F47 1993. US$34.95.

Faulkner:1979:UTA

[FDMS79] R. A. Faulkner, D. S. DeJager, J. A. McGuire, and A. L. Sabsevitz. UNIX/TS augmented task force report. Technical Memorandum 1382, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 19, 1979. ?? pp.

Fraser:1977:FIS

[FE77] A. G. Fraser and R. J. Elliott. File integrity and space accounting for a UNIX based system. Technical Memorandum 1164 (TM 77-1273-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 4, 1977. ?? pp.

Feder:1977:MFL

[Fed77] Jerome Feder. More flexible low speed I/O or UNIX. Technical Memorandum 1202, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1977. ?? pp.

Feder:1978:FTH

[Fed78a] Jerome Feder. A faster terminal handler for UNIX. Technical Memorandum 1265 (TM 78-8234-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 10, 1978. ?? pp.

Feder:1978:THU

[Fed78b] Jerome Feder. Terminal handling under UNIX and MERT/ UNIX. Technical Memorandum 1316 (TM 78-8234-11), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1, 1978. ?? pp.

Feder:1979:RCP

[Fed79] Jerome Feder. Relative CPU performance on C language programs. Technical Memorandum 1480 (TM 79-3624-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 28, 1979. ?? pp.

Feder:1983:EUS

[Fed83] Jerome Feder. Evolution of UNIX system performance. In Association [Ass83], pages 110– 111. Abstract only.

Feder:1984:EUS

[Fed84] Jerome Feder. Evolution of UNIX system performance. AT&T Bell Laboratories Technical Journal, 63(8 part 2):

1791–1814, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Fedor:1988:GAM

[Fed88a] Mark S. Fedor. Gated: a multirouting protocol daemon for UNIX. In Association [Ass88d], pages 365–376.

Fedor:1988:GMR

[Fed88b] Mark S. Fedor. Gated: a multi-routing protocol daemon for UNIX. In USENIX Association [USE88g], pages 365–376. ISBN ???? LCCN ????

Feiler:2001:MXC

[Fei01] Jesse Feiler. Mac OS X: the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212663-9. xxxv + 763 pp. LCCN QA76.76.O63 F43 2001 Computer.

Feiler:2002:MXD

[Fei02] Jesse Feiler. Mac OS X developer’s guide. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2002. ISBN 0-12-251341-X. xxiv + 594 pp. LCCN QA76.76.O63 F435 2002.

Feldman:1975:ASC

[Fel75] Stuart I. Feldman. An application of symbolic computation to crystal physics. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 9(2):19–24, May 1975.

CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic). URL https:

//dl.acm.org/doi/10.1145/
1088301.1088307.

Feldman:1978:MPMa

[Fel78]  Stuart I. Feldman.  Make

— a program for maintaining computer programs. Technical Memorandum 1078, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1978. 9 pp.

Feldman:1979:PLEb

[Fel79a]  S. I. Feldman.  The pro-

gramming language EFL. In ACM [ACM79], pages 76–

79.  URL https://dl.
acm.org/doi/abs/10.1145/
800208.806435.

Feldman:1979:IPF

[Fel79b] Stuart I. Feldman. Implementation of a portable Fortran 77 compiler using modern tools. Technical Memorandum 1354 (TM 79-1273-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1, 1979. ?? pp.

Feldman:1979:MPM

[Fel79c] Stuart I. Feldman. Make — a program for maintaining computer programs. Software— Practice and Experience, 9(4): 255–265, April 1979. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Feldman:1979:PLEa

[Fel79d] Stuart I. Feldman. The programming language EFL. Technical Memorandum 1410 (TM 79-1273-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 4, 1979. ?? pp.

Felton:1979:UFS

[Fel79e] W. A. Felton. UNIX/370: A feasibility study. Technical Memorandum 1352 (TM 79-8224-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 23, 1979. ?? pp.

Feldman:1984:AHU

[Fel84] Stuart I. Feldman. An architecture history of the UNIX system. In Software Tools Users Group [Sof84], pages xi–xvi. LCCN QA76.8.U65 U83 1984.

Feldman:1994:TUP

[Fel94] Jonathan Feldman. A transparent UNIX/PC connection via NFS. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):85–??, July/August 1994. CODEN SYADE7. ISSN 1061-2688.

Feldman:1996:BLW

[Fel96a] Jonathan Feldman. Building a Linux Web server. Sys Admin: The Journal for UNIX Systems Administrators, 5(10): 41–??, October 1996. CODEN SYADE7. ISSN 1061-2688.

Feldman:1996:ULR

[Fel96b]  Jonathan Feldman.  Using

Linux as a router. Sys Admin:

The Journal for UNIX Systems Administrators, 5(1):8–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Feldman:1974:ASC

[Fel4 ] Stuart I. Feldman. An application of symbolic computation to crystal physics. Computing Science Technical Report 15, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974 (??). ?? pp.

Feldman:1978:PLE

[Fel8 a] S. I. Feldman. The programming language EFL. Computing Science Technical Report 78, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1978 (??). ?? pp.

Feldman:1978:MPMb

[Fel8 b] Stuart I. Feldman. Make — a program for maintaining computer programs. Computing Science Technical Report 57, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1978 (??). ?? pp.

Ferrin:1982:PIV

[Fer82] Thomas E. Ferrin. Performance issues of VMUNIX revisited. ;login: the USENIX Association newsletter, 7(5):3– 6, November 1982. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Ferrin:1985:REP

[Fer85] Thomas Ferrin. A recipe for establishing point-to-point TCP/ IP network links with 4.2 BSD UNIX. In USENIX Association [USE85d], pages 113–117. LCCN QA76.8.U65 U8 1985.

Fernando:2004:GGP

[Fer04] Randima Fernando, editor. GPU gems: programming techniques, tips, and tricks for realtime graphics, volume 1 of GPU gems. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-321-22832-4. xvv + 765 pp. LCCN T385 .G6879 2004. US$45.99.

Frankston:1999:LHN

[FES^{+}99] Bob Frankston, Alec Effrat, Fred L. Schneider, Ed Viguerie, William D. Jackson, Jim Geringer, Kenneth Nichols, Alexander Fronk, Pearl Brereton, David Budgen, and Geoff Hamilton. Letters: The home network will be an IP network; pining for the perfect job; what the dickens make Thompson fault Linux?; UNIX Systems Laboratory; presto! A true test of AI; are software patents really different?; support for hypertext maintenance. Computer, 32(6): 6–9, June 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r6006.pdf.

Feuer:1978:STA

[Feu78] A. R. Feuer. STAT: A tool for analyzing data. Technical Memorandum 1252 (TM 78-9442-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 21, 1978. ?? pp.

Feuer:1979:PGO

[Feu79a] A. R. Feuer. PWB/Graphics overview. Technical Memorandum 1398 (TM 79-3782-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 11, 1979. ?? pp.

Feuer:1979:TIG

[Feu79b] A. R. Feuer. A tutorial introduction to the graphical editor (GED). Technical Memorandum 1441 (TM 79-3782-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 30, 1979. ?? pp.

Feuer:1985:SIC

[Feu85] Alan R. Feuer. si — an interpreter for the C language. In USENIX Association [USE85d], pages 47–55. LCCN QA76.8.U65 U8 1985.

Frakes:1988:CES

[FF88] William B. Frakes and Christopher J. Fox. CEST: an expert system function library and workbench for UNIX system/ C language. AT&T Technical Journal, 67(2):95–106, March/ April 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Fenart:1986:OTI

[FFH^{+}86] Jean Marc Fenart, Marc Fievet, Christian Huitema, Bernard Martin, Annie Remille, and Guy Vaysseix. OSI and TCP/ IP protocols on a UNIX system. In USENIX Association [USE86c], pages 46–58.

Fernandez-Fuentes:2022:DFA

[FFPC22] Xosé Fernández-Fuentes, Tomás F. Pena, and José C. Cabaleiro. Digital forensic analysis methodology for private browsing: Firefox and Chrome on Linux as a case study. Computers & Security, 115(??):Article 102626, April 2022. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167404822000256.

Feuer:1979:CPL

[FG79] A. R. Feuer and N. H. Gehani. A comparison of the programming languages C and Pascal — part one: Language concepts. Technical Memorandum 1465, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 12, 1979. ?? pp.

Fateman:1980:AMM

[FG80] Richard J. Fateman and Mathlab Group. Addendum to the Mathlab/MIT MACSYMA reference manual for VAX/UNIX “VAXIMA”. Computing Services, University of California, Berkeley, Berkeley, CA, USA, 1980. 6 pp.

Faulkner:1991:PFS

[FG91] Roger Faulkner and Ron Gomes. The process file system and process model in UNIX System V. In USENIX [USE91d], pages 243–252. LCCN QA 76.76 O63 U84 1992.

URL http://obits.mlive.
com/obituaries/grandrapids/
obituary.aspx?pid=180588279;
http://thenewstack.io/remembering-
roger-faulkner/;   https:
//www.usenix.org/memoriam-
roger-faulkner;   https:/
/www.usenix.org/sites/default/
files/usenix_winter91_faulkner.
pdf.

Forouzan:2003:USP

[FG03] Behrouz A. Forouzan and Richard F. Gilberg. UNIX and Shell programming: a textbook. Brooks/Cole-Thomson Learning, Pacific Grove, CA, 2003. ISBN 0-534-95159-7 (paperback). xix + 875 pp. LCCN QA76.76.O63 F59715 2003.

Fourer:1989:AMP

[FGK89] Robert Fourer, David M. Gay, and Brian W. Kernighan. AMPL: A mathematical programming language. Computing Science Technical Report 133, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1989. ii + 63 pp. Revision of January 1987 edition.

Feldman:1990:FCC

[FGMS90] S. I. Feldman, David M. Gay, Mark W. Maimone, and N. L.

Schryer. A Fortran-to-C converter. Computing Science Technical Report 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 16, 1990. i + 24 + 2 pp. URL ftp:/

/www.digiater.nl/openvms/
decus/lt90a/translators/
vms-f2c/f2c_technical_report.
ps.

Feldman:1995:FCC

[FGMS95] S. I. Feldman, David M. Gay, Mark W. Maimone, and N. L. Schryer. A Fortran-to-C converter. Computing Science Technical Report 149, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1995. i + 25 pp. Revision of edition of 18 May 1990.

Farber:1964:SSM

[FGP64] D. J. Farber, R. E. Griswold, and I. P. Polonsky. SNOBOL, A string manipulation language. Journal of the ACM, 11 (1):21–30, January 1964. CO-DEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Furht:1991:RSH

[FGP^{+}91] B. Furht, D. Gluch, J. Parker, P. Matthews, and D. Joseph. Real/star 2000. A high performance multiprocessor computer for telemetry applications. International Telemetering Conference (Proceedings), 27:365–373 (of 932), 1991. CO-DEN ITCOD6. ISBN 1-55617-329-6. ISSN 0884-5123 (print), 1546-2188 (electronic).

Funkenhauser:1988:UTU

[FH88] M. J. Funkenhauser and R. C.

Holt. Using TUNIS, A UNIX compatible kernel, as a basis for security. In USENIX Association [USE88e], pages 70– 77. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

Fleckenstein:1989:UGN

[FH89] Charles J. Fleckenstein and David Hemmendinger. Using a global name space for parallel execution of UNIX tools. Communications of the Association for Computing Machinery, 32(9):1085–1090, September 1989. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/66456.html.

Farrow:1992:PUE

[FH92] Rik Farrow and Frank Hayes.

The PC UNIX Explosion. UNIX/World, 9(10):48–??, October 1992. ISSN 0739-5922.

Fraser:1996:CCF

[FH96] Christopher W. Fraser and David R. Hanson. Compile C faster on Linux. Linux Journal, 25:32–33, May 1996. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic). URL

http://www.cs.princeton.
edu/software/lcc/doc/linux.
html.

Fritz:1984:NCR

[FHR84] T. E. Fritz, J. E. Hefner, and T. M. Raleigh. Network of computers running the UNIX system. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1877–1896, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Fox:1977:PMS

[FHS77] P. A. Fox, A. D. Hall, and N. L. Schryer. The PORT mathematical subroutine library. Computing Science Technical Report 47, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1977. ii + 50 pp. URL http://

                                        history.siam.org/%5C/sup/
                                        Fox_bell_subroutine.pdf.

Fox:1978:AFP

[FHS78a] P. A. Fox, A. D. Hall, and N. L. Schryer. Algorithm 528: Framework for a portable library [Z]. ACM Transactions on Mathematical Software, 4 (2):177–188, June 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remarks [Fox79, GG99].

Fox:1978:PMS

[FHS78b] P. A. Fox, A. D. Hall, and N. L. Schryer. The PORT mathematical subroutine library. ACM Transactions on Mathematical Software, 4(2):104–126, June 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fox:197x:BUP

[FHS7x] P. A. Fox, A. D. Hall, and N. L. Schryer. Basic utilities for portable FORTRAN libraries. Computing Science Technical Report 37, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? 197x. 53 pp. URL https:/

/archive.computerhistory.
org/resources/access/text/
2022/08/102805039-05-01-
acc.pdf.

Fiedler:1991:USV

[FHS91] David Fiedler, Bruce Hunter, and Ben Smith. UNIX System V Release V Administration. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, second edition, 1991. ISBN 0-672-22810-6. 436 pp. LCCN QA76.76.O63 F54 1991. US$29.95. Highly recommended. A must for any intermediate/advanced users of UNIX.

Fiedler:1988:UIE

[Fie88] David Fiedler. USENET: An informal but extensive Comm-Net for UNIX and XENIX systems and users. C Users Journal, 6(1):54–??, January 1988. ISSN 0898-9788.

Fildes:1979:CAV

[Fil79]  N. R. Fildes.  The Craft

Access Voice Editing System (CAVES). Technical Memorandum 1416 (TM 79-9314-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 20, 1979. ?? pp.

Filipski:1985:LSB

[Fil85] Alan Filipski. Latent source bugs and UNIX system portability. In USENIX Association [USE85c], pages 125–130.

Finseth:1991:CTE

[Fin91] Craig A. Finseth. The Craft of Text Editing: Emacs for the Modern World. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97616-7 (New York), 3-540-97616-7 (Berlin). xii + 220 pp. LCCN QA76.76.T49 F56 1991. Contains extensive discussion of design issues for text editors, with examples from Emacs. Appendix B gives sources of numerous Emacs implementations. Appendix D summarizes the TECO command set.

Fioretti:2004:RMU

[Fio04] Marco Fioretti. Review: Mastering UNIX Shell Scripting. Linux Journal, 2004(127):??, November 2004. CODEN LI-JOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Fischer:1973:PAA

[Fis73] P. C. (Patrick C.) Fischer, editor. Proceedings of the 1st annual ACM SIGACT-SIGPLAN symposium on Principles of programming languages — POPL ’73 [Boston, MA, 1–3 October 1973. ACM Press, New York, NY, USA, 1973. ISBN 1-4503-7349-6. LCCN ????

Fischer:1986:ACU

[Fis86a] Herman Fischer. Ada, “C”, and UNIX. In USENIX [USE86b], pages 225–240.

Fischer:1986:AU

[Fis86b] Herman Fischer. Ada, “C”, and UNIX. In USENIX Association [USE86e], pages 225–240. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Fischer:1986:SBC

[Fis86c] Herman Fischer. SVID as a basis for CAIS implementation. In USENIX [USE86b], pages 294– 301.

Fischer:2002:OEL

[Fis02] Todd Fischer. Optimizing Embedded Linux. Dr. Dobb’s Journal of Software Tools, 27 (5):51–57, May 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_05/emlinux.
txt.

Fitzhorn:1981:CTC

[FJ81] Patrick A. Fitzhorn and Gearold R. Johnson. C: toward a concise syntactic description. ACM SIGPLAN Notices, 16(12):14–21, December 1981. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster-Johnson:1997:UPT

[FJ97] Eric Foster-Johnson. UNIX programming tools. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1997. ISBN 1-55851-482-1. xviii + 346 pp. LCCN QA 76.76 O63 J627 1997.

Fox:1993:BLR

[FJPR93] Richard Fox, Alan Joch, Leonard Presberg, and Leslie Reisz. BYTE Lab report: Desktop dynamite: 116 fast 486s: We tested 116 50-and 66-MHz 486s to tell you which is best for your DOS, Windows, and UNIX applications. how we tested; poor quality; bus choices. BYTE Magazine, 18 (7):156–??, June 1993. CODEN BYTEDJ. ISSN 0360-5280. Foster-Johnson:2005:BSS

[FJWA05] Eric Foster-Johnson, John C. Welch, and Micah Anderson. Beginning shell scripting. Programmer to programmer. Wiley, New York, NY, USA, 2005. ISBN 0-7645-8320-4. xviii + 510 pp. LCCN QA76.76.O63 F59717 2005. URL ftp:

//uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/toc/
ecip056/2005002075.html.

Feldmeier:1990:UPS

[FK90] David C. Feldmeier and Philip R. Karn. UNIX password security — ten years later (invited). Lecture Notes in Computer Science, 435:44–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/0435/04350044.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/0435/04350044.
pdf.

Ferguson:2008:REC

[FKL^{+}08] Justin Ferguson, Dan Kaminsky, Jason Larsen, et al., editors. Reverse engineering code with IDA Pro. Syngress Publishing, Inc., Rockland, MA, USA, 2008. ISBN 1-59749-237-X. xii + 316 pp. LCCN QA76.76.D57 R49 2008.

Fournier:1995:VMS

[FKM95] Robert Fournier, Norbert Kajler, and Bernard Mourrain. Visualization of mathematical surfaces: the IZIC server approach. Journal of Symbolic Computation, 19(1/2/3): 159–174 (or 159–173??), January, February, March 1995. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic). Design and implementation of symbolic computation systems (Gmunden, 1993).

Finger:1985:MCV

[FKN85] Eric J. Finger, Michael M.

Krueger, and Alan F. Nugent. A Multi-CPU version of the UNIX kernel — technical aspects and market need. In USENIX Association [USE85c], pages 11–22.

Ferrin:1980:ICG

[FL80] T. E. Ferrin and R. Langridge. Interactive computer graphics with the Unix time-sharing system. Computer Graphics, 13(4):321–331, February 1980. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Flatters:1990:TUP

[Fla90] Chris Flatters. UNIX Programming: Methods and Tools, by James F. Peters III. C Users Journal, 8(11):119–??, November 1990. ISSN 0898-9788. See [Pet88].

Flanagan:1991:PSR

[Fla91] David Flanagan. Programmer’s Supplement for Release 5 of the X Window System, Version 11. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-86-2. xx + 367 pp. LCCN QA76.76.W56 F5 1991.

Flanagan:1994:MTS

[Fla94] David Flanagan. Motif Tools: Streamlined GUI Design and Programming with the Xmt Library. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1994. ISBN 1-56592-044-9. xxxvii + 984 pp. LCCN QA76.76.W56 F53 1994. US$54.95 (CD-ROM included).

Flannery:2000:IH

[Fla00] Ron Flannery. The Informix handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-012247-5. lxvii + 1362 pp. LCCN QA76.73.I22 F53 2000.

Flegel:2002:PUL

[Fle02] Ulrich Flegel. Pseudonymizing Unix log files. Lecture Notes in Computer Science, 2437:162– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2437/24370162.htm;
http://link.springer.de/
link/service/series/0558/
papers/2437/24370162.pdf.

Flickenger:2003:LSH

[Fli03] Rob Flickenger. Linux server hacks: 100 industrial-strength tips and tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00461-3. xvii + 221 pp. LCCN QA76.9.C55 F58 2003.

URL http://www.oreilly.
com/catalog/9780596004613.

Florence:1990:UTF

[Flo90] Ronald Florence. UNIX ‘termcap’ facility improves portability by hiding terminal dependencies. C Users Journal, 8 (1):93–??, January 1990. ISSN 0898-9788.

Forin:1994:MFU

[FM94] Alessandro Forin and Gerald Malan. An MS-DOS filesystem for UNIX. In USENIX [USE94f], pages 337– 354. ISBN 1-880446-58-8. LCCN QA 76.76 O63 U84

1994.  URL http://www.
usenix.org/publications/
library/proceedings/sf94/
forin.html.

Flautner:2002:VAP

[FM02] Krisztián Flautner and Trevor Mudge. Vertigo: automatic performance-setting for Linux. Operating Systems Review, 36 (5S):105–116, December 2002. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Felton:1984:USI

[FMM84] William A. Felton, Gerald L. Miller, and J. Michael Milner. A UNIX system implementation for System/370. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1751–1767, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic). URL https:

//www.bell-labs.com/usr/
dmr/www/otherports/ibm.html;
 https://www.bell-labs.com/
usr/dmr/www/otherports/ibm.
pdf.

Fostel:1983:DUV

[FN83] Gary Fostel and Alison Naylor. Developing a UNIX validation set. In Software Tools Users Group [Sof83], pages 329–339. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Fogel:1999:OSD

[Fog99] Karl Fogel. Open source development with CVS. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN 1-57610-490-7. xviii + 316 pp. LCCN QA76.76.D47 F63 1999.

Foley:1977:DAD

[Fol77]  G. B. Foley.  A descrip-

tion of automatic dialing using the UNIX operating system. Technical Memorandum 1180, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1977. ?? pp.

Forsyth:2000:MTL

[For00] C. H. Forsyth. More taste:

Less greed? or Sending UNIX to the fat farm. Report, Department of Computer Science, University of York, Heslington, York YO1 5DD, England, October 31, 2000. 11 pp.

URL http://www.collyer.
net/who/geoff/taste.pdf;
 http://www.terzarima.net/
doc/taste.pdf.

Fosner:1996:OPW

[Fos96] Ron Fosner. OpenGL programming for Windows 95 and Windows NT. Addison-Wes-ley, Reading, MA, USA, 1996. ISBN 0-201-40709-4. xv + 259 pp. LCCN T385 .F663 1996. US$39.76. Includes CD-ROM.

Fowler:1985:FGM

[Fow85] Glenn S. Fowler. The fourth generation Make. In USENIX Association [USE85d], pages 159–174. LCCN QA76.8.U65 U8 1985.

Fox:1979:RFP

[Fox79] Phyllis Fox. Remark on “Algorithm 528: Framework for a portable library [Z]”. ACM Transactions on Mathematical Software, 5(4):524, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [FHS78a].

Foxley:1985:US

[Fox85] Eric Foxley. UNIX for superusers. International computer science series. Addison-Wes-ley, Reading, MA, USA, 1985. ISBN 0-201-14228-7. xiv + 213 pp. LCCN QA76.76.O63 F69 1985 Bar. UK£11.95.

Furht:1989:PRI

[FPG89] Borko Furht, J. Parker, and D. Grostick. Performance of REAL/IX^{T M}-fully preemptive real time UNIX. Operating Systems Review, 23(4):45–52, October 1989. CODEN OSRED8.

ISSN 0163-5980 (print), 1943-586X (electronic).

Fraser:1975:COM

[FR75] A. G. Fraser and Dennis M.

Ritchie. C-language oriented microprogram for the HP-21 MX. Technical Memorandum 1058 (TM 75-1273-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 25, 1975. ?? pp.

Fraser:1977:CDA

[FR77] A. G. Fraser and G. G. Riddle. Circuit design aids on UNIX. Technical Memorandum 1165 (TM 77-1273-5, TM 77-9155-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 29, 1977. ?? pp.

Fraser:1974:SEDa

[Fra74a]  A. G. Fraser.  Spider —

an experimental data communications system. Computing Science Technical Report 23, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1974. ??

pp.  URL https://www.
cst.cam.ac.uk/news/death-
computer-science-pioneer-
sandy-fraser;   https://
www.fraserresearch.org/images/
report_2004.pdf.

Fraser:1974:SEDb

[Fra74b] A. G. Fraser. Spider — an experimental data communications system. In IEEE, editor, Proceedings of the IEEE International Conference on Communications (ICC’74), Minneapolis (June 17–9, 1974), pages 21F–1–21F–10. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1974. IEEE Catalog Number 74CH0859-9-CSCB.

Fraser:1976:PSF

[Fra76] A. G. Fraser. The present status and future trends in computer/communication technology. Communications Society, 14(5):10–19, September 1976. ISSN 0094-5579. URL

https://ieeexplore.ieee.
org/abstract/document/1089243.

Fraser:1978:UTS

[Fra78] A. G. Fraser. UNIX timesharing system: Circuit design aids. The Bell System Technical Journal, 57(6): 2233–2249, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2233.
pdf.

Fraser:1979:CLR

[Fra79] A. G. Fraser. C language routines for multi-thread computations. Technical Memorandum 1388 (TM 79-1273-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 7, 1979. ?? pp.

Fraser:1980:GTE

[Fra80] C. W. Fraser. A generalized text editor. Communications of the Association for Computing Machinery, 23(3): 154–158, March 1980. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Frase:1983:HFB

[Fra83] L. T. Frase. Human factors and behavioral science: The UNIX Writer’s Workbench software: Philosophy. The Bell System Technical Journal, 62(6):1883–1890, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1883.
pdf.

Francis:1990:OFS

[Fra90a] B. Francis. Oil firms seek UNIX standards. Datamation, 36(18): 63–64, September 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Francis:1990:UCF

[Fra90b] B. Francis. UNIX checks in at four seasons. Datamation, 36 (8):47–49, April 1990. CODEN DTMNAT. ISSN 0011-6963.

Franklin:1992:TVM

[Fra92] James W. Franklin. Tiled virtual memory for UNIX. In USENIX [USE92b], pages 99– 106. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

Frazier:1999:PCC

[Fra99]  George F. Frazier.  Port-

ing C++ code from NT to UNIX: Using the MainWin XDE toolkit. Dr. Dobb’s Journal of Software Tools, 24(4):84, 86–89, April 1999. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/1999/1999_04/nt2unix.
txt;   http://www.ddj.com/
ftp/1999/1999_04/nt2unix.
zip.

Frazelle:2019:OSF

[Fra19] Jesse Frazelle. Open source firmware. Communications of the Association for Computing Machinery, 62(10):34–38, October 2019. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:

//cacm.acm.org/magazines/
2019/10/239673/fulltext.

Fraser:1975:PSF

[Fra5 ] A. G. Fraser. The present status and future trends in computer/communication technology. Computing Science Technical Report 40, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1975 (??). ?? pp.

Fajardo:1985:UOS

[FRAC85] Ray M. Fajardo, Andrew L. Rood, James R. Andreas, and Robert C. Cline. A UNIX operating system adapted for a technical personal computer. Hewlett–Packard Journal, 36 (10):22–28, October 1985. CO-DEN HPJOAX. ISSN 0018-1153.

Freeman:1978:PCL

[Fre78] R. B. Freeman. Parse() — a C language routine to parse a string into words. Technical Memorandum 1271, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 15, 1978. ?? pp.

Freda:1990:UIC

[Fre90] William J. Freda. UNIX interprocess communications. C Users Journal, 8(11):49–??, November 1990. ISSN 0898-9788.

Freedman:1991:EBP

[Fre91] Dan Freedman. Experience building a process migration subsystem for UNIX. In USENIX [USE91d], pages 349– 356. LCCN QA 76.76 O63 U84 1992.

Freeze:1998:SPR

[Fre98] Wayne S. Freeze. The SQL programmer’s reference: Windows 95/NT and UNIX. Ventana Press, Chapel Hill, NC, USA, 1998. ISBN 1-56604-760-9. liv + 298 pp. LCCN QA76.73.S67 F74 1998.

Frisch:1991:ESA

[Fri91] Æleen Frisch. Essential System Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-74-9 (paperback), 0-937175-80-3 (paperback). xxiii + 440 pp. LCCN QA76.76.O63 F78 1992.

URL http://www.oreilly.
com/catalog/9780937175804.

Friedl:1997:MRE

[Fri97] Jeffrey E. F. Friedl. Mastering regular expressions: powerful techniques for Perl and other tools. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-257-3. xxiv + 342 pp. LCCN QA76.73.P22 F75 1997; QA76.9.D3G728 1997. US$29.95. URL http://www.

ora.com/catalog/regex/;
http://www.oreilly.com/catalog/
9781565922570;   http://
www.oreilly.com/catalog/
regex.

Friedl:2002:MRE

[Fri02] Jeffrey E. F. Friedl. Mastering Regular Expressions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00289-0. xxii + 460 pp. LCCN QA76.73.P22 F75 2002; QA76.9.T48 F75 2002. US$39.95. URL http:/

/www.oreilly.com/catalog/
9780596002893;   http://
www.oreilly.com/catalog/
regex2.

Frisch:2003:ESAb

[Fri03] Æleen Frisch. Essential System Administration Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00449-4. vi + 137 pp. LCCN QA76.76.O63 F782 2003. US$14.95. URL http:/

/www.oreilly.com/catalog/
esapr.

Friedl:2006:MRE

[Fri06] Jeffrey E. F. Friedl. Mastering regular expressions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2006. ISBN 0-596-52812-4 (paperback). xxiv + 515 pp. LCCN QA76.9.T48 F75

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007272426-d.html;
http://www.oreilly.com/
catalog/9780596528126.

Frost:1990:GTB

[Fro90] Lyle Frost. Get thee behind me, spawn of UNIX! Computer Language Magazine, 7(7): 51–??, July 1990. CODEN COMLEF. ISSN 0749-2839.

Fei:2007:EOS

[FRRJ07] Yunsi Fei, Srivaths Ravi, Anand Raghunathan, and Niraj K. Jha. Energy-optimizing source code transformations for operating system-driven embedded software. ACM Transactions on Embedded Computing Systems, 7(1):2:1–2:26, December 2007. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Finkel:1999:EUS

[FSS99] Raphael A. Finkel, Brian Sturgill, and Harlan Stenn. Experience with a Unix systemadministration tool. Software—Practice and Experience, 29(11):953–971, September 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
63501202;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=63501202&
PLACEBO=IE.pdf.

Fyfe:1991:CSS

[FST91] Alastair Fyfe, Ivan Soleimanipour, and Vijay Tatkar. Compiling from saved state: Fast incremental compilation with traditional UNIX compilers. In USENIX [USE91d], pages 161– 172. LCCN QA 76.76 O63 U84 1992.

Fink:2002:LPT

[FSW02] Jason Fink, Matt Sherer, and Kurt Wall. Linux performance tuning and capacity planning. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32081-9. xi + 317 pp. LCCN QA76.76.O63 F557 2002.

Franke:1983:EEU

[FT83] Deborah L. Franke and Thomas R. Truscott. Early experiences using UNIX on the Gould SEL concept computers. In Software Tools Users Group [Sof83], pages 341–347. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Foxwell:2009:PON

[FT09] Harry Foxwell and Christine Tran. Pro OpenSolaris: a New Open Source OS for Linux Developers and Administrators. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1891-6. xxi + 254 pp. LCCN QA76.76.O63 F59733 2009.

Funk:1987:CPM

[Fun87] Susan A. Funk. CAS perspective on the maturation of UNIX. In USENIX Association [USE87d], pages 95–104.

Furuta:TB6-1-15

[Fur85] Richard Furuta. Device drivers included on the UNIX TEX distribution tape. TUGboat, 6(1):15, March 1985. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb06-1/
tb11output.pdf.

Feldman:1978:PFC

[FW78] Stuart I. Feldman and Peter J.

Weinberger. A portable Fortran 77 compiler. Technical Memorandum 1248, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1978. ?? pp.

Feldman:1990:PFC

[FW90] S. I. Feldman and P. J. Weinberger. A portable Fortran 77 compiler. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Figgin:2003:LWS

[FWS^{+}03] Stephen Figgin, Aaron Weber, Ellen Siever, Matt Welsh, George Reese, Ben Laurie, Peter Laurie, Rasmus Lerdorf, Stas Bekman, and Eric Cholet, editors. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, version 2.0. edition, 2003. ISBN 0-596-00529-6. LCCN QA76.76.O63. URL http:/

/www.oreilly.com/catalog/
9780596005290.

Farfour:2018:SUG

[FY18] Mohammed Farfour and Wang Jung Yoon. Seismic UNIX and GNU Octave for VSP Data Processing and Interpretation, pages 73–92. Wiley, New York, NY, USA, 2018. ISBN 1-119-22751-8.

Gilly:1998:UN

[G^{+}98a] Daniel Gilly et al. UNIX in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, deluxe edition, July 1998. ISBN 1-56592-406-1. 444 (est.) pp. LCCN ???? US$69.95. URL http:/

/www.oreilly.com/catalog/
9781565924062;   http://

www.oreilly.com/catalog/
unixcd.

Gilly:1998:UND

[G^{+}98b] Daniel Gilly et al. UNIX in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, deluxe edition, July 1998. ISBN 1-56592-406-1. 444 (est.) pp. LCCN ???? US$69.95.

URL http://www.oreilly.
com/catalog/unixdeluxe.

Geiselhart:2006:IZV

[G^{+}06a] Gregory Geiselhart et al., editors. IBM z/VM and Linux on IBM System z: virtualization cookbook for Red Hat Enterprise Linux 4. Number SG24-7272-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2006. ISBN 0-7384-9495-X (paperback). xiv + 218 pp. LCCN QA76.76.O63 I28 2006.

Gerner:2006:PLL

[G^{+}06b] Jason Gerner et al., editors.

Professional LAMP: Linux, Apache, MySQL, and PHP Web development. Wiley, New York, NY, USA, 2006. ISBN 0-7645-9723-X. xxii + 379 pp. LCCN TK5105.888 .P677

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0623/2005026487-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0654/2005026487-
b.html;   http://www.loc.
gov/catdir/enhancements/
fy0654/2005026487-t.html.

Glass:1999:UPU

[GA99] Graham Glass and King Ables. UNIX for programmers and users. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-681685-1. xviii + 622 pp. LCCN QA76.76.O63 G583 1999.

Ganten:2004:DGLa

[GA04] Peter H. Ganten and Wulf Alex. Debian GNU/Linux: Grundlagen, Installation, Administration und Anwendung. (German) [Debian GNU/Linux: Basis, Installation, Administration, and Use]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-43267-1. xxii + 946 pp. LCCN ????

Gabarro:2007:WAD

[Gab07]  Steven A. Gabarró.  Web

application design and implementation: Apache 2, PHP5, MySQL, JavaScript, and Linux/Unix. Quantitative software engineering series. Wiley, New York, NY, USA, 2007. ISBN 0-471-77391-3 (cloth). xv + 295 pp. LCCN TK5105.8883 .G33 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0613/2006014999.html.

Gagne:2001:LSA

[Gag01] Marcel Gagné. Linux System Administration — a User’s Guide. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-71934-7. xxi + 532 pp. LCCN QA76.76.O63 G34 2002.

Gagne:2006:MUL

[Gag06] Marcel Gagné. Moving to Ubuntu Linux. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-42722-X (paperback). xxii + 463 pp. LCCN QA76.76.O63 .G3455 2006. US$34.99.

URL http://www.aw-bc.com/
catalog/academic/product/
0,1144,032142722X,00.html.

Includes CD-ROM.

Gagne:2007:MUL

[Gag07] Marcel Gagné. Moving to Ubuntu Linux. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-321-42722-X (paperback). xxii + 463 pp. LCCN QA76.76.O63 G3455 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0616/2006021595.html.

Gait:1983:CLS

[Gai83] J. Gait. Command level software tools. ACM SIGPLAN Notices, 18(11):45–53, November 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gait:1984:SOK

[Gai84] Jason Gait. Semaphores outside the kernel. ACM SIG-PLAN Notices, 19(10):12–21, October 1984. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gallmeister:1991:PPR

[Gal91a]  Bill Gallmeister.  Portable

POSIX in real time. UNIX Review, 9(4):32–??, April 1991. CODEN UNRED5. ISSN 0742-3136.

Gallmeister:1991:RUA

[Gal91b] William O. Gallmeister. Reconciling UNIX, Ada, and realtime processing. Dr. Dobb’s Journal of Software Tools, 16 (6):56, 58, 60, 62, 64, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Gallmeister:1992:RP

[Gal92] Bill O. Gallmeister. Real-time POSIX. Embedded Systems Programming, 5(10):28–??, October 1992. CODEN EYPRE4. ISSN 1040-3272.

Gallmeister:1993:UP

[Gal93] B. Gallmeister. Understanding POSIX 4 and 4.a. In Anonymous [Ano93x], pages 211–222. ISBN 0-87930-306-9. LCCN TK7895.E42 E53 1993.

Gallmeister:1995:PPR

[Gal95] Bill Gallmeister. POSIX.4: Programming for the Real World. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1995. ISBN 1-56592-074-0. xviii + 548 pp. LCCN QA76.76.O63 G34 1995. US$29.95.

Galli:2000:DOS

[Gal00] Doreen L. Galli. Distributed Operating Systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-079843-6. xix + 463 pp. LCCN QA76.76.O63 G35 2000.

URL http://www.phptr.com/
ptrbooks/esm_0130798436.
html.

Galvin:2005:SC

[Gal05] Peter Baer Galvin. Solaris 10 Containers. ;login: the USENIX Association newsletter, 30(5):??, October 2005. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic). URL http://www.

usenix.org/publications/
login/2005-10/pdfs/galvin.
pdf.

Gancarz:1995:UP

[Gan95] Mike Gancarz. The UNIX philosophy. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1995. ISBN 1-55558-123-4. xix + 151 pp. LCCN QA76.76.O63G365 1995.

Gancarz:2003:LUP

[Gan03] Mike Gancarz. Linux and the Unix Philosophy. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-273-7. xxvii + 220 pp. LCCN QA76.76.O63G364 2003. US$34.99. URL http://

www.loc.gov/catdir/description/
els031/2003051482.html;
http://www.loc.gov/catdir/
toc/els031/2003051482.html.

Ganten:2004:DGLb

[Gan04] Peter H. Ganten. Debian-GNU-Linux-Powerpack. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-66384-3. ???? pp. LCCN ????

Ganti:2008:PAL

[Gan08] Ashwin Ganti. Plan 9 authentication in Linux. Operating Systems Review, 42(5):27–33, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Gart:1986:TAU

[Gar86] Mitchell Gart. Targeting Ada to 68000/Unix. In USENIX Association [USE86c], pages 261– 274.

Gart:1990:ABU

[Gar90] Mitchell Gart. Ada and binary UNIX standards. In USENIX [USE90d], pages 353– 362. LCCN QA76.8.U65 U82 1990.

Gardner:1991:LU

[Gar91]  James Gardner.  Learn-

ing UNIX. Howard W. Sams, Indianapolis, IN 46268, USA, 1991. ISBN 0-672-30001-X. xxii + 595 pp. LCCN QA76.76.O63G37 1991. US$39.95. With disks containing MSDOS stimulation of UNIX (MSK Tools) …. A good tutorial/reference book for those without constant access to UNIX ….

Garber:1997:NBC

[Gar97]  Lee Garber.  News briefs:

Crucial compromise launches digital TV. US encryption agreement in jeopardy. warning issued about UNIX flaw. WIPO discusses cyberspace copyrights. IT issues could threaten European Monetary Union. COBOL programmers in demand again. chip alliance formed. semiconductor film grown in space. survey reveals security fears and vulnerability. taxing the Internet. Computer, 30(2):18, 19, 22, February 1997. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garber:1999:NBL

[Gar99]  Lee Garber.  News briefs:

Linux support ranges from GUI to Big Blue; new XML standards; partnership for advanced computational infrastructure. Computer, 32(5): 20–22, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r5020.pdf.

Gaskins:1992:PPMb

[Gas92a] Tom Gaskins. PEXlib Programming Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-028-7. xlv + 1105 pp. LCCN QA76.76.W56 G37 1992. US$44.95.

Gaskins:1992:PPMa

[Gas92b] Tom Gaskins. PHIGS Programming Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-85-4 (paperback), 0-937175-92-7 (hardcover). xxxix + 908 pp. LCCN QA76.76.W56 G37 1992. US$42.95 (softcover), US$52.95 (hardcover).

                                        URL http://www.oreilly.
                                        com/catalog/9780937175859.

Gatwood:2001:CCL

[Gat01] David A. Gatwood. Cleaning the cleaner for the Linux logstructured file system. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2001.

Gauthier:1981:UUS

[Gau81] Richard Gauthier. Using the Unix System. Reston Publishing Company, Reston, VA, USA, 1981. ISBN 0-8359-8162-2 (paperback), 0-8359-8164-9 (hardcover). xiv + 297 pp. LCCN QA76.6 .G37.

                                        URL http://books.google.
                                        com/books?id=LIFQAAAAMAAJ.

Gaur:1998:TSU

[Gau98] Nalneesh Gaur. Time synchronization in a UNIX/NT environment. Sys Admin: The Journal for UNIX Systems Administrators, 7(4):27–30, 33, April 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

                                        www.samag.com/.

Gay:1987:PKL

[Gay87] David M. Gay. Pictures of Karmarkar’s linear programming algorithm. Computing Science Technical Report 136, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 30, 1987. i + 46 pp.

Gay:1990:USS

[Gay90] David M. Gay. Usage summary for selected optimization routines. Computing Science Technical Report 153, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 16, 1990. i + 21 pp.

Gayley:1993:TSX

[Gay93] Todd Gayley. Tech support:

The X front end and UNIX installation. Mathematica Journal, 3(4):??, Fall 1993. CO-DEN ???? ISSN 1047-5974 (print), 1097-1610 (elec-

tronic). URL http://www.
mathematica-journal.com/
issue/v3i4/columns/techsupport/
index.html;   http://www.
mathematica-journal.com/
issue/v4i2/columns/techsupport/
54-61_gayley42.mj.pdf.

Gay:1999:STY

[Gay99] Warren W. Gay. Sams’ teach yourself Linux programming in 24 hours: complete starter kit. Macmillan Computer Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31582-3. xviii + 526 pp. LCCN QA76.76.O63 G398 1999. Edited by Brian Gill, Ron Gallagher, and Gus Miklos.

Gay:2000:AUP

[Gay00a] Warren Gay. Advanced UNIX programming. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31990-X. xvii + 604 pp. LCCN QA76.76.O63 G3886 2000.

Gay:2000:LSP

[Gay00b] Warren W. Gay. Linux Socket Programming by Example. Que Corporation, Indianapolis, IN, USA, 2000. ISBN ???? ???? pp. LCCN ????

Galvin:1992:SAU

[GB92] James M. Galvin and David M. Balenson. Security aspects of a UNIX PEM implementation. In USENIX [USE92e], pages 119–131. ISBN 1-880446-46-4. LCCN ????

Greenspan:2001:MPD

[GB01] Jay Greenspan and Brad Bulger. MySQL/PHP applications. Professional mindware. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-3537-4 (paperback). xxi + 596 pp. LCCN QA76.73.S67 G73

                                        2001. URL http://www.loc.
                                        gov/catdir/bios/wiley044/
                                        00053995.html;   http://
                                        www.loc.gov/catdir/description/
                                        wiley038/00053995.html;
                                        http://www.loc.gov/catdir/
                                        toc/wiley021/00053995.html.

Gunter:1997:WNU

[GBG97] David Gunter, Steven Burnett, and Lola Gunter. Windows NT and UNIX Integration Guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882395-1. xxvi + 450 pp. LCCN QA76.76.O63G868 1997. Includes CD-ROM.

Gill:1986:CUC

[GBH86] Helen Gill, Rebecca Bowerman, and Chuck Howell. A comparison of UNIX and CAIS system facilities. In USENIX Association [USE86e], pages 275–293. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Goodheart:1994:MGE

[GC94a] Berny Goodheart and James Cox. The Magic Garden Explained: The Internals of UNIX System V Release 4, an Open Systems Design. Prentice-Hall, 1994. ISBN 0-13-098138-9. xxviii + 664 pp. LCCN QA76.76.O63 G6633 1994. US$38.00. Probably a good companion to [Bac86a] …. Covering the internals, system calls, kernal of System V Release 4 ….

Goodheart:MGE94

[GC94b] Berny Goodheart and James Cox. The Magic Garden Explained: The Internals of UNIX System V Release 4, an Open Systems Design. Prentice-Hall, 1994. ISBN 0-13-098138-9. xxviii + 664 pp. LCCN QA76.76.O63 G6633 1994. US$38.00. Probably a good companion to Bach’s book, The Design of the UNIX Operating System, …. Covering the internals, system calls, kernal of System V Release 4 ….

Gebai:2018:SAK

[GD18] Mohamad Gebai and Michel R. Dagenais. Survey and analysis of kernel and userspace tracers on Linux: Design, implementation, and overhead. ACM Computing Surveys, 51 (2):26:1–26:??, June 2018. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Golub:1990:UAP

[GDFR90] David Golub, Randall Dean, Alessandro Forin, and Richard Rashid. UNIX as an application program. In USENIX [USE90c], pages 87–96. LCCN QA76.8.U65 U81 1990.

Gehani:1984:AAI

[Geh84] Narain Gehani. Ada, an advanced introduction: including reference manual for the Ada programming language. Prentice-Hall software series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-003997-7. various pp. LCCN QA76.73.A35 G43 1984; QA76.73.A35G43 1984.

Gehani:1986:DFT

[Geh86] Narain Gehani. Document Formatting and Typesetting on the UNIX System. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1986. ISBN 0-9615336-0-9. xv + 364 pp. LCCN Z52.5.U54 G43 1986. US$33.95.

Gehani:1987:DFT

[Geh87a]  Narain Gehani.  Document

Formatting and Typesetting on the UNIX System. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, second edition, 1987. ISBN 0-13-938325-5, 0-9615336-2-5. xv + 377 pp. LCCN Z52.5.U54G43 1987. US$40.90.

Gehani:1987:UAP

[Geh87b] Narain Gehani. UNIX Ada programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-938325-5 (paperback). xix + 310 pp. LCCN QA76.73.A35 G437 1987. Revised edition of [Geh84].

Gehani:2003:BLL

[Geh03] Narain Gehani. Bell Labs: life in the crown jewel. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 2003. ISBN 0-929306-27-9. xi + 258 pp. LCCN TK415.B45

G44 2003. URL http://www.
loc.gov/catdir/toc/fy043/
2002012730.html.

Gellis:1978:MMD

[Gel78] H. S. Gellis. MRCS and the MRCS 2 Data Base System. Technical Memorandum 1308 (TM 78-9442-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 26, 1978. ?? pp.

Gemmell:2005:MMA

[Gem05] Keith Gemmell. Making music on the Apple Mac. PC, Merton, UK, 2005. ISBN 1-870775-95-3 (paperback). 103 pp.

Gerkin:1982:IUV

[Ger82] Fred Gerkin. Introduction to UNIX — videotape. In Usr Group [Usr82], pages 108–?? Abstract only.

Gerber:1985:TMR

[Ger85] A. J. Gerber. The trouble with mutual recursion in concurrent Euclid. ACM SIGPLAN Notices, 20(8):64–70, August 1985. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

German:1992:CLC

[Ger92] Hallett German. Command language cookbook for mainframes, minicomputers, and PC’s: DOS/OS/2 batch language, Clist, DCL, Perl, and REXX. Van Nostrand Reinhold, New York, NY, USA, 1992. ISBN 0-442-00801-5. xiv + 352 pp. LCCN QA 76.7 G45 1992.

Gerbarg:2002:ASM

[Ger02] Louis G. Gerbarg. Advanced synchronization in Mac OS X: Extending UNIX to SMP and real-time. In USENIX [USE02], pages 37–45. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736 2002. URL http://www.

usenix.org/publications/
library/proceedings/bsdcon02/
gerbarg.html.

German:2003:GPC

[Ger03]  Daniel M. German.  The

GNOME project: a case study of open source, global software development. Software Process: Improvement and Practice, 8(4):201–215, October 2003. CODEN SPIPFL. ISSN 1077-4866 (print), 1099-1670 (electronic).

Gomberg:1998:CLS

[GES98] Michail Gomberg, Rémy Evard, and Craig Stacey. A comparison of large-scale software installation methods on NT and UNIX. In USENIX [USE98c], pages 37–47. ISBN 1-880446-96-0. LCCN TK5105.5 .L37

1998.  URL http://www.
usenix.org/publications/
library/proceedings/lisa-
nt98/gomberg.html.

Gettys:1986:PIW

[Get86]  James Gettys.  Problems

implementing window systems in UNIX. In USENIX Association [USE86e], pages 89–97. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Gettys:1991:XWS

[Get91] James Gettys. X Windows System design principles, 1991. VHS format. Title on container: X window system design principles. Recorded on February 11, 1991. The X Window System has become widely accepted by many manufacturers and users of computer systems. X provides network transparent access to display servers, allowing local and remote client programs to access a user’s display. X is used on high-performance workstation displays as well as X terminals, and client programs run on everything from micro- to supercomputers. This talk emphasizes system design principles, using X as an example.

MathlabGroup:1980:IMV

[GF80] Mathlab Group and Richard J. Fateman. An introduction to MACSYMA for VAX/UNIX. Computing Services, University of California, Berkeley, Berkeley, CA, USA, 1980. 13 pp.

Griswold:1983:IPL

[GG83] Ralph E. Griswold and Madge T. Griswold. The Icon Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-449777-5. xviii + 313 pp. LCCN QA76.73.I19 G74 1983.

Griswold:1986:IIP

[GG86] Ralph E. Griswold and Madge T.

Griswold. The Implementation of the Icon Programming Language. Princeton University Press, Princeton, NJ, USA, 1986. ISBN 0-691-08431-9. x + 336 pp. LCCN QA76.73.I19 G76 1986. US$39.50.

Griswold:1990:IPL

[GG90] Ralph E. Griswold and Madge T.

Griswold. The Icon Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1990. ISBN 0-13-447889-4. xv + 367 pp. LCCN QA76.73.I19 G74 1990.

Gaman:1991:PE

[GG91] William A. Gaman and W. A.

Giovinazzo. PHIGS by Example. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97555-1, 3-540-97555-1. viii + 218 pp. LCCN T385 .G35 1991.

Gay:1999:SAF

[GG99] David M. Gay and Eric Grosse. Self-adapting Fortran 77 machine constants: Comment on Algorithm 528. ACM Transactions on Mathematical Software, 25(1):123– 126, March 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://cm.bell-labs.
com/who/ehg/mach/d1mach.
ps;   http://www.acm.org/
pubs/citations/journals/
toms/cgi-bin/TOMSbibget?
Gay:1999:SAF;   http://www.
acm.org/pubs/citations/journals/
toms/cgi-bin/TOMScitation?
Fox:1978:AFP;   http://www.
acm.org:80/pubs/citations/
journals/toms/1999-25-1/

p123-gay/. See [FHS78a].

Gaur:1989:EPE

[GGJ89] Yogesh Gaur, Vincent A. Guarna, and David Jablonowski. An environment for performance experimentation on multiprocessors. Technical Report CSRD 865, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, April 1989. 8 pp.

Ghemawat:2003:GFS

[GGL03] Sanjay Ghemawat, Howard Gobioff, and Shun-Tak Leung. The Google File System. In ACM [ACM03], pages 29–43. ISBN 1-58113-757-

5. ISSN 0163-5980 (print), 1943-586X (electronic). URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://uclibs.org/PID/34720.

ACM order number 534030.

Goodman:1993:SPG

[GGM93] Michele Ann Goodman, Manoj Goyal, and Robert A. Massoudi. Solaris porting guide.

SunSoft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-030396-8. lxxiv + 790 pp. LCCN QA76.76.O63 G66347 1993. ????

Gunther:2002:LUR

[GGZ02] Karsten Günther, Kester Grelck, and Thorsten Zilm, editors. Linux User Referenz: [Linux Drucksysteme: BSD und CUPS: Steuerung des Prozess-Systems: CDs erstellen und Textdateien bearbeiten: Pager, Formatierung und die glimpse-Tools]. (German) [Linux User Reference: Linux Printing Systems: BSD and CUPS: System Process Control: CD creation and text file processing; Pagination, Formatting and glimpse tools]. mitp Verlag, Bonn, Germany, 2002. ISBN 3-8266-0709-0. 1055 (est.) pp. LCCN ????

Goldstein:1975:SPP

[GH75] A. J. Goldstein and A. D.

Hall. Solutions to a problem in power series reversion. SIAM Journal on Mathematical Analysis, 6(1):192–198, February 1975. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic). URL https:

//dl.acm.org/doi/10.1137/
0506019.

Hall:1974:SPEb

[GH4 ] A. J. Goldstein and A. D. Hall.

Solutions to a problem in power series reversion. Computing Science Technical Report 30b, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974 (??). ?? pp.

Graham:1982:ETD

[GHS82] Susan L. Graham, Robert R. Henry, and Robert A. Schulman. An experiment in table driven code generation. ACM SIGPLAN Notices, 17(6):32– 43, June 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gianone:1990:UMK

[Gia90] Christine M. Gianone. Using MS-DOS KERMIT: connecting your PC to the Electronic World. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 1-55558-048-3. xxv + 244 pp. LCCN TK5105.9 .G5 1990.

Giampaolo:1999:PFS

[Gia99] Dominic Giampaolo. Practical file system design with the BE file system. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1999. ISBN 1-55860-497-9. x + 237 pp. LCCN QA76.9.F5 G49 1999. US$34.95. Includes comparison with Apple Macintosh, Linux, and Microsoft Windows file systems.

Gigante:2006:HPS

[Gig06] Mike Gigante. High-performance storage solutions on IA-64 Linux. In Anonymous [Ano06a], page ?? ISBN ???? LCCN

                                        ????  URL http://www.

ice.gelato.org/;   http:/
/www.ice.gelato.org/about/
oct06_presentations.php.

Gilmore:1988:PBU

[Gil88]  John Gilmore.  Porting

Berkeley Unix through the GNU C compiler. Submitted to USENIX 1988, but rejected as “not research”. Posted by its author to the TUHS mailing list on 21 May 2020., February 19, 1988.

URL http://mcvoy.com/lm/
papers/porting-berkeley.
pdf;   https://minnie.tuhs.
org/pipermail/tuhs/2020-
May/021204.html.

Gill:1991:FBU

[Gil91] P. J. Gill. Finally, a binary UNIX PC standard. Datamation, 36(24):59–60, December 1991. CODEN DTMNAT. ISSN 0011-6963.

Gillies:1993:RLU

[Gil93a] Betsy Gillies. Review: Learning the UNIX Operating System. ;login: the USENIX Association newsletter, 18(6):48– ??, November/December 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Gilster:1993:INE

[Gil93b]  Paul Gilster.  The Inter-

net Navigator — The Essential Guide to Network Exploration for the Individual Dialup User. Wiley-Interscience, New York, NY, USA, 1993.

ISBN 0-471-59782-1. xxiv + 470 pp. LCCN TK5105.875.I57 G55 1993. US$24.95.

Gill:1996:UTH

[Gil96a] Philip J. Gill. UNIX takes the high road — while NT looks to be a great success in the low-end of the server market, more UNIX-based enterprise servers are being used as IS managers turn away from mainframes and legacy applications. Datamation, 42(12):58– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Gill:1996:UVG

[Gil96b] Philip J. Gill. UNIX vendors get server fever — as NT moves in on the PC server market, commercial UNIX vendors move upscale. Datamation, 42 (11):90–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

GDT:2007:GUM

[GIM07] GIMP Documentation Team. GIMP user manual: GNU image manipulation program user manual. SoHoBooks, ????, USA, 2007. ISBN 1-4414-1932-2. 653 pp. LCCN T385 .G5426 2009.

Gingrich:1983:HFB

[Gin83] P. S. Gingrich. Human factors and behavioral science: The UNIX Writer’s Workbench software: Results of a field study. The Bell System Technical Journal, 62(6): 1909–1921, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1909.
pdf.

Gircys:1988:UUC

[Gir88]  Gintaras R. Gircys.  Un-

derstanding and Using COFF. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-31-5. xvii + 176 pp. LCCN QA76.8.U65 G57 1988. US$21.95.

Gajewska:1975:SSU

[GJ75] H. Gajewska and Steven C.

Johnson. Some statistics on the usage of the C language. Technical Memorandum 1102 (TM 75-1273-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 24, 1975. ?? pp.

Gift:2008:PUL

[GJ08] Noah Gift and Jeremy M.

Jones. Python for Unix and Linux system administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2008. ISBN 0-596-51582-0. xix + 433 pp. LCCN QA76.73.P98 G54 2008.

Ganapathy:2005:APA

[GJJ05] Vinod Ganapathy, Trent Jaeger, and Somesh Jha. Automatic placement of authorization hooks in the Linux security modules framework. In Meadows and Syverson [MS05], pages 330–339. ISBN 1-59593-226-7. LCCN QA76.9.A25. ACM order number 459050.

Gabbe:1977:QNA

[GJLB77] J. D. Gabbe, C. N. Judice, T. B. London, and J. P. Beyer. QUIK: A novel approach to computerized directory assistance services. Technical Memorandum 1245, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 12, 1977. ?? pp.

Gillett:1997:UMC

[GK97] Richard Gillett and Richard Kaufmann. Using the Memory Channel Network — using a cluster of standard PCI-based servers with a low-cost network to improve communication performance. IEEE Micro, 17(1): 19–25, January/February 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gettys:1990:XWS

[GKM90] J. Gettys, P. L. Karlton, and S. McGregor. The X Window System, version 11. Software —Practice and Experience, 20 (S2):35–67, October 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Greenberg:2021:USP

[GKV21] Michael Greenberg, Konstantinos Kallas, and Nikos Vasilakis. Unix shell programming. In ACM, editor, HotOS ’21: Workshop on Hot Topics in Operating Systems, Ann Arbor, Michigan, 1 June, 2021– 3 June, 2021, pages 104–110. ACM Press, New York, NY, USA, June 2021. ISBN 1-4503-

8438-2. URL https://sigops.
org/s/conferences/hotos/
2021/papers/hotos21-s06-
greenberg.pdf.

Goldstein:1975:CFT

[GL75] A. J. Goldstein and A. B. Lesk.

Common feature techniques for discrete optimization. Computing Science Technical Report 27, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1975. ?? pp.

Goldstein:1976:CFT

[GL76] A. J. Goldstein and A. B. Lesk.

Common feature techniques for discrete optimization. In The proceedings of the thirteenth design automation conference on Design automation — DAC ’76, No. 13, pages 232–244. ACM Press, New York, NY, USA, 1976.

Gehani:1988:DFT

[GL88] Narain Gehani and Steven Lally. Document formatting and typesetting on the UNIX system. Vol. 2: grap, mv, ms, and troff. Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1988. ISBN 0-9615336-3-3. xiii + 304 pp. LCCN Z52.5.U54 G431 1988. US$30.95.

Gallmeister:1991:EEP

[GL91] Bill O. Gallmeister and Chris Lanier. Early experience with POSIX 1003.4 and POSIX 1003.4 A. Proceedings — Real-Time Systems Symposium, pages 190–198 (of ix + 307), 1991. CODEN PRSYEA. ISBN 0-8186-2450-7. IEEE catalog number 91CH3090-8.

Gropp:1994:SUT

[GL94] W. Gropp and E. Lusk. Scalable Unix tools on parallel processors. In IEEE [IEE94a], pages 56–62. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994.

Lehey:2002:WBB

[gL02] Greg “groggy” Lehey. Why BSD is better than Linux. IBM Linux Technology Center, Ozlabs, Brisbane, Queensland, Australia, February 7, 2002.

                                        URL http://www.lemis.com/
                                        linux.conf.au.pdf. 197-slide

conference talk.

Gelphman:2006:PQP

[GL06] David Gelphman and Bunny Laden. Programming with Quartz: 2D and PDF graphics in Mac OS X. Elsevier Morgan Kaufmann, Amsterdam, The Netherlands, 2006. ISBN 0-12-369473-6. xxviii + 668 pp. LCCN T385 .G392

                                        2006. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0625/2005052186-d.html;
                                         http://www.loc.gov/catdir/

enhancements/fy0625/2005052186-
t.html.

Glasser:1977:GPL

[Gla77a] A. L. Glasser. A generalpurpose library for PWB/ UNIX (B.11). Technical Memorandum 1157 (TM 77–9144-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 27, 1977. 7 pp.

Glasser:1977:SSC

[Gla77b] A. L. Glasser. A style sheet for C programming, issue I. Technical Memorandum 1143, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 28, 1977. ?? pp.

Glasser:1978:ESC

[Gla78] A. L. Glasser. The evolution of a source code control system. Technical Memorandum 1304 (TM 78-3111-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 12, 1978. ?? pp.

Glasser:1979:DAU

[Gla79] A. L. Glasser. A debugging aid for UNIX C programmers. Technical Memorandum 1413 (MF 79-3116-8), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 28, 1979. ?? pp.

Glass:1991:PUA

[Gla91]  David N. Glass.  Porting

UNIX applications to DOS. Dr. Dobb’s Journal of Software Tools, 16(11):68, 70–72, 76, 78, 80, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Glazier:1992:BPW

[Gla92] Bill Glazier. The “Best principle”: Why OpenGL is emerging as the 3D graphics standard. Computer Graphics World, 15 (4):116, April 1992.

Glahn:1997:FJT

[Gla97] Kay Glahn. Free 4 Java: Tools für das Internet. JDKs für Windows, Mac und UNIX. Computer und Literaturverlag, ????, July 1997. ISBN 3-932311-30-2 (??invalid ISBN??). ???? pp. LCCN ???? 39 DM.

Glass:1999:NBL

[Gla99] Robert L. Glass. News briefs: Loyal opposition: Of open source, Linux . . . . . . and hype. IEEE Software, 16(1):128–??, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s1128.pdf.

Glass:2002:LMB

[Gla02]  Brett Glass.  Log moni-

tors in BSD UNIX. In USENIX [USE02], pages 131– 142. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

                                        2002.  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/bsdcon02/
                                        glass.html.

Guerrero:1993:IEA

[GLG93] R. Guerrero, L. Leguizamon, and R. Gallard. Implementation and evaluation of alternative process schedulers in MINIX. Operating Systems Review, 27(1):79–100, January 1993. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Glick:1990:GLP

[Gli90] P. Glick. A guide to the Lp printer spooler. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Glines:1992:DU

[Gli92] Steven Glines. Downsizing to UNIX. New Riders Publishing, Carmel, IN, USA, 1992. ISBN 1-56205-074-5. xix + 506 pp. LCCN QA76.76.O63 G588 1992. A book for people involved or interested in downsizing mainframe data-processing to a distributed UNIX network.

Guillemont:1991:SGM

[GLOR91] Marc Guillemont, Jim Lipkis, Doug Orr, and Marc Rosier. A second-generation micro-kernel based UNIX: Lessons in performance and compatibility. In USENIX [USE91d], pages 13– 22. LCCN QA 76.76 O63 U84

Gropp:2003:BCC

[GLS03] William Gropp, Ewing Lusk, and Thomas Lawrence Sterling, editors. Beowulf cluster computing with Linux. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, second edition, 2003. ISBN 0-262-69292-9 (paperback). xxxix + 618 pp. LCCN QA76.58.B46 2003; QA76.58.B46.

Grampp:1984:UOS

[GM84] Frederick T. Grampp and Robert H. Morris. UNIX operating system security. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1649– 1672, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

                                        URL https://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=6771912.

Grampp:1990:USS

[GM90] F. T. Grampp and R. H. Morris. UNIX system security. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

                                        1990. URL http://doc.cat-
                                        v.org/unix/v10/10thEdMan/
                                        v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Garfinkel:2002:BCA

[GM02] Simson Garfinkel and Michael K.

Mahoney. Building Cocoa Applications: a Step-by-Step Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00235-

1. xxv + 620 pp. LCCN QA76.76.O63 G37 2002 Stacks. US$44.95. URL http://

safari.oreilly.com/0596002351;
http://www.oreilly.com/
catalog/buildcocoa.

Gregg:2011:DTO

[GM11] Brendan Gregg and Jim Mauro.

Dynamic tracing in Oracle Solaris, Mac OS X, and FreeBSD. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2011. ISBN 0-13-209151-8. x + 285 pp. LCCN QA76.9.D43 G74 2011.

Giering:1993:IAF

[GMB93] E. W. Giering, F. Mueller, and T. P. Baker. Implementing Ada 9X features using POSIX threads: Design issues. In ACM [ACM93], pages 214–228. ISBN 0-89791-621-2. LCCN ???? ACM Order No. 825930.

Gajewska:1990:WXO

[GMM90] Hania Gajewska, Mark S. Manasse, and Joel McCormack. Why X is not our ideal window system. Software—Practice and Experience, 20(S2): 137–171, October 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ganger:2000:SUS

[GMSP00] Gregory R. Ganger, Marshall Kirk McKusick, Craig A. N. Soules, and Yale N. Patt. Soft updates: a solution to the metadata update problem in file systems. ACM Transactions on Computer Systems, 18(2):127–153, May 2000. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocs/
2000-18-2/p127-ganger/.

Guniguntala:2008:RCU

[GMTW08] D. Guniguntala, P. E. McKenney, J. Triplett, and J. Walpole. The read-copy-update mechanism for supporting real-time applications on shared-memory multiprocessor systems with Linux. IBM Systems Journal, 47(2):221–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/472/guniguntala.html.

Guanciale:2016:PSM

[GNDB16] Roberto Guanciale, Hamed Nemati, Mads Dam, and Christoph Baumann. Provably secure memory isolation for Linux on ARM. Journal of Computer Security, 24(6):793– 837, ???? 2016. CODEN JC-SIET. ISSN 0926-227X (print), 1875-8924 (electronic).

Gansner:1988:DPD

[GNV88] E. R. Gansner, S. C. North, and K. P. Vo. DAG: a program that draws directed graphs. Software—Practice and Experience, 18(11):1047– 1062, November 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Gansner:1990:DPD

[GNV90] E. R. Gansner, S. C. North, and K. P. Vo. Dag — a program for drawing directed graphs. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Groening:1986:NNM

[GO86] K. Groening and C. Ohsendoth.

NEMO: a nicely modified YACC. ACM SIGPLAN Notices, 21(4):58–66, April 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goerzen:2000:LPB

[Goe00] John Goerzen. Linux Programming Bible. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-4657-0. 837 pp. LCCN QA76.76.O63 G6343 2000. US$39.99.

Goelker:2007:GPI

[Goe07] Klaus Goelker. GIMP 2 for photographers: image editing with open source software. Rockynook, Santa Barbara, CA, USA, 2007. ISBN 1-933952-03-2. ix + 186 pp. LCCN TR267.5.G56.

Goldsmith:1979:VTP

[Gol79] L. D. Goldsmith. A virtual terminal protocol for UNIX. Technical Memorandum 1451 (TM 79-5412-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 6, 1979. ?? pp.

Goldberg:1995:LCR

[Gol95] Howard L. Goldberg. The Linux commands and resources directory. Cruising the Cutting Edge Pub., Houston, TX, USA, 1995. ISBN 0-00-017192-1. 360 pp. LCCN QA76.76.O63 G6373 1995. Version 1.2.4.

Goldman:1998:USC

[Gol98] Ross Goldman. UNIX serial communications. Sys Admin: The Journal for UNIX Systems Administrators, 7(3):8, 10, 12– 13, March 1998. CODEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/

.

Gomez:1985:UIP

[Gom85] Julian E. Gomez. A UNIX image production pipeline. ;login: the USENIX Association newsletter, 10(4):40–42, October/November 1985. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Goos:1990:IPS

[Goo90] Anke Goos. Is Plan 9 scifi or UNIX for the future? UNIX/World, 7(10):61–??, October 1, 1990. ISSN 0739-5922.

Gookin:2007:PGN

[Goo07] Dan Gookin. Programmer’s guide to NCurses. Wiley, New York, NY, USA, 2007. ISBN 0-470-10759-6. xx + 556 pp. LCCN X07.E03449.

Gopin:1979:DPP

[Gop79] A. M. Gopin. Development of a program for parts inventory control. Technical Memorandum 1478 (TM 79-3237-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 19, 1979. ?? pp.

Gorin:1981:CNS

[Gor81] Ralph E. Gorin. Computer networking at Stanford: Current status and future plans for Ethernet at Stanford University. Report, Computer Science Department, Stanford University, Stanford, CA 94305-2192, USA, January 1981. 20

pp. URL https://stacks.
stanford.edu/file/druid:
yx961bt1370/yx961bt1370.
pdf.

Gosling:1981:UE

[Gos81] James Gosling. UNIX EMACS.

Report, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, December 1981.

Gottleber:2002:BU

[Got02] Tim Gottleber. Bulletproof UNIX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-093028-8. 600 pp. LCCN QA76.76.O63 G69 2003.

Gough:2004:IGG

[Gou04] Brian Gough. An introduction to GCC for the GNU compilers gcc and g++. Network Theory Ltd., Bristol, UK, 2004. ISBN 0-9541617-9-3. iv + 116 pp. Foreword by Richard M. Stallman.

Gove:2008:SAP

[Gov08] Darryl Gove. Solaris application programming. Sun Microsystems Press, Palo Alto, CA, USA, 2008. ISBN 0-13-813455-3 (hardcover). xxii + 468 pp. LCCN QA76.76.O63 G688 2008.

Gray:1988:PAL

[GP88] W. H. Gray and A. K. Powers. Project accounting on a large-scale UNIX system. In USENIX Association [USE88d], pages 7–12. ISBN ???? LCCN ????

Griswold:1968:SPL

[GPP68] Ralph E. Griswold, James F. Poage, and Ivan P. Polonsky. SNOBOL4 programming language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1968. ISBN 0-13-815357-4. x + 221 pp. LCCN QA76.73.S6 G75 1968.

Griswold:1971:SPL

[GPP71] R. E. Griswold, J. E. Poage, and I. P. Polonsky. The SNOBOL4 programming language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1971. ISBN 0-13-815373-6. xii + 256 pp. LCCN QA76.73.S6 G75 1971.

Greenberg:1976:NPR

[GR76] M. E. Greenberg and C. S.

Roberts. Numeric pattern recognition data base for UNIX systems. Technical Memorandum 1085 (TM 76-1353-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 5, 1976. ?? pp.

Grampp:1979:CPT

[Gra79a] Frederick T. Grampp. cq — a program for testing C compilers. Technical Memorandum 1426 (TM 79-2524-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 14, 1979. ?? pp.

Grampp:1979:PTW

[Gra79b] Frederick T. Grampp. Program testing — work to be done. Technical Memorandum 1439, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 14, 1979. ?? pp.

Grampp:1979:SCI

[Gra79c] Frederick T. Grampp. Support for C on IBM computers.

Technical Memorandum 1348, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 25, 1979. ?? pp.

Graham:1984:UPM

[Gra84]  Gig Graham.  UNIX for

the people! A manifesto for appealing to the masses. UNIX Review, 2(9):20–22, 24, December 1984. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/unixreview_1984dec.
pdf.

Gray:1987:MDE

[Gra87] David F. Gray. A microprocessor development environment based on the Amsterdam compiler kit, EMACS and UNIX. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 19 (4):30–35, December 1987. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Graham:1988:PHB

[Gra88] Ken Graham. Pssssttt! Hey buddy, you wanta buy UNIX source for $89? C Users Journal, 6(6):43–??, June 1988. ISSN 0898-9788.

Gosling:1989:NBI

[GRA89a] James Gosling, David S. H. Rosenthal, and Michelle Arden. The NeWS Book: an introduction to the Network/extensible Window System. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-96915-2. vi + 235 pp. LCCN QA76.76.W56 A731 1989.

Graham:1989:TUS

[Gra89b] Ken Graham. UNIX System Programming, by Keith Haviland and Ben Salama. C Users Journal, 7(10):36–??, October 1989. ISSN 0898-9788.

Graff:1994:PRM

[Gra94] Mark Graff. PEXlib: a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-176066-1. 425 pp. LCCN QA76.76.W56 G83 1993.

Graham:1995:SXI

[Gra95] John R. Graham. Solaris 2.x: internals and architecture. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-911876-3. xv + 222 pp. LCCN QA76.76.O63 G72 1995. ????

Gray:1997:ICU

[Gra97] John Shapeley Gray. Interprocess Communications in UNIX: The Nooks and Crannies. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-186891-8. x + 364 pp. LCCN QA76.76.O63 G729 1997. URL http://www.

phptr.com.

Gray:1998:LSC

[Gra98a] Bob Gray. Loading source code UNIX on the PC. ;login: the USENIX Association newsletter, 23(5):??, August 1998. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1998-8/source.html.

Gray:1998:PHS

[Gra98b] Bob Gray. PC hardware for source code UNIX. ;login: the USENIX Association newsletter, 23(4), June 1998. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1998-6/source.html.

Gray:1998:SCUc

[Gra98c]  Bob Gray.  Source code

UNIX: Help a friend get online, cheap. ;login: the USENIX Association newsletter, 23(7): ??, December 1998. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1998-12/sourcecode.
html.

Gray:1998:SCUb

[Gra98d] Bob Gray. Source code UNIX on the PC. ;login: the USENIX Association newsletter, 23(6):??, October 1998. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic). URL http://www.

usenix.org/publications/
login/1998-10/sourcecode.
html.

Gray:1998:ICU

[Gra98e]  John Shapeley Gray.  In-

terprocess Communications in UNIX: The Nooks and Crannies. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-899592-3. xi + 462 pp. LCCN QA76.76.O63G729 1998. US$45.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0138995923.html.

Gray:1998:SCUa

[Gra98f] Robert Gray. Source code UNIX for PCs. ;login: the USENIX Association newsletter, 23(2):??, April 1998. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1998-4/source.html.

Gray:1999:SCUd

[Gra99a]  Bob Gray.  Source code

UNIX. ;login: the USENIX Association newsletter, 24(6): ??, December 1999. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1999-12/features/sourcecode.
html.

Gray:1999:SCUa

[Gra99b] Bob Gray. Source code UNIX:

Embedding source code UNIX in the product. ;login: the USENIX Association newsletter, 24(2):??, April 1999. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1999-4/sourcecode.
html.

Gray:1999:SCUb

[Gra99c] Bob Gray. Source code UNIX: Security on a source code UNIX system. ;login: the USENIX Association newsletter, 24(3):??, June 1999. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1999-6/features/sourcecode.
html.

Gray:1999:SCUc

[Gra99d]  Bob Gray.  Source code

UNIX: What’s your data worth? ;login: the USENIX Association newsletter, 24(4): ??, August 1999. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1999-8/features/sourcecode.
html.

Gray:2000:PTS

[Gra00a] Bob Gray. Performance tuning with source code UNIX. ;login: the USENIX Association newsletter, 25(2): ??, April 2000. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2000-4/features/performance.
html.

Gray:2000:SCU

[Gra00b]  Bob Gray.  Source code

UNIX. ;login: the USENIX Association newsletter, 25(1): ??, February 2000. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2000-2/features/sourcecode.
html.

Gray:2000:TOS

[Gra00c]  Bob Gray.  Teaching op-

erating systems with source code UNIX. ;login: the USENIX Association newsletter, 25(4):??, July 2000. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2000-7/features/teaching.
html.

Gray:2001:LTF

[Gra01a] Bob Gray. A logging and tracing facility for an embedded source code UNIX product. ;login: the USENIX Association newsletter, 26 (2):??, April 2001. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2001-04/pdfs/gray.
pdf.

Gray:2001:SOR

[Gra01b] Peter D. Gray. SUS, an object reference model for distributing UNIX super user privileges.

In USENIX [USE01a], page ?? ISBN 1-880446-05-7. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa2001/tech/
gray.html.

Granor:2003:OTY

[Gra03] Tamar E. Granor. OOoSwitch: 501 Things You Wanted to Know About Switching to OpenOffice.org from Microsoft Office. Hentzenwerke Publishing, Milwaukee, WI 53217-5361, USA, 2003. ISBN 1-930919-36-0. 310 (est.) pp. LCCN ???? US$49.95. URL

http://www.hentzenwerke.
com/catalogpricelists/ooo501.

htm. Edited by Scott Carr and Sam Hiser.

Grant:2004:LNG

[Gra04] Rickford Grant. Linux for non-geeks: a hands-on, projectbased, take-it-slow guidebook. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-034-8. xx + 335 pp. LCCN QA76.76.O63 G723 2004eb. URL http:

//www.loc.gov/catdir/toc/
ecip0411/2003024732.html;
 http://www.oreilly.com/
catalog/1593270348/;   http:
//www.oreilly.com/catalog/
9781593270346. Includes two

CD-ROMs with Fedora.

Grant:2005:LME

[Gra05] Rickford Grant. Linux made easy: the official guide to Xandros 3 for everyday users. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-057-7. xxvi + 463 pp. LCCN QA76.76.O63 G7245 2005; QA76.76.O63 .G7245 2005eb.

URL http://www.oreilly.
com/catalog/9781593270575.

Granneman:2006:LP

[Gra06a]  Scott Granneman.  Linux

Phrasebook. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32909-3; 0-672-32838-0 (paperback). xvii + 382 pp. LCCN QA76.76.O63 G72 2006. EUR 13.90.

Grant:2006:ULN

[Gra06b]  Rickford Grant.  Ubuntu

Linux for non-geeks: a painfree, project-based, get-things-done guidebook. No Starch Press, San Francisco, CA, USA, 2006. ISBN 1-59327-118-2. xx + 334 pp. LCCN QA76.76.O63 G7246 2006eb.

URL http://site.ebrary.
com/lib/ucmerced/Doc?id=
10137836;   http://site.
ebrary.com/lib/ucsc/Doc?
id=10137836;   http://www.
loc.gov/catdir/enhancements/
fy0715/2006015576-d.html;
 http://www.loc.gov/catdir/
toc/ecip0613/2006015576.
html.

Greenberg:1981:UOS

[Gre81] Robert B. Greenberg. The Unix operating system and the Xenix standard operating environment. BYTE Magazine, 6 (6):248–264, June 1981. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Greenberg:1982:IUS

[Gre82a] Robert B. Greenberg. Is UNIX as a standard doomed? In Usr Group [Usr82], pages 280–281. Abstract only.

Greenberg:1982:USD

[Gre82b] Robert B. Greenberg. Is UNIX as a standard doomed? In USENIX [USE82], pages 280– 281. Abstract only.

Greenbaum:1992:FUC

[Gre92a] Joshua Greenbaum. France’s UNIX Champion. UNIX/World, 9(9):61–??, September 1992. ISSN 0739-5922.

Gregory:1992:PM

[Gre92b] Keith D. Gregory. Programming with Motif. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 0-387-97877-1 (New York), 3-540-97877-1 (Berlin). xxii + 664 pp. LCCN QA76.76.W56 G84 1992.

Greenbaum:1993:RUU

[Gre93] Joshua Greenbaum. Riding UNIX into the U.S. Market. UNIX/World, 10(3):51–??, March 1993. ISSN 0739-5922.

Greening:1996:TLE

[Gre96] Tony Greening. Teaching and learning essential computer science skills: the UNIX example. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 28 (2):21–24, June 1996. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Gregory:2000:SS

[Gre00] Peter H. Gregory. Solaris security. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-096053-5. xliii + 291 pp. LCCN QA76.9.A25G75 1999. US$39.99. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130960535.html;   http:
//www.sun.com/books/catalog/
gregory/index.html.

Gregory:2002:SCS

[Gre02] Peter H. Gregory. Sun Certified System Administrator for Solaris 8 Study Guide. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-040933-2. xxviii + 416 pp. LCCN QA76.3.G753 2002.

Greengard:2011:MRM

[Gre11] Samuel Greengard. In memoriam: Robert Morris, 1932– 2011. Communications of the Association for Computing Machinery, 54(9):17, September 2011. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Gregg:2019:BPT

[Gre19] Brendan Gregg. BPF Performance Tools: Linux System and Application Observability. Addison-Wesley Professional, ????, 2019. ISBN 0-13-655482-2, 0-13-658887-5. LCCN ???? Foreword by Alexei Starovoitov, creator of the new BPF.

Griffith:2000:GGP

[Gri00] Arthur Griffith. GNOME/GTK+ Programming Bible. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-4640-6. xxvi + 850 pp. LCCN QA76.9.U83 G75

2000.  US$39.99.  URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/wiley022/00101008.html.

Griffith:2001:KQP

[Gri01a] Arthur Griffith. KDE/QT programming bible. I D G Books Worldwide, Indianapolis, IN, USA, 2001. ISBN 0-7645-4682-1. xxv + 753 pp. LCCN QA76.73.C153 .G7426 2001.

Grinzo:2001:PBP

[Gri01b] Lou Grinzo. Programmer’s bookshelf: Professional Linux Programming. Dr. Dobb’s Journal of Software Tools, 26 (4):151–152, April 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

Grimmett:2013:BRP

[Gri13] Richard Grimmett. Beagle-Bone robotic projects: create complex and exciting robotic projects with the BeagleBone Black. Community experience distilled. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78355-932-2, 1-78355-933-0 (e-book), 1-306-28021-4 (e-book). iii + 228 pp. LCCN TJ211.45 .G75 2013.

URL http://proquest.tech.
safaribooksonline.de/9781783559329.

Grimmett:2014:MBR

[Gri14] Richard Grimmett. Mastering BeagleBone robotics: master the power of the Beagle-Bone Black to maximize your robot-building skills and create awesome projects. Community experience distilled. Packt Publishing, Birmingham, UK, 2014. ISBN 1-78398-890-8 (paperback), 1-78398-891-6 (ebook), 1-322-55407-2 (e-book). iii + 217 pp. LCCN TJ211.45 .G7566 2014.

Grottola:1993:UAU

[Gro93] Michael G. Grottola. The UNIX audit: using UNIX to audit UNIX. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-025127-4. xii + 174 pp. LCCN QA76.76.O63 G77 1993.

Grune:1986:GPC

[Gru86] Dick Grune. Generic packages in C. ACM SIGPLAN Notices, 21(8):31–39, August 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garwood:1989:US

[GS89] Michael W. Garwood and Andrew E. Schweig. UNIX streams. Dr. Dobb’s Journal of Software Tools, 14(1):50, 53, 55, 56, 58, 60, January 1989. CODEN DDJOEB. ISSN 1044-789X.

Garfinkel:1991:PUS

[GS91] Simson Garfinkel and Gene Spafford. Practical UNIX Security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-72-2. xxvii + 483 pp. LCCN QA76.76.O63 G38 1992. US$29.95. Highly recommended. Simply the best book in this field.

Garfinkel:1996:PUI

[GS96] Simson Garfinkel and Gene Spafford. Practical UNIX & Internet Security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1996. ISBN 1-56592-148-8. xxix + 971 pp. LCCN QA76.76.O63G38 1996. US$39.95. URL http://www.

                                        ora.com/gnn/bus/ora/item/
                                        pus2.html;   http://www.
                                        oreilly.com/catalog/puis.

Glaeser:1999:OGO

[GS99] Georg Glaeser and Hellmuth Stachel. Open geometry: OpenGL + advanced geometry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98599-9. xii + 377 pp. LCCN T385 .G576 1998. US$39.95. Includes CD-ROM.

Geiger:2000:AUS

[GS00a] Robert Geiger and John Schweitzer. Automating UNIX security monitoring. Sys Admin: The Journal for UNIX Systems Administrators, 9(11): 32, 34–36, November 2000. CO-DEN SYADE7. ISSN 1061-

2688.  URL http://www.
samag.com/.

Grinzo:2000:PBG

[GS00b] Lou Grinzo and Jacques Surveyer. Programmer’s bookshelf: GNU tools and process patterns. Dr. Dobb’s Journal of Software Tools, 25(12):159– 160, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Garg:2001:TOA

[GS01] Rajat P. Garg and Ilya Sharapov. Techniques for Optimizing Applications: High Performance Computing. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-093476-3. xliv + 616 pp. LCCN QA76.88 .G37

2002. URL books/apt.pdf;
http://www.sun.com/books/
catalog/garg.html/index.

html. Part No. 806-6380-10 June 2001, Revision 01.

Garg:2002:TOA

[GS02] Rajat P. Garg and Ilya Sharapov. Techniques for optimizing applications: high performance computing. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-093476-3. xliii + 616 pp. LCCN QA76.88 .G37 2002. URL

http://www.sun.com/books/
catalog/garg.html/index.
html.

Giaglis:2012:DEP

[GS12] G. M. Giaglis and D. Spinellis. Division of effort, productivity, quality, and relationships in FLOSS virtual teams: Evidence from the FreeBSD project. J.UCS: Journal of Universal Computer Science, 18(19):2625– ??, ???? 2012. CODEN ???? ISSN 0948-6968. URL

http://www.jucs.org/jucs_
18_19/division_of_effort_
productivity.

Gine:2003:DMC

[GSHL03] Francesc Giné, Francesc Solsona, Porfidio Hernández, and Emilio Luque. Dealing with memory constraints in a nondedicated Linux cluster. The International Journal of High Performance Computing Applications, 17(1):39–48, Spring 2003. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Gettys:1990:XCL

[GSN90] James Gettys, Robert W. Scheifler, and Ron Newman. Xlib: C Language X Interface (X version 11, release 4). Silicon Press, 25 Beverly Road, Summit, NJ 07901, USA, 1990. ISBN 0-929306-03-1. x + 295 pp. LCCN QA76.73.C15 G47 1990. US$29.95.

Garfinkel:2003:PUI

[GSS03] Simson Garfinkel, Gene Spafford, and Alan Schwartz. Practical Unix & Internet Security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 0-596-00323-4. xxix + 954 pp. LCCN QA76.76.O63 G38 2003. US$54.95. URL http://www.

oreilly.com/catalog/puis3.

Glasser:1978:CRU

[GT78a] A. L. Glasser and D. W. Ting.

A core-resident UNIX system for PDP-11’s. Technical Memorandum 1264 (TM 78-3111-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 5, 1978. ?? pp.

Glasser:1978:EPU

[GT78b] A. L. Glasser and D. W. Ting.

Enlarging PDP-11 UNIX table space using supervisor mode. Technical Memorandum 1257 (TM 78-3111-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 20, 1978. ?? pp.

Glasser:1979:RFS

[GT79] A. L. Glasser and D. W. Ting.

A remote file system driver for UNIX. Technical Memorandum 1411 (TM 79-3111-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 31, 1979. ?? pp.

Gilly:1992:UN

[Gt92a] Daniel Gilly and the staff of O’Reilly & Associates. UNIX in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1992. ISBN 1-56592-001-5. LCCN QA76.76.O63 G55 1992.

Gilly:1992:UNB

[Gt92b] Daniel Gilly and the staff of O’Reilly & Associates, Inc. UNIX in a Nutshell. Berkeley edition — a Desktop Quick Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-20-X. 272 pp. LCCN QA76.76.O63 U544 1989.

Gilly:1992:UNS

[GT92c] Daniel Gilly and The staff of O’Reilly and Associates. UNIX in a Nutshell: System V Edition. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, June 1992. ISBN 1-56592-001-5. 444 pp. LCCN QA76.76.O63 G55 1992. US$9.95. URL http:/

/www.oreilly.com/catalog/
unutv.

Guenther:1987:REU

[Gue87] G. R. Guenther. Running 7th edition UNIX programs on a VAX in compatibility mode. Operating Systems Review, 21 (1):30–33, January 1987. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Guerrieri:2019:HSP

[Gue19] Alex Guerrieri. Hands-on system programming with Go: build modern and concurrent applications for Unix and Linux systems using Golang. Packt Publishing, Birmingham, 2019. ISBN 1-78980-336-5 (e-book), 1-78980-407-8. xii + 437 pp. LCCN QA76.73.G63. URL http://

proquest.safaribooksonline.
com/?fpi=9781789804072.

Gunther:2007:BKG

[Gün07] Karsten Günther. Bash: kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 3-89721-533-0. 144 pp. LCCN ???? EUR 9.90.

IEN168

[Gur81] Robert F. Gurwitz. VAX-UNIX networking support project implementation description, January 1981. URL http://www.

cis.ohio-state.edu/htbin/
ien/ien168.html.

Guyot:2000:GBL

[Guy00] Paul Guyot. Are GPL and BSD License really compatible? World-Wide Web document., 2000. URL

http://www.kallisys.org/
bsd-lite/bsd-gpl/?lg=en.

Gerndt:2003:LSC

[GVK^{+}03] Andreas Gerndt, Thomas Van Reimersdahl, Torsten Kuhlen, Christian Bischof, Ingolf Hörschler, Matthias Meinke, and Wolfgang Schröder. Large-scale CFD data handling in a VR-based otorhinolaryngological CAS-system using a Linux-cluster. The Journal of Supercomputing, 25(2):143–154, June 2003. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//ipsapp009.kluweronline.
com/content/getfile/5189/
44/5/abstract.htm;   http:
//ipsapp009.kluweronline.
com/content/getfile/5189/
44/5/fulltext.pdf;   http:
//www.springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=25&issue=2&spage=
143.

Gertner:1998:TOD

[GW98] Abigail S. Gertner and Bonnie L. Webber. TraumaTIQ: Online decision support for trauma management. IEEE expert: intelligent systems and their applications, 13(1):32–39, January–February 1998. CO-DEN IEEXE7. ISSN 0885-9000.

Gwozdz:1978:PMB

[Gwo78] P. W. Gwozdz. Proposal for a microprocessor based systems development laboratory. Technical Memorandum 1305 (TM 78-4562-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 29, 1978. ?? pp.

Garfinkel:1994:UHH

[GWS94] Simson Garfinkel, Daniel Weise, and Steven Strassmann, editors. The UNIX-haters handbook. IDG Books, San Mateo, CA, USA, 1994. ISBN 1-56884-203-1. xxxvii + 329 pp. LCCN QA 76.76 .O63 U54518 1994. URL http:/

/research.microsoft.com/~
daniel/uhh-download.html;
http://research.microsoft.
com/~daniel/unix-haters.
html.

Gwyn:1990:BRL

[Gwy90] Douglas A. Gwyn. Book review: Life with UNIX — A Guide For Everyone. ;login: the USENIX Association newsletter, 15(3):13–??, May/ June 1990. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Gusella:1984:TAN

[GZ84a] Riccardo Gusella and Stefano Zatti. TEMPO — a network time controller for a distributed Berkeley UNIX system. In Software Tools Users Group [Sof84], pages 78–85. LCCN QA76.8.U65 U83 1984.

Gusella:1984:TNT

[GZ84b] Riccardo Gusella and Stefano Zatti. TEMPO — a network time controller for a distributed Berkeley UNIX system. In USENIX [USE84a], pages 78–

Gusella:1989:ACS

[GZ89] R. Gusella and S. Zatti. The accuracy of the clock synchronization achieved by TEMPO in Berkeley UNIX 4.3BSD. IEEE Transactions on Software Engineering, 15(7):847– 853, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=29484.

Gatla:2023:UPM

[GZXZ23] Om Rameshwar Gatla, Duo Zhang, Wei Xu, and Mai Zheng. Understanding persistentmemory-related issues in the Linux kernel. ACM Transactions on Storage, 19(4):36:1– 36:??, November 2023. CO-DEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

URL https://dl.acm.org/
doi/10.1145/3605946.

Hedrick:1979:RAM

[H^{+}79] E. L. Hedrick et al. RCM AS/ MERT performance study and general design guidelines. Technical Memorandum 1360 (TM 79-4145-1, TM 79-2524-1, TM 79-3451-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 20, 1979. ?? pp.

Hill:2007:OUBa

[H^{+}07a] Benjamin Mako Hill et al., editors. The official Ubuntu book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-243594-2 (paperback). xxxiv + 412 pp. LCCN QA76.76.O63 O34348 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006016172.html.

Hill:2007:OUBb

[H^{+}07b] Benjamin Mako Hill et al., editors. The official Ubuntu book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2007. ISBN 0-13-235413-6 (paperback). xlii + 463 pp. LCCN QA76.76.O63 O34348 2007a. URL http:

//www.loc.gov/catdir/toc/
ecip0717/2007018467.html.

Hemenway:1984:PSS

[HA84] Kathleen Hemenway and Helene Armitage. Proposed syntax standard for UNIX system commands. In USENIX Association [USE84b], pages 89–98. ISBN none. LCCN QA76.8.U65 U55 1984.

Heslop:1990:MS

[HA90] Brent Heslop and David Angell. Mastering SunOS. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1990. ISBN 0-89588-683-9. xxxii + 588 pp. LCCN QA76.76.O63G37 1991. US$29.95. A good comprehensive hands-on text to SunOS and OpenWindows.

Heslop:1993:MS

[HA93] Brent D. Heslop and David F.

Angell. Mastering Solaris 2. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1993. ISBN 0-7821-1072-X. xl + 899 pp. LCCN QA76.76.O63 H478 1993.

Hodel:1995:BSAa

[HA95a] Alan E. Hodel and /AIXtra. The best of /AIXtra: an eclectic UNIX anthology: Volume I. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, August 1995. ISBN 0-13-328626-6 (paperback). xvi + 295 pp. LCCN QA76.76.O63 B475 1995. US$55.00. URL

http://www.prenhall.com/
ptrbooks/ptr_0133286266.
html.

Hodel:1995:BSAb

[HA95b] Alan E. Hodel and /AIXtra. The best of /AIXtra: an eclectic UNIX anthology: Volume II. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, November 1995. ISBN 0-13-339839-0. xvi + 311 pp. LCCN QA76.76.O63B475

1995.  US$41.00.  URL

http://www.prenhall.com/
ptrbooks/ptr_0133398390.
html.

Hodel:1996:BSA

[HA96] Alan E. Hodel and /AIXtra. The best of /AIXtra: an eclectic UNIX anthology: Volume III. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, August 1996. ISBN 0-13-494444-5. 288 pp. LCCN QA76.76.O63B475

1995.  US$40.00.  URL

http://www.prenhall.com/
ptrbooks/ptr_0134944445.
html.

Hitz:1998:MNU

[HAB^{+}98] Dave Hitz, Bridget Allison, Andrea Borr, Rob Hawley, and Mark Muhlestein. Merging NT and UNIX filesystem permissions. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

usenix.org/publications/
library/proceedings/usenix-
nt98/hitz.html.

Habraken:2001:SCH

[Hab01]  Joe Habraken.  StarOffice

5.2 Calc Handbook. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-029389-X. xix + 346 pp. LCCN HF5548.2. US$29.99.

URL http://www.phptr.com/
ptrbooks/ptr_013029389X.
html;   http://www.sun.com/
books/catalog/habraken/index.
html.

Haddad:2004:RUS

[Had04]  Ibrahim Haddad.  Review:

UNIX systems programming: Communication, concurrency and theory. Linux Journal, 2004(118):??, February 2004. CODEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Haemer:1989:UUC

[Hae89] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 14(6):31– 45, November/December 1989. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1990:UUCa

[Hae90a] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(1): 34–43, January/February 1990. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1990:UUCb

[Hae90b] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(3):25– 54, May/June 1990. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1990:UUCc

[Hae90c] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(4):21– 58, July/August 1990. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1990:UUCd

[Hae90d] Jeffrey S. Haemer. An update on UNIX and C standards activity. ;login: the USENIX Association newsletter, 15(5):16– 24, September/October 1990. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1990:UUR

[Hae90e] Jeffrey S. Haemer. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 15(6):17–39, November/ December 1990. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1991:UURa

[Hae91a] Jeffrey S. Haemer. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(1): 24–34, January/February 1991. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1991:UURb

[Hae91b] Jeffrey S. Haemer. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(2):22– 25, March/April 1991. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Haemer:1992:CUD

[Hae92] Jeff Haemer. Can UNIX designers learn anything from PCs? ;login: the USENIX Association newsletter, 17(3): 17–18, May/June 1992. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Hagen:1983:CSN

[Hag83] Teus Hagen. Cookbook for setting up a national UNIX systems users group. EUUG Newsletter, 3(2):7–10, Summer 1983. CODEN EONLE8. ISSN 1011-4211.

VonHagen:2000:IRL

[Hag00] Bill Von Hagen. Installing Red Hat Linux 7. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31826-1. 198 pp. LCCN QA76.76.O63 V66 2000.

Hahn:1993:SGU

[Hah93] Harley Hahn. A Student’s Guide to UNIX. McGraw-Hill, New York, NY, USA, 1993. ISBN 0-07-025511-3. xxxi + 633 pp. LCCN QA76.76.O63 H34 1993. Highly recommended. A superb introduction to UNIX … In a clear and lively language, the author tells the novice users everything they want to know about UNIX and the Internet — covering UNIX commands, utilities, shells, vi, X-Window, e-mail, netnews, ftp, gopher, etc …. It should be an excellent textbook for any UNIX introductory course … “No experience necessary!” “UNIX is fun.”.

Hahn:1994:OCU

[Hah94a] Harley Hahn. Open Computing’s UNIX Unbound. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-882050-2. xxxi + 792 pp. LCCN QA76.76.O63 H343 1994. US$27.95.

Hahn:1994:UU

[Hah94b] Harley Hahn. UNIX unbound.

Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-882050-2 (paperback). xxxi + 792 pp. LCCN QA76.76.O63 H343 1994.

Hahn:2009:HHG

[Hah09] Harley Hahn. Harley Hahn’s guide to Unix and Linux. McGraw-Hill Higher Education, Boston, MA, USA, 2009. ISBN 0-07-313361-2 (paperback). xxxiv + 926 pp. LCCN QA76.76.O63 H3378

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0808/2007052525-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0808/2007052525-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0810/2007052525-t.html.

Haight:1977:AAP

[Hai77] R. C. Haight. Administrative advice for PWB/UNIX (1.2). Technical Memorandum 1195, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1977. ?? pp.

Haight:1978:PRU

[Hai78a] R. C. Haight. Pre-release UNIX/TS. Technical Memorandum 1274 (MF 78-8234-57), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 4, 1978. ?? pp.

Haight:1978:UFS

[Hai78b] R. C. Haight. UNIX file system conversion. Technical Memorandum 1319 (MF 78-8234-83), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 30, 1978. ?? pp.

Haight:1978:UTF

[Hai78c]  R. C. Haight.  UNIX/TS

file system backup. Technical Memorandum 1273 (MF 78-8234-42), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 11, 1978. ?? pp.

Haight:1978:UTR

[Hai78d] R. C. Haight. UNIX/TS Release 1.0. Technical Memorandum 1321 (MF 78-8234-82), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 30, 1978. ?? pp.

Haight:1979:UTD

[Hai79] R. C. Haight. UNIX/TS development plan. Technical Memorandum 1484, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 10, 1979. ?? pp.

Hall:1969:PFI

[Hal69] A. D. Hall, Jr. A portable FOR-TRAN IV subset. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1969. ?? pp.

Hall:1972:MMP

[Hal72] Andrew D. Hall. The M6 macro processor. Computing Science Technical Report 2, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 12, 1972. ii + 13 pp.

Hall:1975:SPE

[Hal75] Andrew D. Hall, Jr. Solving a problem in eigenvalue approximation with a symbolic algebra system. SIAM Journal on Computing, 4(2):163– 174, ???? 1975. CODEN SMJ-CAT. ISSN 0097-5397 (print), 1095-7111 (electronic).

Hall:1979:DMW

[Hal79] J. T. Hall. Department 3621 MAXIUNIX workplan. Technical Memorandum 1395, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1, 1979. ?? pp.

Halpern:2000:CSL

[Hal00] Pablo Halpern. The C++ Standard Library from scratch. Jesse Liberty’s from scratch programming series. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2128-7. viii + 351 pp. LCCN QA76.73.C153 H347 2000.

Haletky:2005:DLD

[Hal05] Edward Haletky. Deploying Linux on the desktop. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2005. ISBN 1-55558-328-8. ???? pp. LCCN QA76.76.O63 H34345 2005.

Hallinan:2006:ELP

[Hal06] Christopher Hallinan. Embedded Linux primer: a practical, real-world approach. Prentice Hall open source software development series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-167984-8 (paperback). xxix + 537 pp. LCCN QA76.76.O63 H34462 2006. US$44.99. URL http:

//www.loc.gov/catdir/toc/
ecip0612/2006012886.html.

Hallinan:2007:ELP

[Hal07] Christopher Hallinan. Embedded Linux primer: a practical, real-world approach. Prentice Hall open source software development series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-167984-8 (paperback). xxix + 537 pp. LCCN QA76.76.O63 H34462 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0612/2006012886.html.

Halvorsen:2011:XIK

[Hal11] Ole Henry Halvorsen. OS X and iOS Kernel Programming. Apress, Berkeley, CA, USA, 2011. ISBN 1-4302-3536-5. 300 (est.) pp. LCCN ????

Hall:1974:SPEa

[Hal4 ] Andrew D. Hall, Jr. Solving a problem in eigenvalue approximation with a symbolic algebra system. Computing Science Technical Report 30a, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974 (??). ?? pp.

Hamilton:1978:VFS

[Ham78] L. L. Hamilton. Virtual file systems on MERT. Technical Memorandum 1343 (TM 78-5523-9), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 22, 1978. ?? pp.

Hamburger:1987:UHC

[Ham87] Sheldon Hamburger. UNIX in health care: Medical laboratories — a case study. In USENIX Association [USE87e], pages 192–195. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Hammerslag:1990:FLB

[Ham90] David H. Hammerslag. Faust library browser: user’s manual. Technical Report CSRD 961, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, January 1990. 15 + 4 pp.

Hammel:2007:AGG

[Ham07]  Michael J. Hammel.  The

artist’s guide to GIMP effects: creative techniques for photographers, artists, and designers. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-121-2 (paperback). xii + 348 pp. LCCN T385 .H329558

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0728/2007001652-d.html;

                                         http://www.loc.gov/catdir/
                                        toc/ecip078/2007001652.html.

Hanson:1977:RES

[Han77] David R. Hanson. RATSNO — an experiment in software adaptability. Software—Practice and Experience, 7(5):625– 630, September/October 1977. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hansen:1986:TVU

[Han86] August Hansen. vi — The UNIX Screen Editor — a User’s Guide. Robert J. Brady Co., Bowie, MD 20715, USA, 1986. ISBN 0-89303-928-4. xvii + 230 pp. LCCN QA76.76.T49 H36 1986.

Hancock:2001:TUF

[Han01] Steven M. Hancock. Tru64 Unix file system administration handbook. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2001. ISBN 1-55558-227-3 (paperback). xxvii + 533 pp. LCCN QA76.76.O63 H34465 2001.

Hansteen:2008:BPN

[Han08] Peter N. M. Hansteen. The book of PF: a no-nonsense guide to the OpenBSD firewall. No Starch Press, San Francisco, CA, USA, 2008. ISBN 1-59327-165-4 (paperback). xix + 159 pp. LCCN TK5105.585 .H385 2008.

Hansteen:2011:BPN

[Han11] Peter N. M. Hansteen. The book of PF: a no-nonsense guide to the OpenBSD firewall. No Starch Press, San Francisco, CA, USA, second edition, 2011. ISBN 1-59327-274-X. xx + 188 pp. LCCN TK5105.585 .H385 2011.

Harris:1979:NP

[Har79] M. O. Harris. The Nspeak program. Technical Memorandum 1419 (TM 79-1227-18), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 3, 1979. ?? pp.

Harland:1980:HSD

[Har80] David M. Harland. High speed data acquisition: Running a realtime process and a timeshared system (UNIX) concurrently. Software—Practice and Experience, 10(4):273–281, April 1980. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harvey:1983:UL

[Har83] Brian Harvey. UNIX logo. In Association [Ass83], pages 145– 150.

Harrison:1990:NNP

[Har90] Bradford T. Harrison. Networking: NAS, POSIX and RISC/VMS on the network. The DEC Professional, 9(12): 94–??, November 1990. CO-DEN DECPDJ. ISSN 0744-

Hare:1992:HUP

[Har92a] Chris Hare. How UNIX password controls work. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 53–??, May/June 1992. CO-DEN SYADE7. ISSN 1061-2688.

Hare:1992:UFS

[Har92b] Chris Hare. The UNIX file system debugger — FSDB. Sys Admin: The Journal for UNIX Systems Administrators, 1(2): ??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Hardin:1993:STF

[Har93] Reid Hardin. SAS/TUTOR: fundamentals of the SAS system course guide, OS/2 and Windows (Release 6.08), Open-VMS and UNIX (X Window Environments), Windows NT (Release 6.09). SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. vi + 139 pp.

Hardenbergh:1994:BAP

[Har94a] Jan “Yon” Hardenbergh. Building Applications with PEXlib. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-012535-0. xiv + 542 pp. LCCN QA76.76.W56 H36 1994.

Hare:1994:UPM

[Har94b] Chris Hare. The UNIX process management system. Sys Admin: The Journal for UNIX Systems Administrators, 3(4):

75–??, July/August 1994. CO-DEN SYADE7. ISSN 1061-2688.

Harrison:1995:UHU

[Har95] Mark Harrison. The USENET Handbook: a User’s Guide to Netnews. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, May 1995. ISBN 1-56592-101-1. xiv + 372 pp. LCCN TK5105.875.I57 H534 1995. US$24.95.

Hare:1996:IU

[Har96] Chris Hare. Inside UNIX. New Riders Publishing, Carmel, IN, USA, second edition, 1996. ISBN 1-56205-491-0. xxix + 1212 pp. LCCN QA76.76.O63 I562 1996.

Harvel:2000:UWH

[Har00] Lonnie Harvel. Unix and Windows 2000 handbook: planning, integration, and administration. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-025493-2. xvii + 679 pp. LCCN QA76.76.O63 U58 2000.

Hare:2001:RUPa

[Har01a] Chris Hare. Revisiting UNIX password controls – part 1. Sys Admin: The Journal for UNIX Systems Administrators, 10(10):30, 32–34, October 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Hare:2001:RUPb

[Har01b] Chris Hare. Revisiting UNIX password controls – part 2. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):35–38, November 2001. CODEN SYADE7. ISSN 1061-2688.

Hart:2001:MXW

[Har01c] David L. Hart. Mac OS X Web server handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-032715-8. xxvi + 395 pp. LCCN QA76.76.O63 H3555 2001.

Harding:2011:GTM

[Har11] Ted Harding. A guide to typesetting mathematics using GNU eqn, January 19,

                                        2011.  URL https://web.
                                        archive.org/web/*/http:/
                                        /www.zen89632.zen.co.uk/
                                        Groff/Eqn/eqnguide.pdf.

Haskell:1978:IGZ

[Has78] B. G. Haskell. Implementation of GR-Z graphics on UNIX. Technical Memorandum 1298, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 8, 1978. ?? pp.

Hauser:1994:SUL

[Hau94] Carl Hauser. Speeding up UNIX login by caching the initial environment. In USENIX [USE94d], pages 117–124. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

                                        URL  http://www.usenix.
                                        org/publications/library/

proceedings/lisa94/hauser.
html.

Haverbeke:2011:EJM

[Hav11] Marijn Haverbeke. Eloquent JavaScript: a modern introduction to programming. No Starch Press, San Francisco, CA, USA, 2011. ISBN 1-59327-282-0. xiv + 205 pp. LCCN QA76.73.J39 HAV 2011; QA76.73.J39 H38

2009. URL http://www.loc.
gov/catdir/enhancements/
fy1012/2010032246-b.html;
http://www.loc.gov/catdir/
enhancements/fy1012/2010032246-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1107/2010032246-t.html.

Hawley:1985:UTP

[Haw85] Michael J. Hawley. UNIX tools for a personal database. In USENIX Association [USE85d], pages 333–341. LCCN QA76.8.U65 U8 1985.

Hawley:1986:MMS

[Haw86] Michael Hawley. MIDI music software for UNIX. In USENIX Association [USE86c], pages 1– 12.

Hawley:1987:MMSa

[Haw87] Michael Hawley. More MIDI software for UNIX. In USENIX Association [USE87a], pages 201–208. ISBN ???? LCCN ????

Hawkins:2000:LDR

[Haw00] Scott Hawkins. Linux desk reference. Open source technology series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-016391-0. xxvii + 545 pp. LCCN QA76.76.O63 H386 2000.

Hawkins:2002:LDR

[Haw02] Scott Hawkins. Linux desk reference. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-13-061989-2. xxi + 583 pp. LCCN QA76.76.O63 H386 2002.

Hayden:1977:SMP

[Hay77] D. F. Hayden. Some MERT performance measurements in transaction oriented processing. Technical Memorandum 1182 (MF 77-3124-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 28, 1977. ?? pp.

Haynes:2006:OM

[Hay06] Tom Haynes. OpenSolaris: The model. ;login: the USENIX Association newsletter, 31(2):5–8, April 2006. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2006-04/pdfs/haynes.
pdf.

Hazlewood:1999:UAM

[Haz99] Victor Hazlewood. UNIX accounting magic. Sys Admin: The Journal for UNIX Systems Administrators, 8(3): 8, 10, 12, 14, 16, 18, March 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Holbrook:1982:HCR

[HB82] Bernard D. Holbrook and W. Stanley Brown. A history of computing research at Bell Laboratories (1937–1975). Computing Science Technical Report 99, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1982. 29 pp. URL https:/

/archive.computerhistory.
org/resources/access/text/
2022/08/102804421-05-01-
acc.pdf;   https://static.
googleusercontent.com/media/
research.google.com/en//
pubs/archive/94.pdf. See

also final book version [BHM84].

Hearn:2004:CGO

[HB04] Donald Hearn and M. Baker.

Computer Graphics with OpenGL. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2004. ISBN 0-13-015390-7. xxii + 857 pp. LCCN T385 .H395 2004. US$80.00.

Hill:2006:OUB

[HBB^{+}06] Benjamin Mako Hill, Jono Bacon, Corey Burger, Jonathan Jesse, and Ivan Krstić, editors. The official Ubuntu book. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-243594-2 (paperback). 320 (est.) pp. LCCN QA76.76.O63 O34348 2006. US$34.99. URL http://

vig.prenhall.com/catalog/
academic/product/0,1144,
0132435942,00.html;   http:
//www.loc.gov/catdir/toc/
ecip0613/2006016172.html.

Harji:2013:OTL

[HBB13] Ashif S. Harji, Peter A. Buhr, and Tim Brecht. Our troubles with Linux Kernel upgrades and why you should care. Operating Systems Review, 47(2): 66–72, July 2013. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Harchol–Balter:1997:EPL

[HBD97] Mor Harchol-Balter and Allen B. Downey. Exploiting process lifetime distributions for dynamic load balancing. ACM Transactions on Computer Systems, 15(3):253–285, August 1997. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tocs/1997-15-3/p253-harchol-
                                        balter/.

Herder:2006:MHR

[HBG^{+}06a] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. MINIX 3: a highly reliable, self-repairing operating system. Operating Systems Review, 40 (3):80–89, July 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Herder:2006:MSP

[HBG^{+}06b] Jorrit N. Herder, Herbert Bos, Ben Gras, Philip Homburg, and Andrew S. Tanenbaum. Modular system programming in MINIX 3. ;login: the USENIX Association newsletter, 31(2):19–28, April 2006. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic). URL http://

www.minix3.org/;   http:/
/www.usenix.org/publications/
login/2006-04/openpdfs/herder.
pdf.

Hu:2018:SCS

[HBH18] Hanyang Hu, Cor-Paul Bezemer, and Ahmed E. Hassan. Studying the consistency of star ratings and the complaints in 1 & 2-star user reviews for top free cross-platform Android and iOS apps. Empirical Software Engineering, 23(6):3442– 3475, December 2018. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-018-9604-y.

Harrington:1996:IPV

[HBMT96] Daniel T. Harrington, James P.

Bound, John J. McCann, and Matt Thomas. Internet Protocol version 6 and the Digital UNIX implementation experience. Digital Technical Journal of Digital Equipment Corporation, 8(3):5–22, December 1996. CODEN DTJOEL. ISSN 0898-901X.

URL  ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v8n3/Internet_Protocol_Version_
6_an_07jan1997DTJN01P8.ps;
 http://www.digital.com:

                                        80/info/DTJN01/DTJN01AH.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJN01/DTJN01HM.
                                        HTM;   http://www.digital.
                                        com:80/info/DTJN01/DTJN01P8.
                                        PS;   http://www.digital.
                                        com:80/info/DTJN01/DTJN01PF.
                                        PDF;   http://www.digital.
                                        com:80/info/DTJN01/DTJN01SC.
                                        TXT.

Hwu:1989:IFE

[HC89] Wen-Mei W. Hwu and Pohua P. Chang. Inline function expansion for compiling C programs. ACM SIGPLAN Notices, 24(7):246–257, July 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        pldi/73141/p246-chang/.

Hecht:1987:US

[HCC^{+}87] M. S. Hecht, M. E. Carson, C. S. Chandersekaran, R. S. Chapman, L. J. Dotterrer, V. D. Gligor, W. D. Jiang, A. Johri, G. L. Luckenbaugh, and N. Vasudevan. UNIX without the superuser. In USENIX Association [USE87d], pages 243–256.

Hagimont:1994:PSO

[HCF^{+}94] Daniel Hagimont, P.-Y. Chevalier, A. Freyssinet, S. Krakowiak, S. Lacourte, J. Mossière, and X. Rousset de Pina. Persistent shared object support in the Guide system: Evaluation and related work. ACM SIGPLAN Notices, 29(10):129–144, October 1994. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hwang:1982:ULC

[HCG^{+}82] K. Hwang, W. J. Croft, G. H.

Goble, B. W. Wah, F. A. Briggs, W. R. Simmons, and C. L. Coates. A Unix-based local computer network with load balancing. Computer, 15(4):55– 66, April 1982. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hewson:1985:IAP

[HCN85] Denise Hewson, Gregory Cullen, and Alan Nugent. Integral array processing in a multiprocessor UNIX environment. In USENIX Association [USE85d], pages 183–187. LCCN QA76.8.U65 U8 1985.

Hazlewood:1997:BLU

[HD97] Victor Hazlewood and Chris Daharsh. Backups for large UNIX installations. Sys Admin: The Journal for UNIX Systems Administrators, 6(11):8, 10, 12, 15, 16, 18, 19, November 1997. CODEN SYADE7. ISSN 1061-2688.

Hutchinson:1998:UUA

[HD98] Thomas W. Hutchinson and Ronald R. Derynck. Updating a UNIX Application Suite for the Windows NT World. Hewlett–Packard Journal, 50(1):21–??, November 1998. CODEN HPJOAX. ISSN 0018-1153. URL http:/

/www.hp.com/hpj/98nov/tc-
11-98.htm.

Howard:2002:DIA

[HD02] John S. Howard and David Deeths. Designing ISP Architectures. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-045496-6. xxvii + 360 pp. LCCN QA76.9.A73N59

2002.  URL books/isp_
bp.pdf;   http://www.sun.
com/books/catalog/nguyen/.

Part No. 816-0917-10 March 2002, Revision 01.

Hawley:1975:MMV

[HdBM75] John Alfred Hawley III and Walter de Briton Meyer. MU-NIX, a multiprocessing version of UNIX. Master’s thesis, Department of Computer Science, Naval Postgraduate School, Monterey, CA, USA, June 1975. 58

pp. URL http://hdl.handle.
net/10945/20959;   https:
//calhoun.nps.edu/handle/
10945/20959.

Heath:1998:NUP

[Hea98] Steve Heath. Newnes UNIX pocket book. Newnes, Boston, MA, USA, third edition, 1998. ISBN 0-7506-4108-8. x + 340 pp. LCCN QA76.76.O63 H43 1998.

Heisel:1995:SUF

[Hei95] M. Heisel. Specification of the Unix file system: a comparative case study. Lecture Notes in Computer Science, 936:475– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hekmatpour:1989:WMU

[Hek89] S. Hekmatpour. A window manager for UNIX. The Computer Journal, 32(1):21– 23, February 1989. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.
uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/
21.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/
22.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_32/Issue_01/tiff/
23.tif.

Hekman:1997:LND

[Hek97] Jessica Perry Hekman. Linux in a Nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-167-4. xi + 424 pp. LCCN QA76.76.O63 H453 1997. US$19.95. URL http:/

/www.oreilly.com/catalog/
9781565921672;   http://
www.oreilly.com/catalog/
linuxnut.

Heller:1990:XPMa

[Hel90a]  Dan Heller.  XView Pro-

gramming Manual, volume 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-38-2. xxviii + 557 pp. LCCN QA76.76.W56 D44 v.7 1990.

Heller:1990:XPMb

[Hel90b] Dan Heller. XView Programming Manual: for Version 11 of the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-52-8. xxviii + 642 pp. LCCN QA76.76.W56 H355 1990.

Heller:1991:MPM

[Hel91a] Dan Heller. Motif Programming Manual, volume 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-70-6. xix + 557 pp. LCCN QA76.76.W56 H465. US$39.95.

Heller:1991:XPM

[Hel91b]  Dan Heller.  XView Pro-

gramming Manual, volume 7A. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, September 1991. ISBN 0-937175-87-0. 779 pp. LCCN QA76.76.W56 H447 1990.

Hellerstein:1992:CCC

[Hel92] Joseph L. Hellerstein. Control considerations for CPU scheduling in UNIX systems. In USENIX [USE92d], pages 359– 374.

Held:1997:INU

[Hel97] Gilbert Held. Integrating NT into a UNIX environment. Sys Admin: The Journal for UNIX Systems Administrators, 6(4): 27–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Heller:2003:CDP

[Hel03] Steve Heller. C++: a dialogue:

programming with the C++ Standard Library. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-009402-1. li + 1062 pp. LCCN QA76.73.C153 H453 2003.

Hennessey:1992:WDE

[Hen92] Wade Hennessey. WCL: Delivering efficient Common Lisp applications under Unix. In ACM [ACM92], pages 260–269. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

Herman:1993:USV

[Her93] Debra Herman, editor. UNIX System V NFS administration. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-016411-9. xiv + 164 pp. LCCN QA76.76.O63 U5521125 1993.

Herold:1999:LUS

[Her99] Helmut Herold. Linux- Unix-Systemprogrammierung. Addison-Wesley, München, Germany, 1999. ISBN 3-8273-1512-3. xi + 1179 pp. LCCN ???? DM 99,90; EUR 51,08.

Hertzog:2004:DHT

[Her04] Raphaël Hertzog. Debian. Eyrolles, Paris, France, 2004. ISBN 2-212-11398-6. xi + 246 pp. LCCN ????

Hesketh:1991:PUB

[Hes91] Richard Hesketh. Perly — UNIX with buttons. Software—Practice and Experience, 21(11):1165–1187, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hesketh:XR-12-1-21

[Hes94] Richard Hesketh. The X administrator: Configuration and administration of a scalable Xbased UNIX service. The X Resource, 12(1):21–49, October 1994. CODEN XRESEA. ISBN 1-56592-069-4. ISSN 1058-5591.

Hewlett-Packard:1990:UGT

[Hew90] Hewlett–Packard Company. The Ultimate Guide to the vi and ex Text Editors. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1990. ISBN 0-8053-4460-8. LCCN QA76.76.O63 U48 1990. Another decent text on vi and ex.

Hewlett–Packard:1993:GPG

[Hew93] Hewlett–Packard Company. GL to PEXlib Porting Guide, part number b3176-90040 edition, 1993.

Hall:1993:CPP

[HG93] Robert J. Hall and Aaron J. Goldberg. Call path profiling of monotonic program resources in UNIX. In USENIX [USE93c], pages 1– 13. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/cinci93/hall.
html.

Henry:1995:SXS

[HG95a] S. Lee Henry and John R. Graham. Solaris 2.x: system administrator’s guide. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-029368-6. xiii, 320 pp. LCCN QA76.76.O63 H475 1995. US$44.95.

Horlick:1995:NVL

[HG95b] Jeffrey Horlick and Martha M.

Gray. National Voluntary Laboratory Accreditation Program: POSIX: Portable Operating System Interface. Technology Administration, National Institute of Standards and Technology, January 1995. various pp. Shipping list no.: 97-0955-M.

Haviland:1999:USP

[HGS99] Keith Haviland, Dina Gray, and Ben Salama. UNIX system programming: a programmer’s guide to software development. Addison-Wesley, Reading, MA, USA, second edition, 1999. ISBN 0-201-87758-9. xx + 350 pp. LCCN QA76.76.O63H383

1999. URL http://www.awl-
he.com/titles/11268.html.

Hitz:1986:MFS

[HH86] David Hitz and Peter Honeyman. A mail file system for Eighth Edition UNIX. In USENIX Association [USE86c], pages 391–394.

Hunter:1991:USA

[HH91] Bruce Hunter and Karen Hunter. UNIX System — Advanced Administration and Management Handbook. Macmillan Publishing Co., Inc., New York, NY, USA, 1991. ISBN 0-02-358950-7. xxiv + 422 pp. LCCN QA76.76.O63 H86 1991. Another good book on system administration.

Hunter:1994:UN

[HH94] Bruce Hunter and Karen Hunter. UNIX Networks. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-089087-1. LCCN TK5105.7 .H84 1994. Discussing common real-life issues/problems with setting up and maintaining networks ….

Hauben:1997:NHI

[HH97a] Michael Hauben and Ronda Hauben. Netizens: on the history and impact of Usenet and the Internet. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7706-6. xvi + 344 pp. LCCN TK5105.875.I57 H386 1997. URL http://

                                        www.columbia.edu/~hauben/
                                        netbook/;   http://www.columbia.

edu/~rh120/;   http://www.
loc.gov/catdir/bios/wiley046/
97005920.html;   http://
www.loc.gov/catdir/description/
wiley039/97005920.html;
http://www.loc.gov/catdir/
toc/wiley031/97005920.html.

Hauben:1997:EHI

[HH97b] Michael Hauben and Ronda Hauben. On the early history and impact of UNIX: Tools to build the tools for a new millennium. In Netizens: on the history and impact of Usenet and the Internet [HH97a], chapter 9, pages 89–114. ISBN 0-8186-7706-

6. LCCN TK5105.875.I57 H386 1997. URL http://

www.columbia.edu/~hauben/
Book_Anniversary/Netizens%
20An%20Anthology%20part%
20II.pdf;   http://www.columbia.
edu/~rh120/ch106.x09.

Hentzen:2004:LTW

[HH04] Whil Hentzen and Chris Herborth. Linux transfer for Windows power users [electronic resource]. Hentzenwerke Publishing, Whitefish Bay, WI, USA, 2004. ISBN 1-930919-42-5. LCCN QA76.76.O63 H468 2004b. URL ftp:

//uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//link.library.utoronto.
ca/eir/EIRdetail.cfm?Resources\
_\_ID=77297&T=resource.

Hudson:2005:RFL

[HH05] Paul Hudson and Andrew Hudson. Red Hat Fedora 5 Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, 2005. ISBN 0-672-32847-X (paperback). ???? pp. LCCN ???? Includes DVD.

Hudson:2006:UU

[HH06] Andrew Hudson and Paul Hudson. Ubuntu Unleashed. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32909-3 (hardcover). 800 (est.) pp. LCCN ???? US$66.99.

Hudson:2007:UU

[HH07] Andrew Hudson and Paul Hudson. Ubuntu unleashed. SAMS Publishing, Indianapolis, IN, USA, 2007. ISBN 0-672-32909-3 (paperback: CD-ROM). xxvi + 879 pp. LCCN QA76.76.O63 H8167 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        fy0707/2006286996.html.

Hudson:2008:ULU

[HH08] Andrew Hudson and Paul Hudson. Ubuntu 7.10 Linux unleashed. SAMS Publishing, Indianapolis, IN, USA, third edition, 2008. ISBN 0-672-32969-7 (paperback). xxiv + 811 pp. LCCN QA76.76.O63 H81665 2008.

Hudson:2009:UU

[HH09] Andrew Hudson and Paul Hudson. Ubuntu unleashed. SAMS Publishing, Indianapolis, IN, USA, fourth edition, 2009. ISBN 0-672-32993-X (paperback). xxiv + 826 pp. LCCN QA76.76.O63 H81666 2009. URL http:

//www.loc.gov/catdir/toc/
ecip0818/2008021150.html.

Includes DVD.

Helmke:2011:UUC

[HHH11] Matthew Helmke, Andrew Hudson, and Paul Hudson. Ubuntu unleashed: covering 10.10 and 11.04. SAMS Publishing, Indianapolis, IN, USA, sixth edition, 2011. ISBN 0-672-33344-9. ???? pp. LCCN QA76.76.O63 U36 2010.

Hudson:2010:UUH

[HHHT10] Andrew Hudson, Paul Hudson, Matthew Helmke, and Ryan Troy, editors. Ubuntu unleashed: covering 9.10 and 10.4. SAMS Publishing, Indianapolis, IN, USA, fifth edition, 2010. ISBN 0-672-33109-8. 864 (est.) pp. LCCN QA76.76.O63 U385 2010.

Howard:1991:PIP

[HHHW91] T. L. J. Howard, W. T. Hewitt, R. J. Hubbold, and K. M. Wyrwas. A Practical Introduction to PHIGS and PHIGS Plus. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-41641-7. xv + 339 pp. LCCN T385.P72 1991. US$37.95.

Hallen:1999:LW

[HHJC99] Jacob Hallén, Anders Hammarqvist, Fredrik Juhlin, and Anders Chrigström. Linux in the workplace. IEEE Software, 16(1):52–57, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

                                        URL http://computer.org/
                                        software/so1999/s1052abs.
                                        htm;   http://dlib.computer.
                                        org/so/books/so1999/pdf/
                                        s1052.pdf.

Hitz:1990:UUO

[HHLS90] David Hitz, Guy Harris, James K. Lau, and Allan M. Schwartz. Using UNIX as one component of a lightweight distributed kernel for microprocessor file servers. In USENIX [USE90d], pages 285– 296. LCCN QA76.8.U65 U82 1990.

Hill:2005:DGL

[HHV05] Benjamin Mako Hill, David B. Harris, and Jaldhar Vyas. Debian GNU/Linux 3.1 bible. Wiley, New York, NY, USA, 2005. ISBN 0-7645-7644-5. xxvi + 640 pp. LCCN QA76.76.O63 H57135 2005. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip054/2004027963.html.

Hibler:1996:HUP

[Hib96] Mike Hibler. HPBSD: Utah’s 4.3bsd port for HP9000 series machines. Web site, June 1996. URL https:/

                                        /www.flux.utah.edu/~mike/
                                        hpbsd/hpbsd.html.

Hidgson:2004:BRL

[Hid04] J. P. E. Hidgson. Book review: Linux on the Mainframe by John Eilert, Maria Eisenhaendler, Dorothea Matthaeu, and Ingol Salm, Prentice Hall Professional Technical Reference, 2003, $49.99, ISBN: 0-13-101415-3. ACM Queue: Tomorrow’s Computing Today, 2 (1):92, March 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Hill:1993:HLU

[Hil93] Russ Hill. How to login to any UNIX system and get your Email. Sys Admin: The Journal for UNIX Systems Administrators, 2(5):79–??, September/October 1993. CODEN SYADE7. ISSN 1061-2688.

Hill:1998:SBU

[Hil98] Russ Hill. 10 steps to becoming a UNIX administrator. Sys Admin: The Journal for UNIX Systems Administrators, 7(7): 55, 57–61, July 1998. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Hillegass:2002:CPM

[Hil02] Aaron Hillegass. Cocoa programming for Mac OS X. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-72683-1 (paperback). xix + 383 pp. LCCN QA76.76.O63 H57145 2002.

Hillegass:2008:CPM

[Hil08] Aaron Hillegass. Cocoa Programming for Mac OS X. Addison-Wesley, Reading, MA, USA, third edition, 2008. ISBN 0-321-50361-9. 496 (est.) pp. US$.

Hinnant:1988:AUB

[Hin88] David F. Hinnant. Accurate Unix benchmarking: art, science, or black magic? IEEE Micro, 8(5):64–75, September/October 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hiraki:2002:DRU

[HIT^{+}02] Kei Hiraki, Mary Inaba, Junji Tamatsukuri, Ryutaro Kurusu, Yukichi Ikuta, Hisashi Koga, and Akira Zinzaki. Data reservoir: Utilization of multigigabit backbone network for data-intensive research. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.
org/paperpdfs/pap.pap327.
pdf.

Hiura:1996:UIU

[Hiu96] H. Hiura. Unicode and internationalization with UNIX and X-Window’s system. In Unicode Consortium [Uni96a], pages C3–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Hiura:1997:UIU

[Hiu97] Hideki Hiura. Unicode and internationalization with UNIX and the X-Window System. In Unicode Consortium [Uni97], page ?? ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc10/program.html.

Hecht:1988:EAC

[HJAW88] Matthew S. Hecht, Abhai Johri, Radhakrishna Aditham, and T. John Wei. Experience adding C2 security features to UNIX. In Association [Ass88d], pages 133–146.

Haggerty:1979:ARF

[HK79] J. P. Haggerty and P. A. Kirslis.

Adding remote files to UNIX — a progress report on the File Server Project. Technical Memorandum 1434, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 7, 1979. ?? pp.

Hancock:1982:CP

[HK82a] Les Hancock and Morris Krieger. The C Primer. Mc-Graw-Hill, New York, NY, USA, 1982. ISBN 0-07-025981-X. ix + 235 pp. LCCN QA76.73.C15 H36 1982.

Holtkamp:1982:FMS

[HK82b] B. Holtkamp and H. Kaestner.

A firmware monitor to support vertical migration decisions in the UNIX operating system. ACM SIGMICRO Newsletter, 13(4):153–162, December 1982. CODEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/
1014194.800945.

Haight:1983:AEU

[HK83] R. C. Haight and D. B. Knudsen. ARIEL: An experimental UNIX-based interactive video information system. In Association [Ass83], pages 167–168. Abstract only.

Hoel:1986:UBO

[HK86] Timothy W. Hoel and Bruce A. Keller. A Unix-based operating system for the Cray 2. In USENIX [USE86b], pages 219– 224.

Hudson:1989:CSA

[HK89] Scott E. Hudson and Roger King. Cactis: a self-adaptive, concurrent implementation of an object-oriented database management system. ACM Transactions on Database Systems, 14(3):291–321, September 1989. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tods/1989-
14-3/p291-hudson/p291-hudson.
pdf;   http://www.acm.org/
pubs/citations/journals/
tods/1989-14-3/p291-hudson/
;   http://www.acm.org/pubs/
toc/Abstracts/tods/68013.
html.

Hannum:1995:NPH

[HK95] Charles Hannum and John Kohl. The NetBSD project: a highly portable UNIX-like system. ;login: the USENIX Association newsletter, 20(6):39– 41, December 1995. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Hildebrandt:2005:BPS

[HK05] Ralf Hildebrandt and Patrick Koetter. The book of Postfix: state-of-the-art message transport. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-001-1. xxviii + 464 pp. LCCN TK5105.74.P66 H55 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
http://www.loc.gov/catdir/
toc/ecip047/2003017563.html.

Hanson:1984:IDM

[HKF84] S. J. Hanson, R. E. Kraut, and J. M. Farber. Interface design and multivariate analysis of UNIX command use. ACM Transactions on Office Information Systems, 2 (1):42–57, January 1984. CO-DEN ATOSDO. ISSN 0734-

2047. URL http://www.acm.
org:80.

Holland:1985:ESS

[HKM^{+}85] Les Holland, Granino Korn, John Matson, Bob Seader, and Phil Wolfe. Engineering support system software. IEEE Micro, 5(5):17–21, September/October 1985. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hoflehner:2004:COT

[HKS^{+}04] Gerolf Hoflehner, Knud Kirkegaard, Rod Skinner, Daniel Lavery, Yong fong Lee, and Wei Li. Compiler optimizations for transaction processing workloads on Itanium Linux systems. In IEEE [IEE04], pages 294–303. ISBN 0-7695-2126-6. ISSN 1072-4451. LCCN QA76.9.A73. IEEE Computer Society Order Number P2126. ACM Order Number 520040.

Hanslien:2005:MPE

[HKT05] Monica Hanslien, Kenneth H. Karlsen, and Aslak Tveito. A maximum principle for an explicit finite difference scheme approximating the Hodgkin– Huxley model. BIT Numerical Mathematics, 45(4):725– 741, December 2005. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0006-
3835&volume=45&issue=4&spage=
725.

Handa:2021:OAD

[HKVR21] Shivam Handa, Konstantinos Kallas, Nikos Vasilakis, and Martin C. Rinard. An orderaware dataflow model for parallel Unix pipelines. Proceedings of the ACM on Programming Languages (PACMPL), 5(ICFP):65:1–65:28, August 2021. CODEN ???? ISSN 2475-1421 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473570.

Hawley:1985:WUA

[HL85a] Michael J. Hawley and Samuel J.

Leffler. Windows for UNIX at Lucasfilm. In USENIX Association [USE85d], pages 393–406. LCCN QA76.8.U65 U8 1985.

Hawley:1985:WUL

[HL85b] Michael J. Hawley and Samuel J.

Leffler. Windows for UNIX at Lucasfilm. In USENIX [USE85a], pages 393–406.

Hornsby:1985:DIF

[HL85c] C. Hornsby and C. H. C. Leung.

The design and implementation of a flexible retrieval language for a Prolog database system. ACM SIGPLAN Notices, 20(9): 43–51, September 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hirschberg:1990:EDP

[HL90] Daniel S. Hirschberg and Debra A. Lelewer. Efficient decoding of prefix codes. Communications of the Association for Computing Machinery, 33(4): 449–459, April 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/77566.html. This is the

first of four key papers behind the bzip2 compression tools. The others are [BW94c, Whe97, SB97].

Hart:1995:DFB

[HL95] Richard O. Hart and Glenn Lupton. DEC FUSE: Building a graphical software development environment from UNIX tools. Digital Technical Journal of Digital Equipment Corporation, 7(2):5– 19, Spring 1995. CODEN DTJOEL. ISSN 0898-901X.

URL  ftp://ftp.digital.
com/pub/Digital/info/DTJ/
v7n2/DEC_FUSE_Building_a_
Graphica_02oct1995DTJI01P8.
ps;   http://www.digital.
com:80/info/DTJI01/DTJI01AH.
HTM;   http://www.digital.
com:80/info/DTJI01/DTJI01P8.
PS;   http://www.digital.
com:80/info/DTJI01/DTJI01PF.
PDF;   http://www.digital.
com:80/info/DTJI01/DTJI01SC.
TXT.

Haworth:1990:EDS

[HLHR90] Guy Haworth, Steve Leunig, Carsten Hammer, and Mike Reeve. The European Declarative System, database, and languages. IEEE Micro, 10 (6):20–23, 83–88, November/ December 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hatch:2001:HLE

[HLK01] Brian Hatch, James Lee, and George Kurtz. Hacking Linux exposed: Linux security secrets and solutions. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212773-2. xxx + 566 pp. LCCN QA76.76.O63 H377 2001.

Holmes:1979:LCS

[HLLP79] V. P. Holmes, B. J. B. Lax, P. M. Lu, and J. L. Pappas. Load characterization for SDS/ UNIX. Technical Memorandum 1436, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 27, 1979. ?? pp.

Haley:1979:AIP

[HLMR79] C. B. Haley, T. B. London, J. F. Maranzano, and Dennis M. Ritchie. Alternate implementation proposal for Unix/

370. Bell Laboratories letter, January 2, 1979. URL http:

//charles.the-haleys.org/
papers/Alternate_Implementation_
Proposal_for_Unix370.pdf.

This letter argues for porting Unix to the IBM mainframe platform inside a virtual machine, rather than as an independent operating system that would require taking over the entire physical machine, or, as proposed in an earlier memo by B. G. Prieve, as a process under IBM’s TSS (Time Sharing System). In the end, the TSS solution was chosen: see [FMM84].

Hoang:2021:PHD

[HLT^{+}21] Thong Hoang, Julia Lawall, Yuan Tian, Richard J. Oentaryo, and David Lo. PatchNet: Hierarchical deep learningbased stable patch identification for the Linux kernel.

IEEE Transactions on Software Engineering, 47(11):2471– 2486, November 2021. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Hecht:1984:DFS

[HLW84] Matthew S. Hecht, John R. Levine, and Justin C. Walker. A distributed file system for UNIX. In USENIX Association [USE84b], pages 23–32. ISBN none. LCCN QA76.8.U65 U55 1984.

Hunt:1975:ADF

[HM75] J. W. Hunt and M. Douglas McIlroy. An algorithm for differential file comparison. Technical Memorandum 1083 (TM 75-1271-11), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1, 1975. ?? pp.

Hunt:1976:ADF

[HM76] J. W. Hunt and M. D. McIlroy. An algorithm for differential file comparison. Computing Science Technical Report 41, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1976. 6 +

                                        3 pp. URL https://www.cs.
                                        dartmouth.edu/~doug/diff.
                                        pdf.

Hudson:1990:GUI

[HM90a] Scott E. Hudson and Shamim P. Mohamed. A graphical user interface server for UNIX. Software—Practice and Experience, 20(12):1227–1239, December 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hume:1990:URS

[HM90b] Andrew G. Hume and M. Douglas McIlroy, editors. UNIX Research System. Computing Science Research Center, AT&T Bell Laboratories, Murray Hill, NJ, USA, tenth edition, 1990. ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). 702 (volume 1), vi + 630 (volume 2) pp. LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/

. Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Hafner:1991:COH

[HM91] Katie Hafner and John Markoff.

Cyberpunk — Outlaws and Hackers on the Computer Frontier. Simon and Schuster, New York, NY, USA, 1991. ISBN 0-671-68322-5. 368 pp. LCCN QA76.9.A25 H34 1991. US$22.95. See also [Sto89, Sto90].

Harrison:1998:ETT

[HM98] Mark Harrison and Michael McLennan. Effective Tcl/ Tk programming: writing better programs with Tcl and Tk. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-63474-0. xv + 405 pp. LCCN QA76.73.T44H37 1998. US$38.50.

Haider:2020:FHF

[HMK^{+}20] Waqas Haider, Nour Moustafa, Marwa Keshk, Amanda Fernandez, Kim-Kwang Raymond Choo, and Abdul Wahab. FGMC-HADS: Fuzzy Gaussian mixture-based correntropy models for detecting zeroday attacks from Linux systems. Computers & Security, 96(??):Article 101906, September 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167404820301826.

Holt:1983:TAP

[HMP83a] R. C. Holt, M. P. Mendel, and S. G. Perelgut. TU-NIS: a portable, UNIX compatible kernel written in Concurrent Euclid. In Software Tools Users Group [Sof83], pages 61– 74. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Holt:1983:TPU

[HMP83b] R. C. Holt, M. P. Mendel, and S. G. Perelgut. TUNIS: a portable, UNIX compatible kernel written in concurrent Euclid. In Association [Ass83], pages 61–74.

Howard:2001:JTE

[HN01] John S. Howard and Alex Noordergraaf. JumpStart Technology: Effective Use in the Solaris Operating Environment. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-062154-4. xvii + 259 pp. LCCN

???? URL books/js.pdf;
http://www.sun.com/books/
catalog/howard/. Part No.

806-6872-10 September 2001, Revision 01.

Honeyman:1984:HDU

[HNR84] Peter Honeyman, Dave Nowitz, and Brian E. Redman. Honey Danber — the UUCP of the future. In ????, editor, EUUG Conference, Nijmegen 1984 Proceedings, page ?? ????, ????, April 1984.

HCC:1978:GUP

[HO 78] HO Computer Center. A guide to the [use] of pseudo-ATS on UNIX. Technical Memorandum 1301, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1978. ?? pp.

Hamilton:1993:SHU

[HO93] Jeffrey W. Hamilton and Eileen M. Ormsby. Simulating hypercubes in UNIX. part II. Dr. Dobb’s Journal of Software Tools, 18(1):66, 68, 70, 99–100, 102–105, January 1993. CO-DEN DDJOEB. ISSN 1044-789X.

Hoalst:1977:SMC

[Hoa77] B. C. Hoalst. String manipulation and conversion for switching control center. Technical Memorandum 1211 (TM 77-5222-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 20, 1977. ?? pp.

Hobby:1989:RCC

[Hob89] John D. Hobby. Rasterizing curves of constant width. Journal of the ACM, 36(2): 209–229, April 1989. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0004-
                                        5411/62045.html.

Hobby:1991:GAT

[Hob91] John D. Hobby. Generating automatically-tuned bitmaps from outlines. Computing Science Technical Report 148, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 19, 1991. i + 42 pp.

Hobby:1993:GAT

[Hob93] John D. Hobby. Generating automatically tuned bitmaps from outlines. Journal of the ACM, 40(1):48–94, January 1993. CODEN JA-COAH. ISSN 0004-5411 (print), 1557-735X (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0004-
                                        5411/138040.html.

Hobby:19xx:DGM

[Hobxxa] John D. Hobby. Drawing graphs with MetaPost. Computing Science Technical Report 164, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. i + 20 pp.

Hobby:19xx:UMM

[Hobxxb] John D. Hobby. A user’s manual for MetaPost. Computing Science Technical Report 162, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 91 pp.

Hofmann:1977:SIU

[Hof77] B. B. Hofmann. 8080 simulator (i80sim) user manual. Technical Memorandum 1224 (TM 77-2265-8), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1, 1977. ?? pp.

Hoffman:1979:EOC

[Hof79] R. E. Hoffman. Experimental office communications services — message system. Technical Memorandum 1421, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 13, 1979. ?? pp.

Hoffman:1997:CPM

[Hof97] Doris Hoffman. Color printing management in a UNIX environment. Sys Admin: The Journal for UNIX Systems Administrators, 6(3):20– ??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Hofinger:2001:LBE

[Höf01]  Siegfried Höfinger.  Load

balancing for the electronic structure program GREM-LIN in a very heterogeneous SSH-connected WAN-cluster of UNIX-type hosts. Lecture Notes in Computer Science, 2074:801–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-
ny.com/link/service/series/
0558/bibs/2074/20740801.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2074/20740801.
pdf.

Holmgren:1975:RNU

[Hol75]  S. Holmgren.  RFC 681:

Network UNIX, March 18,

1975.  URL ftp://ftp.
internic.net/rfc/rfc681.
txt;   ftp://ftp.math.utah.
edu/pub/rfc/rfc681.txt. Sta-

tus: UNKNOWN. Not online.

Holzmann:1980:DCS

[Hol80] Gerard J. Holzmann. The design of coordination schemes. Computing Science Technical Report 87, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 7, 1980. ii + 25 pp. URL https:/

/spinroot.com/gerard/pdf/
coordination80.pdf.

Holt:1982:TUL

[Hol82] R. C. Holt. Tunis: a Unix lookalike written in concurrent Euclid (abstract). Operating Systems Review, 16(1):4–5, January 1982. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Holub:1987:CUW

[Hol87a] Allen I. Holub. Curses: UNIXcompatible windowing output functions. Dr. Dobb’s Journal of Software Tools, 12(7):94– 104, 74–93, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:curses

[Hol87b] Allen I. Holub. Curses: UNIXcompatible windowing output functions. Dr. Dobb’s Journal of Software Tools, 12(7):94– 104, 74–93, July 1987. CODEN DDJOEB. ISSN 1044-789X.

Holub:1987:CWU

[Hol87c] Allen I. Holub. On Command:

Writing a Unix-Like Shell for MS-DOS. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1987. ISBN 0-934375-29-1. 319 pp. LCCN QA76.76.O63H65 1987.

Holzmann:1987:PPE

[Hol87d] G. J. Holzmann. Pico — a picture editor. AT&T Technical Journal, 66(2):2–13, 1987. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Holzmann:1987:MPA

[Hol87e] Gerard J. Holzmann. Manual for the protocol analyzer ‘Trace’. Computing Science Technical Report 134, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 11, 1987. ii + 19 pp. URL https:/

/spinroot.com/gerard/pdf/
trace87.pdf.

Holzmann:1990:PLC

[Hol90a] G. J. Holzmann. Pico — a language for composing digital images. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990.

URL http://doc.cat-v.org/
unix/v10/10thEdMan/pico.
pdf;   http://doc.cat-v.
org/unix/v10/10thEdMan/v2contents.

html. Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Holzmann:1990:SPA

[Hol90b] G. J. Holzmann. Spin — a protocol analyzer. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Holzmann:1991:TDV

[Hol91] Gerard J. Holzmann. Tutorial: Design and validation of protocols. Computing Science Technical Report 157, AT&T Bell Laboratories, Murray Hill, NJ, USA, May ??, 1991. ii + 39 pp.

Holsberg:1992:UDG

[Hol92] Pete Holsberg. UNIX Desktop Guide to Tools. Howard W. Sams, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30202-0. xvii + 476 pp. LCCN QA76.76.O63 H649 1992. US$27.95. A new and comprehensive guide to numerous UNIX utilities.

Holoch:1997:RSU

[Hol97] Rich Holoch. RPTRTR — a simple UNIX report router. Sys Admin: The Journal for UNIX Systems Administrators, 6(5): 63–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

Hollabaugh:2002:ELH

[Hol02] Craig Hollabaugh. Embedded Linux: Hardware, Software, and Interfacing. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-672-32226-9 (paperback). viii + 419 pp. LCCN QA76.76.O63 H6475 2002.

Holm:2012:PCC

[Hol12] Nils M. Holm. Practical Compiler Construction: A Nononsense Tour through a C Compiler. Lulu Press, ????, 2012. ISBN ???? 390 (est.) pp. LCCN ???? URL https:/

/www.t3x.org/reload.

Hong:2006:BIS

[Hon06] Bryan Hong. Building an Internet Server with FreeBSD 6. Lulu.com, ????, 2006. ISBN 1-4116-9574-7. 228 (est.) pp. LCCN ???? US$29.95.

Hong:2008:BSF

[Hon08] Bryan Hong. Building a Server with FreeBSD 7: a modular approach. No Starch Press, San Francisco, CA, USA, 2008. ISBN 1-59327-145-X. xx + 264 pp. LCCN QA76.76.O63 H6694 2008. US$29.95.

URL http://www.loc.gov/
catdir/enhancements/fy0715/
2007000276-d.html;   http:
//www.loc.gov/catdir/toc/
ecip078/2007000276.html.

Hooker:1978:ARU

[Hoo78] J. W. Hooker. Audio Response Unit (ARU) subsystem design for Real-Time Retrieval (RTLR)/ARU prototype system. Technical Memorandum 1306 (TM 78-9156-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 15, 1978. ?? pp.

Hooker:1979:RTL

[Hoo79] J. W. Hooker. Real Time Listing Retrieval (RTLR)/Audio Response Unit (ARU) prototype system — performance measurements. Technical Memorandum 1333 (TM 79-9156-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 10, 1979. ?? pp.

Hoover:1983:UID

[Hoo83] Clyde W. Hoover. A user information data base for UNIX (what to do when /etc/passwd just isn’t enough). In Software Tools Users Group [Sof83], pages 121–134. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Hook:2005:WPC

[Hoo05] Brian Hook. Write portable code: an introduction to developing software for multiple platforms. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-056-9. xx + 248 pp. LCCN QA76.76.C64 H66 2005.

Horton:1979:PVU

[Hor79] M. R. Horton. A portable version of the UNIX text editor Ed. Technical Memorandum 1456 (TM 79-3723-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 14, 1979. ?? pp.

Horbal:1982:ATL

[Hor82] Mark T. Horbal. ATLAS test language — a real time application under UNIX. In Usr Group [Usr82], pages 165–176.

Horton:1990:PCS

[Hor90a] Mark R. Horton. Portable C Software. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-868050-7. ix + 372 pp. LCCN QA76.73.C15 H67 1990. US$53.00.

Horwitz:1990:IST

[Hor90b] Susan Horwitz. Identifying the semantic and textual differences between two versions of a program. ACM SIG-PLAN Notices, 25(6):234–245, June 1990. CODEN SINODQ. ISBN 0-89791-364-7. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/93542/p234-horwitz/.

Horspool:1992:BUE

[Hor92] Nigel Horspool. The Berkeley UNIX Environment. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1992. ISBN 0-13-089368-

4. xii + 379 pp. LCCN QA76.73.C15H67 1992. Highly recommended. An excellent book on C programming for Berkeley UNIX system … A companion text intended for use in college and university courses concerned with Compiler Construction, Software Engineering and Operating Systems.

Horvath:1998:UME

[Hor98] David B. Horvath. Unix for the mainframer: the essential reference for commands, conversions, and TCP/IP. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-632837-7 (paperback). xxi + 392 pp. LCCN QA76.76.O63 H6744 1998.

Horwitz:2002:USM

[Hor02] Jeff Horwitz. Unix system management: primer plus. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32372-

9. xvii + 531 pp. LCCN QA76.76.O63 H675 2002.

Hosler:1983:MCU

[Hos83] Jay R. Hosler. Meeting the coming UNIX training challenge. In Association [Ass83], pages 177–184.

Hosler:1984:IPU

[Hos84]  Jay Hosler.  Interactivity

in packaged UNIX training: a modest proposal. In Software Tools Users Group [Sof84], pages 346–349. LCCN QA76.8.U65 U83 1984.

Hoskins:2005:UOS

[Hos05] Matthew Hoskins. Unix: Old school. Linux Journal, 2005 (140):??, December 2005. CO-DEN LIJOFX. ISSN 1075-3583 (print), 1938-3827 (electronic).

Howard:2001:BDM

[How01] John S. Howard. Boot Disk Management: a Guide for the Solaris Operating System. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-062153-6. xvi + 204 pp. LCCN QA76.76.O63H687

2002. URL books/BootDisk.
pdf;   http://www.sun.com/
books/catalog/bootdisk/;
http://www.sun.com/books/
catalog/howardbp/. Part No.

816-0240-10 December, 2001.

Anonymous:1993:HU

[HP-93] HP-UX, 1993. ISSN 1075-0703.

Interex (the International Association of Hewlett–Packard Computer Users), Sunnyvale, CA, USA.

Henderson:1989:MIR

[HP89a] Robert L. Henderson and Alan Poston. MSS-II and RASH — a mainframe UNIX based mass storage system with a rapid access storage hierarchy file management system. In USENIX [USE89b], pages 65–84.

Henderson:1989:MRA

[HP89b] Robert L. Henderson and Alan Poston. MSS-II and RASH — a mainframe UNIX based mass storage system with a rapid access storage hierarchy file management system. In USENIX Association [USE89f], pages 65–84.

Holderbaugh:1989:MPM

[HP89c] Mark C. Holderbaugh and Scott E. Preece. Minimalist physical memory control in UNIX. In USENIX Association [USE89f], pages 245–256.

Husain:1995:LU

[HP^{+}95] Kamran Husain, Tim Parker, et al. Linux Unleashed. SAMS Publishing, Indianapolis, IN, USA, 1995. ISBN 0-672-30705-7. xxxv + 1058 pp. LCCN QA76.76.O63L5547. US$49.99. Includes CD ROM.

Hamlin:1998:IAA

[HP98] J. H. Hamlin and W. D. Potter. An intelligent agent to aid in Unix system administration. Lecture Notes in Computer Science, 1416:252–260, 1998. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hall:2001:JLG

[HP01] Michael Hall and Brian Proffitt. The joy of Linux: a gourmet guide to open source. Prima Publishing, Roseville, CA, USA, 2001. ISBN 0-7615-3151-3. 340 pp. LCCN QA76.76.O63 H34354 2001.

Hanrot:1987:KBC

[HQZ^{+}87] S. Hanrot, P. Quintrand, J. Zoller, E. Chouraqui, P. Dugerdil, P. Francois, and M. Ricard. A knowledge based CAD system in architecture on UNIX. In USENIX Association [USE87a], pages 169–181. ISBN ???? LCCN ????

Hall:1973:IFV

[HR73] A. D. Hall, Jr. and B. G. Ryder.

Installation of the FORTRAN verifier. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973. ?? pp.

Hahner:1977:DM

[HR77] I. A. Hahner and T. M.

Raleigh. Description of MK-CONF. Technical Memorandum 1140 (MF 77-8234-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 17, 1977. ?? pp.

Haverty:1978:ICS

[HR78a] J. F. Haverty and R. D. Rettburg. Interprocess communications for a server in UNIX. In ????, editor, Proceedings of Compcon 78: Computer Communications Networks, pages 312–315. ????, ????, September 1978.

Hedrick:1978:PRR

[HR78b] E. L. Hedrick and T. M. Raleigh. Preliminary report on RCMAS and MERT performance problems. Technical Memorandum 8, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 10, 1978. ?? pp.

Hedrick:1979:PDE

[HR79a] E. L. Hedrick and T. M. Raleigh. PDP-11… /70 design error: Handling segmentation violations different from PDP-11/45 implementation. Technical Memorandum 1380 (TM 79-4145-3, TM 79-2524-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 10, 1979. ?? pp.

Huber:1979:DIV

[HR79b] D. J. Huber and W. D. Roome. DIRECT II VAX performance improvement estimates. Technical Memorandum 1422, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 30, 1979. ?? pp.

Holbrook:1991:SSH

[HR91] Paul Holbrook and Joyce Reynolds. Site security handbook. A guide to setting computer security policies and procedures for sites that have systems on the Internet …. This is rfc1244.txt which is available by anonymous ftp from nic. ddn.mil (in rfc directory).

Also, rfc1281.txt — Guidelines for the Secure Operation of the Internet …., 1991.

Hindin:1994:IAO

[HR94] Harvey Hindin and Wendy Rauch. IBM’s AS/400 openness strategy. 3X/400 Systems Management, 22(12):54–??, December 1, 1994. ISSN 1070-6097.

Hays:1982:IMP

[HRO82] A. V. Hays, Jr., B. J. Richmond, and L. M. Optican. Implementing a multiple-process real-time system under UNIX. In USENIX [USE82], pages 15– ?? Abstract only.

Hicks:2010:LSA

[HRS^{+}10] Boniface Hicks, Sandra Rueda, Luke St. Clair, Trent Jaeger, and Patrick McDaniel. A logical specification and analysis for SELinux MLS policy. ACM Transactions on Information and System Security, 13 (3):26:1–26:??, July 2010. CO-DEN ATISBQ. ISSN 1094-9224 (print), 1557-7406 (electronic).

Haight:1977:SPU

[HRW77] R. C. Haight, W. D. Roome, and L. A. Wehr. Setting up PWB/UNIX (I.1). Technical Memorandum 1194, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1977. ?? pp.

Hanley:1982:SUU

[HS82] James R. Hanley and Jeffry A.

Scott. A survey of UNIX usage in scientific and business applications. In Usr Group [Usr82], pages 281–?? Abstract only.

Hearn:1983:RIG

[HS83] A. C. Hearn and L. R. Seward. REDUCE installation guide for the DEC VAX-11 series running UNIX, vers. 3.0. Report Rand Publication CP84 (4/83), The Rand Corporation, Santa Monica, CA, USA, April 1983.

Harbison:1984:CAR

[HS84] Samuel P. Harbison and Guy L. Steele Jr. C — a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-110008-4. x + 352 pp. LCCN QA76.73.C15 H38 1984.

Harbison:1987:CAR

[HS87a] Samuel P. Harbison and Guy L. Steele Jr. C — a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1987. ISBN 0-13-109802-0. xii + 404 pp. LCCN QA76.73.C15 H38 1987.

Haviland:1987:USP

[HS87b] Keith Haviland and Ben Salama. UNIX System Programming. International computer science series. Addison-Wesley, Reading, MA, USA, 1987. ISBN 0-201-12919-1. xv + 339 pp. LCCN QA76.76.O63 H38 1987. UK£15.95.

Hume:1990:FSS

[HS90] Andrew Hume and Daniel Sunday. Fast string searching.

Computing Science Technical Report 156, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1990. ?? pp.

Harbison:1991:CRM

[HS91a] Samuel P. Harbison and Guy L.

Steele Jr. C: a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 1991. ISBN 0-13-110933-2 (paperback), 0-13-110941-3 (hardcover), 0-13-116807-X (international). viii + 392 pp. LCCN QA76.73.C15 H38 1991.

Hume:1991:FSS

[HS91b] Andrew Hume and Daniel Sunday. Fast string searching. Software—Practice and Experience, 21(11):1221–1248, November 1991. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Harbison:1995:CRM

[HS95] Samuel P. Harbison and Guy L.

Steele Jr. C — a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 1995. ISBN 0-13-326232-4 (hardback), 0-13-326224-3 (paperback). xx + 455 pp. LCCN QA76.73.C15 H38 1995. US$33.95.

Hall:1999:RLD

[HS99] Jon Hall and Paul G. Sery. Red Hat Linux for Dummies with CDROM: The Complete Version of Red Hat Linux 6.1 on 2 CD-ROMs. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-0663-3. xxiv + 374 pp. LCCN QA76.76.O63 H34349 2000.

Hall:2000:RLD

[HS00] Jon Hall and Paul G. Sery. Red Hat Linux 7 for Dummies with CDROM. IDG Books, San Mateo, CA, USA, 2000. ISBN 0-7645-0795-8. xxii + 393 pp. LCCN QA76.76.O63 H34348 2001. US$29.99. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
description/wiley038/00108205.
html;   http://www.loc.gov/
catdir/toc/wiley021/00108205.
html.

Harbison:2002:CRM

[HS02] Samuel P. Harbison III and Guy L. Steele Jr. C — a Reference Manual. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fifth edition, 2002. ISBN 0-13-089592-X. xviii + 533 pp. LCCN QA76.73.C15 H38 2002. US$45.00. URL http:

//www.CAReferenceManual.
com/;   http://www.phptr.
com/ptrbooks/ptr_013089592X.
html.

Heuer:2004:USL

[HS04] Konrad Heuer and Reinhard Sippel. UNIX-Systemadministration: Linux, Solaris, AIX, FreeBSD, Tru64-UNIX. X.systems.press. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-43418-6, 3-642-62370-0. xxi + 710 pp.

LCCN ???? URL http://

digitool.hbz-nrw.de:1801/
webclient/DeliveryManager?
pid=1488760\%26custom\_att\
_2=simple\_viewer;   http:
//swbplus.bsz-bw.de/bsz107827026cov.
htm;   http://swbplus.bsz-
bw.de/bsz107827026inh.htm.

Howes:2003:UDL

[HSG03] Timothy A. Howes, Mark C.

Smith, and Gordon S. Good. Understanding and deploying LDAP directory services. Addison-Wesley, Reading, MA, USA, second edition, 2003. ISBN 0-672-32316-8. xxxiv + 899 pp. LCCN TK5105.595 .H69 2003.

Hsiao:2001:STY

[Hsi01] Aron Hsiao. Sams teach yourself Linux Security basics in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32091-6 (paperback). xv + 412 pp. LCCN QA76.76.O63 H755 2001.

Henry-Stocker:2003:SSS

[HSM03] Sandra L. Henry-Stocker and Evan R. Marks. Solaris Solutions for System Administrators: Time-Saving Tips, Techniques, and Workarounds. Wiley, New York, NY, USA, second edition, 2003. ISBN 0-471-43115-X. xx + 604 pp. LCCN QA76.76.O63 H476 2003.

Hart:2008:RTL

[HST08] D. Hart, J. Stultz, and T. Ts’o.

Real-time Linux in real time. IBM Systems Journal, 47(2):

207–??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/472/hart.html.

Heydon:1992:SCU

[HT92] Allan Heydon and J. D. Tygar. Specifying and checking UNIX security constraints. In USENIX [USE92e], pages 211–226. ISBN 1-880446-46-4. LCCN ????

Heydon:1994:SCU

[HT94a] Allan Heydon and J. D. Tygar. Specifying and checking UNIX security constraints. In USENIX [USE94a], pages 91– 112.

Heydon:COMPSYS-7-1-91

[HT94b] Allen Heydon and J. D. Tygar. Specifying and checking UNIX security constraints. Computing Systems, 7(1):91–112, Winter 1994. CODEN CMSYE2. ISSN 0895-6340.

Hagan:1996:UCS

[HTU96] Kerry Hagan, Chris Alan Thyberg, and Carnegie Mellon University. UNIX: computing skills workshop, 1995–1996. College custom series. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-064681-3. 146 pp. LCCN ????

Hecht:1973:ASA

[HU73] Matthew S. Hecht and Jeffrey D. Ullman. Analysis of a simple algorithm global data flow problems. In Fischer [Fis73]. ISBN 1-4503-7349-6. LCCN ????

Hu:1986:MFA

[Hu86] Irene Hu. Measuring file access patterns in UNIX. ACM SIG-METRICS Performance Evaluation Review, 14(2):15–20, August 1986. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Huang:2006:CLK

[Hua06] Feilong Huang. Compiling the Linux kernel with the Intel compiler. In Anonymous [Ano06a], page ?? ISBN ????

LCCN ???? URL http://www.
ice.gelato.org/;   http:/
/www.ice.gelato.org/about/
oct06_presentations.php.

Huber:1979:ULD

[Hub79] D. J. Huber. UNIX line discipline for IBM 2740-1 protocol. Technical Memorandum 1481 (TM 79-3781-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 28, 1979. ?? pp.

Hubbard:1996:BDM

[Hub96] Jordan K. Hubbard, editor.

BSD Docs: the most complete collection of BSD-UNIX related articles and documents from the Internet. Walnut Creek CDROM, Walnut Creek, CA, USA, 1996. ISBN 1-57176-126-8. 4 pp. LCCN QA76.76.O63. Includes CD-ROM.

Hubbard:2002:BCR

[Hub02] John F. Hubbard. Binary code reuse in a Linux environment. C/C++ Users Journal, 20(3): ??, March 2002. CODEN CCU-JEX. ISSN 1075-2838.

Hudson:2006:PN

[Hud06]  Paul Hudson.  PHP in a

nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10067-1 (paperback). xv + 352 pp. LCCN QA76.73.P224 H83 2006. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596100674.

Hughes:1988:MIU

[Hug88] Larry Hughes. A multicast interface for UNIX 4.3. Software —Practice and Experience, 18 (1):15–27, January 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hughes:1998:LDQ

[Hug98] Phil Hughes. Linux for dummies quick reference. For dummies. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 1998. ISBN 0-7645-0422-3 (paperback). xvi + 219 pp. LCCN QA76.76.O63 H844 1998.

Hulse:2005:RBC

[Hul05] Paul Hulse. Review: Beowulf Cluster Computing with Linux, second edition. The Computer Journal, 48(3):379–380, May 2005. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/reprint/48/3/379.

Hume:1988:FMI

[Hum88a] Andrew Hume. The file motel — an incremental backup system for Unix. In Association [Ass88d], pages 61–72.

Hume:1988:TTG

[Hum88b] Andrew Hume. A tale of two greps. Software—Practice and Experience, 18(11):1063– 1072, November 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Hume:1990:FMO

[Hum90a] A. G. Hume. The file motel: An owner’s manual. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Hume:1990:I

[Hum90b] A. G. Hume. Introduction. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Hume:1990:MSM

[Hum90c] A. G. Hume. Mk — a successor to Make. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Hunt:1985:CT

[Hun85] William James Hunt. The C Toolbox. Addison-Wes-ley, Reading, MA, USA, 1985. ISBN 0-201-11111-X. xviii + 413 pp. LCCN QA76.73.C15 H85 1985. US$19.95. An excellent introduction to the use of C for the writing of software tools, with particular emphasis (but not limited to) the IBM PC. Tools developed include a screen-based file viewing utility, sorting programs, a B-tree module, assembly language primitives, and a terminal emulator.

Hunter:1988:PCA

[Hun88a] Chad Hunter. Process cloning: a system for duplicating UNIX processes. In USENIX Association [USE88g], pages 373–379. ISBN ???? LCCN ????

Hunter:1988:PCS

[Hun88b] Chad Hunter. Process cloning:

a system for duplicating UNIX processes. In Association [Ass88d], pages 373–379.

Hunt:1992:TIN

[Hun92a] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-82-X (paperback). xxii + 471 pp. LCCN TK5105.9 .H86.

Hunter:1992:AGU

[Hun92b] Bruce Hunter. An administrator’s guide to the UNIX software jungle. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):68– ??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Hunter:1992:PAP

[Hun92c] Bruce H. Hunter. Printer administration potpourri for large-scale UNIX systems. Sys Admin: The Journal for UNIX Systems Administrators, 1(2): ??, July/August 1992. CODEN SYADE7. ISSN 1061-2688.

Hunter:1992:UNU

[Hun92d] Bruce H. Hunter. UNIX networking in the 90s — DOS/ UNIX connectivity. Sys Admin:

The Journal for UNIX Systems Administrators, 1(3):35– ??, September/October 1992. CODEN SYADE7. ISSN 1061-2688.

Hunt:1999:LNS

[Hun99] Craig Hunt. Linux network servers 24 × seven. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2506-9. xxvi + 626 pp. LCCN QA76.9.C55 H86 1999.

Hunt:2000:LDS

[Hun00a] Craig Hunt. Linux: DNS Server Administration. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2736-3. xxii + 423 pp. LCCN QA76.76.O63 H863 2000. US$39.99. URL http://

www.sybex.com/cgi-bin/rd_
bookpg.pl?2736back.html.

Hunt:2000:LSS

[Hun00b] Craig Hunt. Linux: Samba Server Administration. The Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2740-1. xxiii + 629 pp. LCCN QA76.76.O63 S58847 2001.

Hunt:2000:LSA

[Hun00c] Craig Hunt. Linux: System Administration. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2735-5. xxviii + 657 pp. LCCN A76.76.O63 S7346 2001.

Hunt:2001:LAW

[Hun01] Craig Hunt. Linux: Apache Web Server Administration. The Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2734-

7. xxiii + 615 pp. LCCN TK5105.8885.A63 A95 2001.

Hunt:2002:LNS

[Hun02a]  Craig Hunt.  Linux net-

work servers. Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2002. ISBN 0-7821-4123-

4. LCCN QA76.76.O63 H86 2002b; **See. URL http:

//library.books24x7.com/
library.asp?B&isbn=0782141234.

Hunt:2002:TIN

[Hun02b] Craig Hunt. TCP/IP Network Administration. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2002. ISBN 0-596-00297-1. xvii + 725 pp. LCCN TK5105.55 .H86 2002. US$44.95. URL http://

safari.oreilly.com/0596002971;
http://www.oreilly.com/
catalog/tcp3.

Hunt:2003:SC

[Hun03] Craig Hunt. Sendmail cookbook.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00471-0. xviii + 388 pp. LCCN TK5105.74.S44 H85 2004.

Hurwicz:1994:HIU

[Hur94] Mike Hurwicz. How To Integrate UNIX Servers and Net-Ware LANs. Datamation, 40 (23):57–??, December 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Husemann:1992:ICL

[Hus92] D. Husemann. ISO CONS in LANs — making it all work. A European contribution to 4.4 BSD Unix. Computer Networks and ISDN Systems, 25 (4–5):411–??, November 1992. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Haight:1978:SUT

[HW78] R. C. Haight and L. A. Wehr. Setting up UNIX/TS. Technical Memorandum 1320 (MF 78-8234-98), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 30, 1978. ?? pp.

Hu:2019:SCS

[HWBH19] Hanyang Hu, Shaowei Wang, Cor-Paul Bezemer, and Ahmed E. Hassan. Studying the consistency of star ratings and reviews of popular free hybrid Android and iOS apps. Empirical Software Engineering, 24 (1):7–32, February 2019. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-018-9617-6.

Haskell:1979:SFU

[HY79] B. G. Haskell and O. C.

Yue. SYSTID on F4P-UNIX — user’s guide. Technical Memorandum 1387 (TM 79-1343-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 4, 1979. ?? pp.

Hughes:2001:PDC

[HZ01] Sterling Hughes and Andrei Zmievski. PHP developer’s cookbook. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31924-1. xvii + 505 pp. LCCN QA76.73.P224 H84 2001.

Hughes:2002:PDC

[HZ02] Sterling Hughes and Andrei Zmievski. PHP developer’s cookbook. SAMS Publishing, Indianapolis, IN, USA, second edition, 2002. ISBN 0-672-32325-7. xvii + 480 pp. LCCN QA76.73.P224 H84 2002.

Ioannidis:1993:AIN

[IB93] John Ioannidis and Matt Blaze.

The architecture and implementation of network layer security in UNIX. In USENIX Association [USE93g], pages 29–39. ISBN 1-880446-55-3. LCCN QA 76.9 A25

U54 1993. URL http://www.
usenix.org/publications/
library/proceedings/sec4/.

Ip:1987:CCW

[IBC87] Chok-Ho Ip, Daniel M. Berry, and Kam Pui Chow. Cwpr, a Chinese/Japanese Word-Processing System for Use with the UNIX Device-Independent Troff System. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0780-7. 570–577 pp. LCCN ?QA75.5 .I5761 1987. IEEE Service Cent. Piscataway, NJ, USA.

IBM:1988:AOS

[IBM88] IBM. AIX operating system: IBM AIX family definition overview. Web document, July

1988. URL https://archive.
org/stream/bitsavers_ibmpcrtaixefinitionOverviewJul88_
4100993/GC23-2002-0_AIX_
Family_Definition_Overview_
Jul88_djvu.txt.

IBM:1989:A

[IBM89] IBM AIX Marketing. AIX. IBM Corporation, 44 South Broadway, White Plains, New York, 10601., 1989. 90

pp. URL https://amaus.net/
static/S100/IBM/RTPC/AIX%
20Family%20Definition%201989.
pdf.

IBM:2005:MAS

[IBM05] IBM Corporation. Mathematical Acceleration Subsystem for Linux. World Wide Web document, 2005. URL http://

www-306.ibm.com/software/
awdtools/mass/linux/mass-
linux.html.

IBM:2009:HAD

[IBM09a] IBM. High Availability and Disaster Recovery Options for DB2 on Linux, Unix, and Windows. IBM redbooks. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 2009. ISBN 0-7384-3138-9. xvi + 856 pp. LCCN ????

IBM:2009:MOI

[IBM09b] IBM. Migrating from Oracle to IBM Informix Dynamic Server on Linux, UNIX, and Windows. Redbooks. IBM Redbooks, 11400 Burnet Road, Austin, TX 78758-3493, USA, 2009. ISBN 0-7384-3302-0. xiv + 378 pp. LCCN ????

Immich:2003:PAF

[IBP03] Patricia K. Immich, Ravi S.

Bhagavatula, and Ravi Pendse. Performance analysis of five interprocess communication mechanisms across UNIX operating systems. The Journal of Systems and Software, 68(1): 27–43, October 15, 2003. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

IEEE:1983:CLC

[IEE83] IEEE, editor. 8th Conference on Local Computer Networks, Hilton Inn, Minneapolis, Minnesota, October 17–19, 1983. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1983. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE catalog no. 83CH1931-5. Computer Society order no. 500.

IEEE:1986:ITU

[IEE86] IEEE. IEEE Trial-Use Standard Portable Operating System for Computer Environments. Wiley-Interscience, New York, NY, USA, IEEE Std 1003.1 edition, April 1986. ISBN 0-471-85027-6. 207 pp. LCCN TK 275 I5 Std 1003.1.

IEEE:1988:ISP

[IEE88] IEEE. IEEE Standard Portable Operating System Interface for Computer Environments. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, IEEE Std 1003.1-1988 edition, 1988. ISBN 1-55937-003-3. 317 pp. LCCN QA76.76.O63 I6 1988. Revision of IEEE Std 1003.1, issued for Trial-Use in April 1986.

IEEE:1991:PIN

[IEE91a]  IEEE, editor.  Proceedings

of the IEEE 1991 National Aerospace and Electronics Conference, NAECON 1991, held at the Dayton Convention Center, May 20–24, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN NASEA9. ISBN 0-7803-0085-8. LCCN TL 693 N37 1991. Three volumes.

IEEE:1991:RTS

[IEE91b] IEEE, editor. REAL TIME ’91: Seventh Conference REAL TIME ’91 on Computer Applications in Nuclear, Particle, and Plasma Physics: June 24– 28, 1991, Julich, Fed. Rep. of Germany: together with short course on new backplane bus architectures and seminar on real time operating systems, Real Time -conference- 1991; 7th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-7803-0458-6, 0-7803-0459-4. LCCN QA76.54.C66 1991.

IEEE:1991:ISI

[IEE91c] IEEE Standards Board. IEEE standard for information technology: test methods for measuring conformance to POSIX. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 1-55937-104-8. x + 47 pp. LCCN QA76.76.O63I57 1991. IEEE Std 1003.3-1991. Approved March 21, 1991, IEEE Standards Board.

IEEE:1992:ISIa

[IEE92a] IEEE Standards Board. IEEE standard for information technology: POSIX Ada language interfaces — Part 1: Binding for system application program interface (API). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 18, 1992. xii + 305 pp. LCCN QA76.76.O63I445 1992. Std 1003.5-1992.

IEEE:1992:ISIb

[IEE92b] IEEE Standards Board. IEEE standard for information technology: POSIX FORTRAN 77 language interfaces — Part 1: Binding for system application program interface (API). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 18, 1992. ISBN 1-55937-230-3. xiii + 181 pp. LCCN QA76.76.O63I454 1992. Std 1003.9-1992.

IEEE:1994:PSH

[IEE94a] IEEE, editor. Proceedings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994.

IEEE:1994:PFI

[IEE94b] IEEE, editor. Proceedings: the First International Conference on Requirements Engineering, April 18–22, 1994, Colorado Springs, Colorado, International Conference on Requirements Engineering 1994; 1st. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5480-5, 0-8186-5481-3. LCCN QA76.758.I5717 1994.

IEEE:1994:ROS

[IEE94c] IEEE, editor. Real-time operating systems and software: RTOSS ’94: 11th Workshop — May 1994, Seattle, WA, IEEE Workshop on Real Time Operating Systems and Software 1994; 11th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5710-3. LCCN QA76.54.I173 1994.

IEEE:2002:STI

[IEE02] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

IEEE:2003:PCI

[IEE03]  IEEE, editor.  Proceedings

of the 2003 CGO: the International Symposium on Code Generation and Optimization; March 23–26, 2003, Fisherman’s Wharf, San Francisco, CA, with special emphasis on feedback-directed and runtime optimization. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1913-X. LCCN ???? ACM Order No. 530033.

IEEE:2004:PIS

[IEE04] IEEE, editor. Proceedings: 37th International Symposium on Microarchitecture, MICRO-37: 4–8 December 2004, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2126-6. ISSN 1072-4451. LCCN QA76.9.A73. IEEE Computer Society Order Number P2126. ACM Order Number 520040.

Israel:1992:XWS

[IF92] Elias Israel and Erik Fortune. The X-Window System Server — X version 11, Release 5. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1992. ISBN 1-55558-096-3. xv + 534 pp. LCCN QA76.76.W56 I75 1992. US$44.95.

Israeli:2010:LKC

[IF10] Ayelet Israeli and Dror G. Feitelson. The Linux kernel as a case study in software evolution. The Journal of Systems and Software, 83(3):485– 501, March 2010. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Itoi:1999:SUF

[IHR99] Naomaru Itoi, Peter Honeyman, and Jim Rees. SCFS: a UNIX filesystem for Smartcards. In USENIX [USE99c], page ?? ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL  http://www.usenix.
org/publications/library/
proceedings/smartcard99/
itoiSCFS.html.

Isaak:1998:MVP

[IJ98] Jim Isaak and Lowell Johnson. Micro view: POSIX/ UNIX standards — foundation for 21st-century growth. IEEE Micro, 18(4):88–??, July/August 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

URL http://dlib.computer.
org/mi/books/mi1998/pdf/
m4088.pdf.

Ito:1983:HFO

[IK83] Tetsuro Ito and Makoto Kizawa. Hierarchical file organization and its application to similar-string matching. ACM Transactions on Database Systems, 8(3):410–433, September 1983. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tods/1983-
8-3/p410-ito/p410-ito.pdf;
 http://www.acm.org/pubs/
citations/journals/tods/
1983-8-3/p410-ito/.

Ince:2019:BOS

[ILG19] M. Numan Ince, Joseph Ledet, and Melih Gunay. Building an open source Linux computing system on RISC-V. In IEEE, editor, 2019 1st International Informatics and Software Engineering Conference (UBMYK), pages 1–4. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2019.

Imagna:1976:UIP

[Ima76] C. P. Imagna. UNIX initialization and process control. Technical Memorandum 1128 (TM 76-3412-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 23, 1976. ?? pp.

Imagna:1979:URA

[Ima79] C. P. Imagna. UNIX/RT as an application system for BOS/ LMOS development. Technical Memorandum 1349, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 2, 1979. ?? pp.

Ince:1983:CST

[Inc83] D. C. Ince. A compatibility software tool for use with separately compiled languages. ACM SIGPLAN Notices, 18(9): 31–34, September 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

InfoMagic:1995:PTT

[Inf95] InfoMagic. Perl and Tcl/Tk. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1995. ISBN ???? ???? pp. LCCN ???? US$35.00.

IEC:1990:ITP

[Int90] International Electrotechnical Commission. Information technology — Portable Operating System Interface (POSIX). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 1-55937-061-0 (vol. 1). various pp. LCCN QA76.76.O63 I539 1990. International standard ISO/IEC 9945. IEEE Std 1003.1-1990 (revision of IEEE Std 1003.1-1988). Contents: pt. 1. System application program interface (API) [C language].

IEC:1993:ITP

[Int93] International Electrotechnical Commission. Information technology: Portable Operating System Interface (POSIX): Part 2, Shell and utilities. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, September 17, 1993. ISBN 1-55937-406-3. xix + 1198 pp. LCCN QA76.76.O63 I56 1993. Two volumes. Approved September 17, 1992, IEEE Standards Board; approved April 5, 1993, American National Standards Institute; approved 1993 by the International Organization for Standardization, and by the International Electrotechnical Commission. This standard … defines the applications interface to a shell command language and a set of utility programs for complex data manipulation. When the User Portability Utilities Option is included, the standard also defines a common environment for generalpurpose time-sharing users on character-oriented display terminals. International Standard ISO/IEC 9945-2: 1993. IEEE Std 1003.2-1992 (includes IEEE Std 1003.2a-1992).

Isaak:1986:IRT

[Isa86]  J. Isaak.  The impact of

real time computer systems standards: VME, POSIX and MAP. In Anonymous [Ano86], pages 98–104.

Isaak:1991:PIW

[Isa91] Jim Isaak. POSIX: An introduction to the world of VMS. The VAX professional, 13(3): 24–??, June 1991. CODEN VAXPEN. ISSN 8750-9628.

Isley:1983:ULN

[Isl83] Larry K. Isley. UNIX licensing and new AT&T product offerings. In Software Tools Users Group [Sof83], pages 3–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

ISO:2005:IID

[ISO05] ISO. ISO/IEC DIS 23360: Draft International Standard: Linux Standard Base Core Specification 2.0.1. International Organization for Standardization, Geneva, Switzerland, 2005. ISBN ???? 1214 pp. LCCN ????

                                        URL http://www.iso.org/
                                        iso/en/CatalogueDetailPage.
                                        CatalogueDetail?CSNUMBER=

41481&scopelist=PROGRAMME;
 http://www.linuxbase.org/
LSBWiki/IsoBallot.

IEEE:1993:DTD

[ITC93] Designing, testing, and diagnostics — join them: International Test Conference 1993 proceedings: October 17– 21, 1993, Convention Center, Baltimore, Maryland, USA, Proceedings of the International Test Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-1429-8. LCCN TK7874.I474 1993. IEEE catalog number 93CH3356-3.

Ivinskis:1995:HAC

[Ivi95] Kestutis Ivinskis. High availability of commercial applications. SIGMOD Record (ACM Special Interest Group on Management of Data), 24(2):433– 434, May 1995. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Innocent:1987:EUI

[IvW87] Peter R. Innocent, Gerrit C.

van der Veer, and Yvonne Waern. Experiments with the user interface for UNIX mail. In USENIX Association [USE87a], pages 73–91. ISBN ???? LCCN ????

Jackowski:1977:GIG

[Jac77a] D. J. Jackowski. Gex — an interactive graphics program. Technical Memorandum 1227 (TM 77-8254-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 7, 1977. ?? pp.

Jackson:1977:IGS

[Jac77b] J. H. Jackson. Interactive graphics system for a UNIX environment. Technical Memorandum 1200 (TM 77-5221-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 29, 1977. ?? pp.

Jacob:1984:ULW

[Jac84a] Robert J. K. Jacob. Userlevel window manager for UNIX. In USENIX Association [USE84b], pages 123–134. ISBN none. LCCN QA76.8.U65 U55 1984.

Jacob:1984:UWM

[Jac84b] Robert J. K. Jacob. Userlevel window manager for UNIX. In USENIX Association [USE84b], pages 123–134. ISBN none. LCCN QA76.8.U65 U55 1984.

Jacobson:1987:TUL

[Jac87] Van Jacobson. Tuning UNIX Lex or it’s NOT true what they say about Lex. In USENIX Association [USE87e], pages 163– 164. LCCN QA 76.76 O63 U84 1987. Abstract only.

Jackson:1991:GGM

[Jac91] Alan Jackson. The growing government market. UNIX/World, 8(2):73–??, February 1, 1991. ISSN 0739-5922.

Jacobs:1999:DDG

[Jac99] Jon Jacobs. Delphi developer’s guide to OpenGL. Wordware Publishing, Plano, TX, USA, 1999. ISBN 1-55622-657-8. ???? pp. LCCN T385 .J334 1999.

Jaeschke:1985:LCN

[Jae85]  Rex Jaeschke.  Let’s C

Now. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1985. ISBN 0-9614729-2-8. LCCN QA76.73.C15 J336 1985.

Jaeschke:1986:SCH

[Jae86] Rex Jaeschke. Solutions in C — Hundreds of Programming Tips by the Author of “Doctor C’s Pointers”. Addison-Wes-ley, Reading, MA, USA, 1986. ISBN 0-201-15042-5. 247 pp. LCCN QA76.73.C15 J34 1986. US$17.95.

Jaeschke:1989:MSC

[Jae89a] Rex Jaeschke. Mastering Standard C — a Self-paced Training Workbook for Modern C Language. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1989. ISBN 0-9614729-8-7. 366 pp. LCCN QA76.73.C15 J337 1989.

Jaeschke:1989:PCL

[Jae89b] Rex Jaeschke. Portability and the C Language. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1989. ISBN 0-672-48428-5. xvi + 382 pp. LCCN QA76.76.C64 J34

Jaeschke:1991:DSC

[Jae91] Rex Jaeschke. The Dictionary of Standard C. Professional Press Books, 101 Witmer Road, Horsham, PA 19044, USA, 1991. ISBN 1-878956-07-8. x + 165 pp. LCCN QA76.73.C15 J335 1991.

Jaeschke:1993:CIE

[Jae93] Rex Jaeschke. C++: An Introduction for Experienced C Programmers. CBM Books, A Division of Cardinal Business Media Inc., 101 Witmer Road, Horsham, PA 19044, USA, 1993. ISBN 1-878956-27-2. xii + 236 pp. LCCN QA76.73.C15 J3354 1993.

Jaeschke:1994:DSCa

[Jae94a] Rex Jaeschke. The Dictionary of Standard C. Harcourt Brace Jovanovich, Boston, MA, USA, Japanese language edition, 1994. ISBN 4-8337-8523-4. 240 pp.

Jaeschke:1994:DSCb

[Jae94b] Rex Jaeschke. The Dictionary of Standard C. Piter, St. Petersburg, Russia, Russian language edition, 1994. ISBN 5-7190-0018-6. 221 pp.

Jalan:1992:CID

[Jal92] Ajay Jalan. Comparison and implementation of dynamic load balancing facilities under the UNIX and Mach operating systems. Thesis (M.S.), Worcester Polytechnic Institute, Worcester, MA, USA, 1992. vi + 69 pp.

James:1997:UE

[Jam97] Scott D. James. UNIX for engineers. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-8053-6488-9. vi + 122 pp. LCCN QA 76.76 O63 J36 1997.

James:2009:FSC

[Jam09] Daniel James. Free Software for Creative People. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4302-1887-8. 450 pp. LCCN ????

Jang:2003:LTW

[Jan03] Michael Jang. Linux Transfer for Windows Network Admins: a roadmap for building a Linux file and print server. Hentzenwerke Publishing, Milwaukee, WI 53217-5361, USA, 2003. ISBN 1-930919-46-8. 310 (est.) pp. LCCN ???? US$49.95. Edited by Elizabeth Zinkann.

Jang:2004:RRC

[Jan04] Michael Jang. RHCE Red Hat certified engineer: Linux study guide (exam RH302). Osborne/McGraw-Hill, Berkeley, CA, USA, fourth edition, 2004. ISBN 0-07-225365-7. 755 pp. LCCN QA76.3 .R46 2002.

Jang:2006:LAG

[Jan06a] Michael H. Jang. Linux annoyances for geeks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-00801-5. xv + 484 pp. LCCN QA76.76.O63; QA76.76.O63 J36 2006eb; QA76.76.O63 J36 2006. URL http:/

/www.oreilly.com/catalog/
9780596008017.

Jang:2006:LPM

[Jan06b]  Michael H. Jang.  Linux

Patch Management: keeping Linux systems up to date. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-236675-4. ???? pp. LCCN QA76.76.O63 J368 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0519/2005028070.html.

Janert:2008:GAU

[Jan08] Philipp Janert. Gnuplot in Action: Understanding Data with Graphs. Manning Publications, Greenwich, CT, USA, 2008. ISBN 1-933988-39-8. 275 (est.) pp. LCCN ???? US$35.00.

Janssens:1986:AUM

[JAV86] M. D. Janssens, J. K. Annot, and A. J. Van De Goor. Adapting UNIX for a multiprocessor environment. Communications of the Association for Computing Machinery, 29(9):895– 901, September 1986. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/6598.html.

Jones:1999:OSL

[JB99] Christopher A. Jones and Drew Batchelor. Open source Linux Web programming. IDG Books, San Mateo, CA, USA, 1999. ISBN 0-7645-4619-8. xxi + 476 pp. LCCN QA76.76.O63 J662 1999. US$39.99.

Johnson:1992:UWS

[JC92] Jim Johnson and Jerry Cashin.

Users will seal fate of Posix plan. Software magazine, 12(3): 82–??, March 1, 1992. CODEN SMWMEQ. ISSN 0897-8085.

Jain:2009:SSA

[JDB09] Puneet Jain, Scott Davenport, and David Bustos. Solaris 10 System Administration Essentials. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2009. ISBN 0-13-700009-X. 448 (est.) pp. LCCN ????

Jensen:1975:UFD

[Jen75] P. D. Jensen. UNIX facilities for the development of programs written in Macro-11. Technical Memorandum 1080 (TM 75-8234-14), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 5, 1975. ?? pp.

Jennings:2000:JQC

[Jen00] Mike Jennings. Java Q&A:

Can you write NT services in Java? Dr. Dobb’s Journal of Software Tools, 25(3):113– 116, 118, March 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

ddj.com/ftp/2000/2000_03/
jqa300.txt;   http://www.
ddj.com/ftp/2000/2000_03/
jqa300.zip.

Jennings:2002:JQ

[Jen02] Mike Jennings. Java Q&A: How do I use the Java plugin tool? Dr. Dobb’s Journal of Software Tools, 27(1): 93–94, January 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_01/jqa0102.
zip.

Jenkins:2005:RLE

[Jen05] G. K. Jenkins. Review of “The Linux Enterprise Cluster”, by Karl Kopper, No Starch Press, 2005, $49.95 ISBN: 1-59327-036-4. ACM Queue: Tomorrow’s Computing Today, 3(9): 58, November 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Jensen:2019:UHP

[Jen19]  Richard Jensen.  Unix at

50: How the OS that powered smartphones started from failure. Today, Unix powers iOS and Android — its legend begins with a gator and a trio of researchers. Web site, August 29, 2019. URL

https://arstechnica.com/
gadgets/2019/08/unix-at-
50-it-starts-with-a-mainframe-
a-gator-and-three-dedicated-
researchers/.

Joy:1981:AIC

[JF81a] William Joy and Robert Fabry.

An architecture for interprocess communication in UNIX. Report CSRG-TR-3, Computer Systems Research Group, Computer Science Division, Department of Electrical Engineering and Computer Science, University of California, Berkeley, CA 94720, June 21, 1981. 48 pp.

Joy:1981:PEU

[JF81b] William Joy and Robert Fabry.

Proposals for enhancement of UNIX on the VAX. Report CSRG-TR-4, Computer Systems Research Group, Computer Science Division, Department of Electrical Engineering and Computer Science, University of California, Berkeley, CA 94720, July 21, 1981. 59 pp. Revised 31-Aug-1981.

Jang:2012:OSS

[JF12] Michael Jang and Harry Foxwell. Oracle Solaris 11 System Administration: The Complete Reference. McGraw-Hill Osborne Media, 2012. ISBN 0-07-179042-X. 768 (est.) pp. LCCN ????

Janakiram:2008:OOW

[JGS^{+}08] D. Janakiram, Ashok Gunnam, N. Suneetha, Vineet Rajani, and K. Vinay Kumar Reddy. Object-oriented wrappers for the Linux kernel. Software —Practice and Experience, 38 (13):1411–1427, November 10, 2008. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Jalics:1983:TPO

[JH83] Paul J. Jalics and Thomas S. Heines. Transporting a portable operating system: UNIX to an IBM minicomputer. Communications of the Association for Computing Machinery, 26(12):1066– 1072, December 1983. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Johnston:1986:UBD

[JH86] William E. Johnston and Dennis E. Hall. UNIX based distributed printing in a diverse environment. In USENIX Association [USE86c], pages 514– 528.

Jones:2000:SC

[JH00] Floyd Jones and Solveig Haugland. StarOffice 5.2 Companion. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-030703-3. xix + 1121 pp. LCCN QA76.76.I57 J655 2001. US$39.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130307033.html;   http:
                                        //www.sun.com/books/catalog/
                                        jones/index.html.

Jambor:2007:ILL

[JHT^{+}07] Martin Jambor, Tomas Hruby, Jan Taus, Kuba Krchak, and Viliam Holub. Implementation of a Linux log-structured file system with a garbage collector. Operating Systems Review, 41(1):24–32, January 2007. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Jolitz:1991:PUBd

[JJ91a] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386. the basic kernel (device autoconfiguration). Dr. Dobb’s Journal of Software Tools, 16(11):46, 48, 50, 52, 54, 56–58, 60, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBb

[JJ91b] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386. the basic kernel multiprogramming and multitasking. I. Dr. Dobb’s Journal of Software Tools, 16 (9):44, 46, 48, 50, 52, 54, 56, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBc

[JJ91c] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386. The basic kernel multiprogramming and multitasking. II. Dr. Dobb’s Journal of Software Tools, 16(10):62, 64, 66, 68, 70, 72, 118–120, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPa

[JJ91d] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(1):

??, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPc

[JJ91e] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(2): ??, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPd

[JJ91f] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16 (3):??, March 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPe

[JJ91g] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(4):??, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPf

[JJ91h] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(5):??, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPg

[JJ91i] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(6):??, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPh

[JJ91j] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(7):??, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPi

[JJ91k] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(8):??, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPj

[JJ91l] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(9): ??, September 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPk

[JJ91m] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(10): ??, October 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPl

[JJ91n] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 16(11):

??, November 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUPb

[JJ91o] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach (designing the software specification). Dr. Dobb’s Journal of Software Tools, 16 (1):16–18, 20, 22–24, 28, 30, 32, 34, 36, 38–40, 42, 46, January 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUS

[JJ91p] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a strippeddown kernel. Dr. Dobb’s Journal of Software Tools, 16(7):32, 34–36, 38, 40, 84–88, 88, 90, July 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUL

[JJ91q] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: Language tools cross support. Dr. Dobb’s Journal of Software Tools, 16 (4):34, 36–38, 42, 44–45, 90, April 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUR

[JJ91r] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386 research and the commercial sector. Dr. Dobb’s Journal of Software Tools, 16(6):50, 52, 54, June 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUBa

[JJ91s] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: the basic kernel. Dr. Dobb’s Journal of Software Tools, 16(8):54, 56– 60, 64–65, 126, August 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUI

[JJ91t] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386. the initial root filesystem. Dr. Dobb’s Journal of Software Tools, 16 (5):46, 48, 50, 52–54, May 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1991:PUT

[JJ91u] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: three initial PC utilities. Dr. Dobb’s Journal of Software Tools, 16(2): 54, 56, 58–61, 111–112, 114– 115, February 1991. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUDb

[JJ92a] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386: device drivers. Dr. Dobb’s Journal of Software Tools, 17(3):44, 46, 48, 50, 90–91, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUDc

[JJ92b] W. F. Jolitz and L. G. Jolitz.

Porting UNIX to the 386: device drivers. Dr. Dobb’s Journal of Software Tools, 17(4):64– 67, 108–109, April 1992. CO-DEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUF

[JJ92c] W. F. Jolitz and L. G. Jolitz. Porting UNIX to the 386: the final step. Dr. Dobb’s Journal of Software Tools, 17(7):50, 52, 54, 56, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUDa

[JJ92d] W. F. Jolitz and L. Greer Jolitz. Porting UNIX to the 386 device drivers. Dr. Dobb’s Journal of Software Tools, 17 (2):30, 32–33, 93, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPa

[JJ92e] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 17(2): ??, February 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPb

[JJ92f] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 17 (3):??, March 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPc

[JJ92g] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 17(4):??, April 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPd

[JJ92h] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 17(5):??, May 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPe

[JJ92i] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 17(6):??, June 1992. CODEN DDJOEB. ISSN 1044-789X.

Jolitz:1992:PUPf

[JJ92j] William Frederick Jolitz and Lynne Greer Jolitz. Porting UNIX to the 386: a practical approach. Dr. Dobb’s Journal of Software Tools, 17(7):??, July 1992. CODEN DDJOEB. ISSN 1044-789X.

Johnson:1973:PLB

[JK73] Steven C. Johnson and Brian W.

Kernighan. The programming language B. Technical Report 8, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1973. ?? pp. URL

https://web.archive.org/
web/20230328114209/https:
//www.bell-labs.com/usr/
dmr/www/bintro.html;   https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/bintro.html.

Jung:1986:KUY

[JK86] Robert S. Jung and Joseph T. Kalash. Kanji UNIX: Yunikkusu wa Nihongo o Hanasemasu (UNIX speaks Japanese). In USENIX Association [USE86c], pages 209–222.

Jacobsen:2015:LCD

[JKS^{+}15] Charles Jacobsen, Muktesh Khole, Sarah Spall, Scotty Bauer, and Anton Burtsev. Lightweight capability domains: Towards decomposing the Linux kernel. Operating Systems Review, 49(2):44–50, December 2015. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Johnson:1978:UTSb

[JL78] Steven C. Johnson and Michael E. Lesk. UNIX time-sharing system: Language development tools. The Bell System Technical Journal, 57(6): 2155–2175, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2155.
                                        pdf.

Johnson:1987:LDT

[JL87a] Steven C. Johnson and Michael E. Lesk. Language development tools. In UNIX System Readings and Applications, pages 245–265. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-938532-0. Reprinted from The Bell System Technical Journal, 1978.

Johnson:tools87

[JL87b] Steven C. Johnson and Michael E.

Lesk. Language development tools. In UNIX System Readings and Applications, pages 245–265. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. Reprinted from The Bell System Technical Journal, 1978.

Jones:1996:GCA

[JL96] Richard Jones and Rafael Lins.

Garbage collection: algorithms for automatic dynamic memory management. Wiley, New York, NY, USA, 1996. ISBN 0-471-94148-4. xxvi + 377 pp. LCCN QA76.9.G37 J66 1996.

URL http://www.loc.gov/
catdir/description/wiley036/
96014601.html;   http://
www.loc.gov/catdir/toc/onix07/
96014601.html.

Jaekl:1998:PBS

[JL98] Chris Jaekl and Paul Lu. Programmer’s bookshelf: Software projects and UNIX programming. Dr. Dobb’s Journal of Software Tools, 23(6):127, 129, June 1998. CODEN DDJOEB. ISSN 1044-789X.

Jude:2016:FMA

[JL16] Allan Jude and Michael W. Lucas. FreeBSD Mastery: Advanced ZFS. Tilted Windmill Press, ????, 2016. ISBN 0-692-68868-4 (paperback). xx + 222 pp. LCCN ????

Jacobson:1984:UTI

[JLSG84] Van Jacobson, Craig Leres, Joseph Sventek, and Wayne Graves. 4BSD UNIX TCP/ IP and VMS DECNET: Experience in negotiating a peaceful coexistence. In Software Tools Users Group [Sof84], pages 323–325. LCCN QA76.8.U65 U83 1984. Abstract only.

Judice:1977:OCR

[JLSS77] C. N. Judice, Heinz Lycklama, D. K. Sharma, and T. P. Sosnowski. Office communication research in lab 135. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 8, 1977. 4 pp.

                                        URL https://www.tuhs.org/
                                        Archive/Documentation/TechReports/
                                        Heinz_Tech_Memos/Office_
                                        Communication_Research_in_
                                        Lab_135_19770208.pdf.

Jude:2019:IKM

[JMR19] Allan Jude, Marshall Kirk McKusick, and Benedict Reuschling. An interview with Kirk McKusick. The FreeBSD Journal, 6(6):40–50, November/ December 2019. ISBN 0-615-88479-2. URL http:/

                                        /issue.freebsdfoundation.
                                        org/publication/?m=33057&
                                        i=217640&p=32.

Joyce:1988:RDU

[JN88] Jim Joyce and Bob Nystrom.

Rescuing data in UNIX file systems (what to do after rm *). In USENIX Association [USE88g], pages 331–334. ISBN ???? LCCN ????

Johnson:1975:YYA

[Joh75] Steven C. Johnson. Yacc — yet another compiler compiler. Computing Science Technical Report 32, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1975. ?? pp.

Johnson:1978:PCT

[Joh78a] S. C. Johnson. A portable compiler: Theory and practice. In ACM, editor, Proceedings of the 5th ACM Symposium on Principles of Programming Languages, pages 97–104. ACM Press, New York, NY, USA, January 1978.

Johnson:1978:LPC

[Joh78b] Steven C. Johnson. Lint, a program checker. Technical Memorandum 1186, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 26, 1978. ?? pp.

Johnson:1978:YYA

[Joh78c] Steven C. Johnson. YACC — yet another compiler compiler. Technical Memorandum 1013 (TM 78-1273-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 31, 1978. 2 + 30 pp.

Johnson:1979:BPD

[Joh79a] Steven C. Johnson. A 32bit processor design. Technical Memorandum 1375 (TM 79-1273-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 2, 1979. ?? pp.

Johnson:1979:TTU

[Joh79b] Steven C. Johnson. A tour through the UNIX C compiler. Technical Memorandum 1409, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Johnson:1979:YYA

[Joh79c] Steven C. Johnson. Yacc: Yet another compiler compiler. In UNIX Programmer’s Manual, volume 2, pages 353–387. Holt, Rinehart, and Winston, New York, NY, USA, 1979. AT&T Bell Laboratories Technical Report July 31, 1978.

Johnson:197x:URB

[Joh7x] S. C. Johnson. Users’ reference to B on MH-TSS. Technical Report 8, AT&T Bell Laboratories, Murray Hill, NJ, USA, 197x. 19 pp. URL

https://web.archive.org/
web/20230326075939/https:
//www.bell-labs.com/usr/
dmr/www/bref.pdf.

Johnson:1980:LDT

[Joh80] Steven C. Johnson. Language development tools on the Unix system. Computer, 13(8):16– 24, August 1980. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1983:TTP

[Joh83] Steven C. Johnson. A tour through the Portable C Compiler. In ATT [?], pages 529–543. ISBN 0-03-061743-X. LCCN QA76.8.U65 B44

1983.  URL http://www.
thphy.uni-duesseldorf.de/
~vieth/subjects/computer/
7thEdMan/vol2/porttour.ps.

gz. On pp. 382–386, there is a yacc grammar for a desk calculator implementing interval arithmetic.

Johnson:1987:ULF

[Joh87] Stephen C. Johnson. UNIX:

The language forms. In USENIX Association [USE87e], pages 16–20. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Johson:1992:UXT

[Joh92] Eric F. Johson. Using X — Troubleshooting the X Window System, Motif, and Open Look. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1992. ISBN 1-55828-212-2. xx + 351 pp. LCCN QA76.76.W56 J639 1992. US$29.95.

Johnson:1995:RBU

[Joh95] Colt Johnson. Real-time between DOS, UNIX, and MVS.

Sys Admin: The Journal for UNIX Systems Administrators, 4(5):8–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Johnson:1996:ICB

[Joh96a] Colt Johnson. Interprocess communication between UNIX and MVS applications using sockets. Sys Admin: The Journal for UNIX Systems Administrators, 5(8):??, August 1996. CODEN SYADE7. ISSN 1061-2688.

Johnson:1996:IUM

[Joh96b] Colt S. Johnson. Interlinking UNIX and MVS. Sys Admin: The Journal for UNIX Systems Administrators, 5(3):??, March 1996. CODEN SYADE7. ISSN 1061-2688.

Johnson:2005:SSR

[Joh05] Chris F. A. Johnson. Shell scripting recipes: a problemsolution approach. The expert’s voice in open source; Books for professionals by professionals. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-471-1 (paperback). xxiv + 421 pp. LCCN QA76.76.O63 J628 2005.

Jones:1984:IIU

[Jon84] Douglas W. Jones. Improved interpretation of UNIX-like file names embedded in data. Communications of the Association for Computing Machinery, 27 (8):782–784, 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Jones:1988:TUS

[Jon88] Darrell Jones. UNIX for Super-Users, by Eric Foxley. C Users Journal, 6(1):42–??, January 1988. ISSN 0898-9788. See [Fox85].

Jones:1989:IXW

[Jon89] Oliver Jones. Introduction to the X Window System. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-499997-5. xii + 511 pp. LCCN QA76.76.W56 J66 1989.

Jones:1998:BWC

[Jon98] Tim Jones. Backing up Windows clients on UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 7(9): NT–1–NT–7, September 1998. CODEN SYADE7. ISSN 1061-

2688.  URL http://www.
samag.com/.

Jones:2002:JMA

[Jon02] M. Tim Jones. Java mobile agents and the Aglets SDK. Dr. Dobb’s Journal of Software Tools, 27(1):42, 44, 46–48, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
aglet.txt.

Jones:2005:GLA

[Jon05] M. Tim Jones. GNU/Linux application programming. Charles River Media programming series. Charles River Media, Inc., Hingham, MA, USA, 2005. ISBN 1-58450-371-8. xxv + 486 pp. LCCN QA76.76.O63 J665 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip052/2004024882.html.

Jordan:2004:ESL

[Jor04] Edmund Jordan. Embedded Systeme mit Linux programmieren: GNU-Softwaretools zur Programmierung ARMbasierender Systeme. (German) [Embedded Systems with Linux Programming: GNU Software Tools for Programming ARMbased Systems]. Franzis, Feldkirchen, Germany, 2004. ISBN 3-7723-5599-4. 384 pp. LCCN ????

Joseph:1995:IPA

[Jos95] Moses Joseph. Is POSIX appropriate for embedded systems? Embedded Systems Programming, 8(7):90–??, ???? 1995. CODEN EYPRE4. ISSN 1040-3272.

Josey:1997:GSA

[Jos97] Andrew Josey. Go Solo 2: the authorized guide to version 2 of the Single UNIX specification. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-575689-8. xxxiv + 578 pp. LCCN QA76.76.O63 G618 1997.

Josuttis:1999:CSL

[Jos99] Nicolai M. Josuttis. The C++ Standard Library: a tutorial and handbook. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-37926-0. xx + 799 pp. LCCN QA76.73.C153 J69 1999.

Josuttis:2012:CSL

[Jos12] Nicolai M. Josuttis. The C++ Standard Library: a tutorial and reference. Addison-Wesley, Reading, MA, USA, second edition, 2012. ISBN 0-321-62321-5. xxvii + 1099 pp. LCCN QA76.73.C153 J69 2012.

Josuttis:2019:CCG

[Jos19] Nicolai M. Josuttis. C++17:

The Complete Guide. ????, ????, 2019. ISBN 3-96730-017-X. xxiii + 420 (est.) pp. LCCN ????

Jouvelot:1986:DNL

[Jou86] P. Jouvelot. Designing new languages or new language manipulation systems using ML. ACM SIGPLAN Notices, 21 (8):40–52, August 1986. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Joy:1997:BU

[Joy97] Mike Joy. Beginning UNIX.

Tutorial guides in computing and information systems. International Thomson Computer Press, 20 Park Plaza Suite 1001, Boston, MA 02116 USA, second edition, 1997. ISBN 1-85032-263-5. ix + 242 pp. LCCN QA76.76.O63 J69 1997.

James:2009:CDM

[JP09] Daniel James and Trevor Parsons. Crafting digital media: Audacity, Blender, Drupal, GIMP, Scribus, and other open source tools. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1888-6 (ebook), 1-4302-1887-8. xvi + 428 pp. LCCN QA76.76.I59 J36 2009eb.

Jourdan:2017:SPC

[JP17] Jacques-Henri Jourdan and François Pottier. A simple, possibly correct LR parser for C11. ACM Transactions on Programming Languages and Systems, 39(4):14:1–14:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jepson:2000:DAP

[JPS00] Brian Jepson, Joan Peckham, and Ram Sadasiv. Database application programming with Linux. Wiley, New York, NY, USA, 2000. ISBN 0-471-35549-6 (paperback). xii + 516 pp. LCCN QA76.9.D26 J48 2000.

Johnson:1978:UTSa

[JR78] Steven C. Johnson and Dennis M. Ritchie. UNIX timesharing system: Portability of C programs and the UNIX system. The Bell System Technical Journal, 57(6):

2021–2048, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2021.
pdf;   https://ieeexplore.
ieee.org/document/6770409.

Johnson:1981:CLC

[JR81] Steven C. Johnson and Dennis M. Ritchie. The C language calling sequence. Computing Science Technical Report 102, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, 1981. 12 pp. URL https:

//www.bell-labs.com/usr/
dmr/www/clcs.pdf.

Johnson:1989:XWA

[JR89] Eric F. Johnson and Kevin Reichard. X Window Applications Programming. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1989. ISBN 1-55828-016-2. xxii + 562 pp. LCCN QA76.76.W56 J64 1989.

Johnson:1990:AXW

[JR90] Eric F. Johnson and Kevin Reichard. Advanced X Window Applications Programming — The Basics and Beyond. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1990. ISBN 1-55828-029-4. xxii + 615 pp. LCCN QA76.76.W56 J63 1990. US$29.95 (book), US$59.95 (book/disk).

Johnson:1992:XWA

[JR92] Eric F. Johnson and Kevin Reichard. X Window Applications Programming. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1992. ISBN 1-55828-178-9 (book), 1-55828-180-9 (book/disk: wrong ISBN in book). xxxi + 564 pp. LCCN QA76.76.W56 J64 1992. US$29.95.

Johnson:1993:PGP

[JR93] Eric F. Johnson and Kevin Reichard. Professional Graphics Programming in the X Window System. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1993. ISBN 1-55828-255-

6. xxxvi + 1195 pp. LCCN T385 .J63 1993. US$34.95.

Jepson:2002:MXU

[JR02] Brian Jepson and Ernest E. Rothman. Mac OS X for Unix Geeks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00356-0. xv + 198 pp. LCCN QA76.76.O63 J47 2002. US$24.95. URL http://

                                        safari.oreilly.com/0596003560;
                                         http://www.oreilly.com/
                                        catalog/mosxgeeks.

Jepson:2004:MXP

[JR04] Brian Jepson and Ernest E. (Ernest Eric) Rothman. Mac OS X Panther for Unix geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, second edition, 2004. ISBN 0-596-00607-1. xvii + 363 pp. LCCN QA76.76.O63 J475 2004. URL http:/

/www.oreilly.com/catalog/
9780596006075.

Jepson:2005:MXT

[JR05] Brian Jepson and Ernest E.

(Ernest Eric) Rothman. Mac OS X Tiger for Unix geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00912-7. xviii + 395 pp. LCCN QA76.76.O63 J47 2005.

URL http://www.oreilly.
com/catalog/9780596009120.

Johnson:1977:CLC

[JRL77] Steven C. Johnson, Dennis M.

Ritchie, and Michael E. Lesk. The C language calling sequence. Technical Memorandum 1207 (TM 77-1273-15, TM 77-1274-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 26, 1977. ?? pp.

Jepson:2008:MXU

[JRR08] Brian Jepson, Richard Rosen, and Ernest E. Rothman. Mac OS X for Unix geeks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2008. ISBN 0-596-52062-X. xviii + 406 pp. LCCN ????

Johnson:1990:YPG

[JS90] S. C. Johnson and R. Sethi.

Yacc: a parser generator. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Juned:2000:KPE

[JSG00] Wasim Juned, Kanwar Sidhu, and Todd Green. KDE programming by example. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2290-9. 400 pp. LCCN QA76.76.O63 L372 2004.

Johnson:1998:LAD

[JT98] Michael K. Johnson and Erik W. Troan. Linux Application Development. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30821-5. 576 pp. LCCN QA76.76.O63J635 1998. US$45.95.

Johnson:2005:LAD

[JT05] Michael K. Johnson and Erik W. Troan. Linux application development. Addison-Wesley, Reading, MA, USA, second edition, 2005. ISBN 0-321-21914-7. xxxi + 702 pp. LCCN QA76.76.O63 J635 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0421/2004017882.html.

Jaenson:1985:PAP

[JTUB85] Richard Jaenson, Gregory Taylor, Cyrus Umrigar, and Alison Brown. A parallel array processing environment under 4.2BSD UNIX. In USENIX Association [USE85d], pages 195– 208. LCCN QA76.8.U65 U8 1985.

Judge:1990:UPE

[Jud90] P. Judge. UNIX ports in Europe. Datamation, 36(3):78–80, February 1990. CODEN DTM-NAT. ISSN 0011-6963.

Jude:2017:FVL

[Jud17]  Allan Jude.  FreeBSD vs.

Linux: ZFS. The FreeBSD Journal, 4(5):8–11, September/October 2017. ISBN 0-615-88479-2. URL http:/

/issue.freebsdfoundation.
org/publication/?m=33057&
i=447089&p=10.

Jung:1985:PAD

[Jun85a] Robert S. Jung. Porting the AT&T demand paged UNIX implementation to microcomputers. In USENIX [USE85a], pages 361–372.

Jung:1985:PAT

[Jun85b] Robert S. Jung. Porting the AT&T demand paged UNIX implementation to microcomputers. In USENIX Association [USE85d], pages 361–372. LCCN QA76.8.U65 U8 1985.

Jepson:1999:OPT

[JW99] Brian Jepson and Larry Wall. O’Reilly Perl Toolkit for Linux and Solaris. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-604-8. 120 pp. LCCN ???? US$49.95. URL http:/

/www.oreilly.com/catalog/
perltkit/.

Jeffery:1998:LAP

[JZTB98] Clinton Jeffery, Wenyi Zhou, Kevin Templer, and Michael Brazell. A lightweight architecture for program execution monitoring. ACM SIG-PLAN Notices, 33(7):67–74, July 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kadhim:1995:BRLb

[Kad95a] Basim Kadhim. Book review: Linux Universe, Stefan Strobel and Thomas Uhl. Operating Systems Review, 29(4):3, October 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kadhim:1995:BRLa

[Kad95b] Basim Kadhim. Book review: Linux: Unleashing the Workstation in Your PC, Stefan Strobel and Thomas Uhl. Operating Systems Review, 29(2): 2–3, April 1995. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kojima:1983:AMI

[KAH83] Tomihiko Kojima, Hidehiko Akita, and Hisashi Hashimoto. An approach to a machine independent UNIX — UNIX on HI-TAC M-series virtual machines. In Software Tools Users Group [Sof83], pages 363–369. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Kalsi:2016:PLS

[Kal16]  Tajinder Kalsi.  Practi-

cal Linux Security Cookbook. Packt Publishing, Birmingham, 2016. ISBN 1-78528-642-0 (paperback), 1-78528-530-0 (e-book). LCCN T55.4-

60.8. URL http://proquest.
safaribooksonline.com/9781785286421.

Kamp:2002:RDD

[Kam02]  Poul-Henning Kamp.  Re-

thinking /dev and devices in the UNIX kernel. In USENIX [USE02], pages 77–

88. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736

2002.  URL http://www.
usenix.org/publications/
library/proceedings/bsdcon02/
kamp.html.

Kanazy:1993:SDP

[Kan93] Y. Kanazy. SGML document processing. Fujitsu, 44(6):521– 526, ???? 1993. CODEN FUJ-TAR. ISSN 0016-2515.

Kaplan:1979:GCM

[Kap79] R. S. Kaplan. Guidelines for converting MERT software to UNIX/RT. Technical Memorandum 1476, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 14, 1979. ?? pp.

Kaplan:1992:HCO

[Kap92] Alan E. Kaplan. A history of the COSNIX operating system: Assembly language UNIX 1970 to July 1991. In USENIX [USE92d], pages 429–438.

Kaplenk:1999:USA

[Kap99]  Joe Kaplenk.  UNIX Sys-

tem Administrator’s Interactive Workbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-081308-7. xxiii + 588 pp. LCCN QA76.76.O63K3645 1999. US$34.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130813087.html.

Karels:1983:IVS

[Kar83] Michael Karels. An implementation of the vfork system call for PDP-11 UNIX. In Association [Ass83], pages 40–?? Abstract only.

Karpinski:1984:KUP

[Kar84] Dick Karpinski. Keeping UNIX in its place: an interview with Bob Marsh. UNIX Review, 2 (9):38–43, 46, 48–49, 76, 78, 80, December 1984. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/unixreview_1984dec.
pdf.

Karlton:1992:IXE

[Kar92] Phil Karlton. Integrating the GL into the X environment: a high performance rendering extension working with and not against X. The X Resource, 1 (1):27–32, January 1992.

Karlsson:2006:BCS

[Kar06] Björn Karlsson. Beyond the C++ Standard Library: an introduction to Boost. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-13354-4. xl + 388 pp. LCCN QA76.73.C153 K365 2006. URL http:

//www.loc.gov/catdir/toc/
fy0612/2005927496.html.

Konishi:2006:LIL

[KAS^{+}06] Ryusuke Konishi, Yoshiji Amagai, Koji Sato, Hisashi Hifumi, Seiji Kihara, and Satoshi Moriai. The Linux implementation of a log-structured file system. Operating Systems Review, 40(3):102–107, July 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kasanen:2014:CLT

[Kas14]  Lauri Kasanen.  Into the

core: A look at Tiny Core Linux. ????, ????, 2014. ISBN 952-93-3391-9. LCCN ????

URL http://tinycorelinux.
net/book.html.

Katseff:1979:SSD

[Kat79] H. P. Katseff. Sdb: A symbolic debugger. Technical Memorandum 1369, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1979. ?? pp.

Katz:1982:LSA

[Kat82a] Fred M. Katz. The logical softshell: a full-screen interface to UNIX. In Usr Group [Usr82], pages 92–?? Abstract only.

Katz:1982:LSF

[Kat82b] Fred M. Katz. The logical softshell: a full-screen interface to UNIX. In USENIX [USE82], pages 92–?? Abstract only.

Katsive:1983:UMC

[Kat83] Bob Katsive. UNIX markets and competition. In Association [Ass83], pages 308–309. Abstract only.

Kathan:1999:PCM

[Kat99] Joseph Kathan. Portable control of multiple daemon processes. C/C++ Users Journal, 17(5):??, May 1999. CODEN CCUJEX. ISSN 1075-2838.

Kaufeld:1978:ICM

[Kau78] J. C. Kaufeld, Jr. Interprocess communication mechanisms in CB-UNIX. Technical Memorandum 1234 (TM 77-5223-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 19, 1978. ?? pp.

Kaufman:1979:ADH

[Kau79a] L. Kaufman. Application of dense Householder transformation to a sparse matrix. ACM Transactions on Mathematical Software, 5(4):442–450, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kaufman:1979:B

[Kau79b] Linda Kaufman. BPSS. Computing Science Technical Report 106e, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 28, 1979. 34 pp.

Kaufman:1990:TPS

[Kau90] Linda Kaufman. TTGU — a package for solving time varying partial differential equations on a union of rectangles. Computing Science Technical Report 154, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 29, 1990. i + 52 pp.

Kaufman:1993:B

[Kau93a] Linda Kaufman. BASS. Computing Science Technical Report 106d, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 11, 1993. 34 pp.

Kaufman:1993:CXD

[Kau93b] Linda Kaufman. CSWAP with X and Y declared complex. Computing Science Technical Report 106f, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 11, 1993. 13 pp.

Kaufman:1993:G

[Kau93c] Linda Kaufman. GESS. Computing Science Technical Report 106b, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 11, 1993. 36 pp.

Kaufman:1993:PSL

[Kau93d] Linda Kaufman. Programs for solving linear equations in the PORT library. Computing Science Technical Report 106a, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 11, 1993. i + 10 pp.

Kaufman:1993:S

[Kau93e] Linda Kaufman. SYSS. Computing Science Technical Report 106c, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 11, 1993. 30 pp.

Kaufman:1977:ADH

[Kau7 ] L. Kaufman. Application of dense Householder transformation to a sparse matrix. Computing Science Technical Report 63, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Kayel:1979:HHD

[Kay79] R. G. Kayel. Hi — Hierarchical Data Management System for UNIX. Technical Memorandum 1390 (TM 79-3444-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 30, 1979. ?? pp.

Kaza:2005:PAE

[Kaz05] Avinash Kaza. Preparation of acoustic emission data for neural network analysis using AWK and C programs. Thesis (M.S.), West Virginia University, Morgantown, WV, USA, 2005. xiv + 189 pp. URL

https://etd.wvu.edu/etd/
controller.jsp?moduleName=
documentdata&jsp%5FetdId=
3896.

Kelly-Bootle:2001:PMDb

[KB01]  Stan Kelly-Bootle.  Post-

mortem debunker: Reply ALL. C/C++ Users Journal, 19(3): 96–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Kernighan:1975:STM

[KC75] Brian W. Kernighan and Lorinda L. Cherry. A system for typesetting mathematics. Communications of the Association for Computing Machinery, 18(3):151–157, March 1975. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Kernighan:1976:TMU

[KC76] Brian W. Kernighan and Lorinda L. Cherry. Typesetting mathematics — user’s guide. Technical Memorandum ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 1976. ?? pp.

Kernighan:1978:STM

[KC78a] Brian W. Kernighan and Lorinda L. Cherry. A system for typesetting mathematics. Computing Science Technical Report 17, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1978. ?? pp.

Kernighan:1978:TMU

[KC78b] Brian W. Kernighan and Lorinda L. Cherry. Typesetting mathematics — user’s guide. Technical Memorandum 1011, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1978. 2 + 11

                                        pp. URL http://www.kohala.
                                        com/start/troff/v7man/eqn/
                                        eqn2e.ps.

Kernighan:1979:STM

[KC79] Brian W. Kernighan and Lorinda L. Cherry. A system for typesetting mathematics. Technical Memorandum 1029, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Kernighan:1990:TM

[KC90] B. W. Kernighan and L. L. Cherry. Typesetting mathematics. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

                                        //doc.cat-v.org/unix/v10/
                                        10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Komarinski:1998:LSA

[KC98] Mark F. Komarinski and Cary Collett. Linux System Administration Handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-680596-5. 416 pp. LCCN QA76.76.O63 K6483 1998. US$39.95. URL http://

www.phptr.com/. Includes CD-

ROM.

Komarinski:2000:RLA

[KC00] Mark (Mark F.) Komarinski and Cary Collett. Red Hat Linux administration handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-025395-2. ix + 405 pp. LCCN QA76.76.O63 K6485 2000.

Kirkland:2006:LTS

[KCTT06] James Kirkland, David Carmichael, Christopher L. Tinker, and Gregory L. Tinker, editors. Linux troubleshooting for system administrators and power users. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-185515-8 (paperback). 624 (est.) pp. LCCN QA76.76.O63 L54875 2006. US$49.99. URL http://

vig.prenhall.com/catalog/
academic/product/0,1144,
0131855158,00.html;   http:
//www.loc.gov/catdir/toc/
ecip066/2006000036.html.

Kanev:1994:FGH

[KD94] Kamen Kanev and Kris Dockx.

A framework for graphicallyoriented human computer interactions in intelligent operator support systems. Computers and Graphics, 18(4):563– 570, July–August 1994. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Kirch:2000:LNA

[KD00] Olaf Kirch and Terry Dawson. Linux network administrator’s guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 1-56592-400-2. xxix + 474 pp. LCCN QA76.76.O63 K566 2000.

Karshmer:1983:NMS

[KDP83] Arthur J. Karshmer, Dirk J. Depree, and James Phelan. The New Mexico State University ring-star system: a distributed UNIX environment. Software—Practice and Experience, 13(12):1157–1168, December 1983. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Kemmerer:1985:UUB

[KE85] Richard A. Kemmerer and Steven T. Eckmann. UNISEX: A UNIX-based symbolic EXecutor for Pascal. Software —Practice and Experience, 15 (5):439–458, May 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Keenan:2005:HUC

[Kee05]  Charles Keenan.  HP-UX

CSE: official study guide and desk reference. Hewlett– Packard professional books.

Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-146396-9 (hardcover). xxxiv + 1664 pp. LCCN QA76.76.O63 K43 2004.

Keller:1990:PGX

[Kel90] Brian J. Keller. A Practical Guide to X Window Programming — Developing Applications with the Xt Intrinsics and OSF/Motif. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1990. ISBN 0-8493-7406-5. xiii + 349 pp. LCCN QA76.76.W56 K45 1990.

Kelly:1996:AIA

[Kel96] David A. Kelly. AIX/6000 internals and architecture. Mc-Graw-Hill, New York, NY, USA, 1996. ISBN 0-07-034061-7. xiii + 271 pp. LCCN QA76.76.O63 K452 1996.

Kempe:1991:CTA

[Kem91] Johann Kempe. CL/TB: an Allegro Common Lisp programming interface for Trans-Base. Technical report TUM-I9106, Mathematisches Institut und Institut für Informatik der Technischen Universitãt München, Munich, Germany, April 1991. 56 pp.

Kempf:1993:IHR

[Kem93]  James Kempf.  Integrat-

ing handwriting recognition into UNIX. In USENIX [USE93c], pages 187–204. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/cinci93/kempf.
                                        html.

Kent:1990:SIX

[Ken90] Christopher A. Kent. Special issue on the X Window System, volume 20(S2) of Software practice and experience. Wiley, New York, NY, USA, 1990. ISBN 0-471-93006-7. S2/181 pp. LCCN QA76.5.A1 S653 v.20 no.S2.

Kepecs:1985:LPU

[Kep85] Jonathan Kepecs. Lightweight processes for UNIX implementation and applications. In USENIX Association [USE85d], pages 299–308. LCCN QA76.8.U65 U8 1985.

Kernighan:1974:PCT

[Ker74] Brian W. Kernighan. Programming in C — a tutorial (B.9). Technical Memorandum 1001 (TM 74-1273-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 5, 1974. 27 pp.

Kernighan:1975:RPR

[Ker75] Brian W. Kernighan. RAT-FOR — a preprocessor for a rational Fortran. Software— Practice and Experience, 5(4): 395–406, October/December 1975. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL https:/

                                        /onlinelibrary.wiley.com/
                                        doi/abs/10.1002/spe.4380050408.

Kernighan:1978:TT

[Ker78]  Brian W. Kernighan.  A

TROFF tutorial. Technical Memorandum 1010, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 4, 1978. i

+ 14 pp. URL http://www.
kohala.com/start/troff/v7man/
trofftut/trofftut.ps. Part

of UNIX Version 7 manual.

Kernighan:1979:AEU

[Ker79a] Brian W. Kernighan. Advanced editing on UNIX. Technical Memorandum 1108, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 28, 1979. ?? pp.

Kernighan:1979:RPR

[Ker79b] Brian W. Kernighan. RATFOR — a preprocessor for a rational Fortran. Technical Memorandum 1193, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1, 1979. 2 + 12 pp.

Kernighan:1979:TIU

[Ker79c] Brian W. Kernighan. A tutorial introduction to the UNIX text editor. Technical Memorandum 1000, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 8, 1979. ?? pp.

Kernighan:1979:UB

[Ker79d] Brian W. Kernighan. UNIX for beginners. Technical Memorandum 1042, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 8, 1979. ?? pp.

Kernighan:1979:UPMa

[Ker79e] Unix programmer’s manual. Web site, January 1979. URL

https://s3.amazonaws.com/
plan9-bell-labs/7thEdMan/

v7vol1.pdf. This is the last of seven editions (see [TRS18]).

Kernighan:1979:UPMb

[Ker79f] Unix programmer’s manual. Web site, January 1979. URL

https://s3.amazonaws.com/
plan9-bell-labs/7thEdMan/

v7vol2a.pdf. This is the last of seven editions (see [TRS18]).

Kernighan:1979:UPMc

[Ker79g] Unix programmer’s manual: Volume 2 — supplementary documents. Web site, January

10, 1979. URL https://s3.
amazonaws.com/plan9-bell-
labs/7thEdMan/v7vol2b.pdf.

This is the last of seven editions (see [TRS18]).

Kernighan:197x:TIL

[Ker7x] B. W. Kernighan. A tutorial introduction to the language B. Technical Report 8, AT&T Bell Laboratories, Murray Hill, NJ, USA, 197x. 17 pp. URL

https://web.archive.org/
web/20230506115024/https:
//www.bell-labs.com/usr/
dmr/www/btut.pdf.

Kernighan:1981:PCG

[Ker81a] Brian W. Kernighan. PIC — a crude graphics language for typesetting. Computing Science Technical Report 85 (TM 80-1272-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, January ??, 1981. ?? pp. See also journal article [Ker81b].

Kernighan:1981:PLT

[Ker81b] Brian W. Kernighan. PIC: a language for typesetting graphics. ACM SIGPLAN Notices, 16(6):92–98, June 1981. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kernighan:1981:WPM

[Ker81c] Brian W. Kernighan. Why Pascal is not my favorite programming language. Computing Science Technical Report 100, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, April 2, 1981. URL https://

github.com/arnoldrobbins/
cstr100.

Kernighan:1982:PLT

[Ker82a] Brian W. Kernighan. PIC: a language for typesetting graphics. Software—Practice and Experience, 12(1):1–21, January 1982. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL https:/

/onlinelibrary.wiley.com/
doi/abs/10.1002/spe.4380120102.

Kernighan:1982:TIT

[Ker82b]  Brian W. Kernighan.  A

typesetter-independent TROFF. Computing Science Technical Report 97, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1982. ii + 13 pp. Revised edition. From the first section: “I will be speaking today about work in progress, instead of completed research; this was not my original intention when I chose the subject of this lecture, but the fact is I couldn’t get my computer programs working in time.” — Donald E. Knuth.

Kernighan:1984:PGL

[Ker84a] Brian W. Kernighan. PIC — a graphics language for typesetting user manual. Computing Science Technical Report 116, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1984. ?? pp. See also revised edition [Ker91].

Kernighan:1984:UDP

[Ker84b] Brian W. Kernighan. The Unix document preparation tools — a retrospective. In Miller [Mil84b], pages vii + 59. ISBN 0-906783-41-0 (hardcover), 0-906783-42-9 (paper). LCCN QA76.9.T48 I591 1984.

Kernighan:1985:RWU

[Ker85] Brian W. Kernighan. Recent work in Unix document preparation tools. In Proceedings of the European UNIX Users Group Conference, September, Copenhagen, Denmark, page ?? European UNIX Users Group, Buntingford, Herts, UK, September 1985.

Kernighan:1989:USD

[Ker89] Brian W. Kernighan. The UNIX system document preparation tools: a retrospective.

AT&T Technical Journal, 68 (4):5–20, August 1989. CO-DEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Kernighan:1990:TT

[Ker90a] B. W. Kernighan. A troff tutorial. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Kernighan:1990:PGL

[Ker90b] Brian W. Kernighan. Pic — a graphics language for typesetting. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/pic.pdf;   https:
//www.tuhs.org/cgi-bin/utree.
pl?file=V10/vol2/pic/pic.

ms. Troff input file is dated 24 September 1986, with previous report numbers M 80-1272-2 39199 39199-11, TR 85, and TR 116.

Kernighan:1991:PGL

[Ker91] Brian W. Kernighan. PIC — a graphics language for typesetting user manual. Computing Science Technical Report 116, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1991. ii + 25 pp.

                                        URL https://web.archive.
                                        org/web/*/http://cm.bell-
                                        labs.com/cm/cs/cstr/116.

ps.gz. Revised edition of December 1984 report.

Kerrigan:1993:MF

[Ker93] James F. Kerrigan. Migrating to Fortran 90. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1993. ISBN 1-56592-049-X. 315 pp. LCCN QA76.73.F28 K47 1993. US$24.95.

Kerrisk:2010:LPI

[Ker10] Michael Kerrisk. The Linux programming interface: a Linux and UNIX system programming handbook. No Starch Press, San Francisco, CA, USA, 2010. ISBN 1-59327-220-0 (hardcover). xli + 1506 pp. LCCN QA76.76.O63 K496 2010.

Kernighan:2012:TDR

[Ker12] Brian Kernighan. Tribute to Dennis Ritchie. Alcatel-Lucent 23-minute video., September 11, 2012. URL https:/

                                        /www.youtube.com/watch?v=

uxtKwJZbYr0. From an event entitled, “The Lasting Legacy of Dennis Ritchie: The Impact of Software on Society.”.

Kernighan:2019:UHM

[Ker19] Brian W. Kernighan. UNIX: A History and a Memoir. Kindle Direct Publishing, ????, 2019. ISBN 1-69597-855-2. xiii + 183 pp. LCCN ????

Kernighan:1974:RPR

[Ker4 ] Brian W. Kernighan. RAT-FOR — a preprocessor for a rational Fortran. Computing Science Technical Report 55, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974 (??). ?? pp.

Ketkar:1999:WNS

[Ket99] Priyadarshan Ketkar. Working with Netscape Server on HP-UX. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-095972-3. xviii + 251 pp. LCCN TK5105.8885.N495K38 1999.

URL http://www.phptr.com/
ptrbooks/ptr_0130959723.
html.

Keuffel:1992:CU

[Keu92] Warren Keuffel. In CASE of UNIX. Computer Language Magazine, 9(1):69–??, January 1992. CODEN COMLEF. ISSN 0749-2839.

Kempf:1997:ORM

[KFO97] Renate Kempf, Chris Frazier, and OpenGL Architecture Review Board. OpenGL reference manual: the official reference document to OpenGL, version 1.1. Addison-Wesley Developers Press, Reading, MA, USA, second edition, 1997. ISBN 0-201-46140-4. x + 490 pp. LCCN T385 .O642 1997.

Klappheck:2000:BLE

[KGG00] Günther Klappheck, Peter Glinsky, and Frank Gehrke. Das Buch — LINUX Edition 2000: [jetzt zur S.u.S.E. 6.4, Installation und Bedienung von LINUX und seinen Werkzeugen, LINUX im Netzwerk und Internet, LINUX und ISDN, der Desktop KDE 1.x, VMware unter LINUX]. Sybex, Düsseldorf, Germany, 2000. ISBN 3-8155-0175-X. xxviii + 906 pp. LCCN ???? DM 69.95.

Kowalski:1990:CIT

[KGP90] T. J. Kowalski, H. H. Goguen, and J. J. Puttress. The C interpreter: a tutorial for Cin version 0.18. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Kroah-Hartman:2007:LKN

[KH07] Greg Kroah-Hartman. Linux kernel in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-10079-5. xiii + 182 pp.

LCCN QA76.76.O63 K786

2007.  URL http://www.
kroah.com/lkn/;   http://
www.loc.gov/catdir/enhancements/
fy0728/2007274361-d.html;
 http://www.loc.gov/catdir/
toc/fy0713/2007274361.html;
 http://www.oreilly.com/
catalog/9780596100797.

Khan:2002:PLP

[Kha02]  Ashfaq A. Khan.  Practi-

cal Linux programming: device drivers, embedded systems, and the Internet. Charles River Media, Inc., Hingham, MA, USA, 2002. ISBN 1-58450-096-4 (paperback). xv + 420 pp. LCCN QA76.76.O63 K497 2002.

Kite:2010:MXS

[KHG10] Robert Kite, Michele Hjörleifsson, and Patrick Gallagher. Mac OS X security and mobility v10.6. Apple training series. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 2010. ISBN 0-321-63535-3 (paperback). xviii + 317 pp. LCCN QA76.76.O63 K58 2010.

Korff:2005:MFO

[KHP05] Yanek Korff, Paco Hope, and Bruce Potter. Mastering FreeBSD and OpenBSD security. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00626-8. xiv + 445 pp. LCCN QA76.76.O63 K67

Kurihara:1982:APE

[KI82] Masatoshi Kurihara and Yukio Ikadai. Application programming environment on UNIX. In Usr Group [Usr82], pages 178– ?? Abstract only.

Kedem:1999:BFA

[KI99] Gershon Kedem and Yuriko Ishihara. Brute force attack on UNIX passwords with SIMD computer. In USENIX [USE99b], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

                                        http://db.usenix.org/publications/
                                        library/proceedings/sec99/
                                        kedem.html.

Kiesling:1999:LYB

[Kie99] Robert Kiesling. Linux and the Y2K bug. Sys Admin: The Journal for UNIX Systems Administrators, 8(1):16, 18–20, January 1999. CODEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/

.

Kiesling:2002:OUE

[Kie02] Robert Kiesling. ODBC in Unix environments. Dr. Dobb’s Journal of Software Tools, 27 (12):16, 18–20, 22, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/documents/s=7718/
                                        ddj2012b/.

Killian:1984:PF

[Kil84] T. J. Killian. Processes as files. In USENIX [USE84a], pages 203–207. URL http:

//lucasvr.gobolinux.org/
etc/Killian84-Procfs-USENIX.
pdf.

Killian:1990:CMU

[Kil90] T. J. Killian. Computer music under the 10th Edition UNIX system. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Kilgard:1993:OXI

[Kil93] Mark J. Kilgard. OpenGL & X: An introduction. The X Journal, 3(2):36–38, 40, 42–44, 46–47, 50–51, November/December 1993.

Kilgard:1994:UOX

[Kil94]  Mark J. Kilgard.  Using

OpenGL with Xlib. The X Journal, 4(1):46–65, January/February 1994.

Kilgard:1996:OPX

[Kil96] Mark J. Kilgard. OpenGL Programming for the X Window System. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48359-9. xxviii + 542 pp. LCCN T385.K487 1996. US$39.76.

Kimball:1995:XC

[Kim95] Paul E. Kimball. The X-Toolkit Cookbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-973132-6. xxi + 668 pp. LCCN QA76.76.W56 K56 1995. ????

King:1983:ICD

[Kin83] Laura L. King. The Informix commercial DBMS for UNIX. In Association [Ass83], pages 245–?? Abstract only.

Kinch:TB10-1-23

[Kin89] Richard Kinch. Turbo - : A new port in C for and MS-DOS. TUGboat, 10(1): 23–24, April 1989. ISSN 0896-

                                        3207. URL https://tug.org/
                                        TUGboat/tb10-1/tb23kinch.
                                        pdf.

King:1993:WGC

[Kin93] Peggy King. When Groupware Comes to UNIX. UNIX/World, 10(3):57–??, March 1993. ISSN 0739-5922.

Kirby:1977:RNM

[Kir77] D. B. Kirby. Reissue no. 1 of MAC-8 assembler user’s manual. Technical Memorandum 1150 (TM 77-2265-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 20, 1977. ?? pp.

Kirton:1984:REG

[Kir84] P. Kirton. RFC 911: EGP Gateway under Berkeley UNIX 4.2, August 22, 1984. URL

ftp://ftp.internic.net/rfc/
rfc911.txt;   ftp://ftp.
math.utah.edu/pub/rfc/rfc911.

txt. Status: UNKNOWN.

Kirch:1995:LNAa

[Kir95a] Olaf Kirch. Linux Network Administrator’s Guide. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1995. ISBN 0-916151-75-1. 289 pp. LCCN ???? US$18.95.

Kirch:1995:LNAb

[Kir95b] Olaf Kirch. Linux Network Administrator’s Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, January 1995. ISBN 1-56592-087-2. xxxi + 335 pp. LCCN QA76.76.O63 K566 1995. US$24.95. URL http:/

/www.oreilly.com/catalog/
9781565920873;   http://
www.oreilly.com/catalog/
linag.

Kirsanov:2009:BID

[Kir09] Dmitry Kirsanov. The book of Inkscape: the definitive guide to the free graphics editor. No Starch Press, San Francisco, CA, USA, 2009. ISBN 1-59327-181-6. xx + 448 + 4 pp. LCCN T385 .K491256

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0913/2009023973-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1001/2009023973-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1010/2009023973-t.html.

Kao:1993:FFI

[KIT93] W. I. Kao, R. K. Iyer, and D. Tang. FINE: A fault injection and monitoring environment for tracing the UNIX system behavior under faults. IEEE Transactions on Software Engineering, 19(11):1105– 1118, November 1993. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=256857.

Kivolowitz:1984:OSM

[Kiv84] Perry S. Kivolowitz. Optical storage management under the UNIX operating system. In Software Tools Users Group [Sof84], pages 297–311. LCCN QA76.8.U65 U83 1984.

Kern:1998:BNE

[KJG^{+}98] Harris Kern, Randy Johnson, Stuart Galup, Dennis Horgan, and Mark Cappel. Building the New Enterprise: People, Processes, and Technology. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1998. ISBN 0-13-079671-9. xvi + 382 pp. LCCN HD30.37.B843

                                        1998. URL http://www.sun.
                                        com/books/catalog/Kern4/.

Kern:1997:NNE

[KJH^{+}97] Harris Kern, Randy Johnson, Michael W. Hawkins, Howie Lyke, William Kennedy, and Mark Cappel. Networking the new enterprise: the proof not the hype. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1997. ISBN 0-13-263427-9. xx + 264 pp. LCCN HD30.37 .N48 1997.

Kern:1996:MNE

[KJHL96] Harris Kern, Randy Johnson, Michael W. Hawkins, and Andrew Law. Managing the New Enterprise: The Proof, Not the Hype. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1996. ISBN 0-13-231184-4. xxiv + 212 pp. LCCN QA76.9.C55 M35 1996. URL

http://www.sun.com/books/
catalog/kern2/index.html.

Kay:1989:CPU

[KK89a] Judy Kay and Bob Kummerfeld. C programming in a UNIX environment. International computer science series. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-13-109760-1 (paperback); 0-201-12912-4 (paperback). xii + 340 pp. LCCN QA76.73.C15 K39 1988. See book review [Nel90].

Kingsbury:1989:JPR

[KK89b] Brent A. Kingsbury and John T. Kline. Job and process recovery in a UNIX-based operating system. In USENIX Association [USE89f], pages 355– 364.

Korth:1984:SUD

[KKF^{+}84] Henry F. Korth, Gabriel M.

Kuper, Joan Feigenbaum, Allen Van Gelder, and Jeffrey D. Ullman. System/U: a database system based on the universal relation assumption. ACM Transactions on Database Systems, 9(3):331–347, September 1984. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tods/1984-
9-3/p331-korth/p331-korth.
pdf;   http://www.acm.org/
pubs/citations/journals/
tods/1984-9-3/p331-korth/.

Kernighan:1970:EHP

[KL70] B. W. Kernighan and S. Lin. An efficient heuristic procedure for partitioning graphs. The Bell System Technical Journal, 49(2):291–307, February 1970. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol49/bstj49-2-291.
pdf;   http://www.alcatel-
lucent.com/bstj/vol49-1970/
articles/bstj49-2-291.pdf;
 https://ieeexplore.ieee.
org/document/6771089.

Kernighan:1973:HSS

[KL73] B. W. Kernighan and S. Lin. Heuristic solution of a signal design optimization. The Bell System Technical Journal, 52(7):1145–1159, September 1973. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/

images/Vol52/bstj52-7-1145.
pdf;   http://www.alcatel-
lucent.com/bstj/vol52-1973/
articles/bstj52-7-1145.pdf.

Kernighan:1979:LPC

[KL79] Brian W. Kernighan and Michael E. Lesk. The LEARN program for computer instruction. Technical Memorandum 1104 (TM 79-1274-6, TM 79-1273-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 31, 1979. ?? pp.

Kernighan:1982:UDP

[KL82] Brian W. Kernighan and Michael E. Lesk. UNIX document preparation. In Nievergelt et al. [NCNS82], pages 1– 20. ISBN 0-444-86493-8. LCCN Z244 .D63 1982. US$46.50.

Kuznetsov:1992:SAU

[KL92] S. D. Kuznetsov and V. V.

Leonas. Soviet association of Unix users and its first conference. Programming and Computer Software; translation of Programmirovaniye (Moscow, USSR) Plenum, 17 (3):??, March 1992. CODEN PCSODA. ISSN 0361-7688 (print), 1608-3261 (electronic).

URL https://www.osti.gov/
biblio/458760.

Keogh:1994:OCG

[KL94] James Keogh and Remon Lapid. Open Computing — Guide to the Best Free UNIX Utilities. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994.

ISBN 0-07-882046-4. xix + 290 pp. LCCN QA76.76.O63 K49 1994. US$34.95. Includes CD-ROM.

Kleiman:1986:VAM

[Kle86] S. R. Kleiman. Vnodes: An architecture for multiple file system types in Sun UNIX. In USENIX Association [USE86c], pages 238–247.

Klecker:2000:DGL

[Kle00] Joel Espy Klecker. Debian GNU-Linux 2.2 Potato. Alcove, Madrid, Spain, 2000. ISBN ???? LCCN ???? Includes four CD-ROMs.

Kim:2002:MCR

[KLK^{+}02] JeongWon Kim, YoungUhg Lho, YoungJu Kim, Kwang-Baek Kim, and SeungWon Lee. A memory copy reduction scheme for networked multimedia service in Linux kernel. Lecture Notes in Computer Science, 2510:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer.de/
                                        link/service/series/0558/
                                        bibs/2510/25100188.htm;
                                        http://link.springer.de/
                                        link/service/series/0558/
                                        papers/2510/25100188.pdf.

Kitalong:1992:ISS

[KLM92] Karla Saari Kitalong, Steven R. Lee, and Paul Marzin. Inside Solaris: SunOS and OpenWindows. New Riders Publishing, Carmel, IN, USA, 1992. ISBN 1-56205-032-X. xviii + 715 pp. LCCN QA76.8.S86 K57 1992. US$29.95, CDN$37.95.

Kernighan:1978:UTS

[KLO78] Brian W. Kernighan, Michael E.

Lesk, and Joseph F. Ossanna, Jr. UNIX time-sharing system: Document preparation. The Bell System Technical Journal, 57(6):2115–2135, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2115.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-2115.pdf;
 https://www.tuhs.org/Archive/
Documentation/Papers/BSTJ/
bstj57-6-2115.pdf.

Kaufer:1988:SCI

[KLP88] Stephen Kaufer, Russell Lopez, and Sesha Pratap. Saber-C — an interpreter-based programming environment for the C language. In USENIX Association [USE88g], pages 161–171. ISBN ???? LCCN ????

Kernighan:1978:UPE

[KM78] Brian W. Kernighan and John R. Mashey. The UNIX programming environment. Technical Memorandum 1447, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1978. ?? pp.

Kernighan:1979:UPE

[KM79a] Brian W. Kernighan and John R. Mashey. The UNIX programming environment. Software—Practice and Experience, 9(1):1–15, January 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). Also in IEEE Computer, Vol. 14 (4), April 1981.

Klein:1979:URD

[KM79b] R. L. Klein and J. F. Maranzano. UNIX/RT on a DEC VAX-11/780 project contract. Technical Memorandum 1372, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 20, 1979. ?? pp.

Kernighan:1981:UPE

[KM81] Brian W. Kernighan and John R. Mashey. The Unix programming environment. Computer, 14(4):12–24, April 1981. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL https:

                                        //www.computer.org/csdl/
                                        mags/co/1981/04/01667315.
                                        pdf.

Kernighan:1982:UOS

[KM82] Brian W. Kernighan and Samuel P. Morgan. The UNIX operating system: a model for software design. Science, 215(4534):779–783, February 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL https:

//www.science.org/doi/10.
1126/science.215.4534.779.

Kridle:1983:PED

[KM83] Bob Kridle and Kirk McKusick. Performance effects of disk subsystem choices for Vax systems running 4.2BSD UNIX. In Software Tools Users Group [Sof83], pages 155–169. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Karels:1986:NVM

[KM86a] M. Karels and M. McKusick. A new virtual memory implementation for Berkeley UNIX. In ????, editor, Proceedings of the European UNIX Users Group Meeting, Manchester, England, September 1986, pages 451– 460. ????, ????, 1986.

Karels:1986:TCF

[KM86b] M. Karels and M. McKusick.

Towards a compatible filesystem interface. In ????, editor, Proceedings of the European UNIX Users Group Meeting, Manchester, England, September 1986, pages 481–496. ????, ????, 1986.

Karels:1986:NPM

[KM86c] Michael J. Karels and Marshall Kirk McKusick. Network performance and management with 4.3BSD and IP/TCP. In USENIX, editor, USENIX Association Summer Conference Proceedings, Atlanta 1986, June 9–13, 1986, Atlanta, Georgia USA, pages 182–188. USENIX Association, Berkeley, CA, USA, 1986. URL http:

                                        //www.gbv.de/dms/tib-ub-
                                        hannover/303820217.pdf.

Kogure:1987:USV

[KM87] Hiromichi Kogure and Richard McGowan. A UNIX System V STREAMS TTY implementation for multiple language processing. In USENIX Association [USE87d], pages 323–336.

Kasper:1995:ASI

[KM95] Paul Anthony Kasper and Alan L. McClellan. Automating Solaris Installations: a Custom JumpStart Guide. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1995. ISBN 0-13-312505-X. xx + 282 pp. LCCN QA76.76.O63K368 1995. URL

                                        http://www.sun.com/books/
                                        catalog/kasper/index.html.

Kataoka:1992:MIO

[KMKO92] Yutaka Kataoka, Masato Morisaki, Hiroshi Kuribayashi, and Hiroyoshi Ohara. A model for input and output of multilingual text in a windowing environment. ACM Transactions on Information Systems, 10(4): 438–451, October 1992. CO-DEN ATISET. ISSN 1046-

                                        8188. URL http://www.acm.
                                        org:80.

Knox:2020:WRH

[KMP^{+}20] Shawn Knox, Steven Moghadam, Kenny Patrick, Anh Phan, and Kim-Kwang Raymond Choo. What’s really ‘ Happning’? A forensic analysis of Android and iOS Happn dating apps. Computers & Security, 94(??):Article 101833, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167404820301073.

Kubasiak:2009:MXI

[KMV^{+}09] Ryan R. Kubasiak, Sean Morrissey, Jesse Varsalone, et al. Macintosh OS X, iPod, and iPhone forensic analysis DVD toolkit. Syngress Publishing, Inc., Rockland, MA, USA, 2009. ISBN 1-59749-297-3 (paperback). xix + 551 pp. LCCN QA76.9.A25 V3854

2009. URL http://www.loc.
gov/catdir/enhancements/
fy0903/2008046538-d.html.

Kirk:2010:PMP

[KmWH10] David B. Kirk and Wen mei W. Hwu. Programming Massively Parallel Processors: a Hands-on Approach. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2010. ISBN 0-12-381472-3. xviii + 258 pp. LCCN QA76.642 .K57 2010. Chapter 7 (pages 125– 140) discusses GPU floatingpoint considerations.

Khalidi:1993:IUO

[KN93] Yousef A. Khalidi and Michael N. Nelson. An implementation of UNIX on an object-oriented operating system. In USENIX Association [USE93f], pages 469–479. ISBN 1-880446-48-0. LCCN QA 76.76 O63

                                        U84 1993. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sd93/.

Kahle:1988:UCM

[KNI88] Brewster U. Kahle, William A. Nesheim, and Marshall Isman. UNIX and the Connection Machine operating system. In USENIX Association [USE88f], pages 93–107. ISBN ???? LCCN ????

Knight:1992:CDU

[Kni92] David M. Knight. Coherent 4.0: a desktop UNIX clone. Sys Admin: The Journal for UNIX Systems Administrators, 1(4):37–??, November/December 1992. CODEN SYADE7. ISSN 1061-2688.

Kobara:1991:VDO

[Kob91] Shiz Kobara. Visual Design with OSF/Motif. Addison-Wes-ley, Reading, MA, USA, 1991. ISBN 0-201-56320-7. xxv + 260 pp. LCCN QA76.76.W56 K63 1991.

Kobert:1999:GHA

[Kob99] Jeannie Johnstone Kobert. Guide to high availability: con-

                                        figuring boot/root/swap. Sun

BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1999. ISBN 0-13-016306-6. xvi + 104 pp. LCCN QA76.8.S86K62 1999. US$24.00. URL books/ha.

pdf;   http://www.phptr.
com/ptrbooks/ptr_0130163066.
html;   http://www.sun.com/
books/catalog/kobert/. Part

No.: 805-7711-05 July 1999, Revision 50.

Koch:1987:DFA

[Koc87] Philip D. L. Koch. Disk file allocation based on the buddy system. ACM Transactions on Computer Systems, 5(4):352– 370, November 1987. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1987-5-4/p352-koch/.

Kochan:2009:POC

[Koc09] Stephen G. Kochan. Programming in Objective-C 2.0. Developer’s library. Addison Wesley Professional, Upper Saddle River, NJ, USA, second edition, 2009. ISBN 0-321-56615-7 (paperback). xv + 600 pp. LCCN QA76.73.O115 K63 2009.

Kodosky:1982:UEA

[Kod82a] Jeffrey L. Kodosky. UNIX etc.

at National instruments. In Usr Group [Usr82], pages 141–149.

Kodosky:1982:UEN

[Kod82b] Jeffrey L. Kodosky. UNIX etc.

at National instruments. In USENIX [USE82], pages 141– 149.

Koenig:1978:CLP

[Koe78a] A. Koenig. A C library performance problem. Technical Memorandum 1267 (MF 78-8234-61), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 8, 1978. ?? pp.

Koenig:1978:CCE

[Koe78b] A. Koenig. Changes in the C environment for UNIX/TS Edition 1.0. Technical Memorandum 1323 (MF 78-8234-84), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 27, 1978. ?? pp.

Koenig:1978:VLA

[Koe78c] A. Koenig. Variable length argument lists in C. Technical Memorandum 1268 (MF 78-8234-64), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 12, 1978. ?? pp.

Koenig:1979:TDC

[Koe79a] A. Koenig. Thoughts on debugging C programs. Technical Memorandum 1486, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 6, 1979. ?? pp.

Koenig:1979:UFC

[Koe79b] A. Koenig. Update files for the C Standard Library. Technical Memorandum 1364, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 6, 1979. ?? pp.

Koenig:1984:ASD

[Koe84] Andrew Koenig. Automatic software distribution. In Software Tools Users Group [Sof84], pages 312–322. LCCN QA76.8.U65 U83 1984.

Koenig:1989:CTP

[Koe89] Andrew Koenig. C traps and pitfalls. Computing Science Technical Report 123, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1989. 29 pp. From the title footnote: “This paper, greatly expanded, is the basis for the book C Traps and Pitfalls (Addison-Wesley, 1989, ISBN 0-201-17928-8); interested readers may wish to refer there as well.”.

Koenig:1990:SPL

[Koe90]  A. Koenig.  The Snocone

programming language. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Kofler:2000:LIC

[Kof00] Michael Kofler. LINUX: installation, configuration, and use. Addison-Wesley, Reading, MA, USA, second edition, 2000. ISBN 0-201-59628-8. xxv + 772 pp. LCCN QA76.76.O63 K64413 2000. Includes CD-ROM.

Kohler:1999:EIS

[Koh99] Walt Kohler. EMC information sharing: direct access to MVS data from UNIX and NT. In Delis et al. [DFG99], pages 523–

524. ISBN ???? ISSN 0163-5808 (print), 1943-5835 (electronic). LCCN QA1 .A87. URL

http://www.acm.org/pubs/
articles/proceedings/mod/
304182/p523-kohler/p523-
kohler.pdf;   http://www.
acm.org/pubs/citations/proceedings/
mod/304182/p523-kohler/.

Kohlhepp:2006:RWL

[Koh06] Bayard Kohlhepp. Review of “Windows and Linux Integration: Hands-on Solutions for a Mixed Environment by Jermy Moskowitz and Thomas Boutell,” Sybex Inc., 2005, $59.99, ISBN 0-7821-4428-4. ACM Queue: Tomorrow’s Computing Today, 4(6): 58, July/August 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [MB05].

Kolettis:1977:ESP

[Kol77] N. J. Kolettis. Extended shell — a potential real time interpreter. Technical Memorandum 1228 (TM 77-4145-01), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 25, 1977. ?? pp.

Kolstad:1993:RUP

[Kol93] Rob Kolstad. Review: UNIX power tools. ;login: the USENIX Association newsletter, 18(3):38–39, May/June 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Kolstad:1994:RQC

[Kol94] Rob Kolstad. Review: a Quarter Century of UNIX. ;login: the USENIX Association newsletter, 19(3):35–??, May/ June 1994. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Komarinski:1996:LCE

[Kom96] Mark F. Komarinski. Linux companion: the essential guide for users and system administrators. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-231838-5. xii + 191 pp. LCCN QA76.76.O63 K648 1996.

Komarinski:2000:CRL

[Kom00] Mark F. Komarinski. The Complete Red Hat Linux Training Course. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-088223-2. xii + 405 pp. LCCN QA76.76.O63 K6485 2000. US$99.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130882232.html.

Komarinski:2002:PIT

[Kom02] Mark Komarinski. PTG Interactive’s Training Course for Red Hat Linux: a Digital Seminar on CD-ROM. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-034799-X. ???? pp. LCCN ???? US$69.99.

Kong:2007:DBR

[Kon07] Joseph Kong. Designing BSD rootkits: an introduction to kernel hacking. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-142-5. xvii + 136 pp. LCCN QA76.76.O63 K649

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007007644-d.html;
 http://www.loc.gov/catdir/
toc/ecip0711/2007007644.
html.

Kong:2012:FDD

[Kon12] Joseph Kong. FreeBSD device drivers: a guide for the intrepid. No Starch Press, San Francisco, CA, USA, 2012. ISBN 1-59327-204-9. xxv + 322 pp. LCCN QA76.76.D49 K66

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1211/2012011862-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1211/2012011862-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1211/2012011862-t.html.

Kopp:2000:MCC

[Kop00] Carlo Kopp. Managing cluster computers. Dr. Dobb’s Journal of Software Tools, 25(7): 21–26, July 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_07/cluster.
txt.

Kopper:2005:LEC

[Kop05] Karl Kopper. The Linux Enterprise Cluster: build a highly available cluster with commodity hardware and free software. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-036-4. xxiv + 430 pp. LCCN QA76.58 .K67 2005; QA76.58 K67 2005eb. URL http:/

/www.oreilly.com/catalog/
9781593270360.

Korty:1989:SLL

[Kor89]  Joseph A. Korty.  Sema:

a Lint-like tool for analyzing semaphore usage in a multithreaded UNIX kernel. In USENIX Association [USE89f], pages 113–123.

Korn:1997:UWN

[Kor97a] David Korn. From UNIX to Windows NT — a long day’s journey into night, 1997.

URL  http://www.usenix.
org/publications/library/
proceedings/lisa97/invited_
talks/scarr/sld028.htm. Un-

published keynote address at the Large Scale System Administration of Windows NT Workshop, 1997, August 14– 16, 1997, Seattle, Washington, USA.

Korn:1997:PUW

[Kor97b]  David G. Korn.  Porting

UNIX to Windows NT. In USENIX [USE97c], pages 43–

57. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/ana97/korn.html;
                                         http://www.usenix.org/publications/
                                        library/proceedings/ana97/
                                        summaries/korn.html.

Korn:1997:UUWa

[Kor97c]  David G. Korn.  UWIN

— UNIX for Windows. In USENIX [USE97a], page ?? ISBN ???? LCCN ????

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/usenix-nt97/
                                        korn.html.

Korn:1997:UUWb

[Kor97d]  David G. Korn.  UWIN

— UNIX for Windows. In USENIX [USE97d], pages 133–

145. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845

                                        1997.  URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt97/korn.html.

Koren:2006:SLK

[Kor06] Oded Koren. A study of the Linux kernel evolution. Operating Systems Review, 40(2):110– 112, April 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Kossmann:1987:GSS

[Kos87] Heinz Kossmann. A graphic SDL support environment. Computer Networks and ISDN Systems, 13(2):91–96, 1987. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic).

Kosko:1992:PRM

[Kos92] Linda Kosko. PHIGS Reference Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-91-9. xv + 1099 pp. LCCN T385 .P487 1992. US$39.95.

Koscielny:1996:WNF

[Kos96] Al Koscielny. A Windows NT fileserver with UNIX clients. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):43–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Koskelin:2000:LG

[Kos00] Al Koskelin. Linux games.

No Starch Press, San Francisco, CA, USA, 2000. ISBN 1-886411-33-6. ???? pp. LCCN QA76.76.O63K676 2000.

Kowalski:1979:FUT

[Kow79] T. J. Kowalski. FSCK — the UNIX/TS file system check program. Technical Memorandum 1463 (TM 79-3624-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1, 1979. ?? pp.

Kowalski:1990:FUF

[Kow90] T. J. Kowalski. Fsck — the UNIX File System Check program. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Kernighan:1984:UPE

[KP84] Brian W. Kernighan and Rob Pike. The UNIX Programming Environment. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1984. ISBN 0-13-937699-2 (hardcover), 0-13-937681-X (paperback). x + 357 pp. LCCN QA76.76.O63 K48 1984; QA76.8.U65 K46 1984.

URL http://wiki.tuhs.org/
doku.php?id=publications:
unix_prog_environment.

Kernighan:1999:REL

[KP99] Brian W. Kernighan and Rob Pike. Regular expressions: Languages, algorithms, software. Dr. Dobb’s Journal of Software Tools, 24(4): 19–22, April 1999. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/1999/1999_04/
regexp.txt;   http://www.
ddj.com/ftp/1999/1999_04/
regexp.zip;   http://www.
drdobbs.com/architecture-
and-design/regular-expressions/
184410904;   http://www.
drdobbs.com/architecture-

and-design/regular-expressions/
architecture-and-design/
sourcecode/regular-expressions/
30200909;   http://www.drdobbs.
com/architecture-and-design/
regular-expressions/architecture-
and-design/sourcecode/regular-
expressions/30200910. See

also [Tho68, Cox07, Cox09, Cox10, Cox12].

Kavka:1993:EDM

[KPG93] C. Kavka, M. Printista, and R. Gallard. Extending device management in Minix. Operating Systems Review, 27(2):35– 43, April 1993. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Korn:1979:FOR

[KPSS79] D. G. Korn, M. A. Pilla, N. M.

Scribner, and B. W. Stuck. Functional operation of the Recent Change Memory Administration (RCMAS). Technical Memorandum 2, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1, 1979. ?? pp.

Knickerbocker:2002:AMM

[KPT^{+}02] J. U. Knickerbocker, F. L.

Pompeo, A. F. Tai, D. L. Thomas, R. D. Weekly, M. G. Nealon, H. C. Hamel, A. Haridass, J. N. Humenik, R. A. Shelleman, S. N. Reddy, K. M. Prettyman, B. V. Fasano, S. K. Ray, T. E. Lombardi, K. C. Marston, P. A. Coico, P. J. Brofman, L. S. Goldmann, D. L. Edwards, J. A. Zitz, S. Iruvanti, S. L. Shinde, and H. P. Longworth. An advanced multichip module (MCM) for high-performance UNIX servers. IBM Journal of Research and Development, 46 (6):779–804, ???? 2002. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        rd/466/knickerbocker.pdf.

Kernighan:1977:MMP

[KR77] Brian W. Kernighan and Dennis M. Ritchie. The M4 macro processor. Technical Memorandum 1160, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1, 1977. 2 + 6 pp.

Kernighan:1978:CPL

[KR78] Brian W. Kernighan and Dennis M. Ritchie. The C Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1978. ISBN 0-13-110163-3. x + 228 pp. LCCN QA76.73 .C15 K47 1978.

Kernighan:1979:UP

[KR79] Brian W. Kernighan and Dennis M. Ritchie. UNIX programming. Technical Memorandum 1073, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 26, 1979. ?? pp.

Kernighan:1988:CPL

[KR88] Brian W. Kernighan and Dennis M. Ritchie. The C Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1988. ISBN 0-13-110362-8 (paperback), 0-13-110370-9 (hardcover), 0-13-115817-1, 0-13-308621-6 (e-book), 0-13-308624-0. xii + 272 pp. LCCN QA76.73.C15 K47 1988.

URL http://wiki.tuhs.org/
doku.php?id=publications:
c_prog_language. This book

is 200% a must for any C programmers …. The answers to the exercises can be found in [TG89].

Kranc:1978:PIU

[Kra78a] Morris Kranc. Performance improvements in UNIX pipes and ports. Report 3970, Bolt Beranek and Newman Inc., Cambridge, MA, USA, November 1978. 22 pp. URL

https://www.tuhs.org/cgi-
bin/utree.pl?file=BBN-V6/
doc/ipc/port.runout.

Kraus:1978:PPI

[Kra78b] J. S. Kraus. A peripheral processor interactive control and data acquisition system. Technical Memorandum 1289 (TM 78-1131-22), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 23, 1978. ?? pp.

Kramer:1983:LLN

[Kra83] Steven M. Kramer. LINUS (Leading Into Noticeable UNIX Security). In Association [Ass83], pages 143–?? Abstract only.

Kramer:1988:IAC

[Kra88a] Steven M. Kramer. On incorporating access control lists into the UNIX operating system. In USENIX Association [USE88e], pages 38–

48. LCCN QA76.8.U65 U55

1988(1)-1990(2)//.

Kramer:1988:RSP

[Kra88b] Steven M. Kramer. Retaining SUID programs in a secure UNIX. In Association [Ass88d], pages 107–118.

Krafft:2005:DSC

[Kra05] Martin F. Krafft. The Debian system: concepts and techniques. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-069-

0. ???? pp. LCCN QA76.76.O63 K68 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0515/2005019963.
                                        html.

Krawetz:2007:HUS

[Kra07] Neal Krawetz. Hacking Ubuntu: serious hacks, mods, and customizations. Wiley, New York, NY, USA, 2007. ISBN 0-470-10872-X. xviii + 388 pp. LCCN QA76.76.O63 K742

                                        2007. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0741/2007003316-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0741/2007003316-
                                        d.html;   http://www.loc.

gov/catdir/toc/ecip079/2007003316.
html.

Kridle:1984:NIT

[Kri84a] Bob Kridle. New 1/2-inch tape options and trade-offs for 4.2BSD UNIX on DEC VAX processors. In USENIX Association [USE84b], pages 168–182. ISBN none. LCCN QA76.8.U65 U55 1984.

Kridle:1984:NTO

[Kri84b] Bob Kridle. New 1/2-inch tape options and trade-offs for 4.2BSD UNIX on DEC VAX processors. In USENIX Association [USE84b], pages 168–182. ISBN none. LCCN QA76.8.U65 U55 1984.

Krieger:1991:NTU

[Kri91] Zander Krieger. Nøglen til UNIX: System V og BSD 4.3. (Danish) [Key to UNIX: System V and BSD 4.3]. Teknisk forlag, Copenhagen, Denmark, 1991. ISBN 87-571-1265-7. 167 pp. LCCN ???? Translated and extended by Jørgen Floes.

Krishnamurthy:1995:PRU

[Kri95] Balachander Krishnamurthy.

Practical Reusable Unix Software. Wiley, New York, NY, USA, 1995. ISBN 0-471-05807-6. xxv + 370 pp. LCCN QA76.76.O63 P724 1995. US$34.95. URL http:

//www.research.att.com/%
7Egsf/publications/prus-
1995-1.pdf.

Krol:1992:WIU

[Kro92] Ed Krol. The Whole Internet User’s Guide & Catalog. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-025-2. xxiv + 376 pp. LCCN TK5105.875.I57 K86 1992. US$24.95. Perhaps the most ambitious of the new crop of books, this one is both a user’s guide and a catalog of resources in one. It is aimed at graduate students who want to use the Internet for research.

Kroeker:1999:NTN

[Kro99] Kirk L. Kroeker. New tools: Net development: Sun’s Java embedded server; MetaCreation’s Web-savvy graphics tool; WebCompiler’s HTML packaging tool. component technology: ProtoVIew Development’s ActiveX tools; Avilon Software’s load balancing component system. software development: Verilog’s test checker; Red Hat and Metrowerks’ development tools for Linux; the Object Factory’s optimization tool; Acumen Systems’s SDK for imaging; Aonix’s process-oriented lifecycle environment; Baan’s embedded software development suite. Computer, 32(5): 103–107, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r5103.pdf.

Krumins:20xx:AOL

[Kruxxa] Peteris Krumins. Awk One-Liners Explained. ????, ????, 20xx. 52 (est.) pp. URL https:

//catonmat.net/awk-book.

Krumins:20xx:SOL

[Kruxxb] Peteris Krumins. Sed One-Liners Explained. ????, ????, 20xx. 90 (est.) pp.

URL https://catonmat.net/
sed-one-liners-explained-
part-one.

Kaufman:1985:TPS

[KS85] Linda Kaufman and Norman L.

Schryer. TTGR — a package for solving partial differential equations in two space variables. Computing Science Technical Report 135, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 5, 1985. i + 73 pp.

Kuehne:2008:OPM

[KS08] Robert P. Kuehne and J. D.

Sullivan. OpenGL programming on Mac OS X: architecture, performance, and integration. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-321-35652-7 (paperback). xxxii + 330 pp. LCCN T385 .K82 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0714/2007011974.html.

Kshetri:2004:ELA

[Ksh04] Nir Kshetri. Economics of Linux adoption in developing countries. IEEE Software, 21(1):74–81, January/February 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Krieger:1994:ASF

[KSU94] Orran Krieger, Michael Stumm, and Ron Unrau. The Alloc Stream Facility: a redesign of application-level stream I/O. Computer, 27(3):75–82, March 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karels:1988:IOT

[KTB^{+}88] Michael J. Karels, Chris Torek, James M. Bloom, Marshall Kirk McKusick, Samuel J. Leffler, and William N. Joy. Installing and operating 4.3BSDtahoe UNIX on the VAX. Report, Computer Systems Research Group, Department of Electrical Engineering and Computer Science, University of California, Berkeley, Berkeley, California 94720, USA, July 14, 1988. 66

pp.  URL http://blog.
livedoor.jp/suzanhud/BSD/
4.3BSD_Tahoe_VAX.pdf.

Kuhn:1991:IPM

[Kuh91] D. Richard Kuhn. IEEE’s Posix: making progress. IEEE Spectrum, 28(12):36–39, December 1991. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kulkarni:2006:SCS

[Kul06] Pankaj Kulkarni. S7 case study: Porting 2 million lines of C++ code to HP-UX Itanium. In Anonymous [Ano06a], page ?? ISBN ???? LCCN ???? URL http://

www.ice.gelato.org/oct06/
pres_pdf/gelato_ICE06oct_
s7study_kulkarni_s7softwaresolutions.
pdf.

Kuo:1998:UUU

[Kuo98a]  Benjamin F. Kuo.  Using

UNIX utilities to create CD-ROMs. Sys Admin: The Journal for UNIX Systems Administrators, 7(12):53–56, December 1998. CODEN SYADE7. ISSN 1061-2688. URL http://

www.samag.com/.

Kuo:1998:SEU

[Kuo98b] Peter Kuo. Special edition using Unix. Que Corporation, Indianapolis, IN, USA, third edition, 1998. ISBN 0-7897-1747-6 (paperback). xvi + 816 pp. LCCN QA76.76.O63 K86 1998.

Kuo:1999:UUL

[Kuo99] Peter Kuo. Unix: das umfassende Lern- und Nachschlagewerk; zum effektiven Umgang mit Unix-Systemen; zur System- und Netzwerkadministration unter AIX, BSD, HP-UX, Linux, SCO-Unix, SVR 4, SVR 5, UnixWare und SunOS Solaris. Markt & Technik, Buch- und Software-Verlag, Munich, Germany, 1999. ISBN 3-8272-5532-5. 981 (est.) pp. LCCN ???? Includes CD-ROM.

Kuras:1990:ESA

[Kur90] Joseph Kuras. An expert systems approach to security inspection of UNIX. In USENIX Association [USE90e], pages 55–56. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Kurani:1994:AUP

[Kur94] Bharat Kurani. Applied Unix programming, volume 2. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-304338-X (vol. 1), 0-13-304346-0 (vol. 2). xviii + 1250 pp. LCCN QA76.76.O63 K86 1994. Two volumes.

Kurani:1997:XO

[Kur97] Bharat Kurani. Applied Unix programming, volume 2. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-304338-X (vol. 1), 0-13-304346-0 (vol. 2). xviii + 1250 pp. LCCN QA76.76.O63 K86 1994. Two volumes.

Kutti:2002:CUP

[Kut02] N. S. Kutti. C and Unix programming: a comprehensive guide incorporating the ANSI and POSIX standards. Lightspeed Books, Mt. Pleasant, SC, USA, 2002. ISBN 1-929175-40-X (hardcover), 1-929175-26-4 (paperback). xviii + 661 pp. LCCN QA76.73.C15 K88 2001.

Kernighan:1989:PMP

[KV89] Brian W. Kernighan and Christopher J. Van Wyk. Page makeup by postprocessing text formatter output. Computing Systems, 2(2):103–132, Spring 1989. CODEN CMSYE2. ISSN 0895-6340.

Kernighan:1990:MMP

[KV90] B. W. Kernighan and C. J. Van Wyk. The -mpm macro package. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Kaluzny:1998:SUM

[KVCS98] Stephen P. Kaluzny, Silvia C.

Vega, Tamre P. Cardoso, and Alice A. Shelly. S+SpatialStats: User’s manual for Windows and UNIX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98226-4. xvi + 327 pp. LCCN QA278.2 .S18 1998.

Kochan:1985:USP

[KW85] Stephen G. Kochan and Patrick H. Wood. UNIX shell programming. Hayden UNIX system library. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1985. ISBN 0-8104-6309-1 (paperback), 0-471-83900-0. 422 pp. LCCN QA76.76.O63 K64 1985. US$22.50 (est.).

Kochan:1989:UN

[KW89] Stephen G. Kochan and Patrick H. Wood, editors. UNIX Networking. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1989. ISBN 0-672-48440-4. viii + 400 pp. LCCN QA76.76.O63 U546 1989.

Kochan:1990:USP

[KW90] Stephen Kochan and Patrick Wood. UNIX Shell Programming. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, revised. edition, 1990. ISBN 0-672-48448-X. xi + 490 pp. LCCN QA76.76.O63 K64 1992. US$29.95. Highly recommended. A classic on using and programming Bourne Shell (and Korn Shell).

Kochan:2003:USP

[KW03] Stephen Kochan and Patrick Wood. UNIX Shell Programming. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, third edition, 2003. ISBN 0-672-32490-3. xiii + 437 pp. LCCN QA76.76.O63 K64 2003.

Keromytis:2006:COS

[KWDB06] Angelos D. Keromytis, Jason L. Wright, Theo De Raadt, and Matthew Burnside. Cryptography as an operating system service: a case study. ACM Transactions on Computer Systems, 24(1):1–38, February 2006. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Ligh:2010:MAC

[L^{+}11] Michael W. Ligh et al. Malware analyst’s cookbook and DVD: tools and techniques for fighting malicious code. Wiley, New York, NY, USA, 2011. ISBN 0-470-61303-3 (paperback). xxvi + 716 pp. LCCN QA76.9.A25 M35 2011.

Legge:1990:UFS

[LA90] Gaynor Legge and Mahir Ali.

UNIX file system behaviour and machine architecture dependency. Software—Practice and Experience, 20(11):1077– 1096, November 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lacheiner:2006:EPB

[Lac06] Hermann Lacheiner. Entwicklung einer auf Python basierenden Rich Client Platform für Linux. (German) [Development of a Python-based Rich Client Platform for Linux]. Diplome-Arbeit, Universität Linz, Linz, Austria, 2006. v + 86 pp.

Lu:2014:SLF

[LADADL14] Lanyue Lu, Andrea C. Arpaci-Dusseau, Remzi H. Arpaci-Dusseau, and Shan Lu. A study of Linux file system evolution. ACM Transactions on Storage, 10(1):3:1–3:??, January 2014. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Lalwanai:1993:PIM

[Lal93] Rajesh Lalwanai. POSIX interface for MPE/iX. Hewlett– Packard Journal, 44(3):41– ??, June 1, 1993. CODEN HPJOAX. ISSN 0018-1153.

Lamb:1983:TUS

[Lam83] J. Eli Lamb. Towards a UNIX system Ada programming support environment. In Association [Ass83], pages 143–?? Abstract only.

Lamb:1990:LTV

[Lam90] Linda Lamb. Learning the vi Editor. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 1990. ISBN 0-937175-67-6. xvi + 173 pp. LCCN QA76.8.U65 L35 1988. US$21.95. A very good guide to vi and ex commands …. With a quick reference card …. Other good vi stuff (including an online tutorial) is obtained by anonymous ftp from cs.

                                        uwp.edu (in pub/vi).

Lankford:1984:USV

[Lan84] Jeffrey P. Lankford. UNIX system V and 4BSD performance. In Software Tools Users Group [Sof84], pages 228–236. LCCN QA76.8.U65 U83 1984.

Langer:2001:MX

[Lan01] Maria Langer. Mac OS X. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 2001. ISBN 0-201-70900-7. xiii + 247 pp. LCCN QA76.76.O63 L3626 2001.

Lapin:1987:PCU

[Lap87a] J. E. Lapin. Portable C and UNIX Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-686494-5. xiv + 249 pp. LCCN QA76.73.C15 L36 1987.

Lapin:PCU87

[Lap87b] J. E. Lapin. Portable C and UNIX Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-686494-5. xiv + 249 pp. LCCN QA76.73.C15 L36 1987.

Larus:1983:PCR

[Lar83] James R. Larus. On the performance of Courier remote procedure calls under 4.1c BSD. Report UCB/CSD 83/123, Computer Science Division, University of California, Berkeley, Berkeley, CA, USA, 1983.

Larkham:1994:ATL

[Lar94] Adrian Larkham. Ada — the technical language of choice for UNIX threaded applications. Ada User, 15(2):66–76, June 1994. CODEN ADUSEB. ISSN 0268-652X.

Pere:2005:GLR

[Lás05] Pere László. GNU/LINUX rendszerek üzemeltetése. (Hungarian) []. Kiskapu, Pécs, Hungary, 2005. ISBN ???? ???? pp. LCCN ????

Lathrop:2002:LSB

[Lat02] John P. Lathrop. Linux in Small Business: a Practical User’s Guide. Apress, Berkeley, CA, USA, 2002. ISBN 1-893115-46-1. 336 (est.) pp. LCCN ???? US$36.95.

Laur:1977:DDM

[Lau77] J. B. Laur. DA11B driver in MERT. Technical Memorandum 1212, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 30, 1977. ?? pp.

Laur:1979:SDS

[Lau79a] J. E. Laur. Spooler and daemon software for a printer network. Technical Memorandum 1448 (TM 79-9457-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 27, 1979. ?? pp.

Laur:1979:STM

[Lau79b] J. E. Laur. Synchronous terminals for MERT. Technical Memorandum 1459 (TM 79-9457-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 27, 1979. ?? pp.

Lavigne:2004:BHI

[Lav04] Dru Lavigne. BSD Hacks: 100 Industrial Strength Tips & Tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00679-9. xvii + 427 pp. LCCN QA76.76.O63 L372 2004. US$24.95, CAN$36.95.

Lavigne:2007:BFB

[Lav07] Dru Lavigne. The Best of FreeBSD Basics. Reed Media Services, Marysville, WA, USA, 2007. ISBN 0-9790342-2-1. 595 pp. LCCN ???? US$35.90.

URL http://reedmedia.net/
books/freebsd-basics/.

Lawson:1983:URA

[Law83a] Jim Lawson. UNIX research at Lucasfilms. In Association [Ass83], pages 167–?? Abstract only.

Lawson:1983:URL

[Law83b] Jim Lawson. UNIX research at Lucasfilms. In USENIX [USE83], pages 167–?? Abstract only.

Lawgon:1998:NBA

[Law98] George Lawgon. News briefs:

Alliance wants to give voice to the net; Intel pushes for Unix standards; groups seek to end browser wars; initiative hopes to drive basic chip research; HTML+TIME promises better multimedia; keyword browsing promises simple Web access. Computer, 31(12):21– 23, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1998/pdf/
rz021.pdf.

Lycklama:1975:SOS

[LB75a] Heinz Lycklama and D. L.

Bayer. A structured operating system for a PDP-11/45.

Technical Memorandum TM-75-1352-4, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 6, 1975. 18 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-75-1352-
4_A_Structured_Operating_
System_for_a_PDP-11.45_19750506.
pdf.

Bayer:1975:SOS

[LB75b] Heinz Lycklama and D. L. Bayer. A structured operating system for a PDP-11/45. Technical Memorandum 1052 (TM 75-1352-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 6, 1975. ?? pp.

Lycklama:1978:MOSa

[LB78a] Heinz Lycklama and D. L. Bayer. The MERT operating system. Technical Memorandum 1246 (TM 78-3114-3, TM 78-1352-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1978. 32 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-78-3114-
3_The_MERT_Operating_System_
19780422.pdf.

Lycklama:1978:MOSb

[LB78b] Heinz Lycklama and D. L. Bayer. The MERT operating system. Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 1978. 27 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/

Heinz_Tech_Memos/The_MERT_
Operating_System.pdf.

Lycklama:1978:UTSa

[LB78c] Heinz Lycklama and D. L.

Bayer. UNIX time-sharing system: The MERT operating system. The Bell System Technical Journal, 57(6): 2049–2086, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2049.
pdf.

Lai:1996:PCU

[LB96a] Kevin Lai and Mary Baker.

A performance comparison of UNIX operating systems on the Pentium. In USENIX [USE96b], pages 265–277 (or 265–278??). ISBN 1-880446-76-6. LCCN QA 76.76 O63

U88 1996. URL http://www.
usenix.org/publications/
library/proceedings/sd96/
lai.html.

Lewis:1996:TPG

[LB96b] Bil Lewis and Daniel J. Berg.

Threads Primer: a Guide to Multithreaded Programming. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1996. ISBN 0-13-443698-9. xxvi + 319 pp. LCCN QA76.642 .L478 1996.

Lewis:1998:MPP

[LB98] Bil Lewis and Daniel J.

Berg. Multithreaded programming with pthreads. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1998. ISBN 0-13-680729-1. xxx + 382 pp. LCCN QA76.76.T55 L49 1998. URL

http://www.sun.com/books/
catalog/lewis2/index.html.

Lee:2014:UBC

[LBN14] Eunji Lee, Hyokyung Bahn, and Sam H. Noh. A unified buffer cache architecture that subsumes journaling functionality via nonvolatile memory. ACM Transactions on Storage, 10(1):1:1–1:??, January 2014. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Luerkens:1997:IWN

[LBSL97] Cameron D. Luerkens, Mark J. Bartelme, Kevin K. Sizer, and James ‘Bud’ Lande. Integrating Windows 95/NT into an existing UNIX Enterprise Network. In USENIX [USE97a], page ?? ISBN ???? LCCN ???? URL http://

www.collins.rockwell.com/
presentation/ntpaper/index.
htm;   http://www.usenix.
org/publications/library/
proceedings/nt-sysadmin97/
luerkens.html.

Lycklama:1973:EUP

[LC73] Heinz Lycklama and C. Christensen. Emulation of UNIX on peripheral processors. Technical Memorandum TM-75-l352-2, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 9, 1973. 23 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-75-1352-
2_Emulation_of_UNIX_on_Peripheral_
Processors_19750109.pdf.

Lycklama:1975:EUP

[LC75] Heinz Lycklama and C. Christensen. Emulation of UNIX on peripheral processors. Technical Memorandum 1045 (TM 75-1352-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 9, 1975. ?? pp.

Lesk:1977:USD

[LC77] M. Lesk and A. Cohen. UNIX software distribution by communications link. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977. ?? pp.

Lesk:1978:USD

[LC78a] Michael E. Lesk and A. S. Cohen. UNIX software distribution by communication link. Technical Memorandum 1279 (TM 78-1274-3, TM 78-8234-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 24, 1978. ?? pp.

Lycklama:1978:MSP

[LC78b] Heinz Lycklama and C. Christensen. A minicomputer satellite processor system. Technical Memorandum 1247 (TM 78-3114-2, TM 78-1359-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1978. 14 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/

Heinz_Tech_Memos/A_Minicomputer_
Satellite_Processor_System.
pdf;   https://www.tuhs.
org/Archive/Documentation/
TechReports/Heinz_Tech_Memos/
TM-78-3114-2_A_Minicomputer_
Satellite_Processor_System_
19780322.pdf.

Lycklama:1978:UTSc

[LC78c] Heinz Lycklama and C. Christensen. UNIX time-sharing system: a minicomputer satellite processor system. The Bell System Technical Journal, 57(6):2103–2113, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2103.
pdf.

Li:1990:SPA

[LC90] Xiaobo Li and Yian-Leng Chang. Simulating parallel architectures in a distributed environment. Journal of Parallel and Distributed Computing, 9 (2):218–223, June 1990. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Low:1993:FGO

[LC93] Marie Rose Low and Bruce Christianson. Fine grained object protection in UNIX. Operating Systems Review, 27(1): 33–50, January 1993. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lee:2006:HLK

[LC06] Yueh-Feng Lee and Ruei-Chuan Chang. Hotswapping Linux kernel modules. The Journal of Systems and Software, 79(2):163–175, February 2006. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Luderer:1981:DUS

[LCH^{+}81] G. W. R. Luderer, H. Che, J. P. Haggerty, P. A. Kirslis, and W. T. Marshall. A distributed UNIX system based on a virtual circuit switch. Operating Systems Review, 15(5): 160–168, December 1981. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Lumens:2005:SLE

[LCJH05] Chris Lumens, David Cantrell, Logan Johnson, and Alan Hicks. Slackware Linux Essentials. Slackware Linux, Inc., 1164 Claremont Drive, Brentwood, CA 94513, USA, second edition, 2005. ISBN 1-57176-338-4. xxii + 262 pp. LCCN ???? URL ftp:

//ftp.slackbook.org/pub/
slackbook/slackbook-2.0.
pdf;   http://www.slackbook.
org/.

Lecarme:2013:BGC

[LD13] Olivier Lecarme and Karine Delvare. The book of GIMP: a complete guide to nearly everything. No Starch Press, San Francisco, CA, USA, 2013. ISBN 1-59327-383-5. xvii + 656 pp. LCCN TR267.5.G56 L4313

                                        2012. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy1214/2012020781-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy1214/2012020781-
                                        d.html.

Leach:1990:BRU

[Lea90] George W. Leach. Book review: UNIX System Software Readings. ;login: the USENIX Association newsletter, 15(1):15–??, January/February 1990. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Leach:1992:BRP

[Lea92] George Leach. Book review: Practical UNIX Security. ;login: the USENIX Association newsletter, 17(3):29–??, May/ June 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Ledesma:1994:PHC

[Led94] Ron Ledesma. PC Hardware Configuration Guide for DOS and Solaris. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1994. ISBN 0-13-124678-X. xxii + 331 pp. LCCN TK7887.5 .L38 1994.

Lee:1989:IUT

[Lee89] Geoffrey M. Lee. Integrating UNIX terminal services into a distributed operating system. In USENIX Association [USE89f], pages 29–42.

Lee:1992:XTB

[Lee92] Ken Lee. X Technical Bibliography, November 4 1992. A good collection of publicly available X window system technical materials. The bibliography is posted periodi-

cally on comp.windows.x and

can be obtained by anonymous

ftp from gatekeeper.dec.
com (in /pub/X11/contrib)
or export.lcs.mit.edu (in /

contrib) — look for the file

Xbibliography.

Leffler:1981:ICP

[Lef81] Samuel Jacob Leffler. An implementation of the C programming language for the Harris/6 minicomputer. Master of Science, Department of Computer Engineering and Science, Case Western Reserve University, Cleveland, OH 44106, USA, January 7, 1981. vii + 161 + 1 pp.

URL https://www.tuhs.org/
Archive/Documentation/Theses/
LefflerSamuel_ImplementationOfCProgrammingLanguage_
1981_thesis.pdf.

Laskodi:1988:UFS

[LEG88] Terry Laskodi, Bob Eifrig, and Jason Gait. A UNIX file system for a write-once optical disk. In Association [Ass88d], pages 51– 60.

Lehey:1995:PUS

[Leh95] Greg Lehey. Porting UNIX Software: From Download to Debug. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, November 1995. ISBN 1-56592-126-7. xix + 515 pp. LCCN QA76.76.O63L4472 1995. US$29.95. URL http:/

/www.ora.com/gnn/bus/ora/
item/port.html;   http://
www.oreilly.com/catalog/
port.

Lehey:1999:CF

[Leh99] Greg Lehey. The Complete FreeBSD. Walnut Creek CD-ROM, Walnut Creek, CA, USA, third edition, 1999. ISBN 1-57176-246-9; 1-57176-079-2. xxxiv + 773 pp. LCCN QA76.76.O63. Includes four CD-ROMs.

Lehey:2003:CFD

[Leh03] Greg Lehey. The Complete FreeBSD: Documentation from the Source. O’Reilly Community Press, Sebastopol, CA, USA, fourth edition, 2003. ISBN 0-596-00516-4. xxxiii + 679 pp. LCCN QA76.76.O63 L43 2003. US$44.09, CAN$69.95, UK£31.95. URL http://

press.oreilly.com/commpress.
html;   http://www.oreilly.
com/catalog/cfreebsd/desc.
html.

Leibowitz:1990:UWA

[Lei90] M. R. Leibowitz. UNIX workstations arrive! Datamation, 36 (11):24–27, 30, June 1990. CO-DEN DTMNAT. ISSN 0011-6963.

Leibfritz:1995:UPS

[Lei95a] David Leibfritz. A UNIX print server for UNIX workstations, PCs, and Macs. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):86– ??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Leininger:1995:SDT

[Lei95b] Kevin E. Leininger. Solaris developer’s tool kit. McGraw-Hill, New York, NY, USA, 1995. ISBN 0-07-911851-8 (hardcover), 0-07-911852-6 (paperback). xiii + 422 pp. LCCN QA76.76.O63 L4477 1995. US$69.00 (hardcover), US$49.95 (paperback).

Leininger:1996:ADT

[Lei96a] Kevin E. Leininger. AIX/6000 developer’s tool kit. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-911992-1. xiii + 496 pp. LCCN QA76.76.O63 L4473 1996.

Leininger:1996:HUD

[Lei96b] Kevin E. Leininger. HP-UX developer’s tool kit. J. Ranade workstation series. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-912174-8 (hardcover), 0-07-912175-6 (paperback). ix + 473 pp. LCCN QA76.76.O63 L4475 1995.

Leisher:1996:AIUa

[Lei96c] M. Leisher. An adventure in implementing Unicode support on UNIX platforms. In Unicode Consortium [Uni96a], pages C6–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Leisher:1996:AIUc

[Lei96d] M. Leisher. An adventure in implementing Unicode support on UNIX platforms. In Unicode Consortium [Uni96b], pages C6–?? ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

Leisher:1996:AIUb

[Lei96e] Mark Leisher. An adventure in implementing Unicode support on Unix platforms. In Unicode Consortium [Uni96b], page 8. ISBN ???? LCCN ????

URL ftp://crl.nmsu.edu/
CLR/multiling/unicode/paper.
ps.gz.

Leighton:1998:NDU

[Lei98] Luke Kenneth Casson Leighton. NT 3.5 / 4.0 domains for UNIX. In USENIX [USE98c], pages 85–88. ISBN 1-880446-96-0. LCCN TK5105.5 .L37

1998.  URL http://www.
usenix.org/publications/
library/proceedings/lisa-
nt98/leighton.html.

Leibovitch:1999:BCL

[Lei99] Evan Leibovitch. The business case for Linux. IEEE Software, 16(1):40–44, January/February 1999. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://computer.org/
software/so1999/s1040abs.
htm;   http://dlib.computer.
org/so/books/so1999/pdf/
s1040.pdf.

Leighton:2000:DRS

[Lei00] Luke Kenneth Casson Leighton.

DCE/RPC over SMB: Samba and Windows NT domain internals. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 1-57870-150-3. xvi + 282 pp. LCCN QA76.76.O63 L44725 2000. US$45.00.

Leler:1990:LMU

[Lel90] W. Leler. Linda meets Unix.

Computer, 23(2):43–54, February 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lennert:1987:HWU

[Len87] Dave Lennert. How to write a UNIX daemon. ;login: the USENIX Association newsletter, 12(4):17–23, July/August 1987. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Lenk:1994:SSU

[Len94] Fred Lenk. Setting up a secure user environment with standard UNIX tools. Sys Admin: The Journal for UNIX Systems Administrators, 3(3): 85–??, May/June 1994. CO-DEN SYADE7. ISSN 1061-2688.

Lendecke:1997:UFI

[Len97] Volker Lendecke. UNIX filesystems without I-nodes — the Linux kernel nfs, smbfs, and ncpfs filesystems make it possible to link Linux machines to file servers across a LAN — even though Microsoft’s SMB protocol is not designed to handle UNIX clients like Linux. Volker presents a workaround for this limitation. Dr. Dobb’s Journal of Software Tools, 22 (2):60–??, February 1997. CO-DEN DDJOEB. ISSN 1044-789X.

Lesk:1975:LLA

[Les75a] M. E. Lesk. Lex — a lexical analyzer generator. Computing Science Technical Report 39, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1975. ?? pp.

Lesk:1975:AAM

[Les75b] Michael E. Lesk. AGEN — an associated memory generator. Technical Memorandum 1064 (TM 75-1274-18), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 2, 1975. ?? pp.

Lesk:1976:TPF

[Les76a] Michael E. Lesk. Tbl — a program to format tables. Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 11, 1976. 18

pp. URL http://www.kohala.
com/start/troff/v7man/tbl/
tbl.ps.

Lesk:1976:TDU

[Les76b] Michael E. Lesk. Typing documents on the UNIX system: using the -ms macros with nroff and troff. Computing Science Technical Report 49, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1976. ?? pp.

Lesk:1977:GPD

[Les77a] Michael E. Lesk. A guide to preparing documents on UNIX. Technical Memorandum 1111, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1977. ?? pp.

Lesk:1977:POS

[Les77b] Michael E. Lesk. A proposal for operating system support. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1977. ?? pp.

Lesk:1977:TPF

[Les77c] Michael E. Lesk. Tbl — a program to format tables. Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 4, 1977. 20

                                        pp. URL http://www.kohala.
                                        com/start/troff/v7man/tbl/
                                        tbl.ps.

Lesk:1978:TDU

[Les78] Michael E. Lesk. Typing documents on the UNIX system: Using the -ms macros with troff and nroff. Technical Memorandum 1028, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 3, 1978. ?? pp.

Lesk:1979:SAI

[Les79a] Michael E. Lesk. Some applications of inverted indexes on the UNIX system. Technical Memorandum 1213, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. 12

                                        pp. URL http://www.kohala.
                                        com/start/troff/v7man/refer/
                                        refer.ps.

Lesk:1979:TPF

[Les79b] Michael E. Lesk. TBL — a program to format tables. Technical Memorandum 1106, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Lesk:1988:CUS

[Les88a] Michael Lesk. Can UNIX survive secret source code? In Association [Ass88a], pages 189– 199.

Lesk:1988:WMO

[Les88b] Michael Lesk. Word manipulation in online catalog searching: Using the UNIX system for library experiments. In USENIX Association [USE88c], pages 135–147. ISBN ???? LCCN ????

Levine:1979:FDS

[Lev79] John R. Levine. Forum: Data stream linkage and the UNIX system. IBM Systems Journal, 18(3):470–472, 1979. CODEN IBMSA7. ISSN 0018-8670. See [Mor78c, Mor79c].

Levy:1983:WTA

[Lev83] Leon S. Levy. A walk through AWK. ACM SIGPLAN Notices, 18(12):69–85, December 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levitt:1988:WIU

[Lev88] J. Levitt. Whither IBM and Unix? BYTE Magazine, 13(11):109–110,114, November 1988. CODEN BYTEDJ. ISSN 0360-5280.

Levine:1994:PGF

[Lev94] John Levine. Programming for Graphics Files in C and C++. Wiley-Interscience, New York, NY, USA, 1994. ISBN 0-471-59856-9. xii + 494 pp. LCCN T385 .L47 1994. US$49.95.

Levine:1996:UDC

[Lev96] John R. Levine. The UNIX dictionary of commands, terms, and acronyms. McGraw-Hill, New York, NY, USA, 1996. ISBN 0-07-037643-3 (hardcover), 0-07-037644-1 (paperback). x + 314 pp. LCCN QA76.76.O63 U545134 1996.

Levin:1998:RAS

[Lev98] Stewart A. Levin. Remark on Algorithm 622: a simple macroprocessor. ACM Transactions on Mathematical Software, 24(3):336–340, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
toms/1998-24-3/p336-levin/

. See [RRW84].

Levi:2002:UAC

[Lev02]  Bozidar Levi.  UNIX ad-

ministration: a comprehensive sourcebook for effective systems and network management. Internet and communications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2002. ISBN 0-8493-1351-1. 743 pp. LCCN QA76.76.O63 L4853 2002.

Levine:2004:Q

[Lev04]  John R. Levine.  Qmail.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 1-56592-628-5. xii + 234 pp. LCCN TK5105.74.Q43 L48 2004.

Levenez:2018:UH

[Lév18] Éric Lévénez. Unix history. Web site, August 11, 2018.

URL https://www.levenez.
com/unix/.

Lewis:1988:GEL

[Lew88] Bil Lewis. GNU Emacs Lisp manual: Emacs version 18 for UNIX users. ????, ????, March 1988. various pp.

Lewine:1991:PPG

[Lew91] Donald A. Lewine. POSIX programmer’s guide: writing portable UNIX programs with the POSIX.1 standard. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-73-0. xxvii + 607 pp. LCCN QA76.76.O63 L487 1991b. US$34.95. March 1994 printing with corrections, updates, and December 1991 Appendix G.

Lewis:1999:BCA

[Lew99a]  Ted Lewis.  Binary critic:

Asbestos pajamas: An open source dialogue. Computer, 32(4):112, 108–111, April 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/
co/books/co1999/pdf/r4112.

pdf. Continues the debate about open source software development, notably in GNU/ Linux [Lew99b].

Lewis:1999:BCO

[Lew99b]  Ted Lewis.  Binary critic:

The open source acid test. Computer, 32(2):128, 125– 127, February 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL http://dlib.computer.
                                        org/co/books/co1999/pdf/

r2128.pdf. See continuation in [Lew99a].

Lewis:2001:SWH

[Lew01] Nancy D. Lewis. StarOffice 5.2 Writer Handbook. Sun BluePrints Program. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-029386-5. xix + 410 pp. LCCN QA76.76.I57 W37 2001. US$29.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130293865.html;   http:
                                        //www.sun.com/books/catalog/
                                        warner/index.html.

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Liskov:1991:RUF

[LGJS91] Barbara Liskov, Robert Gruber, Paul Johnson, and Liuba Shrira. A replicated Unix file system (extended abstract). Operating Systems Review, 25 (1):60–64, January 1991. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Look:1984:REU

[LH84a] Byron Look and Gary Ho. Real-time extensions to the UNIX operating system. In USENIX Association [USE84b], pages 293–299. ISBN none. LCCN QA76.8.U65 U55 1984.

Look:1984:RTE

[LH84b] Byron Look and Gary Ho.

Real-time extensions to the UNIX operating system. In USENIX Association [USE84b], pages 293–299. ISBN none. LCCN QA76.8.U65 U55 1984.

Liberty:2000:STY

[LH00] Jesse Liberty and David B.

Horvath. Sams teach yourself C++ for Linux in 21 days. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31895-4. xxix + 1110 pp. LCCN QA76.73.C153 L526 2000.

Limoncelli:2001:PSN

[LH01a] Thomas A. Limoncelli and Christine J. Hogan. The Practice of System and Network Administration. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70271-1 (paperback). xxv + 774 pp. LCCN TK5105.5 .L53 2002.

LSI:2001:PLG

[LH01b] Loki Software, Inc. and John Hall. Programming Linux Games: Learn to Write the Games Linux People Play. Linux Journal Press, San Francisco, CA, USA, 2001. ISBN 1-886411-49-2. xviii + 415 pp. LCCN QA76.76.C672 L65 2001. US$39.95.

Lima:1983:DCF

[LHM^{+}83] Isabel Gouveia Lima, Richard Hopkins, Lindsay Marshall, David Mundy, and Philip Treleaven. Decentralised control flow — BASed on unIX BASIX. ACM SIGPLAN Notices, 18 (6):192–201, June 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Li:1998:DUG

[Li98] Yan Li. Developing a Unicode GIS application for Microsoft Windows and UNIX. In Unicode Consortium [Uni98a], page ?? ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc12/program.html.

Libes:1987:MPO

[Lib87]  Don Libes.  Multiple pro-

grams in one UNIX process. ;login: the USENIX Association newsletter, 12(4):7–13, July/August 1987. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Libes:1993:OCO

[Lib93] Don Libes. Obfuscated C and Other Mysteries. Wiley-Inter-science, New York, NY, USA, 1993. ISBN 0-471-57805-3. xii + 413 pp. LCCN QA76.73.C15 L5 1993. With practical C programming hints and winning programs from the Obfuscated C Code Contests … With a disk containing the source codes.

Libes:1994:EET

[Lib94] Don Libes. Exploring Expect:

A Tcl-based Toolkit for Automating Interactive Programs. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, December 1994. ISBN 1-56592-090-2. xxxiii + 566 pp. LCCN QA76.755 .L52 1995. US$29.95. URL http:/

/www.ora.com/gnn/bus/ora/
item/expect.html;   http:
//www.oreilly.com/catalog/
9781565920903;   http://
www.oreilly.com/catalog/
expect.

Licwinko:1979:BFM

[Lic79a] J. S. Licwinko. BANCS facility for MERT/UNIX. Technical Memorandum 1442 (TM 79-9156-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 11, 1979. ?? pp.

Licwinko:1979:RHU

[Lic79b] J. S. Licwinko. Record handling under MERT/UNIX — the queue facility. Technical Memorandum 1443 (TM 79-9156-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 25, 1979. ?? pp.

Limoncelli:2005:ATM

[Lim05] Tom Limoncelli. The Art of Time Management for System Administrators. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00783-3 (paperback), 0-596-51950-8 (e-book), 0-596-55303-X (ebook). xxiii + 200 pp. LCCN HD69.T54 L546 2006.

Lin:1965:CST

[Lin65] Shen Lin. Computer solutions of the traveling salesman problem. The Bell System Technical Journal, 44(10): 2245–2269, December 1965. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol44/bstj44-10-2245.
pdf;   http://www.alcatel-
lucent.com/bstj/vol44-1965/
articles/bstj44-10-2245.
pdf.

Linderman:1979:DSU

[Lin79a] J. P. Linderman. DIRECT 11 sort utilities. Technical Memorandum 1454 (TM 79-9456-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 31, 1979. ?? pp.

Linderman:1979:DID

[Lin79b] J. P. Linderman. DIRECT II data representation and manipulation. Technical Memorandum 1455 (TM 79-9456-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 8, 1979. ?? pp.

Linderman:1979:DII

[Lin79c] J. P. Linderman. DIRECT II interprocess communication. Technical Memorandum 1368, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 30, 1979. ?? pp.

Lindberg:1984:LIU

[Lin84] Jeff Lindberg. A layered implementation of the UNIX kernel on the HP9000 series 500 computers. In USENIX Association [USE84b], pages 183–194. ISBN none. LCCN QA76.8.U65 U55 1984.

Linnell:1993:WNC

[Lin93] Dennis Linnell. Windows NT:

Can Microsoft make the jump from the desktop to distributed computing? Data communications, 22(6):68–??, April 1993. CODEN DACODM. ISSN 0363-6399.

Young:1994:LJ

[Lin94] Robert F. Young, editor. Linux journal, 1994. ISSN 1075-3583 (print), 1938-3827 (electronic). Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155.

Linthicum:1995:WUB

[Lin95] David S. Linthicum. What UNIX Branding Means to You. Datamation, 41(13):53– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Lingmann:2002:DSK

[Lin02a] Thomas Lingmann. Datenverschlüsselung: sichere Kommunikation mit Linux und BSD: Security mit Open Source. (German) [Data encoding: Secure communication with Linux and BSD: Security with Open Source]. C & L, Böblingen, Germany, 2002. ISBN 3-932311-87-8 (??invalid checksum??). 476 (est.) pp. LCCN ????

Anonymous:2002:LSU

[Lin02b] Anonymous, editor. Die Linuxzeitung: das unabhängige Monatsblatt für GNU/Linux, freie Software und Open Source. (German) [The Linux Times: the independent monthly for GNU/Linux, Free Software, and Open Source], 2002. ISSN ????

Anonymous:2002:LUM

[Lin02c] Anonymous, editor. Die Linuxzeitung: das unabhängige Monatsblatt für GNU/Linux, freie Software und Open Source. (German) [The Linux Times: the independent monthly for GNU/Linux, Free Software, and Open Source], 2002. URL

http://www.die.linuxzeitung.

de/. ????, Berlin, Germany. On-line journal.

Lions:1978:SSF

[Lio78] J. Lions. Shell subprogram facility. Technical Memorandum 1327, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 8, 1978. ?? pp.

Lions:1979:CFA

[Lio79a] John Lions. Checking file access permissions in UNIX systems. Technical Memorandum 1385, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 28, 1979. ?? pp.

Lions:1979:EUTa

[Lio79b] John Lions. Experiences with the UNIX time-sharing system. Technical Memorandum 1477, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1979. ?? pp.

Lions:1979:EUTb

[Lio79c] John Lions. Experiences with the UNIX time-sharing system. Software—Practice and Experience, 9(9):701–709, September 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Lions:1979:MAC

[Lio79d] John Lions. Macros for analyzing C program arguments. Technical Memorandum 1384, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 28, 1979. ?? pp.

Lions:1979:SPP

[Lio79e] John Lions. The second pass of the Portable C Compiler. Technical Memorandum 1430, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1979. ?? pp.

Lions:1996:LCU

[Lio96]  John Lions.  Lions’ Com-

mentary on UNIX 6th Edition, with Source Code. Computer classics revisited. Peerto-Peer Communications, San Jose, CA 95164-0218, USA, 1996. ISBN 1-57398-013-7. 254 pp. LCCN QA76.76.O63 L562 1996. US$29.96; CAN$41.95.

URL http://wiki.tuhs.org/
doku.php?id=publications:
lions_commentary;   http:
//www.lemis.com/grog/Documentation/
Lions/;   http://www.peer-
to-peer.com/catalog/opsrc/
lions.html;   https://github.
com/kanner/lions-book. With

forewords by Dennis M. Ritchie and Ken Thompson. Prefatory notes by Peter H. Salus and Michael Tilson; a Historical Note by Peter H. Salus; and Appreciations by Greg Rose, Mike O’Dell, Berny Goodheart, Peter Collinson, and Peter Reintjes. Originally circulated as two restricted-release volumes: “UNIX Operating System Source Code Level Six”, and “A Commentary on the UNIX Operating System”.

Lippman:1991:CP

[Lip91] Stanley B. Lippman. C++ Primer. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-54848-8 (paperback). xvi + 614 pp. LCCN QA76.73.C15 L57 1991.

Lipkin:1999:LLV

[Lip99] Bernice Sacks Lipkin. L^{A}TEX for Linux: a Vade Mecum. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98708-8 (paperback). xxxi + 568 pp. LCCN Z253.4 L38 L56 1999. US$49.95.

Lischner:2002:BCC

[Lis02]  Ray Lischner.  Borland’s

CLX component framework. Dr. Dobb’s Journal of Software Tools, 27(2):77, 79–81, February 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_02/clx.txt.

Litwin:1980:LHN

[Lit80] W. Litwin. Linear hashing: a new tool for file and table addressing. In Lochovsky and Taylor [LT80], pages 212–223 (or 260–276??). Reprinted in [Sto88, p. 570–581].

Litman:1988:DDO

[Lit88] Ami Litman. The DUNIX distributed operating system. Operating Systems Review, 22(1): 42–51, January 1988. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Liu:1996:LEC

[Liu96] Huiqi Liu. Low-cost email connection between UNIX and Macs. Sys Admin: The Journal for UNIX Systems Administrators, 5(6):37–??, June 1996. CODEN SYADE7. ISSN 1061-2688.

Liu:2011:UOS

[Liu11] Yukun Liu. Unix operating system: the development tutorial via Unix Kernel Services. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-20431-7. ???? pp. LCCN ????

Lai:1984:MMU

[LJ84] Clara S. Lai and Chris Peer Johnson. Memory management units and the UNIX kernel. In Software Tools Users Group [Sof84], pages 208–213. LCCN QA76.8.U65 U83 1984.

Lucas:2015:FM

[LJ15] Michael W. Lucas and Allan Jude. FreeBSD Mastery: ZFS. Tilted Windmill Press, ????, 2015. ISBN 0-692-45235-4 (paperback). xx + 212 pp. LCCN ????

Lin:1973:EHAa

[LK73a] S. Lin and B. W. Kernighan. An effective heuristic algorithm for the travelingsalesman problem. Computing Science Technical Report 1, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973. ?? pp.

Lin:1973:EHAb

[LK73b] S. Lin and B. W. Kernighan. An effective heuristic algorithm for the travelingsalesman problem. Operations Research, 21(2):498–516, April 1973. CODEN OPREAI. ISSN 0030-364X (print), 1526-5463 (electronic). URL https:

                                        //www.cs.princeton.edu/~
                                        bwk/btl.mirror/tsp.pdf.

Lesk:1976:CTTa

[LK76a] Michael E. Lesk and Brian W. Kernighan. Computer typesetting of technical journals on UNIX. Technical Memorandum 1089 (TM 76-1274-2, TM 76-1273-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 29, 1976. ?? pp.

Lesk:1976:CTTb

[LK76b] Michael E. Lesk and Brian W.

Kernighan. Computer typesetting of technical journals on Unix. Computing Science Technical Report 44, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1976. ?? pp.

Lesk:1977:CTT

[LK77] Michael E. Lesk and Brian W.

Kernighan. Computer typesetting of technical journals on Unix. In Robert R. Korfhage and Portia Isaacson, editors, AFIPS conference proceedings 1977: National computer conference, June 13–16, 1977 Dallas, Texas, volume 46, pages 879–888. AFIPS Press, Montvale, NJ, USA, 1977.

URL https://dl.acm.org/
doi/abs/10.1145/1499402.
1499562.

Langer:2000:SCI

[LK00] Angelika Langer and Klaus Kreft. Standard C++ IOStreams and locales: advanced programmer’s guide and reference. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-18395-1. xxvi + 640 pp. LCCN QA76.73.C153 L37 2000. URL http:

//www.angelikalanger.com/
IOStreams/errata1st.htm.

Leffler:1984:MIP

[LKM84] Sam Leffler, Michale Karels, and Marshall Kirk McKusick. Measuring and improving the performance of 4.2BSD. In USENIX [USE84a], pages 237– 252.

Lakshman:1993:DIM

[LKY93] K. Lakshman, G. Kuthethoor, and R. Yavatkar. Design and implementation of a multimedia protocol suite in a BSD UNIX kernel. In USENIX [USE93c], pages 129– 146. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/cinci93/.

Lomuto:1983:UP

[LL83a] Ann Nicols Lomuto and Nico Lomuto. A UNIX Primer. Prentice-Hall Software Series, Editor: Brian W. Kernighan. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-938886-9. xvi + 239 pp. LCCN QA76.8.U65 L65 1983.

Lomuto:UP83

[LL83b] Ann Nicols Lomuto and Nico Lomuto. A UNIX Primer. Prentice-Hall Software Series, Editor: Brian W. Kernighan. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1983. ISBN 0-13-938886-9. xvi + 239 pp. LCCN QA76.8.U65 L65 1983.

Leathrum:1993:ATO

[LL93a] J. F. Leathrum and K. A.

Liburdy. Automated testing of open software standards. In ITC’93 [ITC93], pages 854–858. ISBN 0-7803-1429-8. LCCN TK7874.I474 1993. IEEE catalog number 93CH3356-3.

Leathrum:1993:ERT

[LL93b] J. F. Leathrum and K. A.

Liburdy. Evolving role of testing in open systems standards. In ITC’93 [ITC93], pages 273– 274. ISBN 0-7803-1429-8. IEEE catalog number 93CH3356-3.

Leathrum:1995:FTS

[LL95] J. F. Leathrum and K. A.

Liburdy. Formal test specifications in IEEE POSIX. Computer Standards and Interfaces, 17(5-6):603–614, September 1995. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Lee:1987:SPP

[LLM87] T. P. Lee, M. W. Luppi, and R. E. Menninger. Solving performance problems on a multiprocessor UNIX system. In USENIX Association [USE87d], pages 399–405.

Langue:1988:PUL

[LM88a] Y. Langue and T. Muntean.

PARX: a UNIX-like operating system for transputer-based parallel supercomputers. In USENIX Association [USE88f], pages 109–120. ISBN ???? LCCN ????

Langue:1988:PUO

[LM88b] Y. Langue and T. Muntean. PARX: A UNIX-like operating system for Transputer-based parallel supercomputers. In USENIX Association [USE88f], pages 109–120. ISBN ???? LCCN ????

Leffler:1991:DIU

[LM91] Samuel J. Leffler and Marshall Kirk McKusick. The design and implementation of the 4.3BSD UNIX operating system answer book. Addison-Wesley series in computer science. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-54629-9. ix + 85 pp. LCCN QA76.76.O63 L445 1991. URL http:

//wiki.tuhs.org/doku.php?
id=publications:4.3bsd_design.

Levine:1992:LY

[LMB92] John R. Levine, Tony Mason, and Doug Brown. lex & yacc. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1992. ISBN 1-56592-000-7. xxii + 366 pp. LCCN QA76.76.U84M37 1992. US$29.95. URL http:/

/www.oreilly.com/catalog/
9781565920002;   http://
www.oreilly.com/catalog/
lex.

Leffler:1989:DIU

[LMKQ89] Samuel J. Leffler, Marshall Kirk McKusick, Michael J. Karels, and John S. Quarterman. The Design and Implementation of the 4.3BSD UNIX Operating System. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-06196-1. xxii + 471 pp. LCCN QA76.76.O63 D4741 1989.

Leffler:1990:BUB

[LMKQ90] Samuel J. Leffler, Marshall Kirk McKusick, Michael J. Karels, and John S. Quarterman. Das 4.3-BSD-Unix-Betriebssystem: Design und Implementierung. (German) [The 4.3 BSD UNIX Operating System: Design and Implementation]. Addison-Wesley, Reading, MA, USA, 1990. ISBN 3-89319-239-5. xvi + 678 pp. LCCN ????

Li:2001:LLF

[LMOT01] Ren-Cang Li, Peter Markstein, Jon P. Okada, and James W. Thomas. The libm library and floating-point arithmetic for HP-UX on Itanium. Technical report, Hewlett– Packard Corporation, Palo Alto, CA, USA, April 2001.

?? pp. URL http://h21007.
www2.hp.com/dspp/ddl/ddl_
Download_File_TRX/1,1249,
942,00.pdf;   http://h21007.
www2.hp.com/dspp/tech/tech_
TechDocumentDetailPage_IDX/
1,1701,981,00.html.

Li:2002:LLF

[LMOT02] Ren-Cang Li, Peter Markstein, Jon P. Okada, and James W. Thomas. The libm library and floating-point arithmetic for HP-UX on Itanium-2. Technical report, Hewlett–Packard Corporation, Palo Alto, CA, USA, 2002. ?? pp. URL ???

                                        ?.

Luderer:1978:UTS

[LMT78] G. W. R. Luderer, J. F. Maranzano, and B. A. Tague. UNIX time-sharing system: The UNIX operating system as a base for applications. The Bell System Technical Journal, 57(6):2201–2207, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2201.
                                        pdf.

Lo:1979:MSP

[Lo79] D. S. Lo. A mechanism of starting a program in a UNIX based operating system. Technical Memorandum 1488, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 24, 1979. ?? pp.

Loukides:1997:PGS

[LO97] Mike Loukides and Andy Oram. Programming with GNU Software. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-112-7. xiv + 244 pp. LCCN QA76.76.O63L65 1997. US$39.95. URL http://www.

                                        ora.com/catalog/prognu/;
                                         http://www.oreilly.com/
                                        catalog/prognu. Includes CD-

ROM.

Lobelle:1985:IDW

[Lob85] Marc C. Lobelle. Integration of diskless workstations in UNIX United. Software—Practice and Experience, 15(10): 997–1010, October 1985. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Loikits:1977:MHC

[Loi77] E. A. Loikits. MERT hardware configuration. Technical Memorandum 1161 (MF 77-8234-037), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 20, 1977. ?? pp.

Loikits:1979:ILA

[Loi79]  E. Loikits.  Implementa-

tion of the large argument list feature in UNIX/RT’s exec. system call. Technical Memorandum 1351 (2524-790110.01MF), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 10, 1979. ?? pp.

Lombardo:2001:EL

[Lom01] John Lombardo. Embedded Linux. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-0998-X. xxvii + 192 pp. LCCN QA76.76.O63 L646 2001. US$39.99.

Long:1977:SHD

[Lon77] P. F. Long. Synthesis of a half-duplex line control procedure for UNIX using state transition matrix techniques. Technical Memorandum 1173 (TM 77-8234-14), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 19, 1977. ?? pp.

Long:1978:BTC

[Lon78] P. F. Long. Beta test of the COMM IOP-DUP synchronous communications controller under …. Technical Memorandum 1240 (TM 78-8234-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 23, 1978. ?? pp.

Lonsford:1990:USF

[Lon90] E. H. Lonsford. UNIX security: fact not fiction. Datamation, 36(4):44–46, 48, February 1990. CODEN DTMNAT. ISSN 0011-6963.

Loukides:1990:SPT

[Lou90a] Mike Loukides. System Performance Tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-60-9. xix + 313 pp. LCCN QA76.76.O63 L66 1992. US$24.95. A fine book for system administrators on how to fine tune your UNIX system(s) to do more work.

Loukides:1990:UFP

[Lou90b] Mike Loukides. UNIX for FORTRAN Programmers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-51-X. xviii + 244 pp. LCCN QA76.76.O63 L67 1990. US$24.95.

Love:2005:LKD

[Lov05] Robert Love. Linux kernel development. Novell Press, Indianapolis, IN, USA, second edition, 2005. ISBN 0-672-32720-1 (paperback). xvi + 401 pp. LCCN QA76.76.O63 L673 2005.

Love:2007:LSP

[Lov07] Robert Love. Linux system programming: system and library calls every programmer needs to know. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-00958-5 (paperback). xvii + 368 pp. LCCN QA76.76.O63 L6735 2007; QA76.76.O63 L69 2007.

URL http://www.oreilly.
com/catalog/9780596009588.

Love:2010:LKD

[Lov10]  Robert Love.  Linux ker-

nel development: a thorough guide to the design and implementation of the Linux kernel. Developer’s library: essential references for programming professionals. Addison-Wesley, Reading, MA, USA, third edition, 2010. ISBN 0-672-32946-8 (paperback), 0-7686-9697-6 (e-book). xx + 440 pp. LCCN NLS PB8.210.508/14; QA76.76.O63 L674 2010.

Loza:2005:UAD

[Loz05a] Boris Loza. Under attack:

Dealing with missing UNIX files. ;login: the USENIX Association newsletter, 30 (3):??, June 2005. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

                                        tronic). URL http://www.
                                        usenix.org/publications/
                                        login/2005-06/pdfs/loza0506.
                                        pdf.

Loza:2005:USL

[Loz05b] Boris Loza. UNIX, Solaris and Linux: a Practical Security Cookbook: Securing UNIX Operating System without Third-Party Applications. Author-House, ????, 2005. ISBN 1-4208-4824-0. 368 (est.) pp. LCCN ????

Lanzillo:1989:IDI

[LP89] Leo Lanzillo and Craig Partridge. Implementation of dialup IP for UNIX systems. In USENIX Association [USE89f], pages 201–207.

Lamb:1995:UEE

[LP95] Linda Lamb and Jerry Peek. Using Email Effectively. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1995. ISBN 1-56592-103-8. 160 pp. LCCN TK5105.73.L36 1995. US$14.95.

Leightley:2018:PIO

[LPG^{+}18] Daniel Leightley, Jo-Anne Puddephatt, Laura Goodwin, Roberto Rona, and Nicola T. Fear. InDEx: Open source iOS and Android software for self-reporting and monitoring of alcohol consumption. Journal of Open Research Software, 6(1):13–??, March 23, 2018. CODEN ???? ISSN 2049-9647. URL https://

openresearchsoftware.metajnl.
com/articles/10.5334/jors.
207/.

LoVerso:1991:OUF

[LPLF91] Susan LoVerso, Noemi Paciorek, Alan Langerman, and George Feinberg. The OSF/1 UNIX filesystem (UFS). In USENIX [USE91d], pages 207– 218. LCCN QA 76.76 O63 U84 1992.

Lyon:1977:IUPb

[LR77] Thomas L. Lyon and Dennis M.

Ritchie. Inter-UNIX portability. Technical Memorandum 1203 (TM 77-1273-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 16, 1977. ?? pp.

London:1978:UOS

[LR78] T. B. London and J. F. Reiser.

A UNIX operating system for the DEC VAX-11/780 computer. Technical Memorandum 1297 (TM 78-1353-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 7, 1978. ?? pp.

Libes:1989:LU

[LR89a] Don Libes and Sandy Ressler.

Life with UNIX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-536657-7. xx + 346 pp. LCCN QA76.76.O63 L52 1989. URL http:/

                                        /minnie.tuhs.org/Archive/
                                        Documentation/Books/.

Libes:LU89

[LR89b] Don Libes and Sandy Ressler. Life with UNIX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-536657-7. xx + 346 pp. LCCN QA76.76.O63 L52 1989. URL http:/

                                        /minnie.tuhs.org/Archive/
                                        Documentation/Books/.

Luby:1989:SPS

[LR89c] Michael Luby and Charles Rackoff. A study of password security. Journal of Cryptology: the journal of the International Association for Cryptologic Research, 1(3):151–158, ???? 1989. CODEN JOCREQ. ISSN 0933-2790 (print), 1432-1378 (electronic).

Lanzatella:1990:SMI

[LR90] Thomas W. Lanzatella and Paul G. Rutherford. Storage management issues for Cray Research. Digest of Papers — IEEE Symposium on Mass Storage Systems, pages 176– 181, 1990. CODEN DPISDX. ISBN 0-8186-2034-x. ISSN 1051-9173. IEEE catalog number 90CH2844-9.

Ladd:1995:LIL

[LR95] D. A. Ladd and J. C. Ramming. A*: a language for implementing language processors. IEEE Transactions on Software Engineering, 21(11):

894–901, November 1995. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=473218.

Li:2019:RIC

[LRB^{+}19] Li Li, Timothée Riom, Tegawendé F.

Bissyandé, Haoyu Wang, Jacques Klein, and Le Traon Yves. Revisiting the impact of common libraries for Androidrelated investigations. The Journal of Systems and Software, 154(??):157–175, August 2019. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0164121219301001.

Lesk:1979:LLAa

[LS79a] Michael E. Lesk and E. Schmidt.

Lex — a lexical analyzer generator. Technical Memorandum 1061, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. 13?? pp.

Lesk:1979:LLAb

[LS79b] Michael E. Lesk and Eric Schmidt. Lex — a lexical analyzer generator. In UNIX Programmer’s Manual, volume 2, pages 388–400. Holt, Rinehart, and Winston, New York, NY, USA, 1979. AT&T Bell Laboratories Technical Report in 1975.

Lutz:1983:RUK

[LS83] Michael Lutz and Michael Shon. Running the UNIX kernel in user mode. In Software Tools Users Group [Sof83], pages 171–176. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Loucks:1987:AIE

[LS87] L. K. Loucks and C. H. Sauer. Advanced Interactive Executive (AIX) operating system overview. IBM Systems Journal, 26(4):326–345, December 1987. CODEN IBMSA7. ISSN 0018-8670. URL http:/

/technologists.com/sauer/
Advanced%20Interactive%20Executive%
20(AIX)%20Operating%20System%
20Overview.pdf.

Lesk:1990:LLA

[LS90a] M. E. Lesk and E. Schmidt. Lex — a lexical analyzer generator. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Levy:1990:DFS

[LS90b] Eliezer Levy and Abraham Silberschatz. Distributed file systems: Concepts and examples. ACM Computing Surveys, 22(4):321–374, December 1990. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0360-
0300/98169.html.

Litzkow:1992:SCP

[LS92] Michael Litzkow and Marvin Solomon. Supporting checkpointing and process migration outside the UNIX kernel. In USENIX [USE92d], pages 283– 290.

Luppi:1988:BET

[LSC^{+}88] Mark Luppi, Mark Seiden, Joseph Collins, Daniel Fisher, Keith Iverson, Charles Marshall, Josef Sachs, and David Shaw. Building an equities trading system in a distributed UNIX environment. In USENIX Association [USE88g], pages 97–104. ISBN ???? LCCN ????

Lencoski:1976:UIM

[LT76] Joseph E. Lencoski and Stuart A. Tartarone. UNIX– INTELLEC MDS interface. Technical Memorandum 1118 (TM 76-3141-1, TM 76-3142-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1, 1976. ?? pp.

Lochovsky:1980:SIC

[LT80] Frederick H. Lochovsky and ?. Taylor, editors. Sixth International Conference on Very Large Data Bases: reprinted from Very large data bases. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1980. ACM order no. 471800. IEEE catalog no. 80CH1534-7C. Long Beach order no. 322.

Leong:1991:UPE

[LT91] Philip Leong and Chris Tham. UNIX password encryption considered unsecure. In USENIX [USE91d], pages 269– 280. LCCN QA 76.76 O63 U84 1992.

Lencoski:1977:CCI

[LTW77] Joseph E. Lencoski, Stuart A. Tartarone, and J. E. Wilkes, Jr. A console controlled interface between an INTELLEC MDS and a time-shared computer system. Technical Memorandum 1168 (TM 77-3141-2, TM 77-3142-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 31, 1977. ?? pp.

Lubkin:1990:PSP

[Lub90] Saul Lubkin. Porting sophisticated programs to your UNIX environment for free. Computer Shopper, pages 642–644, March 1990.

Lucas:2000:LUF

[Luc00] Michael Lucas. Linux under FreeBSD. Sys Admin: The Journal for UNIX Systems Administrators, 9(1):26, 29–30, 32, January 2000. CODEN SYADE7. ISSN 1061-2688.

                                        URL http://www.samag.com/

.

Lucas:2002:ABU

[Luc02]  Michael Lucas.  Absolute

BSD: The Ultimate Guide to FreeBSD. No Starch Press, San Francisco, CA, USA, July 2002. ISBN 1-886411-74-3. xlii + 565 pp. LCCN QA76.76.O63 L83 2002. US$39.95. Foreword by Jordan Hubbard.

Lucas:2003:AOU

[Luc03a] Michael W. Lucas. Absolute OpenBSD: Unix for the practical paranoid. No Starch Press, San Francisco, CA, USA, 2003. ISBN 1-886411-99-9. xxviii + 489 pp. LCCN QA76.76.O63 L835 2003. US$39.95.

Lucas:2003:FLU

[Luc03b] Michael W. Lucas. FreeBSD de Luxe: UNIX-Serveradministration; Installation, Upgrading, Systemoptimierung; Mail, DNS, Web, FTP: die Konfiguration von Internetdiensten; aktuelle-Informationen zur Version 5 . mitp Verlag, Bonn, Germany, 2003. ISBN 3-8266-1343-0. 657 pp. LCCN ????

Lucke:2005:BCL

[Luc05]  Robert W. Lucke.  Build-

ing clustered Linux systems. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-144853-6. xxxviii + 606 pp. LCCN QA76.76.O63 L838 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;

                                         http://www.loc.gov/catdir/
                                        toc/ecip0418/2004014016.
                                        html.

Lucas:2008:AFC

[Luc08]  Michael Lucas.  Absolute

FreeBSD: the complete guide to FreeBSD. No Starch Press, San Francisco, CA, USA, second edition, 2008. ISBN 1-59327-151-4 (paperback). xxx + 709 pp. LCCN QA76.76.O63 L83 2008.

Lucas:2012:SMO

[Luc12] Michael Lucas. SSH Mastery: OpenSSH, PuTTY, tunnels and keys. Tilted Windmill Press, ????, 2012. ISBN 1-4700-6971-7. ix + 128 pp. LCCN ????

Lucas:2013:AOU

[Luc13a]  Michael Lucas.  Absolute

OpenBSD: UNIX for the practical paranoid. No Starch Press, San Francisco, CA, USA, second edition, 2013. ISBN 1-59327-476-9 (paperback). xliii + 491 pp. LCCN QA76.76.O63 L833 2013.

Lucas:2013:SMU

[Luc13b] Michael W. (Michael Warren) Lucas. Sudo mastery: user access control for real people. Tilted Windmill Press, ????, 2013. ISBN 1-4936-2620-5. vii + 120 pp. LCCN TK5105.59 .S83 2013.

Lucas:2016:PM

[Luc16] Michael Lucas. PAM mastery. Tilted Windmill Press, ????, 2016. ISBN 1-5376-5770-4. 182 pp. LCCN ????

Lucas:2018:AF

[Luc18a] Michael W. Lucas. Absolute FreeBSD. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2018. ISBN 1-59327-892-6. LCCN ????

Lucas:2018:SMO

[Luc18b] Michael W. Lucas. SSH Mastery: OpenSSH, PuTTY, Tunnels and Keys, volume 12 of IT Mastery. Tilted Windmill Press, ????, second edition, 2018. ISBN 1-64235-022-2 (hardcover), 1-64235-002-8 (paperback). 242 pp. LCCN ????

Lucas:2019:FMJ

[Luc19] Michael W. Lucas. FreeBSD Mastery: Jails. Tilted Windmill Press, ????, 2019. ISBN 1-64235-023-0. ???? pp. LCCN ????

Luderer:1977:CPT

[Lud77a] G. W. R. Luderer. Cut and paste — two UNIX word processing commands. Technical Memorandum 1205 (MF 77-8234-086), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 30, 1977. ?? pp.

Luderer:1977:MRD

[Lud77b] G. W. R. Luderer. MERT Release 0 from Department 8234. Technical Memorandum 1183 (MF 77-8234-067), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 30, 1977. ?? pp.

Luderer:1978:CUR

[Lud78a] G. W. R. Luderer. Contents of UNIX/RT. Technical Memorandum 1231 (MF 78-8234-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 10, 1978. ?? pp.

Luderer:1978:SFT

[Lud78b] G. W. R. Luderer. Standardizing file types in the inodes of UNIX Version 7 descendants. Technical Memorandum 1256 (MF 78-8234-48), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 28, 1978. ?? pp.

Luderer:1979:DWN

[Lud79a] G. W. R. Luderer. Department 3624’s workplan for networking in 1979/80. Technical Memorandum 1435, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 16, 1979. ?? pp.

Luderer:1979:PCE

[Lud79b] G. W. R. Luderer. Preliminary considerations for exploratory UNIX operating system work. Technical Memorandum 1391, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 11, 1979. ?? pp.

Lumme:2013:BHA

[Lum13]  Juha Lumme.  BeagleBone

home automation live your sophisticated dream with home automation using BeagleBone. Community experience distilled. Packt Publishing, Birmingham, UK, 2013. ISBN 1-78328-573-7, 1-306-27005-7 (ebook), 1-78328-574-5 (e-book). LCCN ???? URL http://

proquest.tech.safaribooksonline.
de/9781783285730.

Lunde:1993:UJI

[Lun93] Ken Lunde. Understanding Japanese Information Processing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-043-0. xxxii + 435 pp. US$29.95.

Lund:1996:IUP

[Lun96]  William Lund.  Integrating

UNIX and PC network operating systems: NetWare, AppleTalk, and LAN Manager on UNIX. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-207374-9. xiv + 172 pp. LCCN QA76.76.O63 L86 1996.

Lunde:1999:CIP

[Lun99] Ken Lunde. CJKV Information Processing: Chinese, Japanese, Korean & Vietnamese Computing. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-224-7. 1174 pp. LCCN PL1074.5 .L85 1999. US$64.95. URL http:/

/www.oreilly.com/catalog/
cjkvinfo/.

Lutz:1999:NBF

[Lut99] Michael J. Lutz. New books:

Folding all things virtual into urban life; performance modeling in network design; practical automated software testing; Java patterns; testing software step by step; working with UNIX shells. Computer, 32(11): 119, November 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/

ry119.pdf. Includes review of [Qui99].

Luu:1990:CCR

[Luu90] J. Luu. Comments on ‘A comparison of RISC architectures’ by R. S. Piepho and W. S. Wu. IEEE Micro, 10(2):5, March/ April 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2003:OSW

[LW03] James Lee and Brent Ware. Open source Web development with LAMP: using Linux, Apache, MySQL, Perl, and PHP. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-77061-X (paperback). xxxiv + 460 pp. LCCN QA76.76.D47 L435 2003.

Liu:2005:PAL

[LWL05] Chun-Ho Liu, Chat-Ming Woo, and Dennis Y. C. Leung. Performance analysis of a Linux PC cluster using a direct numerical simulation of fluid turbulence code. The International Journal of High Performance Computing Applications, 19(4):365–374, Winter 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

sagepub.com/content/19/4/
365.full.pdf+html.

Liu:2006:PAP

[LWL06] Chun-Ho Liu, Chat-Ming Woo, and Dennis Y. C. Leung. Performance analysis of a parallel finite element solution to the direct numerical simulation of fluid turbulence on Linux PC clusters. Applied Mathematics and Computation, 172(2):731– 743, January 15, 2006. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Levine:1993:UD

[LY93] John Levine and Margaret Levine Young. UNIX for Dummies. IDG Books, San Mateo, CA, USA, 1993. ISBN 1-878058-58-4. xxvi + 369 pp. LCCN QA76.76.O63 L486 1993. US$19.95. A fairly informal (funny) and non-technical introduction to UNIX.

Levine:1997:UD

[LY97] John R. Levine and Margaret Levine Young. UNIX for dummies. I D G Books Worldwide, Indianapolis, IN, USA, third edition, 1997. ISBN 0-7645-0130-5. xxvi + 376 pp. LCCN QA76.76.O63 L486

Levine:1998:UD

[LY98] John Levine and Margaret Levine Young. UNIX for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, fourth edition, 1998. ISBN 0-7645-0419-

3. xxviii + 376 pp. LCCN xxviii

+ 376. US$21.99.

Levine:2004:UD

[LY04] John Levine and Margaret Levine Young. UNIX for Dummies. Wiley, New York, NY, USA, fifth edition, 2004. ISBN 0-7645-4147-1. xxii + 380 pp. LCCN QA76.76.O63 L486 2004. US$21.99.

Lycklama:1971:PSM

[Lyc71] Heinz Lycklama. Performance simulation and measurement of a virtual memory multiprogramming system for a small computer. Memorandum MM71-1383-3, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 20, 1971. 67 pp.

                                        URL https://www.tuhs.org/
                                        Archive/Documentation/TechReports/
                                        Heinz_Tech_Memos/MM-71-1383-
                                        3_Performance_Simulation_
                                        and_Measurement_of_a_Virtual_
                                        Memory_Multi-progamming_
                                        System_for_a_Small_Computer_
                                        19710120.pdf.

Lycklama:1972:BIS

[Lyc72] Heinz Lycklama. Bus interference in a single bus multiprocessor environment. Report MM72-1353-16, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 20, 1972. 61

pp. URL https://www.tuhs.
org/Archive/Documentation/
TechReports/Heinz_Tech_Memos/
MM-72-1353-16_Bus_Interface_
in_a_Single_Bus_Multi-processor_
Environment_19720920.pdf.

Lycklama:1973:VMM

[Lyc73] Heinz Lycklama. A virtual memory mini-computer system. Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1973. 2 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/A_Virtual_
Memory_Mini-Computer_System_
516-TSS.pdf.

Lycklama:1974:ILC

[Lyc74a] Heinz Lycklama. Implementation of large contiguous files and asynchronous I/0 in UNIX. Technical Memorandum 1003 (TM 74-1352-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 4, 1974. 16 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-74-1352-
1_Implementstion_of_Large_
Contiguous_Files_and_Asynchronous_
IO_in_UNIX_19740104.pdf.

Lycklama:1974:PFM

[Lyc74b] Heinz Lycklama. Plotting facilities for mini-computer systems. Technical Memorandum TM-74-1352-7, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 14, 1974. 32 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/

                                        Heinz_Tech_Memos/TM-74-1352-
                                        7_Plotting_Facilities_for_
                                        Mini-Computer_Systems_19740614.
                                        pdf.

Lycklama:1975:GTU

[Lyc75] Heinz Lycklama. GLANCE terminals on UNIX time-sharing. Technical Memorandum TM-75-1352-3, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 3, 1975. 30 pp.

                                        URL https://www.tuhs.org/
                                        Archive/Documentation/TechReports/
                                        Heinz_Tech_Memos/TM-75-1352-
                                        3_GLANCE_Terminals_on_UNIX_
                                        Time-Sharing_19750303.pdf.

Lycklama:1976:LUS

[Lyc76a] Heinz Lycklama. LSI-UNIX system. Technical Memorandum 1116 (TM 76-1352-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 6, 1976. 10 pp.

Lycklama:1976:MUR

[Lyc76b] Heinz Lycklama. MINI-UNIX regenerating system software. Technical Memorandum 1125, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1976. ?? pp.

Lycklama:1976:MUS

[Lyc76c] Heinz Lycklama. MINI-UNIX summary. Technical Memorandum 1123, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1976. ?? pp.

Lycklama:1976:SMU

[Lyc76d] Heinz Lycklama. Setting-up MINI-UNIX. Technical Memorandum 1124, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1976. ?? pp.

Lycklama:1977:LUS

[Lyc77a] Heinz Lycklama. LSI-UNIX summary. Technical Memorandum 1158, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1977. ?? pp.

Lycklama:1977:LS

[Lyc77b] Heinz Lycklama. LSX software.

Technical Memorandum 1129, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1977. ?? pp.

Lycklama:1977:MUS

[Lyc77c] Heinz Lycklama. The MINI-UNIX system. Technical Memorandum 1135 (TM 77-1352-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 13, 1977. 10 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-77-1352-
1_The_MINI-UNIX_19770103.
pdf.

Lycklama:1977:RLS

[Lyc77d] Heinz Lycklama. Regenerating LSX software. Technical Memorandum 1131, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1977. ?? pp.

Lycklama:1977:RMS

[Lyc77e] Heinz Lycklama. Regenerating MERT software. Technical Memorandum 1147, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1977. ?? pp.

Lycklama:1977:SLU

[Lyc77f] Heinz Lycklama. Setting up LSI-UNIX. Technical Memorandum 1130, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1977. ?? pp.

Lycklama:1977:SM

[Lyc77g] Heinz Lycklama. Setting-up MERT. Technical Memorandum 1146, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1977. ?? pp.

Lycklama:1977:UMP

[Lyc77h] Heinz Lycklama. UNIX on a micro-processor. Technical Memorandum 1154, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1977. ?? pp.

Lycklama:1978:FSS

[Lyc78a] Heinz Lycklama. File system structures for real-time applications. Technical Memorandum 1276 (TM 78-3114-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 20, 1978. 24 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-78-3114-
5_File_System_Structures_
for_Real-Time_Applications_
19780420.pdf.

Lycklama:1978:MUS

[Lyc78b] Heinz Lycklama. The MERT/ UNIX supervisor. Technical Memorandum 1278 (TM 78-3114-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 20, 1978. 17 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-78-3114-
4_The_MERT-UNIX_Supervisor_
19780420.pdf.

Lycklama:1978:UMa

[Lyc78c]  Heinz Lycklama.  UNIX

on a microprocessor. Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1978. 27 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/UNIX_on_
a_Microprocessor_19780322.
pdf.

Lycklama:1978:UMb

[Lyc78d] Heinz Lycklama. UNIX on a microprocessor. Technical Memorandum TM-78-3114-1, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 22, 1978. 15 pp.

URL https://www.tuhs.org/
Archive/Documentation/TechReports/
Heinz_Tech_Memos/TM-78-3114-
1_UNIX_on_a_Microprocessor_
19780322.pdf.

Lycklama:1978:UTSb

[Lyc78e] Heinz Lycklama. UNIX timesharing system: UNIX on a microprocessor. The Bell System Technical Journal, 57(6):

2087–2101, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2087.
                                        pdf.

Lycklama:1985:UMY

[Lyc85] Heinz Lycklama. UNIX on a microprocessor — 10 years later. In USENIX Association [USE85d], pages 5–16. LCCN QA76.8.U65 U8 1985.

Lyons:1976:CUA

[Lyo76a] T. G. Lyons. Cautions on the use of the alarm signal. Technical Memorandum 1094, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 26, 1976. ?? pp.

Lyons:1976:GUR

[Lyo76b] T. G. Lyons. Guide to UNIVAC Remote Job Entry for Programmer’s Workbench users. Technical Memorandum 1217, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 16, 1976. ?? pp.

Lyon:1977:IUPa

[Lyo77] Thomas L. Lyon. Inter-UNIX portability. Report, Bell Laboratories, Murray Hill, NJ 07974, August 1977. 8 pp.

                                        URL https://akapugsblog.
                                        files.wordpress.com/2018/
                                        05/inter-unix_portability.
                                        pdf.

Lyon:1979:URU

[Lyo79] R. B. Lyon. Use of the revised UNIX-to-UNIX copy command for Lab 323. Technical Memorandum 1361, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 17, 1979. ?? pp.

Lyon:2008:NNS

[Lyo08] Gordon Fyodor Lyon. Nmap network scanning: official Nmap project guide to network discovery and security scanning. Insecure.Com, LLC, Sunnyvale, CA, USA, 2008. ISBN 0-9799587-1-7 (paperback). xxix + 434 pp. LCCN TK5105.59 .L96 2008.

Lyon:2018:RUIa

[Lyo18a] Tom Lyon. The roots of UNIX for the IBM mainframe (part 1). Web site, May 12, 2018.

URL https://akapugs.blog/
2018/05/12/370unixpart1/.

Lyon:2018:RUIb

[Lyo18b] Tom Lyon. The roots of UNIX for the IBM mainframe (part 2). Web site, May 12, 2018.

URL https://akapugs.blog/
2018/05/12/370unixpart2/.

Lyon:2018:RUIc

[Lyo18c] Tom Lyon. The roots of UNIX for the IBM mainframe (part 3). Web site, May 12, 2018.

URL https://akapugs.blog/
2018/05/12/370unixpart3/.

Lin:2023:VML

[LZAH23] Jiahuei Lin, Haoxiang Zhang, Bram Adams, and Ahmed E.

Hassan. Vulnerability management in Linux distributions. Empirical Software Engineering, 28(2):??, March 2023. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s10664-022-10267-7.

Liu:2011:STM

[LZHS11] Chang Liu, Qing Zhu, Kenneth A. Holroyd, and Elizabeth K. Seng. Status and trends of mobile-health applications for iOS devices: a developer’s perspective. The Journal of Systems and Software, 84(11):2022–2033, November 2011. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0164121211001610.

Macdonald:1983:HFB

[Mac83] N. H. Macdonald. Human factors and behavioral science: The UNIX Writer’s Workbench software: Rationale and design. The Bell System Technical Journal, 62(6): 1891–1908, July/August 1983. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol62/bstj62-6-1891.
pdf.

MacKay:TB7-1-13

[Mac86] Pierre MacKay. UNIX Site Report. TUGboat, 7(1):13– 14, March 1986. ISSN 0896-

3207. URL https://tug.org/
TUGboat/tb07-1/tb14short.
pdf.

MacKay:TB11-3-453

[Mac90]  Pierre A. MacKay.  TEX

site report. TUGboat, 11 (3):453–454, September 1990. ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb11-3/
tb29site.pdf.

MacLachlan:1991:CCL

[Mac91] Rob MacLachlan. CMU Common Lisp user’s manual. Research paper CMU-CS-91-108, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, February 1991. vi + 168 pp. This is a revised version of Technical Report CMU-CS-87-156.

MacLachlan:1992:CCL

[Mac92] Rob MacLachlan. CMU Common Lisp user’s manual. Research paper CMU-CS-92-161, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, July 1992. v + 142 pp. Supersedes Technical Reports CMU-CS-87-156 and CMU-CS-91-108.

MacGuire:1997:BBW

[Mac97a] Sean MacGuire. Big brother:

a Web-based UNIX system and network monitor. Sys Admin: The Journal for UNIX Systems Administrators, 6(3): 43–??, March 1997. CODEN SYADE7. ISSN 1061-2688.

Macpherson:1997:TSP

[Mac97b] P. A. Macpherson. A technique for student program submission on UNIX systems. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 29(4):54–56, December 1997. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Madruga:1994:BRI

[Mad94] Ewerton Longoni Madruga. Book review: Internetworking with TCP/IP, vol. III: Client-Server programming and applications (BSD Sockets version) by Douglas E. Comer and David L. Stevens (Prentice-Hall, 1993). ACM SIGARCH Computer Architecture News, 22(3):29–30, June 1994. CO-DEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Madell:1997:DFM

[Mad97] Tom Madell. Disk and file management tasks on HP-UX. Hewlett–Packard Professional Books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-518861-X. xvii + 176 pp. LCCN QA76.9.H35 M33 1997.

Maginnis:2000:SLG

[Mag00] Tobin Maginnis. SAIR Linux and GNU certification level I: installation and configuration. Wiley, New York, NY, USA, 2000. ISBN 0-471-36978-0 (paperback). xix + 284 pp. LCCN QA76.3 .M3235 2000.

Maginnis:2004:GLZ

[Mag04] Tobin Maginnis. GNU-Linux-Zertifizierung: Prüfungsvorbereitung zum Sair Linux and GNU Certified Professional/Administrator (LCP/LCA). (German) [GNU/ Linux Certification: Examination Preparation for Sair Linux and GNU Certified Professional/Administrator (LCP/ LCA)]. dpunkt-Verlag, Heidelberg, Germany, 2004. ISBN ???? ???? pp. LCCN ????

Mahoney:1989:IMD

[Mah89] Michael S. Mahoney. Interview with M. D. McIlroy. Web site, August 18, 1989. URL https:

//www.princeton.edu/~hos/
mike/transcripts/mcilroy.
htm.

Maher:2007:MPU

[Mah07] Tim Maher. Minimal Perl:

for UNIX and Linux people. Manning Publications, Greenwich, CT, USA, 2007. ISBN 1-932394-50-8 (paperback). xxxviii + 450 pp. LCCN QA76.73.P22 M34 2007.

Mahoney:20xx:OHU

[Mahxx] Michael Mahoney. An oral history of Unix. Web site, 20xx. URL https:

//www.princeton.edu/~hos/
Mahoney/unixhistory.

Majernik:1977:NSP

[Maj77] J. F. Majernik. Notes on the selection, procurement, installation, and usage of UNIX by department. … Technical Memorandum 1201 (5254-771021.01EN), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 21, 1977. ?? pp.

Mansfield:1991:XWS

[Man91] Niall Mansfield. The X Window System: a user’s guide. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-56344-4. xviii + 344 pp. LCCN QA76.76.W56 .M35 1991.

Mann:1992:UAM

[Man92]  Daniel Mann.  Unix and

the Am29000 microprocessor. IEEE Micro, 12(1):23–31, January/February 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mansfield:1993:JXO

[Man93a] Niall Mansfield. The Joy of X — An Overview of the X Window System. Addison-Wes-ley, Reading, MA, USA, 1993. ISBN 0-201-56512-9. xi + 368 pp. LCCN QA76.76 .W56 M45x 1993. US$22.75.

Mansfield:1993:XWS

[Man93b] Niall Mansfield. The X Window System — a User’s Guide. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-54438-5. Another good tutorial-type book to using X.

Mancill:2001:LRP

[Man01a] Tony Mancill. Linux routers: a primer for network administrators. Prentice Hall series in computer networking and distributed systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-086113-8. xx + 345 pp. LCCN QA76.76.O63 M348 2001.

Mansfield:2001:OBU

[Man01b] Bill Mansfield. Oracle backup for UNIX administrators. Sys Admin: The Journal for UNIX Systems Administrators, 10(5): 24, 26, 28–29, May 2001. CO-DEN SYADE7. ISSN 1061-

2688.  URL http://www.
samag.com/.

Martin:2022:TLA

[MAP^{+}22] Hugo Martin, Mathieu Acher, Juliana Alves Pereira, Luc Lesoil, Jean-Marc Jézéquel, and Djamel Eddine Khelladi. Transfer learning across variants and versions: The case of Linux kernel size. IEEE Transactions on Software Engineering, 48(11):4274–4290, November 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Maranzano:1975:DUF

[Mar75] J. F. Maranzano. A description of the UNIX file system. Technical Memorandum 1065 (MF 75-8234-32), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 16, 1975. ?? pp.

Maranzano:1978:EUH

[Mar78a] J. F. Maranzano. Effects on UNIX host CPU resources by the use of intelligent terminals …. Technical Memorandum 1329, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 13, 1978. ?? pp.

Maranzano:1978:NPI

[Mar78b]  J. F. Maranzano.  New

permuted index program on UNIX. Technical Memorandum 1263 (MF 78-8234-59), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1, 1978. ?? pp.

Maranzano:1978:NTR

[Mar78c] J. F. Maranzano. New trouble reporting procedures. Technical Memorandum 6, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 2, 1978. ?? pp.

Maranzano:1978:USP

[Mar78d] J. F. Maranzano. On understanding the support of the Portable C Compiler. Technical Memorandum 1269 (MF 78-8234-49), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 26, 1978. ?? pp.

Maranzano:1978:SSS

[Mar78e] J. F. Maranzano. Sdiff, a side-by-side difference program. Technical Memorandum 1328, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 13, 1978. ?? pp.

Marcus:1978:PWF

[Mar78f] H. Marcus. A PROM writer facility for UNIX improved version. Technical Memorandum 1281, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1978. ?? pp.

Marione:1978:ALU

[Mar78g] P. Marione. ACS LEAP user’s guide. Technical Memorandum 11 (TM 78-3112-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 25, 1978. ?? pp.

Marca:1982:RCU

[Mar82] David Marca. A repetition construct for UNIX version 6. ACM SIGPLAN Notices, 17(9): 72–75, September 1982. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:1984:USP

[Mar84] R. L. Martin. The UNIX system: Preface. AT&T Bell Laboratories Technical Journal, 63 (8 part 2):1571–1572, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Marsh:2001:PRU

[Mar01a] Matthew G. Marsh. Policy routing using Linux. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32052-5. vii + 205 pp. LCCN TK5105.543 .M37 2001.

Marxer:2001:AUD

[Mar01b] Roman Marxer. Automatic UNIX documentation with unixdoc. Sys Admin: The Journal for UNIX Systems Administrators, 10(6):??, June 2001. CODEN SYADE7. ISSN 1061-2688. URL http://www.

                                        samag.com/. Web only.

Marchesin:2004:ULR

[Mar04] Armand Marchesin. Using Linux for real-time applications. IEEE Software, 21 (5):18–20, September/October 2004. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Marsh:2006:PRU

[Mar06] Matthew G. Marsh. Policy Routing Using Linux. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32052-5. 205 (est.) pp. LCCN ???? EUR 35.90.

Marsh:2010:NCF

[Mar10] Nicholas Marsh. Nmap Cookbook: The Fat-free Guide to Network Scanning. CreateSpace, Scotts Valley, CA, USA, 2010. ISBN 1-4499-0252-9. 198 (est.) pp. LCCN ????

Mashey:1977:PWD

[Mas77a] John R. Mashey. Programmer’s Workbench documentation roadmap (B.1). Technical Memorandum 1121 (TM 77-8234-18), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 30, 1977. 7 pp.

Mashey:1977:PUS

[Mas77b] John R. Mashey. PWB/UNIX shell tutorial (B.5). Technical Memorandum 1189, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1977. ii + 25 pp.

Mashey:1987:ULP

[Mas87] John R. Mashey. UNIX: Leverage — past, present, future. In USENIX Association [USE87e], pages 1–8. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Mashey:1988:UPP

[Mas88] John R. Mashey. UNIX past, present, and future: Changing roles, changing technologies. In USENIX Association [USE88c], pages 7–13. ISBN ???? LCCN ????

Mason:2005:PVC

[Mas05] Michael (Michael G.) Mason.

Pragmatic version control using subversion. Pragmatic starter kit. Pragmatic Bookshelf, Raleigh, NC, USA, 2005. ISBN 0-9745140-6-3. xii + 207 pp. LCCN QA76.76.D4 M3 2005 itb.

MathlabGroup:1980:AMM

[Mat80] Mathlab Group. Addendum to the Mathlab/MIT MACSYMA reference manual for VAX/ UNIX version and VAX/VMS version. Computing Services, University of California, Berkeley, Berkeley, CA, USA, 1980. 5 + [3] pp.

Matthews:1985:PMP

[Mat85a] D. C. J. Matthews. Poly manual: Poly and standard ML. ACM SIGPLAN Notices, 20(9): 52–76, September 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merritt:1985:FLU

[MAT85b] Doug Merritt, Ken Arnold, and Bob Toxen. Fear and loathing on the UNIX trail ’76. UNIX Review, 3(1):44, 108–110, 112, 114, January 1985. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/Berkeley_Unix_History.
pdf.

MathWorks:1993:GST

[Mat93] MathWorks, Inc. Guide to services: technical support, product updates, Internet services, newsletter, short courses: for UNIX and VMS users. Technical report, MathWorks, Inc., Natick, MA, USA, September 1993. 16 pp.

Mateosian:1997:MRA

[Mat97]  Richard Mateosian.  Micro

review: Adobe Framemaker-5.5 for Windows, Macintosh, and UNIX systems. IEEE Micro, 17(6):86–88, November/December 1997. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

URL http://dlib.computer.
org/mi/books/mi1997/pdf/
m6086.pdf.

Matzan:2007:FCC

[Mat07a] Jem Matzan. The FreeBSD 6.2 crash course. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51016-0. LCCN QA76.754.T53; QA76.754.T53 M38 2007eb.

URL http://www.oreilly.
com/catalog/9780596510169.

Matzan:2007:OCC

[Mat07b] Jem Matzan. The OpenBSD 4.0 crash course. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-51015-2. ???? pp. LCCN QA76.76.O63 M38

2007. URL http://proquest.
safaribooksonline.com/0596510152;
http://proquest.safaribooksonline.
com/9780596510152.

Mahmoud:2008:PAS

[MAT08] Feras A. Mahmoud and Mohammad H. Al-Towaiq. Parallel algorithm for the solutions of PDEs in Linux clustered workstations. Applied Mathematics and Computation, 200(1): 178–188, June 15, 2008. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

Mauerer:1997:TLU

[Mau97] Wolfgang Mauerer. Textverarbeitung mit L^{A}TEX 2ε unter UNIX. UNIX easy. Carl Hanser, München, Germany, 1997. ISBN 3-446-18909-2. xiv + 361 pp. LCCN ???? DM 49,80.

Mauerer:2008:PLK

[Mau08] Wolfgang Mauerer. Professional Linux kernel architecture. Wrox professional guides. Wiley, New York, NY, USA, 2008. ISBN 0-470-34343-5 (paperback). xxx + 1337 pp. LCCN QA76.9.A73 M38

2008eb. URL http://site.
ebrary.com/lib/princeton/
Doc?id=10257612.

Maxwell:2000:RLN

[Max00] Steven Maxwell. Red Hat Linux network management tools. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-212260-9, 0-07-212261-7 (CD-ROM), 0-07-212602-7 (CD-ROM), 0-07-212262-5 (set). xxviii + 683 pp. LCCN QA76.76.O63 M373339 2000.

Maxwell:2001:LCK

[Max01] Scott Andrew Maxwell. Linux Core Kernel commentary. Coriolis Group Books, Scottsdale, AZ, USA, 2001. ISBN 1-58880-149-7. xviii + 717 pp. LCCN QA76.76.O63 M373337 2001. US$49.99.

Maxwell:2002:USA

[Max02] Steven Maxwell. UNIX system administration: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2002. ISBN 0-07-219486-3. xx + 675 pp. LCCN QA76.76.O63 M39 2002.

Maybach:1978:VRP

[May78] R. L. Maybach. A version of Ratfor for the PDP-11’s RT-11 operating system. Technical Memorandum 1249 (TM 78-4392-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 10, 1978. ?? pp.

Mayer:1979:UUC

[May79] W. J. Mayer. Using UNIX capabilities more effectively. Technical Memorandum 1370, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 30, 1979. ?? pp.

Mayer:1984:EPP

[May84] Herbert Mayer. Experiences porting Pascal source from a micro computer to a Vax. ACM SIGPLAN Notices, 19(9):16– 20, September 1984. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maynard:1997:USA

[May97] Jack Maynard. UNIX security auditing: a practical guide. Sys Admin: The Journal for UNIX Systems Administrators, 6(6):67–72, June 1997. CODEN SYADE7. ISSN 1061-2688.

May:2000:PHP

[May00] John M. May. Parallel I/O for High Performance Computing. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2000. ISBN 1-55860-664-5. xvii + 366 pp. LCCN QA76.88.

M39 2001. US$54.95. URL

http://www.mkp.com/books_
catalog/catalog.asp?ISBN=
1-55860-664-5.

Mayer:2001:SPU

[May01]  Alexander Mayer.  Shell-

Programmierung in Unix: [das Lehrbuch zu Bourne-Shell, Korn-Shell, Z-Shell, Bash für Linux, Unix, BSD und MacOS X]. (German) [Shell programming in Unix: The textbook for the Bourne Shell, Korn Shell, Z Shell, Bash for Linux, Unix, BSD and MacOS X]. Computer-&-Literatur-Verlag, Böblingen, Germany, 2001. ISBN 3-932311-78-7. 767 pp. LCCN ????

Maranzano:1977:TIA

[MB77] J. F. Maranzano and Stephen R. Bourne. A tutorial introduction to ADB. Technical Memorandum 1185 (TM 77-8234-11, TM 77-1273-10), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 5, 1977. ?? pp.

Mason:1990:TLT

[MB90] Tony Mason and Doug Brown. lex & yacc. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-49-8. xviii + 216 pp. LCCN QA76.76.O63 M37 1990.

Mazlakowski:2000:STY

[MB00] Mark Mazlakowski and Tony Butcher. Sams Teach Yourself MySQL in 21 Days.

SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31914-4. xix + 532 pp. LCCN QA76.73.S67 M328 2000. US$39.99. URL http:/

/www.mcp.com/sams/detail_
sams.cfm?item=0672319144.

McReynolds:2003:AGP

[MB03] Tom McReynolds and David Blythe. Advanced Graphics Programming with OpenGL. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2003. ISBN 1-55860-659-9. 600 (est.) pp. LCCN ???? US$55.00, UK£36.95.

Moskowitz:2005:WLI

[MB05] Jeremy Moskowitz and Thomas Boutell. Windows and Linux integration: hands-on solutions for a mixed environment. Wiley, New York, NY, USA, 2005. ISBN 0-7921-4428-6, 0-7821-4428-4. xix + 539 pp. LCCN ????

Mudge:1991:DM

[MBB^{+}91] Trevor N. Mudge, Richard B.

Brown, William P. Birmingham, Jeffrey A. Dykstra, Ayman I. Kayssi, Ronald J. Lomax, Oyekunle A. Olukotun, Karem A. Sakallah, and Raymond A. Milano. The design of a microsupercomputer. Computer, 24(1):57–64, January 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mackay:1986:UBS

[MBC^{+}86] D. Mackay, G. Ball, M. Crowe, M. Hughes, D. Jenkins, and C. Nicol. A UNIX-based system for software configuration management. The Computer Journal, 29(6):527–530, December 1986. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/content/29/6/527.full.
pdf+html;   http://www3.
oup.co.uk/computer_journal/
hdb/Volume_29/Issue_06/tiff/
527.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_06/tiff/
528.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_06/tiff/
529.tif;   http://www3.oup.
co.uk/computer_journal/hdb/
Volume_29/Issue_06/tiff/
530.tif.

McKusick:1996:DIO

[MBKQ96] Marshall Kirk McKusick, Keith Bostic, Michael J. Karels, and John S. Quarterman. The Design and Implementation of the 4.4BSD Operating System. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-54979-4. xxvi + 580 pp. LCCN QA76.76.O63D4743 1996. US$44.95. URL http:

//wiki.tuhs.org/doku.php?
id=publications:4.4bsd_design.

Muster:1989:UPU

[MBL89] John C. C. Muster, Peter M. Birns, and Lurnix. UNIX Power Utilities for Power Users. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1989. ISBN 1-55828-000-6. 420 pp. LCCN QA76.76 .O63 M87 1989. US$24.95.

Morris:1974:CDT

[MC74] Robert Morris and Lorinda L.

Cherry. Computer detection of typographical errors. Technical Memorandum 1019 (TM 73-1271-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 3, 1974. ?? pp.

Morris:1975:ACT

[MC75a] Robert Morris and Lorinda L.

Cherry. Appendix A. Common technical English words, and Appendix B. Possible typo’s and spelling errors. IEEE Transactions on Professional Communication, PC-18(1):57– 64, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). See [MC75b].

Morris:1975:CDT

[MC75b] Robert Morris and Lorinda L.

Cherry. Computer detection of typographical errors. IEEE Transactions on Professional Communication, PC-18(1):54–56, March 1975. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic). URL

https://ieeexplore.ieee.
org/abstract/document/6593963.

See supplemental word list [MC75a].

Morris:1978:DID

[MC78] Robert Morris and Lorinda Cherry. DC — an interactive desk calculator. Technical Memorandum 1056, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 15, 1978. 8 pp.

Mahler:1985:AAP

[MC85a] Stephen J. Mahler and David A. Curry. Access — a program to interpret pathname access permissions for the UNIX operating system. In USENIX Association [USE85c], pages 59–64.

Mahler:1985:API

[MC85b] Stephen J. Mahler and David A. Curry. Access — a program to interpret pathname access permissions for the UNIX operating system. In USENIX [USE85b], pages 59–64.

McCarron:1988:UUS

[McC88] Shane P. McCarron. An update on UNIX standards activities. ;login: the USENIX Association newsletter, 13(5):18–22, September/October 1988. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

McCormack:1990:WFX

[McC90] Joel McCormack. Writing fast X servers for dumb color frame buffers. Software—Practice and Experience, 20(S2):83–108, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McClaughry:1992:PPT

[McC92] Patrick E. McClaughry. PTOPP:

a practical toolset for the optimization of parallel programs. Thesis (M.S.), University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, May 1992. vii + 22 pp.

McCusker:1995:USR

[McC95] Tom McCusker. UNIX Systems Rule Now — but Make Way for SMP. Datamation, 41(10): 57–??, June 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

McCarthy:1998:ISS

[McC98] Linda McCarthy. Intranet security: stories from the trenches. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1998. ISBN 0-13-894759-7. xxv + 260 pp. LCCN TK5105.59 .M33 1998.

McCarty:1999:LDG

[McC99a] Bill McCarty. Learning Debian GNU/Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-705-2. xiii + 343 pp. LCCN QA76.76.O63 M3758 1999. US$34.95. URL http:/

/www.oreilly.com/catalog/
9781565927056;   http://
www.oreilly.com/catalog/

debian. Includes CD-ROM.

McCarty:1999:LRL

[McC99b] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-627-7. xiii + 378 pp. LCCN QA76.76.O63 M379 1999. US$34.95. URL http:/

/www.oreilly.com/catalog/
9781565926271;   http://
www.oreilly.com/catalog/
redhat.

McCune:2001:ILW

[McC01]  Mike McCune.  Integrating

Linux and Windows. Prentice Hall PTR open source technology series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-030670-3. xxii + 344 pp. LCCN QA76.9.U83 M46 2000. US$39.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130306703.html.

McCarty:2002:LRL

[McC02] Bill McCarty. Learning Red Hat Linux: a Guide to Red Hat Linux for New Users. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00071-5. xx + 346 pp. LCCN QA76.76.O63 M376 2002. US$34.95. URL http://

safari.oreilly.com/0596000715;
 http://www.oreilly.com/
catalog/9780596000714;
http://www.oreilly.com/catalog/

redhat2. Includes CD-ROM with Red Hat Linux 7.2.

McCarty:2003:LRL

[McC03] Bill McCarty. Learning Red Hat Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 0-596-00469-9. xvi + 319 pp. LCCN QA76.76.O63 M376 2003. URL http:/

/www.oreilly.com/catalog/
9780596004699.

McCarty:2004:LRE

[McC04] Bill McCarty. Learning Red Hat Enterprise Linux and Fedora. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2004. ISBN 0-596-00589-X. xvii + 326 pp. LCCN QA76.76.O63 M376 2004. URL http:/

/www.oreilly.com/catalog/
9780596005894.

McCarty:2005:SNO

[McC05]  Bill McCarty.  SELINUX:

NSA’s open source Security Enhanced Linux. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00716-7. xiii + 238 pp. LCCN TK5105.59 .M37 2005.

URL http://www.oreilly.
com/catalog/9780596007164.

McDowell:1984:UBC

[McD84a] Rex McDowell. A UNIX-Based color graphics workstation. In USENIX Association [USE84b], pages 115–122. ISBN none. LCCN QA76.8.U65 U55 1984.

McDowell:1984:UCG

[McD84b] Rex McDowell. A UNIX-Based color graphics workstation. In USENIX Association [USE84b], pages 115–122. ISBN none. LCCN QA76.8.U65 U55 1984.

McDonald:1987:FFU

[McD87] Chris S. McDonald. fsh — a functional UNIX command interpreter. Software—Practice and Experience, 17(10): 685–700, October 1987. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McElhearn:2005:MXC

[McE05] Kirk McElhearn. The Mac OS X command line: Unix under the hood. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2005. ISBN 0-7821-4354-7, 0-470-11385-5. xxv + 438 pp. LCCN QA76.76.O63 M387152 2005eb; QA76.8.M3 M395 2005.

McGuire:1979:BBP

[McG79] J. A. McGuire. Battery backup for PDP 11/70 MOS memory. Technical Memorandum 1418, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 2, 1979. ?? pp.

McGibbon:1997:UAM

[McG97] Barry McGibbon. UNIX application management: Legacy reengineering with Java and the Web. UNIX Developer, 1(1):

31–34, January/February 1997. ISSN 1090-2279.

McDougall:1999:RM

[MCH^{+}99] Richard McDougall, Adrian Cockcroft, Evert Hoogendoorn, Enrique Vargas, and Tom Bialaski. Resource Management. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1999. ISBN 0-13-025855-5. xxviii + 334 pp. LCCN QA76.9.D3R472 1999. US$40.00. URL books/

rm.pdf;   http://www.sun.
com/books/catalog/mcdougall3/

. Part No. 805-7268-10 July 1999, Revision A.

McIlroy:1960:MIE

[McI60] M. Douglas McIlroy. Macro instruction extensions of compiler languages. Communications of the Association for Computing Machinery, 3(4):214–220, April 1960. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McIlroy:1963:VMF

[McI63] M. D. McIlroy. A variant method of file searching. Communications of the Association for Computing Machinery, 6 (3):101, March 1963. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McIlroy:1969:LES

[McI69a] M. D. McIlroy. Letter to the Editor: A scheme for specifying running titles. Communications of the Association for Computing Machinery, 12 (7):413, July 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McIlroy:1969:AGS

[McI69b] M. Douglas McIlroy. Algorithm 354: Generator of spanning trees [H]. Communications of the Association for Computing Machinery, 12(9):511, September 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McIlroy:1971:AFC

[McI71] M. D. McIlroy. ANS FOR-TRAN charts. Computing Science Technical Report 13, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1971. ?? pp.

McIlroy:1972:MTC

[McI72] M. Douglas McIlroy. A manual for the TMG compiler-writing language. Technical Memorandum 1009 (MM 72-1271-8), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 13, 1972. ?? pp.

McIlroy:1973:RTF

[McI73a] M. D. McIlroy. Roff text formatter. Technical Report MHCC-005 (revision 1), AT&T Bell Laboratories, Murray Hill, NJ, USA, 1973. ?? pp.

McIlroy:1973:SES

[McI73b] M. Douglas McIlroy. Synthetic English speech by rule. Technical Memorandum 1025 (TM 73-1271-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 7, 1973. ?? pp.

McIlroy:1974:SES

[McI74] M. Douglas McIlroy. Synthetic English speech by rule. Computing Science Technical Report 14, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1974. ii + 38 pp. URL

https://www.cs.dartmouth.
edu/~doug/speak.tar.

McIlroy:1982:DSL

[McI82]  M. Douglas McIlroy.  De-

velopment of a spelling list. IEEE Transactions on Communications, COM-30(1 pt 1): 91–99, January 1982. CO-DEN IECMBT. ISSN 0096-

1965. URL http://www.cs.
dartmouth.edu/~doug/spell.
pdf.

McIlroy:1983:BAC

[McI83] M. McIlroy. Best approximate circles on integer grids. ACM Transactions on Graphics, 2(4): 237–263, October 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McIlroy:1987:RUR

[McI87] M. Douglas McIlroy. A research UNIX reader: Annotated excerpts from the programmer’s manual, 1971–1986. Technical Report CSTR 139, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1987. 16 pp. URL http:

//www.cs.dartmouth.edu/~
doug/reader.pdf.

McIlroy:COMPSYS-2-2-173

[McI89] M. Douglas McIlroy. Virology 101 (UNIX system virus). Computing Systems, 2(2):173– 181, Spring 1989. CODEN CM-SYE2. ISSN 0895-6340.

McIlroy:1990:SPS

[McI90] M. Douglas McIlroy. Squinting at power series. Software —Practice and Experience, 20 (7):661–683, July 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McIlroy:1992:GRE

[McI92] M. Douglas McIlroy. Getting raster ellipses right. ACM Transactions on Graphics, 11 (3):259–275, July 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/toc/Abstracts/
0730-0301/130892.html.

McIlroy:1999:KAQ

[McI99] M. D. McIlroy. A killer adversary for quicksort. Software—Practice and Experience, 29(4):341–344, April 10, 1999. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
55001843;   http://www3.
interscience.wiley.com/cgi-
bin/fulltext?ID=55001843&
PLACEBO=IE.pdf.

McIlroy:19xx:TNR

[McIxx] M. Douglas McIlroy. There is no royal road to programs: A trilogy on raster ellipses and programming methodology. Computing Science Technical Report 155, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 29 pp.

McIlroy:2012:TDR

[McI12] Doug McIlroy. Tribute to Dennis Ritchie. Alcatel-Lucent 26-minute video., September 11, 2012. URL https:/

/www.youtube.com/watch?v=

Y-TrkYL0VtI. From an event entitled, “The Lasting Legacy of Dennis Ritchie: The Impact of Software on Society.”.

McIlroy:1975:US

[McI5 ] M. Douglas McIlroy. UNIX summary. Technical Memorandum 1043, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1975 (??). ?? pp.

McKee:1982:CUC

[McK82] Roger McKee. The coming UNIX crash. In Usr Group [Usr82], pages 281–??

McKusick:1984:UGT

[McK84] Marshall Kirk McKusick. Using gprof to tune the 4.2BSD kernel. In ????, editor, Proceedings of the European UNIX Users Group Meeting, Nijmegen, Netherlands. April 1984, page ?? ????, ????, 1984.

McKusick:1985:BOT

[McK85a] Marshall Kirk McKusick. A Berkeley odyssey: Ten years of BSD history. UNIX Review, 3(1):30, 32, 34, 36, 38, 40, 42, January 1985. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/Berkeley_Unix_History.
pdf;   http://www.tuhs.org/
Archive/Documentation/Unix_
Review/unixreview_1985jan.
pdf.

McKusick:1985:UUU

[McK85b] Marshall Kirk McKusick. UNIX unleashed: The university role in maintaining system integrity. UNIX Review, 3(1): 28–33, January 1985. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/Berkeley_Unix_History.
pdf.

McKusick:1991:BUK

[McK91] Marshall Kirk McKusick. 4.3 BSD UNIX kernel internals: implementation, tuning, and networking: October 28– November 1, 1991, Engineering 819.188, lecture notes. Los Angeles, CA, USA, 1991. various pp.

McKusick:1992:BUK

[McK92a] Marshall Kirk McKusick. 4.3 BSD UNIX kernel internals: implementation, tuning, and networking: April 6–10, 1992, Engineering 819.188: lecture notes. Los Angeles, CA, USA, 1992. various pp.

McKusick:1992:PBU

[McK92b] Marshall Kirk McKusick. A Preview of 4.4 BSD UNIX kernel intervals: implementation, tuning, and networking: October 26–30, 1992, Engineering 819.235: lecture notes. Los Angeles, CA, USA, 1992. various pp.

McKusick:1992:UVB

[McK92c] Marshall Kirk McKusick. UNIX variants: BSD. UNIX Review, 10(4):46–47, April 1992. CO-DEN UNRED5. ISSN 0742-3136.

McKay:1993:USS

[McK93a] Calum D. McKay. UNIX security symposium report. ;login: the USENIX Association newsletter, 18(6):4–8, November/December 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

McKusick:1993:PUP

[McK93b] Marshall Kirk McKusick. Le prospettive di UNIX su PC. Informatica Oggi and UNIX, 13 (90):34–35, ???? 1993.

McKusick:1994:NCU

[McK94a] Marshall Kirk McKusick. Networking comes to UNIX. UNIX Review, 12(5):40, May 1994. CODEN UNRED5. ISSN 0742-

McKusick:1994:WN

[McK94b] Marshall Kirk McKusick. What’s new in 4.4BSD. UNIX Review, 12(1):51–56, January 1994. CODEN UNRED5. ISSN 0742-3136.

McKusick:1996:SSF

[McK96a] Marshall Kirk McKusick. Secondary storage and filesystems. ACM Computing Surveys, 28(1):217–219, March 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL

http://www.acm.org/pubs/
citations/journals/surveys/
1996-28-1/p217-mckusick/;
 http://www.acm.org/pubs/
toc/Abstracts/surveys/234404.
html.

McKusick:1996:UKIb

[McK96b] Marshall Kirk McKusick. UNIX kernel internals: October 14– 18, 1996, Engineering 819.235. Los Angeles, University Extension, Dept. of Engineering, Information Systems, and Technical Management, Short Course Program, 1996. various pp.

McKusick:1997:UKIb

[McK97a] Marshall Kirk McKusick. UNIX kernel internals: October 20– 24, 1997, Engineering 819.235. Los Angeles, University Extension, Dept. of Engineering, Information Systems, and Technical Management, Short Course Program, 1997. various pp.

McKusick:1997:UKIa

[McK97b] Marshall Kirk McKusick. UNIX kernel internals: September 22-26, 1997, Engineering 819.235. Los Angeles, University Extension, Dept. of Engineering, Information Systems, and Technical Management, Short Course Program, 1997. various pp.

McKusick:1998:CSC

[McK98] Kirk McKusick. Complete set of CSRG BSD releases available. ;login: the USENIX Association newsletter, 23(6): ??, October 1998. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1998-10/mailing.html.

McKusick:2002:RFB

[McK02] Marshall Kirk McKusick. Running fsck in the background. In ????, editor, Proceedings of the Usenix BSDCon 2002 Conference, February 2002, pages 55–64. USENIX Association, Berkeley, CA, USA, 2002.

McKusick:2003:EFF

[McK03] Marshall Kirk McKusick. Enhancements to the fast filesystem to support multi-terabyte storage systems. In ????, editor, Proceedings of the Usenix BSDCon 2003 Conference, September 2003, pages 79–90. USENIX Association, Berkeley, CA, USA, 2003.

McKusick:2006:DPF

[McK06] Marshall Kirk McKusick. Disks from the perspective of a file system. ;login: the USENIX Association newsletter, 31(3): 18–19, June 2006. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

McKusick:2007:BHB

[McK07] Marshall Kirk McKusick. A brief history of the BSD Fast File System. ;login: the USENIX Association newsletter, 32(3):9–18, June 2007. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL https://www.
usenix.org/publications/
login/june-2007-volume-32-
number-3/brief-history-bsd-
fast-file-system.

McKusick:2012:DPFa

[McK12a] Marshall Kirk McKusick. Disks from the perspective of a file system. Communications of the Association for Computing Machinery, 55(11):53–55, November 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

McKusick:2012:DPFb

[McK12b] Marshall Kirk McKusick. Disks from the perspective of a file system. ACM Queue: Tomorrow’s Computing Today, 10 (9):1–3, September 2012. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McKusick:2013:IPF

[McK13] Marshall Kirk McKusick. Improving the performance of fsck in FreeBSD. ;login: the USENIX Association newsletter, 38(2):20–22, April 2013. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

McKusick:2016:BHB

[McK16] Marshall Kirk McKusick. A brief history of the BSD Fast Filesystem. The FreeBSD Journal, 3(2):12–23, March/ April 2016. ISBN 0-615-88479-2. URL http://

issue.freebsdfoundation.
org/publication/?m=33057&
i=297010&p=14.

McLaren:1983:UDG

[McL83] Wayne McLaren. UNIX à la Data General. In Software Tools Users Group [Sof83], pages 79–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

McLeod:1984:IPU

[McL84] Bubette McLeod. Introducing people to UNIX. In Software Tools Users Group [Sof84], pages 344–345. LCCN QA76.8.U65 U83 1984.

McMahon:1978:SNI

[McM78a] L. E. McMahon. SED — a noninteractive text editor. Technical Memorandum 1307 (TM 78-1270-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1978. ?? pp.

McMahon:1978:UTS

[MCM78b] L. E. McMahon, Lorinda L. Cherry, and Robert Morris. UNIX time-sharing system: Statistical text processing. The Bell System Technical Journal, 57(6):2137–2154, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2137.
pdf.

McMahon:1979:SNT

[McM79] L. E. McMahon. Sed — a noninteractive text editor. Computing Science Technical Report 77, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1979. ?? pp.

McMahon:1990:SNI

[McM90] L. E. McMahon. Sed — a non-interactive text editor. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

McMinds:1992:MOM

[McM92] Donald L. McMinds. Mastering OSF/Motif Widgets. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56342-8. xxxii + 731 pp. LCCN QA76.76.W56 M52 1992.

McMinds:1993:MOM

[McM93] Donald L. McMinds. Mastering OSF/Motif Widgets. Addison-Wesley, Reading, MA, USA, second edition, 1993. ISBN 0-201-63335-3. xxxii + 731 pp. LCCN QA76.76.W56 M52 1993.

McMullen:1999:UUI

[McM99] John Harvey McMullen. UNIX User’s Interactive Workbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-099820-6. xxii + 598 pp. LCCN QA76.76.O63M3998 1999. US$34.99.

Miller:2006:ESR

[MCM06] Barton P. Miller, Gregory Cooksey, and Fredrick Moore. An empirical study of the robustness of MacOS applications using random testing. In Mayer and Merkel [MM06a], pages 77–86. ISBN 1-59593-457-X. LCCN QA76.76.T48 I583 2006. URL http:

//portal.acm.org/toc.cfm?
id=1145735.

Miller:2007:ESR

[MCM07] Barton P. Miller, Gregory Cooksey, and Fredrick Moore. An empirical study of the robustness of MacOS applications using random testing. Operating Systems Review, 41(1):78– 86, January 2007. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

McNutt:1992:SAP

[McN92] Dinah McNutt. System administration: POSIX 1003.7. UNIX Review, 10(9):34–??, September 1992. CODEN UNRED5. ISSN 0742-3136.

McNamarra:1997:UNS

[McN97] Roy McNamarra. Unix network security. Information Security Technical Report, 2(3):66–79, ???? 1997. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1363412797897129.

McNab:1999:BPL

[McN99a] A. D. McNab. BSD Portals for LINUX 2.0. Technical report, ????, ????, 1999. 12 pp. NAS2-14303 ; RTOP 509-10-61.

McNally:1999:AP

[McN99b] Jim McNally. Awk programming. DDC Publishers, New York, NY, USA, 1999. ISBN 1-56243-981-2. 124 pp. LCCN ???? US$.

McNab:2000:EUA

[McN00] A. David McNab. Extensible UNIX access control lists. Thesis (M.S.), University of California, Santa Cruz, Santa Cruz, CA, USA, 2000.

McSkimin:1978:RRD

[McS78] J. R. McSkimin. REDAS — a relational data access system for real-time applications. Technical Memorandum 1344 (TM 78-5221-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 9, 1978. ?? pp.

McSkimin:1979:LFS

[McS79] J. R. McSkimin. The Logical File System — a fast-access file system using UNIX raw I/O. Technical Memorandum 1473 (TM 79-9471-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 14, 1979. ?? pp.

McShaffry:2005:GCC

[McS05] Mike McShaffry. Game coding complete. Paraglyph Press, Scottsdale, AZ, USA, second edition, 2005. ISBN 1-932111-91-3. xxix + 906 pp. LCCN QA76.76.C672 M35 2005.

Murrel:1986:WUM

[MD86] S. L. Murrel and D. De Baer.

WVtbl user manual: bitmaps illustrating how to build tables. Report 11229-881116-16TMS, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 16, 1986.

Miller:2009:MHH

[MD09] Charles Miller and Dino A. Dai Zovi. The Mac Hacker’s Handbook. Wiley, New York, NY, USA, 2009. ISBN 0-470-39536-2 (paperback). xvi + 368 pp. LCCN ???? US$26.99.

Ma:2014:FFF

[MDAD^{+}14] Ao Ma, Chris Dragga, Andrea C. Arpaci-Dusseau, Remzi H. Arpaci-Dusseau, and Marshall Kirk McKusick. Ffsck: The fast file-system checker. ACM Transactions on Storage, 10(1):2:1–2:??, January 2014. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Meadow:1988:AUU

[Mea88] Anthony Meadow. Apple’s A/UX: UNIX for the rest of us? Dr. Dobb’s Journal of Software Tools, 13(??):44–??, ???? 1988. CODEN DDJOEB. ISSN 1044-789X.

Meadhra:1999:SLD

[Mea99] Michael Meadhra. StarOffice for LINUX for Dummies. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-0576-9. xxiv + 404 pp. LCCN A76.76.I57 M427 1999. US$19.99. URL http://www.

idgbooks.com/.

Mecklenburg:2005:MPG

[Mec05] Robert Mecklenburg. Managing Projects with GNU Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2005. ISBN 0-596-00610-1. xviii + 280 pp. LCCN QA76.76.U84 O73 2005. US$29.95.

Meek:1995:SGR

[Mee95] Brian L. Meek. Seven golden rules for producing languageindependent standards. Proceedings of the IEEE International Software Engineering Standards Symposium, pages 250–256, 1995. ISSN 1082-3670.

Meissner:1984:EVA

[Mei84] Loren P. Meissner. From the Editor: Vote on ACM constitutional amendment; whither ForTec?; Fortran 77 on micros?; UNIX Fortran 77 rumor. ACM ForTec Forum, 3(1):1–2, March 1984. CODEN ???? ISSN 0735-3731.

Meibner:1992:AAM

[Mei92] K. Meibner. Architectural aspects of multimedia CD-I integration in UNIX/X-Windows workstations. Lecture Notes in Computer Science, 614:314– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Melamed:1987:PAU

[Mel87] Anna S. Melamed. Performance analysis of Unix-based network file systems. IEEE Micro, 7(1):25–38, January/ February 1987. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). CSZ108.

Melford:1995:SUE

[Mel95] Robert J. Melford. Secure UNIX for Enterprise Computing. Datamation, 41(4):55– ??, March 01, 1995. CODEN DTMNAT. ISSN 0011-6963.

Mellander:2002:UFS

[Mel02] Jim Mellander. Unix filesystem security. Information Security Technical Report, 7(1):11–25, March 31, 2002. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S1363412702001036.

Mercurio:1982:UMS

[Mer82] Philip J. Mercurio. The UCSD MSG system: Iterative design in the UNIX environment. In Usr Group [Usr82], pages 151– 163.

Merusi:1997:WNU

[Mer97] Donald Merusi. Windows NT/ 95 for Unix professionals. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1997. ISBN 1-55558-181-1 (paperback). ix + 177 pp. LCCN QA76.76.O63 M465 1997.

Merusi:2001:PWA

[Mer01] Don E. Merusi. Programming the Win32 API and UNIX System Services. Prentice Hall PTR Microsoft technologies series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-025969-1. xv + 291 pp. LCCN QA76.66 .M47 2001. US$49.99. URL http:

                                        //www.phptr.com/ptrbooks/
                                        ptr_0130259691.html.

Meyer:2018:UUL

[Mey18] Jan Max Meyer. UniCC: A universal LALR(1) Parser Generator User Manual. Phorward Software Technologies, Castroper Str. 65, 44357 Dortmund, Germany, 2018. URL

https://www.math.utah.edu/
pub/tex/bib/gnu.bib;   https:
//phorward.info/products/
unicc/unicc.pdf.

Mankins:1983:SWM

[MF83] David Mankins and Daniel Franklin. A simple window management facility for the UNIX timesharing system. In Software Tools Users Group [Sof83], pages 203–228. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Minnich:1999:MSD

[MF99] Ronald G. Minnich and David J.

Farber. The Mether system: Distributed shared memory for SunOS 4.0. Report MS-CIS-93-24, Supercomputing Research Center, University of Pennsylvania, Philadelphia, PA 19104-6309, USA, February 1999. 13 pp. URL

http://repository.upenn.
edu/cis   reports/332;
https://www.researchgate.
net/publication/2411547_
The_Mether_System_Distributed_
Shared_Memory_for_SunOS_

McKenney:2020:RUL

[MFBWW20] Paul E. McKenney, Joel Fernandes, Silas Boyd-Wickizer, and Jonathan Walpole. RCU usage in the Linux kernel: Eighteen years later. Operating Systems Review, 54(1):47–63, August 2020. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL https:

//dl.acm.org/doi/10.1145/
3421473.3421481.

Miller:1990:ESR

[MFS90] Barton P. Miller, Lars Fredriksen, and Bryan So. An empirical study of the reliability of UNIX utilities. Communications of the Association for Computing Machinery, 33(12): 32–44, December 1990. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

tronic). URL ftp://grilled.
cs.wisc.edu/technical_papers/
fuzz-revisited.ps;   ftp:
//grilled.cs.wisc.edu/technical_
papers/fuzz.ps;   http://
www.acm.org/pubs/toc/Abstracts/
0001-0782/96279.html. This

is a fascinating paper on what happens when random input streams are fed into important UNIX utilities on several commercial UNIX systems. In some cases, the tests were able to crash the entire operating system. In 1995, a (sadly, unpublished) followup study showed that many of the failures diagnosed in 1990 still had not been repaired in the commercial systems, and that the GNU implementations were generally more robust. Both 1990 and 1995 papers, and the fuzzgenerating software, are available at the authors’ FTP site

at ftp://grilled.cs.wisc.
edu/technical_papers/fuzz.
ps and ftp://grilled.cs.
wisc.edu/technical_papers/
fuzz-revisited.ps.

McKusick:1999:SUT

[MG99] Marshall Kirk McKusick and G. Ganger. Soft updates: A technique for eliminating most synchronous writes in the fast filesystem. In ????, editor, Proceedings of the Freenix Track at the 1999 Usenix Annual Technical Conference, January 1999, pages 1–17. USENIX Association, Berkeley, CA, USA, 1999.

Mellen:2003:SSU

[MGK^{+}03] Daniel Mellen, Jess Garcia, Joe Keegan, Michael Gauthier, and Michael Royds. Securing Solaris 8 & 9 Using the Center for Internet Security Benchmark. The SANS Institute, Bethesda, MD, USA, 2003. ISBN 0-9724273-9-2. 180 (est.) pp. LCCN ???? US$39.

Mauro:2010:DDT

[MGMMI10] Jim Mauro, Brendan Gregg, Chad Mynhier, and Tariq Magdon-Ismail, editors. Dtrace: dynamic tracing in Solaris, Mac OS X and FreeBSD. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2010. ISBN 0-13-706183-8 (hardcover), 0-13-209151-8 (paperback). x + 285 pp. LCCN QA76.76.A63 .D822

                                        2010. URL http://proquest.
                                        safaribooksonline.com/?fpi=
                                        9780137061839.

Maranzano:1978:EHS

[MH78] J. F. Maranzano and R. C. Haight. Experience with hardware service contractors on PDP-11’s. Technical Memorandum 1261 (TM 78-8234-9), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 10, 1978. ?? pp.

Marcie:1988:GPT

[MH88] S. G. Marcie and R. L. Holt. General purpose transaction support features for the UNIX operating system. In USENIX Association [USE88c], pages 179–184. ISBN ???? LCCN ????

Morris:1996:WPD

[MH96] Mary E. S. Morris and Randy J. Hinrichs. Web Page Design: a Different Multimedia. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1996. ISBN 0-13-239880-X. xxix + 306 pp. LCCN QA76.76.H94 M65 1996. URL

                                        http://www.sun.com/books/
                                        catalog/morris2/index.html.

Manning:1976:UBL

[MHO^{+}76] E. G. Manning, R. Howard, C. G. O’Donnell, K. Pammett, and E. Chang. A Unix-based local processor and network access machines. Computer Networks: The International Journal of Distributed Informatique, 1(2):139–142, September 1976. CODEN CNETDP. ISSN 0376-5075 (print), 1878-3120 (electronic).

Miao:2002:TDM

[MHS02] Yu-Ben Miao, Wen-Shyang Hwang, and Ce-Kuen Shieh. A transparent deployment method of RSVP-aware applications on UNIX. Computer Networks (Amsterdam, Netherlands: 1999), 40(1):45–56, September ??, 2002. CODEN ???? ISSN 1389-1286 (print), 1872-7069 (electronic). URL

http://www.elsevier.com/
gej-ng/10/15/22/97/27/30/
abstract.html.

Michael:1979:CDD

[Mic79a] S. Michael. Character device driver development for CB-UNIX operating system. Technical Memorandum 1424 (TM 79-5252-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 13, 1979. ?? pp.

Michael:1979:UBT

[Mic79b] S. Michael. UNIX based TSPS hotel–motel traffic simulator. Technical Memorandum 1397 (TM 79-5252-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 18, 1979. ?? pp.

Michael:2003:MUS

[Mic03] Randal K. Michael. Mastering Unix shell scripting. Wiley, New York, NY, USA, 2003. ISBN 0-471-21821-9. xxii + 680 pp. LCCN QA76.76.O63 M488

                                        2003. URL http://www.loc.
                                        gov/catdir/bios/wiley044/
                                        2003544699.html;   http:/
                                        /www.loc.gov/catdir/description/
                                        wiley036/2003544699.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley032/2003544699.
                                        html.

Michael:2008:MUS

[Mic08] Randal K. Michael. Mastering Unix Shell Scripting: Bash, Bourne, and Korn Shell Scripting for Programmers, System Administrators, and UNIX Gurus. Wiley, New York, NY, USA, second edition, 2008. ISBN 0-470-18301-2 (paperback). xxx + 1002 pp. LCCN QA76.76.O63 M488 2008. US$50.00.

Mikes:1989:UMP

[Mik89] Steven Mikes. UNIX for MS-DOS Programmers. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17219-4. xxviii + 474 pp. LCCN QA76.76.O63 M525 1989.

Mikes:1990:XWS

[Mik90] Steven Mikes. X Window System Technical Reference. Addison-Wesley, Reading, MA, USA, 1990. ISBN 0-201-52370-

1. xiv + 786 pp. LCCN

QA76.76.W56 M55 1990.

Mikes:1992:XWS

[Mik92] Steven Mikes. X Window System Program Design and Development. Addison-Wes-ley, Reading, MA, USA, 1992. ISBN 0-201-55077-6. viii + 296 pp. LCCN QA76.76.W56 M56 1992. US$26.95.

Miller:1977:COF

[Mil77] R. E. Miller. Computerized office filing system. Technical Memorandum 1175 (TM 77-1353-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 10, 1977. ?? pp.

Miller:1978:UPO

[Mil78]  Richard Miller.  UNIX: a

portable operating system? Operating Systems Review, 12 (3):32–37, July 1978. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Miller:1979:UAI

[Mil79] L. E. Miller. UNIX assembler for the Intel 804X microprocessors. Technical Memorandum 1482 (TM 79-3231-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1979. ?? pp.

Miller:1984:CUS

[Mil84a] Joaquin Miller. Connecting a UNIX system to an X.25 network. In USENIX Association [USE84b], pages 47–56. ISBN none. LCCN QA76.8.U65 U55

Miller:1984:PPF

[Mil84b] John J. H. Miller, editor. PRO-TEXT I: Proceedings of the First International Conference on Text Processing Systems held in Dublin from 24th to 26th October, 1984. Boole Press, Dublin, Ireland, 1984. ISBN 0-906783-41-0 (hardcover), 0-906783-42-9 (paper). LCCN QA76.9.T48 I591 1984.

Millman:1984:HES

[Mil84c] S. (Sidney) Millman, editor. A History of Engineering and Science in the Bell System: Volume 5. Communications Sciences: (1925–1980). Bell Telephone Laboratories, New York, NY, 1984. ISBN 0-932764-06-1. xxii + 521 pp. LCCN ????

Mills:1986:MMB

[Mil86] Philip M. Mills. A multiuser multiprocessor benchmark to compare UNIX systems. In USENIX Association [USE86c], pages 59–71.

Milenkovic:1990:MMM

[Mil90a]  Milan Milenkovic.  Micro-

processor memory management units. IEEE Micro, 10(2):70– 85, March/April 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Miller:1990:OLA

[Mil90b] John David Miller. An OPEN LOOK at UNIX. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1990. ISBN 1-55851-057-5. 482 pp. LCCN QA76.76.U84 M55 1990.

Miller:1990:OLU

[Mil90c] John David Miller. An OPEN LOOK at UNIX. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1990. ISBN 1-55851-057-5. 482 pp. LCCN QA76.76.U84 M55 1990.

Miller:OLU90

[Mil90d] John David Miller. An OPEN LOOK at UNIX. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 1990. ISBN 1-55851-057-5. 482 pp. LCCN QA76.76.U84 M55 1990.

Milazzo:1991:SVU

[Mil91]  Paul G. Milazzo.  Shared

video under UNIX. In USENIX [USE91c], pages 369– 384. LCCN QA 76.76 O63 U83 1991.

Miles:1993:BVP

[Mil93] Douglas Miles. Beyond vector processing: parallel programming on the Cray APP. In 1993 IEEE Compcon Spring (Feb 22–26 1993: San Francisco, CA, USA), pages 321– 328. IEEE, Piscataway, NJ, USA, 1993. ISBN 0-7803-1294-5. LCCN ???? IEEE catalog number 93CH3251-6.

Millard:1998:EPU

[Mil98]  Steven P. Millard.  Envi-

ronmentalStats for S-PLUS: user’s manual for Windows and UNIX. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. ISBN 0-387-98486-0. xii + 381 pp. LCCN GE45.S73 M55 1998.

Miller:1999:AUP

[Mil99] Bonnie L. Miller. AIX for UNIX Professionals. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-757246-8. vi + 184 pp. LCCN QA76.76.O63M74519 1999.

URL http://www.phptr.com/
ptrbooks/ptr_0137572468.
html.

Millard:2000:EPU

[Mil00]  Steven P. Millard.  En-

vironmentalStats for S-Plus: User’s Manual for Windows and UNIX, Versions 1.0 & 1.1. Probability, Statistics & Information, Seattle, WA, USA, March 2000. x + 381 pp. LCCN ????

Miller:2001:LWA

[Mil01] Michael Joseph Miller. Linux for Windows addicts: a 12step program for habitual Windows users. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213081-4. xvii + 414 pp. LCCN QA76.76.O63 M57 2001.

Miller:2016:ECO

[Mil16] Lawrence C. Miller. Enterprise Computing with Oracle Solaris for Dummies. Wiley, New York, NY, USA, 2016. ISBN ???? ???? pp. LCCN ????

Minter:1982:HCS

[Min82a]  Charles Minter.  A high-

performance computer system suited to UNIX. In Usr Group [Usr82], pages 107–?? Abstract only.

Minter:1982:HPC

[Min82b]  Charles Minter.  A high-

performance computer system suited to UNIX. In USENIX [USE82], pages 107–?? Abstract only.

Minnich:2001:PNL

[Min01]  Ronald G. Minnich.  Pri-

vate namespaces for Linux. Dr. Dobb’s Journal of Software Tools, 26(12):23–24, 26, 28, 30, December 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Mione:1997:CMP

[Mio97] Antonino N. Mione. CDE… Motif: a practical primer. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-760828-4. 523 pp. LCCN QA76.9.U83M57 1997. US$49.00.

Mitchell:1984:ISU

[Mit84] William H. Mitchell. An Icon subsystem for UNIX emacs. Technical report TR 84-8, Dept. of Computer Science, The University of Arizona, Tucson, AZ, USA, 1984. 17 pp.

Mittelstaedt:2001:FCN

[Mit01] Ted Mittelstaedt. The FreeBSD corporate networker’s guide. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-70481-

1. xxv + 401 pp. LCCN QA76.754 .M58 2001.

Miya:1986:URU

[Miy86] E. N. Miya. User requirements for UNIX on “Big iron”. In USENIX Association [USE86e], pages 104–109. URL https://

                                        bitsavers.org/pdf/usenix/
                                        USENIX_1986_Winter_Technical_
                                        Conference_Proceedings_198601.
                                        pdf.

McKusick:1984:FFS

[MJLF84] Marshall Kirk McKusick, William N. Joy, Sam J. Leffler, and Robert S. Fabry. A fast file system for UNIX. ACM Transactions on Computer Systems, 2 (3):181–197, August 1984. CO-DEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Murrel:1984:OMH

[MK84a] S. L. Murrel and T. J. Kowalski. Overview of Monk 0.2: High-level text compiler. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 10, 1984. ?? pp.

Murrel:1984:OMT

[MK84b] S. L. Murrel and T. J. Kowalski. Overview of Monk 0.2: Typographical database. Report 11229-841210-12TMS, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 10, 1984. ?? pp.

Mankovich:1985:PUV

[MK85a] James E. Mankovich and Robert B. Kolstad. Porting the 4.2BSD UNIX virtual memory subsystem. In USENIX Association [USE85c], pages 4–10.

Murrel:1985:TDU

[MK85b] S. L. Murrel and T. J. Kowalski. Typing documents on the UNIX system: Using Monk 0.3. Report 11229-851210-11TMS, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1985 10, 1985. ?? pp.

McKusick:1987:DUB

[MK87] Marshall Kirk McKusick and Michael Karels. Directions of UNIX at Berkeley. In ????, editor, Digest of Papers of the Thirty-second IEEE Computer Society International Conference, Compcon, San Francisco, February 23–27, 1987, pages 196–199. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987.

McKusick:1988:DGP

[MK88] Marshall Kirk McKusick and Michael J. Karels. Design of a general purpose memory allocator for the 4.3BSD UNIX kernel. In Association [Ass88d], pages 295–303.

Murrel:1990:TDU

[MK90] S. L. Murrel and T. J. Kowalski. Typing documents on the UNIX system: Using Monk 0.6. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

McVoy:1991:ELP

[MK91] L. W. McVoy and S. R. Kleiman. Extent-like performance from a UNIX file system. In USENIX [USE91f], pages 33–43. FTP - ? ucbarpa.berkeley.edu:/pub/mcvoy.clust.usenix.ps.Z.

Miller:1992:AFMa

[MK92a] Ethan L. Miller and Randy H. Katz. An analysis of file migration in a Unix supercomputing environment. Technical Report UCB/CSD 92/712, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, November 1992. 12 pp. Supported in part by University Corporation for Atmospheric Research. S9128 Supported in part by NSF.

Miller:1992:AFMb

[MK92b] Ethan L. Miller and Randy H. Katz. An analysis of file migration in a Unix supercomputing environment. NASA contractor report NASA CR-192908, Computer Science Division (EECS), University of California Berkeley, Berkeley, CA, USA, 1992. ?? pp. Distributed to depository libraries in microfiche. Shipping list no.:93-1028-M. Microfiche. [Washington, DC: National Aeronautics and Space Administration, 1993] 1 microfiche.

Miller:1993:AFM

[MK93] Ethan L. Miller and Randy H.

Katz. An analysis of file migration in a UNIX supercomputing environment. In USENIX Association [USE93f], pages 421–433. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

McKusick:1996:UKIa

[MK96] Marshall Kirk McKusick and Michael J. Karels. UNIX kernel internals: implementation, tuning, and networking: May 13–17, 1996, Engineering 819.235. Los Angeles, University Extension, Dept. of Engineering, Information Systems, and Technical Management, Short Course Program, 1996. various pp.

Mohror:2004:PTS

[MK04] Kathryn Mohror and Karen L.

Karavanic. Performance tool support for MPI-2 on Linux. In ACM [ACM04], page 28. ISBN 0-7695-2153-3. LCCN ????

McKusick:1989:RE

[MKB89] Marshall Kirk McKusick, Michael Karels, and Keith Bostic. The release engineering of 4.3BSD. In ????, editor, Proceedings of the New Orleans Usenix Workshop on Software Management, April 1989, pages 95–100. USENIX Association, Berkeley, CA, USA, 1989.

McKusick:1990:PMB

[MKB90a] Marshall Kirk McKusick, Michael Karels, and Keith Bostic. A pageable memory-based filesystem. In ????, editor, Proceedings of the Anaheim Usenix Conference, June 1990, pages 137–144. USENIX Association, Berkeley, CA, USA, 1990. Reprinted in Proceedings of the United Kingdom UNIX Users Group Meeting, London, England, pp 109–115, July 1990.

McKusick:1990:R

[MKB90b] Marshall Kirk McKusick, Michael Karels, and Keith Bostic. A retrospective of 4.3BSD. Bit Magazine, ??(??):4–9, September 1990.

McKusick:1985:PIFa

[MKL85a] M. Kirk McKusick, Mike Karels, and Sam Leffler. Performance improvements and functional enhancements in 4.3BSD. Report UCB/CSD 245, Computer Science Division, University of California, Berkeley, Berkeley, CA, USA, 1985. 12 pp.

McKusick:1985:PIFb

[MKL85b] Marshall Kirk McKusick, Michael J.

Karels, and Sam Leffler. Performance improvements and functional enhancements in 4.3BSD’. In USENIX, editor, Proceedings of the Portland Usenix Conference, June 1985, pages 519–531. USENIX Association, Berkeley, CA, USA, 1985.

Miller:1995:FRR

[MKL^{+}95] Barton P. Miller, David Koski, Cjin Pheow Lee, Vivekananda Maganty, Ravi Murthy, Ajitkumar Natarajan, and Jeff Steidl. Fuzz revisited: A re-examination of the reliability of UNIX utilities and services. Technical Report 1264, University of Wisconsin-Madison, Madison, WI, USA, February 18, 1995. 23

pp.  URL ftp://ftp.cs.
wisc.edu/pub/techreports/
1995/TR1268.pdf;   ftp://
grilled.cs.wisc.edu/technical_
papers/fuzz-revisited.pdf.

German translation as Empirische Studie zur Zuverlasskeit von UNIX-Utilities: Nichts dazu Gerlernt, iX, September 1995.

McKusick:1989:CRC

[MKS^{+}89] Marshall Kirk McKusick, Michael Karels, Keith Sklower, Kevin Fall, M. Teitelbaum, and Keith Bostic. Current research by the Computer Systems Research Group of Berkeley. In Proceedings of the European UNIX Users Group Meeting, Brussels, Belgium, April 1989, page ?? ????, ????, 1989.

McKusick:1990:LA

[MKS^{+}90] Marshall Kirk McKusick, Michael Karels, Keith Sklower, Kevin Fall, M. Teitelbaum, and Keith Bostic. Looking ahead to 4.4BSD. Bit Magazine, ??(??): 10–19, September 1990.

Mbongue:2022:DMT

[MKSB22] Joel Mandebi Mbongue, Danielle Tchuinkou Kwadjo, Alex Shuping, and Christophe Bobda. Deploying multi-tenant FPGAs within Linux-based cloud infrastructure. ACM Transactions on Reconfigurable Technology and Systems (TRETS), 15(2):19:1– 19:31, June 2022. CODEN ???? ISSN 1936-7406 (print), 1936-7414 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3474058.

Musumeci:2002:SPT

[ML02a] Gian-Paolo D. Musumeci and Mike Loukides. System Performance Tuning. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00284-X. xiv + 334 pp. LCCN QA76.76.O63 L66 2002 Stacks. US$39.95. URL http://

                                        safari.oreilly.com/059600284X;
                                         http://www.oreilly.com/
                                        catalog/spt2.

Myers:2002:CNA

[ML02b] Dan Myers and Jim Lorenz.

Cisco Networking Academy Program: fundamentals of UNIX companion guide. Macmillan Publishing Co., Inc., New York, NY, USA, 2002. ISBN 1-58713-044-0. xxx + 578 pp. LCCN QA76.76.O63 M9438

2002.  URL http://www.
loc.gov/catdir/toc/fy033/
2001091176.html.

Mark:2009:BID

[ML09] Dave Mark and Jeff LaMarche.

Beginning iPhone 3 Development: Exploring the iPhone SDK. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-2459-2. 584 (est.) pp. LCCN ????

Madany:1988:CCH

[MLRC88] Peter W. Madany, Douglas E.

Leyens, Vincent F. Russo, and Roy H. Campbell. A C++ class hierarchy for building UNIX-Like file systems. In USENIX Association [USE88h], pages 65–79.

Mlynski:2009:IIP

[Mly09]  Maciej Mlynski.  The in-

fluence of the IBM pSeries servers virtualization mechanism on dynamic resource allocation in AIX 5L. Scalable Computing: Practice and Experience, 10(2):189–199, June 2009. CODEN ???? ISSN 1895-1767. URL http://www.

scpe.org/vols/vol10/no2/
SCPE_10_2_05.pdf;   http:

                                        //www.scpe.org/vols/vol10/
                                        no2/SCPE_10_2_05.zip.

McGilton:1983:IUS

[MM83a] Henry McGilton and Rachel Morgan. Introducing the UNIX System. McGraw-Hill, New York, NY, USA, 1983. ISBN 0-07-045001-3. xix + 556 pp. LCCN QA76.8.U65 M38 1983. US$18.95.

McGilton:IUS83

[MM83b] Henry McGilton and Rachel Morgan. Introducing the UNIX System. McGraw-Hill, New York, NY, USA, 1983. ISBN 0-07-045001-3. xix + 556 pp. LCCN QA76.8.U65 M38 1983. US$18.95.

Manis:1986:USP

[MM86] Rod Manis and Marc H. Meyer. The UNIX shell programming language. SAMS Publishing, Indianapolis, IN, USA, 1986. ISBN 0-672-22497-6 (paperback). xiii + 303 pp. LCCN QA76.76.O63 M35 1986. US$24.94.

Meyer:1988:UTO

[MM88] Veronika Meyer and Walter Meyer. The UNIX timesharing operating system. Computer Physics Communications, 50(1–2):51–57, July 1988. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0010465588901154.

McGilton:1990:TTU

[MM90a] Henry McGilton and Mary Mc-Nabb. Typesetting Tables on the UNIX System. Trilithon Press, Los Altos, CA, USA, 1990. ISBN 0-9626289-0-5. xxii + 282 pp. LCCN Z253.4.U53 M33 1990. US$22.00.

McGilton:TTU90

[MM90b] Henry McGilton and Mary Mc-Nabb. Typesetting Tables on the UNIX System. Trilithon Press, Los Altos, CA, USA, 1990. ISBN 0-9626289-0-5. xxii + 282 pp. LCCN Z253.4.U53 M33 1990. US$22.00.

Moore:1996:DGF

[MM96] Johanna D. Moore and Vibhu O.

Mittal. Dynamically generated follow-up questions. Computer, 29(7):75–86, July 1996. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mann:2000:LSS

[MM00] Scott Mann and Ellen L.

Mitchell. Linux system security: an administrator’s guide to open source security tools. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-015807-0. xxxvii + 564 pp. LCCN QA76.76.O63 M3515 1999. US$49.00.

Mauro:2001:SIC

[MM01a] Jim Mauro and Richard Mc-Dougall. Solaris Internals: Core Kernel Architecture. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-022496-0. xli + 657 pp. LCCN QA76.76.O63 M37195

2001. URL http://www.sun.
com/books/catalog/mauro/
index.html.

Miles:2001:EL

[MM01b] Kathy Miles and Ethan Metsger. Everyday Linux. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-091762-1. xxxiv + 369 pp. LCCN QA76.76.O63 M534 2001. US$29.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130917621.html.

Mayer:2006:PIW

[MM06a] Johannes Mayer and Robert Merkel, editors. Proceedings of the 1st International Workshop on Random Testing: 2006, Portland, Maine, July 20–20, 2006. ACM Press, New York, NY, USA, 2006. ISBN 1-59593-457-X. LCCN QA76.76.T48 I583 2006. URL http:

//portal.acm.org/toc.cfm?
id=1145735.

McIlwain:2006:TCL

[MM06b] Sean McIlwain and Barton P. Miller. A tool for converting Linux device drivers into Solaris compatible binaries. Software—Practice and Experience, 36(7):689–710, June 2006. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McDougall:2007:SIS

[MM07] Richard McDougall and Jim Mauro. Solaris internals: Solaris 10 and OpenSolaris kernel architecture. Sun Microsystems Press/Prentice Hall, Upper Saddle River, NJ, USA, second edition, 2007. ISBN 0-13-148209-2 (hardback). xlvi + 1020 pp. LCCN QA76.76.O63 M37195 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006015114.html.

Mayer:2006:SEU

[MMC06] Frank Mayer, Karl MacMillan, and David Caplan. SELinux by example: using Security Enhanced Linux. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-196369-4 (paperback). 460 (est.) pp. LCCN QA76.76.O63 M3738 2006. US$44.99. URL http://

vig.prenhall.com/catalog/
academic/product/0,1144,
0131963694,00.html;   http:
//www.loc.gov/catdir/toc/
ecip0612/2006012657.html.

McDougall:2007:SPT

[MMG07] Richard McDougall, Jim Mauro, and Brendan Gregg. Solaris performance and tools: DTrace and MDB techniques for Solaris 10 and OpenSolaris. Sun Microsystems Press/ Prentice Hall, Upper Saddle River, NJ, 2007. ISBN 0-13-156819-1 (hardback). xl + 444 pp. LCCN QA76.76.O63 M3957 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0615/2006020138.html.

Miller:1986:DPM

[MMS86] Barton P. Miller, Cathryn Macrander, and Stuart Sechrest. A distributed programs monitor for Berkeley UNIX. Software—Practice and Experience, 16(2):183–200, February 1986. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mirin:1994:FTD

[MMTK94] A. A. Mirin, M. G. Mc-Coy, G. P. Tomaschke, and J. Killeen. FPPAC94: a twodimensional multispecies nonlinear Fokker–Planck package for UNIX systems. Computer Physics Communications, 81(3):403–408, July 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0010465594900876.

Mitchell:1988:ISC

[MMTW88] Mike Mitchell, Kent Moat, Tom Truscott, and Bob Warren. Invoking system calls from within the UNIX kernel. In USENIX Association [USE88g], pages 277–282. ISBN ???? LCCN ????

Milton:1969:SCP

[MN69] Roy C. Milton and John Ashworth Nelder, editors. Statistical Computation: proceedings of a conference held at the University of Wisconsin, Madison, April 28–30, 1969. Academic Press, New York, NY, USA, 1969. ISBN 0-12-498150-X. LCCN QA276.4 .C6 1969.

Myer:2009:MXU

[MNC09] Thomas Myer, Christopher Negus, and François Caen. Mac OS X UNIX toolbox: 1000+ commands for Mac OS X power users. Wiley, New York, NY, USA, 2009. ISBN 0-470-47836-5. xxiv + 260 pp. LCCN QA76.76.O63.

McKusick:2004:TSF

[MNN04] Marshall Kirk McKusick and George V. Neville-Neil. Thread scheduling in FreeBSD 5.2. ACM Queue: Tomorrow’s Computing Today, 2(7):58– 64, October 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McKusick:2005:DIF

[MNN05] Marshall Kirk McKusick and George V. Neville-Neil. The Design and Implementation of the FreeBSD Operating System. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-201-70245-2. xxviii + 683 pp. LCCN QA76.76.O63 M398745 2005.

URL http://www.mckusick.
com/FreeBSDbook.html.

McKusick:2015:DIF

[MNNW15] Marshall Kirk McKusick, George V.

Neville-Neil, and Robert N. M. Watson. The Design and Implementation of the FreeBSD Operating System. Addison-Wes-ley, Reading, MA, USA, second edition, 2015. ISBN 0-201-70245-2 (hardcover), 0-321-96897-2 (hardcover), 0-321-68005-7, 0-13-376180-0, 0-13-376183-5. xxx + 886 pp. LCCN QA76.774.F74 M35 2015.

Minamoto:2001:SAP

[MO01] Teruya Minamoto and Hiroshi Okumura. Sheru ando paru nyumon: basshu tishisheru gureppu seddo oku paru. Saienssusha, Tokyo, Japan, 2001. ISBN 4-7819-0997-3. 268 pp. LCCN ????

Mohr:1985:GSU

[Moh85] August Mohr. The genesis story: An unofficial, irreverent, incomplete account of how the UNIX operating systems developed. UNIX Review, 3 (1):18–20, 22, 24, 26, 28, 117, …, January 1985. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/Berkeley_Unix_History.
pdf.

Mohr:1998:LUR

[Moh98] James Mohr. Linux user’s resource: developer’s resource. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-842378-

4. xxxv + 795 pp. LCCN QA76.76.O63 M7453 1998.

URL http://www.prenhall.
com/allbooks/ptr_0138423784.
htm.

Mohr:1999:UWS

[Moh99]  James Mohr.  UNIX Web

Server Administrator’s Interactive Workbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-020065-4. xxi + 587 pp. LCCN TK5105.888.M642 1999. US$34.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130200654.html.

Molinelli:1977:UAI

[Mol77] J. J. Molinelli. UNIX assembler for the Intel 8080 microprocessor. Technical Memorandum 1149 (TM 77-3233-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 17, 1977. ?? pp.

Molloy:2014:EBT

[Mol14] Derek Molloy. Exploring BeagleBone: tools and techniques for building with embedded Linux. Wiley, New York, NY, USA, 2014. ISBN 1-118-93512-8 (paperback), 1-118-93513-6 (e-book). xxx + 564 pp. LCCN ????

Montgomery:1995:UGU

[Mon95] John Montgomery. The underground guide to UNIX: slightly askew advice from a UNIX guru. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-40653-5. xiii + 343 pp. LCCN QA76.76.O63 M7454 1995. US$24.95, CDN$31.95.

Mondal:2000:LBU

[Mon00a] Abdul Sakib Mondal. Load balancing for UNIX and Win32. Dr. Dobb’s Journal of Software Tools, 25(7):32, 34, 36, 38–40, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
load.txt;   http://www.ddj.
com/ftp/2000/2000_07/load.
zip;   http://www.ddj.com/
ftp/2000/2000_07/xyalb.zip.

Mondal:2000:PAU

[Mon00b] Abdul Sakib Mondal. Porting across UNIX and Win32. Dr. Dobb’s Journal of Software Tools, 25(12):30, 32, 34, 36, December 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_12/
portuw.txt;   http://www.
ddj.com/ftp/2000/2000_12/
portuw.zip.

Moody:2001:RCL

[Moo01a] Glyn Moody. Rebel code: Linux and the Open Source Revolution. Allen Lane, London, UK and New York, NY, USA, 2001. ISBN 0-7139-9520-3. viii + 334 pp. LCCN QA76.76.O63 M645 2001.

Moody:2001:RCI

[Moo01b] Glyn Moody. The rebel code: the inside story of Linux and the open source revolution. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0333-5. viii + 334 pp.

LCCN QA76.76.O63 M663 2001. US$27.50.

Morris:1975:LRS

[Mor75] Robert Morris. A library of reference standard mathematical subroutines. Technical Memorandum 1074 (TM 75-1271-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1, 1975. ?? pp.

Morgan:1978:USM

[Mor78a] S. P. Morgan. The Unix system:

making computers easier to use. Bell Laboratories Record, 56 (11):308–313, December 1978. CODEN BLRCAB. ISSN 0005-8564.

Morris:1978:CLN

[Mor78b] Robert Morris. Counting large numbers of events in small registers. Communications of the Association for Computing Machinery, 21(10):840–842, October 1978. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morrison:1978:DSL

[Mor78c] J. Paul Morrison. Data stream linkage mechanism. IBM Systems Journal, 17(4):383–408, 1978. CODEN IBMSA7. ISSN 0018-8670. See letters [Lev79, Mor79c].

Morris:1979:FSE

[Mor79a] Robert Morris. A fast software encryption algorithm. Technical Memorandum 1334 (TM 79-1271-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 5, 1979. ?? pp.

Morris:1979:UFS

[Mor79b] Robert Morris. UNIX file security. Technical Memorandum 1335 (TM 79-1271-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 19, 1979. ?? pp.

Morrison:1979:FAR

[Mor79c] J. Paul Morrison. Forum: Author’s response. IBM Systems Journal, 18(3):473–475, 1979. CODEN IBMSA7. ISSN 0018-8670. See [Mor78c, Lev79].

Morris:1981:VU

[Mor81]  Robert Morris.  on UNIX.

TUGboat, 2(3):29, November 1981. ISSN 0896-3207. URL

                                        https://tug.org/TUGboat/
                                        tb02-3/tb04site.pdf.

Morris:1985:WUT

[Mor85] Robert T. Morris. A weakness in the 4.2BSD Unix TCP/IP software. Computing Science Technical Report 117, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 25, 1985. ii + 2 pp.

Morris:1986:ATU

[Mor86] Robert Tappan Morris. Another try at Uucp. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1986. ?? pp.

Morris:1988:AUA

[Mor88] Robert Morris. Adventures in UNIX arithmetic. In USENIX Association [USE88c], pages 157–159. ISBN ???? LCCN ????

MKS:1989:MA

[Mor89] Mortice Kern Systems. MKS AWK. Mortice Kern Systems, Waterloo, ON, Canada, 1989. ISBN 1-895033-01-2. LCCN A76 .73 A95; QA76.73. Version: 3.1, System: DOS 2.0+.

Morin:1996:MUV

[Mor96] Richard Morin. Mach-based UNIX variants. SunExpert Magazine, 7(9):30, 32, 33, September 1996. ISSN 1053-9239. Discusses the University of Utah’s Mach 4 project (see http://www.cs.utah.edu/projects/flux/mach4/html/) the GNU Hurd project (see http://www.cs.pdx.edu/ trent/gnu/hurd/) and the commercial Mach Ten implementation that runs Mach on top of MacOS (see http://www.tenon.com/).

Morin:1997:MML

[Mor97] Rich Morin. MkLinux: Microkernel Linux for the Power Macintosh. Prime Time Freeware, 370 Altair Way, Suite 150, Sunnyvale, CA, USA, 1997. ISBN 1-881957-24-1. 322 pp. LCCN ???? US$50.00. Includes CD-ROM.

Mitchell:2001:ALP

[MOS01] Mark Mitchell, Jeffrey Oldham, and Alex Samuel. Advanced Linux programming. New Riders Publishing, Carmel, IN, USA, 2001. ISBN 0-7357-1043-0. xxiii + 340 pp. LCCN QA76.76.O63 M58

2001.  URL http://www.
advancedlinuxprogramming.
com/.

Meertens:1985:DB

[MP85] L. Meertens and S. Pemberton. Description of B. ACM SIGPLAN Notices, 20(2):58– 76, February 1985. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mui:1992:XWS

[MP92] Linda Mui and Eric Pearce. X Window System Administrator’s Guide for X11 Release 4 and Release 5. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-83-8 (paperback), 1-56592-052-X (with CD ROM). xxiv + 346 pp. LCCN QA76.76.W56 D44 v.8 1992. US$29.95 (without CD ROM), US$59.95 (with CD ROM). URL http:/

/www.oreilly.com/catalog/
9780937175835;   http://
www.oreilly.com/catalog/

v8. Contains CD ROM with X11R4 and X11R5 source code, plus compiled versions for Sun 3 SunOS 4.1.1, Sun 4 SunOS 4.1.1, DECstation ULTRIX 4.2, and IBM RS/6000 AIX 3.2.

Malkin:1993:IUG

[MP93] Gary Malkin and Tracy LaQuey Parker. Internet Users’ Glossary, January 1993. A glossary that concentrates on terms that are special to the Internet. This document (rfc1392.txt) can be obtained by anonymous ftp

from /rfc at nic.ddn.mil.

McIlroy:1978:UTS

[MPT78] M. Douglas McIlroy, E. N.

Pinson, and B. A. Tague. UNIX time-sharing system: Forward. The Bell System Technical Journal, 57(6): 1899–1904, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1899.
pdf.

McKusick:2009:CSG

[MQ09] Marshall Kirk McKusick and Sean Quinlan. Case study: GFS: Evolution on fastforward. ACM Queue: Tomorrow’s Computing Today, 7 (7):10, August 2009. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

McKusick:2010:GEF

[MQ10] Kirk McKusick and Sean Quinlan. GFS: evolution on fastforward. Communications of the Association for Computing Machinery, 53(3):42–49, March 2010. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Meyer:1967:CLP

[MR67] Albert R. Meyer and Dennis M. Ritchie. The complexity of loop programs. In Solomon Rosenthal, editor, ACM ’67: Proceedings of the 1967 22nd national conference, pages 465– 569. ACM Press, New York, NY, USA, January 1967. ISBN 1-4503-7494-8. URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/800196.806014. Cited in

[Bro20].

McIlroy:1992:MSU

[MR92] M. Douglas McIlroy and James A. Reeds. Multilevel security in the UNIX tradition. Software—Practice and Experience, 22(8):673–694, August 1992. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

McIlroy:19xx:TI

[MRxxa] Doug McIlroy and Jim Reeds. A tour of IX. Computing Science Technical Report 163e, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 11 pp.

McIlroy:19xx:DI

[MRxxb] M. D. McIlroy and J. A. Reeds. The design of IX. Computing Science Technical Report 163d, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 12 pp.

McIlroy:19xx:MSU

[MRxxc] M. D. McIlroy and J. A. Reeds. Multilevel security in the UNIX tradition. Computing Science Technical Report 163c, AT&T Bell Laboratories, Murray Hill, NJ, USA, ?? ??, 19xx. 19 pp.

McIlroy:19xx:MWS

[MRxxd] M. D. McIlroy and J. A. Reeds.

Multilevel windows on a singlelevel terminal. Computing Science Technical Report 163f, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 3 pp.

McKusick:2010:JSU

[MR10] Marshall Kirk McKusick and J. Roberson. Journaled softupdates. In ????, editor, Proceedings of the 2010 European BSD Conference, Karlsruhe, Germany, October 2010, page ?? ????, ????, 2010.

McKusick:2014:FUS

[MR14a] Marshall Kirk McKusick and J. Roberson. The FreeBSD ULE scheduler. The FreeBSD Journal, 1(5):20–26, September/October 2014. ISBN 0-615-88479-2. URL http:/

/issue.freebsdfoundation.
org/publication/?m=33057&
i=230897&p=22.

McKusick:2014:JSU

[MR14b] Marshall Kirk McKusick and J. Roberson. Journaled soft-updates. The FreeBSD Journal, 1(2):30–36, March/ April 2014. ISBN 0-615-88479-2. URL http://

issue.freebsdfoundation.

org/publication/?m=33057&
i=217640&p=32.

McKusick:2017:EFG

[MR17] Marshall Kirk McKusick and Benno Rice. The evolution of FreeBSD governance. The FreeBSD Journal, 4(5):12–23, September/ October 2017. ISBN 0-615-88479-2. URL http:/

/issue.freebsdfoundation.
org/publication/?m=33057&
i=447089&p=28.

Markas:1990:DFS

[MRK90] Tassos Markas, Mark Royals, and Nick Kanopoulos. On distributed fault simulation. Computer, 23(1):40–52, January 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marshall:2006:ASV

[MRM06] David (David W.) Marshall, Wade A. Reynolds, and Dave McCrory. Advanced server virtualization: VMware and Microsoft platforms in the virtual data center. Taylor and Francis, Boca Raton, FL, USA, 2006. ISBN 0-8493-3931-6. ???? pp. LCCN QA76.76.O63 M3646 2006.

Mashey:1977:PMP

[MS77] John R. Mashey and D. W. Smith. PWB/MM — Programmer’s Workbench Memorandum Macros (T.2). Technical Memorandum 1098 (TM 77-9144-5, TM 77-8234-23), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1977. ?? pp.

Morgan:1984:SUF

[MS84] C. Morgan and B. Sufrin.

Specification of the UNIX filing system. IEEE Transactions on Software Engineering, SE-10(2):128–142, March/April 1984. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

URL  http://ieeexplore.
ieee.org/stamp/stamp.jsp?
arnumber=5010215.

McJones:1987:EUS

[MS87] Paul R. McJones and Garret Frederick Swart. Evolving the UNIX system interface to support multithreaded programs: The Topaz Operating System programmer’s manual, volume 21 of Systems Research Center. Digital Systems Research Center, Palo Alto, CA, USA, September 28, 1987. 100 pp. LCCN QA76.76.O63M42 1987.

Mandelberg:1988:PMU

[MS88] K. I. Mandelberg and V. S.

Sunderam. Process migration in UNIX networks. In USENIX Association [USE88g], pages 357–363. ISBN ???? LCCN ????

McJones:1989:EUS

[MS89] Paul R. McJones and Garret F.

Swart. Evolving the UNIX system interface to support multithreaded programs. In USENIX Association [USE89f], pages 393–404.

Murakami:1990:TCP

[MS90] Gary J. Murakami and Ravi Sethi. Terminal call processing in Esterel. Computing Science Technical Report 150, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 15, 1990. 10 pp.

McClennan:1992:IMS

[MS92] Michael McClennan and Stuart Sechrest. Introducing multistructured file naming into UNIX. In USENIX [USE92c], pages 151–152. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Mullet:1995:DVI

[MS95] Kevin Mullet and Darrell Sano. Designing visual interfaces: communication oriented techniques. SunSoft Press, Englewood Cliffs, NJ, USA, 1995. ISBN 0-13-303389-9. xv + 273 pp. LCCN QA76.9.U83 M84

                                        1995. URL http://www.sun.
                                        com/books/catalog/mullet/.

Matthew:2000:PLP

[MS^{+}00] Neil Matthew, Richard Stones, et al. Professional Linux programming. Wrox Press, Chicago, IL, USA, 2000. ISBN 1-86100-301-3. xviii + 1155 pp. LCCN QA76.76.O63 P754 2000.

Meadows:2005:CHE

[MS05] Catherine Meadows and Paul Syverson, editors. CCS ’05: proceedings of the 12th ACM Conference on Computer and Communications Security: November 7–11, 2005, Alexandria, Virginia, USA. ACM Press, New York, NY, USA, 2005. ISBN 1-59593-226-7. LCCN QA76.9.A25. ACM order number 459050.

Maillart:2008:ETZ

[MSSvK08] T. Maillart, D. Sornette, S. Spaeth, and G. von Krogh. Empirical tests of Zipf’s Law mechanism in open source Linux distribution. Physical Review Letters, 101(21): 218701, November 19, 2008. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/
10.1103/PhysRevLett.101.
218701.

Mendoza:2006:ULP

[MSW06] Alfredo Mendoza, Chakarat Skawratananond, and Artis Walker. Unix to Linux porting: a comprehensive reference. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-187109-9 (paperback). ???? pp. LCCN QA76.76.O63 M437 2006. URL http:

//www.loc.gov/catdir/toc/
ecip066/2006000234.html.

Morris:1974:WSH

[MT74] Robert Morris and Ken Thompson. Webster’s second on the head of a pin. Technical Memorandum 1082 (TM 74-1271-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 15, 1974. ?? pp.

Morris:1978:PSC

[MT78] Robert Morris and Ken Thompson. Password security: A case history. Technical Memorandum 1259 (TM 78-1271-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 3, 1978. ?? pp.

Morris:1979:PSC

[MT79] Robert Morris and Ken Thompson. Password security — a case history. Communications of the Association for Computing Machinery, 22(11):594– 597, November 1979. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Morrisett:1993:PLP

[MT93] J. Gregory Morrisett and Andrew P. Tolmach. Procs and locks: a portable multiprocessing platform for Standard ML of New Jersey. ACM SIG-PLAN Notices, 28(7):198–207, July 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McIntosh:2003:MXN

[MTS03] Jason McIntosh, Chuck Toporek, and Chris Stone. Mac OS X in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00370-6. xxii + 801 pp. LCCN QA76.76.O63 M38 2003.

Murai:1983:KSU

[MTT83] Jun Murai, Mario Tokoro, and Fumio Teraoka. KEIO S&Tnet: a UNIX campus network. In IEEE [IEE83], pages 14–23. CODEN CLCPDN. ISBN 0-8186-0500-6 (paperback), 0-8186-8500-X. LCCN TK 5105.5 C66 1983. IEEE Service Cent. Piscataway, NJ, USA.

Masui:2001:DUS

[MTY01] Kenji Masui, Masahiko Tomoishi, and Naoki Yonezaki. Design of UNIX system for the prevention of damage propagation by intrusion and its implementation based on 4.4BSD. Lecture Notes in Computer Science, 2200:536– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2200/22000536.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2200/22000536.
pdf.

Mudawwar:1997:MTM

[Mud97] Muhammed F. Mudawwar.

Multicode: a truly multilingual approach to text encoding: Unicode was designed to extend ASCII for encoding text in different languages, but it still have several important drawbacks. multicode overcomes those drawbacks. Computer, 30(4):37–43, April 1997. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). See also response and rebuttal [DM97] and letter [CPJ^{+}98].

Mueller:1978:NDD

[Mue78a] M. B. Mueller. A new DV-11 driver for use with LEAP. Technical Memorandum 1312 (TM 78-3112-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 22, 1978. ?? pp.

Muenzer:1978:MTU

[Mue78b] T. B. Muenzer. Monitor terminals under the UNIX operating system. Technical Memorandum 1332 (TM 78-9442-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 24, 1978. ?? pp.

Mueller:1979:UBS

[Mue79] M. B. Mueller. A UNIX based system for automated testing of asynchronous terminals and hosts. Technical Memorandum 1489 (TM 79-3162-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 10, 1979. ?? pp.

Mueller:1993:LIP

[Mue93] Frank Mueller. A library implementation of POSIX threads under UNIX. In USENIX Association [USE93f], pages 29–41. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

Mueller:1994:UUW

[Mue94] Scott Hazen Mueller. Understanding UNIX workstation performance. ;login: the USENIX Association newsletter, 19(5):32–34, October 1994. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Mui:1995:WYC

[Mui95] Linda Mui. When You Can’t Find Your UNIX System Administrator. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, April 1995. ISBN 1-56592-104-6. xiii + 139 pp. LCCN QA76.76.O63M845 1995. US$17.95.

Mulrooney:1979:ISS

[Mul79a] T. J. Mulrooney. INTEL 8048 Simulator (SIM48) user manual. Technical Memorandum 1445 (TM 79-2211-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1979. ?? pp.

Mulrooney:1979:ICC

[Mul79b] T. J. Mulrooney. An IN-TEL compatible 8048-49 cross assembler. Technical Memorandum 1383 (TM 79-2211-10), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 29, 1979. ?? pp.

Mulrooney:1979:UIB

[Mul79c] T. J. Mulrooney. An (UNIX) interface between a stand alone operating system (ISIS) and a time shared computer system. Technical Memorandum 1433 (TM 79-2211-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 15, 1979. ?? pp.

Mullen:1987:UNS

[Mul87] John R. Mullen. UNIX and networking: a separate peace. In USENIX Association [USE87e], pages 21–?? LCCN QA 76.76 O63 U84 1987. Abstract only.

Murphy:1983:CU

[Mur83] Timothy Murphy. Circular UNIX. EUUG Newsletter, 3(3): 28–32, Autumn 1983. CODEN EONLE8. ISSN 1011-4211.

Muster:1996:UME

[Mus96] John Muster. UNIX made easy. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 1996. ISBN 0-07-882173-8. xviii + 1061 pp. LCCN QA76.76.O63 U545 1996.

Muster:2002:UME

[Mus02] John Muster. UNIX made easy: Unix and Linux basics and beyond. Osborne/McGraw-Hill, Berkeley, CA, USA, third edition, 2002. ISBN 0-07-219314-X. xviii + 1011 pp. LCCN QA76.76.O63 M877 2002.

Murray:1994:EGF

[Mv94] James D. Murray and William vanRyper. Encyclopedia of Graphics File Formats. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1994. ISBN 1-56592-058-9. xxxii + 894 pp. LCCN T385 .M87 1994. US$59.95.

Mullender:1990:ADO

[MvT^{+}90] Sape J. Mullender, Guido van Rossum, Andrew S. Tanenbaum, Robbert van Renesse, and Hans van Staveren. Amoeba: a distributed operating system for the 1990s. Computer, 23(5):44–53, May 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKenney:2008:ITL

[MW08] Paul E. McKenney and Jonathan Walpole. Introducing technology into the Linux kernel: a case study. Operating Systems Review, 42(5):4–17, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Manheimer:1990:DFS

[MWCR90] Kenneth Manheimer, Barry A.

Warsaw, Stephen N. Clark, and Walter Rowe. The depot: a framework for sharing software installation across organizational and UNIX platform boundaries. In USENIX [USE90b], pages 37–46. LCCN QA76.76.O63 L37 1990.

Manheimer:1991:DFS

[MWCR91] Ken Manheimer, Barry Warsaw, Steve Clark, and Walter Rowe. The depot: a framework for sharing software installation across organizational and UNIX platform boundaries. In USENIX [USE91a], page ?? ISBN ???? LCCN ????

Meadhra:2000:KLD

[MWM00] Michael Meadhra, Kate Wrightson, and Joe Merlino. KDE for Linux for dummies. I D G Books Worldwide, Indianapolis, IN, USA, 2000. ISBN 0-7645-0658-7. xxii, 344 pp. LCCN QA76.9.U83 M46 2000.

Minasi:2003:LWA

[MY03] Mark Minasi and Dan York. Linux for Windows administrators. Mark Minasi Windows administrator library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4119-6. xxxiii + 522 pp. LCCN QA76.76.O63 M57385

                                        2003. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0613/2002113840-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0613/2002113840-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0613/2002113840-t.html.

Minasi:2000:LWN

[MYH00] Mark Minasi, Dan York, and Craig Hunt. Linux for Windows NT/2000 Administrators: The Secret Decoder Ring. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN ???? ???? pp. LCCN ???? URL http://

www.sybex.com/cgi-bin/rd_
bookpg.pl?2730back.html.

Murphy:1996:APM

[MZA96] B. J. Murphy, S. Zeadally, and C. J. Adams. An analysis of process and memory models to support highspeed networking in a UNIX environment. In USENIX [USE96b], pages 239–251 (or 239–252??). ISBN 1-880446-76-6. LCCN QA 76.76 O63

U88 1996. URL http://www.
usenix.org/publications/
library/proceedings/sd96/
murphy.html.

Mihajlovic:2014:DIQ

[MZG14] Bojan Mihajlović, Zeljko Zilić, and Warren J. Gross. Dynamically instrumenting the QEMU emulator for Linux process trace generation with the GDB debugger. ACM Transactions on Embedded Computing Systems, 13(5s):167:1– 167:??, November 2014. CO-DEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Miller:2022:RCF

[MZH22] Barton P. Miller, Mengxiao Zhang, and Elisa R. Heymann. The relevance of classic fuzz testing: Have we solved this one? IEEE Transactions on Software Engineering, 48(6):2028–2039, June 2022. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL https://

                                        arxiv.org/abs/2008.06537;
                                         https://ieeexplore.ieee.
                                        org/document/9309406.

Nadkami:1992:PFS

[Nad92] Ashok V. Nadkami. The “processor” file system in UNIX SVR4.2. In USENIX [USE92c], pages 131–132. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

Nadelson:1998:RTE

[Nad98] Mark Nadelson. Real-time error processing on a Unix network. C/C++ Users Journal, 16(3):??, March 1998. CODEN CCUJEX. ISSN 1075-2838.

Naecker:1990:PP

[Nae90] Philip A. Naecker. POSIX and portability. The DEC Professional, 9(6):46–??, June 1990. CODEN DECPDJ. ISSN 0744-9216.

Nagle:2005:BRH

[Nag05] Dan Nagle. Book review: High Performance Linux Clusters, by A. Joseph and D. Sloan. Scientific Programming, 13(2):173–175, ???? 2005. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nassal:1997:STN

[Nas97] Mark Nassal. Samba — tuning the NT/UNIX dance. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):

27–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

NIST:1988:PPO

[Nat88] National Institute of Standards and Technology (U. S.). POSIX: portable operating system interface for computer environments. Washington, DC, USA, September 12, 1988. 7 pp. Shipping list no.: 88-752-P.

NIST:1990:XWS

[Nat90] National Institute of Standards and Technology (U. S.). X Window System: version 11, release 3. Gaithersburg, MD, USA, 1990. various pp. Shipping list no.: 90-0657-P. Contents: X Window System protocol / Robert W. Scheifler – Xlib-C language X interface / James Gettys, Robert W. Scheifler, Ron Newman – X toolkit intrinsics-C language interface / Joel McCormack, Paul Asente, Ralph R. Swick – Bitmap distribution format 2.1.

NIST:1994:XWS

[Nat94] National Institute of Standards and Technology (U. S.). X Window System: version 11, release 5. Gaithersburg, MD, USA, 1994. various pp. Shipping list no.: 95-0136-P. Contents: X Window System protocol / Robert W. Scheifler – Xlib-C language X interface / James Gettys, Robert W. Scheifler – X toolkit intrinsics-C language interface / Joel McCormack, Paul Asente, Ralph R. Swick – Bitmap distribution format, version 2.1.

Nedginn:1984:CAP

[NB84] Preet J. Nedginn and Trebor L. Bworn. CLOG: An Ada package for automatic footnote generation in UNIX (April 1984 special section). Communications of the Association for Computing Machinery, 27 (4):351–??, 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Noordergraaf:2002:ESS

[NBB^{+}02] Alex Noordergraaf, Tony M. Benson, Glenn Brunette, Vasanthan Dasan, Mark Hashimoto, Dina Kurktchi, Richard Lau, Lou Ordorica, Will Osser, and Keith Watson. Enterprise Security: Solaris Operating Environment Security Journal, Solaris OE v2.5.1, 2.6, 7, and 8. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-100092-6. xxx + 416 pp. LCCN QA76.9.A25N66

                                        2002. URL books/816-5040-
                                        10.pdf;   http://www.sun.
                                        com/books/catalog/noord2/

. Part No. 816-5040-10 June 2002, Revision 01.

Negus:2008:BUT

[NC08] Christopher Negus and François Caen. BSD UNIX Toolbox: 1000+ Commands for FreeBSD, OpenBSD and NetBSD. Wiley, New York, NY, USA, 2008. ISBN 0-470-37603-1. xxvi + 309 pp. LCCN QA76.76.O63 N4166 2008. US$24.99.

URL http://www.loc.gov/
catdir/enhancements/fy0829/
2008016821-b.html;   http:
//www.loc.gov/catdir/enhancements/
fy0829/2008016821-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0829/2008016821-
t.html.

Nievergelt:1982:DPS

[NCNS82] J. Nievergelt, G. Coray, J.-D.

Nicoud, and A. C. Shaw, editors. Document Preparation Systems: a Collection of Survey Articles. Elsevier North-Holland, Inc., New York, NY, USA, 1982. ISBN 0-444-86493-

8. xiv + 274 pp. LCCN Z244 .D63 1982. US$46.50.

Neider:1993:OPG

[NDW93] Jackie Neider, Tom Davis, and Mason Woo. OpenGL Programming Guide — The Official Guide to Learning OpenGL, Release 1. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-63274-8. xiii + 516 pp. LCCN T385.N435 1993. US$34.95. The source code examples are available

at ftp://sgigate.sgi.com/
pub/opengl/opengl.tar.Z. An

HTML version of the book is

available at http://arctic.
eng.iastate.edu:88/SGI_Developer/
OpenGL_PG/ and http://fly.
cc.fer.hr/~unreal/index.
html.

Neely:1990:SDV

[Nee90] Richard Neely. System design and verification for secure applications under UNIX. In USENIX Association [USE90e], pages 125–128. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Nee:1991:GYP

[Nee91] Eric Nee. Get Your PC Software on UNIX. UNIX/World, 8 (10):54–??, October 1991. ISSN 0739-5922.

Negus:2001:RLB

[Neg01] Christopher Negus. Red Hat Linux 7 Bible. IDG Books, San Mateo, CA, USA, 2001. ISBN ???? ???? pp. LCCN ????

Negus:2002:RLB

[Neg02] Chris Negus. Red Hat Linux 8 bible. Wiley, New York, NY, USA, 2002. ISBN 0-7645-4968-5. xviii + 1063 pp. LCCN QA76.73.O63 N44 2002.

Negus:2006:LBB

[Neg06a] Chris Negus. Linux bible: boot up to Fedora, KNOPPIX, Debian, SUSE, Ubuntu, and 7 other distributions. Wiley, New York, NY, USA, 2006 edition, 2006. ISBN 0-471-75489-7 (paper/DVD). xxix + 870 pp. LCCN QA76.76.O63 N42143 2006.

Negus:2006:LTI

[Neg06b] Chris Negus. Linux toys II: 9 cool new projects for home, office, and entertainment. ExtremeTech. Wiley, New York, NY, USA, 2006. ISBN 0-7645-7995-9 (paper/CD-ROM + e-book). xxii + 397 pp. LCCN QA76.76.O63 N4233

2006. URL http://www.loc.
gov/catdir/enhancements/
fy0623/2005025375-d.html;
 http://www.loc.gov/catdir/
enhancements/fy0662/2005025375-
b.html;   http://www.loc.
gov/catdir/toc/ecip0518/
2005025375.html.

Negus:2006:LLC

[Neg06c]  Chris Negus.  Live Linux

CDs: building and customizing bootables. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-243274-9 (paperback). 448 (est.) pp. LCCN QA76.76.O63 N4245 2006. URL http:

//www.loc.gov/catdir/toc/
ecip0619/2006027573.html.

Negus:2007:LBB

[Neg07a] Chris Negus. Linux bible: boot up Ubuntu, Fedora, KNOP-PIX, Debian, SUSE, and 11 other distributions. Wiley, New York, NY, USA, 2007 edition, 2007. ISBN 0-470-08279-8 (paper/DVD). xxxi + 848 pp. LCCN QA76.76.O63 N422 2007. URL http:

//www.loc.gov/catdir/toc/
ecip078/2007000458.html.

Negus:2007:LLC

[Neg07b] Chris Negus. Live Linux CDs:

building and customizing bootables. Negus live Linux series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2007. ISBN 0-13-243274-9 (paperback). xix + 430 pp. LCCN QA76.76.O63 N4245 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0619/2006027573.html.

Nelson:1979:MIC

[Nel79] N-P. Nelson. Minutes, ISCC Common Command Subcommittee. Technical Memorandum 1340, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 25, 1979. ?? pp.

Nelson:1990:TCP

[Nel90] Phyllis Nelson. C Programming In A UNIX Environment, by Judy Kay and Bob Kummerfeld. C Users Journal, 8 (8):119–??, August 1990. ISSN 0898-9788. See [KK89a].

Neugebauer:2000:ULP

[Neu00] Rolf Neugebauer. A Unix-like personality supporting qualityof-service. Operating Systems Review, 34(2):39, April 2000. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Neuner:2006:ILS

[Neu06]  Steve Neuner.  An inside

look at scaling Linux to 1024 processors. In Anonymous [Ano06a], page ?? ISBN ???? LCCN ???? URL http://

                                        www.ice.gelato.org/oct06/
                                        pres_pdf/gelato_ICE06oct_
                                        scaling1024_neuner_sgi.pdf.

Newmarch:1992:XWS

[New92] Jan Newmarch. The X Window System and Motif — a Fast Track Approach. Addison-Wes-ley, Reading, MA, USA, 1992. ISBN 0-201-53931-4. xiv + 221 pp. LCCN QA76.76.W56 N59 1992.

Neyer:1983:UTM

[Ney83a] James A. Neyer. UNIX timesharing menu-driven office system for terminals (UTMOST). In Association [Ass83], pages 116–?? Abstract only.

Neyer:1983:UTS

[Ney83b] James A. Neyer. UNIX timesharing menu-driven office system for terminals (UTMOST). In USENIX [USE83], pages 116–?? Abstract only.

Ng:1977:MGP

[Ng77] S. W. Ng. MICA — a general purpose microassembler: Functional specification/user’s manual — issue II. Technical Memorandum 1220 (TM 77-5514-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 15, 1977. ?? pp.

Nguyen:2008:GG

[Ngu08a] Hubert Nguyen, editor. GPU gems 3, volume 3 of GPU gems. Addison-Wesley, Reading, MA, USA, 2008. ISBN 0-321-51526-9. l + 942 pp. LCCN T385 .G6882 2008. URL http:

//www.loc.gov/catdir/toc/
ecip0720/2007023985.html.

Nguyen:2008:SLC

[Ngu08b] Quan Nguyen, editor. Scripting languages: a collection of Perl, Ruby, Python, TCL and Unix. Ramacad, San Jose, CA, USA, 2008. ISBN 0-9777812-3-2. ???? pp. LCCN ????

Norton:1991:PNG

[NH91] Peter Norton and Harley Hahn. Peter Norton’s Guide to UNIX. Bantam Books, New York, NY, USA, 1991. ISBN 0-553-35260-1. xxiv + 560 pp. LCCN QA76.76.O63 N6817 1991. US$26.95. One of the many books by Peter Norton.

Nadelson:1999:MUW

[NH99] Mark Nadelson and Thomas Hagan. Making UNIX and Windows NT Talk: Object-Oriented Inter-Platform Communication. R&D Books, Lawrence, KS, USA; Berkeley, CA, USA, 1999. ISBN 0-87930-584-3. 512 pp. LCCN ???? US$44.95. URL http:/

/www.rdbooks.com/scripts/
store/vsc/store/products/
rd3104.htm?L+/htdocs/rdbooks/
config/store+dwie8022. In-

cludes CD-ROM.

Nadelson:2000:COM

[NH00] Mark Nadelson and Thomas G. Hagan. C++ Objects for Making UNIX and WinNT Talk. CMP Books, Lawrence, KS, USA, 2000. ISBN 1-929629-07-9. xii + 558 pp. LCCN QA76.76.O63 N33 2000. US$49.95.

Nieh:1994:SUS

[NHNW94] J. Nieh, J. G. Hanko, J. D. Northcutt, and G. A. Wall.

SVR4 UNIX scheduler unacceptable for multimedia applications. Lecture Notes in Computer Science, 846:41–??, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowitz:1984:EIU

[NHR84] D. A. Nowitz, P. Honeyman, and B. Redman. Experimental implementation of UUCP: Security aspects. In ????, editor, UniForum Conference, Washington 1984 Proceedings, page ?? ????, ????, January 1984.

Ni:1997:ETT

[Ni97] David Chi-Liang Ni. Enumeration and traceability tools for UNIX^{T M} and WINDOWS^{T M} environments. The Journal of Systems and Software, 39(1): 15–25, October 1997. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

NIST:1992:NCS

[NIS92] NIST, editor. 15th National Computer Security Conference: October 13–16, 1992, Baltimore Convention Center, Baltimore, MD: information systems security, building blocks to the future. National Institute for Standards and Technology, Gaithersburg, MD, USA, 1992. ISBN ???? LCCN QA76.9.A25 N38 1992. Two volumes.

Nisley:2001:ESP

[Nis01a] Ed Nisley. Embedded space: Penguin specs. Dr. Dobb’s Journal of Software Tools, 26 (6):139–141, June 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

Nisley:2001:ESRc

[Nis01b] Ed Nisley. Embedded space: Rating real time: Count the ways. Dr. Dobb’s Journal of Software Tools, 26(9):113– 115, September 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/.

Nisley:2001:ESS

[Nis01c] Ed Nisley. Embedded space: SEU meets Embedded Linux. Dr. Dobb’s Journal of Software Tools, 26(3):129–131, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/.

Nisley:2003:ESRb

[Nis03] Ed Nisley. Embedded space: Real-Time Linux. Dr. Dobb’s Journal of Software Tools, 28 (4):79–81, April 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7827/
                                        ddj0304j/.

Nowitz:1979:DNU

[NL79] D. A. Nowitz and Michael E. Lesk. A dial-up network of UNIX systems. Technical Memorandum 1437 (TM 79-3782-2, TM 79-1274-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1, 1979. ?? pp.

Nowitz:1982:IUN

[NL82] D. A. Nowitz and M. E. Lesk.

Implementation of a UNIX network. Computer Communications, 5(1):30–34, February 1982. CODEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

Nieh:2000:EV

[NL00] Jason Nieh and Ozgur Can Leonard. Examining VMware. Dr. Dobb’s Journal of Software Tools, 25(8):70, 72–74, 76, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Nielsen:1984:EOB

[NLR84a] Erik Reeh Nielsen, Soren Lauesen, and Vilhelm Rosenqvist. An expandable object-based UNIX kernel. In USENIX [USE84a], pages 193–202.

Nielsen:1984:EOU

[NLR84b] Erik Reeh Nielsen, Soren Lauesen, and Vilhelm Rosenqvist. An expandable object-based UNIX kernel. In Software Tools Users Group [Sof84], pages 193–202. LCCN QA76.8.U65 U83 1984.

Nakamura:1983:LMS

[NM83a] Osamu Nakamura and Jun Murai. On-line manual system for software development on UNIX. In Association [Ass83], pages 19–30.

Nakamura:1983:OMS

[NM83b] Osamu Nakamura and Jun Murai. On-line manual system for software development on UNIX. In Software Tools Users Group [Sof83], pages 19– 30. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Northlich:1982:EUP

[NMP82] William R. Northlich, Jr., T. D. McCreery, and P. M. Powers. Embedding UNIX in a product (or, is ‘real-time’ real?). In Usr Group [Usr82], pages 1–14.

Neelands:1983:UN

[NMS83] Paul Neelands, Richard Miller, and Chris Sturgess. UNIX for the National 16032. In Association [Ass83], pages 269–272. Extended abstract.

Neville-Neil:1992:BRD

[NN92] George Neville-Neil. Book review: The Design and Implementation of the 4.3 BSD UNIX OS Answer Book. ;login: the USENIX Association newsletter, 17(2):35–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Nye:1990:XTIa

[NO90a] Adrian Nye and Tim O’Reilly. X Toolkit Intrinsics Programming Manual, volume 4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-34-X. xxxi + 543 pp.

Nye:1990:XTIb

[NO90b] Adrian Nye and Tim O’Reilly.

X Toolkit Intrinsics Reference Manual: OSF/Motif 1.1 Edition for X11, Release 4, volume 4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, December 1990. ISBN 0-937175-62-5 (vol. 4), 0-937175-66-8 (set). xxxi + 632 pp. LCCN QA76.9.W56N94 1990.

URL http://www.oreilly.
com/catalog/9780937175620.

Nye:1992:XTI

[NO92] Adrian Nye and Tim O’Reilly.

X Toolkit Intrinsics Programming Manual: Motif Edition, volume 4M. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, August 1992. ISBN 1-56592-013-9. 674 pp.

Nye:1993:XTI

[NO93] Adrian Nye and Tim O’Reilly.

X Toolkit Intrinsics Programming Manual, volume 4. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, April 1993. ISBN 1-56592-003-1. 567 pp.

Noer:1998:CFW

[Noe98]  Geoffrey J. Noer.  Cyg-

win32: a free Win32 porting layer for UNIX(R). applications. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

                                        usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/noer.html;   http://
                                        www.usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/noer_slides.

Norden:1979:VCU

[Nor79] R. A. Norden. VM/CMS– UNIX file transfer facility. Technical Memorandum 1339, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 10, 1979. ?? pp.

Norman:1981:TAU

[Nor81] Donald Arthur Norman. The truth about UNIX. Datamation, 27(12):139–150, November 1981. CODEN DTM-NAT. ISSN 0011-6963. URL

                                        https://www.researchgate.
                                        net/profile/Donald_Norman/
                                        publication/202165676_The_
                                        trouble_with_UNIX_The_user_
                                        interface_is_horrid/links/
                                        54a2b6090cf267bdb9042331/
                                        The-trouble-with-UNIX-The-
                                        user-interface-is-horrid.
                                        pdf.

Norwood:1991:NUO

[Nor91] Peter Norwood. A NonStop UNIX operating system. In USENIX [USE91d], pages 325– 334. LCCN QA 76.76 O63 U84 1992.

Norusis:1993:SUA

[Nor93a] Marija J. Norušis. SPSS for UNIX: Advanced Statistics, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-88-5. xiv + 574 pp. LCCN HA32 .N65166 1993.

Norusis:1993:SUB

[Nor93b]  Marija J. Norušis.  SPSS

for UNIX: Base System User’s Guide, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-86-9. xvii + 700 pp. LCCN HA32 .N65167 1993.

Norusis:1993:SUP

[Nor93c] Marija J. Norušis. SPSS for UNIX: Professional Statistics, Release 5.0. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1993. ISBN 0-923967-87-7. xi + 345 pp. LCCN ??

Northrup:1996:PUT

[Nor96] Charles J. Northrup. Programming with UNIX Threads. Wiley, New York, NY, USA, 1996. ISBN 0-471-13751-0 (paperback). xv + 399 pp. LCCN QA76.76.O63 N674 1996.

Northcutt:1999:NID

[Nor99a] Stephen Northcutt. Network Intrusion Detection: An Analysts’ Handbook. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7357-0868-1. 267 pp. LCCN TK5105.59 .N475 1999. US$45.00.

Norton:1999:CGL

[Nor99b] Peter Norton. Peter Norton’s Complete guide to Linux. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-672-31573-4. xv + 581 pp. LCCN QA76.76.O63 N67792 2000. Edited by Bryan Gambrel.

Norton:2002:ISP

[Nor02]  Roger Norton.  IT sys-

tems perspective: Using virtual Linux servers. Computer, 35(11):106–107, November 2002. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://csdl.computer.org/
dl/mags/co/2002/11/ry106.
htm;   http://csdl.computer.
org/dl/mags/co/2002/11/ry106.
pdf.

Nowitz:1978:UID

[Now78] D. A. Nowitz. Uucp — implementation description. Technical Memorandum 1293, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1978. ?? pp.

Nowitz:1990:UA

[Now90] D. A. Nowitz. Uucp administration. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483 1990. URL http:

//doc.cat-v.org/unix/v10/
10thEdMan/v2contents.html.

Two volumes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Nagelberg:1978:UTS

[NP78] E. R. Nagelberg and M. A.

Pilla. UNIX time-sharing system: RBCS/RCMAS — converting to the MERT operating system. The Bell System Technical Journal, 57(6): 2275–2287, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2275.
pdf.

Nelson:1994:UNS

[NR94] Michael N. Nelson and Sanjay R. Radia. A uniform name service for Spring’s UNIX environment. In USENIX [USE94f], pages 201–209. ISBN 1-880446-58-8. LCCN QA 76.76 O63

U84 1994. URL http://www.
usenix.org/publications/
library/proceedings/sf94/
nelson.html.

Newham:1995:LBS

[NR95] Cameron Newham and Bill Rosenblatt. Learning the bash Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1995. ISBN 1-56592-147-X. xvi + 292 pp. LCCN QA76.76.O63N458 1995. US$27.95. URL http:/

                                        /www.ora.com/gnn/bus/ora/
                                        item/bash.html.

Newham:1998:LTB

[NR98] Cameron Newham and Bill Rosenblatt. Learning the bash Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second (revised and updated) edition, 1998. ISBN 1-56592-347-2. xvi + 318 pp. LCCN QA76.76.O63N458 1998. US$29.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        bash2.

Narduzzo:2003:MAG

[NR03] A. Narduzzo and A. Rossi. Modularity in action: GNU/ Linux and Free/Open Source software development model unleashed. World-Wide Web document., May 12, 2003.

                                        URL  http://opensource.
                                        mit.edu/papers/narduzzorossi.
                                        pdf.

Newham:2005:LBS

[NR05] Cameron Newham and Bill Rosenblatt. Learning the bash shell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2005. ISBN 0-596-00965-8 (paperback). xvi + 333 pp. LCCN QA76.76.O63 N458 2005. UK£24.95.

Negus:1992:GUD

[NS92] Chris Negus and Larry Schumer. Guide to the UNIX Desktop.

UNIX Press, Summit, NJ, USA, 1992. ISBN 1-56205-114-8. xxiv + 709 pp. LCCN QA76.76.O63 N42 1992. A fine tutorial/reference text on SVR4.2.

Negus:1994:NGU

[NS94a] Chris Negus and Larry Schumer.

Novell’s Guide to UnixWare 1.1. Novell Press, Sybex, San Jose, CA, USA, 1994. ISBN 0-7821-1292-7. xliii + 817 pp. LCCN QA76.76.O63 N425 1994. This is a comprehensive reference to UnixWare 1.1.

Norusis:1994:SBS

[NS94b] Marija J. Norušis and SPSS, Inc. SPSS 6.1 Base System User’s Guide, Part 2: UNIX Version. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1994. ISBN 0-13-438870-4. xiii + 547 pp.

Nemeth:2002:LAH

[NSH02] Evi Nemeth, Garth Snyder, and Trent Hein. Linux Administration Handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-008466-2. xxi + 890 pp. LCCN QA76.76.O63 N448 2002. US$49.99.

Nemeth:2006:LAH

[NSH06] Evi Nemeth, Garth Snyder, and Trent R. Hein. Linux administration handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2006. ISBN 0-13-148004-9 (paperback). xxxvii + 1001 pp. LCCN QA76.76.O63 N448 2006. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0620/2006030150.html.

Nemeth:2010:ULS

[NSHW10] Evi Nemeth, Garth Snyder, Trent R. Hein, and Ben Whaley. UNIX and Linux system administration handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, fourth edition, 2010. ISBN 0-13-148005-7. xlvii + 1279 pp. LCCN QA76.76.O63 N45 2010.

Nagasaka:1992:TCL

[NSI^{+}92] Atsushi Nagasaka, Yoshihiro Shintani, Tanji Ito, Hiroshi Gomi, and Junichi Takahashi. Tachyon Common Lisp: An efficient and portable implementation of CLtL2. In ACM [ACM92], pages 270–277. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

Nemeth:1989:USA

[NSS89] Evi Nemeth, Garth Snyder, and Scott Seebass. UNIX System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-933441-6. xxx + 593 pp. LCCN QA76.76.O63 N45 1989.

Nemeth:2001:USA

[NSS^{+}01] Evi Nemeth, Garth Snyder, Scott Seebass, Trent R. Hein, Adam Boggs, Rob Braun, Ned McClain, Dan Crawl, Lynda McGinley, and Todd Miller. UNIX System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2001. ISBN 0-13-020601-6. xxxv + 853 pp. LCCN QA76.76.O63 N45 2001. US$68.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0130206016.html.

Nemeth:1995:USA

[NSSH95] Evi Nemeth, Garth Snyder, Scott Seebass, and Trent R. Hein. UNIX System Administration Handbook. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1995. ISBN 0-13-151051-7. xxxii + 779 pp. LCCN QA76.76.O63 N45 1995.

URL http://www.phptr.com/
ptrbooks/ptr_0131510517.

html. Includes CD ROM.

Navarro:2005:LBM

[NT05] Gonzalo Navarro and Jorma Tarhio. LZgrep: a Boyer– Moore string matching tool for Ziv–Lempel compressed text. Software—Practice and Experience, 35(12):1107–1130, October 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL

http://www.dcc.uchile.cl/
~gnavarro/pubcode/.

Nutt:2001:KPL

[Nut01] Gary J. Nutt. Kernel projects for Linux. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 2001. ISBN 0-201-61243-7. xvi + 239 pp. LCCN QA76.76.O63 N885 2001.

Nieh:2006:ETO

[NV06] Jason Nieh and Chris Vaill. Experiences teaching operating systems using virtual platforms and Linux. Operating Systems Review, 40(2):100–104, April 2006. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Negus:2003:LTC

[NW03] Chris Negus and Chuck Wolber. Linux toys: 13 cool projects for home, office, and entertainment. Extreme-Tech. Wiley, New York, NY, USA, 2003. ISBN 0-7645-2508-5. xxv + 330 pp. LCCN QA76.76.O63 N423

                                        2003. URL http://www.loc.
                                        gov/catdir/bios/wiley046/
                                        2003101901.html;   http:/
                                        /www.loc.gov/catdir/description/
                                        wiley039/2003101901.html;
                                         http://www.loc.gov/catdir/
                                        toc/wiley041/2003101901.
                                        html.

Nye:1988:XPM

[Nye88a] Adrian Nye. Xlib Programming Manual for Version 11, volume 1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-26-9, 0-937175-89-7. xxxiii + 615 pp. LCCN QA76.76.W56 D44 v.1 1988. URL http:/

/www.oreilly.com/catalog/
9780937175262.

Nye:1988:XRM

[Nye88b] Adrian Nye. Xlib Reference Manual for Version 11, volume 2. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-27-7. xiv + 701 pp. LCCN QA76.76.W56 D44 v.2 1988.

Nye:1990:XPR

[Nye90a]  Adrian Nye.  X Protocol

Reference Manual, volume 0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1990. ISBN 0-937175-50-1. LCCN QA76.76.W56 X215 1990.

Nye:1990:XPM

[Nye90b] Adrian Nye. Xlib Programming Manual: for Version 11 of the X Window System, volume 1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, July 1990. ISBN 0-937175-11-0. xxxi + 635 pp. LCCN QA76.76.W56 N93 1990.

Nye:1990:XRM

[Nye90c] Adrian Nye. Xlib Reference Manual: for Version 11 of the X Window System, volume 1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, July 1990. ISBN 0-937175-12-9. xxiii + 765 pp. LCCN QA76.76.W56 X52 1990.

Nye:1992:XPR

[Nye92a]  Adrian Nye.  X Protocol

Reference Manual, volume 0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1992. ISBN 1-56592-008-2. LCCN QA76.76.W56 X215 1990. URL http:/

                                        /www.oreilly.com/catalog/
                                        9781565920088.

Nye:1992:XPM

[Nye92b] Adrian Nye. Xlib Programming Manual, volume 1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, July 1992. ISBN 1-56592-002-3. 824 pp. LCCN QA76.76.W56N93 1990. US$34.95. URL http://www.

                                        oreilly.com/catalog/v1.

Nye:1995:PSR

[Nye95] Adrian Nye, editor. Programmer’s Supplement for Release 6 of the X Window System, version 11. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 1995. ISBN 1-56592-089-9. xxi + 430 pp. LCCN QA76.76.W56F55 1995. US$29.95. URL http://www.

                                        oreilly.com/catalog/r6.

Northcutt:2002:INP

[NZW^{+}02] Stephen Northcutt, Lenny Zeltser, Scott Winters, Karen Fredrick, and Ronald W.

Ritchey. Inside Network Perimeter Security: The Definitive Guide to Firewalls, Virtual Private Networks (VPNs), Routers, and Intrusion Detection Systems. Que Corporation, Indianapolis, IN, USA, 2002. ISBN 0-7357-1232-8. xxvii + 678 pp. LCCN TK5105.59 .I53 2003. US$49.99.

OReilly:1998:CWN

[O^{+}98] O’Reilly et al. The Complete Windows NT & UNIX System Administration Pack. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN ???? ?? + ?? pp. LCCN ???? US$149.00. URL http:/

/www.oreilly.com/catalog/
hp12pack.

OBrien:1982:PUP

[O’B82] Gregory J. O’Brien. Porting UNIX to a personal computer. In Usr Group [Usr82], pages 247–?? Abstract only.

Ottaviano:2023:CLL

[OBSB23] Alessandro Ottaviano, Thomas Benz, Paul Scheffler, and Luca Benini. Cheshire: a lightweight, Linux-capable RISC-V host platform for domain-specific accelerator plug-in. IEEE Transactions on Circuits and Systems II: Express Briefs, 70(10): 3777–3781, 2023. ISSN 1549-7747 (print), 1558-3791 (electronic).

ODell:1983:BUA

[O’D83a] Michael O’Dell. Berkeley UNIX after 4.2BSD: Where is it going and why do we want it to get there? In Software Tools Users Group [Sof83], pages 292–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

ODell:1983:UEM

[O’D83b] Michael D. O’Dell. UNIX and electronic mail: Trials, tribulations, and proposals. In Software Tools Users Group [Sof83], pages 327–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

ODell:1983:PUW

[O’D83c] Mike O’Dell. Portability in the UNIX world — what UNIX can learn from the software tools. In Association [Ass83], pages 314–?? Abstract only.

OReilly:1986:UNS

[OD^{+}86] Tim O’Reilly, Jean Diaz, et al. UNIX in a Nutshell: System V Edition: a Desktop Quick Reference. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1986. ISBN 0-937175-19-6 (paperback). 260 (est.) pp. LCCN QA76.76.O63 U5453 1990.

ODell:1987:UWV

[O’D87]  Mike O’Dell.  UNIX: The

world view. In USENIX Association [USE87e], pages 35– 45. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

ODell:1990:PUV

[O’D90] Michael D. O’Dell. Putting UNIX on very fast computers. In USENIX [USE90c], pages 239–246. LCCN QA76.8.U65 U81 1990.

Ousterhout:1985:TDA

[ODH^{+}85] John K. Ousterhout, Hervé Da Costa, David Harrison, John A. Kunze, Mike Kupfer, and James G. Thompson. A trace-driven analysis of the UNIX 4.2 BSD file system. Operating Systems Review, 19(5): 15–24, December 1985. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

OGorman:2001:OSL

[O’G01] John O’Gorman. Operating systems with Linux. Cornerstones of computing. Palgrave, Basingstoke, UK, 2001. ISBN 0-333-94745-2. xv + 462 pp. LCCN QA76.76.O63 O336 2001.

OnWord:1993:FSS

[OH93a] OnWord Press development team and Clint Hicks. Five steps to SunSoft Solaris 2.*. OnWord Press, Santa Fe, NM, USA, 1993. ISBN 0-934605-80-7. xv + 195 pp. LCCN QA76.76.O63 F57 1993. ????

OnWord:1993:SSQ

[OH93b] OnWord Press development team and Clint Hicks. The Sun Solaris 2.* quick reference. OnWord Press, Santa Fe, NM, USA, 1993. ISBN 0-934605-76-9. xvi + 208 pp. LCCN QA76.8.S86 S85 1993. ????

OBrien:1997:SMR

[OH97] Timothy O’Brien and Douglas Heise. Sun/Microsoft rivalry: Will Java fragment like UNIX? Object Magazine, 6(12):14–16, February 1997. CODEN OB-MAFO. ISSN 1055-3614.

Oglesby:2005:VES

[OH05] Ron Oglesby and Scott Herold. VMware ESX Server: Advanced Technical Design Guide. Brian Madden Publishing, Silver Spring, MD, USA, 2005. ISBN 0-9711510-6-7. LCCN QA76.9.V5 O35 2005eb; **See. URL

http://library.books24x7.
com/library.asp?B&bookid=
12048.

Ossanna:1990:TUM

[OK90] J. F. Ossanna and B. W. Kernighan. Troff user’s manual. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Ossanna:1992:TUM

[OK92] Joseph F. Ossanna and Brian W.

Kernighan. Troff user’s manual. Computing Science Technical Report 54, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1992. ii + 35 pp.

URL https://troff.org/54.

pdf. Revised edition of [Oss79].

OnWord:1994:SSU

[OK94] OnWord Press development team and Sam Kimery. SunSoft Solaris 2 user’s guide. OnWord Press, Santa Fe, NM, USA, 1994. ISBN 0-934605-74-2. xxii + 304 pp. LCCN QA76.76.O63 S865 1994. ????

Okrent:1978:TOS

[Okr78] H. F. Okrent. TBL options summary. Technical Memorandum 1314, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 6, 1978. ?? pp.

Olczak:1991:KSQ

[Olc91] Anatole Olczak. The Korn Shell Quick Reference Guide. ASP, San Jose, CA, USA, 1991. ISBN 0-935739-21-1. 52 pp. LCCN QA76.73.K67 O42 1991.

Olczak:1992:KSU

[Olc92] Anatole Olczak. The Korn Shell User and Programming Manual. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56548-X. xiv + 384 pp. LCCN QA76.73 K67 O43 1992. An everythingyou-want-to-know-about-Korn-Shell book.

Olczak:2001:KSU

[Olc01] Anatole Olczak. The Korn shell: Unix and Linux programming manual. Addison-Wesley, Reading, MA, USA, third edition, 2001. ISBN 0-201-67523-4. xxi + 446 pp. LCCN QA76.73.K67 O38 2001.

Olker:2003:ONP

[Olk03] Dave Olker. Optimizing NFS performance: tuning and troubleshooting NFS on HP-UX systems. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-042816-7. xxxvi + 328 pp. LCCN TK5105.574 .O45 2003.

Olsson:1973:GCC

[Ols73] S. B. Olsson. Guidelines for C compiler distribution. Technical Memorandum 1229 (TM 78-8234-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 3, 1973. ?? pp.

Olsson:1978:IUC

[Ols78] S. B. Olsson. On the inconsistencies of UNIX commands. Technical Memorandum 1325, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 26, 1978. ?? pp.

Ooms:2013:RPE

[Oom13] Jeroen Ooms. The RAppArmor package: Enforcing security policies in R using dynamic sandboxing on Linux. Journal of Statistical Software, 55(7): ??, November 2013. CODEN JSSOBK. ISSN 1548-7660.

URL http://www.jstatsoft.
org/v55/i07.

OBrien:2001:KSP

[OP01] Dennis O’Brien and David Pitts. Korn shell programming by example. Que Corporation, Indianapolis, IN, USA, 2001. ISBN 0-7897-2465-0. xiv + 431 pp. LCCN QA76.73.K67 O33 2001. US$29.99.

OSF:1991:OMPb

[Ope91a] Open Software Foundation.

OSF/Motif Programmer’s Reference, Revision 1.1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-640681-5. xii + 1212 + 15 pp. LCCN QA76.76.W56 O7 1991.

OSF:1991:OMS

[Ope91b] Open Software Foundation.

OSF/Motif Style Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1991. ISBN 0-13-640616-5. LCCN QA76.76.W56 O833 1991.

OpenGL:1993:ORM

[Ope93] OpenGL Architecture Review Board. OpenGL Reference Manual: The Official Reference Document for OpenGL, Release 1. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-63276-4. ix + 388 pp. LCCN T385 .O64 1993. US$32.95, CDN$42.95.

OpenInventor:1994:OIC

[Ope94] Open Inventor Architecture Group. Open Inventor C++ Reference Manual: The Official Reference Document for Open Systems. Addison-Wes-ley, Reading, MA, USA, 1994. ISBN 0-201-62491-5. vi + 767 pp. LCCN QA76.64 .O6 1994.

OReilly:1988:XWS

[OQL88] Tim O’Reilly, Valerie Quercia, and Linda Lamb. X Window System User’s Guide for Version 11, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1988. ISBN 0-937175-29-3. xviii + 344 pp. LCCN QA76.76.W56 D44 v.3 1988.

OReilly:1990:BUN

[O’R90a] Tim O’Reilly. Das BSD-Unix-Nutshell-Buch. (German) [The BSD UNIX Nutshell Book]. Addison-Wesley, Reading, MA, USA, 1990. ISBN 3-89319-219-0. xii + 318 pp. LCCN ????

OReilly:1990:XTI

[O’R90b] Tim O’Reilly. X Toolkit Intrinsics Reference Manual, volume 5. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1990. ISBN 0-937175-35-8. xii + 543 pp. LCCN QA76.76.W56 D44 v.5 1990.

OReilly:1993:COB

[O’R93]  O’Reilly staff.  Catalog of

O’Reilly Books, 1993. You

                                        get can it (book.catalog.Z) by

anonymous ftp from ftp.ora. com where you can also find source codes for examples in many books in the Nutshell and X series … Or read the catalog on the O’Reilly gopher server

(telnet gopher.ora.com, lo-
gin: gopher or gophergopher.
ora.com (if a gopher client is

installed) …). And to get a hard copy of their catalog-plus-magazine called ora.com, mail

letters@ora.com.

Oram:2021:APP

[Ora21]  Andy Oram.  Awk: The

power and promise of a 40year-old language. FOSSlife Web site, May 19, 2021. URL

https://www.fosslife.org/
awk-power-and-promise-40-
year-old-language.

Oxer:2006:UH

[ORC06] Jonathan Oxer, Kyle Rankin, and Bill Childers. Ubuntu Hacks. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52720-9. xix + 426 pp. LCCN QA76.76.O63 2006. EUR 29.00.

Orwant:1996:PIC

[Orw96] John Orwant. Perl 5 interactive course. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-064-6. xli + 938 pp. LCCN QA76.73.P22 O79 1996.

Ozment:2006:MWD

[OS06] Andy Ozment and Stuart E.

Schechter. Milk or wine:

Does software security improve with age? In Angelos D. Keromytis, editor, Proceedings of the 15th USENIX Security Symposium, pages 93–104. USENIX Association, Berkeley, CA, USA, July 2006.

                                        URL https://www.usenix.
                                        org/conference/15th-usenix-
                                        security-symposium/milk-
                                        or-wine-does-software-security-
                                        improve-age.

Osborne:1992:TCU

[Osb92] Lawrence Osborne. Teaching C with UNIX for college credit to professional programmers. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 24 (4):43–48, December 1992. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Osman:1996:SCC

[Osm96] Mohamed (Mohamed S.) Osman. Successful C for commercial UNIX developers. The Artech House computer science library. Artech House Inc., Boston, MA, USA, 1996. ISBN 0-89006-642-6. xiv + 369 pp. LCCN QA76.73.C15 O817 1996.

Ossanna:1974:TUM

[Oss74] J. F. Ossanna. TROFF user’s manual. Internal memorandum ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974. ?? pp.

Ossanna:1976:NTU

[Oss76] Joseph F. Ossanna. NROFF/ TROFF user’s manual. Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1976. 34 pp.

Ossanna:1979:NTU

[Oss79] Joseph F. Ossanna. NROFF/ TROFF user’s manual. Technical Memorandum 1119, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

OReilly:1988:MUU

[OT88] Tim O’Reilly and Grace Todino. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1988. ISBN 0-937175-09-9. xvi + 256 pp. LCCN QA76.8.U65 O64 1988.

Oram:1991:MPM

[OT91] Andrew Oram and Steve Talbott. Managing Projects with Make. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1991. ISBN 0-937175-90-0. xiv + 136 pp. LCCN QA76.76.O63 T35 199. URL http:/

/www.oreilly.com/catalog/
9780937175903;   http://
www.oreilly.com/catalog/
make2.

Ortiz-Tapia:2007:RRL

[OT07] Arturo Ortiz-Tapia. A recommended resource for Linuxbased companies. IEEE Distributed Systems Online, 8 (8):??, August 2007. CO-DEN ???? ISSN 1541-4922 (print), 1558-1683 (electronic).

URL http://csdl.computer.
org/comp/mags/ds/2007/08/
mds2007080004.pdf.

Oualline:1991:PCP

[Oua91] Steve Oualline. Practical C Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-65-X. xxii + 396 pp. LCCN QA76.73.C15 O84 1991. US$24.95. Yet another good C book describing how to create programs that are easy to read, maintain and debug.

Oualline:1992:CES

[Oua92] Steve Oualline. C elements of style: the programmers style manual for Elegant C and C++ programs. M and T Books, San Matteo, CA, USA, 1992. ISBN 1-55851-291-8. xi + 265 pp. LCCN QA76.73.C153 O83 1992.

Oualline:1997:DL

[Oua97] Steve Oualline. Discover Linux. I D G Books Worldwide, Indianapolis, IN, USA, 1997. ISBN 0-7645-3105-0. xxvi + 438 pp. LCCN QA76.76.O63O83 1997. US$24.99.

Ousterhout:1994:TTT

[Ous94] John K. Ousterhout. Tcl and the Tk Toolkit. Addison-Wes-ley, Reading, MA, USA, 1994.

ISBN 0-201-63337-X. xx + 458 pp. LCCN QA76.73.T44 O97 1994. US$36.75.

Outerbridge:1982:CFU

[Out82] Richard Outerbridge. Cryptographic features of the UNIX operating system. Cryptologia, 6(3):253–257, July 1982. CODEN CRYPE6. ISSN 0161-1194 (print), 1558-1586 (electronic). URL http:

//www.cs.bell-labs.com/~
dmr/crypt.html;   http://
www.informaworld.com/smpp/
content~content=a741903079~
db=all~order=page;   https:
//www.tandfonline.com/doi/
abs/10.1080/0161-118291857055.

Ozimek:1978:UDE

[Ozi78] L. C. Ozimek. Use of display editors. Technical Memorandum 1292, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 21, 1978. ?? pp.

Pritchard:2006:LLC

[P^{+}06] Steven Pritchard et al., editors.

LPI Linux certification in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-00528-8. xviii + 961 pp. LCCN QA76.76.O63; QA76.76.O63 L65 2006eb. URL http:/

/www.oreilly.com/catalog/
9780596005283.

Parziale:2008:ZVL

[P^{+}08] Lydia Parziale et al., editors.

z/VM and Linux on IBM System z: the virtualization cookbook for RHEL 5.2. Number SG24-7492-00 in IBM redbooks. IBM Corporation, San Jose, CA, USA, 2008. ISBN 0-7384-3181-8. xvi + 250 pp. LCCN QA76.76.O63 Z867

2008. URL http://proquest.
safaribooksonline.com/?fpi=
0738431818.

Pabrai:1996:UI

[Pab96] Uday O. Pabrai. UNIX internetworking. The Artech House telecommunications library. Artech House Inc., Boston, MA, USA, second edition, 1996. ISBN 0-89006-778-3. xix + 454 pp. LCCN TK5105.5 .P32 1996.

Paciello:2000:COM

[Pac00] Michael G. Paciello. C++ Objects for Making UNIX and Windows NT Talk. CMP Books, ????, 2000. ISBN ???? ???? pp. LCCN ????

Packer:2002:CTD

[Pac02] Allan N. Packer. Configuring & Tuning Databases on the Solaris Platform. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-083417-3. xxxiv + 502 pp. LCCN QA76.9.D26P33 2002.

Padovano:1993:NAU

[Pad93] Michael Padovano. Networking Applications on UNIX System V. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-613555-2. xvi + 544 pp. LCCN QA76.76.O63 P33 1993. US$44.

Pajari:1988:MUW

[Paj88] George E. Pajari. Of mice and UNIX — writing a UNIX device driver for the Microsoft bus mouse. C Users Journal, 6(9): 54–??, September 1988. ISSN 0898-9788.

Pajari:1992:WUD

[Paj92] George Pajari. Writing UNIX Device Drivers. Addison-Wes-ley, Reading, MA, USA, 1992. ISBN 0-201-52374-4. xii + 323 pp. LCCN QA76.76.D49 P35 1992. An easy-to-understand book covering character, block, terminal, and stream drivers.

Palmer:1987:DEOb

[Pal87] James E. Palmer. The design and evaluation of online help for UNIX Emacs: Capturing the user in menu design. CDC technical report series 39, Communications Design Center, Carnegie Mellon University, Pittsburgh, PA, USA, November 1987. 17 pp.

Palnitkar:1996:VHG

[Pal96] Samir Palnitkar. Verilog HDL:

a Guide to Digital Design and Synthesis. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1996. ISBN 0-13-451675-3. xxxviii + 396 pp. LCCN TK7885.7.P35 1996. URL

http://www.sun.com/books/
catalog/palnitkar/index.
html.

Panoff:1988:RPR

[Pan88] Robert M. Panoff. Real productivity for real science without real UNIX. In USENIX Association [USE88f], pages 35–?? ISBN ???? LCCN ???? Abstract only.

Panattoni:1994:PSB

[Pan94] Dale Panattoni. Printer sharing between UNIX and Windows for Workgroups. Sys Admin: The Journal for UNIX Systems Administrators, 3(6):41– ??, November/December 1994. CODEN SYADE7. ISSN 1061-2688.

Para:1978:CRD

[Par78] P. S. Para. CLIO — a relational data base system. Technical Memorandum 1243 (TM 78-5522-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 9, 1978. ?? pp.

Partridge:1988:UIH

[Par88] Craig Partridge. A UNIX implementation of HEMS. In USENIX Association [USE88g], pages 89–96. ISBN ???? LCCN ????

Parker:1990:UIN

[Par90] Tim Parker. UNIX is no halfbaked operating system. Computer Language Magazine, 7(7): 83–??, July 1990. CODEN COMLEF. ISSN 0749-2839.

Parrette:1993:MPX

[Par93] William A. Parrette. Motif programming in the X Window System environment. Mc-Graw-Hill, New York, NY, USA, 1993. ISBN 0-07-031722-4 (hardcover), 0-07-031723-2 (softcover). xxii + 466 pp. LCCN QA76.76.W56 P37 1993. US$44.95 (hardcover), US$34.95 (softcover).

Parker:2000:LSA

[Par00] Tim Parker. Linux system administrator’s survival guide. SAMS Publishing, Indianapolis, IN, USA, second edition, 2000. ISBN 0-672-31793-1. xx + 740 pp. LCCN QA76.76.O63 P36 2000.

Pass:1988:UST

[Pas88] E. M. Pass. UNIX system 5 terminal raw mode setting. C Users Journal, 6(2):69–??, February 1988. ISSN 0898-9788.

Patriquin:1983:FSC

[Pat83] Ed Patriquin. File system considerations in a multiple processor UNIX environment. In Software Tools Users Group [Sof83], pages 118–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Pate:1996:UIP

[Pat96] Steve D. Pate. UNIX Internals: a Practical Approach. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 1996. ISBN 0-201-87721-X. xxii + 654 pp. LCCN QA76.76.O63P3776 1996. US$45.75. This books discusses UNIX internals for SCO UNIX, derived from AT&T System V.3 UNIX, running on Intel x86 hardware.

Patsis:1999:UAS

[Pat99]  Peter Patsis.  UNIX awk

and sed Programmer’s Interactive Workbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-082675-8. xix + 622 pp. LCCN QA76.76.O63P3777 1999. US$34.99.

Patterson:2012:TDR

[Pat12] David Patterson. Tribute to Dennis Ritchie. Alcatel-Lucent 24-minute video., September 11, 2012. URL https:/

/www.youtube.com/watch?v=

Bx8PL5sMD6o. From an event entitled, “The Lasting Legacy of Dennis Ritchie: The Impact of Software on Society.”.

Pavlidis:1981:CFR

[Pav81] Theo Pavlidis. Contour filling in raster graphics. Computer Graphics, 15(3):29– 36, August 1981. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/
800224.806786.

Pavlidis:1980:CFR

[Pav0 ] Theo Pavlidis. Contour filling in raster graphics. Computing Science Technical Report 96, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1980 (??). ?? pp.

Paxson:1988:FFL

[Pax88] Vern Paxson. flex — fast lexical analyzer generator. Free Software Foundation, Inc. 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel: (617) 876-3296, 1988. Elec-

tronic mail: vern@lbl-csam.
arpa or vern@lbl-rtsg.arpa.

Software also available via ANONYMOUS FTP to lbl-

csam.arpa, lbl-rtsg.arpa,
or prep.ai.mit.edu. See also

[DS88].

Payne:1984:DAPa

[Pay84a] Kendall Robert Payne. The development and analysis of a portable runtime library accessible to all FORTRAN, COBOL and PASCAL compilers under the UNIX system 5 operating system. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1984. 70 pp.

Payne:1984:DAPb

[Pay84b] Kendall Robert Payne. The development and analysis of a portable runtime library accessible to all FORTRAN, COBOL and PASCAL compilers under the UNIX system 5 operating system. Thesis (M.S.), Kansas State University, Manhattan, KS, USA, 1984. 70 pp.

Polyak:1984:LUR

[PB84] Steven T. Polyak and Jeffrey S.

Barr. Life with UNIX in realtime. In USENIX Association [USE84b], pages 285–292. ISBN none. LCCN QA76.8.U65 U55 1984.

Park:1991:MPB

[PB91] Arvin Park and Jeffrey C. Becker. Measurements of the paging behavior of UNIX. ACM SIGMETRICS Performance Evaluation Review, 19 (1):216–217, May 1991. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Petreley:2005:LDH

[PB05] Nick Petreley and Jono Bacon. Linux desktop hacks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00911-9. xxi + 318 pp. LCCN QA76.76.O63 P528645 2005. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596009113.

Piper:2009:LXT

[PB09] Ian Piper and James Bucanek. Learn Xcode tools for Mac OS X and iPhone development. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-7221-X. LCCN QA76.76.O63 P56 2009.

Paloque-Berges:2021:RPM

[PB21] Camille Paloque-Bergès. The real “poor man’s Arpanet”? A conversation about Unix networks with Kilnam Chon, godfather of the Asian Internet. Internet Histories, 5(3–4):376– 388, 2021. CODEN ???? ISSN 2470-1483. URL http://www.

tandfonline.com/doi/full/
10.1080/24701475.2021.1982167.

Plank:1995:LTC

[PBKL95] James S. Plank, Micah Beck, Gerry Kingsley, and Kai Li. Libckpt: Transparent checkpointing under UNIX. In USENIX [USE95b], pages 213–

224. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL  http://www.usenix.
org/publications/library/
proceedings/neworl/plank.
html.

Probert:1986:SIH

[PBL86] Dave Probert, Jeff Berkowitz, and Mark Lucovsky. A straightforward implementation of 4.2BSD on a highperformance multiprocessor. In USENIX Association [USE86e], pages 141–156. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Pagani:2022:LBS

[PBM^{+}22] Marco Pagani, Alessandro Biondi, Mauro Marinoni, Lorenzo Molinari, Giuseppe Lipari, and Giorgio Buttazzo. A Linuxbased support for developing real-time applications on heterogeneous platforms with dynamic FPGA reconfiguration. Future Generation Computer Systems, 129(??):125– 140, April 2022. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167739X21004362.

Paas:1998:WAE

[PBS98] Sven M. Paas, Thomas Bemmerl, and Karsten Scholtyssik. Win32 API emulation on UNIX for software DSM. In USENIX [USE98a], page ?? ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://www.

                                        usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/paas.html;   http://
                                        www.usenix.org/publications/
                                        library/proceedings/usenix-
                                        nt98/paas_slides.pdf.

Peachey:1984:EIS

[PBWB84] Darwyn R. Peachey, Richard B. Bunt, Carey L. Williamson, and Tim B. Brecht. An experimental investigation of scheduling strategies for UNIX. ACM SIGMETRICS Performance Evaluation Review, 12 (3):158–166, August 1984. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Presotto:1990:USA

[PC90] D. L. Presotto and W. R. Cheswick. Upas — a simpler approach to network mail. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

                                        1990. URL http://doc.cat-
                                        v.org/unix/v10/10thEdMan/
                                        v2contents.html. Two vol-

umes. Volume 2 is edited by A.

G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Phillips:2001:TIL

[PC01] Iain Phillips and Jon Crowcroft.

TCP/IP and the Linux protocol implementation: systems code for the Linux Internet. Wiley Networking Council series. Wiley, New York, NY, USA, 2001. ISBN 0-471-40882-4 (cloth). 925 pp. LCCN TK5105.585 .T34 2002.

Pausch:1992:LLS

[PCD92] Randy Pausch, Matthew Conway, and Robert DeLine. Lessons learned from SUIT, the Simple User Interface Toolkit. ACM Transactions on Information Systems, 10(4):320– 344, October 1992. CO-DEN ATISET. ISSN 1046-

8188. URL http://www.acm.
org:80.

Pilato:2008:VCS

[PCSF08] C. Michael Pilato, Ben Collins-Sussman, and Brian W. Fitzpatrick. Version control with Subversion. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2008. ISBN 0-596-51033-0 (paperback). xxiv + 404 pp. LCCN QA76.6 .C6274 2008.

URL http://svnbook.red-
bean.com/.

Palmer:1987:DEOa

[PDG^{+}87] J. E. Palmer, T. M. Duffy, K. Gomoll, T. Gomoll, J. Palmquist-Richards, and J. A. Trumble. The design and evaluation of online help for UNIX EMACS: access mechanisms. In Bullinger et al. [BSK87], pages 461–466. ISBN 0-444-70304-7. LCCN QA76.9.S88I325 1987.

Palmer:1988:DEO

[PDG^{+}88] J. Palmer, T. Duffy, K. Gomoll, T. Gomoll, J. Richards-Palmquist, and J. A. Trumble. The design and evaluation of online help for Unix EMACS: capturing the user in menu design. IEEE Transactions on Professional Communication, 31(1):44–51, March 1988. CODEN IEPCBU. ISSN 0361-1434 (print), 1558-1500 (electronic).

Palmer:2002:GUU

[PDG02] Michael Palmer, Jack Dent, and Tony Gaddis. Guide to UNIX using Linux. Networking. Thomson/Course Technology, Boston, MA, USA, second edition, 2002. ISBN 0-619-12147-5. xx + 540 pp. LCCN QA76.76.O63 P35 2002. Previous edition written by Jack Dent and Tony Gaddis. Includes a copy of Red Hat Linux 7.2 publisher’s edition.

Pepple:2004:MSO

[PDL04] Ken Pepple, Brian Down, and David Levy. Migrating to the Solaris Operating System: The Discipline of UNIX-to-UNIX Migrations. Sun BluePrints Program. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-150263-8. xix + 251 pp. LCCN QA76.76.O63 P445 2004.

Pearlman:1977:PUO

[Pea77]  M. E. Pearlman.  PWB/

UNIX operations manual (1.3). Technical Memorandum 1196, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1977. ?? pp.

Peachey:1980:BDU

[Pea80] Darwyn Peachey. Buffer deadlock in UNIX. ;login: the USENIX Association newsletter, 5(8):10–??, October 1980. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Peachey:1983:ISN

[Pea83] Darwyn Peachey. Improved schedulers for non-paged UNIX systems. In Association [Ass83], pages 39–?? Abstract only.

Peck:2008:BGN

[Pec08]  Akkana Peck.  Beginning

GIMP: from novice to professional. The expert’s voice in open source. Apress, Berkeley, CA, USA, second edition, 2008. ISBN 1-4302-1070-2 (paperback). xxvi + 557 pp. LCCN T385 .P395 2008.

Pekarich:1978:UTS

[Pek78] S. P. Pekarich. UNIX timesharing system: No. 4 ESS diagnostic environment. The Bell System Technical Journal, 57(6):2265–2274, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2265.
                                        pdf.

Perez:1977:UUE

[Per77] C. D. Perez. On using the UNIX error logger. Technical Memorandum 1176 (MF 77-8234-82), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 31, 1977. ?? pp.

Perez:1978:REC

[Per78] C. D. Perez. A revised ERRPT command. Technical Memorandum 1270 (MF 78-8234-60), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 12, 1978. ?? pp.

Perez:1979:GCL

[Per79] C. D. Perez. A guide to the C library for UNIX users. Technical Memorandum 1446, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 15, 1979. ?? pp.

Perlman:1982:DAP

[Per82a] Gary Perlman. Data analysis programs on CSL UNIX. In Usr Group [Usr82], pages 211–224.

Perlman:1982:MIU

[Per82b] Gary Perlman. MENUNIX: An interface to UNIX files and programs. In Usr Group [Usr82], pages 225–234.

Perlman:1984:UCD

[Per84] Gary Perlman. UNIX—STAT:

Compact data analysis package. The American Statistician, 38(2):155, May 1984. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/2683258.

Perkins:1987:MDP

[Per87] Lawrence B. Perkins. Managing the development of performance-constrained UNIX-Based software on microcomputers. In USENIX Association [USE87e], pages 46–59. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Perkins:1993:SYN

[Per93] Edwin C. Perkins, Jr. Should Your Next UNIX Workstation be a PC? UNIX/World, 10(2): 90–??, February 1993. ISSN 0739-5922.

Peters:1988:UPM

[Pet88] James F. Peters. UNIX programming: methods and tools. Harcourt Brace Jovanovich, Boston, MA, USA, 1988. ISBN 0-15-593021-4 (paperback). xv + 447 pp. LCCN QA76.76.O63 P52 1988. See book review [Fla90]. System requirements for computer disk: IBM PC.

Petersen:1998:LPR

[Pet98a] Richard Petersen. Linux programmer’s reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1998. ISBN 0-07-882587-3. xiv + 303 pp. LCCN QA76.76.O63 P525 1998. US$16.99.

Petersen:1998:LCR

[Pet98b] Richard Petersen. Linux: The Complete Reference. McGraw-Hill, New York, NY, USA, second edition, 1998. ISBN 0-07-882461-3. xxvi + 1059 pp. LCCN QA76.76.O63 P523 1998. US$49.99.

Petersen:1998:UCE

[Pet98c]  Richard Petersen.  UNIX

Clearly Explained. AP Professional, Boston, MA, USA, 1998. ISBN 0-12-552130-8. xviii + 691 pp. LCCN QA76.76.O63P524 1999. US$39.95.

Petersen:1999:UCE

[Pet99] Richard Petersen. UNIX clearly explained. AP Professional, Boston, MA, USA, 1999. ISBN 0-12-552130-8. xviii + 691 pp. LCCN QA76.76.O63 P524 1999.

Petersen:2000:LPR

[Pet00a] Richard Petersen. Linux programmer’s reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, second edition, 2000. ISBN 0-07-212355-9. xix + 443 pp. LCCN QA76.8.U65 P484 2000. US$19.99.

Petersen:2000:LCR

[Pet00b] Richard Petersen. Linux: The Complete Reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fourth edition, 2000. ISBN 0-07-212940-9. xxxv + 1257 pp. LCCN QA76.76.O63 P523 2001. US$19.99.

Petron:2000:LER

[Pet00c] Ed Petron. Linux essential reference. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0852-5. xxiii + 332 pp. LCCN QA76.76.O63 P5286 2000.

Petersen:2001:LPB

[Pet01a] Richard Petersen. Linux programming: a beginner’s guide. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-212743-0. xxiii + 422 pp. LCCN QA76.76.O63 P5228 2001.

Petersen:2001:LCR

[Pet01b] Richard Petersen. Linux: the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fourth edition, 2001. ISBN 0-07-212940-9. xxxv + 1257 pp. LCCN QA76.76.O63 P523 2001. Includes CD-ROM.

Petersen:2002:LCR

[Pet02a] Richard Petersen. Linux: the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fifth edition, 2002. ISBN 0-07-222505-X, 0-07-222506-8 (book), 0-07-222507-6 (DVD). xxvii + 911 pp. LCCN QA76.76.O63 P523 2002. DVD-ROM includes complete 5 CD-ROM set for the entire Red Hat distributions and comprehensive set of Linux software applications, including the GNU software packages (graphics, communications, publishing, editing, programming, games), as well as development tools, and Internet servers (FTP, Web, mail, news, and DNS).

Petersen:2002:RLC

[Pet02b] Richard Petersen. Red Hat Linux: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2002. ISBN 0-07-219178-3. xxx + 1117 pp. LCCN QA76.76.O63 P5237 2002.

Peters:2009:ESS

[Pet09] Ron Peters. Expert shell scripting. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2009. ISBN 1-4302-1841-X. xxii + 293 pp. LCCN QA76.76.O63 P48 2009.

Pew:1993:GS

[Pew93] John A. Pew. Guide to Solaris. Ziff-Davis Press, Emeryville, CA, USA, 1993. ISBN 1-56276-087-4. xxv + 625 pp. LCCN QA76.76.O63 P534 1993. US$34.95, CDN$48.95.

Pawlowski:1984:DSL

[PF84a] Brian Pawlowski and Alan Filipski. The dynamics of a semilarge software project with specific reference to a UNIX system port. In USENIX [USE84a], pages 332–342.

Pawlowski:1984:DSS

[PF84b] Brian Pawlowski and Alan Filipski. The dynamics of a semilarge software project with specific reference to a UNIX system port. In Software Tools Users Group [Sof84], pages 332–342. LCCN QA76.8.U65 U83 1984.

Probst:1996:PLL

[PF96] Stefan Probst and Ralf Flaxa.

The power Linux: Linux 1.2, LST-distribution 2.1. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-14556-7. vi + 196 pp. LCCN QA76.76.O63P758 1996.

Pharr:2005:GGP

[PF05] Matt Pharr and Randima Fernando, editors. GPU gems 2: programming techniques for high-performance graphics and general-purpose computation, volume 2 of GPU gems. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-33559-7 (hardcover). xlix + 814 pp. LCCN T385 .G688

2005. URL http://www-docs.
tu-cottbus.de/bibliothek/
public/katalog/420569.PDF;

                                         http://www.loc.gov/catdir/
                                        toc/ecip055/2004030181.html.

Pfaffenberger:2000:LCR

[Pfa00] Bryan Pfaffenberger. Linux command reference. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2000. ISBN 0-7821-2748-7. 640 pp. LCCN QA76.76 .P435 2000.

Pfaffenberger:2001:LNC

[Pfa01] Bryan Pfaffenberger. Linux networking clearly explained. Academic Press, New York, NY, USA, 2001. ISBN 0-12-533171-1. ix + 390 pp. LCCN QA76.76.O63 P533 2001.

Petrisko:2020:BAO

[PGW^{+}20] D. Petrisko, F. Gilani, M. Wyse, D. C. Jung, S. Davidson, P. Gao, C. Zhao, Z. Azad, S. Canakci, B. Veluri, T. Guarino, A. Joshi, M. Oskin, and M. B. Taylor. BlackParrot: An agile open-source RISC-V multicore for accelerator SoCs. IEEE Micro, 40(4):93– 102, July/August 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Perlman:1986:UDA

[PH86] Gary Perlman and Frederick L. Horan. UNIX—STAT 5.2: Data analysis programs for MS–DOS and UNIX. The American Statistician, 40(3): 229–230, August 1986. CO-DEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (elec-

tronic). URL http://www.
jstor.org/stable/2684546.

Potmesil:1987:FST

[PH87] Michael Potmesil and Eric M.

Hoffert. FRAMES: Software tools for modeling, rendering and animation of 3D scenes. Computer Graphics, 21(4):85– 93, July 1987. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic). URL http://www.

acm.org:80/pubs/citations/
proceedings/graph/37401/
p85-potmesil/.

PeytonJones:1991:FIS

[PH91] Simon L. Peyton Jones and Mark S. Hardie. A Futurebus interface from off-the-shelf parts. IEEE Micro, 11(1):38– 41, 84–93, January/February 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pavlidis:1998:UCX

[PH98] Theo Pavlidis and Kevin Hunter. Using color in the X Window System versus Microsoft Windows: Part 1. IEEE Computer Graphics and Applications, 18(6):64–73, November/December 1998. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

URL http://computer.org/
cga/cg1998/g6064abs.htm;
 http://dlib.computer.org/
cg/books/cg1998/pdf/g6064.
pdf.

Pavlidis:1999:UCX

[PH99] Theo Pavlidis and Kevin Hunter. Using color in the X Window System versus Microsoft Windows: Part 2. IEEE Computer Graphics and Applications, 19(1):75–83, January/February 1999. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

                                        URL http://computer.org/
                                        cga/cg1999/g1075abs.htm;
                                         http://dlib.computer.org/
                                        cg/books/cg1999/pdf/g1075.
                                        pdf.

Phillips:2000:LMS

[Phi00] Dave Phillips. Linux Music and Sound. No Starch Press, San Francisco, CA, USA, 2000. ISBN 1-886411-34-4. xx + 408 pp. LCCN MT723 .P53 2000. US$39.95.

Pickard:1983:PNU

[Pic83] Monte Pickard. The Plexus networked UNIX. In Association [Ass83], pages 51–?? Abstract only.

Pike:1982:MBG

[Pik82] Rob Pike. Merging bitmap graphics and UNIX. In Usr Group [Usr82], pages 61–?? Abstract only.

Pike:1983:GOB

[Pik83a] R. Pike. Graphics in overlapping bitmap layers. ACM Transactions on Graphics, 2 (2):135–160, April 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pike:1983:USC

[Pik83b] Rob Pike. UNIX style, or cat -v considered harmful. In Software Tools Users Group [Sof83], pages 263–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Pike:1984:BMG

[Pik84a]  Rob Pike.  The Blit: a

multiplexed graphics terminal. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1607–1631, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Pike:1984:TOT

[Pik84b] Rob Pike. A text-oriented terminal multiplexor for blits. In USENIX [USE84a], pages 173– ?? Abstract only.

Pike:1987:TES

[Pik87] Rob Pike. The text editor sam.

Software—Practice and Experience, 17(11):813–845, November 1987. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:1988:WSS

[Pik88] Rob Pike. Window systems should be transparent. In Association [Ass88b], pages 279–

Pike:1989:CWS

[Pik89] Rob Pike. A concurrent window system. In Association [Ass89a], pages 133–153.

Pike:1990:TES

[Pik90] R. Pike. The text editor Sam. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Pike:1993:NLC

[Pik93] Rob Pike. Newsqueak: A language for communicating with mice. Computing Science Technical Report 143, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 16, 1993. ii + 13 pp.

Pike:19xx:PWS

[Pikxx] Rob Pike. \frac{81}{2}, the Plan 9 window system. Computing Science Technical Report 158d, AT&T Bell Laboratories, Murray Hill, NJ, USA, ?? ??, 19xx. 9 pp.

Pike:2006:SRE

[Pik06] Rob Pike. Structural regular expressions. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 23, 2006. 7

pp. URL https://doc.cat-v.
org/bell_labs/structural_
regexps/se.pdf.

Pintar:1992:PEV

[Pin92] Douglas L. Pintar. Practice and experience: Voices and FAXes and UNIX (oh, my!). ;login: the USENIX Association newsletter, 17(2):30–??, March/April 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Piner:1999:RMT

[Pin99] Mary-Louise G. Piner. Report to members: With two new awards, we honor Unix, RISC innovators. Computer, 32(5): 11–13, May 1999. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://dlib.computer.
org/co/books/co1999/pdf/
r5011.pdf.

Pirzada:1988:SEE

[Pir88] Shamim Sharifuddin Pirzada.

A Statistical Examination of The Evolution of the UNIX System. Ph.D. thesis, Imperial College of Science, Technology and Medicine, University of London, London, UK, September 1988. 182 pp. URL

https://spiral.imperial.
ac.uk/bitstream/10044/1/
7942/1/Shamim_Sharfuddin_
Pirzada-1988-PhD-Thesis.
pdf.

Piskorik:1977:PUM

[Pis77] E. M. Piskorik. PWB/UNIX manual page macros (I.9). Technical Memorandum 1152 (TM 77-9141-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1, 1977. ?? pp.

Phillipson:1989:PFS

[PJ89a] Steven H. Phillipson and Stefan Jeffers. Part-task flight simulation on a UNIX graphics workstation. In USENIX Association [USE89h], pages 115–128. ISBN ???? LCCN ????

Phillipson:1989:PTF

[PJ89b] Steven H. Phillipson and Stefan Jeffers. Part-task flight simulation on a UNIX graphics workstation. In USENIX Association [USE89h], pages 115–128. ISBN ???? LCCN ????

Pike:1984:PDU

[PK84] Rob Pike and Brian W. Kernighan. Program design in the UNIX system environment. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1595–1605, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Polk:1988:FUD

[PK88] Jeff Polk and Rob Kolstad. A faster UNIX dump program. In USENIX Association [USE88g], pages 125–129. ISBN ???? LCCN ????

Patel:1993:UTD

[PK93a] Dorab Patel and Scott D.

Kalter. A UNIX toolkit for distributed synchronous collaborative applications. In USENIX [USE93a], pages 105–133.

Patel:COMPSYS-6-2-105

[PK93b] Dorab Patel and Scott D.

Kalter. A UNIX toolkit for distributed synchronous collaborative applications. Computing Systems, 6(2):105–133, Spring 1993. CODEN CMSYE2. ISSN 0895-6340.

Prinz:2005:CEP

[PKP05] Peter Prinz and Ulla Kirch-Prinz. C — Einführung und professionelle Anwendung: [auf Basis des Standards ANSI C 99; Anwendungen: Windows-Programmierung, Grafik und hardwarenahe Programmierung; auf CD: Microsoft C/C++- Compiler Book edition, GNU C/C++-Compiler, Beispiele, Musterlöungen]. (German) [C — Introductory and Professional Use [based on ANSI C 99; Use: Windows Programming, Graphics and Hardware Programming; with CD: Microsoft C/C++-Compiler Book edition, GNU C/C++- Compiler, Examples, and Solutions]]. mitp-Verlag, Bonn, Germany, 2005. ISBN 3-8266-1580-8. 800 pp. LCCN ???? Includes one CD-ROM.

Placeway:1989:BDB

[Pla89] Paul W. Placeway. A better dump for BSD UNIX. In USENIX Association [USE89d], pages 99–107. ISBN ???? LCCN ????

Plauger:1992:SCL

[Pla92] P. J. Plauger. The Standard C Library. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-838012-0. xiv + 498 pp. LCCN QA76.73.C15 P563 1991.

Plauger:1992:UVM

[PLBL92] P. J. Plauger, Marty Leisner, Tim Berens, and Andy Levinson. UNIX versus MS-DOS — some letters. C Users Journal, 10(6):130–??, June 1992. ISSN 0898-9788. See additional comments, November 1992, p. 140; April 1993, p. 126.

Padioleau:2008:DAC

[PLHM08] Yoann Padioleau, Julia Lawall, René Rydhof Hansen, and Gilles Muller. Documenting and automating collateral evolutions in Linux device drivers. Operating Systems Review, 42 (4):247–260, May 2008. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Padioleau:2006:UCE

[PLM06] Yoann Padioleau, Julia L. Lawall, and Gilles Muller. Understanding collateral evolution in Linux device drivers. Operating Systems Review, 40(4):59– 71, October 2006. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Pike:1985:HST

[PLR85] Rob Pike, Bart Locanthi, and John Reiser. Hardware/ software trade-offs for bitmap graphics on the Blit. Software—Practice and Experience, 15(2):131–151, February 1985. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Pike:2000:HST

[PLR00] Rob Pike, Bart Locanthi, and John Reiser. Hardware/ software tradeoffs for bitmap graphics on the Blit. Report, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, March 30, 2000. URL

https://9p.io/cm/cs/doc/
87/archtr.ps.gz. Published

in [PLR85].

Prata:1987:USV

[PM87] Stephen Prata and Donald Martin. Unix System V Bible: commands and utilities. Howard W. Sams, Indianapolis, IN 46268, USA, 1987. ISBN 0-672-22562-X. ix + 516 pp. LCCN QA76.76.O63 P74 1987.

Pendry:1995:UML

[PM95] J. Pendry and Marshall Kirk McKusick. Union mounts in 4.4BSD-Lite. In ????, editor, Proceedings of the New Orleans Usenix Conference, January 1995, pages 25–33. USENIX Association, Berkeley, CA, USA, 1995.

Perla:2010:GKE

[PM10] Enrico Perla and Oldani Massimiliano. A guide to kernel exploitation: attacking the core. Syngress Publishing, Inc., Rockland, MA, USA, 2010. ISBN 1-59749-486-0 (paperback). xxi + 442 pp. LCCN QA76.76.O63 P5168 2010.

Papadimitriou:2018:MVF

[PM18] Stergios Papadimitriou and Lefteris Moussiades. Mac OS versus FreeBSD: A comparative evaluation. Computer, 51(2): 44–53, February 2018. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

                                        URL https://www.computer.
                                        org/csdl/mags/co/2018/02/
                                        mco2018020044-abs.html.

Pato:1988:UAR

[PMD88] Joseph N. Pato, Elizabeth Martin, and Betsy Davis. A user account registration system for a large (heterogeneous) UNIX network. In USENIX Association [USE88g], pages 155–161. ISBN ???? LCCN ????

Prata:1999:UPP

[PMW^{+}99] Stephen Prata, Donald (Donald W.) Martin, Mike Wessler, Dan Wilson, and Mitchell Waite. Unix primer plus. Macmillan Computer Pub., Indianapolis, IL, third edition, 1999. ISBN 0-672-31797-4. x + 350 pp. LCCN QA76.76.O63 P737 1999.

Prata:2000:UPP

[PMW^{+}00] Stephen Prata, Donald Martin, Michael Wessler, Daniel Wilson, and Mitchell Waite. UNIX primer plus. Waite Group Press, Corte Madera, CA, USA, third edition, 2000. ISBN 1-57169-165-0 (paperback). x + 350 pp. LCCN QA76.6 .W3185 2000.

Palmer:2004:SAO

[PN04] Brandon Palmer and Jose Nazario. Secure Architectures with OpenBSD. Addison-Wes-ley, Reading, MA, USA, 2004. ISBN 0-321-19366-0 (paperback). xix + 519 pp. LCCN QA76.76.O63 P335 2004. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0413/2004001163.
html.

Park:2002:RTM

[PNK02] Nam-Sup Park, Sang-Jun Nam, and Tai-Yun Kim. Realtime multimedia data transmission module based on Linux. Lecture Notes in Computer Science, 2343:504– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2343/23430504.htm;
http://link.springer.de/
link/service/series/0558/
papers/2343/23430504.pdf.

Poepping:1987:BRU

[Poe87] Mark Poepping. Backup and restore for UNIX systems. In USENIX Association [USE87c], pages 10–11. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

Pogue:2002:MX

[Pog02a]  David Pogue.  Mac OS

X. Missing manual. Pogue Press/O’Reilly and Associates, Sebastopol, CA, USA, second edition, 2002. ISBN 0-596-00450-8. xii + 712 pp. LCCN QA76.76.O63 P634 2002.

Pogue:2002:MXM

[Pog02b] David Pogue. Mac OS X: the missing manual. Missing manual. Pogue Press/O’Reilly and Associates, Sebastopol, CA, USA, 2002. ISBN 0-596-00082-0. xii + 583 pp. LCCN QA76.76.O63 P634 2002.

Pogue:2005:MXT

[Pog05] David Pogue. Mac OS X, Tiger edition: the missing manual. Missing manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00941-0 (paperback). xvi + 847 pp. LCCN QA76.76.O63 P634 2005.

Pogue:2007:MXL

[Pog07] David Pogue. Mac OS X Leopard edition: the missing manual. Missing manual. Pogue Press/O’Reilly, Sebastopol, CA, USA; Beijing, China, 2007. ISBN 0-596-52952-X. xvi + 893 pp. LCCN QA76.76.O63 P634428

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2004558625-d.html;
http://www.oreilly.com/
catalog/9780596529529.

Pogue:2017:TID

[Pog17] David Pogue. TechnoFiles:

The iOS 11 to-do list. Scientific American, 316(6):26, June 2017. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL  http://www.nature.
com/scientificamerican/journal/
v316/n6/full/scientificamerican0617-
26.html;   http://www.nature.
com/scientificamerican/journal/
v316/n6/pdf/scientificamerican0617-
26.pdf.

Polytron:1988:P

[Pol88] Polytron Corporation, Beaverton, OR, USA. PolyAWK, 1988.

Polk:1993:ACW

[Pol93] W. Timothy Polk. Approximating Clark-Wilson access triples with basic UNIX commands. In USENIX Association [USE93g], pages 145–

154. ISBN 1-880446-55-3. LCCN QA 76.9 A25 U54

1993.  URL http://www.
usenix.org/publications/
library/proceedings/sec4/.

Peek:1994:UPT

[POL94] Jerry D. Peek, Tim O’Reilly, and Michael Kosta Loukides.

UNIX power tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1994. ISBN 1-56592-260-3. xliii + 1073 pp. LCCN QA76.76.O63 P44 1993. URL http:/

/www.oreilly.com/catalog/
9781565922600.

Peek:1997:UPT

[POL^{+}97] Jerry Peek, Tim O’Reilly, Mike Loukides, Linda Mui, et al. UNIX Power Tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1997. ISBN 1-56592-260-3. xliii + 1073 pp. LCCN QA76.76.O63 P44 1997. US$59.95. URL http://www.

ora.com/catalog/upt2/;
http://www.oreilly.com/catalog/
upt2.

Peek:1993:UPT

[POLo93] Jerry Peek, Tim O’Reilly, Mike Loukides, and other authors of the Nutshell handbooks. UNIX Power Tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 0-553-35402-7, 0-679-79073-X (with CD ROM). xlii + 1119 pp. LCCN QA76.76.O63 P44 1993. US$59.95 (with CD ROM). Highly recommended. Simply great!!! “[It] contains literally thousands of tips, scripts, and techniques that make using UNIX easier, more effective, and even more fun.” With a CD-ROM disk containing PD programs and shell scripts …. The shell scripts can also be obtained by anonymous ftp from

ftp.uu.net (as /published/
oreilly/power_tools/unix/
upt.mar93.tar.Z).

Pomerantz:2004:LHU

[Pom04] Gregory M. Pomerantz. The legal history of UNIX and free software. Web document., June 19, 2004. URL http:

//wiki.tuhs.org/lib/exe/
fetch.php?media=publications:
theses:gmp_thesis.pdf.

Poniatowski:1994:HUS

[Pon94] Marty Poniatowski. The HP-UX systems administrator’s “how to” book. Hewlett– Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-099821-4. xvi + 245 pp. LCCN QA76.8.H48 P66 1994.

Poniatowski:1996:HUX

[Pon96] Marty Poniatowski. HP-UX 10.x system administration: “how to” book. Hewlett Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-125873-7. xx + 383 pp. LCCN QA76.8.H48 P658 1996.

Ponder:1997:OUD

[Pon97a] Carl Ponder. Organizing UNIX directories as lattices. Operating Systems Review, 31(4):72– 77, October 1997. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Poniatowski:1997:LHU

[Pon97b] Marty Poniatowski. Learning the HP-UX operating system. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-258534-0. xxi + 313 pp. LCCN QA76.76.O63 P65 1997.

Poniatowski:1997:WNH

[Pon97c] Marty Poniatowski. The Windows NT and HP-UX system administrator’s “how-to” book. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-861709-0 (paperback). xxvi + 576 pp. LCCN QA76.76.O63 P655 1997.

Poniatowski:1998:HUS

[Pon98] Marty Poniatowski. HP-UX System Administration Handbook and Toolkit. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-905571-1. xxxii + 691 pp. LCCN QA76.76.O63P648 1998.

                                        URL http://www.phptr.com/
                                        ptrbooks/ptr_0139055711.
                                        html.

Poniatowski:1999:HUXb

[Pon99a] Marty Poniatowski. The HP-UX 11.x System Administration Handbook and Toolkit. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-012514-8. xxxvi + 1147 pp. LCCN QA76.76.O63P647 1999.

URL http://www.phptr.com/
ptrbooks/ptr_0130125148.

html. Includes CD-ROM.

Poniatowski:1999:HUXa

[Pon99b] Marty Poniatowski. HP-UX 11.x System Administration “How To” Book. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-012515-6. xxxi + 817 pp. LCCN QA76.76.O63 P65 1999. URL http:

//www.phptr.com/ptrbooks/
ptr_0130125156.html.

Poniatowski:2000:UUH

[Pon00] Marty Poniatowski. UNIX User’s Handbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-027019-9. xxxix + 1366 pp. LCCN QA76.76.O63 P654 2000. Includes CD-ROM.

Poniatowski:2001:HUS

[Pon01]  Marty Poniatowski.  HP-

UX 11i System Administration Handbook and Toolkit. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN ???? 1872 (est.) pp. LCCN 0-13-060081-4. US$54.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0130600814.html.  In-

cludes CD-ROM.

Poniatowski:2002:UUH

[Pon02a] Martin Poniatowski. UNIX user’s handbook. Hewlett– Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2002. ISBN 0-13-065419-1 (paperback). xxxix + 1416 pp. LCCN QA76.76.O63 P654 2002.

Poniatowski:2002:HUV

[Pon02b] Marty Poniatowski. HP-UX Virtual Partitions. Hewlett– Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-035212-8. xxxii + 1012 pp. LCCN QA76.76.O63 P649 2002. US$49.99.

Poniatowski:2003:HUS

[Pon03] Marty Poniatowski. HP-UX 11i system administration handbook and toolkit. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2003. ISBN 0-13-101883-3 (paperback). xxxvi + 1390 pp. LCCN QA76.76.O63 P647 2003.

Poniatowski:2005:HUV

[Pon05a] Marty Poniatowski. HP-UX 11i Version 2 system administration: HP Integrity and HP 9000 servers. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-192759-0. xxiv + 643 pp. LCCN QA76.76.O63 P6477 2005.

Poniatowski:2005:LHI

[Pon05b] Marty Poniatowski. Linux on HP Integrity Servers: system administration for Itaniumbased systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-140000-2. xxvi + 332 pp. LCCN QA76.76.O63 P652 2005.

Poole:1991:TUC

[Poo91] Gary Andrew Poole. The Top 10 UNIX Companies of 1991. UNIX/World, 8(12):51–??, December 1991. ISSN 0739-5922.

Poston:1988:HPF

[Pos88] Alan Poston. A High Performance File System for UNIX. In USENIX Association [USE88f], pages 215–226. ISBN ???? LCCN ????

Povey:1999:EWF

[Pov99] Dean Povey. Enforcing wellformed and partially formed transactions for UNIX. In USENIX [USE99b], page ?? ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999. URL

http://db.usenix.org/publications/
library/proceedings/sec99/
povey.html.

Powell:1984:POC

[Pow84a]  Michael L. Powell.  A

portable optimizing compiler for Modula-2. ACM SIGPLAN Notices, 19(6):310–318, June 1984. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Powell:1984:UMS

[Pow84b]  Michael L. Powell.  Using

Modula-2 for system programming with UNIX. In Software Tools Users Group [Sof84], pages 119–132. LCCN QA76.8.U65 U83 1984.

Powell:2000:PK

[Pow00] Dennis E. Powell. Practical KDE. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2216-X. xi + 702 pp. LCCN QA76.9.U83 P68 2000. Technical editor: Bob Bernstein.

Pozgaj:1983:UCA

[Poz83] Steve Pozgaj. UNIX for the Computer Automation 4/95. In Association [Ass83], pages 307– ?? Abstract only.

Pike:1985:FN

[PP85] Rob Pike and David L. Presotto. Face the nation. In USENIX Association [USE85d], pages 81–86. LCCN QA76.8.U65 U8 1985.

Powers:2002:UPT

[PPO^{+}02] Shelley Powers, Jerry Peek, Tim O’Reilly, Mike Loukides, et al. Unix Power Tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2002. ISBN 0-596-00330-7. xxxviii + 1116 pp. LCCN QA76.76.O63 P44 2003. US$69.95. URL http://www.

                                        oreilly.com/catalog/upt3.

Powers:2003:UPT

[PPO^{+}03] Shelley Powers, Jerry Peek, Tim O’Reilly, Mike Loukides, et al. Unix Power Tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2003. ISBN 0-596-00330-7. xxxviii + 1116 pp. LCCN QA76.76.O63 P44 2003. US$69.95. URL http://www.

oreilly.com/catalog/upt3.

Pike:19xx:PEP

[PPT^{+}xx] Rob Pike, Dave Presotto, Ken Thompson, Howard Trickey, Tom Duff, and Gerard Holzmann. Plan 9: The early papers. Computing Science Technical Report 158a, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 1 pp.

Pike:19xx:PSW

[PPTHxx] Rob Pike, Dave Presotto, Ken Thompson, and Gerard Holzmann. Process sleep and wakeup on a shared-memory multiprocessor. Computing Science Technical Report 158f, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 6 pp.

Pike:19xx:PBL

[PPTTxxa] Rob Pike, Dave Presotto, Ken Thompson, and Howard Trickey. Plan 9 from Bell Labs. Computing Science Technical Report 158b, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 9 pp.

Presotto:19xx:PDS

[PPTTxxb] Dave Presotto, Rob Pike, Ken Thompson, and Howard Trickey. Plan 9, a distributed system. Computing Science Technical Report 158c, AT&T Bell Laboratories, Murray Hill, NJ, USA, ?? ??, 19xx. 7 pp.

Passos:2021:SFS

[PQM^{+}21] L. Passos, R. Queiroz, M. Mukelabai, T. Berger, S. Apel, K. Czarnecki, and J. A. Padilla. A study of feature scattering in the Linux kernel. IEEE Transactions on Software Engineering, 47(1):146–164, 2021. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Perez:1976:USC

[PR76] C. D. Perez and T. M. Raleigh. UNIX system call measurements. Technical Memorandum 1113 (MF 76-8234-79), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 16, 1976. ?? pp.

Panzieri:1985:IUD

[PR85a] F. Panzieri and B. Randell. Interfacing UNIX to data communications networks. IEEE Transactions on Software Engineering, SE-11(10): 1016–1032, October 1985. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1701916.

Presotto:1985:ICE

[PR85b] D. L. Presotto and D. M. Ritchie. Interprocess communication in the Eighth Edition Unix system. In USENIX Association [USE85d], pages 309– 316. LCCN QA76.8.U65 U8 1985.

Presotto:1985:ICN

[PR85c] David L. Presotto and Dennis M. Ritchie. Interprocess communication in the Ninth Edition Unix system. Report, ATT Bell Laboratories, Murray Hill, NJ 07974, USA, 1985. URL https:

//www.bell-labs.com/usr/
dmr/www/ipcpaper.html.

Presotto:1990:ICN

[PR90] D. L. Presotto and D. M.

Ritchie. Interprocess communication in the Ninth Edition UNIX system. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Palmer:1996:PDA

[PR96a] Michael Palmer and Jeffrey M.

Russo. Parasight: Debugging and analyzing realtime applications under Digital UNIX. Digital Technical Journal of Digital Equipment Corporation, 7(4):101– 108, March 1996. CODEN DTJOEL. ISSN 0898-901X.

URL  ftp://ftp.digital.
com/pub/Digital/info/DTJ/

                                        v7n4/Parasight_Debugging_
                                        and_Analy_20apr1996DTJK07P8.
                                        ps;   http://www.digital.
                                        com:80/info/DTJK07;   http:
                                        //www.digital.com:80/info/
                                        DTJK07/DTJK07AH.HTM;   http:
                                        //www.digital.com:80/info/
                                        DTJK07/DTJK07P8.PS;   http:
                                        //www.digital.com:80/info/
                                        DTJK07/DTJK07PF.PDF;   http:
                                        //www.digital.com:80/info/
                                        DTJK07/DTJK07SC.TXT.

Purcell:1996:LBG

[PR96b] John Purcell and Amanda Robinson, editors. The Linux Bible: The GNU Testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, fourth edition, 1996. ISBN 1-883601-10-X (??), 1-883601-20-7. 1176 (??) pp. LCCN QA76.76.O63 L5458 1996. US$39.95. Includes CD ROM.

Pagano:2023:SSE

[PRC^{+}23] Francesco Pagano, Andrea Romdhana, Davide Caputo, Luca Verderame, and Alessio Merlo. SEBASTiAn: a static and extensible black-box application security testing tool for iOS and Android applications. SoftwareX, 23(??): ??, July 2023. CODEN ???? ISSN 2352-7110. URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S2352711023001449.

Preston:1999:UBR

[Pre99] W. Curtis Preston. UNIX Backup & Recovery. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-642-0. xix + 707 pp. LCCN QA76.9.D348 P73 1999. US$36.95. URL http:/

/www.oreilly.com/catalog/
unixbr.

Presotto:19xx:MSP

[Prexx] David Leo Presotto. Multiprocessor streams for Plan 9. Computing Science Technical Report 158e, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 10 pp.

Preston:2007:BR

[Pre07] W. Curtis Preston. Backup and recovery. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-10246-1 (paperback). xxviii + 729 pp. LCCN QA76.9.B32 P74

2007. URL http://www.loc.
gov/catdir/enhancements/
fy0715/2007273944-d.html;
 http://www.oreilly.com/
catalog/9780596102463/index.
html.

Prieve:1978:UP

[Pri78] B. G. Prieve. UNIX/370 — a proposal. Technical Memorandum 1310, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 28, 1978. ?? pp.

Prince:1983:SU

[Pri83] V. Prince. Le système Unix.

Editests, Paris, France, 1983. ISBN 2-86688-003-X. 128 (est.) pp. LCCN ????

Pritchard:1999:LRC

[Pri99] Kara J. Pritchard. Linux Red Hat Certified Engineer Exam Cram. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN 1-57610-487-7. xxvii + 380 pp. LCCN QA76.3 .P75 1999. US$29.99.

Pruett:2002:RUC

[Pru02] Mark Pruett. RDB — a UNIX command-line database. Sys Admin: The Journal for UNIX Systems Administrators, 11(1): 33–36, January 2002. CODEN SYADE7. ISSN 1061-2688.

Pearlman:1977:PUB

[PS77] M. E. Pearlman and S. H. Strauss. PWB/UNIX beginner’s course (B.2). Technical Memorandum 1188, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1977. ?? pp.

Petrella:1979:PUE

[PS79] M. J. Petrella and A. L. Sabsevitz. PWB/UNIX Edition 2.0 preparation. Technical Memorandum 1403, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1, 1979. ?? pp.

Patel:1982:UEA

[PS82] Sanand Patel and Richard Sniderman. UNIX emulation, again. In Usr Group [Usr82], pages 248–?? Abstract only.

Pabrai:1994:XWS

[PS94] Uday O. Pabrai and Hemant T. Shah. X Window System user’s guide. The Artech House telecommunications library. Artech House Inc., Boston, MA, USA, 1994. ISBN 0-89006-740-6. xv + 236 pp. LCCN QA76.76.W56P33 1994.

Pandey:2007:SCM

[PS07] Nirved Pandey and G. K.

Sharma. Startup comparison for message passing libraries with DTM on Linux clusters. The Journal of Supercomputing, 39(1):59–72, January 2007. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=39&issue=1&spage=
59.

Pittman:2009:SOS

[PS^{+}09] Gregory Pittman, Christoph Schäfer, et al., editors. Scribus: open source desktop publishing: the official manual. FLES Books, Lostwithiel, Cornwall, UK, 2009. ISBN 0-9560780-0-1 (paperback). 439 pp. LCCN Z253.532 S33 S434 2009.

Paludo:2022:NAL

[PS22] Rogério Paludo and Leonel Sousa. NTT architecture for a Linux-ready RISC-V fullyhomomorphic encryption accelerator. IEEE Transactions on Circuits and Systems I: Regular Papers, 69(7):2669–2682, 2022. ISSN 1549-8328 (print), 1558-0806 (electronic).

Pike:1993:HWD

[PT93] Rob Pike and Ken Thompson. Hello world (distributed OS text format). In USENIX Association [USE93f], pages 43– 50. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

Peek:1998:LUO

[PTS98] Jerry D. Peek, Grace Todino, and John Strang. Learning the UNIX operating system. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 1998. ISBN 1-56592-390-1. xi + 92 pp. LCCN QA76.76.O63 T62 1998. US$10.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        9781565923904;   http://
                                        www.oreilly.com/catalog/

lunix4. Additions and revisions by Jerry Peek.

Peek:2001:LUO

[PTS01] Jerry D. Peek, Grace Todino, and John Strang. Learning the UNIX operating system. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2001. ISBN 0-596-00261-0. xiii + 157 pp. LCCN QA76.76.O63 T62 2002.

Palix:2011:FLT

[PTS^{+}11] Nicolas Palix, Gaël Thomas, Suman Saha, Christophe Calvès, Julia Lawall, and Gilles Muller. Faults in Linux: ten years later. ACM SIGARCH Computer Architecture News, 39(1):

305–318, March 2011. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Palix:2014:FL

[PTS^{+}14] Nicolas Palix, Gael Thomas, Suman Saha, Christophe Calvès, Gilles Muller, and Julia Lawall. Faults in Linux 2.6. ACM Transactions on Computer Systems, 32(2):4:1–4:??, June 2014. CODEN ACSYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

Publishers:2011:LIQ

[Pub11] Vibrant Publishers. Linux interview questions you’ll most likely be asked. Interview questions you’ll most likely be asked. Vibrant Publishers, Erie, CO, USA, 2011. ISBN 1-4564-7381-6. ???? pp. LCCN ????

Pugh:1988:ASL

[Pug88a] Kenneth Pugh. ANSI standards list for UNIX. C Users Journal, 6(1):22–??, January 1988. ISSN 0898-9788.

Pugh:1988:IHT

[Pug88b] Kenneth Pugh. INIT.C in HOC6 from The UNIX Programming Environment. C Users Journal, 6(11):10–??, November 1988. ISSN 0898-9788.

Pugh:1988:UAO

[Pug88c] Kenneth Pugh. UNIX alternatives to overlays. C Users Journal, 6(2):9–??, February 1988. ISSN 0898-9788.

Pugh:1990:LUF

[Pug90a] Kenneth Pugh. Locking UNIX files. C Users Journal, 8(9): 109–??, September 1990. ISSN 0898-9788.

Pugh:1990:TMP

[Pug90b] Kenneth Pugh. malloc() problems under UNIX disappear on PC. C Users Journal, 8(2):69– ??, February 1990. ISSN 0898-9788.

Pugh:1990:RMF

[Pug90c] Kenneth Pugh. Reading MS-DOS files into a UNIX system. C Users Journal, 8(6):75–??, June 1990. ISSN 0898-9788.

Pugh:1992:SFS

[Pug92] Kenneth Pugh. Setting file sizes under UNIX. C Users Journal, 10(2):109–??, February 1992. ISSN 0898-9788.

Purdy:2004:LIP

[Pur04]  Gregor N. Purdy.  Linux

iptables: pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2004. ISBN 0-596-00569-5. iii + 91 pp. LCCN QA76.76.O63 P873 2004. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596005696.

Puttress:1986:CB

[Put86] J. J. Puttress. The C browser. Report 11229-861017-19TMS, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 17, 1986. ?? pp.

Pate:2000:UFE

[PV00] Steve D. Pate and Fred Van Den Bosch. UNIX Filesystems: Evolution, Design, and Implementation, volume 10 of Veritas Series. Wiley, New York, NY, USA, 2000. ISBN 0-471-16483-6, 0-471-45675-6 (e-book), 1-280-36600-1. xxv + 443 pp. LCCN QA76.76.O63P3773 200.

URL http://proquest.tech.
safaribooksonline.de/9780471456759.

Petschenik:1977:STT

[PW77] N. H. Petschenik and R. P.

Woodall. A system testing technique using a universal LEAP script. Technical Memorandum 1238 (TM 77-9426-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 31, 1977. ?? pp.

Petschenik:1979:TTR

[PW79] N. H. Petschenik and R. P.

Woodall. TERSE — a technique for reporting small data bases. Technical Memorandum 1496 (TM 79-9252-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 31, 1979. ?? pp.

Pike:1985:HN

[PW85] Rob Pike and P. J. Weinberger.

The hideous name. In USENIX Association [USE85d], pages 563–568. LCCN QA76.8.U65 U8 1985.

Pendry:1994:AAR

[PW94] Jan-Simon Pendry and Nick Williams. AMD: The 4.4BSD Automounter Reference Manual. USENIX Association, Berkeley, CA, USA, 1994. ISBN ???? ???? pp. LCCN ????

Patsis:2000:SJS

[PWC00] Peter Patsis, Zhonghua Wu, and Xingzhi Chen. Shi jian da shi: UNIX awk ho sed pien ch‘eng p‘ien. (Chinese) [UNIX awk and sed programmer’s interactive workbook]. Dian zi gong ye chu ban she, Beijing, China, 2000. ISBN 7-5053-5632-1. 5 + 6 + 405 pp. LCCN QA76.76.O63.

Petrovsky:2001:LDB

[PWN01] Michele Petrovsky, Stephen Wysham, and Mojo Nichols. Linux database bible. Hungry Minds, 909 Third Avenue, New York, NY 10022, USA, 2001. ISBN 0-7645-4641-4. xxi + 715 pp. LCCN QA76.76.O63 P5288 2001. URL http:

//www.loc.gov/catdir/toc/
wiley024/2001092731.html.

Pyne:1984:MUB

[PY84a] T. Scott Pyne and Joseph S. D. Yao. MIPS: a UNIX-Based microcomputer message switching system. In USENIX Association [USE84b], pages 99–114. ISBN none. LCCN QA76.8.U65 U55 1984.

Pyne:1984:MUM

[PY84b] T. Scott Pyne and Joseph S. D. Yao. MIPS: A UNIX-Based microcomputer message switching system. In USENIX Association [USE84b], pages 99–114. ISBN none. LCCN QA76.8.U65 U55 1984.

Quinter:2011:IPS

[Q^{+}11] Dino Quinter et al. IBM PowerHA SystemMirror 7.1 for AIX. IBM redbooks. IBM Corporation, Poughkeepsie, NY, USA, 2011. ISBN 0-7384-3512-0. xvi + 534 pp. LCCN QA76.77 .I36 2011.

Querel:1987:SIP

[QBdR87] Stephane Querel, Bruno Borghi, and Daniel de Rauglaudre. SM-SCRIPT: An interpreter for the PostScript language under UNIX. In Earnshaw [Ear87], chapter 2.2, pages 27–39. ISBN 0-387-96527-0, 3-540-96527-0. LCCN Z286.E43 W67 1987. US$29.50.

Quigley:2001:CLS

[QH01] Ellie Quigley and Scott Hawkins.

The Complete Linux Shell Programming Training Course. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-040676-7. ???? pp. LCCN ???? US$109.99. URL http:

//www.phptr.com/ptrbooks/
ptr_0130406767.html.

Quong:1991:LPI

[QL91] Russell W. Quong and Mark A.

Linton. Linking programs incrementally. ACM Transactions on Programming Languages and Systems, 13(1):1– 20, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0164-
0925/102804.html.

Quercia:1989:XWS

[QO89] Valerie Quercia and Tim O’Reilly. X Window System user’s guide: for Version 11 of the X Window System, volume 3 of The definitive guides to the X Window System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second, revised edition, July 1989. ISBN 0-937175-36-6. xxi + 546 pp. LCCN QA76.76.W56 D44 v.3 1989.

Quercia:1990:XWS

[QO90] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1990. ISBN 0-937175-14-5. xxvi + 723 pp. LCCN QA76.76.W56 Q83 1990. URL http://

proquestcombo.safaribooksonline.
com/9780937175149;   http:
//www.oreilly.com/catalog/
9780937175149.

Quercia:1991:XWS

[QO91] Valerie Quercia and Tom O’Reilly. X Window System User’s Guide: OSF/Motif edition. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-61-7. xxvi + 709 pp. LCCN QA76.76.W56 Q4 1991.

Quercia:1993:XWSa

[QO93a] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide, volume 3. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, May 1993. ISBN 1-56592-014-7. xxx + 835 pp. LCCN QA76.76.W56D43 1993.

Quercia:1993:XWSb

[QO93b] Valerie Quercia and Tim O’Reilly. X Window System User’s Guide: OSF/Motif 1.2 Edition, volume 3M. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, May 1993. ISBN 1-56592-015-5. xxx + 835 pp.

Quarterman:1985:EUS

[QSP85] John S. Quarterman, Abraham Silberschatz, and James L. Peterson. 4.2BSD and 4.3BSD as examples of the UNIX system. ACM Computing Surveys, 17(4):379–418, December 1985. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0360-
0300/6043.html.

Quarterman:1993:RIB

[Qua93] John S. Quarterman. Recent Internet Books, March 1993. This is a list of books related to using the Internet. The document (rfc1432.txt) can be obtained by anonymous ftp

                                        from /rfc at nic.ddn.mil. See

[Bee93a].

Quigley:1995:PE

[Qui95] Ellie Quigley. Perl by Example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-122839-0. xv + 358 pp. LCCN QA76.73.P22 Q53 1995. US$26.95.

Quigley:1997:USE

[Qui97] Ellie Quigley. UNIX shells by example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1997. ISBN 0-13-460866-6. x + 644 pp. LCCN QA76.76.O63 Q54 1997.

Quigley:1999:USE

[Qui99] Ellie Quigley. UNIX Shells by Example. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-021222-9. xiv + 654 pp. LCCN QA76.76.O63 Q54 1999. US$49.99.

Quigley:2000:LSE

[Qui00] Ellie Quigley. Linux shells by example. Open source technology series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-014711-7. xviii + 761 pp. LCCN QA76.76.O63 Q538 2000.

Quigley:2002:USE

[Qui02] Ellie Quigley. UNIX Shells by Example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2002. ISBN 0-13-066538-X (paperback). xix + 1015 pp. LCCN QA76.76.O63 Q54 2002.

Quigley:2005:USE

[Qui05] Ellie Quigley. UNIX shells by example. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, fourth edition, 2005. ISBN 0-13-147572-X. xxxv + 1150 pp. LCCN QA76.76.O63 Q54 2004.

Quarterman:1993:UPO

[QW93] John S. Quarterman and Susanne Wilhelm. UNIX, POSIX, and Open Systems — the Open Standards Puzzle. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-52772-3. xxix + 416 pp. LCCN QA76.76.O63 Q37 1993.

Russell:2002:SCI

[R^{+}02] Steve Russell et al., editors.

Server consolidation with the IBM eServer xSeries 440 and VMware ESX Server. IBM redbooks. IBM, International Technical Support Organization, ????, November 2002. ISBN 0-7384-2684-9. xiv + 222 pp. LCCN QA76.9.C55 S378 2002. Publication number SG24-6852-00.

Radnik:1979:IIC

[Rad79] M. K. Radnik. An interactive interface for C language programs. Technical Memorandum 1377 (TM 79-8224-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 14, 1979. ?? pp.

Raffenetti:1979:CSF

[Raf79] Richard C. Raffenetti. A comparative study of the Fortran development environment provided by the VAX/VMS and VAX/UNIX operating systems. Technical memorandum 346, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1979. iii + 29 pp. Reproduced by the National Technical Information Service.

Rago:1990:LNE

[Rag90] S. A. Rago. A look at the Ninth Edition Network file system. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Rago:1993:USV

[Rag93] Steven A. Rago. UNIX System V network programming. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56318-5. xv + 784 pp. LCCN 92-45276, QA76.76.O63 R34 1993. US$45.50.

Rain:1989:PIV

[Rai89] Mark Rain. Portable IPC on Vanilla Unix. ACM SIG-PLAN Notices, 24(5):47–56, May 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rajagopal:1998:WNU

[Raj98] Raj Rajagopal. Windows NT, UNIX, NetWare migration and coexistence: a professional’s guide. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 1998. ISBN 0-8493-1669-3. 249 pp. LCCN QA76.76.O63 R343 1998.

Rajagopal:2000:MOS

[Raj00] Raj Rajagopal, editor. Multi-Operating System Networking: Living with UNIX, NetWare, and NT. Best practices series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2000. ISBN 0-8493-9831-2. 1360 (est.) pp. LCCN QA76.76.O63 M8455 2000. US$79.95.

Raleigh:1975:EAC

[Ral75a] T. M. Raleigh. Explanation of abnormal conditions within the UNIX operating system. Technical Memorandum 1047 (MF 75-8234-28), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 17, 1975. ?? pp.

Raleigh:1975:ISS

[Ral75b] T. M. Raleigh. Introduction to scheduling and switching under UNIX. Technical Memorandum 1069 (TM 75-8234-7), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 20, 1975. ?? pp.

Raleigh:1976:DPA

[Ral76a] T. M. Raleigh. DEAD — a program for analyzing UNIX system crashes. Technical Memorandum 1093 (MF 76-8234-30), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 30, 1976. ?? pp.

Raleigh:1976:SSU

[Ral76b] T. M. Raleigh. Scheduling and switching under UNIX algorithms and implementation. Technical Memorandum 1126 (TM 76-8234-18), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1, 1976. ?? pp.

Raleigh:1976:TPC

[Ral76c] T. M. Raleigh. TCMP — a program for comparing program text after a system crash. Technical Memorandum 1092 (MF 76-8234-29), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 27, 1976. ?? pp.

Raleigh:1977:LUS

[Ral77] T. M. Raleigh. Lecture on UNIX scheduler and switcher given by T. M. Raleigh at Whippany. Technical Memorandum 1214 (MF 77-8234-99), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 21, 1977. ?? pp.

Raleigh:1978:UMR

[Ral78] T. M. Raleigh. Upgrading MERT Release 0 with Modification 0.1 to MERT Release 0.1. Technical Memorandum 1287 (MF 78-8234-56), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 6, 1978. ?? pp.

Raleigh:1980:STM

[Ral80] T. M. Raleigh. Some timing measurements of Emts, message transfers and context switches for UNIX/RT. Technical Memorandum 1500, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 25, 1980. ?? pp.

Ramamurthy:1988:AMU

[Ram88] Gopalakrishnan Ramamurthy.

An analytical model for Unix systems. AT&T Technical Journal, 67(5):86–99, September/October 1988. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Ramsey:1994:AAA

[Ram94] Rick Ramsey. All About Administering NIS+. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1994. ISBN 0-13-309576-2. 500 pp. LCCN QA76.9.D3 R24 1993b. US$42.00. URL

http://www.sun.com/books/
catalog/ramsey/index.html.

Reizer:1994:UFM

[RAMP94] Neal R. Reizer, Gregory D. Abowd, B. Craig Meyers, and Patrick R. H. Place. Using formal methods for requirements specification of a proposed POSIX standard. In IEEE [IEE94b], pages 118–125. ISBN 0-8186-5480-5, 0-8186-5481-3. LCCN QA76.758.I5717 1994.

Rannenberg:1996:SUU

[Ran96] W. Rannenberg. Supporting Unicode on a UNIX system. In Unicode Consortium [Uni96a], pages C5–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

Rankin:2004:KHI

[Ran04] Kyle Rankin. Knoppix hacks: 100 industrial-strength tips and tools. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00787-6 (paperback). xx + 314 pp. LCCN QA76.76.O63 R368 2004.

Rankin:2005:KPR

[Ran05] Kyle Rankin. Knoppix pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-10075-2. vii + 84 pp. LCCN QA76.76.O63 R36 2005.

Rankin:2006:LMH

[Ran06] Kyle Rankin. Linux multimedia hacks: tips and tools for taming images, audio and video. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10076-0. xviii + 310 pp. LCCN QA76.575; QA76.575 .R345 2006eb. URL http:/

/www.oreilly.com/catalog/
9780596100766.

Rash:2007:LFA

[Ras07] Michael Rash. Linux firewalls: attack detection and re-

sponse with iptables, psad,

and fwsnort. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-141-7. xix + 308 pp. LCCN QA76.9.A25 R36 2007. URL http:/

/www.oreilly.com/catalog/
9781593271411.

Rautiainen:2001:TLM

[Rau01] Sami Rautiainen. Travelling with Linux malware: Is Linux security for real? Information Security Technical Report, 6(4):58–64, December 1, 2001. CODEN ISTRFR. ISSN 1363-4127 (print), 1873-605X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S136341270100406X.

Raymond:1991:NHD

[Ray91]  Eric Raymond.  The New

Hacker’s Dictionary. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-68069-6. xx + 433 pp. LCCN PN6231.E4 H3 1991. This book corresponds to version 2.9.6 of the on-line jargon file. The latest (at the time of writing) is version 2.9.12 (jargon2912. txt.z) which is available by anonymous ftp from prep.

ai.mit.edu (in /pub/gnu)
or wuarchive.wustl.edu (in

mirrors/gnu). Changes since the publication of this book can be found in the file jargonupd.z. (*.z are files compressed by GNU zip (gzip)).

Raymond:1998:LU

[Ray98] Eric Raymond, editor. Linux Undercover: Linux Secrets from the LDP. Red Hat Software, Inc., Westport, CT, USA, 1998. ISBN 1-888172-05-3. 2018 pp. LCCN ???? UK£34.99, US$39.99. URL http://

genesis.ukpost.com/undercover.
html.

Ray:1999:STY

[Ray99a] William Ray. Sams teach yourself Unix in 10 minutes. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-672-31523-8. vii + 210 pp. LCCN QA76.8.U65 R39 1999.

Raymond:1999:ILO

[Ray99b]  Eric Raymond.  Interview:

Linux and open-source success. IEEE Software, 16(1):85–89, January/February 1999. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1999/pdf/
s1085.pdf.

Raymond:1999:CBM

[Ray99c] Eric S. Raymond. The Cathedral & the Bazaar: Musings on Linux and Open Source by an Accidental Revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-724-9. xi + 268 pp. LCCN QA76.76.O63 R396 1999. US$19.95. URL http:/

/www.oreilly.com/catalog/
9781565927247;   http://
www.oreilly.com/catalog/
cb.

Ray:2001:MLS

[Ray01a] John Ray. Maximum Linux security. SAMS Publishing, Indianapolis, IN, USA, second edition, 2001. ISBN 0-672-32134-3. 800 pp. LCCN QA76.9.A25 M387 2001.

Raymond:2001:CBM

[Ray01b] Eric S. Raymond. The cathedral and the bazaar: musings on Linux and open source by an accidental revolutionary. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, revised edition, 2001. ISBN 0-596-00131-2. xiv + 241 pp. LCCN QA76.76.O63 R397 2001. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596001315.

Raymond:2004:AUP

[Ray04] Eric Steven Raymond. The Art of UNIX Programming. Addison-Wesley, Reading, MA, USA, 2004. ISBN 0-13-124085-4, 0-13-142901-9. xxxii + 525 pp. LCCN QA76.76.O63 R395 2003. US$39.99, CAN$60.99.

Raymond:2008:UHH

[Ray08] Eric Raymond. The Unix Hater’s Handbook, reconsidered. Web site, September

                                        30, 2008. URL http://esr.
                                        ibiblio.org/?p=538.

Rowe:1982:LNB

[RB82] L. A. Rowe and K. P. Birman. A local network based on the UNIX operating system. IEEE Transactions on Software Engineering, SE-8(2):137– 146, March/April 1982. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1702921.

Roberts:1992:UDG

[RB92a] Ralph Roberts and Mark Boyd. UNIX Desktop Guide to Emacs. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30171-7. xxiii + 504 pp. LCCN QA76.76.T49 R62 1992. US$27.95.

Roberts:UDG92

[RB92b] Ralph Roberts and Mark Boyd. UNIX Desktop Guide to Emacs. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-30171-7. xxiii + 504 pp. LCCN QA76.76.T49 R62 1992. US$27.95.

Robbins:2005:CSS

[RB05a] Arnold Robbins and Nelson H. F. Beebe. Classic Shell Scripting. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00595-4. xxii + 534 pp. LCCN QA76.76.O63 R633 2005. URL http:/

/www.oreilly.com/catalog/

shellsrptg/. Also available in Chinese [RB08], French [RB05b], German [RB06a], Japanese [RB06c], and Polish [RB06b] translations.

Robbins:2005:ISS

[RB05b] Arnold Robbins and Nelson H. F. Beebe. Introduction aux scripts shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 2-84177-375-2. xxii + 558 pp. LCCN ????

URL http://www.silicon.
fr/getarticle.asp?id=14015.

French translation of [RB05a] by Eric Jacoboni.

Robbins:2006:KSP

[RB06a] Arnold Robbins and Nelson H. F. Beebe. Klassische Shell-Programmierung: [automatisieren Sie Ihre Unix/Linux-Tasks]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 3-89721-441-

5. xxiii + 572 pp. LCCN QA76.76.O63 R563 2005. EUR 44.00; EUR 45.65 (AT).

URL  http://www.gbv.de/
dms/hebis-darmstadt/toc/
17645067X.pdf.  German

translation of [RB05a] by Kathrin Lichtenberg.

Robbins:2006:PSP

[RB06b] Arnold Robbins and Nelson H. F. Beebe. Programowanie skryptów powłoki. Helion, Gliwice, Poland, 2006. ISBN 83-246-0131-7. 557 + 2 pp. LCCN ???? URL http:/

/www.empik.com/b/o/19/f1/
19f16b85e0d75ae1d3a1e7062569fbb0.
jpg;   http://www.empik.
com/programowanie-skryptow-
powloki-ksiazka,360529,p.

Polish translation of [RB05a] by Przemysław Szeremiota.

Robbins:2006:SSS

[RB06c] Arnold Robbins and Nelson H. F. Beebe. Shōkai shieru sukuriputo. Orairī Japan, Tōkyō, Japan, 2006. ISBN 4-87311-267-2. 345 pp. LCCN ???? Japanese translation of

[RB05a] by Aoi Hyūga.

Robbins:2008:SJB

[RB08] Arnold Robbins and Nelson H. F. Beebe. Shell Jiao Ben Xue Xi Zhi Nan = Shell Script Study Guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2008. ISBN 7-111-25504-

6. vi + 494 pp. LCCN ???? 79.00 renminbi. Simplified Chinese translation of [RB05a].

Ricart:1999:STY

[RBHB99] Manuel Alberto Ricart, Grace Buechlein, Gregory Harris, and Laura Bulcher. Sams Teach Yourself Linux in 10 minutes. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31524-6. vii + 212 pp. LCCN QA76.76.O63R3918 1999. US$12.99.

Raves:1983:DDS

[RC83] William Raves and James Cassidy. Development of a digital simulation system in a UNIX environment. In Association [Ass83], pages 169–176.

Rubini:2001:LDD

[RC01] Alessandro Rubini and Jonathan Corbet. Linux Device Drivers. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, June 2001. ISBN 0-596-00008-1. xix + 564 pp. LCCN QA76.76.D49 R92 2001. US$39.95. URL http://

safari.oreilly.com/0596000081;
 http://www.oreilly.com/
catalog/9780596000080;
http://www.oreilly.com/catalog/
linuxdrive2.

Roylance:1975:BDT

[RDK75] G. Roylance, M. T. Dolan, and J. F. Kaiser. BLOSIM — a discrete time block simulator. Technical Memorandum 1079 (TM 75-1352-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 15, 1975. ?? pp.

Reagen:1991:FFS

[Rea91] Jeff Reagen. FS: a file status utility for UNIX. Dr. Dobb’s Journal of Software Tools, 16 (12):56–62, 96–98, December 1991. CODEN DDJOEB. ISSN 1044-789X.

Read:2011:OSC

[Rea11] Tim Read. Oracle Solaris Cluster essentials. Oracle Solaris system administration series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2011. ISBN 0-13-248940-6, 0-13-248622-9 (paperback). 434 pp. LCCN QA76.76.I57 R43 2011.

                                        URL http://proquestcombo.
                                        safaribooksonline.com/9780132489409.

Redfern:1997:GJP

[Red97] Heather Redfern. GO-Joe provides Java users access to UNIX X apps. SunServer, 11(9):6, September 1997. ISSN 1091-4986.

Reeves:1982:UAL

[Ree82a] Bill Reeves. UNIX at Lucasfilm Ltd. or does Darth Vader code in C? In Usr Group [Usr82], pages 29–?? Abstract only.

Reeves:1982:ULL

[Ree82b] Bill Reeves. UNIX at Lucasfilm Ltd. or does darth Vader code in C? In USENIX [USE82], pages 29–?? Abstract only.

Reeds:1988:SBU

[Ree88]  J. Reeds.  /bin/sh: the

biggest UNIX security loophole. Report 11217-840302-04TM, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1988. ?? pp.

Reeds:19xx:SIN

[Reexx] Jim Reeds. Secure IX network. Computing Science Technical Report 163g, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 8 pp.

Reed:2006:OPP

[Ree06a] Jeremy C. Reed, editor. The OpenBSD PF Packet Filter Book: PF for NetBSD, FreeBSD, DragonFly, and OpenBSD. Reed Media Services, ????, 2006. ISBN 0-9790342-0-5. 193 (est.) pp. LCCN ???? US$19.90. URL

http://www.reedmedia.net/
books/pf-book/.

Reese:2006:CSL

[Ree06b] Greg Reese. C++ Standard Library practical tips. Charles River Media, Inc., Hingham, MA, USA, 2006. ISBN 1-58450-400-5 (paperback). xviii + 542 pp. LCCN QA76.73.C153 R44

2006. URL http://www.loc.
gov/catdir/toc/ecip0513/

2005015845.html. Includes

CD-ROM.

Regelson:1978:SSU

[Reg78] K. L. Regelson. Simulator SM80 user manual. Technical Memorandum 1275 (TM 78-8234-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 15, 1978. ?? pp.

Regupathy:2011:BYL

[Reg11] Rajaram Regupathy. Bootstrap yourself with Linux-USB strap: design, develop, debug, and validate embedded USB. Cengage Learning, Boston, MA, USA, 2011. ISBN 1-4354-5786-2. ???? pp. LCCN ????

Rehman:2000:HCH

[Reh00] Rafeeq Ur Rehman. HP certified: HP-UX system administration. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-018374-1. xli + 789 pp. LCCN QA76.76.O63 R435 2000.

Rehman:2003:IDS

[Reh03] Rafeeq Ur Rehman. Intrusion Detection with SNORT: Advanced IDS Techniques Using SNORT, Apache, MySQL, PHP, and ACID. Bruce Perens’ Open source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2003. ISBN 0-13-140733-3. xii + 263 pp. LCCN TK5105.59 .R44 2003. US$39.99.

Rehman:2005:HUC

[Reh05] Rafeeq Ur Rehman. HP-UX CSA: official study guide and desk reference. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2005. ISBN 0-13-144854-4. xxxviii + 1031 pp. LCCN QA76.76.O63 R4352 2005.

Reiser:1977:EDO

[Rei77] John F. Reiser. EDTV — a display-oriented text editor for UNIX. Technical Memorandum 1177 (TM 77-1353-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1, 1977. ?? pp.

Reiser:1978:CPP

[Rei78] John F. Reiser. Cumulative program profiling. Technical Memorandum 1286 (TM 78-1353-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1, 1978. ?? pp.

Reid:1986:LUB

[Rei86] Brian Reid. Lessons from the UNIX breakins at Stanford. ACM SIGSOFT Software Engineering Notes, 11(5):29–35, October 1986. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Reichard:1993:PLD

[Rei93] Kevin Reichard. Power of … Desqview/X. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1993. ISBN 1-55828-257-2. 436 pp. US$27.95.

Reid:1997:WUM

[Rei97]  Jim Reid.  When UNIX

met air traffic control, 1997.

URL  http://www.usenix.
org/publications/library/
proceedings/lisa97/invited_
talks/reid/reid.pdf;   http:
//www.usenix.org/publications/
library/proceedings/lisa97/
invited_talks/reid/reid.

ps. Unpublished talk at the Eleventh Systems Administration Conference (LISA ’97) October 26–31, 1997 San Diego, California, USA.

Reichard:1998:UB

[Rei98] Kevin Reichard. UNIX: the basics. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1998. ISBN 1-55828-583-0. xiii + 273 pp. LCCN QA76.76.O63 R446 1998.

Reid:2003:SSE

[Rei03] Jason Reid. Secure Shell in the Enterprise. Sun blueprints. Sun Microsystems Press, Palo Alto, CA, USA, 2003. ISBN 0-13-142900-0 (paperback). xxiii + 198 pp. LCCN QA76.76.O63 R448 2003. US$39.00.

Remde:1973:CCS

[Rem73] J. R. Remde. A C compiler for the SEL computer. Technical Memorandum 1237 (TM 78-1228-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 13, 1973. ?? pp.

Remy:1997:PSH

[Rem97] Martin Remy. Portable signal handling under UNIX. C/C++ Users Journal, 15(3):23–??, March 1997. CODEN CCU-JEX. ISSN 1075-2838.

Requa:1985:UKN

[Req85] Joseph E. Requa. UNIX kernel networking support and the LINCS communications architecture. In USENIX Association [USE85c], pages 98–103.

Reznick:1994:TU

[Rez94] Larry Reznick. Time for UNIX.

Sys Admin: The Journal for UNIX Systems Administrators, 3(3):65–??, May/June 1994. CODEN SYADE7. ISSN 1061-2688.

Reznick:1995:HUA

[Rez95] Larry Reznick. Hiding UNIX applications in utility wrappers. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):68–??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Rinaldo:1993:ERH

[RF93] Frank J. Rinaldo and Matthew R.

Fausey. Event reconstruction in high-energy physics. Computer, 26(6):68–77, June 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reichard:1997:UPE

[RFJ97] Kevin Reichard and Eric Foster-Johnson. UNIX in plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1997. ISBN 1-55828-549-0 (paperback). v + 438 pp. LCCN QA76.76.O63 R4448 1997.

Reichard:1999:UPE

[RFJ99] Kevin Reichard and Eric Foster-Johnson. UNIX in plain English. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, third edition, 1999. ISBN 0-7645-7011-0. xx + 379 pp. LCCN QA76.76.O63 R44 1999.

Rodriguez:2005:LKP

[RFS05] Claudia Salzberg Rodriguez, Gordon Fischer, and Steven Smolski. The Linux Kernel primer: a top-down approach for x86 and PowerPC architectures. Prentice Hall open source software development series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2005. ISBN 0-13-118163-7 (paperback). ???? pp. LCCN QA76.76.O63 R633 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip0514/2005016702.
html;   http://www.phptr.
com/title/0131181637.

Rodriguez:2006:LKP

[RFS06] Claudia Salzberg Rodriguez, Gordon Fischer, and Steven Smolski. The Linux Kernel primer: a top-down approach for X86 and PowerPC architectures. Prentice Hall open source software development series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-118163-7 (paperback). xxvii + 616 pp. LCCN QA76.76.O63 R633 2006.

Russell:1991:CSB

[RG91] Deborah Russell and G. T.

Gangemi Sr. Computer Security Basics. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-71-4. xx + 441 pp. LCCN QA76.9.A25 R8 1991. US$29.95. A clear overview on many different security issues.

Rehn-Gostenmeier:2005:LE

[RGRG05] Gudrun Rehn-Göstenmeier and Ralph Rehn-Göstenmeier. Linux — Das Einsteigerseminar. (German) [Linux — The Beginner Seminar]. bhv, Bonn, Germany, 2005. ISBN 3-8266-7375-1. 448 pp. LCCN ????

Rauch:1995:IMM

[RH95] Wendy Rauch and Harvey Hindin. IBM makes more openness moves. 3X/400 Systems Management, 23(1):62–??, January 1, 1995. ISSN 1070-6097.

Rosen:2007:UCR

[RHK^{+}07] Kenneth H. Rosen, Douglas Host, Rachel Klee, James Farber, and Richard Rosinski. UNIX: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2007. ISBN 0-07-226336-9. LCCN

????  URL http://www.
unixreview.com/documents/
s=10125/ur0704c/.

Rossbach:2007:TUM

[RHP^{+}07] Christopher J. Rossbach, Owen S. Hofmann, Donald E. Porter, Hany E. Ramadan, Bhandari Aditya, and Emmett Witchel. TxLinux: using and managing hardware transactional memory in an operating system. Operating Systems Review, 41(6): 87–102, December 2007. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ribar:1989:SSD

[Rib89] John Ribar. A survey of system differences affecting ports between UNIX and VAX environments. C Users Journal, 7 (1):97–??, January 1989. ISSN 0898-9788.

Rice:1977:MSI

[Ric77] John R. Rice, editor. Mathematical software III: Proceedings of a symposium conducted by the Mathematics Research Center, the University of Wisconsin–Madison, March 28–30, 1977, number 39 in Publication of the Mathematics Research Center, the University of Wisconsin, Madison. Academic Press, New York, NY, USA, 1977. ISBN 0-12-587260-7. LCCN QA3 .U45 no. 39; QA297 .M36 1977. URL https:

//www.sciencedirect.com/
book/9780125872607/mathematical-
software.

Richards:1985:BDA

[Ric85] Robert R. Richards. A basic direct access method for UNIX. In USENIX Association [USE85c], pages 176–182.

Richardson:1990:LCL

[Ric90] C. Richardson. LispWorks:

a Common Lisp programming environment for Unix workstations. In Steels [Ste90a], pages 127–134. ISBN ???? LCCN ????

Rice:1993:FSH

[Ric93] Jim Rice, editor. Five steps to HP-UX. OnWord Press, Santa Fe, NM, 1993. ISBN 0-934605-24-6. xvi + 120 pp. LCCN QA76.8.H48 F58 1993.

Ricciuti:1994:MSI

[Ric94a] Mike Ricciuti. The mainframe as server: Is IBM totally bonkers — or brilliant? Datamation, 40(10):61–??, May 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Ricciuti:1994:USA

[Ric94b] Mike Ricciuti. UNIX Servers Approach PC Prices. Datamation, 40(14):65–??, July 15, 1994. CODEN DTMNAT. ISSN 0011-6963.

Rice:1994:HUQ

[Ric94c] Jim Rice, editor. The HP-UX quick reference. OnWord Press, Santa Fe, NM, 1994. ISBN 0-934605-28-8. xix + 211 pp. LCCN QA76.76.O63 H69 1994.

Rice:1995:HUU

[Ric95] Jim Rice. HP-UX user’s guide. OnWord Press, Santa Fe, NM, 1995. ISBN 0-934605-21-1. xii + 266 pp. LCCN QA76.76.O63 R52 1995.

Rice:1996:KTS

[Ric96] Ernest H. Rice, III. Kernel tuning for system I/O under SCO UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 5(1):33–??, January 1996. CODEN SYADE7. ISSN 1061-2688.

Richmond:1998:IUC

[Ric98] Charles M. Richmond. Internationalization of UNIX commands and utilities. In Unicode Consortium [Uni98b], page ?? ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc13/program.html.

Richardson:2013:GSB

[Ric13]  Matt Richardson.  Getting

started with BeagleBone. Maker Media, Sebastopol, CA, 2013. ISBN 1-4493-4537-9 (paperback), 1-4493-4536-0, 1-4493-4535-2 (e-book), 1-4493-4533-6 (e-book). xiii + 126 pp. LCCN TK7895.E42 R43 2013.

Richardson:2014:BEL

[Ric14] Matt Richardson. BeagleBone für Einsteiger: [Linux-basierte Elektronik-Projekte mit Python und JavaScript]. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2014. ISBN 3-95561-409-

3. xii + 134 pp. LCCN ????

Rieken:1993:USF

[Rie93] Bill Rieken. UNIX security facilities. Sys Admin: The Journal for UNIX Systems Administrators, 2(6):33–??, November/December 1993. CODEN SYADE7. ISSN 1061-2688.

Rieken:1994:UAF

[Rie94] Bill Rieken. UNIX accounting facilities. Sys Admin: The Journal for UNIX Systems Administrators, 3(1):107–??, January/February 1994. CODEN SYADE7. ISSN 1061-2688.

Ritchie:1971:DUT

[Rit71] Dennis M. Ritchie. Draft:

The UNIX time-sharing system. Web document, 1971.

URL https://www.tuhs.org/
Archive/Distributions/Research/
McIlroy_v0/UnixEditionZero-
Threshold_OCR.pdf. Undated

typescript, but believed to have been written in mid-1971. Neither references, nor date mentions, are given.

Ritchie:1974:CRM

[Rit74] D. M. Ritchie. C reference manual. Internal memorandum ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1974. ?? pp.

Ritchie:1977:CRM

[Rit77a] Dennis M. Ritchie. C reference manual. Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1, 1977. 32 pp.

Ritchie:1977:NIO

[Rit77b] Dennis M. Ritchie. A new input–output package (B.10). Technical Memorandum 1159, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1977. 6 pp.

Ritchie:1978:UTSb

[Rit78] Dennis M. Ritchie. UNIX time-sharing system: a retrospective. The Bell System Technical Journal, 57(6): 1947–1969, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1947.
                                        pdf.

Ritchie:1979:CPL

[Rit79a] Dennis M. Ritchie. The C programming language — reference manual. Technical Memorandum 1002, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Ritchie:1979:SU

[Rit79b] Dennis M. Ritchie. On the security of UNIX. Technical Memorandum 1138, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1979. ?? pp.

Ritchie:1979:PDF

[Rit79c] Dennis M. Ritchie. Protection of data file content. U.S.

Patent 4135240., January 16,

1979. URL https://patents.
google.com/patent/US4135240A/

. Patent filed 9 July 1973. Expires 16 January 1996.

Ritchie:1980:EUT

[Rit80]  Dennis W. Ritchie.  The

evolution of the Unix timesharing system. In Tobias [Tob80], pages 25–35. CODEN LNCSD9. ISBN 0-387-09745-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .S94 1979. URL http:

//cm.bell-labs.com/cm/cs/
who/dmr/hist.html. Spon-

sored by the Australian Atomic Energy Commission and the University of New South Wales.

Ritchie:1984:EUTa

[Rit84a] Dennis M. Ritchie. Evolution of the UNIX time-sharing system. Report, Bell Laboratories, Murray Hill, NJ 07974, USA, 1984. 13 pp. URL http:

//www.read.seas.harvard.
edu/~kohler/class/aosref/
ritchie84evolution.pdf;
https://www.bell-labs.com/
usr/dmr/www/hist.html.

Ritchie:1984:EUTb

[Rit84b] Dennis M. Ritchie. Evolution of the UNIX time-sharing system. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1577–1593, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ritchie:1984:RSR

[Rit84c] Dennis M. Ritchie. Reflections on software research. Communications of the Association for Computing Machinery, 27 (8):758–760, August 1984. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Ritchie:1984:SIO

[Rit84d] Dennis M. Ritchie. Stream input-output system. AT&T Bell Laboratories Technical Journal, 63(8 part 2):1897– 1910, October 1984. CODEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic).

Ritchie:1985:RSR

[Rit85] Dennis M. Ritchie. Reflections on software research: Can the circumstances that nurtured the UNIX project be produced again? UNIX Review, 3(1):28, 118–120, 122, January 1985. CODEN UNRED5. ISSN 0742-3136.

                                        URL http://www.tuhs.org/
                                        Archive/Documentation/Unix_
                                        Review/Berkeley_Unix_History.
                                        pdf.

Ritchie:1988:EUC

[Rit88a]  Dennis M. Ritchie.  Ex-

perience with Unicos on the Cray X-MP. Report, Bell Laboratories, Murray Hill, NJ 07974, September 1988. 4 pp. URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/earlyunicos.pdf.

Ritchie:1988:ECX

[Rit88b] Dennis M. Ritchie. Experiences with the Cray X/MP. Web site, 1988. URL https:

//www.bell-labs.com/usr/
dmr/www/cray.html.

Ritchie:1988:GFU

[Rit88c] Dennis M. Ritchie. A guest facility for Unicos. Report, Bell Laboratories, Murray Hill, NJ 07974, September 1988. 4 pp. URL https:

//www.bell-labs.com/usr/
dmr/www/unicos.pdf.

Ritchie:198x:TTU

[Rit8x] Dennis M. Ritchie. A tour through the UNIX C compiler. Report ??, Bell Laboratories, Murray Hill, NJ 07974, USA,

198x.  URL https://www.
jslite.net/cgi-bin/9995/
doc/tip/doc/old-ctour.pdf.

Undated report.

Ritchie:1990:SIO

[Rit90] D. M. Ritchie. A stream input-output system. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Ritchie:1993:DCL

[Rit93]  Dennis M. Ritchie.  The

development of the C language. ACM SIGPLAN Notices, 28(3):201–208, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        plan/154766/p201-ritchie/.

Ritchie:2002:OCS

[Rit02]  Dennis M. Ritchie.  Odd

comments and strange doings in Unix. Bell Labs Web site., 2002. URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/odd.html.

Ritchie:2012:RSR

[Rit12] Dennis M. Ritchie. Reflections on software research. Resonance, 17(8):810–816, August 2012. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Ritchie:1972:UAR

[Rit2 ] Dennis M. Ritchie. UNIX assembler reference manual. Technical Memorandum 1012, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1972 (??). ?? pp.

Ritchie:1975:US

[Rit5 ] Dennis M. Ritchie. The UNIX I/O system. Technical Memorandum 1005, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1975 (??). ?? pp.

Ritchie:1977:UTS

[Rit7 ] Dennis M. Ritchie. The UNIX time-sharing system — a retrospective (G.3). Technical Memorandum 1187, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1977 (??). ?? pp.

Ritchie:1979:EOU

[Rit9 ] Dennis M. Ritchie. The evolution of the Unix time-sharing system. Technical Memorandum 1470, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? 1979 (??). ?? pp.

Ritchie:1978:UTSc

[RJLK78] Dennis M. Ritchie, Steven C.

Johnson, Michael E. Lesk, and B. W. Kernighan. UNIX time-sharing system: The C programming language. The Bell System Technical Journal, 57(6):1991–2019, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1991.
pdf;   http://www.alcatel-
lucent.com/bstj/vol57-1978/
articles/bstj57-6-1991.pdf.

Rothman:2008:MXU

[RJR08] Ernest E. (Ernest Eric) Rothman, Brian Jepson, and Rich Rosen. Mac OS X for Unix geeks. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2008. ISBN 0-596-52062-X. [380] pp. LCCN QA76.76.O63 R68 2008.

Rodriguez:1988:DUO

[RKPP88] Robert Rodriguez, Matt Koehler, Larry Palmer, and Ricky Palmer. A dynamic UNIX operating system. In Association [Ass88d], pages 305–319.

Radhika:2020:CAF

[RKSV20] B. S. Radhika, N. V. Narendra Kumar, R. K. Shyamasundar, and Parjanya Vyas. Consistency analysis and flow secure enforcement of SELinux policies. Computers & Security, 94(??):Article 101816, July 2020. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0167404820300948.

Regnier:2008:EIH

[RLB08] Paul Regnier, George Lima, and Luciano Barreto. Evaluation of interrupt handling timeliness in real-time Linux operating systems. Operating Systems Review, 42(6):52–63, October 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Reeds:1992:IMS

[RM92] James A. Reeds and M. Douglas McIlroy. The IX multilevelsecure UNIX system. Computing Science Technical Report 163, AT&T Bell Laboratories, Murray Hill, NJ, USA, January ??, 1992. 1 pp.

Reeds:19xx:IMS

[RMxx] James A. Reeds and M. Douglas McIlroy. The IX multilevelsecure UNIX system. Computing Science Technical Report 163b, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 1 pp.

Reghenzani:2019:RTL

[RMF19] Federico Reghenzani, Giuseppe Massari, and William Fornaciari. The real-time Linux kernel: a survey on PRE-EMPT RT. ACM Computing Surveys, 52(1):18:1–18:??, February 2019. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3297714.

Russell:2001:LDR

[RMZ01] Steve Russell, Kathleen Mc-Givney, and Zeljka Zoranovic. Lotus Domino R5 Clustering. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-060836-X. xii + 255 pp. LCCN HF5548.4.L673 R87 2001. US$39.99. URL http:

//www.phptr.com/ptrbooks/
ptr_013060836X.html.

Rosenblum:1992:DIL

[RO92] Mendel Rosenblum and John K.

Ousterhout. The design and implementation of a logstructured file system. ACM Transactions on Computer Systems, 10(1):26–52, February 1992. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1992-10-1/p26-rosenblum/

.

Rodriguez:2001:EIL

[RÓ01] Ameneiros Rodríguez and Ibán Óscar. Estudio e implementación de una LAN para PYMES utilizando GNU/ LINUX como sistema operativo. (Spanish) [Study and implementation of a LAN with PYMES using GNU/ Linux as operating system]. E.U. Politécnica, Ferrol, Spain, 2001. ISBN ???? LCCN ???? Includes one CD-ROM.

Roberts:1975:UIU

[Rob75] C. S. Roberts. UNIXLIST — an IBM/370 utility program to list a UNIX file stored on a 9track magnetic tape. Technical Memorandum 1068, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 29, 1975. ?? pp.

Roberts:1979:ITN

[Rob79a] C. S. Roberts. Implementing and testing new versions of a good 48-bit pseudo-random number generator. Technical Memorandum 1453 (TM 79-1353-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 4, 1979. ?? pp.

Roberts:1979:PMR

[Rob79b] C. S. Roberts. Partial-match retrieval via the method of superimposed codes. Proceedings of the IEEE, 67(12):1624– 1642, December 1979. CO-DEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

URL https://ieeexplore.
ieee.org/abstract/document/
1455812.

Roberts:1982:ITN

[Rob82] C. S. Roberts. Implementing and testing new versions of a good, 48-bit, pseudorandom number generator. The Bell System Technical Journal, 61(8):2053–2063, October 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol61/bstj61-8-2053.
pdf;   http://www.alcatel-
lucent.com/bstj/vol61-1982/
articles/bstj61-8-2053.pdf.

Robboy:1984:UBO

[Rob84a] David Robboy. UNIX block I/O optimization on microcomputers. In USENIX Association [USE84b], pages 223–230. ISBN none. LCCN QA76.8.U65 U55 1984.

Robins:1984:AUD

[Rob84b] Charles M. Robins. Adapting UNIX for data communications. In USENIX Association [USE84b], pages 39–46. ISBN none. LCCN QA76.8.U65 U55

Roberts:1995:LLC

[Rob95] Rick Roberts. Linux — the low cost UNIX. C/C++ Users Journal, 13(1):43–??, January 1995. CODEN CCUJEX. ISSN 1075-2838.

Robbins:1999:VEP

[Rob99a] Arnold Robbins. vi Editor Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 1-56592-497-5. v + 66 pp. LCCN QA76.76.T49 R63 1999.

                                        US$6.95. URL http://www.
                                        oreilly.com/catalog/vipr/.

Robbins:1999:UND

[Rob99b] Arnold Robbins. UNIX in a Nutshell: a Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1999. ISBN 1-56592-427-4 (paperback). xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        unixnut3.

Robbins:1999:UNS

[Rob99c] Arnold Robbins. UNIX in a Nutshell: a Desktop Quick Reference for SVR4 and Solaris 7. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1999. ISBN 1-56592-427-4 (paperback). xvi + 598 pp. LCCN QA76.76.O63 R623 1999. US$24.95. URL http:/

/www.oreilly.com/catalog/
unixnut3.

Robbins:1999:VEK

[Rob99d] Arnold Robbins. Vi-Editor:

kurz and gut. O’Reillys Taschenbibliothek. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1999. ISBN 3-89721-213-7 (paperback). 62 pp. LCCN ???? DM 12.80, S 93.00.

Robbins:2000:HUC

[Rob00] Arnold Robbins. HP-UX: a companion to Unix in a Nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-760-5. ???? pp. LCCN QA76.76.O63 R565 2000.

Robbins:2001:EAP

[Rob01]  Arnold Robbins.  Effective

AWK Programming. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2001. ISBN 0-596-00070-7 (paperback), 0-596-80537-3 (e-book). xxiv + 421 pp. LCCN QA76.73.A95 R63 2001. US$39.95. URL http:/

/www.oreilly.com/catalog/
awkprog3/.

Robbins:2002:SAP

[Rob02a] Arnold Robbins. sed & awk Pocket Reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00352-

8. 52 (est.) pp. LCCN ???? US$12.95. URL http://

                                        safari.oreilly.com/0596003528;
                                         http://www.oreilly.com/
                                        catalog/sedawkrepr2.

Robbins:2002:SAK

[Rob02b] Arnold Robbins. Sed und awk. Kurz und gut. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 3-89721-246-3. 62 pp. LCCN ????

Roberson:2003:UMS

[Rob03] Jeff Roberson. ULE: A modern scheduler for FreeBSD. Report ??, The FreeBSD Project, ????, August 25, 2003. 13

                                        pp. URL http://www.usenix.
                                        org/event/bsdcon03/tech/
                                        full_papers/roberson/roberson.
                                        pdf.

Robbins:2004:LPE

[Rob04] Arnold Robbins. Linux Programming by Example: The Fundamentals. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-142964-7. xxxii + 687 pp. LCCN QA76.76.O63 R568 2004. US$39.99, CAN$57.99.

Robbins:2005:UN

[Rob05]  Arnold Robbins.  Unix in

a Nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-10029-

9. xviii + 885 pp. LCCN ???? US$34.95, CAN$48.95, UK£24.95. URL http:/

/www.oreilly.com/catalog/
unixnut4/index.html.

Robbins:2006:UN

[Rob06a]  Arnold Robbins.  Unix in

a Nutshell. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2006. ISBN 0-596-10029-9. xviii + 885 pp. LCCN QA76.76.O63 R566 2006. US$34.95, CAN$48.95, UK£24.95. URL http:/

/www.oreilly.com/catalog/
9780596100292;   http://
www.oreilly.com/catalog/
unixnut4/index.html.

Robbins:2006:UCS

[Rob06b] Steven Robbins. A UNIX concurrent I/O simulator. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 38(1):303– 307, March 2006. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Roberts:1978:PMR

[Rob8 ] C. S. Roberts. Partial-match retrieval via the method of superimposed codes. Computing Science Technical Report 64, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1978 (??). ?? pp.

Rochkind:1977:GDE

[Roc77] M. J. Rochkind. A generalized data entry system. Technical Memorandum 1156 (TM 77-9156-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1, 1977. ?? pp.

Rochkind:1978:NTP

[Roc78] M. J. Rochkind. NFTS: Tools for processing data with named fields. Technical Memorandum 1346 (TM 78-4565-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1, 1978. ?? pp.

Rochkind:1979:YYA

[Roc79] M. J. Rochkind. YADS — yet another database system. Technical Memorandum 1461 (TM 79-9323-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1, 1979. ?? pp.

Rochkind:1982:DSS

[Roc82] M. J. Rochkind. Database systems: Structure of a database file system for the UNIX operating system. The Bell System Technical Journal, 61(9): 2387–2405, November 1982. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol61/bstj61-9-2387.
                                        pdf.

Rochkind:1985:AUP

[Roc85] Marc J. Rochkind. Advanced UNIX Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-011818-4 (hardcover), 0-13-011800-1 (paperback). xv + 265 pp. LCCN QA76.76.O63 R63 1985. US$32.95 (hardcover), US$24.95 (paperback).

Rochkind:2004:AUP

[Roc04] Marc J. Rochkind. Advanced UNIX Programming. Addison-Wesley, Reading, MA, USA, second edition, 2004. ISBN 0-13-141154-3. xiii + 719 pp. LCCN QA76.76.O63 R63 2004. US$44.99. URL http:

//www.aw-bc.com/catalog/
academic/product/0,1144,
0131411543,00.html.

Rodriguez:1986:SCT

[Rod86] R. Rodriguez. A system call tracer for UNIX. In USENIX Association [USE86c], pages 72–80.

Roddy:1987:UNT

[Rod87]  Kevin P. Roddy.  UNIX

NROFF/TROFF — a User’s Guide. Holt, Rinehart, and Winston, New York, NY, USA, 1987. ISBN 0-03-000167-6. xii + 362 pp. LCCN QA76.76.T49 R634 1987.

Rodgers:1990:UDM

[Rod90a] Ulka Rodgers. UNIX Database Management Systems. Yourdon Press, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-945593-0. xiv + 338 pp. LCCN QA76.9.D3 R65 1990.

Rodgers:UDM90

[Rod90b] Ulka Rodgers. UNIX Database Management Systems. Yourdon Press, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-945593-0. xiv + 338 pp. LCCN QA76.9.D3 R65 1990.

Roelle:2002:HFS

[Roe02] Harald Roelle. A hot-failover state machine for gateway services and its application to a Linux firewall. Lecture Notes in Computer Science, 2506:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2506/25060181.htm;
http://link.springer.de/
link/service/series/0558/
papers/2506/25060181.pdf.

Rogan:1979:CFM

[Rog79] R. E. Rogan. COSNIX file management (FM) commandsoftware development and control …. Technical Memorandum 1379 (TM 79-4552-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 15, 1979. ?? pp.

Rolfe:2003:BBP

[Rol03]  Timothy Rolfe.  Bargain-

basement parallelism. Dr. Dobb’s Journal of Software Tools, 28(2):46, 48, 50, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302f/.

Ronneburg:2005:DGLb

[Ron05a]  Frank Ronneburg.  Debian

GNU/Linux Anwenderhandbuch: für Einsteiger, Umsteiger und Fortgeschritten. (German) [Debian GNU/Linux User Handbook: for Beginning, Switching, and Advanced Users]. Addison Wesley in Pearson Education Deutschland, München, Germany, 2005. ISBN 3-8273-2148-4. 744 pp. LCCN ???? EUR49.95. Includes one DVD.

Ronneburg:2005:DGLa

[Ron05b]  Frank Ronneburg.  De-

bian GNU/Linux: installation, administration, exploitation. (French) [Debian GNU/ Linux: installation, administration, use]. Campus Press, Paris, France, 2005. ISBN 2-7440-1941-0. 624 pp. LCCN ???? Includes one CD-ROM.

Roome:1979:DIC

[Roo79a] W. D. Roome. DIRECT II incore tables: User’s guide. Technical Memorandum 1412 (TM 79-9156-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 7, 1979. ?? pp.

Roome:1979:PWN

[Roo79b] W. D. Roome. Programmer’s Workbench: new tools for software development. Bell Laboratories Record, 57(1):19–25, January 1979. CODEN BLRCAB. ISSN 0005-8564.

Rosin:1977:ISV

[Ros77a] R. F. Rosin. “Introduction to software” — a videotape course. Technical Memorandum 1199 (TM 77-3233-9), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 3, 1977. ?? pp.

Rosin:1977:RSP

[Ros77b] R. F. Rosin. Reference sheet for PWB/MM macros. Technical Memorandum 1174, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 1977. ?? pp.

Rose:1978:PEU

[Ros78] Greg Rose. Performance evaluation under Unix and a study of PDP-11 instruction usage. Operating Systems Review, 12(3): 38–45, July 1978. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Rosenthal:1979:UUS

[Ros79] V. H. Rosenthal. The USG UNIX stockroom. Technical Memorandum 1381, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 23, 1979. ?? pp.

Rosenthal:1984:UGT

[Ros84] Steve Rosenthal. The UNIX glossary: Text processing utilities. UNIX Review, 2(9): 64, 66, 68, 70, 72, December 1984. CODEN UNRED5. ISSN 0742-3136.

                                        URL http://www.tuhs.org/
                                        Archive/Documentation/Unix_
                                        Review/unixreview_1984dec.
                                        pdf.

Rost:1988:PIO

[Ros88] Randi J. Rost. PEX introduction and overview. Technical Report Version 3.20, Digital Equipment Corporation, Workstation Systems Engineering, April 1988. This document is present in the X Window System Version 11 Release 3 in the

file X11/X11/doc/extensions/
pex/doc/intro/doc.ms.

Rost:1990:XMQ

[Ros90] Randi J. Rost. X and Motif — Quick Reference Guide. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1990. ISBN 1-55558-052-1. 369 pp. LCCN QA76.76.W56 R67 1990.

Rosenberg:1991:KSP

[Ros91] Barry Rosenberg. Korn Shell Programming Tutorial. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-56324-X. xxi + 324 pp. LCCN QA76.73.K67 R67 1991. A good tutorial on creating Korn shell scripts.

Rosenberry:1992:UD

[Ros92] Ward Rosenberry. Understanding DCE. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, October 1992. ISBN 1-56592-005-8. xxv + 233 pp. LCCN QA76.9.D5 R67 1992. US$24.95.

Rosenblatt:1993:LKS

[Ros93a] Bill Rosenblatt. Learning the Korn Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1993. ISBN 1-56592-054-6. xxii + 338 pp.

LCCN QA76.73.K67 R68 1993. US$29.95. URL http://www.

                                        oreilly.com/catalog/korn.

Rost:1993:XMQ

[Ros93b] Randi J. Rost. X and Motif — Quick Reference Guide. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1993. ISBN 1-55558-116-1, 1-55558-118-8. vii + 398 pp. LCCN QA76.76.W56 R67 1993. US$24.95.

Rosenblatt:1994:LKS

[Ros94a] Bill Rosenblatt. Learning the Korn Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1994. ISBN 1-56592-054-6. xxii + 336 pp. LCCN QA76.73.K67 R68 1994. US$29.95. URL http://www.

                                        oreilly.com/catalog/korn.

Rosenblatt:1994:URN

[Ros94b] Bill Rosenblatt. Unix RDBMS: the next generation: what are the Unix relational-database vendors doing to survive in the next generation of client/ server environments. SIG-MOD Record (ACM Special Interest Group on Management of Data), 23(4):91–103, December 1994. CODEN SRECD8. ISSN 0163-5808 (print), 1943-5835 (electronic).

Rose:1995:UTC

[Ros95] Greg Rose. UNIX tip: I can never remember those pesky options. ;login: the USENIX Association newsletter, 20(4): 36–38, August 1995. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Rosen:1996:USV

[Ros96] Kenneth H. Rosen. UNIX System V, release 4: an introduction. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 1996. ISBN 0-07-882130-4. xliii + 1175 pp. LCCN QA76.76.O63 R67 1996.

Rossmeyer:1998:BIL

[Ros98]  Brian Rossmeyer.  Book-

shelf: Intermediate level Linux guide. IEEE Software, 15(4): 90, July/August 1998. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/
s4087.pdf.

Rosenberg:1999:HKP

[Ros99]  Barry Rosenberg.  Hands-

On KornShell93 Programming. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 1999. ISBN 0-201-31018-X. xxvi + 444 pp. LCCN QA76.73.K67 R67 1999.

Ross:2000:USS

[Ros00] Seth Ross. UNIX system security tools. McGraw-Hill Unix series. McGraw-Hill, New York, NY, USA, 2000. ISBN 0-07-913788-1. xviii + 444 pp. LCCN QA76.9.A25 R665 2000. CD-ROM contains a wide selection of UNIX tools.

Rosenthal:2006:ACG

[Ros06] Hanaan Rosenthal. Apple-Script: the comprehensive guide to scripting and automation on Mac OS X. Expert’s voice in Mac development. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 1-59059-653-6 (paperback). xxxiv + 772 pp. LCCN QA76.73.A66 R67 2006.

Rovegno:1977:MCB

[Rov77] H. D. Rovegno. MAC-8 C for beginners. Technical Memorandum 1166, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1, 1977. ?? pp.

Rovegno:1978:UTS

[Rov78]  H. D. Rovegno.  UNIX

time-sharing system: a support environment for MAC-8 systems. The Bell System Technical Journal, 57(6): 2251–2263, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-2251.
                                        pdf.

Rowland:1979:CCP

[Row79a] B. R. Rowland. Changes to C and the Portable C Compiler — issue 1. Technical Memorandum 1467, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 2, 1979. ?? pp.

Rowland:1979:RCC

[Row79b] B. R. Rowland. Recent changes to C. Technical Memorandum 1498 (TM 79-3621-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 2, 1979. ?? pp.

Rowe:1995:EPD

[Row95] K. Rowe. Embedded POSIX for DSP. In Anonymous [Ano95h], pages 1719–1723.

Roy:1993:UFA

[Roy93] Paul J. Roy. UNIX file access and caching in a multicomputer environment. In USENIX [USE93d], pages 21– 37. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/mach3/roy.html.

Redman:1984:BEB

[RP84] Brian E. Redman and Pat E.

Parseghian. Behind every binary license ins the UNIX heritage. In USENIX Association [USE84b], pages 75–82. ISBN none. LCCN QA76.8.U65 U55 1984.

Rehman:2002:LDP

[RP02] Rafeeq Rehman and Christopher Paul. The Linux Development Platform. Bruce Perens’ Open source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-009115-4. xxii + 294 pp. LCCN QA76.76.O63 R435 2003. US$49.99.

Reiss:1992:XWI

[RR92] Levi Reiss and Joseph Radin. X Window Inside & Out. Mc-Graw-Hill, New York, NY, USA, 1992. ISBN 0-07-881796-X. xxv + 698 pp. LCCN QA76.76.W56 R45 1992. A recent text on using and programming X Windows.

Rogers:1993:PSF

[RR93] D. Rogers and J. Ross. POSIX security framework. In Anonymous [Ano93g], pages 432–441. ISBN 1-85617-211-2. LCCN QA76.9.A25 W68 1993.

Ram:1995:IKS

[RR95a] P. Ram and D. K. Rand. Internet kiosk: Satan: double-edged sword. Computer, 28(6):82–83, June 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reiss:1995:OCG

[RR95b] Levi Reiss and Joseph Radin. Open Computing’s Guide to Mosaic. Osborne/McGraw-Hill, Berkeley, CA, USA, 1995. ISBN 0-07-882088-X. xvii + 262 pp. LCCN TK5105.875.I57 R44 1995. US$19.95.

Robbins:1996:PUP

[RR96] Kay A. Robbins and Steven Robbins. Practical UNIX Programming: a Guide to Concurrency, Communication, and Multithreading. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-443706-3. xiv + 658 pp. LCCN QA76.76.O63 R615 1996.

Ray:1999:UVQ

[RR99] Deborah S. Ray and Eric J.

Ray. UNIX: Visual QuickStart Guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1999. ISBN 0-201-35395-4. xi + 354 pp. LCCN QA76.76.O63 R37 1998. US$17.99. URL http:

//www.peachpit.com/books/
catalog/K5852.html.

Ray:2002:MXU

[RR02a] John Ray and William Ray.

Mac OS X unleashed. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32229-3. xxi + 1464 pp. LCCN QA76.76.O63 R391 2002.

Rosenblatt:2002:LKS

[RR02b] Bill Rosenblatt and Arnold Robbins. Learning the Korn Shell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00195-9 (paperback), 1-4493-7127-2 (e-book), 1-4493-7128-0 (e-book). xviii + 412 pp. LCCN QA76.73.K67 R68 2002. US$34.95.

Ray:2003:MXM

[RR03a] John Ray and William C. Ray.

Mac OS X Maximum Security: a Hacker’s Guide to Protecting Your Mac OS X Workstation and Server. SAMS Publishing, Indianapolis, IN, USA, 2003. ISBN 0-672-32381-8. xiii + 747 pp. LCCN QA76.76.O63 R39175 2003. US$44.95.

Robbins:2003:USP

[RR03b] Kay A. Robbins and Steven Robbins. UNIX Systems Programming: Communication, Concurrency, and Threads. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2003. ISBN 0-13-042411-0 (hardcover). xvii + 893 pp. LCCN QA76.76.O63 R6215 2003.

Ray:2015:UL

[RR15] Deborah S. Ray and Eric J. Ray. Unix and Linux. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fifth edition, 2015. ISBN 0-321-99754-9 (paperback). xii + 396 pp. LCCN QA76.76.O63 R3915

                                        2015. URL http://proquest.
                                        safaribooksonline.com/9780133793871.

Robbins:2003:UCB

[RRD^{+}03] Arnold Robbins, Bill Rosenblatt, Dale Dougherty, Jerry Peek, Linda Lamb, and Brian Jepson, editors. The UNIX CD bookshelf: 7 bestselling books on CD-ROM: Unix power tools. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00392-7. ???? pp. LCCN ????

Rosen:1990:USV

[RRF90] Kenneth Rosen, Richard Rosinski, and James Farber. UNIX System V Release 4: An Introduction for New and Experienced Users. McGraw-Hill, New York, NY, USA, 1990. ISBN 0-07-881552-5. xlix + 1211 pp. LCCN QA76.76.O63 R68 1990. A very comprehensive text targeted to novice users.

Rosen:1993:UT

[RRF93] Kenneth Rosen, Richard Rosinski, and James Farber. 1001 UNIX Tips. Osborne/McGraw-Hill, Berkeley, CA, USA, 1993. ISBN 0-07-881924-5.

Rosen:1994:OCB

[RRH94] Kenneth Rosen, Richard Rosinski, and Douglas A. Host. Open Computing’s Best UNIX Tips Ever. Osborne/McGraw-Hill, Berkeley, CA, USA, 1994. ISBN 0-07-881924-5. various pp. LCCN QA76.76.O63 R678 1994. US$29.95.

Rossbach:2008:TMT

[RRH^{+}08] Christopher J. Rossbach, Hany E.

Ramadan, Owen S. Hofmann, Donald E. Porter, Aditya Bhandari, and Emmett Witchel. TxLinux and MetaTM: transactional memory and the operating system. Communications of the Association for Computing Machinery, 51(9):83–91, September 2008. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Radin:1995:OCG

[RRN95] Joseph Radin, Levi Reiss, and Steven Nameroff. Open Computing Guide to UnixWare. Osborne/McGraw-Hill, Berkeley, CA, USA, 1995. ISBN 0-07-882027-8. xix + 359 pp. LCCN QA76.76.O63R32 1995. US$29.95.

Ramadan:2007:MTT

[RRP^{+}07] Hany E. Ramadan, Christopher J. Rossbach, Donald E. Porter, Owen S. Hofmann, Aditya Bhandari, and Emmett Witchel. MetaTM/ TxLinux: transactional memory for an operating system. ACM SIGARCH Computer Architecture News, 35(2):92–103, May 2007. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Ramadan:2008:MTT

[RRP^{+}08] Hany E. Ramadan, Christopher J. Rossbach, Donald E. Porter, Owen S. Hofmann, Aditya Bhandari, and Emmett Witchel. MetaTM/TxLinux: Transactional memory for an operating system. IEEE Micro, 28(1):42–51, January/February 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ray:2009:UL

[RRR09] Deborah S. Ray, Eric J. Ray, and Deborah S. Unix Ray. Unix and Linux. Visual quickstart guide. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, fourth edition, 2009. ISBN 0-321-63678-3 (paperback). xiv + 393 pp. LCCN QA76.76.O63 R3915 2009.

Rice:1984:ASM

[RRW84] John R. Rice, Calvin Ribbens, and William A. Ward. Algorithm 622: a simple macroprocessor. ACM Transactions on Mathematical Software, 10(4): 410–416, December 1984. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Lev98].

Robbins:1978:PSC

[RS78] F. Robbins and R. Stensgaard.

Power supply controller using a MAC 8 micro processor. Technical Memorandum 1299 (TM 78-1131-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 23, 1978. ?? pp.

Rabinowitz:1990:PC

[RS90] Henry Rabinowitz and Chaim Schaap. Portable C. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-685967-4. xi + 269 pp. LCCN QA76.73 C15 R33 1990. US$29.95.

Raymond:1993:NHD

[RS93] Eric Raymond and Guy L.

Steele. The New Hacker’s Dictionary. MIT Press, Cambridge, MA, USA, second edition, 1993. ISBN 0-262-18154-1. xxi + 505 pp. LCCN PN6231.E4 H3 1993. The New Hacker’s Dictionary is a great book for learning about the various slang, jargon and customs and folklore of the ’net (as well as other lairs of the hacker.) Very silly and highly recommended. An FT-Pable version, called the Jargon File version 3.0, is available from rtfm.mit.edu, but the bound book makes great bathroom reading and contains silly cartoons and stuff.

Rubanov:2014:EPL

[RS14] Vladimir Rubanov and Denis Silakov. Ensuring portability of Linux applications through standardization and knowledge base driven analysis. Science of Computer Programming, 91 (part B)(??):234–248, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0167642314000240.

Rozman:2006:CPL

[RšT06] Igor Rozman, Marjan šterk, and Roman Trobec. Communication performance of LAM/ MPI and MPICH on a Linux cluster. Parallel Processing Letters, 16(3):323–334, September 2006. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Ritchie:1970:QTE

[RT70] D. M. Ritchie and K. L. Thompson. QED text editor. Memorandum TM-70-1373-3, TM-70-1371-2, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 22, 1970. i + 19 + 6 pp. URL

https://web.archive.org/
web/20230223030724/https:
//www.bell-labs.com/usr/
dmr/www/qedman.pdf.

Ritchie:1973:UTSa

[RT73a] Dennis M. Ritchie and Ken Thompson. The UNIX timesharing system. In ACM, editor, Fourth ACM Symposium on Operating Systems Principles, IBM Thomas J. Watson Research Center, Yorktown Heights, New York, October 15–17, 1973, page 27. ACM Press, New York, NY, USA, 1973. URL https:

//www.bell-labs.com/usr/
dmr/www/cacm.html. Abstract

only.

Ritchie:1973:UTSb

[RT73b] Dennis M. Ritchie and Ken Thompson. The UNIX timesharing system. Operating Systems Review, 7(4):27, October 1973. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic). URL http:

//citeseerx.ist.psu.edu/
viewdoc/summary?doi=10.1.
1.100.7314.

Ritchie:1974:UTS

[RT74] Dennis W. Ritchie and Ken Thompson. The UNIX timesharing system. Communications of the Association for Computing Machinery, 17(7):

365–375, July 1974. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        labouseur.com/courses/os/
                                        UNIX-Time-Sharing-System-
                                        by-Ritchie-and-Thompson.
                                        pdf;   https://www2.cs.duke.
                                        edu/courses/cps210/spring16/
                                        resources/papers/p365-ritchie.
                                        pdf.

Ritchie:1978:UTSa

[RT78] Dennis M. Ritchie and Ken Thompson. The UNIX timesharing system. The Bell System Technical Journal, 57(6): 1905–1929, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

                                        /bstj.bell-labs.com/BSTJ/
                                        images/Vol57/bstj57-6-1905.
                                        pdf;   http://www.alcatel-
                                        lucent.com/bstj/vol57-1978/
                                        articles/bstj57-6-1905.pdf.

Ritchie:1979:UTS

[RT79a] Dennis W. Ritchie and Ken Thompson. The UNIX timesharing system. In UNIX Programmer’s Manual, volume 2, pages 20–35. Holt, Rinehart, and Winston, New York, NY, USA, 1979. ISBN 0-03-061743-X. LCCN QA76.8.U65 B44 1983.

Ritchie:time-sharing79

[RT79b] Dennis W. Ritchie and Ken Thompson. The UNIX timesharing system. In UNIX Programmer’s Manual, volume 2, pages 20–35. Holt, Rinehart, and Winston, New York, NY, USA, 1979.

Ritchie:1983:UTS

[RT83a] Dennis M. Ritchie and Ken Thompson. The UNIX timesharing system (reprint). Communications of the Association for Computing Machinery, 26 (1):84–89, January 1983. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). Reprint of [RT74].

Ross:1983:USG

[RT83b] Douglas J. Ross and M. Martin Taylor. UNIX support for guaranteed real-time processing. In Software Tools Users Group [Sof83], pages 137–153. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Rosenberry:1993:DAA

[RT93] Ward Rosenberry and Jim Teague. Distributing Applications Across DCE and Windows NT. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, November 1993. ISBN 1-56592-047-3. 302 pp. LCCN QA76.9.D5 R65 1993. US$24.95.

Ritchie:2001:UTS

[RT01] Dennis M. Ritchie and Ken Thompson. The Unix timesharing system. In Classic Operating Systems: from Batch Processing to Distributed Systems [Bri01], pages 195– 220. ISBN 0-387-95113-X. LCCN QA76.76.O63 B7425

                                        2001. URL https://link.
                                        springer.com/chapter/10.
                                        1007/978-1-4757-3510-9_11.

Ritchie:2021:PPB

[RT21] Dennis Ritchie and Ken Thompson. The Unix time-sharing system (1974). In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 37, pages 357–372. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Thompson:1975:UTS

[RT5 ] Dennis M. Ritchie and Ken Thompson. The UNIX timesharing system. Technical Memorandum 1008, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1975 (??). ?? pp.

Rubin:1979:FUZ

[Rub79] S. M. Rubin. A filesystemless UNIX for the Z80 microprocessor. Technical Memorandum 1394 (TM 79-1355-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1, 1979. ?? pp.

Rubini:1998:LDD

[Rub98]  Alessandro Rubini.  Linux

Device Drivers. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1998. ISBN 1-56592-292-1. xviii + 421 pp. LCCN QA76.76.D49 R92 1998; QA87.76.O63 R82 1998. US$29.95. URL http://www.

ora.com/catalog/linuxdrive/
;   http://www.oreilly.com/
catalog/9781565922921;
http://www.oreilly.com/catalog/
linuxdrive.

Rugaber:1983:USU

[Rug83] Spencer Rugaber. A uniform and simple user interface to UNIX. In Association [Ass83], pages 113–115.

Rule:1996:GFF

[Rul96] Keith Rule. 3D Graphics File Formats: a Programmer’s Reference. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-48835-3. xii + 530 pp. LCCN T385 .R86 1996. US$39.95. Includes CD-ROM.

Russo:1991:NUG

[Rus91] Michael Russo. The New User’s Guide to the Sun Workstation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97249-8. 203 pp. LCCN QA76.8.S86 R87 1991. US$40.00.

Roskos:1981:TUS

[RW81] J. Eric Roskos and Robert I.

Winner. Toward user sharing of the microprogramming level under UNIX on the Perkin-Elmer 3220. ACM SIG-MICRO Newsletter, 12(4):67– 73, December 1981. CO-DEN SIGMDJ. ISSN 0163-5751, 1050-916X. URL https:

//dl.acm.org/doi/10.1145/
1014192.802436.

Reeds:1984:FSU

[RW84] James A. Reeds and Peter J. Weinberger. File security and the UNIX system crypt command. AT&T Bell Laboratories Technical Journal, 63(8 part 2): 1673–1683, October 1984. CO-DEN ABLJER. ISSN 0748-612X (print), 2376-7162 (electronic). Reprinted in [?, pp. 93– 103].

Rieken:1986:HUBb

[RW86a] Bill Rieken and Jim Webb. HoneyDanBer UUCP — bringing UNIX systems in the information age, part 2: Error handling, administrative aids, and user enhancements. ;login: the USENIX Association newsletter, 11(4):10–35, July/August 1986. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Rieken:1986:HUBa

[RW86b] Bill Rieken and Jim Webb. HoneyDanBer UUCP — bringing UNIX systems into the information age, part 1: Performance, security, and networking facilities. ;login: the USENIX Association newsletter, 11(3):27–36, May/June 1986. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Russell:1987:VUP

[RW87] Channing H. Russell and Pamela J. Waterman. Variations on UNIX for parallelprocessing computers. Communications of the Association for Computing Machinery, 30(12):1048–1055, December 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
33450.html.

Riekan:1992:AUN

[RW92] Bill Riekan and Lyle Weiman.

Adventures in UNIX Network Applications Programming. Wiley-Interscience, New York, NY, USA, 1992. ISBN 0-471-52858-7. xv + 448 pp. LCCN QA76.76.O63 R546 1992. A fine book on network programming.

Ruemmler:1993:UDA

[RW93] Chris Ruemmler and John Wilkes. UNIX disk access patterns. In USENIX Association [USE93f], pages 405– 420. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84

1993.  URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

Ramamurthy:1987:PCP

[RWNA87] G. Ramamurthy, Y. T. Wang, Hank Nichols, and Mike Andrews. A prototype capacity planning and configuration modeling tool for UNIX systems. In USENIX Association [USE87e], pages 103–110. LCCN QA 76.76 O63 U84 1987.

Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Ryder:1972:FVM

[Ryd72] B. G. Ryder. The FORTRAN Verifier: Motivation and implementation. Technical Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1972. ?? pp.

Ryder:1974:PV

[Ryd74]  Barbara G. Ryder.  The

PFORT verifier. Software— Practice and Experience, 4(4): 359–377, October/December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ryder:1979:CCG

[Ryd79] B. G. Ryder. Constructing the call graph of a program. IEEE Transactions on Software Engineering, SE-5(3):216– 226, May/June 1979. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

                                        URL  http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        arnumber=1702621;   https:
                                        //ieeexplore.ieee.org/abstract/
                                        document/1702621.

Ryder:1973:FVU

[Ryd3 ] B. G. Ryder. The FORTRAN Verifier: User’s guide. Computing Science Technical Report 12, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 1973 (??). ?? pp.

Ryder:1978:CCG

[Ryd8 ] B. G. Ryder. Constructing the call graph of a program. Computing Science Technical Report 73, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1978 (??). ?? pp.

Rzehak:1993:RUW

[Rze93] H. Rzehak. Real-time UNIX:

What performance can we expect? Control Engineering Practice, 1(1):65–70, February 1993. CODEN COEPEL. ISSN 0967-0661 (print), 1873-6939 (electronic).

Siever:2000:LND

[S^{+}00] Ellen Siever et al. Linux in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2000. ISBN 0-596-00025-1. xv + 797 pp. LCCN QA76.76.O63 L5459 2000.

Siever:2005:LND

[S^{+}05] Ellen Siever et al., editors.

Linux in a nutshell: a desktop quick reference. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2005. ISBN 0-596-00930-5. xiv + 925 pp. LCCN QA76.76.O63 L56 2005; QA76.76.O63 L5459 2005.

URL http://www.oreilly.
com/catalog/9780596009304.

Segal:1993:OGI

[SA93a] Mark Segal and Kurt Akeley. The OpenGL graphics interface. Technical report, Silicon Graphics Computer Systems, Mountain View, CA,USA, 1993.

Segal:1993:OGS

[SA93b] Mark Segal and Kurt Akeley. The OpenGL graphics system: a specification. Technical report, Silicon Graphics Computer Systems, Mountain View, CA,USA, 1993.

Soques:1995:UDF

[SA95] Kimbol Soques and Bruce Altman. UNIX as a DOS file and print server. Sys Admin: The Journal for UNIX Systems Administrators, 4(5):45– ??, September/October 1995. CODEN SYADE7. ISSN 1061-2688.

Spinellis:2021:EUS

[SA21] Diomidis Spinellis and Paris Avgeriou. Evolution of the Unix system architecture: An exploratory case study. IEEE Transactions on Software Engineering, 47(6):1134–1163, June 2021. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Sabsevitz:1976:GIR

[Sab76a] A. L. Sabsevitz. Guide to IBM Remote Job Entry for PWB/ UNIX users. Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1976. 7 + 2 pp.

Sabsevitz:1976:REM

[Sab76b] A. L. Sabsevitz. A regular expression matcher for the C language. Technical Memorandum 1101 (TM 76-9141-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 2, 1976. ?? pp.

Sabsevitz:1977:PUR

[Sab77]  A. L. Sabsevitz.  PWB/

UNIX RJE administrator’s guide (I.5). Technical memorandum, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1977. ?? pp.

Sabsevitz:1978:RJE

[Sab78] A. L. Sabsevitz. Remote job entry for PWB/UNIX users (B.12). Technical Memorandum 1087, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1978. ?? pp.

Saers:2003:PMF

[Sae03] Niklas Saers. A project model for the FreeBSD project. Ph.D. thesis, Universitetet i Oslo, Oslo, Norway, 2003. 232 pp.

Sakaguchi:1997:UUS

[Sak97] Takeyuki Sakaguchi. UNISEX-C: a UNIx-based symbolic EXecutor for standard C. Thesis (M.S.), University of California, Santa Barbara, Santa Barbara, CA, USA, 1997.

Singhal:1993:DOO

[SAL93] Anoop Singhal, Robert M. Arlein, and Chi-Yuan Lo. DDB: an object oriented design data manager for VLSI CAD. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):467–470, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Salus:1994:QCU

[Sal94] Peter H. Salus. A Quarter Century of UNIX. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-54777-5. xii + 256 pp. LCCN QA76.76.O63 S342 1994. URL http:

//wiki.tuhs.org/doku.php?
id=publications:quarter_
century_of_unix.

Salus:1997:TYA

[Sal97] Peter H. Salus. Twenty years ago in UNIX NEWS. ;login: the USENIX Association newsletter, 22(5):??, October 1997. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Salus:1999:YAUa

[Sal99] Peter H. Salus. 20 years ago in UNIX. ;login: the USENIX Association newsletter, 24(1):??, February 1999. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic). URL http://www.

usenix.org/publications/
login/1999-2/20yearsago.
html.

Salus:2000:YAUb

[Sal00a] Peter Salus. 20 years ago in UNIX. ;login: the USENIX Association newsletter, 25 (2):??, April 2000. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2000-4/20yearsago.
html.

Salus:2000:YAUc

[Sal00b] Peter Salus. 20 years ago in UNIX. ;login: the USENIX Association newsletter, 25 (3):??, June 2000. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2000-6/usenixnews.
html#20.

Salus:2003:BRBd

[Sal03] Peter H. Salus. Book reviews:

The bookworm; twenty-five years ago in UNIX. ;login: the USENIX Association newsletter, 28(4):??, August 2003. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/2003-08/openpdfs/salus.
pdf.

Salus:2008:DGP

[Sal08] Peter H. Salus. The daemon, the gnu, and the penguin: how free and open software is changing the world. Reed Media Services, Keller, TX, USA, 2008.

ISBN 0-9790342-3-X. 204 (est.) pp. LCCN ????

Samadi:1987:KBS

[Sam87] Behrokh Samadi. A knowledgebased system for performance tuning of the UNIX operating system. In USENIX Association [USE87e], pages 110– 123. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Samadi:1989:TKB

[Sam89]  B. Samadi.  TUNEX: a

knowledge-based system for performance tuning of the UNIX operating system. IEEE Transactions on Software Engineering, 15(7):861–874, July 1989. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/
                                        stamp.jsp?arnumber=29486.

Sanger:1989:UC

[San89] Colston Sanger. UNIX clinic. EUUG Newsletter, 9(1):63– 68, Spring 1989. CODEN EONLE8. ISSN 1011-4211.

Santifaller:1991:TIN

[San91] Michael Santifaller. TCP/IP and NFS: internetworking in a UNIX environment. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-54432-6. x + 235 pp. LCCN TK5105.7 .S27 1991. Translated by Stephen S. Wilson from the German edition, “TCP/IP und NFS in Theorie und Praxis”.

Sand:1993:FLU

[San93] Paul A. Sand. A form letter utility for UNIX. Sys Admin: The Journal for UNIX Systems Administrators, 2(3): 49–??, May/June 1993. CO-DEN SYADE7. ISSN 1061-2688.

Santifaller:1994:TIO

[San94] Michael Santifaller. TCP/IP and ONC/NFS: internetworking in a UNIX environment. Addison-Wesley, Reading, MA, USA, second edition, 1994. ISBN 0-201-42275-1. xii + 288 pp. LCCN TK5105.7 .S2713 1994. US$39.76. Translated by Stephen S. Wilson from the German edition, “TCP/IP und NFS in Theorie und Praxis”.

Sanders:1998:ELO

[San98] James Sanders. Edgewatch:

Linux, open source, and software’s future. IEEE Software, 15(5):88–91, September/October 1998. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://dlib.computer.
org/so/books/so1998/pdf/
s5088.pdf.

Sanghera:2007:SEQ

[San07] Paul Sanghera. SCSA Exam Quicklet: Sun Certified System Administrator for Solaris 10 Practice Exams. Infonential, Inc., ????, 2007. ISBN 0-9791797-0-X. 274 (est.) pp. LCCN ????

Shingledecker:2008:ODS

[SAN08] Robert Shingledecker, John Andrews, and Chris Negus. The official Damn Small Linux book: the tiny adaptable Linux that runs on anything. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-233869-6 (paperback). xx + 426 pp. LCCN QA76.76.O63 S555554 2007. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip0718/2007020589.html.

SAS:1988:STRe

[SAS88a] SAS Institute. SAS technical report: P-177 system administrator’s guide to the SAS system, release 6.03, under UNIX. Technical report, SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-097-5. 96 pp. US$6.00.

SAS:1988:USS

[SAS88b] SAS Institute. Using the SAS System: release 6.03, under UNIX operating systems and derivatives. SAS technical report; P-176. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1988. ISBN 1-55544-098-3. viii + 91 pp. LCCN QA276.4 .U75 1988. Includes index.

SAS:1990:SCU

[SAS90] SAS Institute. SAS companion for the UNIX environment and derivatives: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1990. ISBN 1-55544-420-2. xxii + 273 pp. LCCN QA276.4 .S222 1990. Includes index.

SAS:1991:GSNa

[SAS91a] SAS Institute. A Guide to the SAS Notes, Sample Library, and online customer support facility for the UNIX environment and derivatives. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-434-2. iv + 34 pp. LCCN QA276.4 .G85 1991.

SAS:1991:SAG

[SAS91b] SAS Institute. System administrator’s guide to the SAS System, release 6.07, for the UNIX environment and derivatives. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1991. ISBN 1-55544-436-9. x + 93 pp. LCCN QA276.4 .S97 1991.

Sasaki:1991:NEP

[Sas91c] Shinsuke Sasaki. Nihongo eidaburyukei puroguramingu tekunikku. (Japanese) []. Mainichi Komyunikeshonzu, Tokyo, Japan, 1991. ISBN 4-89563-165-6. 255 pp. LCCN ????

SAS:1993:ICS

[SAS93a] SAS Institute. International character support for UNIX environments: release 6.09. SAS technical report; P-257. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-569-1. x + 22 pp. LCCN ????

SAS:1993:SCUb

[SAS93b] SAS Institute. SAS companion for the UNIX environments: language: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-565-9. xviii + 256 pp. LCCN ????

SAS:1993:SCUc

[SAS93c] SAS Institute. SAS companion for UNIX environments: user interfaces: version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1993. ISBN 1-55544-560-8. xiv + 156 pp. LCCN ????

SAS:1994:SSTc

[SAS94a] SAS Institute. SAS/SHARE technical report for UNIX environments: release 6.09. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. ISBN 1-55544-609-4. 46 pp. LCCN ????

SAS:1994:STF

[SAS94b] SAS Institute. SAS/TUTOR: fundamentals of the SAS system: course guide: release 6.08: OS/2, Windows; release 6.09: OpenVMS for AXP, Windows NT, UNIX Environments; release 6.10: Microsoft Windows Environments, OS/2, UNIX environments. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1994. xvi + 131 + 1 pp.

SAS:1995:SSTc

[SAS95] SAS Institute. SAS/SHARE technical report for UNIX environments: release 6.11. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, 1995. ISBN 1-55544-261-7. x + 73 pp. LCCN ????

Satyanarayanan:1989:ISL

[Sat89]  M. Satyanarayanan.  Inte-

grating security in a large distributed system. ACM Transactions on Computer Systems, 7(3):247–280, August 1989. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1989-7-3/p247-satyanarayanan/

.

Satdeva:1990:PSI

[Sat90a] Bjorn Satdeva. Perspectives and solutions for increasing security in UNIX system administration. In USENIX Association [USE90e], pages 131–136. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Satyanarayanan:1990:SSH

[Sat90b] Mahadev Satyanarayanan. Scalable, secure, and highly available distributed file access. Computer, 23(5):9–18, 20–21, May 1990. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Satdeva:1993:SVC

[Sat93] Bjorn Satdeva. SAGE views: Communication: An important aspect of UNIX system administration. ;login: the USENIX Association newsletter, 18(1): 12–15, January/February 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Saunders:1988:AGB

[Sau88]  D. Saunders.  An adden-

dum to G. S. Blair, J. A. Mariani and W. D. Shepherd: ‘Practical Extension to UNIX for Interprocess Communication’. Software—Practice and Experience, 18(12):1191– 1192, December 1988. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Sauer:2017:LSV

[Sau17]  Charles H. Sauer.  Let’s

start at the very beginning … 801, ROMP, RT/PC, AIX versions. Web site., March

                                        2017. URL https://notes.
                                        technologists.com/notes/
                                        2017/03/08/lets-start-at-
                                        the-very-beginning-801-romp-
                                        rtpc-aix-versions/.

Savio:2011:OCL

[Sav11] Hubert Savio. Oracle Certified Linux Expert exam cram: OCE exam: 1Z0-046: managing Oracle on Linux Certified Expert, volume 38 of Oracle in-focus series. Rampant TechPress, Kittrell, NC, USA, 2011. ISBN 0-9844282-1-6. ???? pp. LCCN ????

Saxer:1985:IUB

[Sax85a] John Saxer. Interpreting UNIX benchmarks. In USENIX Association [USE85c], pages 78–89.

Saxon:1985:UGA

[Sax85b] Michael S. Saxon. Using gsck — a guide to the UNIX file system check program. ;login: the USENIX Association newsletter, 10(3):13–26, August 1985. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Saxon:1985:UGG

[Sax85c] Michael S. Saxon. Using gsck — a guide to the UNIX file system check program. ;login: the USENIX Association newsletter, 10(3):13–26, August 1985. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Schaefer:1996:CUF

[SB96] Ed Schaefer and Fred Brunet.

Converting UNIX filenames to MS-DOS. Sys Admin: The Journal for UNIX Systems Administrators, 5(2):81– ??, February 1996. CODEN SYADE7. ISSN 1061-2688.

Sedgewick:1997:FAS

[SB97] R. Sedgewick and J. Bentley.

Fast algorithms for sorting and searching strings. In ACM [ACM97], pages 360–369. CO-DEN PAAAF2. ISBN 0-89871-390-0. LCCN ???? URL http:

//www.cs.princeton.edu/~
rs/talks/strings.ps. This

is the fourth of four key papers behind the bzip2 compression tools. The others are [HL90, BW94c, Whe97].

Sieh:2002:UVS

[SB02] Volkmar Sieh and Kerstin Buchacker. UMLinux — a versatile SWIFI tool. Lecture Notes in Computer Science, 2485:159–??, 2002. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.de/link/service/
series/0558/bibs/2485/24850159.
htm;   http://link.springer.
de/link/service/series/0558/
papers/2485/24850159.pdf.

So:2008:UHS

[SB08] Hayden Kwok-Hay So and Robert Brodersen. A unified hardware/software runtime environment for FPGA-based reconfigurable computers using BORPH. ACM Transactions on Embedded Computing Systems, 7(2):14:1–14:??, February 2008. CODEN ???? ISSN 1539-9087 (print), 1558-3465 (electronic).

Schoonover:1992:GEU

[SBA92] Michael A. Schoonover, John S. Bowie, and William R. Arnold. GNU Emacs: UNIX Text Editing and Programming. Hewlett-Packard Press series. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56345-2. xxvii + 609 pp. LCCN QA76.76.T49S36.

Seltzer:1993:ILS

[SBMS93] Margo Seltzer, Keith Bostic, Marshall Kirk McKusick, and Carl Staelin. An implementation of a log-structured file system for UNIX. In USENIX Association [USE93f], pages 307–326. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/

seltzer.pdf. Best presentation winner.

Samson:2001:STY

[SBR01] Judith Samson, Jason Byars, and Dallas Releford. Sams teach yourself Red Hat Linux in 24 hours. Sams teach yourself in 24 hours series. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31845-8 (paperback). xv + 429 pp. LCCN QA76.76.O63 S354 2001.

Schuyler:1978:SAP

[SC78] S. T. Schuyler and M. M.

Chodrow. A software architecture for providing a graphics system development workbench. Technical Memorandum 1342, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1, 1978. ?? pp.

Schryer:1975:UGD

[Sch75] Norman L. Schryer. A user’s guide to DODES, a double precision ordinary differential equation solver. Computing Science Technical Report 33, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 1975. ii +

                                        18 pp. URL https://www.
                                        legacy.com/funeral-homes/
                                        obituaries/name/norman-schryer-
                                        obituary?pid=183271335&v=
                                        batesville.

Schryer:1976:NST

[Sch76a] Norman L. Schryer. Numerical solution of time-varying partial differential equations in one space variable. Computing Science Technical Report 53, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 19, 1976. ii + 44 pp.

Schryer:1976:TGM

[Sch76b] Norman L. Schryer. A tutorial on Galerkin’s method, using on B-splines, for solving differential equations. Computing Science Technical Report 52, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 17, 1976. ii + 36 pp.

Schuyler:1977:TTU

[Sch77] S. T. Schuyler. Timing tests of the UNIGRAPHIX graphics system running on a DEC PDP-11/70 under the UNIX operating system. Technical Memorandum 1233 (TM 77-8254-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 14, 1977. ?? pp.

Schan:1979:MPD

[Sch79a] E. P. Schan. Multi-processor DMERT. Technical Memorandum 1472, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 8, 1979. ?? pp.

Schryer:1979:UEPa

[Sch79b] N. L. Schryer. UNIX^{T M} as an environment for producing numerical software. ACM SIGNUM Newsletter, 14(1):49– 52, 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Schryer:1979:UEPb

[Sch79c] N. L. Schryer. UNIX^{T M} as an environment for producing numerical software. In ACM [ACM79], pages 49–52.

Scheiderman:1980:UB

[Sch80] C. L. Scheiderman. UNIX bibliography. Technical Memorandum MH X6058, AT&T Bell Laboratories, Murray Hill, NJ 07974, USA, March 1980. 59

pp. URL https://minnie.
tuhs.org/wktcloud/index.
php/s/JH3bGTrStKsA6bY.

Schryer:1981:TCF

[Sch81] Norman L. Schryer. A test of a computer’s floating-point arithmetic unit. Computing Science Technical Report 89, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 4, 1981. ii + 64 pp.

Schreiner:1985:UEa

[Sch85a] Axel Schreiner. UNIX exchange. Dr. Dobb’s Journal of Software Tools, 10(6):??, June 1985. CODEN DDJOEB. ISSN 1044-789X.

Schreiner:1985:UEb

[Sch85b] Axel Schreiner. UNIX exchange. Dr. Dobb’s Journal of Software Tools, 10(9): ??, September 1985. CODEN DDJOEB. ISSN 1044-789X.

Schaffler:1988:CPU

[Sch88]  G. Schaffler.  Connecting

PEACE to UNIX. Parallel Computing, 7(3):335–339, September 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Scheifler:1989:XPR

[Sch89] Robert W. Scheifler. X Protocol Reference Manual, volume 0. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1989. ISBN 0-937175-40-4. xv + 398 pp. US$30.00.

Schreiner:1990:UCC

[Sch90] Axel T. Schreiner. Using C with curses, lex, and yacc: building a window shell for UNIX System V. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-932864-5. ix + 257 pp. LCCN QA76.76.O63 S39 1990. US$44.95. Translation of original German edition: C-Praxis mit curses, lex und yacc.

Schildt:1991:CCRa

[Sch91a] Herbert Schildt. C: The Complete Reference. McGraw-Hill, New York, NY, USA, 1991. ISBN 0-07-881538-X. xv + 823 pp. LCCN QA76.73.C15 S34 1990. Another classic on the C language.

Schildt:1991:CCRb

[Sch91b] Herbert Schildt. C++: The Complete Reference. McGraw-Hill, New York, NY, USA, 1991. ISBN 0-07-881654-8. xxi + 594 pp. LCCN QA76.73.C15 S36 1991. Another good text on C++.

Schmitt:1991:RUT

[Sch91c] David A. Schmitt. Reviving the UNIX sbrk function. C Users Journal, 9(2):97–??, February 1991. ISSN 0898-9788.

Schaefer:1993:UUP

[Sch93a] Ed Schaefer. Using the UNIX pipe in C. Sys Admin: The Journal for UNIX Systems Administrators, 2(4):39– ??, July/August 1993. CODEN SYADE7. ISSN 1061-2688.

Schmidt:1993:CWU

[Sch93b] D. C. Schmidt. A C++ wrapper for UNIX I/O multiplexing: the object-oriented design and implementation of the Reactor. C++ Report, 5(7):32– 43, September 1993. CODEN CRPTE7. ISSN 1040-6042.

Schmidt:1993:EOS

[Sch93c] D. C. Schmidt. Encapsulating operating system IPCs. an object-oriented interface for event-driven UNIX I/O multiplexing. C++ Report, 5(2):43– 50, February 1993. CODEN CRPTE7. ISSN 1040-6042.

Schwartz:1993:LP

[Sch93d] Randal L. Schwartz. Learning Perl. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-042-2. xxv + 246 pp. LCCN QA76.73.P224 S39 1993. US$24.95.

Schaefer:1994:CIS

[Sch94a] Marc Schaefer. Courte information sur le système UNIX et les réseaux. ALPHANET NF, Battieux 6c 2013 Colombier, CH, 1994. ISBN none. 60 pp. 10 SFr. Beginner introduction to UNIX and introduction to networking tools (uucp, mail, news, www, ftp, etc.).

Schimmel:1994:USM

[Sch94b] Curt Schimmel. UNIX Systems for Modern Architectures: Symmetric Multiprocessing and Caching for Kernel Programmers. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63338-8. xxiv + 396 pp. LCCN QA76.76.063S3756 1994. US$47.50.

Scheer:1995:IUS

[Sch95a] Randall J. Scheer. Internationalizing UNIX software projects.

AT&T Technical Journal, 74 (3):85–94, 1995. CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Schimmel:1995:RUS

[Sch95b] Curt Schimmel. Review: UNIX systems for modern architectures, symmetric multiprocessing, and caching for kernel programmers. ;login: the USENIX Association newsletter, 20(2): 54–55, April 1995. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Schulman:1995:UDY

[Sch95c] Rand R. Schulman. The UNIX desktop you’ve been waiting for: CDE: a common desktop environment for UNIX delivers interoperability, platform independence, and a unified environment. The X Journal, 4(3): 38–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

Schaefer:1997:RSC

[Sch97] Ed Schaefer. Returning a single character in a UNIX shell script. Sys Admin: The Journal for UNIX Systems Administrators, 6(4):53–??, April 1997. CODEN SYADE7. ISSN 1061-2688.

Schwartz:1998:MML

[Sch98]  Alan Schwartz.  Manag-

ing Mailing Lists: Majordomo, LISTSERV, Listproc, and SmartList. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1998. ISBN 1-56592-259-X. xiv + 282 pp. LCCN ZA4480 .S39 1998. US$29.95. URL http:/

/www.oreilly.com/catalog/
mailing;   http://www.oreilly.
com/catalog/mailing/noframes.
html.

Schwartz:1999:IU

[Sch99] David I. Schwartz. Introduction to UNIX. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-095135-

8. 150 pp. LCCN xvi + 133. US$20.00.

Schildt:2000:CCR

[Sch00] Herbert Schildt. C, the complete reference. Osborne/ McGraw-Hill, Berkeley, CA, USA, fourth edition, 2000. ISBN 0-07-212124-6. xxvi + 805 pp. LCCN QA76.73.C15 S353 2000. URL http:/

/www.loc.gov/catdir/bios/
mh042/00269273.html;   http:
//www.loc.gov/catdir/description/
mh021/00269273.html;   http:
//www.loc.gov/catdir/toc/
mh021/00269273.html.

Schweitzer:2004:BRL

[Sch04] Martin Schweitzer. Book review: Linux in a Nutshell, 4th ed., by Ellen Siever, Stephen Figgins, and Aaron Weber, O’Reilly Books, 2003, $39.95, ISBN: 0-596-00482-6. ACM Queue: Tomorrow’s Computing Today, 1(10):88, February 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Schroder:2005:LC

[Sch05] Carla Schroder. Linux cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00640-3 (paperback). xxiii + 553 pp. LCCN QA76.76.O63; QA76.76.O63 S377 2004eb; QA76.76.O63 S377 2004. URL http:/

/www.oreilly.com/catalog/
9780596006402.

Schroder:2007:LNC

[Sch07] Carla Schroder. Linux networking cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2007. ISBN 0-596-10248-8 (paperback). xxiii + 612 pp. LCCN TK5105.5 .S384 2007.

URL http://www.oreilly.
com/catalog/9780596102487.

Schroder:2008:LNC

[Sch08] Carla Schroder. Linux networking cookbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2008. ISBN 0-596-10248-8. 612 pp. LCCN ????

Schmidt:2012:TDR

[Sch12] Eric Schmidt. Tribute to Dennis Ritchie. Alcatel-Lucent 40-minute video., September 11, 2012. URL https:/

/www.youtube.com/watch?v=

3-gTXWhx1y8. From an event entitled, “The Lasting Legacy of Dennis Ritchie: The Impact of Software on Society.”.

Shelly:1996:IIC

[SCJ96] Gary B. Shelly, Thomas J. Cashman, and Kurt A. Jordan. The Internet: introductory concepts and techniques (UNIX). Shelly Cashman series Shelly, Gary B. Shelly Cashman series. boyd & fraser, One Corporate Place, Suite 205, Danvers, MA, USA 01923, USA, 1996. ISBN 0-7895-0003-5. viii + 112 pp. LCCN TK5105.875.I57S52 1996.

Sum:1986:AOS

[SCK86] R. N. Sum, Jr., R. H. Campbell, and W. J. Kubitz. An approach to operating system testing. The Journal of Systems and Software, 6(3):273– 284, August 1986. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Scott:2009:PLP

[Sco09] Michael L. Scott. Programming Language Pragmatics. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, third edition, 2009. ISBN 0-12-374514-4. xxx + 910 pp. LCCN ???? Many sections of the book are relegated to the accompanying CD-ROM.

Scrocca:1976:NGS

[Scr76] Carmela Scrocca. New graphic symbols for EQN and NEQN (T.8). Technical Memorandum 1112 (TM 76-7133-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 15, 1976. 2 + 8 pp.

Scussel:1978:BP

[Scu78] C. Scussel. Basic-16 prospectus. Technical Memorandum 13 (TM 78-2522-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 6, 1978. ?? pp.

Stephens:1987:ESA

[SD87] Lindsey E. Stephens and Lawrence W. Dowdy. Experimental sensitivity analysis of performance in a UNIX system. In USENIX Association [USE87e], pages 60–72. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Sanderson:1993:S

[SD93] David Sanderson and Dale Dougherty. Smileys. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1993. ISBN 1-56592-041-4. 93 pp. US$5.95. A ‘dictionary’ of the smileys (aka: emoticons) …: -).

Spitz:1995:IAP

[SD95] A. L. Spitz and A. . Dengel, editors. International Association for Pattern Recognition Workshop on Document Analysis Systems, volume 14 of Series in machine perception and artificial intelligence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1995.

ISBN 981-02-2122-3, 981-02-2122-3. LCCN TK7882.P3I55 1994.

Smith:2005:DL

[SD05a] Roderick Smith and Jeff Duntemann. Degunking Linux. Paraglyph Press, Inc., Scottsdale, AZ, USA, 2005. ISBN 1-933097-04-3. xx + 332 pp. LCCN QA76.76.O63 S58 2005.

                                        URL http://www.oreilly.
                                        com/catalog/9781933097046.

Smith:2005:DLH

[SD05b] Roderick Smith and Jeff Duntemann. Degunking Linux. Paraglyph Press, Scottsdale, AZ, USA, 2005. ISBN ???? xx + 332 pp. LCCN QA76.76.O63 S65 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://site.ebrary.com/
                                        lib/ucsc/Doc?id=10080003.

Strobel:1996:LPW

[SE96] Stefan Strobel and Volker Elling. Linux — vom PC zur Workstation: Grundlagen, Installation und praktischer Einsatz. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-60557-6, 3-540-58097-2. xiv + 521 pp. LCCN ???? Also available in English [SE97] and Hungarian [SU96] translations.

Strobel:1997:LUW

[SE97] Stefan Strobel and Volker Elling. Linux, unleashing the workstation in your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94880-5. xv + 587 pp. LCCN QA76.5 .S78513 1997. US$30. Foreword by Jurgen Gulbins. Translated by Robert Bach and Aileen Darling from the German edition [SE96].

Seacord:2005:SCC

[Sea05] Robert C. Seacord. Secure coding in C and C++. Addison-Wesley, Reading, MA, USA, 2005. ISBN 0-321-33572-4 (paperback). xxiv + 341 pp. LCCN QA76.9.A25 S368

2005. URL http://www.cert.
org/books/secure-coding/;
 http://www.loc.gov/catdir/
toc/ecip0513/2005015012.
html.

Seacord:2009:CCS

[Sea09] Robert C. Seacord. The CERT C secure coding standard. Addison-Wesley, Reading, MA, USA, 2009. ISBN 0-321-56321-2 (paperback). xxxiii + 682 pp. LCCN QA76.73.C15 S4155 2008.

Seacord:2014:CCC

[Sea14] Robert C. Seacord. The CERT C coding standard: 98 rules for developing safe, reliable, and secure systems. The SEI series in software engineering. Addison-Wesley, Reading, MA, USA, second edition, 2014. ISBN 0-321-98404-8 (paperback). xliv + 517 pp. LCCN QA76.73.C153 S433 2014.

                                        URL http://proquestcombo.
                                        safaribooksonline.com/9780133812275.

Seacord:2020:ECI

[Sea20] Robert C. Seacord. Effective C: An Introduction to Professional C Programming. No Starch Press, San Francisco, CA, USA, 2020. ISBN 1-71850-104-8. 272 (est.) pp. LCCN

                                        ???? URL https://nostarch.
                                        com/Effective_C.

Sebern:1994:BOM

[Seb94] Mark J. Sebern. Building OSF/ Motif applications — a Practical Introduction. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-122409-

3. xvii + 717 pp. LCCN

QA76.9.U83 S43 1994.

Seebach:2008:BPS

[See08]  Peter Seebach.  Beginning

portable shell scripting: from novice to professional. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2008. ISBN 1-4302-1043-5 (paperback). xix + 352 pp. LCCN QA76.76.O63 S4668 2009.

Sefton:1993:USN

[Sef93] Laurie Sefton. UNIX security in a networked environment. Sys Admin: The Journal for UNIX Systems Administrators, 2(1): 38–??, January/February 1993. CODEN SYADE7. ISSN 1061-2688.

Seifried:2000:HHU

[Sei00] Kurt Seifried. How to hack UNIX: Part 2. Sys Admin:

The Journal for UNIX Systems Administrators, 9(12):77– 80, December 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Szydlowski:2012:CDA

[SEKV12] Martin Szydlowski, Manuel Egele, Christopher Kruegel, and Giovanni Vigna. Challenges for dynamic analysis of iOS applications. Lecture Notes in Computer Science, 7039:65–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.
com/chapter/10.1007/978-
3-642-27585-2_6/.

Sellens:1994:SLT

[Sel94]  John Sellens.  SPP —

low tech, practical, UNIX software portability. In USENIX [USE94c], pages 57– 66. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

URL  http://www.usenix.
org/publications/library/
proceedings/appdev94/sellens.
html.

Semich:1996:FPU

[Sem96] J. William Semich. FRAME-WORK — the promise of UNIX — who could possibly need 16 qbytes of memory? Qbytes — you know, “Quintillions of bytes.”. Datamation, 42(11):5– ??, ???? 1996. CODEN DTM-NAT. ISSN 0011-6963.

Sery:2000:RLN

[Ser00] Paul G. Sery. Red Hat Linux Network Toolkit with CD. I D G Books Worldwide, Indianapolis, IN, USA, second edition, 2000. ISBN 0-7645-4656-2. xxviii + 439 pp. LCCN QA76.76.O63 S4693 2000.

Sethi:1979:SCPa

[Set79a] Ravi Sethi. Semantics of the C programming language, part 0: Prelude. Technical Memorandum 1353 (TM 79-1271-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 2, 1979. ?? pp.

Sethi:1979:SCPb

[Set79b] Ravi Sethi. Semantics of the C programming language, part 1: Statements. Technical Memorandum 1376 (TM 79-1271-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 2, 1979. ?? pp.

Sethi:1979:SCPc

[Set79c] Ravi Sethi. Semantics of the C programming language, part 2: Declarations. Technical Memorandum 1427, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 2, 1979. ?? pp.

Sethi:1981:CFA

[Set81]  Ravi Sethi.  Control flow

aspects of semantics-directed compiling. Computing Science Technical Report 98, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1981. ?? pp.

Sethi:1982:CFA

[Set82]  Ravi Sethi.  Control flow

aspects of semantics-directed compiling (summary). In ACM [ACM82], pages 245–260. ISBN 0-89791-074-5. LCCN QA76.7 .A8 v.17, no.6. URL https:

//dl.acm.org/doi/10.1145/
800230.806999.

Sethi:1983:CFA

[Set83]  Ravi Sethi.  Control flow

aspects of semantics-directed compiling. ACM Transactions on Programming Languages and Systems, 5(4):554–595, October 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/69575.357227.

Schultz:2001:MME

[SEZ^{+}01] Matthew G. Schultz, Eleazar Eskin, Erez Zadok, Manasi Bhattacharyya, and Salvatore J. Stolfo. MEF, Malicious Email Filter — a UNIX mail filter that detects malicious Windows executables. In USENIX [USE01b], page ?? ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001.

URL  http://www.usenix.
org/publications/library/
proceedings/usenix01/freenix01/
schultz.html.

Schreiner:1985:ICC

[SF85] Axel T. Schreiner and H. George Friedman, Jr. Introduction to Compiler Construction Under UNIX. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1985. ISBN 0-13-474396-2. viii + 194 pp. LCCN QA76.76.C65 S37 1985. URL ftp://ftp.

                                        cs.uiuc.edu/pub/faculty/
                                        friedman/tar.

Sarna:1995:PSD

[SF95] D. Sarna and G. Febish. Paradigm shift — does Windows do UNIX? That’s what the UNIX folks want to know. Datamation, 41(21):33– ??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Shabtai:2010:SAP

[SFE10] Asaf Shabtai, Yuval Fledel, and Yuval Elovici. Securing Android-powered mobile devices using SELinux. IEEE Security & Privacy, 8(3):36–44, May/June 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Stevens:2004:UNP

[SFR04] W. Richard Stevens, Bill Fenner, and Andrew M. Rudoff. UNIX Network Programming: The Sockets Networking API, volume 1. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2004. ISBN 0-13-141155-1 (hardcover). xxiii + 991 pp. LCCN QA76.76 .O63 S75 2004. US$69.99.

Siever:2003:LN

[SFW03] Ellen Siever, Stephen Figgins, and Aaron Weber. Linux in a nutshell. In a nutshell.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2003. ISBN 0-596-00482-6. xiv + 928 pp. LCCN QA76.76.O63 S5582 2003. URL http:/

/www.oreilly.com/catalog/
9780596004828.

Scheifler:1990:XWSa

[SG90] Robert W. Scheifler and James Gettys. The X Window System. Software—Practice and Experience, 20(S2):5–34, October 1990. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Siemann:1995:WWI

[SG95] Robert Siemann and Shirley Ann Gartmann. Welcome to the world of the Internet: an Internet simulator for PC-DOS and Macintosh, UNIX (version 2.0). Technical report, ????, ????, 1995. 62 pp.

Scheifler:1997:XWSa

[SG97a] Robert W. Scheifler and James Gettys. X Window System: core and extension protocols: X version 11, releases 6 and 6.1. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1997. ISBN 1-55558-148-X. xix + 996 pp. LCCN QA76.76.W56S342 1997. Edited by Al Mento and Donna Converse.

Scheifler:1997:XWSb

[SG97b] Robert W. Scheifler and James Gettys. X Window System: core library and standards: X version 11, releases 6 and 6.1. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1997. ISBN 1-55558-154-4. xix + 798 pp. LCCN QA76.76.W56S342 1997. Edited by Al Mento and Donna Converse.

Serral:2004:LNS

[SG04] René Serral and Marisa Gil. A Linux networking study. Operating Systems Review, 38(3): 1–11, July 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sanghera:2006:SCS

[SG06] Paul Sanghera and Nalneesh Gaur. Sun Certified System Administrator for Solaris 10 Study Guide (Exams CX-310-200 and CX-310-202). Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 2006. ISBN 0-07-222959-4. 599 (est.) pp. LCCN ????

Seager:2009:CHS

[SG09] Mark Seager and Brent Gorda. The case for a hierarchical system model for Linux clusters. The International Journal of High Performance Computing Applications, 23(4):350– 354, November 2009. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (elec-

tronic). URL http://hpc.
sagepub.com/content/23/4/
350.full.pdf+html.

Scheifler:1990:XWSb

[SGFR90] Robert W. Scheifler, James Gettys, Jim Flowers, and David Rosenthal. X Window System: The Complete Reference to Xlib, X Protocol, IC-CCCM, XLFD. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1990. ISBN 1-55558-050-5. xxiv + 851 pp. LCCN QA76.76.W56 S34 1990. US$44.95.

Scheifler:1992:XWS

[SGFR92] Robert W. Scheifler, James Gettys, Jim Flowers, and David Rosenthal. X Window System: The Complete Reference to Xlib, X Protocol, ICCCCM, XLFD, X Version 11, Release

5. Digital Press and Prentice-Hall, 12 Crosby Drive, Bedford, MA 01730, USA and Upper Saddle River, NJ 07458, USA, third edition, 1992. ISBN 1-55558-088-2 (DP), 0-13-971201-1 (PH). xxviii + 1000 pp. LCCN QA76.76.W56 S34 1992. US$49.95.

Silberschatz:2005:OSC

[SGG05] Abraham Silberschatz, Peter B.

Galvin, and Greg Gagne. Operating System Concepts. Wiley, New York, NY, USA, seventh edition, 2005. ISBN 0-471-69466-5 (hardcover). xxii + 921 pp. LCCN QA76.76.O63 S5583

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0618/2004059084-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0618/2004059084-

                                        d.html;   http://www.loc.
                                        gov/catdir/toc/wiley051/
                                        2004059084.html.

Silberschatz:2009:OSC

[SGG09] Abraham Silberschatz, Peter Baer Galvin, and Greg Gagne. Operating System Concepts. Wiley, New York, NY, USA, eighth edition, 2009. ISBN 0-470-12872-0. xx + 972 pp. LCCN QA76.76.O63.

Seltzer:2000:JVS

[SGM^{+}00] M. Seltzer, G. Ganger, Marshall Kirk McKusick, K. Smith, C. Soules, and C. Stein. Journaling versus soft updates: Asynchronous meta-data protection in file systems. In ????, editor, Proceedings of the San Diego Usenix Conference, June 2000, pages 71–84. USENIX Association, Berkeley, CA, USA, 2000.

Scheifler:1988:XWS

[SGN88] Robert W. Scheifler, James Gettys, and Ron Newman. X Window System: C Library and Protocol Reference. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1988. ISBN 1-55558-012-2. xxix + 701 pp. LCCN QA76.76.W56 S34 1988.

S:2023:DMC

[SGP^{+}23] Sajin S, Shubham Sunil Garag, Anuj Phegade, Deepshikha Gusain, and Kuruvilla Varghese. Design of a multi-core compatible Linux bootable 64bit out-of-order RISC-V processor core. In IEEE, editor, 2023 36th International Conference on VLSI Design and 2023 22nd International Conference on Embedded Systems (VLSID), pages 42–47. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023.

Shalaby:2002:SSN

[SGWP02] Nadia Shalaby, Yitzchak Gottlieb, Mike Wawrzoniak, and Larry Peterson. Snow on silk: a NodeOS in the Linux kernel. Lecture Notes in Computer Science, 2546:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.de/
link/service/series/0558/
bibs/2546/25460001.htm;
http://link.springer.de/
link/service/series/0558/
papers/2546/25460001.pdf.

Shattan:1985:DUB

[SH85] Ariel Shattan and Jenny Hecker. Documenting UNIX: Beyond man pages. In USENIX Association [USE85d], pages 437–454. LCCN QA76.8.U65 U8 1985.

Siyan:1995:IFN

[SH95] Karanjit Siyan and Chris Hare.

Internet Firewalls and Network Security. New Riders Publishing, Carmel, IN, USA, 1995. ISBN 1-56205-437-6. xv + 410 pp. LCCN TK5105.875.I57 H36 1995. US$35.00, CDN$47.95, UK£32.49.

Sery:1999:RLS

[SH99a] Paul G. Sery and Eric Harper. Red Hat Linux in Small Business. I D G Books Worldwide, Indianapolis, IN, USA, 1999. ISBN 0-7645-3335-

5. xxv + 392 pp. LCCN QA76.76.O63 S469 1999. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        bios/wiley044/99038069.html;
                                         http://www.loc.gov/catdir/
                                        description/wiley038/99038069.
                                        html;   http://www.loc.gov/
                                        catdir/toc/wiley031/99038069.
                                        html.

Siever:1999:LND

[SH99b] Ellen Siever and Jessica Perry Hekman. Linux in a nutshell: a desktop quick reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 1999. ISBN 1-56592-585-8. xiv + 612 pp. LCCN QA76.76.O63 .S4 1999; QA76.76.O63 H453 1999. US$24.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        9781565925854.

Salah:2013:PIF

[SH13] K. Salah and M. Hamawi. Performance of IP-forwarding of Linux hosts with multiple network interfaces. Journal of Network and Computer Applications, 36(1):452–465, January 2013. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1084804512001063.

Shannon:1977:CAC

[Sha77] M. J. Shannon, Jr. A computer assisted communication system. (CACS). Technical Memorandum 1184 (TM 77-1352-5), AT&T Bell Laboratories, Murray Hill, NJ, USA, July 22, 1977. ?? pp.

Shanahan:1979:FRS

[Sha79a] M. Shanahan. Final report:

Secretarial word processing facility enhancement. Technical Memorandum 1386, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 30, 1979. ?? pp.

Shannon:1979:SPI

[Sha79b] M. J. Shannon. SPS protocol implemented for DZ11 under MERT Release 0.1. Technical Memorandum 1469, AT&T Bell Laboratories, Murray Hill, NJ, USA, June 1, 1979. ?? pp.

Shannon:1981:DPU

[Sha81] William Arthur Shannon. A demand paged UNIX system for the Harris/6 minicomputer. Master of Science, Department of Computer Engineering and Science, Case Western Reserve University, Cleveland, OH 44106, USA, January 7, 1981. vii + 115 + 1 pp.

URL https://www.tuhs.org/
Archive/Documentation/Theses/
ShannonWilliam_DemandPagedUNIXSystem_
1981_thesis.pdf.

Sharp:1993:DLU

[Sha93] Oliver Sharp. Dynamic linking under Berkeley UNIX. Dr. Dobb’s Journal of Software Tools, 18(5):40, 42, 44, 86, 88, May 1993. CODEN DDJOEB. ISSN 1044-789X.

Sharp:1994:OUR

[Sha94] Bill Sharp. It’s Official: UNIX Rules on the Server! Datamation, 40(22):50–??, November 15, 1994. CODEN DTM-NAT. ISSN 0011-6963.

Shah:2000:UWI

[Sha00a] Rawn Shah. UNIX and Windows 2000 Integration Toolkit: a Complete Guide for System Administrators and Developers. Wiley, New York, NY, USA, 2000. ISBN 0-471-29354-7. xxiv + 487 pp. LCCN QA76.76.O63 S524 2000. US$49.99. URL http:/

                                        /www.wiley.com/compbooks/
                                        catalog/29354-7.htm;   http:
                                        //www.wiley.com/compbooks/

shah. Includes CD-ROM.

Shah:2000:LAB

[Sha00b] Steve Shah. Linux administration: a beginner’s guide. Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212229-3. xxx + 542 pp. LCCN QA76.76.O63 S523 2000.

Shah:2001:LAB

[Sha01] Steve Shah. Linux administration: a beginner’s guide. Network professional’s library. Osborne/McGraw-Hill, Berkeley, CA, USA, second edition, 2001. ISBN 0-07-213136-5. xxvi + 843 pp. LCCN QA76.76.O63 S524 2001. Includes CD-ROM.

Simon:2023:CAH

[SHB23] Marek Simon, Ladislav Huraj, , and Nicolas Búcik. A comparative analysis of high availability for Linux container infrastructures. Future Internet, 15(8):253, July 28, 2023. CO-DEN ???? ISSN 1999-5903.

URL https://www.mdpi.com/
1999-5903/15/8/253.

Shein:1993:RUS

[She93a] Barry Shein. Review: UNIX System V network programming. ;login: the USENIX Association newsletter, 18(4):40– ??, July/August 1993. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Sheldon:1993:NNC

[She93b] Thomas Sheldon. Novell Net-Ware 4: The Complete Reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 1993. ISBN 0-07-881909-1. xxii + 920 pp. LCCN QA76.76.O63 S5537 1993. Probably the best book on the latest version of Novell NetWare (which provides better support for global networking) ….

Sherlock:1993:UDM

[She93c]  Margie Sherlock.  Using

DECwindows Motif for Open-VMS. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1993. ISBN 1-55558-114-5. xiii + 350 pp. LCCN QA76.76.W56 S497 1993.

Sheets:2001:WGA

[She01]  John R. Sheets.  Writing

Gnome Applications. Addison-Wesley, Reading, MA, USA, 2001. ISBN 0-201-65791-0. xxv + 449 pp. LCCN QA76.76.D47 S4885 2000.

Sheer:2002:LRU

[She02] Paul Sheer. Linux: rute users tutorial and exposition. Prentice Hall PTR open source technology series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-033351-4. xxxi + 630 pp. LCCN QA76.76.O63 S5527 2002. URL http://

                                        vig.prenhall.com/catalog/
                                        academic/product/1,4096,
                                        0130333514,00.html. Includes

CD-ROM with a complete HTML version of the book.

Sherer:2003:RTS

[She03] Matt Sherer. Real-time signal analysis and Real-Time Linux: Part 1. Dr. Dobb’s Journal of Software Tools, 28 (7):62–65, July 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2003/2003_07/
                                        rtlp1.txt.

Shermerhorn:2006:HOL

[She06] Lee Shermerhorn. HP/OSLO Linux scalability tracking and investigations. In Anonymous [Ano06a], page ?? ISBN ???? LCCN ???? URL http://

www.ice.gelato.org/oct06/
pres_pdf/gelato_ICE06oct_
scaltracking_shermerhorn_
hp.pdf.

Shirtz:1977:ULM

[Shi77] A. L. Shirtz. UNIX link to MBS line printer and Computer Center services. Technical Memorandum 1181 (MF 77-8234-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, August 11, 1977. ?? pp.

Shirley:1992:GWD

[Shi92] John Shirley. Guide to Writing DCE Applications. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, July 1992. ISBN 1-56592-004-X. xxv + 251 pp. LCCN QA76.9.D5 S5 1992. US$29.95.

Sherman:1980:ACG

[SHLR80] Mark Sherman, Andy Hisgen, David Alex Lamb, and Jonathan Rosenberg. An Ada code generator for VAX 11/780 with Unix. ACM SIGPLAN Notices, 15(11):91–100, November 1980. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1997:TUS

[SHN97] Garth Snyder, Trent R. Hein, and Evi Nemeth. Tools for UNIX system administrators. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1997.

ISBN 0-13-665431-2. LCCN QA76.76.O63 T45 1997. Companion CD-ROM to UNIX system administration handbook, 2nd ed..

Shotts:2012:LCL

[Sho12] William E. Shotts, Jr. The Linux command line: a complete introduction. No Starch Press, San Francisco, CA, USA, 2012. ISBN 1-59327-389-4 (paperback). xxix + 446 pp. LCCN QA76.76.O63 S5556 2012.

Smith:2007:LAD

[SHPP07] Bob Smith, John Hardin, Graham Phillips, and Bill Pierce. Linux appliance design: a hands-on guide to building Linux appliances. No Starch Press, San Francisco, CA, USA, 2007. ISBN 1-59327-140-9. xxi + 356 pp. LCCN QA76.76.O63 L545115 2007. URL http:/

                                        /www.oreilly.com/catalog/
                                        9781593271404.

Shultis:1983:FS

[Shu83] J. Shultis. A functional shell. ACM SIGPLAN Notices, 18 (6):202–211, June 1983. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shumba:2006:THL

[Shu06] Rose Shumba. Teaching handson Linux host computer security. ACM Journal on Educational Resources in Computing (JERIC), 6(3):5:1–5:??, September 2006. CODEN ???? ISSN 1531-4278.

Simon-Hurtado:2005:LUF

[SHVP05] Ma Aránzazu Simón-Hurtado and Carlos Vivaracho-Pascual. Learning UNIX in first year of computer engineering. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 37(3):392, September 2005. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Siever:2005:LN

[Sie05]  Ellen Siever.  Linux in a

nutshell. In a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fifth edition, 2005. ISBN 0-596-00930-5. xiv + 928 pp. LCCN QA76.76.O63; QA76.76.O63 H453 2005eb; QA76.76.O63 H453 2005; QA76.76.O63 L5459 2003. URL http:/

/www.oreilly.com/catalog/
9780596009304.

Silvester:1984:USG

[Sil84] Peter P. Silvester. The Unix System Guidebook: an Introductory Guide for Serious Users. Springer books on professional computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 0-387-90906-0 (New York), 3-540-90906-0 (Berlin). ISSN 0177-7718. xi + 207 pp. LCCN QA76.6 .S564 1984.

Simonson:1991:SRA

[Sim91] John Simonson. System resource accounting on UNIX. In USENIX [USE91b], pages 61– 72.

Simmons:1993:SBR

[Sim93] Steve Simmons. SAGE book review: UNIX System Performance Tuning. ;login: the USENIX Association newsletter, 18(1):17–??, January/ February 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Simpson:1995:USN

[Sim95] David Simpson. A UNIX Server Is No Mainframe. Datamation, 41(23):40–??, 1995. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:CSU

[Sim96a]  David Simpson.  COVER

STORY: UNIX. Datamation, 42(11):??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Simpson:1996:HCN

[Sim96b] David Simpson. How to cross the NT, UNIX divide — more IS shops are finding themselves with a mix of UNIX and NT apps. here are some integration and migration options you should consider. Datamation, 42(10):68–??, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Simpson:1996:UNT

[Sim96c] David Simpson. UNIX? NT? Time to decide! For enterpriselevel servers, solving the UNIX vs NT conundrum may be one of your biggest decisions over the next three to. Datamation, 42(2):24–??, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Simpson:1997:DUC

[Sim97] David Simpson. Do UNIX clusters pass muster? UNIX clusters are three steps ahead of NT clusters, but a few features are still missing. Datamation, 43 (2):104–??, ???? 1997. CODEN DTMNAT. ISSN 0011-6963.

Singer:1997:TAD

[Sin97]  Daniel E. Singer.  Tool-

Man’s approach to documenting UNIX directories. ;login: the USENIX Association newsletter, 22(5):45–48, June 1997. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Singh:2007:MXI

[Sin07] Amit Singh. Mac OS X internals: a systems approach. Addison-Wesley, Reading, MA, USA, 2007. ISBN 0-321-27854-2 (hardback). xxxiii + 1641 pp. LCCN QA76.76.O63 S5645 2007. URL http:

//www.loc.gov/catdir/toc/
ecip0613/2006014901.html.

Sivula:1990:HOS

[Siv90] C. Sivula. How open system suppliers differ (UNIX). Datamation, 36(10):57–58, 60, 62, May 1990. CODEN DTMNAT. ISSN 0011-6963.

Sivonen:2001:SBC

[Siv01] Timo Sivonen. Setting up BIND8 in a change-rooted environment on Solaris. ;login: the USENIX Association newsletter, 26(4):??, July 2001. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

                                        tronic). URL http://www.
                                        usenix.org/publications/
                                        login/2001-07/pdfs/sivonen.
                                        pdf.

Sivonen:2003:ICS

[Siv03] Timo Sivonen. IPv6 configuration on Solaris 9 and FreeBSD-4.x. ;login: the USENIX Association newsletter, 28(2):??, April 2003. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

                                        tronic). URL http://www.
                                        usenix.org/publications/
                                        login/2003-04/pdfs/sivonen.
                                        pdf.

Skinner:1983:UNS

[SJ83] Glenn C. Skinner and Bill Jolitz. UNIX on the National Semiconductor NS16032. In Association [Ass83], pages 291– 306.

Stevenson:1995:MUG

[SJ95a] J. Mark Stevenson and Daniel P. Julin. Mach-US: UNIX on generic OS object servers. In USENIX [USE95b], pages 119– 130. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

                                        URL  http://www.usenix.
                                        org/publications/library/

proceedings/neworl/index.
html.

Stevenson:1995:MUU

[SJ95b] J. Mark Stevenson and Daniel P.

Julin. Mach-US: UNIX on generic OS object servers. In USENIX [USE95b], pages 119– 130. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL  http://www.usenix.
org/publications/library/
proceedings/neworl/stevenson.
html.

Solter:2009:OB

[SJM09] Nicholas Solter, Jerry Jelinek, and David Miner. OpenSolaris bible. Wiley, New York, NY, USA, 2009. ISBN 0-470-38548-0 (paperback). xxxv + 971 pp. LCCN QA76.76.O63. US$26.99.

Seo:2009:CTR

[SJP^{+}09] Euiseong Seo, Jinkyu Jeong, Seonyeong Park, Jinsoo Kim, and Joonwoon Lee. Catching two rabbits: adaptive real-time support for embedded Linux. Software—Practice and Experience, 39(5):531–550, April 10, 2009. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Schweikert:1972:PMP

[SK72] D. G. Schweikert and B. W.

Kernighan. A proper model for partitioning of electrical circuits. In IEEE, editor, Proceedings of the ACM-IEEE Design Automation Workshop, pages 57–62. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1972.

Sery:2001:SLS

[SK01] Paul G. Sery and Mohammed J. Kabi. The SuSE Linux server. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2001. ISBN 0-7645-4765-8. xxviii + 609 pp. LCCN QA76.9.C55 S39 2000.

Shankar:2004:COS

[SK04] K. S. Shankar and Helmut Kurth. Certifying open source—the Linux experience. IEEE Security & Privacy, 2(6):28–33, November/ December 2004. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic). URL

http://csdl.computer.org/
dl/mags/sp/2004/06/j6028.
htm;   http://csdl.computer.
org/dl/mags/sp/2004/06/j6028.
pdf.

Salah:2010:PEC

[SK10a] K. Salah and A. Kahtani. Performance evaluation comparison of Snort NIDS under Linux and Windows Server. Journal of Network and Computer Applications, 33(1):6–15, January 2010. CODEN JN-CAF3. ISSN 1084-8045 (print), 1095-8592 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1084804509001040.

Sanders:2010:CEI

[SK10b] Jason Sanders and Edward Kandrot. CUDA by Example: an Introduction to Generalpurpose GPU Programming. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-13-138768-5. xix + 290 pp. LCCN QA76.76.A65.

Skazinski:1994:PAR

[Ska94] Joseph G. Skazinski. Porting Ada: a report from the field. Computer, 27(10):58–64, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stringfellow:2000:BRP

[SKB00] Stan Stringfellow, Miroslav Klivansky, and Michael Barto. Backup and Restore Practices for Sun Enterprise Servers. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-089401-X. xvi + 166 pp. LCCN QA76.9.D348 S77 2000. URL

books/brbp.pdf;   http://
www.sun.com/books/catalog/
stringfellow/. Part No. 806-

2894-10 May 2000.

Skjellum:1982:SFS

[Skj82] Anthony Skjellum. Shell of five subroutines to expand wildcards under UNIX. Dr. Dobb’s Journal of Software Tools, 7 (11):12–??, November 1982. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1983:CUPa

[Skj83a] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 8(10):??, October 1983. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1983:CUPb

[Skj83b] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 8(12):??, December 1983. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPa

[Skj84a] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(2):??, February 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPb

[Skj84b] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(4):??, April 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPc

[Skj84c] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(6):??, June 1984. CO-DEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPd

[Skj84d] Anthony Skjellum. C/UNIX programmer’s notebook. Dr.

Dobb’s Journal of Software Tools, 9(8):??, August 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPe

[Skj84e] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(9):??, September 1984. CODEN DDJOEB. ISSN 1044-789X.

Skjellum:1984:CUPf

[Skj84f] Anthony Skjellum. C/UNIX programmer’s notebook. Dr. Dobb’s Journal of Software Tools, 9(10):??, October 1984. CODEN DDJOEB. ISSN 1044-789X.

Sklower:1991:TBP

[Skl91] Keith Sklower. A tree-based packet routing table for Berkeley UNIX. In USENIX [USE91d], pages 93–104. LCCN QA 76.76 O63 U84 1992.

Skocovski:1998:UPP

[Sko98]  Ludek Skocovski.  UNIX,

POSIX, Plan 9. ????, Brno, Czech Republic, 1998. ISBN 80-902612-0-5. 394 pp. In Czech.

Speer:1994:IUK

[SKP94] Steven E. Speer, Rajiv Kumar, and Craig Partridge. Improving UNIX kernel and networking performance using profile based optimization. In USENIX [USE94f], pages 181–188. ISBN 1-880446-58-8. LCCN QA 76.76 O63

                                        U84 1994. URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/sf94/
                                        partridge.html.

Skrivervik:1989:DUB

[Skr89] Helge Skrivervik. Driftshåndbok for Unix BSD. Yrkesopplæring, Oslo, Norway, 1989. ISBN 82-585-0699-4. 199 (est.) pp. LCCN ????

Sarwar:2001:UTS

[SKS01] Syed Mansoor Sarwar, Robert Koretsky, and Syed Aqeel Sarwar. Unix, the textbook. Addison-Wesley Longman, Harlow, Essex CM20 2JE, England, 2001. ISBN 0-201-61260-7. various pp. LCCN QA76.76.O63 S3555 2001.

Skubiszewski:1988:SEU

[Sku88] Marcin Skubiszewski. Security of Ethernet under UNIX and Internet protocol. EUUG Newsletter, 8(1):2–10, Spring 1988. CODEN EONLE8. ISSN 1011-4211.

Lycklama:1977:SED

[SL76] E. W. Stark and Heinz Lycklama. System for entering data through computer displayed forms. Technical Memorandum 1117 (TM 76-1352-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 22, 1976. ?? pp.

Schaefer:1995:CDB

[SL95] Ed Schaefer and Charles Leonard. Converting data between UNIX and windows. Sys Admin: The Journal for UNIX Systems Administrators, 4(6): 67–??, November/December 1995. CODEN SYADE7. ISSN 1061-2688.

Slater:1994:CW

[Sla94] A. F. Slater. Controlled by the Web. Computer Networks and ISDN Systems, 27(2):289– 295, November 1, 1994. CO-DEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/comnet/cas_sub/
browse/browse.cgi?year=1994&
volume=27&issue=2&aid=1395.

Sloan:2005:HPL

[Slo05] Joseph D. (Joseph Donald)

Sloan. High performance Linux clusters with OSCAR, Rocks, openMosix, and MPI. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2005. ISBN 0-596-00570-9. xv + 350 pp. LCCN QA76.58; QA76.58 .S56 2005eb; QA76.58 .S56 2005; QA76.58 .S58 2005; QA76.58 .S595 2005. URL http:/

/www.oreilly.com/catalog/
9780596005702.

Smith:1977:PMT

[SM77a] D. W. Smith and J. R.

Mashey. PWB/MM tutorial (T.4). Technical Memorandum 1191, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1977. 56 pp.

Smith:1977:SPM

[SM77b] D. W. Smith and John R. Mashey. Summary of PWB/ MM macros, strings, and registers (T.2). Technical Memorandum 1216, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1977. ?? pp.

Snyder:1979:UTD

[SM79] G. A. Snyder and John R. Mashey. UNIX/TS documentation road map. Technical Memorandum 1485 (TM 79-3646-1, TM 79-9322-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 9, 1979. ?? pp.

Symbolics:1985:VUM

[SM85] Symbolics, Inc. and Mathlab Group. VAX UNIX MAC-SYMA reference manual. Symbolics, Inc., 11 Cambridge Center, Cambridge MA 02142, USA, version 11 edition, 1985. vii + 479 pp. Document number SMI0501030.011, corresponding to MACSYMA release 309.1.

Smith:1988:ECM

[SM88a] Jonathan M. Smith and Gerald Q. Maguire, Jr. Effects of copy-on-write memory management on the response time of UNIX fork operations. In Association [Ass88b], pages 255– 278.

Smith:1988:ECW

[SM88b] Jonathan M. Smith and Gerald Q. Maguire, Jr. Effects of copy-on-write memory management on the response time of UNIX fork operations. In Association [Ass88b], pages 255– 278.

Smith:COMPSYS-1-3-255

[SM88c] Jonathan M. Smith and Gerald Q. Maguire, Jr. Effects of copy-on-write memory management on the response time of UNIX fork operations. Computing Systems, 1(3):255–278, Summer 1988. CODEN CM-SYE2. ISSN 0895-6340.

Spencer:1989:FAA

[SM89] Jon F. Spencer and Jackie McAlexander. Factors affecting application portability to a B1 level trusted UNIX. In USENIX Association [USE89e], pages 239–255. LCCN QA 76.76 O63 U83 1989.

Smaha:1988:PVF

[Sma88] Steve Smaha. PC/VI faithfully replicates capabilities of UNIX original. C Users Journal, 6(5): 64–??, May 1988. ISSN 0898-9788.

Stein:1994:SUG

[SMD^{+}94] Lincoln Stein, Andre Marquis, Robert Dredge, Mary Pat Reeve, Mark Daly, Steve Rozen, and Nathan Goodman. Splicing UNIX into a genome mapping laboratory. In USENIX [USE94e], pages 221–229. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

URL  http://www.usenix.
org/publications/library/

                                        proceedings/bos94/stein.
                                        html.

Smith:1976:NTF

[Smi76] D. W. Smith. NROFF/TROFF formatting codes for departmental organization directories on PWB/UNIX. Technical Memorandum 1115, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 1, 1976. ?? pp.

Smith:1977:NTF

[Smi77] D. W. Smith. New text formatting software for the PWB systems. Technical Memorandum 1179, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 2, 1977. ?? pp.

Smith:1978:APM

[Smi78] N. W. Smith. ATS to PWB/ MM converter. Technical Memorandum 1300, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 29, 1978. ?? pp.

Smith:1983:UWW

[Smi83]  Charles R. Smith.  UNIX

writer’s workbench. In Software Tools Users Group [Sof83], pages 321–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Smith:1989:DRA

[Smi89] Griffith G. Smith, Jr. A distributed resource allocator for UNIX systems. In USENIX Association [USE89e], pages 95– 108. LCCN QA 76.76 O63 U83 1989.

Smith:1991:OPX

[Smi91] Jerry D. Smith. Object-oriented programming with the X Window System toolkits. Wiley, New York, NY, USA, 1991. ISBN 0-471-53260-6, 0-471-53259-2 (paperback). xiv + 576 pp. LCCN QA76.64 .S58 1991.

Smith:1992:DXC

[Smi92] Jerry D. Smith. Designing X clients with Xt/Motif. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1992. ISBN 1-55860-255-0. xvi + 341 pp. LCCN QA76.76.W56 S56 1992.

Smith:1999:UCW

[Smi99] Roderick W. Smith. Using Corel WordPerfect 8 for Linux. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-7897-2032-9. xiii + 841 pp. LCCN Z52.5.W65 S6 1999. Edited by Jim Minatel.

Smith:2000:LHH

[Smi00a] Roderick W. Smith. Linux Hardware Handbook. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31918-7. xix + 695 pp. LCCN QA76.76.O63 S58846 2000.

Smith:2000:LNY

[Smi00b] Roderick W. Smith. Linux:

networking for your office. SAMS Publishing, Indianapolis, IN, USA, 2000. ISBN 0-672-31792-3. xv + 618 pp. LCCN QA76.76.O63 S5885 2000.

Smith:2001:LSS

[Smi01] Roderick W. Smith. Linux Samba server administration. The Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2740-1 (paperback). xxiii + 629 pp. LCCN QA76.76.O63 S58847 2001.

Smith:2002:LOT

[Smi02a] Kevin Smith. Linux, OpenBSD, and Talisker: a comparative complexity analysis. Master’s thesis, Naval Postgraduate School, Monterey, CA, USA, 2002. 153 pp.

Smith:2002:ALN

[Smi02b] Roderick W. Smith. Advanced Linux networking. Addison-Wesley, Reading, MA, USA, 2002. ISBN 0-201-77423-2. xviii + 752 pp. LCCN QA76.76.O63 S58845 2002.

Smith:2003:FCR

[Smi03a] Roderick W. Smith. FreeBSD: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, 2003. ISBN 0-07-222409-6. xxxvi + 869 pp. LCCN QA76.76.O63 S588455 2003. URL ftp:

//uiarchive.cso.uiuc.edu/
pub/etext/gutenberg/;   http:
//www.loc.gov/catdir/bios/
mh042/2003272231.html;
http://www.loc.gov/catdir/
description/mh041/2003272231.
html;   http://www.loc.gov/

catdir/toc/mh041/2003272231.

html. Includes CD-ROM.

Smith:2003:LPT

[Smi03b] Roderick W. Smith. Linux power tools. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2003. ISBN 0-7821-4226-5 (paperback). xxiv + 752 pp. LCCN QA76.76.O63 S5886

2003. URL http://www.loc.
gov/catdir/enhancements/
fy0615/2003101647-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0615/2003101647-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy0615/2003101647-t.html.

Smith:2005:LNS

[Smi05a] Peter G. Smith. Linux network security. Administrator’s advantage series. Charles River Media, Inc., Hingham, MA, USA, 2005. ISBN 1-58450-396-3 (pbk. with CD-ROM). xviii + 541 pp. LCCN TK5105.59 .S59 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip056/2005000312.html.

Smith:2005:LWW

[Smi05b] Roderick W. Smith. Linux in a Windows world. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2005. ISBN 0-596-00758-2. xiv + 478 pp. LCCN QA76.76.O63 S548 2005. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596007584.

Smith:2006:BSE

[Smi06a] Larry L. Smith. BASH Shell: Essential Programs for Your Survival at Work. BookSurge Publishing, ????, 2006. ISBN 1-4196-4833-0. 590 (est.) pp. LCCN ???? US$32.99.

Smith:2006:HUU

[Smi06b] Larry L. Smith. How To Use the UNIX-LINUX vi Text Editor: Tips, Tricks, and Techniques (And Tutorials Too!). BookSurge Publishing, Charleston, SC, USA, 2006. ISBN 1-4196-4834-9. 56 pp. LCCN ????

Smith:2006:KSK

[Smi06c] Larry L. Smith. Korn Shell / ksh: Essential Programs for Your Survival at Work. Book-Surge Publishing, Charleston, SC, USA, 2006. ISBN 1-4196-4831-4. 588 pp. LCCN ????

Smith:2012:LE

[Smi12] Roderick W. Smith. Linux essentials. Wiley, New York, NY, USA, 2012. ISBN 1-118-10679-2 (paperback). xxiii + 344 pp. LCCN QA76.76.O63 S5884584

                                        2012. URL http://proquest.
                                        safaribooksonline.com/9781118106792.

Satyanarayanan:1994:LRV

[SMK^{+}94] M. Satyanarayanan, Henry H. Mashburn, Puneet Kumar, David C. Steere, and James J. Kistler. Lightweight recoverable virtual memory. ACM Transactions on Computer Systems, 12(1):33–57, February 1994. CODEN AC-SYEC. ISSN 0734-2071 (print), 1557-7333 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
tocs/1994-12-1/p33-satyanarayanan/

.

Schroeder:2006:VTO

[SML06] Will Schroeder, Ken Martin, and Bill Lorensen. The visualization toolkit: an objectoriented approach to 3D graphics [visualize data in 3D — medical, engineering or scientific; build your own applications with C++, Tcl, Java or Python; includes source code for VTK (supports UNIX, Windows and Mac)]. Kitware, Clifton Park, NY, fourth edition, 2006. ISBN 1-930934-19-X. xvi + 512 pp. LCCN ????

Strang:1988:TT

[SMO88] John Strang, Linda Mui, and Tim O’Reilly. termcap & terminfo. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, April 1988. ISBN 0-937175-22-6. xv + 253 pp. LCCN QA76.76.O63 S765 1990; QA76.8.U65 S79 1988. US$21.95. URL http:/

/www.oreilly.com/catalog/
9780937175224;   http://
www.oreilly.com/catalog/
term.

Snevely:2001:EDC

[Sne01] Rob Snevely. Enterprise Data Center: Design and Methodology. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-047393-6. xxii + 198 pp. LCCN TK5103.S59 2002. US$49.95. URL books/

                                        edcdesign.pdf;   http://
                                        www.sun.com/books/catalog/
                                        snevely/index.html. Part

No. 816-2765-10 December 2001, Revision 01.

Schumer:1995:UU

[SNG95] Larry Schumer, Chris Negus, and Dave Gunter. Using UNIX. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0253-3. xxiv + 948 pp. LCCN QA76.76.O63 S438 1995.

Saarinen:2022:DRV

[SNM22] Markku-Juhani O. Saarinen, G. Richard Newell, and Ben Marshall. Development of the RISC-V entropy source interface. Journal of Cryptographic Engineering, 12(4): 371–386, November 2022. CO-DEN ???? ISSN 2190-8508 (print), 2190-8516 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s13389-021-00275-6.

Snyder:1975:PCL

[Sny75]  Alan Snyder.  A portable

compiler for the language C. M.S. thesis, Project MAC, Massachusetts Institute of Technology, Cambridge, MA 02139, USA, May 1975. 74 pp. URL http:

//publications.csail.mit.
edu/lcs/specpub.php?id=717.

Based on a thesis submitted to the Department of Electrical Engineering at MIT on 10 May 1974 in partial fulfillment of the requirements for the degree of Bachelor of Science and Master of Science.

Snyder:1978:STS

[Sny78] G. A. Snyder. Shcvt — a tool for shell conversion. Technical Memorandum 1244 (MF 78-8234-32), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 7, 1978. ?? pp.

Seltzer:1992:LPM

[SO92] Margo Seltzer and Michael Olson. LIBTP: Portable, modular transactions for UNIX. In USENIX [USE92d], pages 9–26.

Sobell:1989:PGU

[Sob89] Mark Sobell. A Practical Guide to the UNIX System. Benjamin/Cummings Pub. Co., Benjamin/Cummings Pub. Co., second. edition, 1989. ISBN 0-8053-0243-3. xxv + 632 pp. LCCN QA76.76.O63 S595 1989. Similar to Christian’s book [Chr88] … slightly easier to read … There is a new edition for System V Release 4 ….

Sobell:1991:PGU

[Sob91] Mark Sobell. A Practical Guide to the UNIX System V Release 4. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, second edition, 1991. ISBN 0-8053-7560-0. xxvii + 700 pp. LCCN QA76.76.O63 S6 1991. A very good tutorial/reference book.

Sobell:1995:PGU

[Sob95] Mark G. Sobell. A Practical Guide to the UNIX System. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, third edition, 1995. ISBN 0-8053-7565-1. xxxii + 800 pp. LCCN QA76.76.063S595 1994. US$48.95.

Sobell:1997:PGL

[Sob97] Mark G. Sobell. A Practical Guide to Linux. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-89549-8. xlvii + 1015 pp. LCCN QA76.76.O63S5948 1997. US$37.61. Forward by Linus Torvalds.

Sobell:1999:PGS

[Sob99] Mark G. Sobell. A Practical Guide to Solaris. Addison-Wes-ley, Reading, MA, USA, 1999. ISBN 0-201-89548-X. 1120 pp. LCCN QA76.76.O63S5949 1999. US$44.95.

Sobell:2003:PGR

[Sob03] Mark G. Sobell. A Practical Guide to Red Hat Linux 8. Addison-Wesley, Reading, MA, USA, 2003. ISBN 0-201-70313-0. xlvii + 1565 pp. LCCN QA76.76.O63 S59485 2002. Includes two CD-ROMs.

Sobell:2006:PGR

[Sob06] Mark G. Sobell. A practical guide to Red Hat Linux: Fedora Core and Red Hat Enterprise Linux. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third edition, 2006. ISBN 0-13-228027-2 (paperback). 1168 (est.) pp. LCCN QA76.76.O63 S59485 2006. URL http://

vig.prenhall.com/catalog/
academic/product/0,1144,
0132280272,00.html;   http:
//www.loc.gov/catdir/toc/
ecip0613/2006014003.html.

Includes DVD.

STUG:1983:PUA

[Sof83] Software Tools Users Group, editor. Proceedings: USENIX Association [and] Software Tools Users Group Summer Conference, Toronto 1983, July 1983, Toronto, Ontario, Canada. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1983. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

STUG:1984:UAS

[Sof84] Software Tools Users Group, editor. USENIX Association [and] Software Tools Users Group Summer Conference, Salt Lake City 1984: proceedings, June 12–15, 1984, Salt Lake City, Utah, USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1984. LCCN QA76.8.U65 U83 1984.

Sonnenschein:1987:GTV

[Son87] Dan Sonnenschein. A guide to vi — visual editing on the UNIX system. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1987. ISBN 0-13-371311-3. xix + 180 pp. LCCN QA76.76.T49 S66 1987.

Sorfa:2001:OSI

[Sor01] Petr Sorfa. Open source IDEs for Linux/Unix. C/C++ Users Journal, 19(3):8–??, March 2001. CODEN CCUJEX. ISSN 1075-2838.

Southerton:1992:MU

[Sou92]  Alan Southerton.  Modern

UNIX. Wiley-Interscience, New York, NY, USA, 1992. ISBN 0-471-54916-9. xvii + 308 pp. LCCN QA76.76.O63 S65497 1993. Covering selected topics like shells, X Windows, networking.

Southerton:1994:SHG

[Sou94] Alan Southerton. The Shell Hacker’s Guide to X and Motif — Custom Power Tools and Window Manager Tricks. Wiley-Interscience, New York, NY, USA, 1994. ISBN 0-471-59723-6 (book/disk), 0-471-59722-8 (paper). xvii + 382 pp. LCCN QA76.76.W56 S66 1994.

Sovereign:1992:PPT

[Sov92] John Sovereign. The power of Posix thinking. UNIX/World, 9(7):93–??, July 1, 1992. ISSN 0739-5922.

Smith:1977:TDP

[SP77] D. W. Smith and E. M. Piskorik. Typing documents with PWB/MM (T.3). Technical Memorandum 1190, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 1977. 8 pp.

Step90a

[SP90] Bruce R. Stephens and John D.

Pryce. The DAPRE/UNIX Preprocessor Users’ Guide v1.2. Royal Military College of Science at Shrivenham, 1990.

Santos:2015:BD

[SP15] Rui Santos and Luis Perestrelo. BeagleBone for dummies. Wiley, New York, NY, USA, 2015. ISBN 1-118-99291-1 (paperback). ???? pp. LCCN ????

Spencer:1996:KYP

[Spe96] Will Spencer. Keep your password to yourself — here are some of the most useful software packages for UNIX security. Datamation, 42(9):76–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Spector:2000:BLC

[Spe00] David H. M. Spector. Building Linux clusters: scaling Linux for scientific and enterprise applications. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2000. ISBN 1-56592-625-0. xviii + 332 pp. LCCN QA76.76.O63 S6647 2000. URL http:/

/www.oreilly.com/catalog/
9781565926257. Includes CD-

ROM with Red Hat Linux clustering software, cluster management tools and scripts, and parallel programming tools.

Spinellis:1994:WUT

[Spi94] Diomidis D. Spinellis. Wux: UNIX tools under Windows. In USENIX [USE94f], pages 325–336. ISBN 1-880446-58-

8. LCCN QA 76.76 O63 U84 1994. URL http:

//kerkis.math.aegean.gr/~
dspin/pubs/conf/1994-Usenix-
Wux/html/wux.html;   http:
//www.usenix.org/publications/
library/proceedings/sf94/
spin.html.

Spinellis:2000:OUT

[Spi00] Diomidis D. Spinellis. Outwit: UNIX tool-based programming meets the Windows world. In USENIX [USE00a], pages 149–158. ISBN 1-880446-22-7. LCCN ???? URL http:

//www.usenix.org/events/
usenix2000/general/spinellis.
html.

Spinellis:2005:WUT

[Spi05]  Diomidis Spinellis.  Work-

ing with Unix tools. IEEE Software, 22(6):9–11, November/December 2005. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Spinellis:2017:RUH

[Spi17] Diomidis Spinellis. A repository of Unix history and evolution. Empirical Software Engineering, 22(3):1372– 1404, June 2017. CO-DEN ESENFW. ISSN 1382-3256 (print), 1573-7616 (elec-

tronic). URL http://link.
springer.com/article/10.
1007/s10664-016-9445-5;
http://www.spinellis.gr/
pubs/jrnl/2016-EMPSE-unix-
history/html/unix-history.
html;   http://www.spinellis.
gr/pubs/jrnl/2016-EMPSE-
unix-history/html/unix-history.
pdf.

Sprang:2006:XVL

[Spr06] Henning Sprang. Xen: Virtualisierung unter Linux. Open Source Press, München, Germany, 2006. ISBN 3-937514-29-

5. 350 pp. LCCN ???? ca. EUR 39.90, EUR 41.35 (AT).

Sprang:2007:XVL

[Spr07] Henning Sprang, editor. Xen:

Virtualisierung unter Linux. (German) [Xen: Virtualization under Linux]. Open Source Press, München, Germany, 2007. ISBN 3-937514-29-5. 350 pp. LCCN ???? URL http:

//deposit.ddb.de/cgi-bin/
dokserv?id=2809360.

SPSS:1990:SUO

[SPS90] SPSS, Inc. SPSS for UNIX:

Operations Guide. SPSS, Inc. and Prentice-Hall, Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA and Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-918469-97-X (SPSS), 0-13-177981-8 (PH). 152 pp. LCCN HA32 .S625 1990 Reserve Controlled Circulation. US$12.00.

SPSS:1993:SUA

[SPS93a] SPSS, Inc. SPSS for UNIX Advanced Statistics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-107137-8. US$18.00.

SPSS:1993:SUB

[SPS93b] SPSS, Inc. SPSS for UNIX Base Systems User’s Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-107145-9. US$24.75.

SPSS:1993:SUP

[SPS93c] SPSS, Inc. SPSS for UNIX Professional Statistics. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1993. ISBN 0-13-107152-1. US$18.00.

SPSS:1995:SBS

[SPS95] SPSS, Inc. SPSS 6.1 Base System User’s Guide, Part 1: UNIX Version. SPSS, Inc., Suite 3300, 444 N. Michigan Ave., Chicago, IL, USA, 1995. ISBN 0-13-459561-0. xvii + 341 pp.

Stallman:2002:DGG

[SPS^{+}02] Richard M. Stallman, Roland Pesch, Stan Shebs, et al. Debugging with GDB: The GNU Source-Level Debugger. GNU Press, Boston, MA, USA, 2002. ISBN 1-882114-88-4. viii + 344 pp. LCCN QA76.9.D43 D422 2003. US$25.00. URL http:

//www.gnupress.org/book7.
html.

Spurgeon:1992:NRL

[Spu92] Charles Spurgeon. Network Reading List: TCP/IP, UNIX and Ethernet, February 1992. This annotated list describes those items that cover the subject areas well. It is obtainable by anonymous ftp from ftp.

uu.net (in /inet/doc).

Shapiro:1977:GGD

[SR77] E. M. Shapiro and M. F.

Robbins. Gf — a graphical data analysis: Program for the MERT-UNIX operating system. Technical Memorandum 1222 (TM 77-1131-44), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 21, 1977. ?? pp.

Scheer:1985:UAR

[SR85a] M. D. Scheer and S. Rajeev. A UNIX-based Ada runtime system. In USENIX Association [USE85c], pages 51–?? Abstract only.

Scheer:1985:UBA

[SR85b] M. D. Scheer and S. Rajeev. A UNIX-based Ada runtime system. In USENIX [USE85b], pages 51–?? Abstract only.

Shirley:1995:MRP

[SR95] John Shirley and Ward Rosenberry. Microsoft RPC Programming Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, March 1995. ISBN 1-56592-070-8. xix + 232 pp. LCCN QA76.9.D5S554 1995. US$24.95.

Stevens:2004:APU

[SR04] W. Richard Stevens and Stephen Rago. Advanced Programming in the UNIX Environment. Addison-Wesley, Reading, MA, USA, second edition, 2004. ISBN 0-201-43307-9. ???? pp. LCCN QA76.76.O63 S754 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.loc.gov/catdir/
toc/ecip059/2005007943.html.

Stevens:2005:APU

[SR05] W. Richard Stevens and Stephen A. Rago. Advanced Programming in the Unix Environment. Addison-Wesley, Reading, MA, USA, second edition, 2005. ISBN 0-201-43307-9 (hardcover). xxviii + 927 pp. LCCN QA76.76.O63 S754 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/
;   http://www.aw-bc.com/
catalog/academic/product/
0,1144,0201433079,00.html;
 http://www.loc.gov/catdir/
toc/ecip059/2005007943.html.

Stevens:2008:APU

[SR08] W. Richard Stevens and Stephen A. Rago. Advanced programming in the UNIX environment. Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, second edition, 2008. ISBN 0-321-52594-9. xxviii + 927 pp. LCCN QA76.76.O63.

Stevens:2013:APU

[SR13] W. Richard Stevens and Stephen A. Rago. Advanced programming in the UNIX environment. The Addison-Wesley professional computing series. Addison-Wesley, Reading, MA, USA, third edition, 2013. ISBN 0-321-63773-9 (paperback). xxxii + 994 pp. LCCN QA76.76.O63 S754 2013.

Smart:1999:UCO

[SRBB99] Allan Smart, Erik Ratcliffe, Tim Bird, and David Bandel. Using Caldera OpenLinux, Special Edition. Que Corporation, Indianapolis, IN, USA, 1999. ISBN 0-7897-2058-2. xxii + 1208 pp. LCCN A76.76.O63 U7118 1999. US$39.99.

Srinivasan:1993:UDP

[Sri93] Balasubramaniam Srinivasan.

UNIX Document Processing and Typesetting. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1993. ISBN 981-02-0605-4. xiv + 445 pp. LCCN Z52.5.U54S65 1993.

Santana:2015:FSS

[SRTH15] Ricardo Santana, Raju Rangaswami, Vasily Tarasov, and Dean Hildebrand. A fast and slippery slope for file systems. Operating Systems Review, 49 (2):27–34, December 2015. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Sauers:2004:HUT

[SRW04] Robert (Robert F.) Sauers, Chris P. Ruemmler, and Peter Weygant. HP-UX 11i tuning and performance. Hewlett– Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-143349-0 (paperback). xxvi + 530 pp. LCCN QA76.76.O63 S35625 2004.

Szczur:1993:TPT

[SS93] Martha R. Szczur and Sylvia B. Sheppard. TAE plus: Transportable applications environment plus: a user interface development environment. ACM Transactions on Information Systems, 11(1):76–101, January 1993. CODEN ATISET. ISSN 1046-8188. URL http://www.

                                        acm.org:80.

Schmidt:1995:UDP

[SS95a] D. C. Schmidt and P. Stephenson. Using design patterns to evolve system software from UNIX to Windows NT. C++ Report, 7(3):47–60, March– April 1995. CODEN CRPTE7. ISSN 1040-6042.

Small:1995:SAB

[SS95b] Christopher Small and Margo Seltzer. Scheduler activations on BSD: sharing thread management between kernel and application. Technical Report 31-95, Center for Research in Computing Technology, Harvard University, Cambridge, MA, USA, 1995. 12 pp.

Stone:2000:UFM

[SS00] Brad Stone and Julie Symons.

UNIX fault management: a guide for system administration. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2000. ISBN 0-13-026525-X. xvi + 353 pp. LCCN QA76.76.O63 S7594 2000 Bar.

Stanfield:2001:LSA

[SS01] Vicki Stanfield and Roderick W. Smith. Linux system administration. The Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2735-5 (paperback). xxviii + 657 pp. LCCN QA76.76.O63 S7346 2001.

Stanfield:2002:LSA

[SS02] Vicki Stanfield and Roderick W. Smith. Linux system administration. Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, second edition, 2002. ISBN 0-7821-4138-2. xxvii + 624 pp. LCCN QA76.76.O63 S7346

                                        2002. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0613/2002106413-b.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0613/2002106413-
                                        d.html;   http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0613/2002106413-t.html.

Schoblick:2005:DGL

[SS05] Gabriele Schoblick and Robert Schoblick. Debian GNU/Linux. bhv, Bonn, Germany, 2005. ISBN 3-8266-8151-7. 767 pp. LCCN ???? EUR22.95. Includes two CD-ROMs.

Sobell:2006:PGU

[SS06] Mark G. Sobell and Peter Seebach. A practical guide to Unix for Mac OS X users. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-186333-9 (paperback). xxxix + 999 pp. LCCN QA76.76.O63 S597 2006.

Shaver:1992:ILF

[SSB92] Dave Shaver, Eric Schnoebelen, and George Bier. An implementation of large files for BSD UNIX. In USENIX [USE92d], pages 61–68.

SSC:1984:SRC

[SSC93] SSC staff. SCC Reference Cards. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1984–1993. These are some good, inexpensive reference/tutorial cards on UNIX commands, Bourne shell, Korn shell, emacs, vi, C, C++, etc …. e.g. the new “UNIX System Command Summary for SVR4.2/Solaris 2.1” (ISBN: 0-916151-61-1) …. Contact Belinda Frazier (bel@

ssc.com) or sales@ssc.com

for more info.

SSC:1995:TTR

[SSC95] SSC. Tcl and Tk Reference Card. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1995. ISBN 0-916151-80-8.

Silverstein:1990:USM

[SST90] Joel D. Silverstein, Steven E.

Sommars, and Yio-Chian Tao. The UNIX system math library, a status report. In USENIX [USE90d], pages 117– 131. LCCN QA76.8.U65 U82 1990.

Shienbrood:1983:UAC

[SSWW83] Eric R. Shienbrood, Carl A.

Soeder, James R. Ward, and Kincade N. Webb. UNIX on Apollo computers (yet another UNIX emulation). In Association [Ass83], pages 133–142.

Smith-Thomas:1989:SML

[ST89a] Barbara Smith-Thomas. Secure multi-level windowing in a B1 certifiable secure UNIX operating system. In USENIX [USE89b], pages 429–439.

Smith-Thomas:1989:SMW

[ST89b] Barbara Smith-Thomas. Secure multi-level windowing in a B1 certifiable secure UNIX operating system. In USENIX Association [USE89f], pages 429– 439. Describes the architecture of the multi-level version of the AT&T 630 graphics terminal. This terminal was evaluated as part of AT&T System V/MLS, which received a B1 rating.

Smith-Thomas:SML89

[ST89c] Barbara Smith-Thomas. Secure multi-level windowing in a B1 certifiable secure UNIX operating system. In Proceedings of the USENIX Winter 1989 Conference. USENIX Association, Berkeley, CA, USA, January 1989. Describes the architecture of the multi-level version of the AT&T 630 graphics terminal. This terminal was evaluated as part of AT&T System V/MLS, which received a B1 rating.

Shimizu:2003:TLS

[ST03] Naohiko Shimizu and Ken Takatori. A transparent Linux super page kernel for Alpha, Sparc64 and IA32: reducing TLB misses of applications. ACM SIGARCH Computer Architecture News, 31(1): 75–84, March 2003. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Straathof:1986:USL

[STA86] Jeffrey H. Straathof, Ashok K. Thareja, and Ashok K. Agrawala. UNIX scheduling for large systems. In USENIX Association [USE86e], pages 111–139. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

Straathof:1987:MRP

[STA87] Jeffrey H. Straathof, Ashok K.

Thareja, and Ashok K. Agrawala. Methodology and results of performance measurements for a new UNIX scheduler. In USENIX Association [USE87e], pages 165–180. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Staubach:1989:UWT

[Sta89] Gottfried Staubach. UNIX-Werkzeuge zur Textmusterverarbeitung: Awk, Lex und Yacc. (German) [UNIX Tools for Text Processing: Awk, Lex, and Yacc]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 3-540-51232-2. x + 157 pp. LCCN ????

Stallman:1993:GEM

[Sta93] Richard M. Stallman. GNU EMACS Manual. Free Software Foundation, Inc., 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA, Tel:

(617) 876-3296, ninth edition, August 1993. ISBN 1-882114-03-5. This is the official manual for GNU Emacs. It is available both as a typeset document, and online in the Emacs info system.

Stanik:2004:NNW

[Sta04] John Stanik. News 2.0: New Web provides battery-powered option; there’s nothing like a map; shall I compare thee to a big squeeze?; Linux support: a black hole? ACM Queue: Tomorrow’s Computing Today, 2 (8):14, November 2004. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NLO

[Sta05a] John Stanik. News 2.0: Losing our edge? The real cost of Linux; say no to crackberries. ACM Queue: Tomorrow’s Computing Today, 3 (5):14, June 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2005:NRW

[Sta05b]  John Stanik.  News 2.0:

“Ransom-ware” on the loose; adopt IPv6 or die; IBM steps up on Linux training. ACM Queue: Tomorrow’s Computing Today, 3(6):10, July 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2006:NML

[Sta06] John Stanik. News 2.0: The mobile Linux challenge; Google not into googling; keeping online video legit. ACM Queue: Tomorrow’s Computing Today, 4(8):8, October 2006. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Stanik:2007:NGD

[Sta07a] John Stanik. News 2.0: Googlebombs away; DST: This year’s Y2K?; Linux through Windows. ACM Queue: Tomorrow’s Computing Today, 5 (2):7, March 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

State:2007:REL

[Sta07b] Radu State. Review of ‘Embedded Linux Primer: A Practical Real-world Approach’ by Christopher Hallinan, Prentice Hall PTR, 2006, $44.99, ISBN 0-13-167984-8. ACM Queue: Tomorrow’s Computing Today, 5(3):44, April 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic). See [Hal07].

Steffen:1985:IEC

[Ste85] Joseph L. Steffen. Interactive examination of a C program with Cscope. In USENIX Association [USE85c], pages 170– 175.

Stewartson:1988:UVB

[Ste88] Ian Stewartson. UNIX V.3 and beyond. In USENIX Association [USE88c], pages 161–177. ISBN ???? LCCN ????

Steels:1990:EEC

[Ste90a] Luc Steels, editor. Europal 90: Proceedings of the First European Conference on the Practical Application of Lisp, Churchill College, Cambridge. Europal, Dorking, UK, 1990. ISBN ???? LCCN ????

Stevens:1990:UNP

[Ste90b] W. Richard Stevens. UNIX Network Programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-949876-1. xi + 772 pp. LCCN QA76.76.O63 S755 1990.

Stern:1991:MNN

[Ste91] Hal Stern. Managing NFS and NIS. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-75-7. xxiv + 410 pp. LCCN TK5105.5 .S74 1991. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780937175750;   http://
                                        www.oreilly.com/catalog/
                                        nfs.

Sterling:1992:HC

[Ste92a] Bruce Sterling. Hacker Crackdown. Bantam Books, New York, NY, USA, 1992. ISBN 0-553-56370-X. xiv + 328 pp. LCCN HV6773.2 .S74 1992. US$23.00. Also available by anonymous ftp from

                                        mrcnext.cso.uiuc.edu (as /
                                        etext/etext94/hack11.zip).

See also [Sto90].

Stevens:1992:APU

[Ste92b]  W. Richard Stevens.  Ad-

vanced Programming in the UNIX Environment. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56317-7. xviii + 744 pp. LCCN QA76.76.O63 S754 1992. US$52.25. Highly recommended. A going-to-be classic on how programs work under UNIX … The source codes and errata list are obtainable by anonymous ftp from

ftp.uu.net (in /published/
books).

Stevens:APU92

[Ste92c]  W. Richard Stevens.  Ad-

vanced Programming in the UNIX Environment. Addison-Wesley, Reading, MA, USA, 1992. ISBN 0-201-56317-7. xviii + 744 pp. LCCN QA76.76.O63 S754 1992. US$52.25. The source codes and errata list are obtainable by anonymous

ftp from ftp.uu.net (in /
published/books).

Stevens:1994:TII

[Ste94] W. Richard Stevens. TCP/IP Illustrated, Volume 1: The Protocols. Addison-Wesley Professional Computing Series. Addison-Wesley, Reading, MA, USA, 1994. ISBN 0-201-63346-9. xix + 576 pp. LCCN TK5105.55 .S74 1994. US$65.95.

Stevens:1996:TII

[Ste96] W. Richard Stevens. TCP/IP Illustrated, Volume 3: TCP for Transactions, HTTP, NNTP, and the UNIX Domain Protocols. Addison-Wesley Professional Computing Series. Addison-Wesley, Reading, MA, USA, 1996. ISBN 0-201-63495-3. xix + 328 pp. LCCN TK5105.55.S74 1994. US$41.47.

Stevens:1998:UNP

[Ste98] W. Richard Stevens. UNIX Network Programming: Networking APIs: Sockets and XTI, volume 1. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1998. ISBN 0-13-490012-X. xx + 1009 pp. LCCN QA76.76.O63S755 1998. US$59.00. URL http:

//www.phptr.com/ptrbooks/
ptr_013490012X.html.

Stephenson:1999:BLW

[Ste99a] Neal Stephenson. In the beginning … was the command line. Avon Books, New York, NY, USA, 1999. ISBN 0-380-81593-1. 151 pp. LCCN QA76.76.O63 S7369 1999. US$10.00.

Stevens:1999:UNP

[Ste99b] W. Richard Stevens. UNIX Network Programming, Interprocess Communications, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1999. ISBN 0-13-081081-9 (hardcover). xvii + 558 pp. LCCN QA76.76.O63S755 1998. US$54.00. URL http:/

/www.phptr.com/ptrbooks/
ptr_0130810819.html.

Stevens:2000:CPG

[Ste00] Al Stevens. C programming:

Going undercover. Dr. Dobb’s Journal of Software Tools, 25 (7):113–117, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
cprog.txt.

Sterling:2001:BCCa

[Ste01a] Thomas L. Sterling, editor. Beowulf Cluster Computing with Linux. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-69274-0. xxxiii + 496 pp. LCCN QA76.58 .B46 2002. US$42.95, UK£28.50.

Stevens:2001:CPMa

[Ste01b] Al Stevens. C programming:

a moving target. Dr. Dobb’s Journal of Software Tools, 26 (4):129–130, 132–133, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Stevens:2001:CPA

[Ste01c] Al Stevens. C programming: I almost get a Linux editor and compiler. Dr. Dobb’s Journal of Software Tools, 26(7):113–116, July 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Stevens:2001:CPG

[Ste01d] Al Stevens. C programming:

It’s good work when you can find it. Dr. Dobb’s Journal of Software Tools, 26(5):121–124, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Stevens:2001:CPR

[Ste01e]  Al Stevens.  C program-

ming: Road rage = editors + drivers. Dr. Dobb’s Journal of Software Tools, 26(6):131–134, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Stevens:2001:CPS

[Ste01f]  Al Stevens.  C program-

ming: Software development, Linux, and the White House. Dr. Dobb’s Journal of Software Tools, 26(3):124–127, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/.

Stevens:2001:CPYb

[Ste01g]  Al Stevens.  C program-

ming: YAPP: Yet Another Programming Platform. Dr. Dobb’s Journal of Software Tools, 26(10):105–107, 109, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/.

Steinberger:2002:VMU

[Ste02a]  Richard Steinberger.  Vul-

nerability management in Unix environments. Information Security Technical Report, 7(1):26–36, March 31, 2002. CODEN ISTRFR.

ISSN 1363-4127 (print), 1873-605X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S1363412702001048.

Sterling:2002:BCC

[Ste02b] Thomas Lawrence Sterling. Beowulf cluster computing with Linux. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 2002. ISBN 0-262-69274-0. xxxiii + 496 pp. LCCN QA76.58 .B46 2002.

Stitt:1983:RDM

[Sti83]  F. W. Stitt.  Research

database management software for UNIX-based microcomputers. In Association [Ass83], pages 201–209.

Storm:1976:IGT

[Sto76] A. R. Storm. Instructions for GRAPH, TEK, TEKSTARE and GSIP graphics packages. Technical Memorandum 1110 (TM 76-1527-35), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 30, 1976. 9 pp.

Stonebraker:1980:RDS

[Sto80] Michael Stonebraker. Retrospection on a database system. ACM Transactions on Database Systems, 5(2):225– 240, June 1980. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic).

URL http://www.acm.org/
pubs/articles/journals/tods/
1980-5-2/p225-stonebraker/
p225-stonebraker.pdf;   http:

//www.acm.org/pubs/citations/
journals/tods/1980-5-2/p225-
stonebraker/. Reprinted in

[Sto88].

Stonebraker:1988:RDS

[Sto88] Michael Stonebraker, editor. Readings in Database Systems. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1988. ISBN 0-934613-65-6. xii + 644 pp. LCCN QA76.9.D3 R4 1988. US$29.95. Contains reprint of [Lit80].

Stoll:1989:CET

[Sto89] Clifford Stoll. The Cuckoo’s Egg: Tracking a Spy Through the Maze of Computer Espionage. Doubleday, New York, NY, USA, 1989. ISBN 0-385-24946-2, 0-307-81942-6 (ebook), 0-7434-1145-5, 0-7434-1146-3, 1-299-04734-3. vi + 326 pp. LCCN UB271.R92 H477 1989; UB271.R92 H4771 1989; UB271.R92 S47 1989.

US$18.95. URL http://vxer.
org/lib/pdf/The%20Cuckoo%
27s%20Egg.pdf.

Stoll:1990:CET

[Sto90] Cliff Stoll. The Cuckoo’s Egg — Tracking a Spy through the Maze of Computer Espionage. Pocket Books, New York, NY, USA, 1990. ISBN 0-671-72688-9. vi + 356 pp. LCCN UB 271 R92 H477 1990. A fascinating real story on computer espionage …. A good alternative to this is [HM91]. Another nonfiction on computer security is [Ste92a].

Stoughton:1993:UURd

[Sto93a] Nicholas Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(4): 34–37, July/August 1993. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stoughton:1993:UURe

[Sto93b] Nicholas Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18 (5):36–45, September/October 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stoughton:1993:UURc

[Sto93c] Nicholas M. Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(3):25–33, May/June 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stoughton:1993:UURf

[Sto93d] Nicholas M. Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(6):39–40, November/ December 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stoughton:1994:UURb

[Sto94a] Nicholas Stoughton. An update on UNIX-related activities. ;login: the USENIX Association newsletter, 19(2):34– 43, March/April 1994. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stoughton:1994:UURa

[Sto94b] Nicholas Stoughton. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 19(1): 39–45, January/February 1994. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stoughton:1994:UURc

[Sto94c] Nick Stoughton. An update in UNIX-related standards activities. ;login: the USENIX Association newsletter, 19(3): 27–30, May/June 1994. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Stone:1995:UTI

[Sto95] Donald C. Stone. UNIX, TCP/ IP, and Macs. Sys Admin: The Journal for UNIX Systems Administrators, 4(3):63–??, May/ June 1995. CODEN SYADE7. ISSN 1061-2688.

Stones:2000:PLP

[Sto00] Richard Stones. Professional Linux programming. Wrox Press, Chicago, IL, USA, 2000.

ISBN 1-86100-301-3. xviii + 1155 pp. LCCN QA76.76.O63 P754 2000.

Strom:1979:CBI

[Str79] B. I. Strom. Computer buffered information exchange user’s guide. Technical Memorandum 1406 (TM 79-5133-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 11, 1979. ?? pp.

Stroustrup:1980:CAD

[Str80a] Bjarne Stroustrup. Classes: an abstract data type facility for the C language. Computing Science Technical Report 84, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 1980. ?? pp.

Stroustrup:1980:SCC

[Str80b] Bjarne Stroustrup. A set of C classes for co-routine style programming. Computer Science Technical Report CSTR-90a, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 1980.

Stroustrup:1982:CAD

[Str82] Bjarne Stroustrup. Classes: an abstract data type facility for the C language. ACM SIG-PLAN Notices, 17(1):42–51, January 1982. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https:

//dl.acm.org/doi/10.1145/
947886.947893.

Stroustrup:1984:CPL

[Str84a] Bjarne Stroustrup. The C++ programming language — reference manual. Computing Science Technical Report 108, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1, 1984. viii + 56 pp.

URL https://cds.cern.ch/
record/169940/files/cer-
000081326.pdf.

Stroustrup:1984:CT

[Str84b] Bjarne Stroustrup. A C++ tutorial. Computing Science Technical Report 113, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 10, 1984. iv + 20 pp. URL

https://web.archive.org/
web/20191029151437/http:/
/web.eah-jena.de/~kleine/
history/languages/Stroustrup-
CplusplusTutorial.pdf.

Stroustrup:1984:DAC

[Str84c] Bjarne Stroustrup. Data abstraction in C. Computing Science Technical Report 109, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 1, 1984. iii + 62 pp. URL

https://web.archive.org/
web/20191029151522/http:/
/web.eah-jena.de/~kleine/
history/languages/Stroustrup-
DataAbstractionInC.pdf.

Stroustrup:1985:CT

[Str85] Bjarne Stroustrup. A C++ tutorial. In Proceedings of the 1985 ACM annual conference on The range of computing: mid-80;s perspective mid-80’s perspective — ACM ’85. ACM Press, 1985. URL https:

//dl.acm.org/doi/10.1145/
320435.320459.

Strang:1986:PTC

[Str86a] John Strang. Programming with curses. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1986. ISBN 0-937175-02-1. 71 pp.

Stroustrup:1986:CPL

[Str86b] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, 1986. ISBN 0-201-12078-X. viii + 327 pp. LCCN QA76.73.C153 S77 1986.

Stroustrup:1988:WOO

[Str88] Bjarne Stroustrup. What is object-oriented programming? IEEE Software, 5(3):10–20, May 1988. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic). URL

https://ieeexplore.ieee.
org/abstract/document/2020.

Strack:1990:EAC

[Str90] Hermann Strack. Extended access controls in UNIX System V: ACLs and context. In USENIX Association [USE90e], pages 87–102. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Stroustrup:1991:CPL

[Str91a] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, second edition, 1991. ISBN 0-201-53992-6. xi + 669 pp. LCCN QA76.73.C15 S79 1991.

Stroustrup:1991:WOO

[Str91b]  Bjarne Stroustrup.  What

is “Object-Oriented Programming”? Computing Science Technical Report 160, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 1991. 21 pp. 1991 revised version.

Strobel:1994:LUJ

[Str94]  Stefan Strobel.  LINUX:

Unix fur jedermann; der PC als Workstation; Installation Schritt fur Schritt, Unix-Grundlagen und TCP/IP, Editieren mit dem Emacs, DOS-Programme unter LINUX, Textverarbeitung mit LaTeX 2e, Sound, Grafik, Netzwerkspiele. Vogel-Verlag, Postfach 67 40, D-8700 Würzburg, Germany, 1994. ISBN 3-8259-1330-9. 100 pp. LCCN ???? 49.00 DM; 358.00 ÖS; 49.00 Sfr.

Stroustrup:19xx:SWS

[Strxx] Bjarne Stroustrup. Sixteen ways to stack a cat. Computing Science Technical Report 161, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx. 19 pp.

Stroustrup:2013:CPL

[Str13] Bjarne Stroustrup. The C++ Programming Language. Addison-Wesley, Reading, MA, USA, fourth edition, 2013.

ISBN 0-321-56384-0 (paperback). xiv + 1342 pp. LCCN QA76.73.C153 S77

2013.  URL http://www.
pearsonhighered.com/educator/
product/C-Programming-Language-
The/9780321563842.

Stroustrup:1986:WOO

[Str6 ]  Bjarne Stroustrup.  What

is object-oriented programming? Computing Science Technical Report 160, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1986 (??).

??  pp.  URL https://
www.softwarepreservation.
org/projects/c_plus_plus/
index.html.

Suzuki:1986:REC

[STT86a] Tatsuo Suzuki, Hideo Taniguchi, and Hisayasu Takada. A realtime electronic conferencing system based on distributed UNIX. In USENIX Association [USE86c], pages 189–199.

Suzuki:1986:RTE

[STT86b] Tatsuo Suzuki, Hideo Taniguchi, and Hisayasu Takada. A realtime electronic conferencing system based on distributed UNIX. In USENIX [USE86a], pages 189–199.

Stuck:1978:EDA

[Stu78a]  B. W. Stuck.  Engineer-

ing and data acquisition system/network management (EADAS/NM) capacity estimation — methodology and data requirements. Technical Memorandum 1449, AT&T Bell Laboratories, Murray Hill, NJ, USA, July 17, 1978. ?? pp.

Stuck:1978:ULC

[Stu78b] B. W. Stuck. Using a line clock to estimate mean disk access time. Technical Memorandum 1282 (EN 3451), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 18, 1978. ?? pp.

Sturman:1978:MSM

[Stu78c] J. N. Sturman. mview — a set of macrocommands for the generation of view graphs. Technical Memorandum 1280 (TM 78-8231-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 21, 1978. ?? pp.

Sturman:1979:LP

[Stu79] J. N. Sturman. LSX performance. Technical Memorandum 1457 (TM 79-9323-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 24, 1979. ?? pp.

Stutz:2001:LCT

[Stu01] Michael Stutz. The Linux Cookbook: Tips and Techniques for Everyday Use. Linux Journal Press, San Francisco, CA, USA, 2001. ISBN 1-886411-48-4. 396 pp. LCCN QA76.76.O63 S788 2000. US$29.95. URL

                                        http://dsl.org/cookbook/.

Prepared with TEXinfo.

Stutz:2004:LCT

[Stu04] Michael Stutz. The Linux cookbook: tips and techniques for everyday use. No Starch Press, San Francisco, CA, USA, second edition, 2004. ISBN 1-59327-031-3. xxxiv + 788 pp. LCCN QA76.76.O63 S788 2004eb; QA76.76.O63 S788 2004. URL http:/

/www.oreilly.com/catalog/
9781593270315.

Sethi:1970:GOC

[SU70] Ravi Sethi and J. D. Ullman. The generation of optimal code for arithmetic expressions. Journal of the ACM, 17(4):715–728, October 1970. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic). Reprinted as pp. 229– 247 in Compiler Techniques, B. W. Pollack, ed., Auerbach, Princeton NJ (1972).

Strobel:1994:LPW

[SU94a] Stefan Strobel and Thomas Uhl. LINUX — vom PC zur Workstation: Grundlagen, Installation und praktischer Einsatz. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58098-0, 3-540-57383-6. xii + 238 pp. LCCN ????

Strobel:1994:LWY

[SU94b] Stefan Strobel and Thomas Uhl. Linux: Unleashing the Workstation in Your PC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 3-540-58077-8 (Berlin), 0-387-58077-8 (New York). xii + 248 pp. LCCN QA76.5 .S78513 1994. DM49.00, US$29.00. Translation by Robert Bach of German title Linux — vom PC zur Workstation. See also [SU95].

Strobel:1995:CLK

[SU95] Stefan Strobel and Thomas Uhl. The Complete Linux Kit. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-14224-X (Berlin), 0-387-14224-X (New York). US$59.95. Package of one CD ROM plus two books, [SU94b, SUM95].

Strobel:1996:L

[SU96] Stefan Strobel and Thomas Uhl. Linux. Kossuth Konyvk, Budapest, Hungary, 1996. ISBN 963-09-3896-0. 366 pp. LCCN ???? Hungarian translation by K. Papp Laszlone and Tarjan Gyorgy of [SE96]. See also [SU95].

Sullivan:2002:SLG

[Sul02] Cary Sullivan, editor. Sair Linux and GNU certification: level II core concepts and practices. Wiley, New York, NY, USA, 2002. ISBN 0-471-40538-8. xvi + 399 pp. LCCN QA76.3 .S253 2002. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/wiley022/2002265408.
                                        html.

Strobel:1995:LUI

[SUM95] Stefan Strobel, Thomas Uhl, and Rainer Maurer. Linux Universe: Installation and Configuration. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 0-387-94506-7. 160 pp. LCCN A76.76.O63 S766 1995. US$34.95. Translation by Robert Bach. Includes CD ROM. See also [SU95].

Sunshine:1977:ICE

[Sun77] Carl A. Sunshine. Interprocess communication extensions for the UNIX operating system: I — design considerations. Report R-2064/1-AF, Rand Corporation, Santa Monica, CA, USA, June 1977. vii + 31

pp. URL https://apps.dtic.
mil/dtic/tr/fulltext/u2/
a044200.pdf.

Sun:1986:UTN

[Sun86]  Sun Microsystems.  Using

nroff and troff on the Sun Workstation. Rev. A. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1986. xvii + 212 pp. LCCN CM.2.5.07. Part number 800-1321-03.

SunSoft:1993:SAD

[Sun93a] SunSoft. Solaris application developer’s guide. SunSoft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-205097-8. xvi + 102 + 6 pp. LCCN QA76.76.O63 S625 1993. ????

SunSoft:1993:SDI

[Sun93b] SunSoft. Solaris desktop integration guide. SunSoft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-035726-X. xviii + 204 pp. LCCN QA76.76.O63 S627 1993. ????

SunSoft:1993:SOD

[Sun93c] SunSoft. Solaris OpenWindows DeskSet reference. SunSoft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-035718-9. xviii + 343 pp. LCCN QA76.76.W56 S64 1993.

SunSoft:1993:SOU

[Sun93d] SunSoft. Solaris OpenWindows user’s guide. SunSoft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-035700-6. xvi + 311 pp. LCCN QA76.9.U83 S65 1993. ????

SunSoft:1995:SPG

[Sun95] SunSoft Developer Engineering. Solaris Porting Guide. SunSoft Press, Englewood Cliffs, NJ, USA, second edition, 1995. ISBN 0-13-443672-5. xlvi + 696 pp. LCCN QA76.76.O63 S64 1995. US$45.00. URL

                                        http://www.sun.com/books/
                                        catalog/sundev1/index.html.

Sun:1996:DIC

[Sun96] Y. Sun. Design and implementation of the CJK unified character set on UNIX system. In Unicode Consortium [Uni96a], pages B6–?? ISBN ???? LCCN ???? A1. Input method design / by Mark Leisher – A2.

Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

SPS:2001:DCB

[Sun01] Sun Professional Services.

Dot-com & beyond: breakthrough Internet-based architectures and methodologies. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2001. ISBN 0-13-062297-4. xvii + 314 pp. LCCN TK5105.5.D68 2001.

Soto:1992:TCC

[SVV92] Antoni Soto, Sebastia Vila, and Alvar Vinacua. A toolkit for constructing command driven graphics programs. Computers and Graphics, 16(4):375– 382, Winter 1992. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Sanford:1982:DCA

[SW82] Curtis Sanford and David Walden. Development of a commercial applications system under UNIX. In Usr Group [Usr82], pages 177–?? Abstract only.

Schaffer:1988:LII

[SW88] Mark A. Schaffer and Geoff Walsh. LOCK/ix: An implementation of UNIX for the LOCK TCB. ;login: the USENIX Association newsletter, 13(3):11–24, May/June 1988. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Shipley:1991:MAL

[SW91a] Carl Shipley and Chingyow Wang. Monitoring activity on a large UNIX network with Perl and Syslogd. In USENIX [USE91b], pages 209–216.

Szabo:1991:DCJ

[SW91b] Bernard I. Szabo and Gregory K. Wallace. Design considerations for JPEG video and synchronized audio in a UNIX workstation environment. In USENIX [USE91c], pages 353– 368. LCCN QA 76.76 O63 U83 1991.

Sauers:1999:HUT

[SW99] Robert F. Sauers and Peter Weygant. HP-UX Tuning and Performance: Concepts, Tools, and Methods. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-102716-6. xxii + 296 pp. LCCN QA76.8.H48S28 1999.

                                        URL http://www.phptr.com/
                                        ptrbooks/ptr_0131027166.
                                        html.

Srinivasan:2012:SSS

[SW12] Avinash Srinivasan and Jie Wu.

SafeCode — safeguarding security and privacy of user data on stolen iOS devices. Lecture Notes in Computer Science, 7672:11–20, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.
springer.com/chapter/10.
1007/978-3-642-35362-8_2/.

Swanson:1974:GFC

[Swa74] G. K. Swanson. GRAFPAC/ FR80 — a C version of deferred GRAFPAC under the UNIX time-sharing system. Technical Memorandum 1018 (TM 74-1353-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 23, 1974. ?? pp.

Swanson:1977:RTP

[Swa77]  G. K. Swanson.  RXD-

11D: Two programs to generate UNIX-compatible magnetic tapes. Technical Memorandum 1215, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 17, 1977. ?? pp.

Swaine:2000:PPe

[Swa00a] Michael Swaine. Programming paradigms: Oh behave! Dr. Dobb’s Journal of Software Tools, 25(5):99–100, 102, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Swan:2000:TSG

[Swa00b]  Tom Swan.  Tom Swan’s

GNU C++ for Linux. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2153-8. xii + 831 pp. LCCN QA76.73.C153 S93 2000. Includes CD-ROM with Mandrake-Linux 6.0, X Class library for C, and the entire source code from the book.

Swaine:2003:PPJ

[Swa03] Michael Swaine. Programming paradigms: Just released. Dr. Dobb’s Journal of Software Tools, 28(6):68, 70–71, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=8213/
ddj0306h/.

Sweet:2001:KD

[Swe01a] David Sweet. KDE 2.0 Development. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN ???? ???? pp. LCCN ????

Sweet:2001:CCU

[Swe01b] Michael R. Sweet. CUPS: Common UNIX Printing System. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-32196-3. xxvi + 622 pp. LCCN QA76.76.O63 S942 2001. US$44.99. URL http:

//www.cups.org;   http://
www.easysw.com/cups/pricing.
html.

Sweetman:2007:SMR

[Swe07] Dominic Sweetman. See MIPS Run. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, second edition, 2007. ISBN 0-12-088421-6. xix + 492 pp. LCCN QA76.9.A73 S88 2007.

Shi:2022:AFL

[SWF^{+}22] Heyuan Shi, Guyu Wang, Ying Fu, Chao Hu, Houbing Song, Jian Dong, Kun Tang, and Kai Liang. Abaci-finder: Linux kernel crash classification through stack trace similarity learning. Journal of Parallel and Distributed Computing, 168(??):70–79, October 2022. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0743731522001320.

Southerton:1992:UTS

[SWG92] Alan Southerton, Andrew D.

Wolfe Jr., and David Granz. DOS and UNIX On a Two-Way Street. UNIX/World, 9(8):48– ??, August 1992. ISSN 0739-5922.

Swift:1977:CUD

[Swi77] R. E. Swift. CM, a UNIX driver for the MINITAPE system. Technical Memorandum 1167 (TM 77-8234-13), AT&T Bell Laboratories, Murray Hill, NJ, USA, June 13, 1977. ?? pp.

Swift:1979:DMP

[Swi79] R. E. Swift. Documentation of minicomputer product trials. Technical Memorandum 1428, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 27, 1979. ?? pp.

Swick:1991:XSF

[Swi91]  Ralph R. Swick.  The X

Window System: fundamentals and applications development: April 15–18, 1991, Engineering 819.209: lecture notes. Los Angeles, University Extension, Dept. of Business, Engineering and Management, Short Course Program, 1991. various pp.

Stonebraker:1976:DII

[SWKH76] Michael Stonebraker, Eugene Wong, Peter Kreps, and Gerald Held. The design and implementation of INGRES. ACM Transactions on Database Systems, 1(3):189–222, September 1976. CODEN ATDSD3. ISSN 0362-5915 (print), 1557-4644 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tods/1976-
1-3/p189-stonebraker/p189-
stonebraker.pdf;   http:/
/www.acm.org/pubs/citations/
journals/tods/1976-1-3/p189-
stonebraker/. Reprinted in

[Sto88]. Also published in/as: UCB, Elec. Res. Lab, Memo No. ERL-M577, Jan. 1976.

Siever:2001:LWS

[SWL^{+}01] Ellen Siever, Matt Welsh, Ben Laurie, Randy Jay Yarger, Alligator Descartes, and Scott Guelich. Linux Web Server CD Bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00208-4. 812 (est.) pp. LCCN ???? US$79.95. URL http:/

/www.oreilly.com/catalog/
linuxwebcdbs.

Seltzer:1991:NHP

[SY91] Margo Seltzer and Ozan Yigit.

A new hashing package for UNIX. In USENIX [USE91d], pages 173–184. LCCN QA 76.76 O63 U84 1992.

Sonnenreich:2000:BLA

[SY00] Wes Sonnenreich and Tom Yates. Building Linux and OpenBSD firewalls. Wiley, New York, NY, USA, 2000. ISBN 0-471-35366-3 (paperback). xxi + 362 pp. LCCN QA76.9.A25 S66 1999. US$44.99. URL http:/

/www.wiley.com/compbooks/
sonnenreich.

Symbolics:1985:IUM

[Sym85] Symbolics, Inc. An introduction to UNIX MACSYMA. Symbolics, Inc., 11 Cambridge Center, Cambridge MA 02142, USA, version 3.0 edition, 1985. 26 pp.

SysAdmin:1997:UPT

[Sys97a] Sys Admin Editors, editor.

UNIX Performance Tuning. R&D Books, Lawrence, KS, USA; Berkeley, CA, USA, 1997. ISBN 0-87930-470-7. viii + 236 pp. LCCN ???? US$29.95. Includes floppy disk.

SysAdmin:1997:US

[Sys97b] Sys Admin Editors, editor.

UNIX Security. R&D Books, Lawrence, KS, USA; Berkeley, CA, USA, 1997. ISBN 0-87930-471-5. viii + 246 pp. LCCN ???? US$29.95.

Smith:1995:IMC

[SZ95] Amy Smith and Betsy Zeller. Indigo Magic: An (un)Common Desktop Environment: The overriding goal of the Indigo Magic Desktop from Silicon Graphics is to provide an integrated, complete work area so that busy professionals can concentrate on using their applications rather than wrestling with UNIX. The X Journal, 4(3):56–??, January/February 1995. CODEN XJOUEA. ISSN 1056-7003.

URL http://www.sigs.com/
publications/docs/txjr/95/
txjr9501s.html.

Taylor:1998:STY

[TA98] Dave Taylor and James C. Armstrong, Jr. Sams teach yourself UNIX in 24 hours. SAMS Publishing, Indianapolis, IN, USA, second edition, 1998. ISBN 0-672-31480-0. ix + 529 pp. LCCN QA76.76.O63 T3888 1998.

Tague:1983:USN

[Tag83] Berkley A. Tague. The UNIX system: New directions. In Association [Ass83], pages 109–?? Abstract only.

Talati:1989:DLT

[Tal89a] Kirit Talati. Distributed online transaction processing on UNIX. In USENIX Association [USE89g], pages 7–12. ISBN ???? LCCN ????

Talati:1989:DOT

[Tal89b] Kirit Talati. Distributed online transaction processing on UNIX. In USENIX Association [USE89g], pages 7–12. ISBN ???? LCCN ????

Talbott:1992:PRM

[Tal92] Steve Talbott. PEXlib Reference Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 1-56592-029-5. xxv + 551 pp. LCCN QA76.76.W56 P4 1992. US$34.95.

Tannenbaum:1984:PHU

[Tan84] Andrew Tannenbaum. Political history of UNIX. In USENIX Association [USE84b], pages 83–88. ISBN none. LCCN QA76.8.U65 U55 1984.

Tanenbaum:1987:MUC

[Tan87a] Andrew S. Tanenbaum. MINIX:

A UNIX clone with source code for the IBM PC. ;login: the USENIX Association newsletter, 12(2):3–9, March/April 1987. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Tanenbaum:1987:UCS

[Tan87b] Andrew S. Tanenbaum. A UNIX clone with source code for operating systems courses. Operating Systems Review, 21 (1):20–29, January 1987. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Tannenbaum:1987:UML

[Tan87c] Andrew Tannenbaum. The UNIX marketplace in 1987: Life, the UNIverse, and everything. In USENIX Association [USE87d], pages 419–424.

Tanenbaum:1988:CN

[Tan88] Andrew S. Tanenbaum. Computer Networks. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1988. ISBN 0-13-162959-X. xv + 658 pp. LCCN TK5105.5 .T36 1988.

Tansley:2000:LUS

[Tan00] David Tansley. Linux and UNIX Shell Programming. Addison-Wesley, Reading, MA, USA, 2000. ISBN 0-201-67472-6. xxiii + 504 pp. LCCN QA76.76.O63 T365 2000. US$44.95. URL http:

                                        //cseng.aw.com/bookpage.
                                        taf?ISBN=0-201-67472-6&ptype=
                                        3017.

Tanenbaum:2001:MOS

[Tan01] Andrew S. Tanenbaum. Modern operating systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-031358-0. xxiv + 951 pp. LCCN QA76.76.O63 T359 2001.

Tanenbaum:2008:MOS

[Tan08] Andrew S. Tanenbaum. Modern Operating Systems. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2008. ISBN 0-13-600663-9. xxvii + 1076 pp. LCCN QA76.76.O63 T359 2008.

Tanenbaum:2010:MOS

[Tan10] Andrew S. Tanenbaum. Modern operating systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, third international edition, 2010. ISBN 0-13-813459-6. xxvii + 1072 pp. LCCN QA76.76.O63 2010.

Tartanella:1980:UTE

[Tar80] C. Tartanella. UNIX Text Editing and Phototypesetting Starter Package. Computing Information Service Libraries and Information Systems Center, Bell Laboratories, Murray Hill, NJ, USA, 1980. ???? pp.

Tare:1989:DPU

[Tar89] Ramkrishna S. Tare. Data processing in the UNIX environment: with INFORMIX-SQL, Embedded-SQL, C-ISAM, and TURBO. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-062885-8. xxi + 438 pp. LCCN QA76.9 .D3 T374 1989.

Taylor:1996:PYU

[Tay96] Martin Taylor. Port your UNIX apps to OS/400 — think it’s a strange idea to port your UNIX apps to the AS/400? Considering the cost is a lot lower, maybe you should be asking yourself why you’re not. Datamation, 42(8):77–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Taylor:2003:STY

[Tay03]  Dave Taylor.  Sams teach

yourself Unix system administration in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 2003. ISBN 0-672-32398-

2. xiii + 508 pp. LCCN QA76.76.O63 T3885 2003. Covers: Unix/Linux/Solaris/Mac OS X.

Taylor:2005:LUM

[Tay05] Dave Taylor. Learning UNIX for Mac OS X Tiger. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-00915-1. xviii + 260 pp. LCCN QA76.76.O63 T388 2005.

Taylor:2006:STY

[Tay06] Dave Taylor. Sams teach yourself Unix in 24 hours. Sams, Indianapolis, Ind., fourth edition, 2006. ISBN 0-672-32814-3. vii + 518 pp. LCCN QA76.76.O63 T3888

                                        2006. URL http://proquest.
                                        safaribooksonline.com/0672328143.

Turk:2005:VLS

[TB05] D. Turk and J. Bausch. Virtual Linux servers under z/VM: Security, performance, and administration issues. IBM Systems Journal, 44(2):341– ??, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/442/turk.html;   http:
                                        //www.research.ibm.com/journal/

sj/442/turk.pdf;   http:/
/www.research.ibm.com/journal/
sj/442/turk.txt.

Tevanian:1987:MTU

[TBG^{+}87] Avadis Tevanian, David Black, David Golub, Richard Rashid, Eric Cooper, and Michael Young. MACH threads and the UNIX kernel: The battle for control. Research paper, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 12 pp.

Tomayko:1989:AWC

[TBGE89] James E. Tomayko, Ralf Bülow, Herbert R. J. Grosch, and John D. Elson. Anecdotes: The Windmill Computer— an eyewitness report of the Scheutz Difference Engine; in Von Braun country; origins of terms [Winchester disk and UNIX]. Annals of the History of Computing, 11(1):43–48, January/March 1989. CODEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1989/pdf/
a1043b.pdf;   http://www.
computer.org/annals/an1989/
a1043babs.htm.

Taudes:1993:UEA

[TBR93] Alfred Taudes, Frank Bensberg, and Jan-Armin Reepmeyer. UNIX: Einstieg für DOS-Anwender. (German) [UNIX: Introduction for DOS users]. Physica-Verlag, Vienna, Austria, 1993. ISBN 3-7908-0700-1. xii + 232 pp. LCCN ????

Tolchin:1987:RWU

[TBS87] Stephen Tolchin, Eric Bergan, and Marvin Schneider. Real world UNIX DBMS applications: Experiences and observations. In USENIX Association [USE87e], pages 210– 222. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

Torrey:2007:CIL

[TCM07] Lisa A. Torrey, Joyce Coleman, and Barton P. Miller. A comparison of interactivity in the Linux 2.6 scheduler and an MLFQ scheduler. Software— Practice and Experience, 37(4): 347–364, April 10, 2007. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Todino:1991:UUU

[TD91] Grace Todino and Dale Dougherty. Using UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1991. ISBN 0-937175-10-2. xv + 194 pp. LCCN QA76.76 O63 T63 1991. US$21.95. Well written on how to use uucp and Netnews.

Torvalds:2001:JFS

[TD01] Linus Torvalds and David Diamond. Just for Fun: the Story of an Accidental Revolutionary. HarperBusiness, New York, NY, USA, 2001. ISBN 0-06-662072-4 (hardcover), 0-06-662073-2, 1-58799-151-9. xxii + 262 pp. LCCN QA76.2.T67 T67 2001. URL http:

//catdir.loc.gov/catdir/
description/hc043/00054199.
html.

Ts:2003:USF

[TECB03] Jay Ts, Robert Eckstein, and David Collier-Brown. Using Samba: a File & Print Server for Linux, Unix & Mac OS X. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00256-4. xiii + 539 pp. LCCN QA76.9.C55 E38 2003. US$39.95 ; EUR 44.00; SFr 72.90. URL http:/

/www.oreilly.com/catalog/
9780596002565;   http://
www.oreilly.com/catalog/
samba2/.

Teer:2005:SSP

[Tee05] Rich Teer. Solaris systems programming. Addison-Wes-ley, Reading, MA, USA, 2005. ISBN 0-201-75039-2. xxxvi + 1211 pp. LCCN QA76.76.O63 T4314 2005.

Teodorovici:2013:BRA

[Teo13] Vasile G. Teodorovici. Book review: Advanced programming in the UNIX environment, third edition by W. Richard Stevens and Stephen A. Rago. ACM SIGSOFT Software Engineering Notes, 38 (6):45, November 2013. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Test:1986:MMC

[Tes86] Jack A. Test. Multi-processor management in the Concentrix operating system. In USENIX Association [USE86e], pages 172–182. URL https://

                                        bitsavers.org/pdf/usenix/
                                        USENIX_1986_Winter_Technical_
                                        Conference_Proceedings_198601.
                                        pdf.

Tevanian:1987:MBF

[Tev87] Avadis Tevanian. MACH: a basis for future UNIX development. Research paper CMU-CS-87-139, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 5 pp.

Tsegaye:2004:CLW

[TF04] Melekam Tsegaye and Richard Foss. A comparison of the Linux and Windows device driver architectures. Operating Systems Review, 38(2):8– 33, April 2004. CODEN OS-RED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Ting:1978:PNI

[TG78] D. W. Ting and A. L. Glasser. PORT — a new inter-process communication facility under UNIX. Technical Memorandum 1230 (TM 78-3124-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 3, 1978. ?? pp.

Ting:1979:PLF

[TG79] D. W. Ting and A. L. Glasser.

A public library facility for the UNIX time-sharing system. Technical Memorandum 1363 (TM 79-3111-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 21, 1979. ?? pp.

Teixeira:1988:SUG

[TG88] Thomas J. Teixeira and Robert F.

Gurwitz. Stellix: UNIX for a graphics supercomputer. In Association [Ass88d], pages 321– 330.

Tondo:1989:CAB

[TG89] Clovis L. Tondo and Scott E.

Gimpel. The C Answer Book — Solutions to the Exercises in The C Programming Language. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1989. ISBN 0-13-109653-2. 208 pp. LCCN QA76.73 C15 T66 1989. This book contains answers to the problems in [KR88].

Tackett:1996:SEU

[TG96] Jack Tackett, Jr. and David Gunter. Special edition using Linux. Que Corp., Indianapolis, IN, USA, second edition, 1996. ISBN 0-7897-0742-X. xxxvii + 792 pp. LCCN QA76.76.O63T28 1996.

Tackett:1995:SEU

[TGB95] Jack Tackett, David Gunter, and Lance Brown. Special Edition Using Linux. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0100-6. xxvi + 861 pp. LCCN QA76.76.O63T28 1995. US$49.95.

Thakkar:1988:BMS

[TGF88] Shreekant Thakkar, Paul Gifford, and Garay Fielland. The Balance multiprocessor system. IEEE Micro, 8(1):57–69, January/February 1988. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tague:1977:CLP

[TH77] B. A. Tague and R. C. Haight. C language portability. Technical Memorandum 1204, AT&T Bell Laboratories, Murray Hill, NJ, USA, September 22, 1977. ?? pp.

Tanenbaum:1983:TPMa

[TH83a] Andrew S. Tanenbaum and Teus Hagen. Two programs, many UNIX systems. EUUG Newsletter, 3(1):12–13, Spring 1983. CODEN EONLE8. ISSN 1011-4211.

Tanenbaum:1983:TPMb

[TH83b] Andrew S. Tanenbaum and Teus Hagen. Two programs, many UNIX systems (reprint). EUUG Newsletter, 3(4):24– 26, Winter 1983. CODEN EONLE8. ISSN 1011-4211.

Thomas:2004:PVC

[TH04] David Thomas and Andrew Hunt. Pragmatic version control: using CVS, volume 1 of Pragmatic starter kit series. Pragmatic Bookshelf, Raleigh, NC, USA, 2004. ISBN 0-9745140-0-4. xiv + 161 pp. LCCN QA76.76.D47 T46 2004.

Tanenbaum:2006:FSD

[THB06] Andrew S. Tanenbaum, Jorrit N. Herder, and Herbert Bos. File size distribution on UNIX systems: then and now. Operating Systems Review, 40(1): 100–104, January 2006. CO-DEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

OReilly:1991:GO

[The91] The Staff of O’Reilly and Associates. Guide to OSF/1. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, June 1991. ISBN 0-937175-78-1. 304 pp. LCCN QA76.755.G85 1991. US$21.95.

MathWorks:1992:MHPb

[The92a] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: External Interface Guide: for UNIX workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? LCCN QA188 .M384 1993.

MathWorks:1992:MHPd

[The92b] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: External Interface Guide: for UNIX workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? LCCN QA188 .M384 1993.

MathWorks:1992:MHPa

[The92c] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: User’s Guide: for UNIX workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1992. ISBN ???? LCCN QA188 .M389 1992.

MathWorks:1993:MHPc

[The93] The MathWorks, Inc. MAT-LAB, High-performance Numeric Computation and Visualization Software: Release Notes, Version 4.1: for UNIX Workstations. The Mathworks, Cochituate Place, 24 Prime Park Way, Natick, MA, USA, 1993. ISBN ???? 59 pp. LCCN QA188 .M388 1993.

OG:2002:SUS

[The02] The Open Group. The Single UNIX Specification: The Authorized Guide to Version 3. The Open Group, Publications Department, Apex Plaza, Forbury Road, Reading, Berkshire RG1 1AX, UK, 2002. ISBN 1-85912-277-9 (UK), 1-931624-13-5 (US). ???? pp. LCCN ???? US$59, UK£36. URL

http://www.unix-systems.
org/version3/theguide.html.

Open Group Document Number G906.

Thirion:1992:EUC

[Thi92] Bernard Thirion. Edison, a Unix and C friendly Rete based production system. ACM SIG-PLAN Notices, 27(1):75–84, January 1992. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thiem:1999:KAD

[Thi99] Uwe Thiem. KDE application development. New Riders Publishing, Carmel, IN, USA, 1999. ISBN 1-57870-201-1. xiii + 189 pp. LCCN QA76.76.A65 T4713 1999. URL http:/

/www.newriders.com/books/
title.cfm?isbn=1578702011.

Thiruvathukal:2004:GLN

[Thi04] George K. Thiruvathukal. Gentoo Linux: The next generation of Linux. Computing in Science and Engineering, 6 (5):66–74, September/October 2004. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL

http://csdl.computer.org/
dl/mags/cs/2004/05/c5066.
htm;   http://csdl.computer.
org/dl/mags/cs/2004/05/c5066.
pdf.

Thompson:1968:PTR

[Tho68] Ken Thompson. Programming techniques: Regular expression search algorithm. Communications of the Association for Computing Machinery, 11(6): 419–422, June 1968. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://patft.

uspto.gov/. See also [KP99, Cox07, Cox09, Cox10, Cox12].

Thompson:1969:BUM

[Tho69] Kenneth L. Thompson. Bon user’s manual. Report, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1, 1969. As told to M. D. McIlroy and R. Morris. Bon is a step between BCPL and B, after which came New B and then C.

Thompson:1972:URB

[Tho72] Ken Thompson. Users’ reference to B. Technical Memorandum MM-72-1271-1, Bell Laboratories, Murray Hill, NJ, USA, January 7, 1972. iv + 27 pp.

                                        URL http://www.bell-labs.
                                        com/usr/dmr/www/kbman.pdf.

Thompson:1976:UCL

[Tho76] Ken Thompson. The UNIX command language. In Structured Programming (Infotech state of the art report), pages 377–384. Infotech Limited, Maidenhead, UK, 1976. ISBN 0-85539-270-3. LCCN QA76.6.

                                        URL https://archive.org/
                                        details/theunixcommandlanguage;
                                         https://github.com/susam/
                                        tucl.

Thompson:1978:UTS

[Tho78] Ken Thompson. UNIX timesharing system: UNIX implementation. The Bell System Technical Journal, 57(6): 1931–1946, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-1931.
pdf.

Thompson:1979:GPA

[Tho79] J. S. Thompson. A general purpose assembler. Technical Memorandum 1357 (TM 79-1365-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 12, 1979. ?? pp.

Thomas:1985:ACA

[Tho85a] Spencer W. Thomas. The Alpha 1 computer-aided geometric design system in the UNIX environment. ;login: the USENIX Association newsletter, 10(4):54–64, October/ November 1985. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Thomas:1985:ACG

[Tho85b] Spencer W. Thomas. The Alpha 1 computer-aided geometric design system in the UNIX environment. ;login: the USENIX Association newsletter, 10(4):54–64, October/ November 1985. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Thomborson:1992:VSD

[Tho92]  Clark Thomborson.  The

V.42bis standard for datacompressing modems. IEEE Micro, 12(5):41–53, September/October 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Thomas:1995:BFI

[Tho95]  C. G. Thomas.  BASAR:

a framework for integrating agents in the World Wide Web. Computer, 28(5):84–86, May 1995. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thompson:19xx:NCC

[Thoxx] Ken Thompson. A new C compiler. Computing Science Technical Report 158h, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? ??, 19xx.

                                        12 pp. URL http://doc.
                                        cat-v.org/bell_labs/new_
                                        c_compilers/new_c_compiler.
                                        pdf.

Thomas:2003:IMF

[Tho03] James W. Thomas. Inlining of mathematical functions in HP-UX for Itanium 2. In IEEE [IEE03], pages 135–144. ISBN 0-7695-1913-X. LCCN ???? ACM Order No. 530033.

Thomas:2005:RBC

[Tho05] Marlin Thomas. Review of Building Clustered Linux Systems by Robert W. Lucke, Prentice Hall PTR, 2004, $49.99, ISBN: 0-13-144853-6. ACM Queue: Tomorrow’s Computing Today, 3 (4):60, May 2005. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Thomas:2006:BUL

[Tho06a]  Keir Thomas.  Beginning

Ubuntu Linux: from novice to professional. Apress, Berkeley, CA, USA, 2006. ISBN 1-59059-627-7. xxxii + 573 pp. LCCN QA76.76.O63 T565 2006.

Thompson:2006:QGM

[Tho06b] R. Scott Thompson. Quartz 2D graphics for Mac OS X developers. Addison-Wesley, Reading, MA, USA, 2006. ISBN 0-321-33663-1 (paperback). xvi + 330 pp. LCCN T385 .T4953995 2006. URL http:

//www.loc.gov/catdir/toc/
ecip064/2005034911.html.

Thornton:2019:UHH

[Tho19a] Adam Thornton. The UNIX-HATERS Handbook, a review. Web site., May 16, 2019.

URL  https://athornton.
dreamwidth.org/14272.html.

Thornton:2019:EOU

[Tho19b]  Adam Thornton.  Explor-

ing old Unixes. Web site., June 2, 2019. URL https://

athornton.dreamwidth.org/
14340.html.

Tilson:1984:TUS

[Til84] Michael Tilson. Towards a UNIX standard. In Software Tools Users Group [Sof84], pages 1–10. LCCN QA76.8.U65 U83 1984.

Tilson:1987:UAT

[Til87a] Michael Tilson. UNIX at the turn of the century. In USENIX Association [USE87d], pages 425–435.

Tilson:1987:UTC

[Til87b] Michael Tilson. UNIX at the turn of the century. In USENIX Association [USE87d], pages 425–435.

Till:1995:TYP

[Til95] Dave Till. Teach yourself Perl in 21 days. Howard W. Sams, Indianapolis, IN 46268, USA, 1995. ISBN 0-672-30586-0. xxxiv + 841 pp. LCCN QA76.73.P22 T55 1995. US$29.99, CDN$39.99.

Timar:1993:FAQ

[Tim93] Ted Timar. The Frequently Asked Questions List, March 18 1993. This is a multi-part list of often-asked UNIX questions (with answers). Read the

                                        postings on news.answers or

get them by anonymous ftp

                                        from rtfm.mit.edu (in /pub/
                                        usenet/news.answers/unix-
                                        faq).

Ting:1978:AFR

[Tin78a] D. W. Ting. Asynchronous I/O facilities for raw block devices under UNIX. Technical Memorandum 1241 (TM 78-3124-9), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 1, 1978. ?? pp.

Ting:1978:IPC

[Tin78b] D. W. Ting. Inter-process communications facility for transaction oriented UNIX systems. Technical Memorandum 1303 (TM 78-3111-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 14, 1978. ?? pp.

Ting:1978:MSB

[Tin78c] D. W. Ting. Moving system I/O buffers of UNIX data address space. Technical Memorandum 1235 (TM 78-3124-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 2, 1978. ?? pp.

Ting:1978:OPC

[Tin78d] D. W. Ting. Optimizing process creation and termination mechanisms in UNIX. Technical Memorandum 1236 (TM 78-3124-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, February 4, 1978. ?? pp.

Ting:1979:DBM

[Tin79a] A. C. Ting. A data-base management system under MERT. Technical Memorandum 1374 (TM 79-4145-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 4, 1979. ?? pp.

Ting:1979:USD

[Tin79b] P. D. Ting. UNITY — a small DBMS for a UNIX environment. Technical Memorandum 1438 (TM 79-1712-12), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 4, 1979. ?? pp.

Tipton:1992:RPC

[Tip92] Bob Tipton. RS/6000: POSIXtively confusing. News 3X/400, ????(????):177–180, June 1992. ISSN 1040-6093.

Taylor:2003:LUM

[TJ03] Dave Taylor and Brian Jepson. Learning Unix for Mac OS X. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2003. ISBN 0-596-00470-2. xiv + 141 pp. LCCN QA76.76.O63 T388 2003. Covers Mac OS X Jaguar.

Taylor:2004:LUM

[TJ04] Dave Taylor and Brian Jepson. Learning UNIX for Mac OS X Panther. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, third edition, 2004. ISBN 0-596-00617-9. xiii + 168 pp. LCCN QA76.76.O63 T3884 2004. URL http:/

/www.oreilly.com/catalog/
9780596006174.

Tjiang:1985:TRM

[Tji85] Steven W. K. Tjiang. Twig reference manual. Computing Science Technical Report 120, AT&T Bell Laboratories, Murray Hill, NJ, USA, ???? 1985. ii + 19 pp.

Tang:1999:MBM

[TLK99] Ty Tang, Vipul Lal, and Shesha Krishnapura. MTEX — a bridge for migrating CAD design environment from UNIX to NT. In USENIX [USE99a], page ?? ISBN 1-880446-29-4. LCCN ????

URL  http://www.usenix.
org/publications/library/
proceedings/usenix-nt99/
tang.html.

Tomayko:1989:ACI

[TLR89] James E. Tomayko, Lawrence W.

Langley, and Juris Reinfelds. Anecdotes: a critical incident; the first port of UNIX. Annals of the History of Computing, 11(3):207–210, July/ September 1989. CODEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1989/pdf/
a3207.pdf;   http://www.
computer.org/annals/an1989/
a3207abs.htm.

Tondo:1992:MMG

[TNY92a] Clovis L. Tondo, Andrew Nathanson, and Eden Yount. Mastering MAKE — a Guide to Building Programs on DOS and UNIX Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-554619-2. x + 143 pp. LCCN QA76.76.U84 T66 1992.

Tondo:MM92

[TNY92b] Clovis L. Tondo, Andrew Nathanson, and Eden Yount. Mastering MAKE — a Guide to Building Programs on DOS and UNIX Systems. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-554619-2. x + 143 pp. LCCN QA76.76.U84 T66 1992. Prepared with L^{A}TEX.

Tondo:1994:MMG

[TNY94a] Clovis L. Tondo, Andrew Nathanson, and Eden Yount. Mastering Make — a Guide to Building Programs on DOS, OS/2, and UNIX Systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1994. ISBN 0-13-121906-5. xvi + 310 pp. LCCN QA76.76.U84T65 1994. US$22.00.

Tondo:MM94

[TNY94b] Clovis L. Tondo, Andrew Nathanson, and Eden Yount. Mastering Make — a Guide to Building Programs on DOS, OS/2, and UNIX Systems. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1994. ISBN 0-13-121906-5. xvi + 310 pp. LCCN QA76.76.U84T65 1994. US$22.00. Prepared with L^{A}TEX.

Todino:1992:MUU

[TO92] Grace Todino and Tim O’Reilly. Managing UUCP and Usenet. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, tenth edition, January 1992. ISBN 0-937175-93-5. 368 pp. LCCN QA76.76 O63 T63. US$27.95. Well written on the setting-up and the maintenance of UUCP and Netnews.

Tobias:1980:LDP

[Tob80] Jeffrey M. Tobias, editor. Language design and programming methodology: proceedings of a symposium held at Sydney, Australia, 10–11 September 1979, volume 79 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1980. CODEN LNCSD9. ISBN 0-387-09745-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.7 .S94 1979. Sponsored by the Australian Atomic Energy Commission and the University of New South Wales.

Tomson:2002:DCR

[Tom02] Phil Tomson. Distributed computing with Ruby. Dr. Dobb’s Journal of Software Tools, 27 (9):16, 18, 22, September 2002. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_09/
taskmas.txt;   http://www.
ddj.com/ftp/2002/2002_09/
taskmas.zip.

Tomaszewska:2009:I

[Tom09] Aleksandra Tomaszewska. Inkscape.

Wydawnictwo Helion, Gliwice, Poland, 2009. ISBN 83-246-1902-X. 135 pp. LCCN ???? US$.

Tommasino:2011:HGR

[Tom11] Damian Tommasino. Handson guide to the Red Hat exams: RHCSA and RHCE cert guide and lab manual. Pearson, Indianapolis, IN, USA, 2011. ISBN 0-321-76795-0. ???? pp. LCCN QA76.76.O63 T6494 2011.

Thomas:2004:LLF

[TOML04] James W. Thomas, Jon P. Okada, Peter Markstein, and Ren-Cang Li. The Libm library and floating-point arithmetic in HP-UX for Itaniumbased systems: Updated for HP-UX 11i v2. Technical report, Hewlett–Packard Corporation, Palo Alto, CA, USA, December 3, 2004. 26

pp. URL http://h21007.
www2.hp.com/dspp/ddl/ddl_
Download_File_TRX/1,1249,
942,00.pdf.

Toomey:2009:REU

[Too09] Warren Toomey. The restoration of early UNIX artifacts. In ????, editor, Proceedings of the 2009 USENIX Annual Technical Conference, June 14– 19, San Diego, CA, USA, page ?? USENIX Association, Berkeley, CA, USA, 2009. ISBN ???? LCCN ????

URL  http://www.usenix.
org/events/usenix09/tech/
full_papers/toomey/toomey.
pdf;   http://www.usenix.
org/events/usenix09/tech/
slides/toomey.pdf;   http:
//www.usenix.org/multimedia/
atc09toomey.

Toomey:2010:FEU

[Too10] Warren Toomey. First Edition Unix: Its creation and restoration. IEEE Annals of the History of Computing, 32(3):74– 82, July/September 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Toomey:2011:SBL

[Too11] Warren Toomey. The strange birth and long life of Unix: The classic operating system turns 40, and its progeny abound. IEEE Spectrum, 48(12):34– 55, December 2011. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

URL http://spectrum.ieee.
org/computing/software/the-
strange-birth-and-long-life-
of-unix.

Toomey:2017:UBD

[Too17] Warren Toomey. Unix: Building a development environment from scratch. Report, The Unix Heritage Society & TAFE Queensland, ????, June 22, 2017. 17 pp. URL http:

//minnie.tuhs.org/Y5/wkt_
hapop_paper.pdf.

Topham:1990:PU

[Top90a] Douglas Topham. Portable UNIX. Wiley-Interscience, New York, NY, USA, 1990. ISBN 0-471-57926-2. xiii + 274 pp. LCCN QA76.76.O63 T666 1992. A pretty good taskoriented quick reference.

Topham:1990:SVG

[Top90b] Douglas W. Topham. A System V Guide to Unix and Xenix.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-97021-5. xxii + 733 pp. LCCN QA76.76.O63 T667 1990. US$50.00.

Toporek:2002:MXPa

[Top02a] Chuck Toporek. Mac OS X pocket guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2002. ISBN 0-596-00458-3. vii + 141 pp. LCCN QA76.76.O63 T66 2002. Covers Jaguar.

Toporek:2002:MXPb

[Top02b] Chuck Toporek. Mac OS X pocket reference. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00346-3. vii + 115 pp. LCCN QA76.6.O63 P634A 2002.

Toporek:2005:MXT

[Top05] Chuck Toporek. Mac OS X Tiger: pocket guide. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, fourth edition, 2005. ISBN 0-596-00914-3 (paperback). xi + 223 pp. LCCN QA76.76.O63 T668 2005.

Torvalds:1999:LE

[Tor99] Linus Torvalds. The Linux edge. Communications of the Association for Computing Machinery, 42(4):38– 39, April 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
cacm/1999-42-4/p38-torvalds/

.

Toth:1999:LNS

[Tot99] Viktor Toth. Linux: a network solution for your office. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31628-5. xxii + 512 pp. LCCN QA76.76.O63 T6778 1999. Edited by Grace Buechlein and Laura Bulcher.

Toxen:2007:SDS

[Tox07] Bob Toxen. The seven deadly sins of Linux security. ACM Queue: Tomorrow’s Computing Today, 5(4):38–47, May– June 2007. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Terek:1991:EAC

[TP91] Robert Terek and Joseph Pasquale. Experiences with audio conferencing using the X Window System, UNIX and TCP/IP. In USENIX [USE91c], pages 405–418. LCCN QA 76.76 O63 U83 1991.

Todino:1998:LUO

[TPS98] Grace Todino, Jerry D. Peek, and John Strang. Learning the UNIX operating system. Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 1998. ISBN 1-56592-390-1. xi + 92 pp. LCCN QA76.76.O63 T62 1998. Additions and revisions by Jerry Peek.

Taylor:2002:LUM

[TPTS02] Dave Taylor, Jerry D. Peek, Grace Todino, and John Strang. Learning Unix for Mac OS X. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00342-0. xiv + 139 pp. LCCN QA76.76.O63 T388 2002 Stacks. US$19.95. URL http:/

                                        /safari.oreilly.com/0596003420;
                                         http://www.oreilly.com/
                                        catalog/lunixmacosx.

Thompson:1971:UPM

[TR71] Ken Thompson and Dennis M. Ritchie. Unix programmer’s manual. Web site, November 3, 1971. URL https:

                                        //www.bell-labs.com/usr/
                                        dmr/www/1stEdman.html. This

is the first of seven editions (see [TRS18]). Lucent Technologies retains copyright on this material.

Thompson:1972:UPM

[TR72] Ken Thompson and Dennis M. Ritchie. Unix programmer’s manual. Web site, June 12,

                                        1972. URL http://www.tuhs.
                                        org/Archive/Distributions/
                                        Research/Dennis_v2/v2man.

pdf. This is the second of seven

editions (see [TRS18]).

Thompson:1973:UPMa

[TR73a] Ken Thompson and Dennis M. Ritchie. Unix Programmer’s Manual. Bell Telephone Laboratories, Inc., Murray Hill, NJ, USA, third edition, February 1973. xxxii + 260 pp.

URL https://dspinellis.
github.io/unix-v3man/v3man.

pdf. This is the third of seven editions (see [TRS18]). The document is formatted in a monospaced typewriter font.

Thompson:1973:UPMb

[TR73b] Ken Thompson and Dennis M.

Ritchie. Unix Programmer’s Manual. Bell Telephone Laboratories, Inc., Murray Hill, NJ, USA, fourth edition, November 1973. 27 + 255 pp.

URL https://dspinellis.
github.io/unix-v4man/v4man.

pdf. This is the fourth of seven editions (see [TRS18]). The document is typeset.

Thompson:1974:UPM

[TR74] Ken Thompson and Dennis M.

Ritchie. Unix programmer’s manual. Web site, June 1974.

URL http://www.tuhs.org/
Archive/Distributions/Research/
Dennis_v5/v5man.pdf. This is

the fifth of seven editions (see [TRS18]).

Thompson:1975:UPMa

[TR75a] Ken Thompson and Dennis M.

Ritchie. Unix programmer’s manual. Web site, May 1975.

URL https://ia800600.us.
archive.org/19/items/v6-
manual/v6-manual.pdf;   https:
//minnie.tuhs.org//cgi-bin/
utree.pl?file=V6/usr/doc/

unix. This is the sixth of seven editions (see [TRS18]).

Thompson:1975:UPMb

[TR75b] Ken Thompson and Dennis M. Ritchie. UNIX programmer’s manual. Technical Memorandum 1014, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 1975. ?? pp.

Thomas:1994:EUK

[TR94] Philip K. Thomas and Shmuel Rotenstreich. Enhancing the UNIX Korn shell using predictor techniques. C Users Journal, 12(3):83–??, March 1994. ISSN 0898-9788.

Taber:1999:MLS

[TR99] Mark Taber and Randi Roger. Maximum Linux security: a hacker’s guide to protecting your Linux server and network. Macmillan Computer Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31670-6. xvii + 743 pp. LCCN QA76.9.A25 M387 2000.

Trammell:1985:CBH

[Tra85] R. D. Trammell. A capability based hierarchic architecture for UNIX window management. In USENIX Association [USE85d], pages 373–379. LCCN QA76.8.U65 U8 1985.

Tranter:1996:LMG

[Tra96] Jeff Tranter. Linux Multimedia Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, Fall 1996.

ISBN 1-56592-219-0. xx + 363 pp. LCCN QA76.575 .T73 1996. US$32.95. URL http:/

/www.oreilly.com/catalog/
9781565922198;   http://
www.oreilly.com/catalog/
multilinux.

Treese:1988:BUW

[Tre88] G. Winfield Treese. Berkeley UNIX on 1000 workstations: Athena changes to 4.3BSD. In USENIX Association [USE88g], pages 175–182. ISBN ???? LCCN ????

Trickey:1990:LUG

[Tri90] H. Trickey. L^{A}TEX user guide.

In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Thompson:2018:TRE

[TRS18] Ken Thompson, Dennis M.

Ritchie, and Diomidis Spinellis. The Third Research Edition Unix programmer’s manual. Web site, 2018.

URL https://github.com/
dspinellis/unix-v3man. The

site contains links to online manuals for the first through seventh editions.

Truscott:1983:EU

[Tru83] Tom Truscott. An enhanced Uucp. Technical Memorandum CDSR005, Research Triangle Institute, Research Triangle Park, NC 27709-2194, USA, December 1983. ?? pp.

Tilbrook:1987:CUS

[TS87] David Tilbrook and Zalman Stern. Cleaning up UNIX source or bringing discipline to anarchy. In USENIX Association [USE87a], pages 275–286. ISBN ???? LCCN ????

Tuthill:1997:SID

[TS97] Bill Tuthill and David Smallberg. Solaris International Developer’s Guide. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, second edition, 1997. ISBN 0-13-494493-3. xxiii + 382 pp. LCCN QA76.76.O63T888 1997. URL

                                        http://www.sun.com/books/
                                        catalog/tuthill/index.html.

Thiruvathukal:2002:WPT

[TSC02] George K. (George Kuriakose) Thiruvathukal, John P. Shafaee, and Thomas W. Christopher. Web programming: techniques for integrating Python, Linux, Apache, and MySQL. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2002. ISBN 0-13-041065-9. xviii + 745 pp. LCCN QA76.625 .T48 2002.

                                        URL http://www.phptr.com/

ptrbooks/ptr_0130410659.
html.

Toporek:2006:MXT

[TSM^{+}06] Chuck Toporek, Chris Stone, Jason McIntosh, et al. Mac OS X Tiger in a nutshell. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 2006. ISBN 0-596-00943-7 (paperback). xv + 498 pp. LCCN QA76.76.O63 M3873 2006.

Thomas:1996:UHA

[TSO^{+}96] Gregory S. Thomas, James O.

Schroeder, Merrilee E. Orcutt, Desiree C. Johnson, Jeffrey T. Simmelink, and John P. Moore. UNIX host administration in a heterogeneous distributed computing environment. In USENIX [USE96a], pages 43– 50. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

URL  http://www.usenix.
org/publications/library/
proceedings/lisa96/gthomas.
html.

Todino:1993:LUO

[TSP93a] Grace Todino, John Strang, and Jerry Peek. Learning the UNIX Operating System. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1993. ISBN 1-56592-060-0. 108 pp. US$9.95.

Todino:LUS93

[TSP93b] Grace Todino, John Strang, and Jerry Peek. Learning the UNIX Operating System.

O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, August 1993. ISBN 1-56592-060-0. 108 pp. US$9.95.

Todino:1994:LUO

[TSP94] Grace Todino, John Strang, and Jerry D. Peek. Learning the UNIX Operating System. A Nutshell handbook. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1994. ISBN 1-56592-060-0. xv + 92 pp. LCCN QA76.76.O63 T62 1994. US$9.95. URL http:/

/www.oreilly.com/catalog/
9781565920606. Minor correc-

tions.

Ting:1979:MDD

[TT79] P. D. Ting and D. C. Tsichritzis. A micro-DBMS for a distributed data base. Technical Memorandum 1350 (TM 79-1712-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 29, 1979. ?? pp.

Toral:2009:MML

[TTB09] S. L. Toral, R. Martínez Torres, and F. Barrero. Modelling mailing list behaviour in open source projects: the case of ARM Embedded Linux. J.UCS: Journal of Universal Computer Science, 15(3): 648–??, ???? 2009. CO-DEN ???? ISSN 0948-

6968. URL http://www.jucs.
org/jucs_15_3/modelling_
mailing_list_behaviour.

Tevanian:1987:UIS

[TTG^{+}87] Avadis Tevanian, Jr., Mary Thompson, David Golub, Richard Rashid, William Bolosky, Michael Young, and Richard Sanzi. A UNIX interface for shared memory and memory mapped files under MACH. Research paper, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. 14 pp. URL http:

//cs.cmu.edu/afs/cs.cmu.
edu/project/mach/public/
doc/published/mapfiles87.
ps.

Tiemann:2002:FU

[TU02] Brian Tiemann and Michael Urban. FreeBSD Unleashed. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32206-4. xxiv + 992 pp. LCCN QA76.754 .T54 2002. US$49.99.

Tiemann:2003:FU

[TU03] Brian Tiemann and Michael C.

Urban. FreeBSD Unleashed. SAMS Publishing, Indianapolis, IN, USA, second edition, 2003. ISBN 0-672-32456-3. xxvii + 974 pp. LCCN QA76.754 .T54 2003.

Tiemann:2006:FU

[TU06] Brian Tiemann and Michael C.

Urban. FreeBSD 6 unleashed. SAMS Publishing, Indianapolis, IN, USA, 2006. ISBN 0-672-32875-5. xxv + 877 pp. LCCN QA76.754 T53 2006.

Tucker:1990:PPV

[Tuc90] Michael Jay Tucker. Paradoxically Posix: Vendors catering to the federal market are eyeing Posix as the key to winning future contracts. UNIX/World, 7 (3):85–??, March 1, 1990. ISSN 0739-5922.

Tucker:1998:SBN

[Tuc98] Michael Jay Tucker. Strange bedfellows: NT, UNIX, and the integration option. Datamation, 44(2):42–??, ???? 1998. CODEN DTMNAT. ISSN 0011-6963.

Tung:2005:FMF

[Tun05] Michael M. Tung. FORM matters: Fast symbolic computation under UNIX. Computers and Mathematics with Applications, 49(7–8):1127– 1137, May/June 2005. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.com/
                                        science/article/pii/S0898122105001276.

Tuori:1982:UBT

[Tuo82] Martin Tuori. A UNIX benchmarking tool with results from the PDP-11/44, VAX 11/780, and Perkin-Elmer 3242. In Usr Group [Usr82], pages 237–246.

Tuori:1983:TUS

[Tuo83] Martin Tuori. Talking to UNIX — some experience with speech input. In Software Tools Users Group [Sof83], pages 179–185. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Turner:1986:OM

[Tur86] D. Turner. An overview of Miranda. ACM SIGPLAN Notices, 21(12):158–166, December 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turner:1987:MAF

[Tur87] David Turner. Miranda — an advanced functional programming system running under UNIX. In USENIX Association [USE87d], pages 459–469.

Turnbull:2005:HL

[Tur05] James Turnbull. Hardening Linux. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-444-4 (paperback). xxvii + 552 pp. LCCN QA76.76.D49 T87

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0663/2005280402-d.html.

Tuthill:1983:TUS

[Tut83] B. Tuthill. Typesetting on the UNIX system. BYTE Magazine, 8(19):253–265, October 1983. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Tuthill:1993:SID

[Tut93] Bill Tuthill. Solaris international developer’s guide. Sun-Soft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 0-13-031063-8. xxiv + 216 pp. LCCN QA76.76.O63 T89 1993. ????

Terpstra:2004:OSH

[TV04] John H. Terpstra and Jelmer R. Vernooij, editors. The Official Samba-3 HOWTO and Reference Guide. Bruce Perens Open Source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2004. ISBN 0-13-145355-6. xxxviii + 685 pp. LCCN QA76.76.O63 O34345 2004. US$49.99, 50.99 EUR. URL

http://vig.prenhall.com/
catalog/academic/product/
0,4096,0131453556,00.html.

Tanenbaum:1983:UTK

[TvK83] Andrew S. Tanenbaum, Hans van Staveren, and E. G. Keizer. A UNIX tool kit for making portable compilers. In Software Tools Users Group [Sof83], pages 255–261. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Tanenbaum:1997:OSD

[TW97] Andrew S. Tanenbaum and Albert S. Woodhull. Operating Systems — Design and Implementation. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-638677-6. xvii + 939 pp. LCCN QA76.76.O63T36 1997. US$62.00. Includes CD-ROM.

Tanenbaum:2006:OSD

[TW06] Andrew S. Tanenbaum and Albert S. Woodhull. Operating systems: design and implementation. Pearson Prentice Hall, Upper Saddle River, NJ 07458, USA, third edition, 2006. ISBN 0-13-142938-8. xvii + 1054 pp. LCCN QA76.76.O63 T36 2006.

Truscott:1986:SUD

[TWM86a] Tom Truscott, Bob Warren, and Kent Moat. A statewide UNIX distributed computing system. In USENIX Association [USE86c], pages 499– 513.

Truscott:1986:SWU

[TWM86b] Tom Truscott, Bob Warren, and Kent Moat. A state-wide UNIX distributed computing system. In USENIX [USE86a], pages 499–513.

Thomas:1982:CU

[TY82a] Rebecca Thomas and Jean Yates. The commercialization of UNIX. In Usr Group [Usr82], pages 281–282. Abstract only.

Thomas:1982:UGU

[TY82b] Rebecca Thomas and Jean Yates. A User Guide to the UNIX System. Osborne/Mc-Graw-Hill, Berkeley, CA, USA, 1982. ISBN 0-931988-71-3. xi + 508 pp. LCCN QA76.8.U65 T45 1982.

Tyler:2006:FL

[Tyl06] Chris Tyler. Fedora Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-52682-2. xiii + 639 pp. LCCN QA76.76.O63; QA76.76.O63 T9 2006eb. URL http:/

/www.oreilly.com/catalog/
9780596526825.

Ubell:1976:FSC

[Ube76] Michael Ubell. FORTED, a syntax checking Fortran editor for the Princeton UNIX Fortran system: research project. Thesis (M.S. in electrical engineering), University of California, Berkeley, Berkeley, CA, USA, May 1976. various pp.

Uhlir:1978:CRH

[Uhl78] S. A. Uhlir. Comnd — a routine to handle command line parsing. Technical Memorandum 1260 (TM 78-3231-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 10, 1978. ?? pp.

Uhler:1987:MWS

[Uhl87]  Stephen A. Uhler.  MGR

— a window system for UNIX. In USENIX Association [USE87b], pages 106–?? ISBN ???? LCCN ???? Abstract only.

Umali:1996:RUT

[Uma96] Rick Umali. Review: UNIX tamed. ;login: the USENIX Association newsletter, 21(4): 57–??, August 1996. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Ungar:1978:UDD

[Ung78] D. M. Ungar. A UNIX device driver for the PCL11-A — a parallel interprocessor communications …. Technical Memorandum 1242 (TM 78-3124-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 25, 1978. ?? pp.

Unger:1986:PJF

[Ung86]  Brian W. Unger.  Project

JADE final report. Report, University of Calgary, Calgary, AB, Canada, April 1,

1986.  URL http://hdl.
handle.net/1880/46070.

USG:1977:UPM

[UNI77] UNIX Support Group (USG).

UNIX programmer’s manual — PG-1C300 issue 3. Technical Memorandum 1088, AT&T Bell Laboratories, Murray Hill, NJ, USA, February 1977. ?? pp.

UCBCC:1982:UFS

[Uni82] University of California, Berkeley.Computing Services. UNIX FORTRAN subroutines. Unx; 1.2.7 unx; 1.2.07., Computing Services, University of California, Berkeley, CA, USA, 1982. various pp.

UniPress:1985:UESc

[Uni85] UniPress Software, Inc., Edison, NJ, USA. UniPress EMACS screen editor: UNIX EMACS reference manual, version 2.0 edition, 1985. various pp.

UI:1993:PMA

[UNI93]  UNIX International.  Per-

formance management activities within UNIX International. ACM SIGMETRICS Performance Evaluation Review, 21 (2):42, December 1993. CO-DEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

UC:1996:PCT

[Uni96a] Unicode Consortium, editor. Pre-conference tutorials proceedings: Software development + the Internet: going global with Unicode: Ninth International Unicode Conference, San Jose, CA, September 4–6 1996. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, 1996. ISBN

???? LCCN ???? A1. Input method design / by Mark Leisher – A2. Weaving the multilingual web: standards and their implementations / by Martin Durst … [et al.] – B1. National language and Unicode support in relational databases and SQL2/3 / by Stefan Buchta – B2. The Unicode Standard: version 2 / by Asmus Freytag – C1/C2. Non-Latin writing systems: characteristics and impacts on multinational product design / by Richard Ishida.

UC:1996:SDI

[Uni96b] Unicode Consortium, editor. Software Development + the Internet: Going Global with Unicode: Ninth International Unicode Conference, 4–6 September, San Jose, California, USA. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1996. ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc9conf/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

UC:1997:ESI

[Uni97] Unicode Consortium, editor.

Europe, Software + the Internet: Going Global with Unicode: Tenth International Unicode Conference, March 10–12, 1997, Mainz, Germany. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1997. ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc10/;   http://www.unicode.
org/unicode/uni2book/u2ord.
html.

UC:1998:ASI

[Uni98a] Unicode Consortium, editor.

Asia, Software + the Internet: Going Global with Unicode: The Twelfth International Unicode/ISO10646 Conference (IUC12) April 8–10, 1998, Tokyo, Japan. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-

5082, E-mail: unicode-inc@

unicode.org, 1998. ISBN ???? LCCN ???? URL http:/

/www.unicode.org/unicode/
iuc12/index.html;   http:
//www.unicode.org/unicode/
uni2book/u2ord.html.

UC:1998:TIU

[Uni98b] Unicode Consortium, editor. Thirteenth International Unicode Conference: Software + the Internet: Going Global with Unicode (R), September 8–11, 1998, San Jose, California. The Unicode Consortium, P.O. Box 700519, San Jose, CA 95170-0519, USA, Phone: +1-408-777-5870, Fax: +1-408-777-5082, E-mail: unicode-

inc@unicode.org, 1998. ISBN

???? LCCN ???? URL http:/

/www.reuters.com/unicode/
iuc13;   http://www.unicode.
org.

Hillebrand:1995:MLM

[uRH95] A. Sorgatz und R. Hillebrand. Mathematik unter Linux: Mu-PAD — Ein Computeralgebra System I. (German) [Mathematics in Linux: MuPAD — a computer algebra system, I]. Linux Magazin, 12/95:11–14, 1995.

Hillebrand:1996:MLM

[uRH96] A. Sorgatz und R. Hillebrand.

Mathematik unter Linux: Mu-PAD — Ein Computeralgebra System II. (German) [Mathematics in Linux: MuPAD — a computer algebra system, II]. Linux Magazin, 2/96, 3/96 (Nachdruck):60–67, 1996.

Uttal:1985:TIU

[URK85] Judi Uttal, Jeff Rothschild, and Charles Kline. Transparent integration of UNIX and MS-DOS. In USENIX Association [USE85c], pages 104–116.

Usas:1978:TCS

[Usa78] A. M. Usas. The Tandem computer system — a first look. Technical Memorandum 1330 (TM 78-1352-11), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 23, 1978. ?? pp.

USENIX:1982:UCP

[USE82] USENIX, editor. USENIX Conference Proceedings, Summer, 1982. Boston, MA. USENIX Association, Berkeley, CA, USA, Summer 1982.

USENIX:1983:UCPa

[USE83] USENIX, editor. USENIX Conference Proceedings, Winter, 1983. San Diego, CA. USENIX Association, Berkeley, CA, USA, Winter 1983.

USENIX:1984:UCP

[USE84a] USENIX, editor. USENIX Conference Proceedings, Summer, 1984. Salt Lake City, UT.

USENIX Association, Berkeley, CA, USA, Summer 1984.

USENIX:1984:UUC

[USE84b] USENIX Association, editor. USENIX UniForum Conference Proceedings, January 17– 20, 1984. Washington DC, USA. USENIX Association, Berkeley, CA, USA, January 17–20, 1984. ISBN none. LCCN QA76.8.U65 U55 1984.

USENIX:1985:UCPa

[USE85a] USENIX, editor. USENIX Conference Proceedings, Summer, 1985. Portland, OR. USENIX Association, Berkeley, CA, USA, Summer 1985.

USENIX:1985:UCPb

[USE85b] USENIX, editor. USENIX Conference Proceedings, Winter, 1985. Dallas, TX. USENIX Association, Berkeley, CA, USA, Winter 1985.

USENIX:1985:PUA

[USE85c] USENIX Association, editor. Proceedings: USENIX Association Winter Conference, January 23–25, 1985, Dallas, Texas, USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1985.

USENIX:1985:SCP

[USE85d] USENIX Association, editor. Summer conference proceedings, Portland 1985: June 11–14, 1985, Portland, Oregon USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1985. LCCN QA76.8.U65 U8 1985.

USENIX:1986:UCPa

[USE86a] USENIX, editor. USENIX Conference Proceedings, Summer, 1986. Atlanta, GA. USENIX Association, Berkeley, CA, USA, Summer 1986.

USENIX:1986:UCPb

[USE86b] USENIX, editor. USENIX Conference Proceedings, Winter, 1986. Denver, CO. USENIX Association, Berkeley, CA, USA, Winter 1986.

USENIX:1986:SCP

[USE86c] USENIX Association, editor.

Summer conference proceedings, Atlanta 1986: June 9–13, 1986, Atlanta, Georgia, USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1986.

USENIX:1986:TCG

[USE86d] USENIX Association, editor. Third Computer Graphics Workshop Proceedings, November 20–21, 1986. Monterey, CA. USENIX Association, Berkeley, CA, USA, November 20–21, 1986. ISBN ???? LCCN ????

USENIX:1986:UAW

[USE86e] USENIX Association, editor. USENIX Association Winter Conference proceedings: January 15–17, 1986, Denver, Colorado USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1986. URL https://

bitsavers.org/pdf/usenix/
USENIX_1986_Winter_Technical_
Conference_Proceedings_198601.
pdf.

USENIX:1987:ECP

[USE87a] USENIX Association, editor. EUUG Conference Proceedings, Autumn, 1987. Dublin, Ireland. European UNIX Users Group, Buntingford, Herts, UK, Autumn 1987. ISBN ???? LCCN ????

USENIX:1987:FCG

[USE87b] USENIX Association, editor. Fourth Computer Graphics Workshop Proceedings, October 8–9, 1987. Cambridge, MA. USENIX Association, Berkeley, CA, USA, October 8–9, 1987. ISBN ???? LCCN ????

USENIX:1987:LIS

[USE87c] USENIX Association, editor. Large Installation System Administrators Workshop Proceedings, April 9–10, 1987. Philadelphia, PA. USENIX Association, Berkeley, CA, USA, April 9–10, 1987. ISBN ???? LCCN TK 5105.5 L374 Bar B5-6(1991-92).B9(1995).

USENIX:1987:PSU

[USE87d] USENIX Association, editor. Proceedings of the Summer 1987 USENIX Conference: June 8–12, 1987, Phoenix, Arizona, USA. USENIX Association, Berkeley, CA, USA, 1987.

USENIX:1987:UAW

[USE87e] USENIX Association, editor.

USENIX Association Winter Conference proceedings, Washington, DC, 1987: January 21– 23, 1987, Washington, District of Columbia, USA. USENIX Association, P.O. Box 7, El Cerrito 94530, CA, USA, 1987. LCCN QA 76.76 O63 U84 1987. Cover title: Conference proceedings. Spine title: Winter 1987 USENIX Association Conference proceedings, Washington, DC.

USENIX:1988:PUS

[USE88a]  USENIX, editor.  Proceed-

ings. UNIX Security Workshop, Portland, OR, USA, August 29–30, 1998. USENIX Association, Berkeley, CA, USA, 1988.

USENIX:1988:USWa

[USE88b] USENIX, editor. UNIX Security Workshop Proceedings, August 29–30, 1988. Portland, OR. USENIX Association, Berkeley, CA, USA, August 29– 30, 1988. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

USENIX:1988:ECP

[USE88c] USENIX Association, editor.

EUUG Conference Proceedings, Spring, 1988. London, England. European UNIX Users Group, Buntingford, Herts, UK, Spring 1988. ISBN ???? LCCN ????

USENIX:1988:LIS

[USE88d] USENIX Association, editor. Large Installation Systems Administration Workshop Proceedings, November 17–18, 1988. Monterey, CA. USENIX Association, Berkeley, CA, USA, November 17–18, 1988. ISBN ???? LCCN ????

USENIX:1988:PFU

[USE88e] USENIX Association, editor. Proceedings of the (First) USENIX Security Workshop, August 29–30, 1988, Portland, OR, USA. USENIX Association, Berkeley, CA, USA, 1988. LCCN QA76.8.U65 U55 1988(1)-1990(2)//.

USENIX:1988:USWb

[USE88f] USENIX Association, editor. UNIX and Supercomputers Workshop Proceedings, September 26–27, 1988. Pittsburgh, PA. USENIX Association, Berkeley, CA, USA, September 26–27, 1988. ISBN ???? LCCN ????

USENIX:1988:UCPb

[USE88g] USENIX Association, editor. USENIX Conference Proceedings (Dallas, TX, USA). USENIX Association, Berkeley, CA, USA, Winter 1988. ISBN ???? LCCN ????

USENIX:1988:UPC

[USE88h] USENIX Association, editor. USENIX proceedings: C++ Conference, Denver, CO, October 17–21, 1988. USENIX Association, Berkeley, CA, USA, 1988.

USENIX:1989:UCPb

[USE89a] USENIX, editor. USENIX Conference Proceedings, Summer, 1989. Baltimore, MD. USENIX Association, Berkeley, CA, USA, Summer 1989.

USENIX:1989:UCPa

[USE89b] USENIX, editor. USENIX Conference Proceedings, Winter, 1989. San Diego, CA. USENIX Association, Berkeley, CA, USA, Winter 1989.

USENIX:1989:DMS

[USE89c] USENIX Association, editor.

Distributed and Multiprocessor Systems Workshop Proceedings, October 5–6, 1989. Fort Lauderdale, FL. USENIX Association, Berkeley, CA, USA, October 5–6, 1989. ISBN ???? LCCN ????

USENIX:1989:LIS

[USE89d] USENIX Association, editor.

Large Installation Systems Administration III Workshop Proceedings, September 7–8, 1989. Austin, TX. USENIX Association, Berkeley, CA, USA, September 7–8, 1989. ISBN ???? LCCN ????

USENIX:1989:PSU

[USE89e] USENIX Association, editor.

Proceedings of the Summer 1989 USENIX Conference: June 12 — June 16, 1989, Baltimore, Maryland USA. USENIX Association, Berkeley, CA, USA, 1989. LCCN QA 76.76 O63 U83 1989.

USENIX:1989:PWU

[USE89f] USENIX Association, editor. Proceedings of the Winter 1989 USENIX Conference: January 30–February 3, 1989, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 1989.

USENIX:1989:UTP

[USE89g] USENIX Association, editor. UNIX Transaction Processing Workshop Proceedings, May 1–2, 1989. Pittsburgh, PA. USENIX Association, Berkeley, CA, USA, May 1–2, 1989. ISBN ???? LCCN ????

USENIX:1989:UFC

[USE89h] USENIX Association, editor. USENIX Fifth Computer Graphics workshop, November 16–17, 1989. Monterey, CA. USENIX Association, Berkeley, CA, USA, November 16– 17, 1989. ISBN ???? LCCN ????

USENIX:1990:CSF

[USE90a] USENIX, editor. Computing Systems, Fall, 1990. USENIX Association, Berkeley, CA, USA, Fall 1990.

USENIX:1990:PFL

[USE90b]  USENIX, editor.  Proceed-

ings of the fourth Large Installation System Administrator’s Conference: October 17– 19, 1990, Colorado Springs, Colorado, USA. USENIX Association, Berkeley, CA, USA, 1990. LCCN QA76.76.O63 L37 1990.

USENIX:1990:PSU

[USE90c] USENIX, editor. Proceedings of the Summer 1990 USENIX Conference: June 11–15, 1990, Anaheim, California, USA. USENIX Association, Berkeley, CA, USA, 1990. LCCN QA76.8.U65 U81 1990.

USENIX:1990:PWU

[USE90d] USENIX, editor. Proceedings of the Winter 1990 USENIX Conference, January 22–26, 1990, Washington, DC, USA. USENIX Association, Berkeley, CA, USA, 1990. LCCN QA76.8.U65 U82 1990.

USENIX:1990:USI

[USE90e] USENIX Association, editor.

UNIX Security II: USENIX workshop proceedings, August 27–28, 1990, Portland, Oregon. USENIX Association, Berkeley, CA, USA, 1990. ISBN ???? LCCN QA 76.9 A25 U55 1990.

USENIX:1991:LIC

[USE91a] USENIX, editor. LISA IV Conference Proceedings, October 24–25, 1991. Colorado Springs, CO. USENIX Association, Berkeley, CA, USA, October 24–25, 1991. ISBN ???? LCCN ????

USENIX:1991:PFL

[USE91b] USENIX, editor. Proceedings of the fifth Large Installation Systems Administration Conference: September 30–October 3, 1991, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 1991.

USENIX:1991:PSU

[USE91c] USENIX, editor. Proceedings of the Summer 1991 USENIX Conference: June 10–14, 1991, Nashville, Tennessee, USA. USENIX Association, Berkeley, CA, USA, 1991. LCCN QA 76.76 O63 U83 1991.

USENIX:1991:PWU

[USE91d] USENIX, editor. Proceedings of the Winter 1991 USENIX Conference: January 21–January 25, 1991, Dallas, TX, USA. USENIX Association, Berkeley, CA, USA, 1991. LCCN QA 76.76 O63 U84 1992.

USENIX:1991:SED

[USE91e] USENIX, editor. Symposium on Experiences with Distributed and Multiprocessor Systems (SEDMS II), March 21–22, 1991. Atlanta, GA. USENIX Association, Berkeley, CA, USA, March 21–22, 1991. ISBN ???? LCCN ????

USENIX:1991:UCPc

[USE91f] USENIX, editor. USENIX Conference Proceedings, Winter, 1991. Dallas, TX. USENIX Association, Berkeley, CA, USA, Winter 1991.

USENIX:1992:PSS

[USE92a]  USENIX, editor.  Proceed-

ings of the sixth Systems Administration Conference (LISA VI): October 19–23, 1992, Long Beach, California, USA. USENIX Association, Berkeley, CA, USA, 1992. ISBN 1-880446-47-2. LCCN ????

USENIX:1992:PSU

[USE92b] USENIX, editor. Proceedings of the Summer 1992 USENIX Conference: June 8–12, 1992, San Antonio, Texas, USA. USENIX Association, Berkeley, CA, USA, Summer 1992. ISBN 1-880446-44-8. LCCN QA 76.76 O63 U83 1992.

USENIX:1992:PUF

[USE92c] USENIX, editor. Proceedings of the USENIX File Systems Workshop, 21–22 May, 1992, Ann Arbor, MI, USA. USENIX Association, Berkeley, CA, USA, 1992. ISBN 1-880446-43-X. LCCN QA76.8.U65 U84 1992.

USENIX:1992:PWU

[USE92d] USENIX, editor. Proceedings of the Winter 1992 USENIX Conference: January 20 — January 24, 1992, San Francisco, California. USENIX Association, Berkeley, CA, USA, 1992.

USENIX:1992:USI

[USE92e] USENIX, editor. UNIX Security III Symposium, September 14–17, 1992. Baltimore, MD. USENIX Association, Berkeley, CA, USA, September 14– 17, 1992. ISBN 1-880446-46-4. LCCN ????

USENIX:1993:CSSa

[USE93a]  USENIX, editor.  Comput-

ing Systems, Spring, 1993. USENIX Association, Berkeley, CA, USA, Spring 1993.

USENIX:1993:PSS

[USE93b] USENIX, editor. Proceedings of the Seventh Systems Administration Conference (LISA VII): November 1–5, 1993, Monterey, CA, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/lisa93/.

USENIX:1993:PSU

[USE93c]  USENIX, editor.  Proceed-

ings of the Summer 1993 USENIX Conference: June 21– 25, 1993, Cincinnati, Ohio, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/cinci93/.

USENIX:1993:PUMa

[USE93d] USENIX, editor. Proceedings of the USENIX Mach III Symposium: April 19–21, 1993, Santa Fe, New Mexico, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-49-9. LCCN QA 76.76 O63 U85 1993.

USENIX:1993:PUMb

[USE93e] USENIX, editor. Proceedings of the USENIX Mobile and Location-Independent Computing Symposium: August 2– 3, 1993, Cambridge, Massachusetts, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-51-0. LCCN QA 76.76 O63 U86 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/mobile93/.

USENIX:1993:PWU

[USE93f] USENIX Association, editor.

Proceedings of the Winter 1993 USENIX Conference: January 25–29, 1993, San Diego, California, USA. USENIX Association, Berkeley, CA, USA, 1993. ISBN 1-880446-48-0. LCCN QA 76.76 O63 U84 1993.

USENIX:1993:USI

[USE93g] USENIX Association, editor.

UNIX Security IV Symposium proceedings: October 4– 6, 1993, Santa Clara, CA, USA. USENIX Association, Berkeley, CA, USA, October 4–6, 1993. ISBN 1-880446-55-3. LCCN QA 76.9 A25

U54 1993. URL http://www.
usenix.org/publications/
library/proceedings/sec4/.

USENIX:1994:CSW

[USE94a]  USENIX, editor.  Comput-

ing Systems, Winter, 1994. USENIX Association, Berkeley, CA, USA, Winter 1994.

USENIX:1994:PUSa

[USE94b] USENIX, editor. Proceedings of the 1994 USENIX Symposium on High-Speed Networking, August 1–3, 1994, Oakland, California, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-63-4. LCCN TK5105.5 .U84 1994.

USENIX:1994:PUU

[USE94c]  USENIX, editor.  Proceed-

ings of the 1994 USENIX UNIX Applications Development Symposium: April 25– 28, 1994, Toronto, Ontario, Canada. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-61-8. LCCN QA76.76.D47 U55 1994.

                                        URL  http://www.usenix.
                                        org/publications/library/
                                        proceedings/appdev94/index.
                                        html.

USENIX:1994:PES

[USE94d] USENIX, editor. Proceedings of the Eighth Systems Administration Conference (LISA VIII): September 19–23, 1994, San Diego, CA, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-64-2. LCCN QA 76.76 O63 L37 1994.

USENIX:1994:PSU

[USE94e] USENIX, editor. Proceedings of the Summer 1994 USENIX Conference: June 6–10, 1994, Boston, Massachusetts, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-62-6. LCCN QA 76.76 O63 U83 1994.

USENIX:1994:PWU

[USE94f] USENIX, editor. Proceedings of the Winter 1994 USENIX Conference: January 17–21, 1994, San Francisco, California, USA. USENIX Association, Berkeley, CA, USA, 1994. ISBN 1-880446-58-8. LCCN QA 76.76 O63

U84 1994. URL http://www.
usenix.org/publications/
library/proceedings/sf94/
index.html.

USENIX:1995:UUS

[USE95a] USENIX, editor. 5th USENIX UNIX Security Symposium, June 5–7, 1995. Salt Lake City, UT. USENIX Association, Berkeley, CA, USA, June 5–7, 1995.

USENIX:1995:PUT

[USE95b] USENIX, editor. Proceedings of the 1995 USENIX Technical Conference: January 16–20, 1995, New Orleans, Louisiana, USA. USENIX Association, Berkeley, CA, USA, 1995. ISBN 1-880446-67-7. LCCN QA 76.76 O63 U88 1995.

URL  http://www.usenix.
org/publications/library/
proceedings/neworl/index.
html.

USENIX:1995:PFUa

[USE95c] USENIX Association, editor. Proceedings of the fifth USENIX UNIX Security Symposium: June 5–7, 1995, Salt Lake City, Utah, USA. USENIX Association, Berkeley, CA, USA, 1995. ISBN 1-880446-70-7. LCCN QA76.8.U65 U55 1992(3)-

                                        1995(5). URL http://www.
                                        usenix.org/publications/
                                        library/proceedings/security95/
                                        index.html.

USENIX:1996:PTS

[USE96a] USENIX, editor. Proceedings of the Tenth Systems Administration Conference (LISA X), September 29–October 4, 1996, Chicago, IL, USA. USENIX Association, Berkeley, CA, USA, 1996. ISBN 1-880446-81-2. LCCN QA76.76.O63 S97 1996.

USENIX:1996:PUA

[USE96b] USENIX, editor. Proceedings of the USENIX 1996 annual technical conference: January 22– 26, 1996, San Diego, California, USA, USENIX Conference Proceedings 1996. USENIX Association, Berkeley, CA, USA, 1996. ISBN 1-880446-76-6. LCCN QA 76.76 O63 U88 1996.

USENIX:1997:DLS

[USE97a] USENIX, editor. Digest of the Large Scale System Administration of Windows NT Workshop, August 14–16, 1997, Seattle, Washington, USA. USENIX Association, Berkeley, CA, USA, 1997. ISBN ???? LCCN ????

USENIX:1997:PES

[USE97b] USENIX, editor. Proceedings of the Eleventh Systems Administration Conference (LISA XI), October 26–31, 1997, San Diego, CA, USA. USENIX Association, Berkeley, CA, USA, 1997. ISBN 1-880446-90-1. LCCN QA76.76.O63 S97 1997.

URL  http://www.usenix.
org/publications/library/
proceedings/lisa97/.

USENIX:1997:PUA

[USE97c] USENIX, editor. Proceedings of the USENIX 1997 annual Technical Conference, January 6–10, 1997, Anaheim, California, USA. USENIX Association, Berkeley, CA, USA, 1997. ISBN 1-880446-84-7. LCCN QA76.76.O63 U88 1997.

USENIX:1997:PUW

[USE97d]  USENIX, editor.  Proceed-

ings of the USENIX Windows NT Workshop: August 11–13, 1997, Seattle, Washington. USENIX Association, Berkeley, CA, USA, 1997. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997.

USENIX:1998:PUWa

[USE98a] USENIX, editor. Proceedings of the 2nd USENIX Windows NT Symposium: August 3–5, 1998, Seattle, Washington. USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-95-2. LCCN QA76.76.O63 U885 1998. URL http://db.

usenix.org/publications/
library/proceedings/usenix-
nt98.

USENIX:1998:PFT

[USE98b] USENIX, editor. Proceedings of the FreeNIX Track: USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USENIX Association, Berkeley, CA, USA, 1998. ISBN ???? LCCN ????

URL  http://www.usenix.
org/publications/library/
proceedings/usenix98/freenix/

.

USENIX:1998:PLI

[USE98c] USENIX, editor. Proceedings of the Large Installation System Administration of Windows NT Conference, August 5– 8, 1998, Seattle, Washington. USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-96-0. LCCN TK5105.5 .L37 1998. URL http://www.

usenix.org/publications/
library/proceedings/lisa-
nt98/.

USENIX:1998:PTS

[USE98d]  USENIX, editor.  Proceed-

ings of the Twelfth Systems Administration Conference (LISA XII). USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-40-5. LCCN QA76.76.O63 S97 1998. URL http://db.

usenix.org/publications/
library/proceedings/lisa98.

USENIX:1998:PUA

[USE98e]  USENIX, editor.  Proceed-

ings of the USENIX 1998 annual technical conference: June 15–19, 1998, New Orleans, LA. USENIX Association, Berkeley, CA, USA, 1998. ISBN 1-880446-94-4. LCCN QA76.76.O63 U88

1998.  URL http://db.
usenix.org/publications/
library/proceedings/usenix98.

USENIX:1999:PUWb

[USE99a] USENIX, editor. Proceedings of the 3rd USENIX Windows NT Symposium: July 12–15, 1999, Seattle, Washington, USA. USENIX Association, Berkeley, CA, USA, 1999. ISBN 1-880446-29-4. LCCN ????

URL  http://www.usenix.
org/publications/library/
proceedings/usenix-nt99/.

USENIX:1999:PEU

[USE99b] USENIX, editor. Proceedings of the eighth USENIX Security Symposium (Security ’99), August 23–26, 1999, Washington, DC, USA. USENIX Association, Berkeley, CA, USA, 1999. ISBN 1-880446-28-6. LCCN QA76.9.A25 U83 1999.

URL  http://www.usenix.
org/publications/library/
proceedings/sec99/.

USENIX:1999:PUWa

[USE99c]  USENIX, editor.  Proceed-

ings of the USENIX Workshop on Smartcard Technology (Smartcard ’99): May 10–11, 1999, Chicago, Illinois, USA. USENIX Association, Berkeley, CA, USA, 1999. ISBN 1-880446-34-0. LCCN TK7895.S62 U84 1999.

URL  http://www.usenix.
org/publications/library/
proceedings/smartcard99/.

USENIX:2000:UAT

[USE00a] USENIX, editor. 2000 USENIX Annual Technical Conference: San Diego, CA, USA, June 18–23, 2000. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-22-7.

LCCN ???? URL http://www.
usenix.org/publications/
library/proceedings/usenix2000.

USENIX:2000:PFSa

[USE00b] USENIX, editor. Proceedings of the Fourteenth Systems Administration Conference (LISA XIV), December 3–8, 2000, New Orleans, Louisiana, USA. USENIX Association, Berkeley, CA, USA, 2000. ISBN 1-880446-13-8. LCCN ????

URL  http://www.usenix.
org/publications/library/
proceedings/lisa2000/technical.
html.

USENIX:2001:PLS

[USE01a] USENIX, editor. Proceedings of LISA 2001: 15th Systems Administration Conference, December 2–7, 2001, Town and Country Resort Hotel, San Diego, CA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-05-7. LCCN

???? URL http://db.usenix.
org/publications/library/
proceedings/lisa2001/tech/

.

USENIX:2001:PFT

[USE01b]  USENIX, editor.  Proceed-

ings of the FREENIX Track: 2001 USENIX Annual Technical Conference: June 25– 30, 2001, Marriott Copley Place Hotel, Boston, Massachusetts, USA. USENIX Association, Berkeley, CA, USA, 2001. ISBN 1-880446-10-3. LCCN QA76.8.U65 U84 2001.

URL  http://www.usenix.
org/publications/library/
proceedings/usenix01/freenix01/
technical.html.

USENIX:2002:PBF

[USE02] USENIX, editor. Proceedings of BSDCon 2002: February 11– 14, 2002, Cathedral Hill Hotel, San Francisco, CA. USENIX Association, Berkeley, CA, USA, 2002. ISBN 1-880446-02-2. LCCN QA76.76.O63 B736 2002. URL http://www.

usenix.org/publications/
library/proceedings/bsdcon02/
tech.html.

USENIX:1982:UAS

[Usr82] Usr Group, editor. /usr/ group, USENIX Association, Software Tools Users Group Joint Conference Proceedings: Boston, July 1982. USENIX Association, Berkeley, CA, USA, 1982.

Urban:2003:STY

[UT03] Michael Urban and Brian Tiemann. Sams teach yourself FreeBSD in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 2003. ISBN 0-672-32424-5. xvii + 433 pp. LCCN QA76.76.O63 U647 2003.

Ullman:1984:BTB

[UTC84] Ellen Ullman, Page Thompson, and Jerry Carlin. Building tunnels and bridges: Constructing a commercial application under UNIX. In USENIX Association [USE84b], pages 259–284. ISBN none. LCCN QA76.8.U65 U55 1984.

Vadala:2002:MRL

[Vad02] Derek Vadala. Managing RAID on Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 1-56592-730-3. 304 (est.) pp. LCCN ???? US$39.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        mraidlinux.

Vahalia:1996:UI

[Vah96] Uresh Vahalia. UNIX Internals. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1996. ISBN 0-13-101908-2. xxxiii + 601 pp. LCCN QA76.76.063V33 1996.

Valley:1992:CPU

[Val92a]  John Valley.  C Program-

ming for UNIX. Howard W. Sams, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-48518-

4. xxvii + 644 pp. LCCN QA76.73.C15 V35 1992. A comprehensive book on C programming on UNIX systems.

Valley:1992:UDG

[Val92b] John Valley. UNIX Desktop Guide to the Korn Shell. Howard W. Sams, Indianapolis, IN 46268, USA, 1992. ISBN 0-672-48513-3. xxxii + 455 pp. LCCN QA76.73.K67V35 1992. This one is, in my humble opinion, even better and easier to read than the authoritative work by Korn and Bolsky [BK89].

VanWyk:1980:LTG

[Van80] Christopher John Van Wyk.

A Language for Typesetting Graphics. Ph.D. thesis, Department of Computer Science, Stanford University, Stanford, CA, USA, June 1980. v +

59 pp. URL https://www.
proquest.com/pqdtglobal/
docview/302984144/7AF5FACEA77C4EA1PQ.

VanWyk:1981:IUM

[Van81] Christopher J. Van Wyk.

IDEAL user’s manual. Computing Science Technical Report 103, AT&T Bell Laboratories, Murray Hill, NJ, USA, December 17, 1981. ii + 25 pp.

vanKatwijk:1983:PYP

[van83] J. van Katwijk. A preprocessor for YACC. A poor man’s approach to parsing attributed grammars. ACM SIGPLAN Notices, 18(10):12–15, October 1983. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Veer:1986:UEO

[van86] Gerrit C. van der Veer. UNIX and the electronic office — cognitive ergonomic reflections. EUUG Newsletter, 6(3):41– 49, Winter 1986. CODEN EONLE8. ISSN 1011-4211.

VanWyk:1988:DSC

[Van88] Christopher J. Van Wyk. Data Structures in C. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-16116-8. x + 387 pp. LCCN QA76.73.C15 V36 1988.

VanWyk:1990:IPS

[Van90] C. J. Van Wyk. Ideal — a picture-specification language. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

1990. URL http://doc.cat-
v.org/unix/v10/10thEdMan/
v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

VanRaalte:1991:XRM

[Van91] Thomas Van Raalte. XView Reference Manual. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, September 1991. ISBN 0-937175-88-9. 291 pp. LCCN QA76.76.W56 X85 1991. US$24.95.

vanLicht:1997:PRR

[van97] Ernst van Licht. Product review: RoboMon UNIX v6.1. Sys Admin: The Journal for UNIX Systems Administrators, 6(5):55–??, May 1997. CODEN SYADE7. ISSN 1061-2688.

VanBon:2005:FIS

[Van05] Jan Van Bon, editor. Foundations of IT service management based on ITIL. Van Haren Publishing, Zaltbommel, The Netherlands, second edition, 2005. ISBN 90-77212-58-2. 233 pp. LCCN ????

Varian:1991:VVC

[Var91] Melinda Varian. VM and the VM community: Past, present, and future. Technical report, Office of Computing and Information Technology, Princeton University, Princeton, NJ 08544, USA, April 1991.

168 pp. URL http://www.
leeandmelindavarian.com/
Melinda/neuvm.pdf. Origi-

nal presented at Australasian SHARE/GUIDE in Melbourne, Victoria, Australia in 1989. This is a detailed history of the development of virtual machine technology on IBM System/360 and later mainframes, and of the opposition by much of IBM to that technology until it was demonstrated that their performance could equal, or even exceed, that of an operating system running on bare hardware, and also allow a single physical host to support multiple operating systems, and software development, simultaneously. There are also several comments about the development of the REXX language, and about the influence of Unix on IBM’s software development.

Varhol:1996:FUB

[Var96a] Peter D. Varhol. The future of UNIX: 64 bits — so far, only Digital and SGI can give you the OS you need to get the main benefit of 64-bit UNIX RISC machines: gazillions of gigabytes of memory. Datamation, 42(11):26–??, ???? 1996. CO-DEN DTMNAT. ISSN 0011-6963.

Varhol:1996:NRD

[Var96b] Peter D. Varhol. NT/RISC: a desktop combo for the future — computing managers looking for an alternative to UNIX on RISC systems may finally have found a good one in Windows NT. it can spread a single look and feel across an enterprise, but for desktop use more general-purpose business apps need to be developed. Datamation, 42(2):27–??, ???? 1996. CODEN DTMNAT. ISSN 0011-6963.

Verbaeten:1983:PU

[VB83] P. Verbaeten and Y. Berbers. Porting UNIX. In Association [Ass83], pages 285–289.

VanEpp:1992:DMC

[VB92] Peter Van Epp and Bill Baines.

Dropping the mainframe without crushing the users: Mainframe to distributed UNIX in nine months. In USENIX [USE92a], pages 39–53 (or 39– 54??). ISBN 1-880446-47-2. LCCN ????

VanEpp:1994:DMC

[VB94] Peter Van Epp and Bill Baines.

Dropping the mainframe without crushing the users. A biased choice: -) …. This is a paper describing the migration at Simon Fraser University from centralized mainframe (MTS) to distributed computing (UNIX) in 9 months ….

This paper (LISA-VI.paper.

ps.Z (PostScript — 16 pages)) can be obtained by anony-

mous ftp from fraser.sfu.ca
in /pub/papers/ucspapers.,

November 1994.

Vargas:2000:SCE

[VBD00] Enrique Vargas, Joseph Bianco, and David Deeths. Sun Cluster Environment: Sun Cluster 2.2. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2000. ISBN 0-13-041870-6. xxx + 390 pp. LCCN QA278.V37 2001. URL books/

sunmicrosystems_clusterbook.
pdf;   http://www.sun.com/
books/catalog/vargas/index.

html. Part No. 806-3345-10 October 2000, Revision 01.

Vargas:2001:SCE

[VBD01] Enrique Vargas, Joseph Bianco, and David Deeths. Sun Cluster Environment: Sun Cluster 2.2. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-041870-6. xxx + 389 pp. LCCN QA278.V37 2001. US$40.00.

URL http://www.phptr.com/
ptrbooks/ptr_0130418706.
html;   http://www.sun.com/
books/catalog/vargas/index.
html.

Vaamonde:2001:ISG

[VD01] Fernández Vaamonde and Manuel David. Implantación de un sistema de gestión centralizada de paquetes deb para su uso en configuraciones Debian GNU/Linux. (Spanish) [Implantation of a system of centralized management of deb packages for its use in configurations of Debian GNU/Linux ]. Technical report, Trabajos fin de carrera de la Facultade de Informática de A Coruña. Departamento de Electrónica e Sistemas, Facultade de Informática Universidade da Coruña, Coruña, Spain, 2001. 111 pp. Final engineering project. Includes one CD-ROM.

vanderLinden:1994:ECP

[vdL94] Peter van der Linden. Expert C Programming: Deep C Secrets. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 1994. ISBN 0-13-177429-8. xxvi + 353 pp. LCCN QA76.73.C15 V356 1994. URL

http://www.sun.com/books/
catalog/vanderlinden/index.
html.

Veach:2010:SSE

[Vea10] Sharon Veach. Solaris 10 security essentials. Solaris system administration series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2010. ISBN 0-13-701233-0 (paperback). xxii + 272 pp. LCCN QA76.9.A25 S65524 2010.

Venkateswaran:2008:ELD

[Ven08] Sreekrishnan Venkateswaran.

Essential Linux device drivers. Prentice Hall open source software development series. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2008. ISBN 0-13-239655-6 (hardback). xxx + 714 pp. LCCN QA76.76.D49 V35 2008. URL http:

//www.loc.gov/catdir/toc/
ecip087/2008000249.html.

Vernon:1989:DCC

[Ver89]  Vaughn Vernon.  Design

and coding considerations for UNIX/MS-DOS application portability. C Users Journal, 7 (1):59–??, January 1989. ISSN 0898-9788.

Vesonder:1988:RBP

[Ves88] Gregg T. Vesonder. Rule-based programming in the Unix system. AT&T Technical Journal, 67(1):69–80, January 1988.

CODEN ATJOEM. ISSN 2376-676X (print), 8756-2324 (electronic).

Vesperman:2006:EC

[Ves06] Jennifer Vesperman. Essential CVS. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2006. ISBN 0-596-52703-9. 428 (est.) pp. LCCN ????

Vaughan:2000:GAA

[VETT00] Gary V. Vaughan, Ben Elliston, Tom Tromey, and Ian Lance Taylor. GNU Autoconf, Automake and Libtool. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 1-57870-190-2. xx + 390 pp. LCCN QA76.76.O63 G598 2000. US$40.00. URL http://

sources.redhat.com/autobook/
;   http://sources.redhat.
com/autobook/autobook/autobook_
toc.html;   http://sources.
redhat.com/autobook/download.
html;   http://www.newriders.
com/books/title.cfm?isbn=
1578701902.

vanGumster:2010:GB

[vGS10] Jason van Gumster and Robert Shimonski. GIMP bible. Wiley, New York, NY, USA, 2010. ISBN 0-470-52397-2 (paperback), 0-470-63642-4 (ebook). xxxv + 722 pp. LCCN T385 .V36 2010.

vonHagen:2002:LF

[vH02] William von Hagen. Linux Filesystems. SAMS Publishing, Indianapolis, IN, USA, 2002. ISBN 0-672-32272-2. xii + 555 pp. LCCN QA76.76.O63 V66 200.

vonHagen:2006:ULB

[vH06] William von Hagen. Ubuntu Linux bible. Wiley, New York, NY, USA, 2006. ISBN 0-470-03899-3 (paperback). xxvi + 904 pp. LCCN QA76.76.O63 V665 2007.

vonHagen:2007:ULB

[vH07] William von Hagen. Ubuntu Linux bible. Wiley, New York, NY, USA, 2007. ISBN 0-470-03899-3, 0-470-12454-7 (electronic book). xxvi + 904 pp. LCCN QA76.76.O63 V685 2007eb.

vonHagen:2003:DGG

[vHW03] William von Hagen and Kurt Wall. The Definitive Guide to GCC. Apress, Berkeley, CA, USA, 2003. ISBN 1-59059-109-7. xxviii + 519 pp. LCCN QA76.76.C65 W36 2004.

vonHagen:2006:DGG

[vHW06] William von Hagen and Kurt Wall. The Definitive Guide to GCC. Apress, Berkeley, CA, USA, second edition, 2006. ISBN 1-59059-585-8. 584 (est.) pp. LCCN ????

Publishers:2011:UIQ

[Vib11] Vibrant Publishers. Unix interview questions you’ll most likely be asked. Interview questions you’ll most likely be asked. Vibrant Publishers, Erie, CO, USA, 2011. ISBN 1-4564-8274-2. ???? pp. LCCN ????

Vickery:1998:USP

[Vic98] Christopher Vickery. UNIX Shell Programmer’s Interactive Workbook. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-020064-6. 496 pp. LCCN QA76.76.O63V44 1998. US$34.99.

VonHagen:2006:LSH

[VJ06] William Von Hagen and Brian K. (Brian Kenneth) Jones. Linux server hacks. Vol. 2. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2006. ISBN 0-596-10082-5. xx + 456 pp. LCCN QA76.76.O63; QA76.76.O63eb. URL http:/

/www.oreilly.com/catalog/
9780596100827.

Volkerding:1996:LP

[VJR96] Patrick Volkerding, Eric F. Johnson, and Kevin Reichard. Linux programming. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1996. ISBN 1-55828-507-5. xxi + 374 pp. LCCN QA76.76.O63 V65 1996. US$34.95. Includes CD-ROM.

VanWyk:1979:GLT

[VK79] Cris J. Van Wyk and Brian W. Kernighan. A graphics language for typesetting. Technical Memorandum 1471 (TM 79-1274-10), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 13, 1979. ?? pp.

vandeGoor:1988:UMS

[vM88] A. J. van de Goor and A. Moolenaar. UNIX I/O in a multiprocessor system. In USENIX Association [USE88g], pages 251–258. ISBN ???? LCCN ????

Venton:2005:ULH

[VMKB05] T. Venton, M. Miller, R. Kalla, and A. Blanchard. Using Linux for hardware bring up, development, and manufacturing. IBM Systems Journal, 44(2): 319–329, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/venton.html;   http:
//www.research.ibm.com/journal/
sj/442/venton.pdf;   http:
//www.research.ibm.com/journal/
sj/442/venton.txt.

vandeGoor:1988:MUS

[vMM88] A. J. van de Goor, A. Moolenaar, and J. M. Mulder. Multiprocessor UNIX: Separate processing of I/O. In USENIX Association [USE88c], pages 123–

134. ISBN ???? LCCN ????

Vogel:1974:RSU

[Vog74] G. C. Vogel. Referencing standard UNIX documentation. Technical Memorandum ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 23, 1974. ?? pp. Memorandum For File.

Vogel:1976:PGC

[Vog76a] G. C. Vogel. Program generic control and documentation on UNIX. Technical Memorandum ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, May 17, 1976. ?? pp.

Vogel:1976:USC

[Vog76b] G. C. Vogel. UNIX support classification for PG-1C300 — issue 2. Technical Memorandum ?? (MF 76-8234-9), AT&T Bell Laboratories, Murray Hill, NJ, USA, January 30, 1976. ?? pp.

Vogel:1977:EPV

[Vog77a] G. C. Vogel. Easy phototypeset view graphs on UNIX. Technical Memorandum 1139 (MF 77-8234-055), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 18, 1977. ?? pp.

Vogel:1977:FPG

[Vog77b] G. C. Vogel. Formatting program generic documents on Unix. Technical Memorandum 1170 (MF 77-8234-57), AT&T Bell Laboratories, Murray Hill, NJ, USA, May 24, 1977. ?? pp.

Vogel:1977:UOS

[Vog77c] G. C. Vogel. The UNIX operating system — PG-1C300. Technical Memorandum 1162 (TM 77-8234-02), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 15, 1977. ?? pp.

Vogel:1978:ZAR

[Vog78a] G. C. Vogel. 8080/Z80 assembler reference manual. Technical Memorandum 1250 (TM 78-8234-6), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 21, 1978. ?? pp.

Vogel:1978:USS

[Vog78b] G. C. Vogel. UNIX standardization subcommittee proposed archiving facility. Technical Memorandum 1294 (MF 78-8234-99), AT&T Bell Laboratories, Murray Hill, NJ, USA, September 27, 1978. ?? pp.

Vogel:1979:SUC

[Vog79a] G. C. Vogel. Standard UNIX configurations — issue 2. Technical Memorandum 1365, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 27, 1979. ?? pp.

Vogel:1979:USF

[Vog79b] G. C. Vogel. UNIX specification — file system related system calls. Technical Memorandum 1493, AT&T Bell Laboratories, Murray Hill, NJ, USA, January 10, 1979. ?? pp.

Vogel:1979:UTA

[Vog79c] G. C. Vogel. UNIX/TS alarm facility. Technical Memorandum 1359, AT&T Bell Laboratories, Murray Hill, NJ, USA, March 7, 1979. ?? pp.

Vogt:1997:VUS

[Vog97]  Carsten Vogt.  Visualizing

UNIX synchronization operations. Operating Systems Review, 31(3):52–64, July 1997. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Vogel:1976:UOS

[Vog6 ] G. C. Vogel. The UNIX operating system — PG-1C300. Technical Memorandum 1162, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1976 (??). ?? pp.

Vollaro:1975:CSR

[Vol75] J. R. Vollaro. A continuous service real time clock for digital computers. Technical Memorandum 1048 (TM 75-1273-4), AT&T Bell Laboratories, Murray Hill, NJ, USA, March 13, 1975. ?? pp.

Volkman:1994:CLU

[Vol94] Victor R. Volkman. CUG417: LIBFTP for UNIX, CUG418: Rasmol molecular graphics. C/C++ Users Journal, 12(10): 119–??, October 1994. CODEN CCUJEX. ISSN 1075-2838.

Volkman:1997:CCSc

[Vol97] Victor R. Volkman. C/C++ sources: UNIX and database resources on the Internet. C/C++ Users Journal, 15(3): 89–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

Volkerding:2000:LSC

[VR00] Patrick Volkerding and Kevin Reichard. Linux system commands. M&T Books, M&T Publishing, Inc., 501 Galveston Drive, Redwood City, CA 94063, USA, 2000. ISBN 0-7645-4669-4. xiv + 462 pp. LCCN QA76.76.O63 V463 2000.

Vrenios:2003:LCA

[Vre03] Alex Vrenios. Linux cluster architecture. SAMS Publishing, Indianapolis, IN, USA, 2003. ISBN 0-672-32368-0. x + 247 pp. LCCN QA76.58 .V74 2002.

Volkerding:1997:LPE

[VRFJ97a] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. Linux in Plain English. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-542-3. iv + 571 pp. LCCN QA76.76.O63V645 1997. US$14.95, CDN$20.95.

URL http://www.mispress.
com/.

Volkerding:1997:LIS

[VRFJ97b] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. The Linux Internet Server. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-545-8. xxi + 530 pp. LCCN QA76.9.C55R444 1997. US$39.95, CDN$55.95. URL

http://www.mispress.com/.

Includes CD-ROM.

Volkerding:1997:LP

[VRFJ97c] Patrick Volkerding, Kevin Reichard, and Eric Foster-Johnson. Linux Programming. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1997. ISBN 1-55828-507-5. xxi + 374 pp. LCCN QA76.76.O63V65 1997. US$39.95, CDN$55.95. URL

                                        http://www.mispress.com/.

Includes CD-ROM.

Volkerding:1995:LCI

[VRJ95] Patrick Volkerding, Kevin Reichard, and Eric F. Johnson. Linux configuration and installation. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, 1995. ISBN 1-55828-426-5. xv + 463 pp. LCCN QA76.76.O63 V64 1995.

Volkerding:1996:LCI

[VRJ96] Patrick Volkerding, Kevin Reichard, and Eric F. Johnson. Linux configuration and installation. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, second edition, 1996. ISBN 1-55828-492-3. xix + 522 pp. LCCN QA76.76.O63 V64 1996. US$39.95. URL http://

                                        www.mispress.com/. Includes

two CD-ROMs.

Volkerding:1997:LCI

[VRJ97] Patrick Volkerding, Kevin Reichard, and Eric F. Johnson.

Linux Configuration and Installation. The MIS:Press slackware series. MIS Press, P. O. Box 5277, Portland, OR 97208-5277, USA, Tel: (503) 282-5215, third edition, 1997. ISBN 1-55828-566-0. xxix + 512 pp. LCCN QA76.76.O63V64 1997. US$39.95, CDN$55.95.

URL  http://smartbooks.
com/bw708linuxconfg.htm;
 http://www.mispress.com/.

Includes two CD-ROMs.

Victor:2010:OSS

[VSC^{+}10] Jeff Victor, Jeff Savit, Gary Combs, Simon Hayler, and Bob Netherton, editors. Oracle Solaris 10 system virtualization essentials. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2010. ISBN 0-13-708188-X (paperback). 384 (est.) pp. LCCN QA76.9.V5 O73 2010.

Vetter:1994:MWW

[VSW94] Ronald J. Vetter, Chris Spell, and Charles Ward. Mosaic and the World Wide Web. Computer, 27(10):49–57, October 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Valente:2023:HVH

[VTS^{+}23] Luca Valente, Yvan Tortorella, Mattia Sinigaglia, Giuseppe Tagliavini, Alessandro Capotondi, Luca Benini, and Davide Rossi. HULK-V: a heterogeneous ultra-low-power Linux capable RISC-V SoC. In IEEE, editor, 2023 Design, Automation & Test in Europe Conference & Exhibition (DATE), pages 1–6. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2023.

Welsh:1995:LBG

[W^{+}95] Matt Welsh et al. The Linux bible: the GNU testament. Yggdrasil Computing, Inc., Berkeley, CA, USA, third edition, 1995. ISBN 1-883601-12-6. 1596 pp. LCCN QA 76.76 O63 L56 1995.

Woo:1999:OPG

[W^{+}99] Mason Woo et al. OpenGL programming guide: the official guide to learning OpenGL, version 1.2. Addison-Wesley, Reading, MA, USA, third edition, 1999. ISBN 0-201-60458-2. xi + 730 pp. LCCN T385 .N435 1999.

Wilensky:1984:TUE

[WAC84] Robert Wilensky, Yigal Arens, and David N. Chin. Talking to UNIX in English: An overview of UC. Communications of the Association for Computing Machinery, 27(6):574–593, June 1984. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wagner:1979:DUH

[Wag79] M. R. Wagner. DPD: Unix– Honeywell interface programs. Technical Memorandum 1371, AT&T Bell Laboratories, Murray Hill, NJ, USA, April 6, 1979. ?? pp.

Waite:1986:CLA

[Wai86] W. M. Waite. The cost of lexical analysis. Software— Practice and Experience, 16 (5):473–488, May 1986. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wainer:1995:IRT

[Wai95] Gabriel A. Wainer. Implementing real-time services in MINIX. Operating Systems Review, 29(3):75–84, July 1995. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Walsh:1982:UUAa

[Wal82a] Daniel Walsh. UTS: UNIX on the Amdahl 470. In USENIX [USE82], pages 210–?? Abstract only.

Walsh:1982:UUAb

[Wal82b] Daniel Walsh. UTS: UNIX on the Amdahl 470. In USENIX [USE82], pages 247–?? Abstract only.

Walsh:1982:UUAc

[Wal82c] Daniel Walsh. UTS: UNIX on the Amdahl 470. In Usr Group [Usr82], pages 210–?? Abstract only.

Walli:1991:UURa

[Wal91a] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(4):18– 38, July/August 1991. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1991:UURb

[Wal91b] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(5):27–42, September/October 1991. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1991:UURc

[Wal91c] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 16(6):18– ??, November/December 1991. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1992:UURc

[Wal92a] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(3):19– 27, May/June 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1992:UURd

[Wal92b] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(4):30– 44, July/August 1992. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1992:UURe

[Wal92c] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(5):32–43, September/October 1992. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1992:UURf

[Wal92d] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 17(6):35– 39, November/December 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1992:UURa

[Wal92e] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 17(1):17–??, January/February 1992. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1992:UURb

[Wal92f] Stephen Walli. An update on UNIX-related standards activity. ;login: the USENIX Association newsletter, 17(2):38– 46, March/April 1992. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walls:1992:PVT

[Wal92g] Keith Walls. POSIX and VMS:

a technical view. The VAX professional, 14(5):27–29, September 1992. CODEN VAXPEN. ISSN 8750-9628.

Walli:1993:UURa

[Wal93a] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(1):27–46, January/February 1993. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1993:UURb

[Wal93b] Stephen Walli. An update on UNIX-related standards activities. ;login: the USENIX Association newsletter, 18(2):36– 45, March/April 1993. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1994:PMM

[Wal94a] S. R. Walli. Posix myths and make-believe: Practical portability and Posix. American Programmer, 7(8):28–??, ???? 1994. CODEN AMPRFD. ISSN 1048-5600.

Walli:1994:UAD

[Wal94b] Stephen Walli. UNIX applications development symposium report. ;login: the USENIX Association newsletter, 19(4): 5–8, August 1994. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walli:1995:EYE

[Wal95a] Stephen R. Walli. Everything you’ve ever wanted to know about X/Open UNIX, and the single UNIX specification. ;login: the USENIX Association newsletter, 20(3):35–39, June 1995. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Walnum:1995:DGP

[Wal95b] Clayton Walnum. 3-D graphics programming with OpenGL. Que Corporation, Indianapolis, IN, USA, 1995. ISBN 0-7897-0277-0. xiii + 511 pp. LCCN T385 .W36 1995.

WalnutCreek:1995:TT

[Wal95c] Walnut Creek CD-ROM. Tcl/Tk.

Publishers Group West, ????, 1995. ISBN 1-57176-023-7. LCCN QA76.73.T44. US$39.95.

WalnutCreek:1995:SL

[Wal95d] Walnut Creek CDROM. Slackware Linux, 1995. ISBN 1-57176-117-9. Two computer laser optical discs.

Walter:1995:PMS

[Wal95e] Stephen Walter. Put multiprocessing systems to work. II. UNIX Review, 13(1):39– ??, January 1995. CODEN UNRED5. ISSN 0742-3136.

Wallace:1996:FPG

[Wal96] C. S. Wallace. Fast pseudorandom generators for normal and exponential variates. ACM Transactions on Mathematical Software, 22(1):119–127, March 1996. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1996-22-1/p119-wallace/.

Walli:1997:WTT

[Wal97a] Stephen Walli. Win32 Tcl/Tk GUIs on UNIX apps on Windows NT. ;login: the USENIX Association newsletter, 22(6): ??, November 1997. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL http://www.
usenix.org/publications/
login/1997-11/guis.html.

Walli:1997:OTU

[Wal97b] Stephen R. Walli. OPENNT^{T M}: UNIX application portability to Windows NT^{T M} via an alternative environment subsystem. In USENIX [USE97d], pages 123–132. ISBN 1-880446-88-X. LCCN QA76.76.O63 U845 1997. URL http://www.

usenix.org/publications/
library/proceedings/usenix-
nt97/walli.html.

Walli:1998:OUA

[Wal98] Stephen R. Walli. OPENNT?: UNIX(R) application portability to Windows NT via an alternative environment subsystem, 1998. URL http://www.

usenix.org/publications/
library/proceedings/usenix98/
invited_talks/walli.ps. Un-

published invited talk at the USENIX 1998 Annual Technical Conference June 15–19, 1998 New Orleans, Louisiana, USA.

Wall:1999:OCL

[Wal99] Larry Wall. The origin of the camel lot in the breakdown of the bilingual Unix. Communications of the Association for Computing Machinery, 42(4): 40–41, April 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:
80/pubs/citations/journals/
cacm/1999-42-4/p40-wall/.

Wall:2000:LPE

[Wal00] Kurt Wall. Linux programming by example. Que Corporation, Indianapolis, IN, USA, 2000. ISBN 0-7897-2215-1. xvii + 533 pp. LCCN QA76.76.O63 W357 2000.

Wall:2001:LPU

[Wal01a] Kurt Wall. Linux programming unleashed. SAMS Publishing, Indianapolis, IN, USA, second edition, 2001. ISBN 0-672-32021-5. ixi + 886 pp. LCCN QA76.76.O63 W3573 2001.

Walters:2001:EGC

[Wal01b] E. Garrison Walters. The essential guide to computing: the story of information technology. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-019469-7. xxv + 499 pp. LCCN QA76.6 .W335 2001.

Walton:2001:LSP

[Wal01c] Sean Walton. Linux socket programming. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31935-7 (paperback). xv + 533 pp. LCCN QA76.76.O63 W358 2001.

Watson:2010:ICP

[WALK10] Robert N. M. Watson, Jonathan Anderson, Ben Laurie, and Kris Kennaway. Introducing Capsicum: Practical capabilities for UNIX. ;login: the USENIX Association newsletter, 35(6): ??, December 2010. CO-DEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (elec-

tronic). URL https://www.
usenix.org/publications/
login/december-2010-volume-
35-number-6/introducing-
capsicum-practical-capabilities-
unix.

Watson:2012:TCP

[WALK12] Robert N. M. Watson, Jonathan Anderson, Ben Laurie, and Kris Kennaway. A taste of Capsicum: practical capabilities for UNIX. Communications of the Association for Computing Machinery, 55(3):97–104, March 2012. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Wambecq:1983:NAD

[Wam83a]  A. Wambecq.  NETIX: a

distributed operating system based on UNIX software and local networking. In Software Tools Users Group [Sof83], pages 295–310. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Wambecq:1983:NDO

[Wam83b]  A. Wambecq.  NETIX: a

distributed operating system based on UNIX software and local networking. In Association [Ass83], pages 295–310.

Wandzilak:1977:RDP

[Wan77] P. D. Wandzilak. Repairing damaged PWB/UNIX file systems (I.4). Technical Memorandum 1197, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 1977. ?? pp.

Wandzilak:1978:ERP

[Wan78] P. D. Wandzilak. Exercises in repairing PWB/UNIX file systems. Technical Memorandum 1302, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 19, 1978. ?? pp.

Wang:1990:UA

[Wan90] Y. E. Gail Wang. UNIVER-SAL FILE NAMES for Ada. ACM SIGADA Ada Letters, 10 (1):111–117, January/February 1990. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wang:1996:IUX

[Wan96] Paul S. Wang. An introduction to UNIX with X and the Internet. PWS Pub. Co., Boston, MA, USA, 1996. ISBN 0-534-94768-9. ???? pp. LCCN QA76.76.O63W365 1996.

Wang:2002:DIR

[Wan02] Yu-Chung Wang. Design and implementation of RED-Linux. Thesis (Ph.D.), Electrical and Computer Engineering, University of California, Irvine, Irvine, CA, USA, 2002.

Wang:2011:ML

[Wan11]  Paul S. Wang.  Master-

ing Linux. Chapman and Hall/CRC, Boca Raton, FL, USA, 2011. ISBN 1-4398-0686-1 (paperback). xviii + 421 pp. LCCN QA76.76.O63 W365143 2011.

Ward:1978:UDF

[War78] J. C. Ward. UNIED — a data file editor. Technical Memorandum 1317, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 16, 1978. ?? pp.

Warnock:1984:PPL

[War84]  Rob Warnock.  Prospects

for personal low cost UNIX systems. UNIX Review, 2 (9):26–28, 30, 32–33, December 1984. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/unixreview_1984dec.
pdf.

Ward:1985:UEN

[War85a] Jean Renard Ward. UNIX as an environment for non-UNIX software development: a case history. ACM SIGSOFT Software Engineering Notes, 10 (3):95–108, July 1985. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Ward:1985:APT

[War85b] Terry A. Ward. Applied Programming Techniques in C. Scott, Foresman and Company, Glenview, IL, USA, 1985. ISBN 0-673-18050-6. xii + 349 pp. LCCN QA76.73.C15 W37 1985. US$19.95.

Ward:1988:SMS

[War88] Robert L. Ward. A simple menu system for MS-DOS and Unix. C Users Journal, 6(3): 49–??, March 1988. ISSN 0898-9788.

Ward:1990:SUB

[War90] Robert L. Ward. Some UNIX book recommendations. C Users Journal, 8(7):138–??, July 1990. ISSN 0898-9788.

Ward:2001:QPL

[War01] Patrick Ward. Qt programming for Linux and Windows 2000. Hewlett–Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2001. ISBN 0-13-027001-6. xvii + 280 pp. LCCN QA76.9.U83 W37 2001.

Ward:2002:BVC

[War02] Brian Ward. The book of VMware: the complete guide to VMware workstation. No Starch Press, San Francisco, CA, USA, 2002. ISBN 1-886411-72-7. xv + 249 pp. LCCN QA76.76.O63 W3653 2002. US$39.95.

Ward:2004:HLW

[War04] Brian Ward. How Linux works: what every super-user should know. No Starch Press, San Francisco, CA, USA, 2004. ISBN 1-59327-035-6. xvi + 347 + 4 pp. LCCN QA76.76.O63 W3654 2004eb; QA76.76.O63 W3654 2004. URL http:/

/www.oreilly.com/catalog/
9781593270353.

Warren:2005:VWH

[War05]  Steven S. Warren.  The

VMWare Workstation 5 Handbook. Charles River Media, Inc., Hingham, MA, USA, 2005. ISBN 1-58450-393-9. 352 (est.) pp. LCCN QA76.76.O63 W3665 2005. URL http:

//www.loc.gov/catdir/toc/
ecip0510/2005010053.html.

Ware:2008:RIE

[War08] Willis H. Ware. RAND and the information evolution: a history in essays and vignettes. Rand Corporation, Santa Monica, CA, 2008. ISBN 0-8330-4513-X, 0-8330-4816-3, 1-282-45123-5. xxvi + 201 pp. LCCN QA76.27. URL http:/

/www.jstor.org/stable/10.
7249/cp537rc;   https://
www.rand.org/content/dam/
rand/pubs/corporate_pubs/
2008/RAND_CP537.pdf.

Ward:2015:HLW

[War15]  Brian Ward.  How Linux

works: what every superuser should know. No Starch Press, San Francisco, CA, USA, second edition, 2015. ISBN 1-59327-567-6 (paperback). xxii + 366 pp. LCCN QA76.76.O63 W3654

2015. URL http://proquest.
safaribooksonline.com/9781457185519.

Watt:1983:LSU

[Wat83] Alan S. Watt. List of sources for UNIX device drivers. ;login: the USENIX Association newsletter, 8(3):14–39, June 1983. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Watkins:1987:SAU

[Wat87] Marvin L. Watkins. Software architecture and the UNIX operating system: an introduction to interprocess communication. Hewlett–Packard Journal, 38(6):26–36, June 1987. CODEN HPJOAX. ISSN 0018-1153.

Watkins:1988:AUL

[Wat88] Marvin L. Watkins. Adapting UNIX logon mechanisms to automation applications. Hewlett–Packard Journal, 39 (2):39–47, April 1988. CODEN HPJOAX. ISSN 0018-1153.

Watters:2002:SAG

[Wat02] Paul A. Watters. Solaris 8 Administrator’s Guide. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2002. ISBN 0-596-00073-1. xv + 282 pp. LCCN QA76.76.O63 W37 2002; QA76.76.O63 W3918 2002. US$39.95. URL http://

safari.oreilly.com/0596000731;
 http://www.oreilly.com/
catalog/9780596000738;
http://www.oreilly.com/catalog/
solaris8.

Watanabe:2010:SZE

[Wat10] Scott Watanabe. Solaris 10 ZFS essentials. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 2010. ISBN 0-13-704963-3 (hardback), 0-13-700010-3 (paperback). xv + 124 pp. LCCN QA76.9.F5. US$39.99.

Wayner:2000:FAH

[Way00] Wayner, Peter. Free for all: how Linux and the free software movement undercut the hightech titans. Harper Business, New York, 2000. ISBN 0-06-662050-3. viii + 340 pp. LCCN QA76.76.O63 W394 2000.

Wood:1985:PVUa

[WB85a] Jean Wood and Hans-Joachim Brede. Parlez-vous l’UNIX? The European perspective, past and future. In USENIX Association [USE85d], pages 17–23. LCCN QA76.8.U65 U8 1985.

Wood:1985:PVUb

[WB85b] Jean Wood and Hans-Joachim Brede. Parlez-vous l’UNIX? The European perspective, past and future. In USENIX [USE85a], pages 17–23.

Willcox:1992:TCS

[WB92] D. A. Willcox and S. R.

Bunch. A tool for covert storage channel analysis of the UNIX kernel. In NIST [NIS92], pages 697–706. ISBN ???? LCCN QA76.9.A25 N38 1992.

URL http://www.cl.cam.ac.
uk/~fapp2/steganography/
bibliography/021148.html.

Two volumes.

Wang:1996:PEI

[WB96] X. Wang and E. K. Blum. Parallel execution of iterative computations on workstation clusters. Journal of Parallel and Distributed Computing, 34(2): 218–226, May 1, 1996. CO-DEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (elec-

tronic). URL http://www.
idealibrary.com/links/doi/
10.1006/jpdc.1996.0058/production;
 http://www.idealibrary.
com/links/doi/10.1006/jpdc.
1996.0058/production/pdf.

Wilding:2006:SSL

[WB06] Mark Wilding and Dan Behman.

Self-service Linux: mastering the art of problem determination. Bruce Perens’ Open Source series. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 2006. ISBN 0-13-147751-X. xxii + 431 pp. LCCN QA76.76.O63 W52

Worsley:2001:GSY

[WBS01] John Worsley, Andrew Brookins, and Kai Staats. Getting started with Yellow Dog Linux. Open-Docs, Salem, OR, USA, 2001. ISBN 0-9700330-3-6. 267 pp. LCCN QA76.76.O63; QA 76.76 .O63W67 2001 SCAR.

Waters:1987:AUE

[WBSK87] F. C. H. Waters, R. G. Bias, and P. L. Smith-Kerker. AIX usability enhancements and human factors. IBM Systems Journal, 26(4):383–394, December 1987. CODEN IBMSA7. ISSN 0018-8670.

Williams:1993:JID

[WC93] Robin Williams and Steve Cummings. Jargon: An Informal Dictionary of Computer Terms. Peachpit Press, Inc., 1085 Keith Avenue, Berkeley, CA 94708, USA, 1993. ISBN 0-938151-84-3 (paperback). LCCN QA 76.15.W56 1993. US$22.00.

Walker:1998:CSP

[WC98] Kathryn M. Walker and Linda Croswhite Cavanaugh. Computer Security Policies and SunScreen Firewalls. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1998. ISBN 0-13-096015-2. xviii + 121 pp. LCCN TK5105.59 .W35 1998. US$32.00.

Wang:2012:UBW

[WCC^{+}12] Xi Wang, Haogang Chen, Alvin Cheung, Zhihao Jia, Nickolai Zeldovich, and M. Frans Kaashoek. Undefined behavior: What happened to my code? In ACM, editor, Proceedings of the Asia–Pacific Workshop on Systems: APSys ’12, July 23–24, 2012, Seoul, South Korea, page 101. ACM Press, New York, NY, USA, 2012. ISBN 1-4503-1669-7. LCCN QA75.5. URL http:/

/dl.acm.org/citation.cfm?
id=2349896.

Woyna:1993:MBS

[WCHS93] Mark A. Woyna, John H. Christiansen, Christopher W. Hield, and Kathy Lee Simunich. Modeling battlefield sensor environments with an object database management system. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):499–501, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Wills:1993:CUC

[WCM93] Craig E. Wills, Kirstin Cadwell, and William Marrs. Customization in a UNIX computing environment. In USENIX [USE93b], pages 43– 49. ISBN 1-880446-56-1. LCCN QA76.76.O63 L37 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/lisa93/wills.
html.

Wright:2006:VUS

[WDG^{+}06] Charles P. Wright, Jay Dave, Puja Gupta, Harikesavan Krishnan, David P. Quigley, Erez Zadok, and Mohammad Nayyer Zubair. Versatility and Unix semantics in namespace unification. ACM Transactions on Storage, 2(1):74–105, February 2006. CODEN ???? ISSN 1553-3077 (print), 1553-3093 (electronic).

Weeks:2004:LU

[WDJ04] Roger Weeks, Edd Dumbill, and Brian Jepson. Linux unwired. O’Reilly Media, Inc., 1005 Gravenstein Highway North, Sebastopol, CA 95472, USA, 2004. ISBN 0-596-00583-0. xii + 297 pp. LCCN QA76.76.O63 W433 2004. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596005832.

Welsh:1999:RL

[WDK99] Matt Welsh, Matthias Kalle Dalheimer, and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, third edition, 1999. ISBN 1-56592-469-X. xix + 730 pp. LCCN QA76.76.O63 W465 1999. US$34.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        runux2.

Weber:1995:CWP

[Web95]  Kate Weber.  Chapter 6,

in which Pooh proposes improvements to Web authoring tools, having seen said tools for the Unix platform. Computer Networks and ISDN Systems, 27(6):823–829, April 3, 1995. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

http://www.elsevier.com/
cas/tree/store/comnet/sub/
1995/27/6/1465.pdf.

Weeks:1999:WUC

[Wee99] John Weeks. Webifying UNIX commands. Sys Admin: The Journal for UNIX Systems Administrators, 8(10):49–50, 52– 54, 56–57, 59–60, October 1999. CODEN SYADE7. ISSN 1061-2688. URL http://www.

samag.com/.

Wehr:1977:ADK

[Weh77a] L. A. Wehr. Assembler for the DEC KMC11 microprocessor. Technical Memorandum 1225 (TM 77-8234-22), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 15, 1977. ?? pp.

Wehr:1977:DAD

[Weh77b] L. A. Wehr. Description and analysis of the DEC KMC11 microprocessor. Technical Memorandum 1226 (TM 77-8234-21), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 15, 1977. ?? pp.

Wehr:1983:UFS

[Weh83] Larry A. Wehr. UNIX file system evolution. In Association [Ass83], pages 110–?? Abstract only.

Wehrli:2001:LEP

[Weh01] Rob Wehrli. Linux Embedded Programming. CMP Books, 6600 Silacci Way, Gilroy, CA 95020, USA, 2001. ISBN 1-57820-085-7. 352 (est.) pp. LCCN ???? US$39.95. URL

                                        http://www.cmpbooks.com/.

Weinberger:1978:DTC

[Wei78a] Peter J. Weinberger. Decentralized transaction control and crash recovery. Technical Memorandum 1251 (TM 78-3444-3), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 4, 1978. ?? pp.

Weinberger:1978:ECS

[Wei78b] Peter J. Weinberger. An essay in computer security: Decrypting a former UNIX crypt. Technical Memorandum 1345 (TM 78-1271-20), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 29, 1978. ?? pp.

Weinberger:1981:UBT

[Wei81] Peter J. Weinberger. Unix Btrees. Report ??, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1981. ?? pp.

Weinberger:1982:DSM

[Wei82] Peter J. Weinberger. Database systems: Making UNIX operating systems safe for databases. The Bell System Technical Journal, 61(9):2407–2422, November 1982. CODEN BST-JAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL

http://bstj.bell-labs.com/
BSTJ/images/Vol61/bstj61-
9-2407.pdf.

Weinberger:1984:VNF

[Wei84] Peter J. Weinberger. The version 8 network file system. In ????, editor, USENIX Summer Conference Proceedings, June 1984, Salt Lake City, UT, page ?? USENIX Association, Berkeley, CA, USA, 1984.

Weinstein:1988:UST

[Wei88] Sydney S. Weinstein. UNIX signals and terminal control parameters are more natural fix for BBS problem. C Users Journal, 6(2):62–??, February 1988. ISSN 0898-9788.

Weinstein:1990:CB

[Wei90] Sydney S. Weinstein. comp.

sources.unix is back! C Users Journal, 8(8):133–??, August 1990. ISSN 0898-9788.

Weinstein:1992:UEC

[Wei92] Sydney Weinstein. USENET ELM: a case study in portability between UNIX systems. Sys Admin: The Journal for UNIX Systems Administrators, 1(1): 32–??, May/June 1992. CO-DEN SYADE7. ISSN 1061-2688.

Weinstein:1993:TWU

[Wei93] Sydney S. Weinstein. Writing a UNIX Device Driver, 2nd edition, by Janet I. Egan and Thomas J. Teixeira. C Users Journal, 11(5):79–??, May 1993. ISSN 0898-9788. See [ET92].

Weinstein:1994:UUB

[Wei94] Sydney S. Weinstein. Using UNIX for a BBS. C/C++ Users Journal, 12(9):101–??, September 1994. CODEN CCUJEX. ISSN 1075-2838.

Weiberger:1995:BRB

[Wei95] Volker Weiberger. Book review: UNIX: Einstieg für DOS-Anwender: Taudes, Alfred, Reepmeyer, Jan-Armin, & Bensberg, Frank (1993) Heidelberg: Physica-Verlag. ISBN 3-7908-0700-1, pp 232, DM40.-. Computational Statistics & Data Analysis, 19(3): 360–361, March 1995. CO-DEN CSDADW. ISSN 0167-9473 (print), 1872-7352 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0167947395900136.

Weidner:1997:GPE

[Wei97] Jeff Robert Weidner. A general purpose extended attribute service as a file system layer: with an example client that implements POSIX.6 access control lists. M.s., UCLA Computer Science Dept., Los Angeles, CA, 1997. 32 pp.

Weitsch:1998:DUS

[Wei98] John R. Weitsch. Distributed UNIX System Administrator. R&D Books, Lawrence, KS, USA; Berkeley, CA, USA, 1998. ISBN 0-87930-540-1. xiv + 317 pp. LCCN ???? US$49.95. Includes floppy disk.

Weinstabl:2004:PAE

[Wei04] Paul Weinstabl. PostgreSQL:

[Administration und Einsatz ; für Linux und Windows]. Computer and Literatur. Computerund Literatur-Verlag, Böblingen, Germany, 2004. ISBN 3-936546-22-3. 589 pp. LCCN ???? EUR 49.90.

Weinberger:197x:FEW

[Weix ] Peter J. Weinberger. File expansion when hashing into bins. Computing Science Technical Report 74, AT&T Bell Laboratories, Murray Hill, NJ, USA, 197x (??). ?? pp.

Wells:1994:WSD

[Wel94a]  Bill Wells.  Writing serial

drivers for UNIX. Dr. Dobb’s Journal of Software Tools, 19 (15):68, 70–71, 97–99, December 1994. CODEN DDJOEB. ISSN 1044-789X.

Welsh:1994:LIG

[Wel94b] Matt Welsh. LINUX installation and getting started. Specialized Systems Consultants, P.O. Box 55549, Seattle, WA 98155, 1994. ISBN 0-916151-71-9. xviii + 231 pp. LCCN ????

Welch:1995:PPT

[Wel95] Brent B. Welch. Practical Programming in Tcl and Tk. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-182007-9. xxxvi + 428 pp. LCCN QA76.73.T44 W45 1995.

Welch:1997:PPT

[Wel97] Brent B. Welch. Practical programming in Tcl and Tk. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 1997. ISBN 0-13-616830-2. xlvi + 630 pp. LCCN QA76.73.T44W45. URL http:

//www.phptr.com/ptrbooks/
ptr_0136168302.html.

Welch:1998:CTT

[Wel98] Brent Welch. The Complete Tcl/Tk Training Course. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-080756-7. xlvi + 630 pp. LCCN QA76.73.T44 W45 1998. URL http:

//www.phptr.com/ptrbooks/
ptr_0130807567.html.

Wells:1999:STYb

[Wel99a] Nicholas Wells. Sams teach yourself KDE in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31608-0. xiii + 457 pp. LCCN QA76.9.U83 W48 1999. Includes CD-ROM.

Wells:1999:LDK

[Wel99b] Nicholas D. Wells. Linux! I Didn’t Know You Could Do That. … Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 1999. ISBN 0-7821-2612-X. xiv + 283 pp. LCCN QA76.76.O63 W4638 2000. US$19.99; CAN $29.95; UK £14.99. URL

http://scooter.sybex.com/
sybexbooks.nsf/2604971535a28b098825693d0053081b/
f68011722d45db5c8825693d0057f9a6?
OpenDocument&Highlight=0,
0-7821-2612-X.

Welch:2000:PPT

[Wel00] Brent B. Welch. Practical Programming in Tcl & Tk. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, third edition, 2000. ISBN 0-13-022028-0. lvi + 772 pp. LCCN QA76.73.T44 W45 2000.

Wendland:1979:DUC

[Wen79a] K. E. Wendland. Division 56 UNIX course. Technical Memorandum 1429, AT&T Bell Laboratories, Murray Hill, NJ, USA, November 20, 1979. ?? pp.

Wendland:1979:IPM

[Wen79b] K. E. Wendland. An introduction to the PWB/MM text formatter. Technical Memorandum 1479, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 23, 1979. ?? pp.

Weste:1978:IFC

[Wes78] N. Weste. An intelligent fast color display. Technical Memorandum 1341 (TM 78-1354-15), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 28, 1978. ?? pp.

Westbury:1981:UGT

[Wes81]  Judith Westbury.  UNIX

guide for text processor users.

Report, RAND Corporation, Santa Monica, CA, 1981. ???? pp.

WEC:1983:USP

[Wes83] Western Electric Company, Inc., ???? UNIX^{T M} System Prospectus System V, January 1983. 12 pp. URL https://

archive.org/details/unix-
system-release-description-
system-v.

Weygant:2001:CHA

[Wey01] Peter S. Weygant. Clusters for High Availability: a Primer of HP Solutions. Hewlett– Packard professional books. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, second edition, 2001. ISBN 0-13-089355-2. xxxiii + 296 pp. LCCN QA76.8.H48 W49 2001. US$39.00. URL http:

//www.phptr.com/ptrbooks/
ptr_0130893552.html.

Widdowson:2001:TAU

[WF01] Liam Widdowson and John Ferlito. Tales from the abyss: UNIX file recovery. Sys Admin: The Journal for UNIX Systems Administrators, 10(11):16, 18, 20, 22, November 2001. CO-DEN SYADE7. ISSN 1061-2688.

Ward:2003:VWH

[WF03] Brian Ward and Gerhard Franken. VMware Workstation: [das Handbuch; Installation, Konfiguration, Anwendung und Troubleshooting; Gast-Systeme: Windows, Linux, BSD, Novell NetWare, Solaris, FreeDOS und Oberon; virtuelle Netzwerke, Netzwerkkonfiguration und -dienste]. mitp-Verl., Landsberg, Germany, 2003. ISBN 3-8266-0964-6. 336 pp. LCCN ???? EUR 32.00 (DE). URL http:

//www.gbv.de/du/services/
agi/FCC0A57071BE8695C125704A0029797F/
FLMA122525.

Williams:1998:WNU

[WG98] G. Robert Williams and Ellen Beck Gardner. Windows NT and UNIX: administration, coexistence,integration, and migration. Addison-Wes-ley, Reading, MA, USA, 1998. ISBN 0-201-18536-9. 784 pp. LCCN QA76.76.O63W5547 1998. US$49.95.

Wagner:2000:PSU

[WG00] David Wagner and Ian Goldberg. Proofs of security for the Unix password hashing algorithm. Lecture Notes in Computer Science, 1976:560– 572, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/1976/19760560.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/1976/19760560.
pdf.

Wolfe:1983:PAP

[WH83] Peter Wolfe and Allen Hustler.

A powerful accounting package for UNIX-Based systems.

In Software Tools Users Group [Sof83], pages 321–?? ISBN ???? LCCN QA76.8.U65 U74 1983. Abstract only.

Wheeler:1997:UBM

[Whe97]  David Wheeler.  Upgrad-

ing bred with multiple tables. Report, The Computer Laboratory, Cambridge University, Cambridge, UK, April 1997. URL http:

                                        //ftp.cl.cam.ac.uk/users/
                                        djw3/bred3.ps. This is the

third of four key papers behind the bzip2 compression tools. The others are [HL90, BW94c, SB97].

White:1975:LNU

[Whi75a] R. C. White, Jr. LKOUT — a new UNIX system call to suspend and resume input from …. Technical Memorandum 1076 (TM 75-4145-2), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 17, 1975. ?? pp.

White:1975:UDD

[Whi75b] R. C. White, Jr. UNIX DH-11 driver to support both teletypewriter and DATASPEED 40 terminals …. Technical Memorandum 1077 (TM 75-4145-1), AT&T Bell Laboratories, Murray Hill, NJ, USA, December 17, 1975. ?? pp.

White:1989:SCU

[Whi89] Eric White. Serial communications for UNIX. C Users Journal, 7(7):41–??, July 1989. ISSN 0898-9788.

Widener:1990:XIC

[Wid90] G. Widener. The X11 Inter-Client Communication Conventions Manual. Software—Practice and Experience, 20(S2): 109–118, October 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Widdowson:2000:DMF

[Wid00] Liam Widdowson. Deploying Microsoft FrontPage 2000 UNIX Web services. Sys Admin: The Journal for UNIX Systems Administrators, 9(2): 16, 18, 21–22, 24, 27, 29–30, 32, February 2000. CODEN SYADE7. ISSN 1061-2688.

URL http://www.samag.com/

.

Wiegandt:1989:UH

[Wie89]  Dietrich Wiegandt.  UNIX

and HEP. Computer Physics Communications, 57(1–3):134– 139, December 2, 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/0010465589902026.

Williams:1982:BFM

[Wil82a] Gary Williams. A businessoriented file manager under UNIX, with contention control and ISAM. In Usr Group [Usr82], pages 269–279.

Williams:1982:BOF

[Wil82b] Gary Williams. A businessoriented file manager under UNIX, with contention control and ISAM. In USENIX [USE82], pages 269–279.

Wilens:1983:SAH

[Wil83a] Michael E. Wilens. SERIX — a high performance implementation of UNIX for the IBM Series/1. In Software Tools Users Group [Sof83], pages 89– 99. ISBN ???? LCCN QA76.8.U65 U74 1983. Sponsored by USENIX Association in cooperation with Software Tools Users Group.

Wilens:1983:SHP

[Wil83b] Michael E. Wilens. SERIX — a high performance implementation of UNIX for the IBM Series/1. In Association [Ass83], pages 89–99.

Wilson:1990:SRU

[Wil90] N. Wilson. Setting up a research UNIX system. In Hume and McIlroy [HM90b], page ?? ISBN 0-03-047532-5 (vol. 1), 0-03-047529-5 (vol. 2). LCCN QA76.76.O63 U5483

                                        1990. URL http://doc.cat-
                                        v.org/unix/v10/10thEdMan/
                                        v2contents.html. Two vol-

umes. Volume 2 is edited by A. G. Hume and M. D. McIlroy, and co-published by Saunders College Publishing, under same ISBN.

Wilent:1994:PDL

[Wil94] Steve Wilent. Parallel DBMS Lets UNIX Challenge Mainframes. Datamation, 40(17): 45–??, September 01, 1994. CODEN DTMNAT. ISSN 0011-6963.

Williams:1998:MLU

[Wil98] James D. Williams. A methodology for Linux as a user process based on Solaris Minix on the SPARC architecture. Thesis (M.S.), New Mexico State University, as Cruces, NM 88003-8001, USA, 1998. xiii + 141 pp.

Williamson:2011:PC

[Wil11] Matt Williamson. PfSense 2 Cookbook: a practical, exampledriven guide to configure even the most advanced features of pfSense 2. Packt Publishing, Birmingham, 2011. ISBN 1-84951-486-0, 1-84951-487-9 (ebook). iii + 234 pp. LCCN QA76.9 .A25; TK5105.59 .W55 2011.

Winheim:1975:DHU

[Win75] I. A. Winheim. A description of how UNIX Fortran works. Technical Memorandum 1063 (TM 75-8234.3), AT&T Bell Laboratories, Murray Hill, NJ, USA, April 10, 1975. ?? pp.

Winsor:1993:SAS

[Win93a] Janice Winsor. Solaris advanced system administrator’s guide. SunSoft Press, Englewood Cliffs, NJ, USA, 1993.

ISBN 1-56276-131-5. xxvi + 477 pp. LCCN QA76.76.O63 W569 1993. US$29.95, CDN$41.95.

Winsor:1993:SSA

[Win93b] Janice Winsor. Solaris system administrator’s guide. Sun-Soft Press, Englewood Cliffs, NJ, USA, 1993. ISBN 1-56276-080-7. xix + 288 pp. LCCN QA76.9.M3 W56 1993. US$29.95, CDN$38.95.

Winsor:1999:SRG

[Win99]  Janice Winsor.  Solaris 7

reference guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-13-020048-4. 1551 pp. LCCN QA76.76.O63W568 1999. US$44.99.

URL http://www.sun.com/
books/catalog/winsor9/index.
html.

Winsor:2000:SSA

[Win00] Janice Winsor. Solaris system administrator’s guide. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, third edition, 2000. ISBN 0-13-027702-9. xxxiii + 461 pp. LCCN QA76.9.M3 W56 2000. URL

http://www.sun.com/books/
catalog/winsor11/index.html.

Winsor:2001:SAS

[Win01]  Janice Winsor.  Solaris 8

Advanced System Administrator’s Guide. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, third edition, 2001. ISBN 0-13-027703-7. l + 789 pp. LCCN QA76.76.O63W5682

2001. URL http://www.sun.
com/books/catalog/winsor12.

Winsor:2002:SOE

[Win02a] Janice Winsor. Solaris 9 operating environment reference. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-100701-7. xxvi + 1601 pp. LCCN QA76.76.O63 W569 2002.

Winsor:2002:SMC

[Win02b] Janice Winsor. Solaris Management Console Tools. Sun BluePrints Program. Sun Microsystems Press, Palo Alto, CA, USA, 2002. ISBN 0-13-046476-7. xx + 330 pp. LCCN ???? URL

http://www.sun.com/books/
catalog/winsormgmt/index.
html.

Wall:1997:PRK

[WIP^{+}97] Larry Wall, Clay Irving, Nate Patwardhan, Ellen Siever, and Brian Jepson. The Perl Resource Kit — UNIX Edition. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1997. ISBN 1-56592-370-7. 1700 (est.) pp. LCCN ???? US$149.95. URL http://www.

ora.com/catalog/prkunix/;
 http://www.oreilly.com/
catalog/prkunix.

Wisniewski:2003:LOI

[Wis03] John Robert Wisniewski. Linux and OpenVMS interoperability: tricks for old dogs, new dogs, and hot dogs with open systems. Compact solutions. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2003. ISBN 1-55558-267-2. xiii + 198 pp. LCCN QA76.76.O63 W584 2003.

                                        URL http://www.loc.gov/
                                        catdir/description/els041/
                                        2003276793.html;   http:/
                                        /www.loc.gov/catdir/toc/
                                        els041/2003276793.html.

Witschorik:1979:SOI

[Wit79] C. A. Witschorik. A shell operator for inter-machine communication. Technical Memorandum 1499, AT&T Bell Laboratories, Murray Hill, NJ, USA, August 2, 1979. ?? pp.

Withers:2001:IWU

[Wit01] Alex Withers. Integrating Windows 2000 and UNIX using Kerberos. Sys Admin: The Journal for UNIX Systems Administrators, 10(12):39, 41–42, 44, December 2001. CODEN SYADE7. ISSN 1061-2688.

Weiner:1982:UPP

[WJ82] James L. Weiner and Brian L. Johnson. UNIX/Prime: Porting the UNIX operating system to Prime machines. In Usr Group [Usr82], pages 247–248. Abstract only.

Welch:2003:PPT

[WJH03] Brent B. Welch, Ken Jones, and Jeffrey Hobbs. Practical Programming in Tcl & Tk. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, fourth edition, 2003. ISBN 0-13-038560-3. lx + 882 pp. LCCN QA76.73.T44W45 2003. Includes CD-ROM.

Wulf:1975:DOC

[WJW^{+}75] William Allan Wulf, Richard K.

Johnsson, Charles B. Weinstock, Steven O. Hobbs, and Charles M. Geschke. The Design of an Optimizing Compiler, volume 2 of Programming languages series. Elsevier Publishing Company, New York, NY, USA, 1975. ISBN 0-444-00158-1 (paperback), 0-444-00164-6 (hardcover). x + 165 pp. LCCN QA76.6 .D47. URL https:

//en.wikipedia.org/wiki/
The_Design_of_an_Optimizing_
Compiler;   https://kilthub.
cmu.edu/articles/journal_
contribution/The_design_
of_an_optimizing_compiler/
6610535.

Wood:1985:USS

[WK85] Patrick H. Wood and Stephen G.

Kochan. UNIX System Security. Hayden UNIX system library. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, 1985. ISBN 0-8104-6267-2. 299 pp. LCCN QA76.76.O63 W66 1985.

Wang:1992:UT

[WK92] H. Wang and A. Kushniruk. The UNIX tutor. Lecture Notes in Computer Science, 608:317– ??, 1992. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welsh:1995:RL

[WK95] Matt Welsh and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, February 1995. ISBN 1-56592-100-3. xxii + 576 pp. LCCN QA76.76.O63W465 1995. US$24.95.

Welsh:1996:RL

[WK96] Matt Welsh and Lar Kaufman. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, Summer 1996. ISBN 1-56592-151-8. xviii + 630 pp. LCCN QA76.76.O63W465 1996. US$29.95.

Welsh:2002:RL

[WKDD02] Matt Welsh, Lar Kaufman, Terry Dawson, and Matthias Kalle Dalheimer. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2002. ISBN 0-596-00272-6. xviii + 672 pp. LCCN QA76.76.O63 R855 2003. US$44.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        runux4.

Walsh:2003:RL

[WKDD03] Matt Walsh, Lar Kaufman, Terry Dawson, and Matthias Kalle Dalheimer, editors. Running Linux. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, fourth edition, 2003. ISBN 0-596-00272-6. xviii + 672 pp. LCCN QA76.76.O63 R855 2003. US$44.95. URL http:/

/www.oreilly.com/catalog/
9780596002725;   http://
www.oreilly.com/catalog/
runux4.

Wingo:1994:PMF

[WL94] Scot Wingo and Louis Lu. Porting Microsoft’s foundation class library to UNIX. C Users Journal, 12(1):55–??, January 1994. ISSN 0898-9788.

Wiil:1997:HHS

[WL97] Uffe K. Wiil and John J.

Leggett. Hyperform: a hypermedia system development environment. ACM Transactions on Information Systems, 15(1): 1–31, January 1997. CODEN ATISET. ISSN 1046-8188.

URL http://www.acm.org:
80/tois/abstracts/wiil.html.

Wang:1995:IEP

[WLC95] Hsiao-Hsi Wang, Pei-Ku Lu, and Ruei-Chuan Chang. An implementation of an external pager interface on BSD UNIX. The Journal of Systems and Software, 29(2):177–189, May 1995. CODEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Wang:1984:HIU

[WLHC84] Scott W. Y. Wang, Jeff B. Lindberg, M. V. Hetrick, and M. L. Connor. HP-UX: implementation of UNIX on the HP 900 Series 500 computer systems. Hewlett– Packard Journal, 35(3):7–8, 10–11, 13–15, March 1984. CODEN HPJOAX. ISSN 0018-1153. URL https://

www.hpl.hp.com/hpjournal/
pdfs/IssuePDFs/1984-03.pdf.

Waldegg:1996:MAU

[WLL96] D. Bourges Waldegg, N. Lagha, and J.-P. Le Narzul. Multimedia applications on a Unix SVR4 kernel: Performance study. Lecture Notes in Computer Science, 1185:232–??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Welsh:2001:LWS

[WLS^{+}01] Matt Welsh, Ben Laurie, Ellen Siever, Randy Jay Yarger, Alligator Descartes, and Scott Guelich. The Linux Web server CD bookshelf. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2001. ISBN 0-596-00208-4. LCCN Disk 2057 Protect. CD-ROM with Running Linux, 3rd ed. – Linux in a nutshell, 3rd ed. – CGI programming with Perl, 2nd ed. – Apache: the definitive guide, 2nd ed. – MySQL and mSQL – Programming the Perl DBI.

Wurmsdobler:2000:MMR

[WM00] Peter Wurmsdobler and Nicholas McGuire. MiniRTL: a minimal real-time Linux. Dr. Dobb’s Journal of Software Tools, 25 (12):48, 50, 52, 54, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_12/
minirtl.txt.

Wrightson:2001:MU

[WM01] Katherine Wrightson and Joseph Merlino. Mastering UNIX. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2817-3. xlii + 897 pp. LCCN QA76.76.O63 W75 2001. Includes index.

Webb:2007:PSR

[WMK^{+}07] A. M. Webb, R. Mansell, J. W.

Knight, S. J. Greenspan, and D. B. Emmes. Practical software reuse for IBM System z I/O subsystems. IBM Journal of Research and Development, 51(1/2):229–??, January /March 2007. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
rd/511/webb.html.

Waite:1984:USV

[WMP84] Mitchell Waite, Donald Martin, and Stephen Prata. UNIX System V Primer. Howard W. Sams, Indianapolis, IN 46268, USA, 1984. ISBN 0-672-22404-6. 431 pp. LCCN QA76.6 .W3186 1985. US$19.95. Highly recommended. A very very good hand-holding tutorialtype book for UNIX/SVR4.

Waite:1992:WGU

[WMP92] Mitchell Waite, Donald Martin, and Stephen Prata. The Waite Group’s UNIX System V Primer. Hayden Books, 4300 West 62nd Street, Indianapolis, IN 46268, USA, second edition, 1992. ISBN 0-672-30194-6. xxv + 564 pp. LCCN QA76.76.O63 W35 1992. US$29.95. Highly recommended. A very very good hand-holding tutorialtype book for UNIX/SVR4.

Waite:1983:UPP

[WMPC83] Mitchell Waite, Donald (Donald W.) Martin, Stephen Prata, and Kevin Caddell. UNIX primer plus. SAMS Publishing, Indianapolis, IN, USA, 1983. ISBN 0-672-22028-8 (paperback). 414 pp. LCCN QA76.76.O63 W34 1983; QA76.774.U64 W34 1983.

Wilfred:2002:SS

[WN02] Ashish D. (Daniel) Wilfred and NIIT. Solaris 9 Security. Premier Press, Cincinnati, OH, USA, 2002. ISBN 1-59200-005-3. xx + 353 pp. LCCN ????

Wang:2015:SDB

[WN15] Yongge Wang and Tony Nicol. On statistical distance based testing of pseudo random sequences and experiments with PHP and Debian OpenSSL. Computers & Security, 53 (??):44–64, September 2015. CODEN CPSEDU. ISSN 0167-4048 (print), 1872-6208 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167404815000693.

Woo:1997:OPG

[WNDO97] Mason Woo, Jackie Neider, Tom Davis, and OpenGL Architecture Review Board. OpenGL programming guide: the official guide to learning OpenGL, version 1.1. Addison-Wesley, Reading, MA, USA, second edition, 1997. ISBN 0-201-46138-2. xxxviii + 650 pp. LCCN T385 .N435 1997.

Woehr:2002:PBE

[Woe02] Jack J. Woehr. Programmer’s bookshelf: Embedded systems programming. Dr. Dobb’s Journal of Software Tools, 27 (10):86, October 2002. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/.

Woehr:2003:PBL

[Woe03] Jack J. Woehr. Programmer’s bookshelf: Linux, XP, and everything in between. Dr. Dobb’s Journal of Software Tools, 28(3):69, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303j/.

Wolfe:1993:TBP

[Wol93] Andrew D. Wolfe, Jr. Transform Brings Popular Data Visualization to UNIX. UNIX/World, 10(1):107–??, January 1993. ISSN 0739-5922.

Wolfe:2004:TSD

[Wol04a] Alex Wolfe. Toolkit: Samba does Windows-to-Linux dance. ACM Queue: Tomorrow’s Computing Today, 2(5):18–21, July 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Wolfe:2004:GTC

[Wol04b] Alexander Wolfe. Toolkit: Grid tools: Coming to a cluster near you. ACM Queue: Tomorrow’s Computing Today, 2(4):20–23, June 2004. CODEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Womack:1994:PT

[Wom94] Paula Womack. PEXlib: a Tutorial. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1994. ISBN 0-13-015843-7. 400 pp. LCCN QA76.76.W56 W653 1993. US$42.00.

Wong:1990:SSU

[Won90] Raymond Wong. A survey of secure UNIX operating systems. In USENIX Association [USE90e], pages 57–60. ISBN ???? LCCN QA 76.9 A25 U55 1990.

Wong:1992:TUC

[Won92] Carolyn W. C. Wong. The Top 10 UNIX Companies of 1992. UNIX/World, IX(12):46– ??, December 1992. ISSN 0739-5922.

Wong:1998:CCP

[Won98] Brian L. Wong. Configuration and capacity planning for Solaris servers. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, 1998. ISBN 0-13-349952-9. vii + 428 pp. LCCN QA76.76.C69W66 1997. URL

http://www.prenhall.com/
allbooks/ptr_0133499529.
html;   http://www.sun.com/
books/catalog/wong/.

Woodman:1994:PLS

[Woo94] Mark Woodman. Programming language standards scene, ten years on paper 10: Modula-2. Computer Standards and Interfaces, 16(5-6):487–494, September 1994. CODEN CSTIEZ. ISSN 0920-5489 (print), 1872-7018 (electronic).

Woznick:1982:MRU

[Woz82] Benjamin J. Woznick. Managing a roomful of UNIX systems. In Usr Group [Usr82], pages 179–185.

Waite:1990:WGN

[WP90] Mitchell Waite and Stephen Prata. The Waite Group’s New Primer C Plus. Howard W. Sams, Indianapolis, IN 46268, USA, 1990. ISBN 0-672-22687-1. xxiv + 731 pp. LCCN QA76.73.C15 W35 1990. A really good introduction to C for beginners.

Witchel:2005:MMI

[WRA05] Emmett Witchel, Junghwan Rhee, and Krste Asanović. Mondrix: memory isolation for Linux using Mondrian memory protection. Operating Systems Review, 39(5):31–44, December 2005. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wright:1990:IXW

[Wri90] Rusty C. Wright. An introduction to the X Window System. University of California, Berkeley, CA, USA, April 12, 1990. vii + 68 pp.

Wright:1993:YAB

[Wri93] Mitch Wright. Yet Another Book List, January 22 1993. This is an excellent compilation of almost all UNIX and C book titles along with info for locating them and short reviews and summaries of book contents. You can get it (yabl) by anony-

mous ftp from ftp.rahul.net
(in /pub/mitch/YABL).

Wonsiewicz:1978:WRU

[WS78] B. C. Wonsiewicz and A. R. Storm. A workshop on rebooting the UNIX time-sharing system. Technical Memorandum 1336 (TM 78-1523-30), AT&T Bell Laboratories, Murray Hill, NJ, USA, November 29, 1978. ?? pp.

Wall:1992:PP

[WS92] Larry Wall and Randal L.

Schwartz. Programming Perl. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 1992. ISBN 0-937175-64-1. xxv + 454 pp. LCCN QA76.73.P43 W35 1990. US$29.95. The authoritative guide to perl — the programming language for any serious UNIX users.

Wright:1995:TII

[WS95] Gary Wright and W. Richard Stevens. TCP/IP Illustrated: Volume 2. The Implementation. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-63354-X. xxii + 1174 pp. LCCN TK5105.55 .S74 1994. US$52.75.

Wright:1996:OSC

[WS96] Richard S. Wright, Jr. and Michael Sweet. OpenGL superbible: the complete guide to OpenGL programming for Windows NT and Windows 95. Waite Group Press, Corte Madera, CA, USA, 1996. ISBN 1-57169-073-5. xxix + 714 pp. LCCN T385 .W73 1996.

Wonsiewicz:1977:UMC

[WSS77] B. C. Wonsiewicz, A. R. Storm, and J. D. Siever. UNIX microcomputer control of apparatus, machinery, and experiments. Technical Memorandum 1219 (TM 77-1523-28, TM 77-1527-114), AT&T Bell Laboratories, Murray Hill, NJ, USA, October 21, 1977. ?? pp.

Wonsiewicz:1978:UTS

[WSS78] B. C. Wonsiewicz, A. R. Storm, and J. D. Sieber. UNIX timesharing system: Microcomputer control of apparatus, machinery, and experiments. The Bell System Technical Journal, 57(6):2209–2232, July/August 1978. CODEN BSTJAN. ISSN 0005-8580 (print), 2376-7154 (electronic). URL http:/

/bstj.bell-labs.com/BSTJ/
images/Vol57/bstj57-6-2209.
pdf.

Wesenberg:1991:LCK

[WT91] David P. Wesenberg and C. Mark Turner. A low cost knowledge based software assistant. In IEEE [IEE91a], pages 640–646. CODEN NASEA9. ISBN 0-7803-0085-8. LCCN TL 693 N37 1991. Three volumes.

Welsh:1995:DL

[Wt95a] Matt Welsh and the Linux Documentation Project. DRX. Linux. Linux System Laboratory, ????, 1995. ISBN ???? 1176 pp. LCCN ???? US$49.95.

Welsh:1995:LGS

[Wt95b] Matt Welsh and the Linux Documentation Project. Linux: Getting Started. Linux System Laboratory, ????, 1995. ISBN ???? 565 pp. LCCN ???? US$59.95. Includes CD-ROM.

Wells:1999:STYa

[WT99] Nicholas Wells and R. Dean Taylor. Sams’ teach yourself StarOffice for Linux in 24 hours. SAMS Publishing, Indianapolis, IN, USA, 1999. ISBN 0-672-31412-6 (paperback). xvi + 505 pp. LCCN QA76.73.I57W4637 1999.

Welling:2001:PMW

[WT01] Luke Welling and Laura Thomson. PHP and MySQL Web Development. SAMS Publishing, Indianapolis, IN, USA, 2001. ISBN 0-672-31784-2. xxii + 867 pp. LCCN QA76.73.P224 W45 2001. US$49.99. URL http:/

/www.mcp.com/sams/detail_
sams.cfm?item=0672317842.

Wong:2008:TAF

[WTKW08] Chee Siang Wong, Ian Tan, Rosalind Deena Kumari, and Fun Wey. Towards achieving fairness in the Linux scheduler. Operating Systems Review, 42(5):34–43, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Wunschiers:2013:CBP

[Wün13] Röbbe Wünschiers. Computational biology: a practical introduction to biodata processing and analysis with Linux, MySQL, and R. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2013. ISBN 3-642-34748-7, 3-642-34749-5 (e-book). xxix + 449 pp. LCCN QH324.2 .W86 2013.

Weerawarana:1992:PCG

[WW92] Sanjiva Weerawarana and Paul S. Wang. A portable code generator for CRAY FORTRAN. ACM Transactions on Mathematical Software, 18(3):241–255, September 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1992-18-3/p241-weerawarana/

.

Wang:2019:FEO

[WWCW19] Pei Wang, Dinghao Wu, Zhaofeng Chen, and Tao Wei. Field experience with obfuscating million-user iOS apps in large enterprise mobile development. Software—Practice and Experience, 49(2):252– 273, February 2019. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Wall:1999:LPU

[WWW99] Kurt Wall, Mark Watson, and Mark Whitis. Linux programming unleashed. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1999. ISBN 0-672-31607-2. xxiv + 818 pp. LCCN QA76.76.O63 W3573 1999. Edited by Brian Gill and Ron Gallagher.

Wu:2008:DNL

[WXX08] Fengguang Wu, Hongsheng Xi, and Chenfeng Xu. On the design of a new Linux readahead framework. Operating Systems Review, 42(5):75–84, July 2008. CODEN OSRED8. ISSN 0163-5980 (print), 1943-586X (electronic).

Xiao:1997:UTA

[Xia97]  Hong Xiao.  Using tem-

plates across UNIX platforms. C/C++ Users Journal, 15(3): 33–??, March 1997. CODEN CCUJEX. ISSN 1075-2838.

XOPEN:1988:XPGb

[X/O88a] X/Open Company, Ltd. X/ Open Portability Guide, Data Management, volume 5. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685876-7. xiii + 204 pp.

XOPEN:1988:XPGd

[X/O88b] X/Open Company, Ltd. X/ Open Portability Guide, Networking Services, volume 7. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685892-9. xiii + 144 pp.

XOPEN:1988:XPGa

[X/O88c] X/Open Company, Ltd. X/ Open Portability Guide, Programming Languages, volume 4. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685868-6. xiii + 198 pp.

XOPEN:1988:XPGc

[X/O88d] X/Open Company, Ltd. X/ Open Portability Guide, Window Management, volume 6.

Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1988. ISBN 0-13-685884-8. xiv + 338 pp.

XOPEN:1989:XPGc

[X/O89a] X/Open Company, Ltd. X/ Open Portability Guide, Supplementary Definitions, volume 3. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-685850-3. xiv + 172 pp.

XOPEN:1989:XPGa

[X/O89b] X/Open Company, Ltd. X/ Open Portability Guide, XSI Commands and Utilities, volume 1. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-685835-X. xii + 340 pp.

XOPEN:1989:XPGb

[X/O89c] X/Open Company, Ltd. X/ Open Portability Guide, XSI System Interface and Headers, volume 2. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-685843-0. xx + 666 pp.

Xu:2019:ACP

[XWRZ19] Yang Xu, Guojun Wang, Ju Ren, and Yaoxue Zhang. An adaptive and configurable protection framework against Android privilege escalation threats. Future Generation Computer Systems, 92 (??):210–224, March 2019. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL https:

//www.sciencedirect.com/
science/article/pii/S0167739X18307775.

Anonymous:1995:XWS

[XWS95] X Window System version 11, release 5. Washington, DC, USA, 1995. various pp. Contents: X Window System protocol, MIT X Consortium standard / Robert W. Schneider. – Xlib, C language X interface, MIT X Consortium standard. 1st revision, August, 1991. – X Toolkit intrinsics, C language interface. 1st revision, August, 1991. – Bitmap distribution format, version 2.1, MIT X Consortium standard.

Xia:2010:ITA

[XZY^{+}10] Liang Xia, Yongxin Zhu, Jun Yang, Jingwei Ye, and Zonghua Gu. Implementing a thermalaware scheduler in Linux kernel on a multi-core processor. The Computer Journal, 53(7):895–903, September 2010. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.
org/cgi/content/abstract/
53/7/895;   http://comjnl.
oxfordjournals.org/cgi/reprint/
53/7/895.

Yager:1991:UPD

[Yag91] Thomas Yager. UNIX program development for IBM PCs — Including OSF/Motif. Addison-Wesley, Reading, MA, USA, 1991. ISBN 0-201-57727-5. xv + 283 pp. LCCN QA76.8.I1015 Y34 1991.

Yaghmour:2003:BEL

[Yag03]  Karim Yaghmour.  Build-

ing Embedded Linux Systems. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2003. ISBN 0-596-00222-X. xxii + 391 pp. LCCN QA76.76.O63 Y35 2003. US$44.95, CDN$69.95, UK£31.95. URL http:/

                                        /www.oreilly.com/catalog/
                                        9780596002220;   http://
                                        www.oreilly.com/catalog/
                                        belinuxsys/.

Yao:1983:UA

[Yao83] Joseph Yao. UNIX APL. In Association [Ass83], pages 330– ?? Abstract only.

Yates:1984:BGU

[YE84] Jean Yates and Sandra L. Emerson. The Business guide to the UNIX system. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-08848-7 (paperback). xxi + 474 pp. LCCN HF5548.4.U53 B87 1984; QA76.8 .U65 .Y32 1984.

Yates:1984:BGX

[YEB84] Jean Yates, Sandra L. Emerson, and Candice Basham. The business guide to the XENIX system. Addison-Wesley, Reading, MA, USA, 1984. ISBN 0-201-08847-9 (paperback). xxi + 474 pp. LCCN HF5548.4.X45 Y38 1984.

Yeoh:2005:BAL

[Yeo05] C. Yeoh. Building applications for the Linux Standard Base. IBM Systems Journal, 44(2):369–??, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/yeoh.html;   http:
//www.research.ibm.com/journal/
sj/442/yeoh.pdf;   http:/
/www.research.ibm.com/journal/
sj/442/yeoh.txt.

Yggdrasil:1994:YLG

[Ygg94] Yggdrasil Linux/GNU/X operating system, 1994. ISSN 1069-3955. Yggdrasil Computing, Inc., Berkeley, CA, USA. CD-ROM. Updated quarterly.

Yang:1992:PIA

[YHR92] Wuu Yang, Susan Horwitz, and Thomas Reps. A program integration algorithm that accommodates semanticspreserving transformations. ACM Transactions on Software Engineering and Methodology, 1(3):310–354, July 1992. CODEN ATSMER. ISSN 1049-331X (print), 1557-7392 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tosem/
1992-1-3/p310-yang/p310-
yang.pdf;   http://www.acm.
org/pubs/citations/journals/
tosem/1992-1-3/p310-yang/.

Yoder:2015:BCS

[YK15] Mark A. Yoder and Jason Kridner. BeagleBone Cookbook:

Software and Hardware Problems and Solutions. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, 2015. ISBN 1-4919-0539-5. ???? pp. LCCN ????

Yang:2009:DBM

[YL09] Chao-Tung Yang and Kuan-Chou Lai. A directive-based MPI code generator for Linux PC clusters. The Journal of Supercomputing, 50(2):177– 207, November 2009. CO-DEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL http://www.
springerlink.com/openurl.
asp?genre=article&issn=0920-
8542&volume=50&issue=2&spage=
177.

Yavatkar:1993:DIM

[YLK93] Raj Yavatkar, K. Lakshman, and Giri Kuthethoor. Design and implementation of a multimedia protocol suite in a BSD UNIX kernel. In USENIX [USE93c], pages 129– 146. ISBN 1-880446-50-2. LCCN QA 76.76 O63 U83 1993.

URL  http://www.usenix.
org/publications/library/
proceedings/cinci93/lakshman.
html.

Yu:2005:MXD

[YLM^{+}05] Yijun Yu, Jianguo Lu, John Mylopoulos, Weiwei Sun, Jing-Hao Xue, and Erik H. D’Hollander. Making XML document markup international. Software—Practice and Experience, 35(1):1–14, January 2005. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Yoo:1996:CAA

[Yoo96] H. Chuck Yoo. Comparative analysis of asynchronous I/O in multithreaded UNIX. Software—Practice and Experience, 26(9):987–997, September 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=
16832.

Yost:1985:CTM

[Yos85] David Yost. The cloned tree method of revision control or A rich person’s revision control system or how I adapted the UNIX file system and tools that manipulate it to perform project revision control. In USENIX Association [USE85d], pages 229–245. LCCN QA76.8.U65 U8 1985.

Young:1989:XWS

[You89] Douglas A. Young. X Window Systems — Programming and Applications with Xt. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1989. ISBN 0-13-972167-3. x + 468 pp. LCCN QA76.76.W56 Y68 1989.

Young:1990:OMR

[You90a] Douglas A. Young. OSF/Motif Reference Guide. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-642786-3. 155 pp. LCCN QA76.76.W56 Y679 1990.

Young:1990:XWS

[You90b] Douglas A. Young. X Window Systems — Programming and Applications with Xt — OSF/Motif Edition. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1990. ISBN 0-13-497074-8. x + 533 pp. LCCN QA76.76.W56 Y67 1990.

Young:1992:OOP

[You92a] Douglas A. Young. Objectoriented programming with C++ and OSF/Motif. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-630252-1. xi + 434 pp. LCCN QA76.64 .Y68 1992.

Yourdon:1992:DFA

[You92b] Edward Yourdon. Decline and Fall of the American Programmer. Yourdon Press, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-203670-3. xvi + 352 pp. LCCN QA76.6 .Y64 1992. Despite the title, this is a very good book on the technical and management aspects of software development ….

Young:1993:PB

[You93] C. R. Young. POSIX benefits. In Anonymous [Ano93t], pages 1327–1332.

Young:1995:MDP

[You95] Douglas A. Young. Motif Debugging and Performance Tuning. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1995. ISBN 0-13-147984-9. xi + 547 pp. LCCN QA76.76.W56Y65 1995. US$36.00.

Young:1992:XWS

[YP92] Douglas A. Young and John A.

Pew. The X Window System: Programming and Applications with Xt: OPEN LOOK Edition. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-982992-X (paperback). vii + 589 pp. LCCN QA76.76 .W56Y67 1992.

Yoo:1993:UKS

[YR93] Hyuck Yoo and Tom Rogers.

UNIX kernel support for OLTP performance. In USENIX Association [USE93f], pages 241–247. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

Yu:2006:MKO

[YSC^{+}06] Liguo Yu, Stephen R. Schach, Kai Chen, Gillian Z. Heller, and Jeff Offutt. Maintainability of the kernels of open-source operating systems: a comparison of Linux with FreeBSD, NetBSD, and OpenBSD. The Journal of Systems and Software, 79 (6):807–815, June 2006. CO-DEN JSSODM. ISSN 0164-1212 (print), 1873-1228 (electronic).

Yeong:1989:UIS

[YSF89] Wengyik Yeong, Martin Lee Schoffstall, and Mark S. Fedor. A UNIX implementation of the simple network management protocol. In USENIX Association [USE89f], pages 209– 217.

Yates:1983:WUD

[YT83] Jean Yates and Rebecca Thomas. Writing user documentation for UNIX systems. In Association [Ass83], pages 117–?? Abstract only.

Yuen:2001:TPS

[Yue01] Andy Yuen. A tiny Perl Server Pages engine. Dr. Dobb’s Journal of Software Tools, 26 (8):71, 73, 77–78, 80, 82, 84, August 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_08/psp.zip;
                                         http://www.ddj.com/ftp/
                                        2001/2001_08/psplist.zip.

Zadok:2001:LNA

[Zad01]  Erez Zadok.  Linux NFS

and Automounter administration. The Craig Hunt Linux library. Sybex, 2021 Challenger Driver, Suite 100, Alameda, CA 94501, USA, 2001. ISBN 0-7821-2739-8. xlvi + 638 pp. LCCN QA76.76.O63 Z34 2001.

Zintz:1984:PPT

[ZB84] Walter Zintz and John Bass. Putting the pieces together: Strategies for affordable UNIX. UNIX Review, 2(9):34–37, December 1984. CODEN UNRED5. ISSN 0742-3136.

URL http://www.tuhs.org/
Archive/Documentation/Unix_
Review/unixreview_1984dec.
pdf.

Ziccardi:2001:DAN

[ZB01] Giovanni Ziccardi and John P.

Barlow. Il diritto d’autore nell’era digitale: evoluzione tecnologica e copyright: Internet, mp3, DivX;-), open source, Gnu/Linux, free software, mezzi di protezione. Diritto. Il sole 24 ore, Milano, Italy, 2001. ISBN 88-324-4459-3. xxi + 343 pp. LCCN KE1809 .Z53 2001.

Zaruba:2019:CAC

[ZB19] Florian Zaruba and Luca Benini. The cost of applicationclass processing: Energy and performance analysis of a Linux-ready 1.7-GHz 64-Bit RISC-V core in 22-nm FDSOI technology. IEEE Transactions on Very Large Scale Integration (VLSI) Systems, 27(11):2629– 2640, 2019. CODEN IEVSE9. ISSN 1063-8210 (print), 1557-9999 (electronic).

Ziegler:2002:LF

[ZC02] Robert L. (Robert Loren)

Ziegler and Carl B. Constantine. Linux firewalls. New Riders Publishing, Carmel, IN, USA, second edition, 2002. ISBN 0-7357-1099-6. xxiii + 562 pp. LCCN QA76.9.A25Z54

Zwicky:2000:BIF

[ZCC00] Elizabeth D. Zwicky, Simon Cooper, and D. Brent Chapman. Building Internet firewalls. O’Reilly & Associates, Inc., 981 Chestnut Street, Newton, MA 02164, USA, second edition, 2000. ISBN 1-56592-871-7. xxi + 869 pp. LCCN TK5105.59 .Z85 2000.

Zhou:1985:FST

[ZDS85] Songnian Zhou, Herve Dacosta, and Alan Jay Smith. A file system tracing package for Berkeley UNIX. In USENIX Association [USE85d], pages 407–419. LCCN QA76.8.U65 U8 1985.

Zdziarski:2005:ESB

[Zdz05] Jonathan A. Zdziarski. Ending spam: Bayesian content filtering and the art of statistical language classification. No Starch Press, San Francisco, CA, USA, 2005. ISBN 1-59327-052-6. 312 (est.) pp. LCCN TK5105.743 .Z35 2005. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/;
                                         http://www.loc.gov/catdir/
                                        toc/ecip0510/2005008221.
                                        html;   http://www.oreilly.
                                        com/catalog/1593270526/.

Zuquete:1996:TAC

[ZG96] Andre Zuquete and Paulo Guedes. Transparent authentication and confidentiality for stream sockets — ensuring private network communications for Unix and Windows systems. IEEE Micro, 16(3):

34–41, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Zuffo:1996:PEH

[ZGdDL^{+}96] Marcelo Knörich Zuffo, Andrew J. Grant, Roseli de Deus Lopes, Eduardo Toledo Santos, and João Antonio Zuffo. A programming environment for high-performance volume visualization applications. Computers and Graphics, 20(3):385– 394, May–June 1996. CO-DEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/cag/cas_sub/browse/
browse.cgi?year=1996&volume=
20&issue=3&aid=9600007.

Zielinski:1993:SVK

[Zie93] Stephan Zielinski. SAGE views:

Know your UNIX system administrator — a field guide. ;login: the USENIX Association newsletter, 18(2):25–27, March/April 1993. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).

Ziegler:2000:LF

[Zie00] Robert L. (Robert Loren)

Ziegler. Linux Firewalls. New Riders professional library. New Riders Publishing, Carmel, IN, USA, 2000. ISBN 0-7357-0900-

9. 470 pp. LCCN TK5105.59 .Z54 2000.

Zientara:2016:MP

[Zie16] David Zientara. Mastering pfSense. Packt Publishing, Birmingham, 2016. ISBN 1-78646-336-9 (e-book), 1-78646-343-1. LCCN T55.4-

                                        60.8. URL http://proquest.
                                        safaribooksonline.com/9781786463432.

Zimmerman:1985:DUK

[Zim85] Steven A. Zimmerman. A debugger for the UNIX kernel. In USENIX Association [USE85d], pages 151–153. LCCN QA76.8.U65 U8 1985.

Zimmer:1998:TTP

[Zim98] J. Adrian Zimmer. Tcl/Tk for Programmers, with Solved Exercises that Work with Unix and Windows. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8515-8. 450 pp. LCCN QA76.73.T44Z56 1998. US$45.00.

Zintz:1992:YWL

[Zin92]  Walter Zintz.  So, You

Wanna Learn About UNIX. UNIX/World, 9(7):65–??, July 1992. ISSN 0739-5922.

Zeller:2005:EOS

[ZK05] Andreas Zeller and Jens Krinke. Essential open source toolset: programming with Eclipse, JUnit, CVS, Bugzilla, Ant, Tcl/Tk and more. Wiley, New York, NY, USA, 2005. ISBN 0-470-84445-0 (paperback). xii + 392 pp. LCCN QA76.76.D47 Z45 2005. URL http:

                                        //www.loc.gov/catdir/toc/
                                        ecip053/2004026271.html.

Zeller:1996:DFG

[ZL96] Andreas Zeller and Dorothea Lütkehaus. DDD — a free graphical front-end for UNIX debuggers. ACM SIGPLAN Notices, 31(1):22–27, January 1996. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zarra:2010:CAS

[ZL10] Marcus Zarra and Matt Long.

Core animation: simplified animation techniques for Mac and iPhone development. Core frameworks series. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-321-61775-4 (paperback). xii + 245 pp. LCCN TR897.7 .Z37 2010.

Zlotnick:1991:PSP

[Zlo91] Fred Zlotnick. The POSIX.1 Standard: a Programmer’s Guide. Benjamin/Cummings Pub. Co., Redwood City, CA, USA, 1991. ISBN 0-8053-9605-5. xi + 379 pp. LCCN QA76.76.063 Z57 1991.

Zeeb:2008:BFD

[ZM08] B. Zeeb and Marshall Kirk McKusick. BSDCan 2008 FreeBSD Developer Summit summary. In ????, editor, Proceedings of the 2008 BSD Canada Conference, Ottawa, Canada, May 2008, page ?? ????, ????, 2008.

Zajcew:1993:OUM

[ZRB^{+}93] Roman Zajcew, Paul Roy, David Black, Chris Peak, Paulo Guedes, Bradford Kemp, John LoVerso, Michael Leibensperger, Michael Barnett, Faramarz Rabii, and Durriya Netterwala. An OSF/1 UNIX for massively parallel multicomputers. In USENIX Association [USE93f], pages 449–468. ISBN 1-880446-48-0. LCCN QA 76.76 O63

U84 1993. URL http://www.
usenix.org/publications/
library/proceedings/sd93/.

Zee:2007:ISZ

[ZST^{+}07] M. Zee, J. W. Stevens, B. L. Thompson, J. A. Fowler, J. Goldman, P. T. Chan, and T. P. McSweeney. IBM System z9 open systems adapter for communication controller for Linux. IBM Journal of Research and Development, 51 (1/2):119–130, January /March 2007. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
rd/511/zee.html.

Zucker:1977:LMS

[Zuc77a] H. Zucker. A lexicographic multikey sort program ZSORT. Technical Memorandum 1209, AT&T Bell Laboratories, Murray Hill, NJ, USA, October 31, 1977. ?? pp.

Zucker:1977:ICE

[Zuc77b] Steven Zucker. Interprocess communication extensions for the UNIX operating system: II — implementation. Report R-2064/2-AF, Rand Corporation, Santa Monica, CA, USA, June 1977. viii + 19 pp.

URL ttps://apps.dtic.mil/
sti/pdfs/ADA044201.pdf.

Zucker:1983:CLM

[Zuc83a] Steve Zucker. Contiguous load modules for UNIX. In Association [Ass83], pages 39–?? Abstract only.

Zucker:1983:CEU

[Zuc83b] Steven Zucker. IS/3: a compatible extension of UNIX system III. In USENIX [USE83], pages 325–329.

Zucker:1983:IAC

[Zuc83c] Steven Zucker. IS/3: a compatible extension of UNIX system III. In Association [Ass83], pages 325–329.

Zwaska:1995:TTU

[Zwa95] Steve Zwaska. Tools and Toys for UnixWare. Prime Time Freeware, 370 Altair Way, Suite 150, Sunnyvale, CA, USA, 1995. ISBN 1-881957-15-2. 136 pp. LCCN ???? US$60.00. Includes two CD ROMs.

Zwicky:1992:SBR

[Zwi92] Elizabeth Zwicky. SAGE book reviews: UNIX for superusers. ;login: the USENIX Association newsletter, 17(6):13– ??, November/December 1992. CODEN LOGNEM. ISSN 1044-6397 (print), 2169-9364 (electronic).